A Complete Bibliography of Computers and Security, 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

24 October 2025 Version 1.43

Title word cross-reference

#31415 [KLA^{+}22]. #Segments [WZX^{+}22].

27701 [CRdMM^{+}25]. 3 [CCS^{+}24]. 5 [BBBB^{+}20]. ^{1} [RCD21]. ^{2} [CG24]. _{diff} [LTCR25]. α [QLZ^{+}25]. i [HKM^{+}23]. K [KE25, LSTL24, XHTZ20, GT24a, Kac25, LS20, LLL^{+}23c, SHK^{+}21, SYG^{+}21, XGS25]. L [RKG^{+}25]. O(1) [ZCS^{+}23]. μ [LQL25b]. N [BDA23, SQL^{+}24b, WLY^{+}24, MT21c, OO21, TQJ^{+}23]. Q [DP25, WHW^{+}24]. S^{3} [OX22]. Z [HZZ20].

*nix [LCSD20].

-Anonymity [SHK^{+}21, GT24a, Kac25, LS20]. -based [AOA22, CRdMM^{+}25]. -confidence [QLZ^{+}25]. -gram [TQJ^{+}23, BDA23, SQL^{+}24b]. -HTLC [MT21c]. -Means [XHTZ20, LSTL24]. -moments [RKG^{+}25]. -network [WHW^{+}24]. -Salp [KE25]. -shot [WLY^{+}24].

/102599 [HuRMMW22]. /compliance [GGK25].

1.2 [TO22]. 102463 [ZHJ^{+}23]. 102757 [GMLB23]. 111 [ZHJ^{+}23]. 112 [Ske24]. 117 [ATMN23, KL23]. 120 [GMLB23]. 124 [DJG^{+}25]. 126 [FFLQ24]. 137 [KLSD24a]. 141 [LCH^{+}24a]. 144 [LWSW25]. 150 [BCGC25a]. 154 [ABAJ25a]. 1553 [DMRV21]. 19 [HvCSJ23, KSVY22, LSN^{+}21, SAJP22, vdWLM24].

2 [CSB^{+}25, CKK^{+}23, DMPT25, WAPA25]. 2.0 [MV23, WLC^{+}22]. 2020 [Ano20a, Ano20b, Ano20c, Ano20o, Ano20r, Ano20q, Ano20s, Ano20t, Ano20u, Ano20v, Ano20p, Ano20w]. 2021 [Ano21a, Ano21b, Ano21c, Ano21p, Ano21r, Ano21q, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21y]. 2022 [Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22y]. 2023 [Ano23a, Ano23b, Ano23c, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x]. 2024 [Ano24a, Ano24b, Ano24c, Ano24p, Ano24q, Ano24s, Ano24r, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Pit25]. 2025 [Ano25a, Ano25b, Ano25c, Ano25p, Ano25q, Ano25s, Ano25r, Ano25t, Ano25u, Ano25v, Ano25w, Ano25x]. 256 [GCN^{+}24]. 2DF [FFB^{+}23]. 2DF-IDS [FFB^{+}23]. 2PC [ZFWL22]. 2PC-based [ZFWL22].

3 [CKK^{+}23].

4 [MLS^{+}23]. 4.0 [BAM^{+}24, BSR^{+}25, GMB^{+}25, MNMS24]. 42001 [MSL^{+}24].

5.0 [CBB25, MVH^{+}21]. 5G [DBB^{+}24, GC22, JB21, LDS24, Mit20, NR22, SDA^{+}24b, WAPA25, YM22, YCWW23]. 5G/6G [SDA^{+}24b].

62443 [DMPT25]. 62443-3-2 [DMPT25]. 6G [SDA^{+}24b, TLHY25, ZHL25]. 6LoWPAN [PCG23, ZMQ21]. 6LoWPANs [ASH^{+}24].

8/102695 [KL23]. 802.11ah [ZM20]. 802.11v [LPS23].

97 [CAT24].

Aadhaar [SS24a]. ABAC [CFT22, MYS^{+}25, SWL24]. ABE [SYH^{+}22]. abnormal [GZ24, WLL^{+}23, ZZW^{+}24]. absolute [ZZJC20]. abstract [XLW24, RKG^{+}25]. abuse [ASG^{+}25]. abusive [XWY21]. AC [LS21]. academia [dVABH20]. Accelerating [SMC^{+}21]. accept [RvdSM23, ZZ25]. Access [HK25, IMN22, MBK^{+}22, TFJ23, AOM24, AJ20, ABK^{+}20, CPT^{+}22, CZZ^{+}21, DFJ^{+}23, DS21, DZSCR24, FMU22, GRKS24, GYZ^{+}23, GWY^{+}25, GKC22, LCSD20, LLZJ22, MKC^{+}21, MKC^{+}23, MVSSZ24, PGC^{+}24, RYS25, SBY^{+}22, SU21, WGS^{+}25a, WP21, Yan23, YLW24a, ZLZC22]. access-control [WP21]. access-driven [GYZ^{+}23]. access-graph [DS21].

Accessible [FHW22, KRN24]. Accomplished [GM24, SL20]. account [PGZB23]. accountability [TGOF22]. Accountable [JP25, ILB20, LXLL25]. Accounting [NCW^{+}20]. Accounts [CHP21, AAAAS20, LXH^{+}23, TG25]. accuracy [CWZ^{+}24b, GTL^{+}22, HHE22, YFO^{+}21]. Accurate [CWW^{+}25, CZM^{+}25, ZX23b, ADJS21, AK24b, MGG^{+}20, SS25]. Achieve [YLZ^{+}21]. Achievements [MAKH21]. Achieving [LZL^{+}22, LZ20, TLPY21, GZG21]. Acknowledgement [Ano21x, Ano22x]. Acoustic [GSE20, YZLS25]. Acquisition [US20, FSN22]. across [GAW^{+}23, TJD23]. Act [Bel25]. ActID [SKS21]. action [KXZ^{+}23, MGZ^{+}20, WZG24, CZE^{+}22]. actions [MGTK22]. activation [LaTL^{+}25]. activations [AKAFQ22]. Active [YWW24, ASI^{+}20a, ASI^{+}20b, CRC^{+}23, DMRV21, FSN22, KL24, LWX^{+}22, MT24, NGX^{+}23, OCMF23, WHW^{+}24, XYH^{+}20, ZT21]. active-risk [ASI^{+}20a, ASI^{+}20b]. actively [LZ25]. actives [CYL^{+}21]. activist [JB22]. activities [CEO^{+}25]. activity [FO25, HG21, IM23, LdGW^{+}22, LLG21, RT20, SKM21, SKS21]. activity-approach [LdGW^{+}22]. Actor [VHMGRR22, CKS^{+}24, DG25, LHD^{+}23, RCD25, XLWC24, DANS20]. actor-critic [CKS^{+}24, DG25, LHD^{+}23]. actor-observer [RCD25]. Ad [GRJV24, KJR^{+}20, MBP20, RKA25]. ad-hoc [KJR^{+}20, MBP20]. Adam [TCYL21]. adaptable [HLZ^{+}21, IJR^{+}25, LSMK25]. Adaptation [DPSN20, BLDB24, CTC^{+}25, CZPX22, LXZ^{+}23, ZZHY25, WCZ^{+}24]. adaption [FK24]. Adaptive [DCS24, HN25, LLD^{+}25, MSL^{+}25, NLW24, Rep23, SSZ^{+}24, ZDH^{+}23, ZXZJ25, AGAP24, ACGP25, CB22a, CPT^{+}22, DBG24, ESTA^{+}25, HZ22, LMD^{+}22, LHL^{+}24a, LGH21, LWS^{+}25, WCZW23, XSK23, XHWL20, ZZP^{+}25, LXW^{+}24]. Add [AMN22]. adding [HX21]. additional [AdSJN25]. additive [AAS24, KEM25b]. address [SPZ^{+}20, WLLS22, ZYH^{+}22, ZZJC20]. Addressing [MNMS24, WLKHD25, BC23a]. ADESSA [NGX^{+}23]. ADF [TLKW20]. ADFL [ZZS^{+}23]. adhoc [RJ21]. AdIoTack [PBG22]. adjustment [PEGA23]. administration [SSRK20]. ADMM [YWZ20]. ADOC [KG20]. Adopting [AMM^{+}21, MEH^{+}24]. adoption [HNH^{+}22, KLC23, NDDH^{+}21]. ADS [FL21, LWS^{+}20, LWLT21, LWT24]. ADS-B [FL21, LWS^{+}20, LWLT21, LWT24]. ADSAD [QMW20b]. AdStop [AA22b]. ADT [YHS24]. Adv [DCK^{+}23]. Adv-Bot [DCK^{+}23]. Advanced [KK25c, AA25, AV25, CZZ^{+}23, CZPX22, HWD^{+}25, ISR^{+}23, NKA^{+}25, QYYH24, RBHW25, WPAG24, WHLS24, XGL20, XLWC24, ZGNZ21, AASWC25, HZ20b]. Advances [QGW^{+}25, AKT21, BBA^{+}21]. Advancing [CNK25, KK25a, MV25, SDA^{+}24a]. advantage [DG25]. AdvCheck [CJC^{+}24]. Adventure [LDW25]. Adversarial [DZZ^{+}22, DV24, HKY^{+}25, JTF^{+}24, JJW^{+}23, KWF25, LYL^{+}23, LCH^{+}24a, LWZ^{+}23, LXY^{+}24, LLPL25, LET^{+}25, MDB^{+}20, QMW^{+}20a, QZT^{+}22, QZP23, RKH^{+}22, XLL23, ZHL^{+}22, AIBKJ24, ALAN25, AESF23, AFAkJS25, AWJB21, AQAK21, AGAS25, BLDB24, CWW^{+}23, CZCX20, CZX^{+}21, CTH^{+}22, CHZ^{+}23, CJC^{+}24, CHP^{+}24, CWX^{+}24, CCS^{+}24, CWZ^{+}24b, CDLC24, DCK^{+}23, DSD^{+}25, DTT^{+}22, DDW^{+}23, DZZ^{+}21, DTK^{+}21, GSSC24, GDM^{+}24, GWL^{+}23, GHZS23, GT24b, HLC^{+}24, HLP22, HMTC22, HDZ^{+}22, HLQD23, HKA24, HLYZ21, HWWB24, ISM22, JLC25, KKK23, KDTR25, KS23a, KN23, LL21a, LMD^{+}22, LLF^{+}23, LYSB23, LGZD23, LFHY23, LZRR23, LHX23, LZNZ24, LCH^{+}24b, LLD^{+}25, LaTL^{+}25, LYLT24, LWW^{+}21, LCL^{+}25, LCL24, LXZ^{+}22, LJJ22, LLX^{+}23, LWH^{+}24, LKLZ24, LQL25b, LLX25, LGXZ22, MPL^{+}25, MLC^{+}23, MCM^{+}24, MBSE22, MBGF20, MLQ^{+}23, MRL22, MKZ^{+}23, NDPC21, PBG22, PLW^{+}21, QJQ22, QWW^{+}22, QMZ^{+}23, QSA24, RKA25, RS23, RZ24, RNSS23, SKD23, SZY^{+}22, SYL^{+}20]. adversarial [SLX^{+}23, SHH^{+}24, TSAG24, VS22, VPM23, WYW^{+}22, WYJ^{+}22, WZZ^{+}22, WMK23, WSW24, WLLL24a, WLLL24b, WS21, WTC^{+}22, XSC^{+}23, XY24, YCCZ21, YL23, YS22, YHH^{+}24, YXL^{+}22, ZDH^{+}23, ZLH22, ZM22, ZQCX24, ZXZJ25, ZFT24, ZHZ^{+}22, ZLH^{+}25, ZZS^{+}23, ZZZ^{+}23, ZFG^{+}24, PAGDG24]. adversarially [AKA^{+}23]. adversaries [GMH^{+}23]. adversary [ABAJ25a, ABAJ25b, DGH21]. advertisement [XZG^{+}25]. advertising [SZL^{+}22]. advice [BvO23, RCD23, SH25]. adware [AA22b]. AECR [CZL^{+}25]. Aerial [SGC23, HMBY23]. affect [FSMM24, KSMW23, SHK^{+}21, XLL^{+}25, vSRW^{+}20]. affected [MGTK22]. affecting [BCP22, EMS23]. affective [FVF21]. AflIot [DCH^{+}22]. AFPr [AZO^{+}23]. AFPr-AM [AZO^{+}23]. after [AABLS24, Fur22, KN20, KSVY22, LLC^{+}23, XLG^{+}23]. AG [MLB^{+}25]. Against [CCO^{+}20, LLL^{+}23a, SYA^{+}21, AMVG24, AA20, AWJB21, AQAK21, BWZ^{+}25, BvtH^{+}23, BGM^{+}23, CWW^{+}23, CGS22, CLS22, CHWH25, DG20, DCK^{+}23, DS20, DMZ24, DSD^{+}25, DDW^{+}23, GR24, GMPMS21, GPAF25, GVJ23, GB23, HSX^{+}24, HCLRM20, HZX^{+}21, HXZ^{+}21, HZ20b, HLYZ21, HZ22, HYZ^{+}24, JYCL25, KKR^{+}24, KL24, KKK23, KDTR25, LT25, LBA25, LKA^{+}21, LL22b, LGZD23, LWYG24, LHZZ24, LX21, LWW^{+}21, LCL^{+}25, LWZ^{+}23, LXZ^{+}22, LZX^{+}23, LKLZ24, LQL25b, LJO^{+}20, LLLC22, MPM23, MPL^{+}25, MCM^{+}24, MT21c, MDB24, NPF^{+}23, NOVJ23, NCW^{+}23, NHC25, PBG22, PAGDG24, QY22, QYYH24, RKA25, RS23, RZ20b, RZW^{+}20, RZY^{+}24, SKD23, SMM22, SKG21, STS25, TWLK22, Wan21, WZX^{+}22, WSW24, XLZ^{+}22, XRW^{+}25, XHWL20, XHW^{+}22, YHW^{+}23, YJK^{+}24, YWWH22, ZDH^{+}23, ZXL^{+}25, ZSL24, ZXL^{+}24, ZTYA25, ZNL^{+}20, ZHZ^{+}22, ZCJ^{+}20, ZLH^{+}25, ZHC25a]. AGDs [RLL^{+}23]. age [LSN^{+}21]. agenda [ATMA23, PNF23, SV25, vdSPR24]. agent [TSW^{+}24, ZLC^{+}23b]. agents [GSG23a, NHFB25, YK22b]. Aggregate [AGBB22, WBZ^{+}25]. aggregated [HHL^{+}22]. Aggregation [WR24, ZJL^{+}20, BS20, LLL^{+}23c, SVA21, WYJ^{+}24]. Aggregation-based [ZJL^{+}20]. aggregator [HYW^{+}20]. aggressive [CCS^{+}24]. agile [LCC^{+}22, LHYW22, NNA^{+}23, Rep23, TCJS22, SJH^{+}21]. aging [LBWV23, LWP^{+}25]. AGLFuzz [ZGG^{+}25]. agnostic [LCH^{+}24a, LCH^{+}24b]. agreement [NM21, AA24b]. agriculture [CMLM25, BAM^{+}24, BSR^{+}25]. ahead [GBS22]. AHP [AZO^{+}23, ÖEÖ22]. AHP-based [ÖEÖ22]. AI [Bel25, CYL^{+}24, GLM^{+}24, GT24b, KLC25, MDK^{+}23, SRY^{+}25, SGJ^{+}25, TLHY25, VBS^{+}25, WFK25]. AI-based [GLM^{+}24]. AI-enabled [VBS^{+}25]. Aided [LZG^{+}22, XCY^{+}25]. aiding [Jan22, PL23]. AIDTF [XLL23]. Aiming [LGJW21]. AIPA [DV24]. Air [IJR^{+}25, CVL22, GSE20, LWT24]. air-gapped [GSE20]. AIR4ICS [SJH^{+}21]. airborne [FL21]. AIS [FWKB24]. AIS-NIDS [FWKB24]. AJSAGE [XZZ^{+}25]. alarm [AES22]. ALERT [DANS20, BS20, CKS^{+}24, HZ22, IE22, KMH^{+}22, MFA20, WYL^{+}24, WYJ^{+}24]. Alerting [MDR20]. AlertPro [WYL^{+}24]. alerts [MWM^{+}25]. Alexa [AR25, SGT24]. algebra [TA25]. algorithm [AA22a, AR21, CDG^{+}23, DMM^{+}23, FYL^{+}24, FNZ^{+}23, GSSC24, GYZ^{+}25b, GSS^{+}20b, HYW^{+}21, KLW^{+}23, KBMM23, LGL21, LHD^{+}23, LSTL24, LLL^{+}25, MZW^{+}23, MS23, MDKR25, NMLN23, TT20, VC25, WGRZ23, WXJ^{+}25, XYH^{+}20, XHB^{+}21, YLL23, YWZW24, ZGM21, ZZL^{+}22, ZHL^{+}20, ZCPB23, DP25, Kac25]. algorithm-based [HYW^{+}21, LSTL24, DP25]. algorithmic [Bel25, MRSP24, MC20]. algorithmically [AKCS20]. algorithms [ARAG23, Bel25, GJB22, KAO^{+}23, LS20, MAP23, MGG^{+}20, NJ22, SB23, TMJ^{+}23, WT23, YFO^{+}21, ZCPB23]. Aligning [Coj25, PFR20]. Alignment [LXW^{+}24, DLQ^{+}21, SAR24]. alive [BTAK21]. allocation [AEA24, JSYK25, LSTL24, ZZW^{+}21]. along [SS23b]. Altering [GBA22]. alternate [WRG^{+}21]. always [ASF^{+}23, BOS^{+}21]. Am [XGS^{+}20, AZO^{+}23]. AMAD [YWW24]. AMalNet [PYT20]. ambidextrous [SSK22]. AMGmal [ZDH^{+}23]. AMI [KTGDE20, KAO^{+}23]. amid [FSMM24]. among [BTA24, CVZHN25, JCS^{+}25, ND23, OJH25, SA22]. amongst [JQYL22, PL23]. Amplification [LCH^{+}24a, DHW24, IHJZ21, LCH^{+}24b]. amplification-based [IHJZ21]. analyses [MAR22, SAM25]. Analysing [TDM^{+}21]. Analysis [ACG20, BVB^{+}23, CMFUA^{+}20, CB22b, DAZ20, GAW^{+}23, GBA22, KUA^{+}25, VDL24, WCOZ23, WHLS24, XLZ^{+}22, tBLLV21, AHS25, AMA22, AM25, AD23, ASF^{+}23, ABAM22, AABE20, AKA^{+}23, AAS24, Alm25, ASB^{+}21, ADAT25, BEFI25, BKS22, BBA^{+}21, BvO23, BD23, BAD^{+}24, BDM^{+}24, BLR23, BWB^{+}21, BSR^{+}25, CCL^{+}24, CMA23, CGBK25, CJL^{+}25, CRP25, CGH^{+}23, CDM^{+}20, CBFH20, CDN21, DPKHP22, DMB21, DWX^{+}23, EFI23, FHSQ20, FZZ^{+}25, FFPC22, FFPC23, FBL20, GAL^{+}20a, GSSP23, GYL^{+}20, GKS24, GYZ^{+}23, HHSL20, HRU24, HWD^{+}25, HvCSJ23, HSL20, HLZ^{+}21, JCS^{+}25, JK21a, JYZ^{+}23, kJLcL24, JYWS25, JYF^{+}25, JFQ^{+}22, JLS22, KKGK25, KKP22, KM22, KAK21, KA23, KMP^{+}20, KLA^{+}22, KTH20, KIAV22, LSN^{+}21, LCL23, LCZ^{+}22, LLL^{+}23b, LHL^{+}24b, LWD^{+}25, LXW^{+}25b, LdSP21, LZZ^{+}21, LFW^{+}22, LP24, MYS^{+}25, MLB^{+}25, MMC24, MVVL^{+}24, MCC^{+}24, MVH^{+}21, MSMH21, MCMMMA23]. analysis [MDB24, NH21, NTBH^{+}22, NFVN^{+}22, NCPVRT23, NDDH^{+}21, NMCRB21, NGC^{+}25, NOVJ23, NZZ^{+}22, OUR^{+}23, Onw20, OXNA25, PLN23, PMF^{+}20, PvSvdB25, QAA23, QMW20b, RKSV20, RGPAF20, RGACPA23, RMS^{+}21, SRN24, SEM^{+}23, SCH^{+}20, SSHP21, SRY^{+}25, SGT24, SLR25, SMvH^{+}21, SVPM21, SWS^{+}25, Spr23, SDG20, SGVA21, SH25, STM^{+}20, SGA^{+}23, SSTRD22, TDYQ23, TKKF24, TD21, TFP21, TLY^{+}24, TK20, WX23, WK23, WBN^{+}20, WPS20, WCL25, WGS^{+}25b, XLH^{+}25, XZY^{+}24, XZZ^{+}24, YZL^{+}20, YGC^{+}23, YPWS20, YFdCdC^{+}24, YSM^{+}21, YP21, ZKL^{+}23, Zen23, ZMF^{+}25, ZTD21, ZYH^{+}21, ZJY^{+}23, ZSGB^{+}22, vdSPR24]. analysis-based [MMC24, Onw20, SWS^{+}25]. analyst [ACRB23]. Analytical [YZW^{+}20, CVZHN25, HRFS24]. analytics [Lee23, MGZ^{+}25, MWR20, NNA^{+}23, UHK^{+}21]. analytics-embedded [NNA^{+}23]. Analyzing [DS23, HP23, KF25, Lee23, MB20b, OCMF23, RW24, SW22, SKS25, WFKH25, GGR^{+}24, LXX^{+}25, PGZB23, QPL^{+}23, SS22, ZYL^{+}20]. anatomy [CPP24]. anchors [PLE^{+}23]. Android [AYS20, AES22, ACG20, BBD23, BCM^{+}25, BM24, CLX21, CQLW25, DLC^{+}24b, EZLC21, FFPC23, FCSP21, GAL^{+}20a, GCZ21, GB21, GZH^{+}24, GRdAS23, GMBN21, GMLB22, GMB22, GMLB23, GM24, HZ20a, JDBB20, JDBB22, KS25a, KDDM20a, KDDM20b, KMP^{+}20, KZYZ22, LHL^{+}24a, LL23, LQL25b, LWP^{+}25, LLG21, MSG22, MRSV21, MMdR24, MCMMMA23, MHLZ22, NCN24, OX22, PDML25, RMCT23, STPB25, SFX24, SLZ^{+}25, ST24, SGVA21, TDYQ23, TTA23, WWL^{+}25a, WPS20, WLZ^{+}23, XLH^{+}25, XSC^{+}23, YMR^{+}22, YWX^{+}22, YWZ^{+}24, ZMF^{+}25, ZTD21, ZSLY25, ZYL^{+}25, ZTY^{+}20a, ZWX^{+}24]. AnForA [ACG20]. angle [JK23]. ANiTW [ALZ^{+}20]. ANN [CVZHN25, LS21]. ANN-based [LS21]. annuli [ZTK^{+}23]. anomalies [DGD23, KF25, MBFIdD21, SPN^{+}25, SRM22, STS25]. Anomalous [CWYW25, JQYL22, GZ24, KS23b, SG21]. Anomaly [CLN^{+}25, DAZ20, DKV^{+}24, HMT^{+}25, KE25, KTGDE20, KDBS22, LSG^{+}24, LWT24, SJMP23, SPL25, SHH^{+}24, TXY^{+}23, WMK23, XLZ24, ZLC^{+}23a, ZHJ^{+}23, AHZL25, ADDH25, AHFL25, AA22a, BBMW21, BTM21, BAC^{+}23, CFZL22, CLL24, CKL^{+}25, CVL22, CSKD22, DVAAM23, EA20, GAZ25, GHMG24, HMS21, HCCS22, KT23, LCZ^{+}24, LZY^{+}25a, LYLT24, LWZ22, LWLT21, LWLL25, MSCJ21, MMSJ25, NGX^{+}23, PSP^{+}22, PHT^{+}25, QMW20b, RDM20, dCFA^{+}23, SAB^{+}25, SLC^{+}21, TLY^{+}24, WWWS24, WWS^{+}25, WQL25, WAPA25, WHLS24, YLL^{+}22, YL23, YHS24, YQL24, YCLZ24, YLY20, YKY24, YKSK25, YWW24, ZHJ^{+}21]. Anomaly-based [KTGDE20, YLL^{+}22]. Anonymity [SHK^{+}21, GT24a, Kac25, LS20, MYH25, XCSZ21, YFO^{+}21]. anonymization [FMA22, JN23, SS23b]. Anonymous [AA24b, ABK^{+}20, GXL25, OO21, SJX^{+}20, XCSZ21]. anormaly [LCZY23]. antecedents [SA22]. Anti [ACCS25, MRSV21, YLL23, BMV24, GMBV21, GMPMS21, JK23, KL24, LL21a, LZX^{+}23, NBBS22, SGSS22, SARG23, VK22, WS21, WTC^{+}22, XSL^{+}24, ZXH^{+}25]. anti-forensics [WS21, WTC^{+}22]. anti-fraud [VK22]. anti-fuzzing [KL24]. anti-malware [GMPMS21, LL21a, NBBS22, XSL^{+}24]. Anti-money [YLL23]. anti-phishing [BMV24, GMBV21, SARG23]. Anti-Repackaging [MRSV21]. anti-spoofing [JK23]. anti-theft [LZX^{+}23, ZXH^{+}25]. anti-virus [SGSS22]. Anticipating [DGH21]. antiphishing [HSL23]. AntiTomo [LXZ^{+}22]. Antivirus [PBdL25, RG23]. antiviruses [BCdGG20, BDC^{+}22]. anxiety [WLKT23, WS24]. anycast [PdARD24]. APELID [VDN24]. API [ADDH25, AZ20, AZES21, CZLZ24, CLMU25, DSN^{+}22, LCZ^{+}22, LLL^{+}22c, SB25, SN24, SLBG21, STM^{+}20, WGS^{+}25b, YLH^{+}24, YWZ^{+}24]. API-based [CLMU25]. APIARY [CLMU25]. Apícula [DSN^{+}22]. APIs [CJD25, LXW^{+}25b]. APLDP [SSZ^{+}24]. Apollon [PAGDG24]. Apologize [MGTK22]. app [FCSP21, LLG20, OMO20a, RMCT23, SLBG21, TTRY20, ZLCA21, HP23]. app-in-the-middle [LLG20]. appeal [JDB^{+}23]. appliances [AL23, LLC^{+}23]. Applicability [Les21, FHW22, KEK^{+}21, PLE^{+}23, vGCHZ24]. Application [LCP^{+}23, CEO^{+}25, CRdMM^{+}25, CBK22, DS21, HD25, JDB^{+}23, KEB23, LLL^{+}22a, LM23, NMCRB21, RSEK20, Sar21, SB23, SLZ^{+}25, VLV25, WK23, vdKWH20]. application-layer [SB23]. Applications [ACG20, NLA^{+}23, SHL^{+}20, APK24, BCM^{+}25, CCVMZ24, CWL^{+}22, FD23, GRdAS23, HZG19, HZG20, KE25, LSWR20, LCA^{+}23, MR23, NCN24, PLE^{+}23, RT20, RKH^{+}22, SSC23, WK23, WPS20, ZXZ^{+}20, ZTD21, ZTY^{+}20a]. applied [NGJ^{+}22, PBdL25]. applies [ARB20]. Applying [AA21a, AMA21, LZY^{+}25a, MKC^{+}23, KS23a]. appraisals [VM21]. Approach [FEA22, LSXJ22, MAK21, NN24, NHC25, ÖEÖ22, YWX^{+}22, AGBT23, AIBKJ24, AJ20, AA22a, AN23, AAAAS20, APK24, ÁFM25, AZES21, AMM^{+}21, BC23a, BKSA25, BDC^{+}23, BAD^{+}25, dCBJSdS23, BBD23, BTAK21, BLR23, CVZHN25, CH25, CGCY21, CKV22, DS20, FXL^{+}20, FZZ^{+}25, FKvdA^{+}24, FJR23, GZS^{+}22, GAZ25, GMP24, GT24b, GL21, HZGM25, HDS21, HMV^{+}25, HHP^{+}24, HZG19, HZG20, HZ20b, HSH24, IB20, IJR^{+}25, JYS21, JK21a, JAAS^{+}23, Kac25, KKRP21, KLR^{+}25, KAM^{+}21, Kem23, KE25, KF25, KT25, KUA^{+}25, LdGW^{+}22, LXH^{+}23, LXW^{+}25b, LFHH22, LQL25b, LFW^{+}25, LHYW22, MRSP24, MNMS24, MBFIdD21, MC20, MSP^{+}20, MHSK21, MDB24, MMIS22, MTD^{+}24, NB25b, NGC^{+}25, NGX^{+}23, OSPP22, QYYH24, RFdS23, RRMSM^{+}20, REHS24, SDA^{+}24a, SU25, SAA20, SYZ^{+}23, SSK22, SPN^{+}25, SS24b, SWS^{+}25, STM^{+}20, SYL^{+}20, SPSK24, TCY^{+}20, TZZ^{+}21, TML^{+}20, TG25, TCYL21, VBS^{+}25, VS24, WNF20]. approach [WTW21, WLLS22, WWL^{+}25a, WTHL25, WRG^{+}21, WJB23, WZG24, WPAG24, WP21, XLH^{+}25, XLS24, YWZ20, YLW^{+}24b, ZCAP23, ZWDE24, ZJL^{+}20, ZLCA21, ZTD21, ZW24, ZZL^{+}22, ZH24, ZSGB^{+}22, ZTW25, dSFG20, CBK22]. Approaches [KMG21, SPZ^{+}20, WFT22, CSRA22, CDMR25, FFLQ23, FFLQ24, LSWR20, LBD^{+}23, LX21, RJ24, VDL24, WTL^{+}24]. approximate [FO25, MBH20, vDK22]. apps [CHWH25, CQLW25, GAL^{+}20a, HP23, HZ20a, KMP^{+}20, LB25, vdSFF20b]. April [Ano20a, Ano21a, Ano22a, Ano23a, Ano24a, Ano25a]. APT [JZY^{+}25, LFC^{+}25, LZY^{+}25a, LdSP21, NZZ^{+}22, RKW^{+}23, RG25, SGSS22, SN24, SHSK20, TNN^{+}22, XLWC24, YLW^{+}24b, ZHC25a]. APT-MMF [XLWC24]. APT-related [LdSP21]. Aquila [NMLN23]. Arabian [ARB20]. Arabic [PA20]. arbitrary [LMW^{+}25, PTD20]. Archimedes [RVS24]. Architecture [SSS^{+}21, AMA22, CRP22, DWX^{+}23, DPSS23, FK22, HAKK21, HZG^{+}23, HXX^{+}24, HZZ^{+}24, JWG^{+}24, JE21, KKY^{+}23, LXC^{+}20, LHYW22, PRTV22, RMVG25, dCFA^{+}23, SMCP21, SCOE23, SZL^{+}24, STZ^{+}22, Tek21, TLY^{+}24, YKG20, ZWF^{+}24, ZLY23, dPRU25, dASJ^{+}23]. architecture-based [HZZ^{+}24]. architectures [BF23, KZFM24, VAW^{+}20, ZYGL25]. Area [HMBY23, CWC^{+}25, DMB21, HAKK21, KJI^{+}24, TLKW20, UWL22, VLV25]. areas [CZCX20, OAA^{+}22]. ARES [CS24]. ARLHNIDS [SS25]. ARLHNIDS-IoT [SS25]. ARM [CJJ^{+}22, CSLK24, FDK^{+}22, JK21b]. ARM-based [FDK^{+}22]. arms [YZLS25]. ARP [MTD^{+}24]. arrival [SU25]. art [AMC23, KDE20, LWZ^{+}23, PLS20, PAA^{+}24, QS22]. artefacts [AN21]. artifacts [NRKK23, VB20]. Artificial [BC23a, GRJV24, ILVR^{+}24, JDBB20, JIG22, SDA^{+}24a, SJMP23, Wan21, XYH^{+}20, ZCZW24]. AS-level [GMH^{+}23]. ASCA [NMLN23]. ASCA-squeeze [NMLN23]. Asian [ÖEÖ22]. ASIDS [YZLS25]. ASIRDetector [ZWL^{+}25]. ASLR [Jan22]. aspect [SWK20]. aspects [SS24a]. assembly [BSG23, GAJ25, JWG^{+}24]. assess [CBK22]. assessed [KIAV22]. Assessing [AVR20, AUIH25, AMVG24, AV25, CVZHN25, FHW22, Fur22, GTP^{+}25, KAFDW22, MZW^{+}25, Olu22, PvSvdB25, RMCT23, SKKG25, SKG21, WCP^{+}23, YCMM20, ZWDE24, AAT23, CRdMM^{+}25, GMSA24, MBB^{+}23, ZZ20]. Assessment [BVB^{+}23, SM23, TCM^{+}25, TRM22, AZO^{+}23, ARS^{+}25, BCGC25a, BCGC25b, BL25, BBJ20, BKV^{+}23, BSO25, CR25, CNK25, DGH21, EFI23, ESBJ20, GB21, GG22, GMB^{+}25, GRdAS23, GKB21, HLL^{+}21, HSH24, JYF^{+}25, JNKZ25, KLJE24, LKW^{+}20, Les21, LCZY23, MYMC20, MWL23, MR23, NDDH^{+}21, OLH^{+}23, PAM^{+}24, SSHP21, SLZP24, SLZ^{+}25, SSRK20, TZZY25, TAES20, US20, WNF20, WLM25, XCWC25, YWL25, YLSJ23, YGB^{+}25, ZLY^{+}20, vdSFF20b, SP20]. assessments [SSHP21, SS23a]. asset [SDG20]. assets [Ade21, HLL^{+}21]. assistance [GYYK22, STH^{+}22]. assistant [LCA^{+}23]. assistants [BMAP24]. assisted [DKSS20, FDK^{+}22, GXL25, SBV24, TP25, TLHY25, TD24, NCM24]. associated [BJZ^{+}23, GSG23b, KMK^{+}24, TJ20]. Association [AD23, NPCM24, NZZ^{+}22, QJQ22, TDYQ23, TW20]. assumption [WS25b]. assurance [AUIH25, KR23, KJB^{+}25]. assured [TW20]. AST [ZLLY25]. ASTNN [ZLLY25]. Asymmetric [WNI20]. asynchronous [TCYL21, ZWL^{+}25]. ATDAD [YL23]. ATHITD [QYW^{+}25]. ATS [LLF^{+}23]. ATS-O2A [LLF^{+}23]. ATSDetector [WWL^{+}25a]. ATT&CK [LHC24, PL23, SSK^{+}24, WS25a, ZFS25]. attached [KLA^{+}22]. Attack [AA23, BP24, CRP25, DV24, GYG^{+}23, KK25c, LLL^{+}20b, LYL^{+}23, MSZ20, PC22, SRT25, VBCM23, WML^{+}23, WCOZ23, YLDW23, YZC25, ZGZL23, Zen23, SHM22a, AIBKJ24, AHFL25, AN23, AEA24, AÇ23, AGAP24, ACGP25, AGBB22, BHO23, BD23, BTS23b, BM24, BGL^{+}20, CJD25, CGS22, CZX^{+}21, CWZ^{+}24a, CWZ^{+}24b, CZL^{+}25, CMCVGH^{+}23, DP25, DHW24, DMB21, DS23, DGH21, DG25, DMZ24, DDW^{+}23, DCSY25, DCS24, DZZ^{+}22, GWF^{+}24, GT24b, GB23, HLC^{+}24, HLQD23, HSX^{+}24, HYW^{+}20, HNNN^{+}24, HZX^{+}21, HXZ^{+}21, HZL^{+}23, ILVR^{+}24, kJLcL24, JYWS25, KLC25, KM24b, KLR^{+}25, KXZ^{+}23, KLSD24a, KLSD24b, KD21, KS23a, KEM25b, KEK^{+}22, KL22, KL23, LLJL23, LKL23, LWS^{+}20, LMD^{+}22, LLF^{+}23, LHX23, LWYG24, LCX^{+}25, LWZL25, LYS^{+}24, LWL21, LLX^{+}23, LKLZ24, LJL^{+}25, LLX25, LPS23, LLLC22, LLL23d, MRSP24, MLM20, MZWM23, MG24a, MPM23, MLB^{+}25, MK24, MQH^{+}23, MLS^{+}23, MSL^{+}25, MT21c, NGJ^{+}22, NH21]. attack [NCW^{+}23, OFIdD^{+}20, PB25, PKK^{+}21, PJL23, PL23, QZP23, QJDD24, RJ21, RDPR25, Rs25, RKRS24, RZY^{+}24, RMVG25, RGP23, SBV24, SWS^{+}25, SSK23b, SKG21, SHSK20, SZCZ24, TSW^{+}24, TLKW20, TCM^{+}25, Tek21, TD21, VS25, WIOO24, WLQ20, WYW^{+}22, WYJ^{+}22, WYL^{+}24, WYJ^{+}24, WHW^{+}25, WCH^{+}25, WLQ^{+}25, WXZ^{+}25, WHE^{+}23, WHW^{+}24, XCY^{+}25, XSC^{+}23, XRW^{+}25, XHWL20, YKN21, YK22b, YZL^{+}20, YHY23b, YLW^{+}24b, ZKS25, ZWF^{+}24, ZLY^{+}20, ZDH^{+}23, ZXL^{+}25, ZLH22, ZM22, ZSL24, ZQCX24, ZXXG24, ZXL^{+}24, ZDM^{+}25, ZLDX25, ZNL^{+}20, ZLY^{+}21, ZD21, ZH24, ZC24, ZYGL25, ZCD^{+}22, ZZZ^{+}23, ZLL24]. attack-defense [TSW^{+}24]. Attack-specific [CRP25]. Attacker [QMZ^{+}25, GHZS23, KKB^{+}25, SS22, SB25]. attackers [LL25]. AttacKG [ZDM^{+}25]. Attacking [AKAFQ22, ZLDX25]. AttackNet [KZFM24]. Attacks [BRV^{+}22, BBR20, CCO^{+}20, HZG19, HZG20, KDBS22, KFS^{+}22, LLL^{+}23a, LBA25, NHC25, SYA^{+}21, SUC23, WLC^{+}22, XLZ^{+}22, AHC22, AFAkJS25, AMC23, ARR^{+}25, AB23, AAS24, ANG^{+}25, AMVG24, AA20, AWJB21, AQAK21, FAAF23, ABC22, AK24a, AGAS25, ADC^{+}20, AXAL24, AMM^{+}21, APPN23, BvtH^{+}23, BAD^{+}25, BMAP24, BKPZ21, BAJ25, CMMST22, CLS22, CTH^{+}22, CCS^{+}24, CDLC24, CZG^{+}22, CJA25, DG20, DCK^{+}23, DDCK22, DVAAM23, DS20, DSD^{+}25, DZZ^{+}21, ESTA^{+}25, FL21, GIO24, GGC24, Gal20b, GPC^{+}22a, GR23, GR24, GMPMS21, GVJ23, GSY^{+}20, GSLS21, GLB^{+}22, GLX^{+}23, GYZ^{+}23, HMK^{+}20, HFLZ23, HY21, HMB^{+}21, HDZ^{+}22, HXZ^{+}24, HCLRM20, HZX^{+}25, HLYZ21, HZ22, HSH24, HYZ^{+}24, HWML25, HKY^{+}25, IHJZ21, JQYL22, JN23, JGS^{+}24, JYS21, JLC25, JJW^{+}23, JYCL25, KKR^{+}24, KJCL20, KJI^{+}24, KKJ^{+}21, KHG24, KJJ^{+}24, KAO^{+}23, KJ23, KT25, KR21, LT25]. attacks [LSN^{+}21, LKA^{+}21, LLL^{+}20a, LLD^{+}25, LHZZ24, LX21, LWW^{+}21, LCL^{+}25, LWZ^{+}23, LRH^{+}21, LHS21, LWZ22, LXY^{+}24, LKLZ24, LPZF24, LGXZ22, LJO^{+}20, LLZJ22, LLPL25, LXW^{+}24, MT24, MPL^{+}25, MLC^{+}23, MCM^{+}24, MKC^{+}21, MBB^{+}23, MCC^{+}24, MH25, MAAA20, MAR^{+}23, MDB24, MMIS22, MKZ^{+}23, MTD^{+}24, NBBS22, NN24, NPF^{+}23, NKA^{+}25, NJ22, NOVJ23, NTH25, PT25, PFG^{+}24, PBG22, PAGDG24, PX21, QJQ22, QMW^{+}20a, QWW^{+}22, QY22, RBHW25, RKA25, RS23, RZ20b, RZW^{+}20, RSÁGVV23, RFdS23, RHH^{+}22, RNSS23, SKD23, SPN^{+}25, SB23, SM24, SMvH^{+}21, SYH^{+}22, SD25, SLX^{+}23, STS25, TSAG24, TA25, TWLK22, TAF24, TTP23, VLV25, VBS^{+}25, VPM23, WIOO24, WW25a, WFKH25, WLC^{+}20, Wan21, WYW^{+}22, WLLS22, WZX23, WYF^{+}24, WZC^{+}24, WW25b, WFY^{+}20, XMK21, XGS^{+}20, XHW^{+}22, YHW^{+}23, YS22, YWWH22, YZC25, YLW^{+}24b, ZHL^{+}22, ZDZD24, ZTYA25, ZLZ^{+}24, ZCJ^{+}20, ZZWF21, ZLH^{+}25, ZHC25a, ZZS^{+}23, ZCPB23]. attacks [ZZ20, dNPM^{+}21]. attainable [GMP24]. attempts [BJZ^{+}23]. Attention [GTAK25, QYW^{+}25, WDJ^{+}23, YZC25, ZJC^{+}25, BLC^{+}24, CXL^{+}24, CLN^{+}25, CCH^{+}23, CZX^{+}21, CWHH23, CXC^{+}24, DH25, GSSC24, GRJV24, HWFW25, HXZ^{+}23, HZ22, JFQ^{+}22, KLZ24, LWH^{+}24, QMW20b, RKRS24, RSÁGVV23, SS23a, SCW^{+}22, SD25, TP25, TLY^{+}24, WML^{+}23, WLY^{+}24, XXZ^{+}21, XLZ24, YLPZ21, YMZ^{+}23, ZM22, ZXXG24, ZWB^{+}25, ZLY^{+}21, ZGM21, ZSWT24]. Attention-based [QYW^{+}25, CCH^{+}23, QMW20b]. Attention-driven [YZC25]. attentional [WYJ^{+}22]. attentiono [WJP^{+}23]. attentive [LLL^{+}22a]. Attestation [BWZ^{+}25, KFZ^{+}20, KFS^{+}22, KBC21, MYW24, SSRM25]. Attribute [BGL^{+}20, XLZ^{+}22, AJ20, CZZ^{+}21, CNTBG21, DG20, FYDH25, OSTO20, TW20, WLLS22, Yan23, ZZP^{+}25]. attribute-based [AJ20, CZZ^{+}21, Yan23]. attributed [SLC^{+}21, ZLY^{+}21]. Attributes [MAAG22, HHZ21, KAM^{+}21]. attribution [Het21, PDWF25, RCD25, SN24, XLWC24]. AttriChain [SJX^{+}20]. auction [DG20, VS20]. auctions [AN20]. Audio [KN23, KKK23, TTA23, VS22]. Audit [AMNR20, CHJ22, KEK^{+}22, RSW^{+}21]. Auditing [GMPMS21, GPK21, HSQH23, HSQ24, LRL^{+}22, LZS^{+}20, RZT^{+}23, RMMP23, TWW21, TWQ^{+}25, TSCM21, YSZ24, ZJY^{+}23, ZSL25]. augmentation [AAAAS20, BFR23, DTT^{+}22, PSP^{+}22, TY22, YLY^{+}23, ZFG^{+}24]. Augmented [DWX^{+}23, FGY^{+}24, VDN24]. augmenting [WPZW25]. AugPersist [WPZW25]. August [Ano20b, Ano21b, Ano22b, Ano23b, Ano24b, Ano25b]. Australia [CZK24]. Australian3 [TRH21]. AuthCODE [SMCP21]. AuthConFormer [HZYT23]. authenticated [LHH^{+}20]. Authenticating [ISD24, UWL22]. Authentication [AA24b, AFAS25, DKSS20, GC22, HZYT23, LCP^{+}23, MAKH21, NJS^{+}24, SM23, SSS^{+}21, AGBT23, AMAA21, ASA23, AA21b, AK20, AFLR25, ABK^{+}20, ASH^{+}24, BC23a, BF20, BBA20, BH25b, BDLP24, CHH25, CRS^{+}22, CHK24, DZSCR24, FHW22, GAS^{+}23, GAZ25, GVJ23, GG21, GYL^{+}24, GLX24, GXL25, He25, HW25, JCJ^{+}21, KK25a, KKRP21, KEM^{+}25a, KR23, KKS^{+}22, KMK^{+}24, KOO24, LCL23, LZHL20, MBK25, MKL21, OO21, PPGC24, PHH^{+}20, PGZB23, PAA^{+}24, RNA22, RGPAF20, RGACPA23, SMCP21, SCBP24, SBY^{+}22, SCCZ20, SCTK23, TSKG23, WZZW20, WHPL21, WZDZ25, WBZ^{+}25, XHWL20, YM22, YWK^{+}22, YMZ^{+}23, YK24, YGG24, YK23, ZSZ^{+}25, ZM20, ZMQ21, ZXH^{+}25, ZZW^{+}25, ZGS24, ZY21]. authentication-based [GVJ23]. authentications [KHG24]. Authenticity [CQL^{+}21]. authoritative [LCZY23]. authorities [BTS23a, LSL20]. authorization [PAA^{+}24, ZCAP23, ZS24]. Authorizations [IMN22].

Authorship [BMH20]. Auto [AAB22, CVL22, TP25, YHZ^{+}22, LCZW20]. Auto-Encoder [AAB22, CVL22, TP25, YHZ^{+}22, LCZW20]. AutoCert [KR23]. Autoencoder [HMS21, KWF25, AGI^{+}25, GHMG24, SCBP24, SKD23, VS25, XKY25, ZCLW23, WQL25]. Autoencoder-based [HMS21, SKD23, VS25]. autoencoders [CPV23, FL21, MMC24, RDM20, VT25]. autoencoding [CLL24]. automata [TA25, ZGG^{+}25]. Automata-Guided [ZGG^{+}25]. Automated [ACG20, BCKN22, BGA24, GRdAS23, ILVR^{+}24, KR23, KLA^{+}22, LHC24, MV23, NKA^{+}25, PCR22, SKL23, SZL^{+}24, SA25, AGM20, AOAA20, AMNR20, AMM^{+}21, BDM^{+}24, DA20, GDK^{+}21, HCCS22, HS25a, HZL^{+}23, JK21a, JAAS^{+}23, KG20, OFIdD^{+}20, RFdS23, SHSK20, WLZ^{+}24, WBNT20, YWL25, YGB^{+}25, vGCHZ24]. Automatic [BPCZ23, CMFUA^{+}20, CZL^{+}25, CJL^{+}25, JLS22, KLSD24a, KLSD24b, MG24b, SDG20, ZQMC20, SHM22a, CLMU25, HZG^{+}23, LHL^{+}24b, MKH24, NPL^{+}23, WGRZ23, XWW^{+}20, ZCZW24]. Automatically [WPZW25, ZYL^{+}20, ZLY^{+}21, TO22]. Automating [MM21, TRM22]. Automation [DMPT25, GCS22, BDC^{+}23]. Automator [ACG20]. AutoMIC [AGI^{+}25]. automotive [KZN^{+}25, LM24, PLE^{+}23, SBJ^{+}24, dASJ^{+}23]. Autonomous [HFK^{+}24, NGC^{+}25, ARR^{+}25, HKY^{+}25, KKJ^{+}21, LQL^{+}25a, MSL^{+}25, PX21, ZC24]. AutoSeg [CJL^{+}25]. AutoVAS [JK21a]. AV [BCdGG20]. AVA [MNMS24]. AVA-MA [MNMS24]. Avaddon [YP21]. Availability [ASB^{+}21]. average [PD25a]. averaging [AGW25, Kem23]. aviation [DCS^{+}22, LSG^{+}24]. avionic [DMRV21]. avoidance [VS24, WLKT23]. AVX [KSC23]. AVX-TSCHA [KSC23]. Aware [TZZY25, AJ20, ABK^{+}20, BQA^{+}23, CPV23, CH25, DJKA25, DZSL25, DZSCR24, DLC^{+}24b, HCL^{+}23, JGS^{+}24, JB21, JLB22, LWX^{+}22, LZW^{+}24, MCS22, MS23, SYZ^{+}23, SZW^{+}25, WYL^{+}24, XWJ^{+}24, ZTY^{+}20b, ZTW25]. Awareness [HMPS20, LDW25, WMC20, AMD^{+}21, ADO23, AMA21, BMV24, BTDH20, CVZHN25, CC20, Cha24a, GMPA23, GSG23a, HHT23, HG21, HSŠ^{+}22, KIMJ23, KGIS21, MGG^{+}25, MBR23, ND23, OK23, OXNA25, RSEK20, SEGD^{+}22, SAM25, SYC^{+}24, TKS21, TAA^{+}25, VBM23, XLL^{+}25, vdSF21].

B [FL21, LWS^{+}20, LWLT21, LWT24]. BA [GB23]. Backdoor [GLB^{+}22, GB23, SYA^{+}21, SUC23, AKWR21, CDLC24, HSX^{+}24, LCX^{+}25, LLLC22, MZWM23, MG24a, MQH^{+}23, QLL^{+}25, RZY^{+}24, WZX23, WYF^{+}24, XMK21, XHW^{+}22, ZXL^{+}25, ZSL24, ZDZD24, ZZS^{+}23]. backdoors [CCC^{+}24, QLL^{+}25, SZY^{+}22]. Background [DDCK22, DHL^{+}23, GPRGMF25]. backup [LS25]. Bad [SGC23, TRH21]. Badaslr [Jan22]. Bagging [ZH24]. balanced [HS25a, MLXC24]. Balancing [FVV23, KDTR25, XCSZ21]. band [PHT^{+}25]. Bangladeshi [SRFC24]. bank [CVZHN25, VK22]. banking [KKS^{+}22, SCCZ20, WASA24]. banks [ADM23, Sma25]. Bard [RCA24]. bare [JKS^{+}25]. bare-metal [JKS^{+}25].

Barriers [GAW^{+}23, HNH^{+}22]. base [LQHZ25, ZZJC20]. Based [AAB22, BDA23, GPRGMF25, GTAK25, LSXJ22, NN24, QYJ22, QXS^{+}22, TXY^{+}23, VAW^{+}20, YWX^{+}22, YLDW23, SHM22a, ANAAK25, AGBT23, ACD25, AAT23, AHSZ21, AZO^{+}23, AHC22, AJ20, AA22a, ABAM22, ADÖU^{+}20, AEA24, AESF23, AA22b, AMM23, AK20, ATMN20, AB23, ACKS22, Alm25, ACLA22, ASAA21, AA20, AAAK23, AYS20, AZ20, AES22, AZDF20, AC25b, AGI^{+}25, ACC^{+}20, ADC^{+}20, ACD^{+}24, AOA22, BS24a, BWZ^{+}25, BH25a, BKS22, BFC^{+}25, BC23b, BEAE25, BAD^{+}25, BL25, BD23, BBD23, BBA20, BBMW21, BA20, BLR23, BGA24, BS20, BSR^{+}25, CLX21, CXL^{+}24, CTC^{+}25, CB22a, CCL^{+}22, CCH^{+}23, CRdMM^{+}25, CDMO24, CDJD24, CJD25, CLHT24, CRP22, CZ24, CYL^{+}21, CZX^{+}21, CZZ^{+}21, CTH^{+}22, CWHH23, CQC^{+}24, CXC^{+}24, CCC^{+}24, CZL^{+}25, CWYW25, CQC25, CXL^{+}25, CZH^{+}25, CSZ^{+}25, CZM^{+}25, CNK25, CKG22, CJA25, CFT22, CL20, CDMR25, CLMU25]. based [DP25, DG20, DHW24, DKSS20, DL21, DGS^{+}23, DG25, DVAAM23, DS20, DA20, DMM^{+}23, DBG24, DCH^{+}22, DLL24, DYW24, DC23, ETAB22, EYYZ20, FWF^{+}23, FHSQ20, FXZ22, FHZ^{+}22, FYL^{+}24, FFLQ23, FFLQ24, FZZ^{+}25, FRF24, FLLY24, FCA25, FD23, FSN22, FFB^{+}23, FMU22, FDK^{+}22, GGC24, GLM^{+}24, GAS^{+}23, GDM^{+}24, GYZ^{+}25a, GWL^{+}23, GAZ25, GTP^{+}25, GSS20a, GVJ23, GYZ^{+}25b, GRJV24, GZH^{+}24, GMBN21, GGZ25, GKS24, GFH^{+}22, GXZ23, GPD^{+}23, GYL^{+}24, GHZ^{+}24, GLC^{+}25, GJB22, GKC22, GYG^{+}23, HMK^{+}20, HYW^{+}21, HRFS24, HHSL20, HCL^{+}23, HHL24, HLZ^{+}25, HRU24, HDS21, HMS21, HuRMMW22, HMMW22, HLL^{+}21, HZM21, HWW^{+}25, HCLRM20, HK25, HMV^{+}25, HHP^{+}24, HWFW25, HZ20a, HZZ20, HXZ^{+}21, HZYT23, HLX^{+}24, HYZ25, HZX^{+}20, HSK22, HZG^{+}23, HXX^{+}24, HZZ^{+}24, HSH24, ISR^{+}23, IE22, IHJZ21, JRT^{+}25, JKR^{+}21, JFQ^{+}22, JWY^{+}25, JIG22, JLQ^{+}20, JZY^{+}25, JCJ^{+}21, JNKZ25, KM24a, Kac25, KLR^{+}25, KXZ^{+}23]. based [KLW^{+}23, KDDM20a, KDDM20b, KKP22, KS20, KMG21, KAK21, KIMJ23, KAU^{+}23, KEM^{+}25a, KE25, KEK^{+}21, KC22, KZYZ22, KTGDE20, KJB^{+}25, KOO24, KRR23, KSSL22, KS23b, KT25, KEK^{+}22, LTL^{+}22, LLL^{+}23a, LCC^{+}22, LLJL23, LFC^{+}25, LCZW20, LWS^{+}20, LLX^{+}20, LGL21, LL21a, LZCS22, LCZ^{+}22, LMD^{+}22, LLL^{+}22c, LLF^{+}23, LTZ23, LHX23, LSG^{+}24, LWYG24, LSTL24, LCCZ24, LHL^{+}24a, LCZ^{+}25, LQHZ25, LXX^{+}25, LPL25, zLXZrxZ25, LQY^{+}20, LS20, LWZC23, LWD^{+}25, LLLB22, LX21, hLHhLfW21, LHG^{+}24, LWZ^{+}24, LCTZ25, LCL^{+}25, LLZY20, LLLZ20, LGJW21, LZDZ21, LZZ^{+}21, LHS21, LS21, LRL^{+}22, LXZ^{+}22, LFHH22, LJJ22, LWL^{+}23, LFHX23, LXY^{+}24, LSLL24, LQL25b, LFW^{+}25, LLH^{+}25, LGG25, LJ22, LZA^{+}23, LZHL20, LZS^{+}20, LHW^{+}20, LLPL25, LWLT21, LLL23d, LWLL25, LZ20, MLM20, MSCJ21, MZW^{+}23, MLXC24, MYW24, MFA20, MZA^{+}20, MS23, MSG22, MMC24, MVVL^{+}24, MSSO^{+}25, MHE24, MBB^{+}23, MGL^{+}20, MKL21, MLQ^{+}23, MKH24, MSL^{+}25, MHC^{+}25, MWR20, MWL23, MAAA20, MCMMMA23]. based [MDB24, MMdSdS23, MMIS22, MZLK25, MHLZ22, NGJ^{+}22, NH21, NK21, NJ22, NHT^{+}25, NPL^{+}23, NCW^{+}23, NGX^{+}23, OS22, OYZ^{+}20, OSPP22, Onw20, OX22, OZW^{+}21, ÖEÖ22, Özt24, PRTV22, PMA^{+}23, PM25, PLW^{+}23, PSN^{+}22, PA20, PSP^{+}22, PMNL21, PYT20, PLW^{+}21, PGP^{+}24, PVFM^{+}21, PFR20, PGC^{+}24, PTD20, PC24, PC22, QYW^{+}25, QZT^{+}22, QMW20b, QY22, QIN25, QMC^{+}22, QMZ^{+}25, RRM20, RYS25, RPR^{+}21, RKH^{+}22, RS23, RGG^{+}24, RKRS24, RLL^{+}23, RG25, RFdS23, RGPAF20, RSEK20, RTBK21, SA20, SAR24, SAA20, SUS25, SCH^{+}20, SMCP21, SCBP24, SU21, SCE21, Sar21, SKD23, SY24, SRY^{+}25, STB^{+}20, SVA21, SWL24, SZL^{+}24, SDCL25, SLCS24, SK24a, SCW^{+}22, SGS^{+}22, SFX24, SLZ^{+}25, SEGD^{+}22, SKM21, SGJ^{+}25, SYH^{+}22, SQL^{+}24a, SQL^{+}24b, SWS^{+}25, SSGM21, SS23b, SYD^{+}25, SSK23b, STM^{+}20, SG21, SPL25, SKS21, SJ21, SZC^{+}21, SCL^{+}21, STH^{+}22]. based [SLX^{+}23, SLL23, SYC^{+}24, SZWH25, SWW^{+}25, TK25, TYAH^{+}24, TSAG24, TSKG23, TML^{+}20, TDYQ23, TQJ^{+}23, TSW^{+}24, TYZ^{+}24, TG25, TAES20, TLK^{+}24, Tek21, TY22, TA25, TW20, TCYL21, TZZ22, TZZ^{+}23, TD21, TC23, TLY^{+}24, VRS^{+}25, VS25, VAAG23, VK22, WFK25, WLQ20, WLHC21, WWH^{+}22, WYJ^{+}22, WJT^{+}22, WGRZ23, WCZW23, WTL^{+}23, WCW^{+}23, WZX23, WML^{+}23, WYP^{+}24, WZW24a, WLH^{+}24, WCZ^{+}24, WSW24, WZC^{+}24, WYJ^{+}24, WPZW25, WCH^{+}25, WW25b, WTHL25, WWS^{+}25, WLQ^{+}25, WQL25, WJB23, WK23, WBC^{+}23, WZY^{+}23, WFLX23, WWZ24, XGL20, XLH^{+}25, XGS^{+}21, XLWC24, XZZ^{+}24, XCWY25, XLY^{+}20, XLW24, XLD25, XYH^{+}20, XSK23, XX20, XWW^{+}20, XZZ^{+}25, XKJG25, XLS24, XHWL20, XKY25, XCWC25, YMR^{+}22, YHY23a, YZL^{+}20, YL20, YLS^{+}22, YWK^{+}22, YHZ^{+}22, YLL^{+}22, Yan23, YK24, YWZW24, YZLS25, YGK^{+}25, YFdCdC^{+}24, YCLZ24, YSM^{+}21, YLZ^{+}21, YKY24, YK23, YWWH22, YWW24, YSZ24, YHC25, YWL^{+}20, YLY^{+}23, YHH^{+}24, YLW^{+}24b, YLZ^{+}25, YXL^{+}22, ZPP20]. based [ZWF^{+}24, ZJL^{+}20, ZLF^{+}20, ZWW^{+}21, ZYH^{+}22, ZJW^{+}22, ZLH22, ZCLW23, ZMF^{+}23, ZX23a, ZXB^{+}24, ZSL24, ZQCX24, ZXXG24, ZTYA25, ZXH^{+}25, ZWB^{+}25, ZZP^{+}25, ZGM21, ZGS24, ZZL^{+}22, ZLY23, ZZW^{+}24, ZSWT24, ZXS^{+}25, ZY21, ZHZ^{+}22, ZZCD20, ZD21, ZFWL22, ZC24, ZSL25, ZLH^{+}25, ZHC25a, ZZJC20, ZJJS^{+}22, ZZZ^{+}23, ZSGB^{+}22, ZTJ^{+}21, ZLL24, dSFG20, dSVST23, KS23b, GZS^{+}22, RXFZ21, AHFL25, LL21b, RKH^{+}22, SPG^{+}25, SRT25, XZG^{+}25, YCCZ21, YKK24]. bases [DPKHP22]. Basic [BF23]. Batch [FYDH25, AGW25]. battle [GRJV24, BS24b]. Bayes [BEAE25, GL21, Vu22, ZX23b]. Bayes-based [BEAE25]. Bayesian [EYYZ20, FXZ22, Kem23, LZZ^{+}21, WNF20, YJK^{+}24, YQL24, dPR23]. BCIs [BCP22]. BD [BS24a, ZY21]. BD-D1Sec [ZY21]. BD-MDLC [BS24a]. BDDR [SYA^{+}21]. be [LYL^{+}23, SL20, SPV20]. beaconing [TNN^{+}22]. beans [ZXL^{+}24]. bearing [TDM^{+}21]. beats [LBD^{+}23]. BEDIM [DYL^{+}23]. Before [BRV^{+}22]. Behavior [BS24a, XLZ^{+}22, AD23, Als20, ASI^{+}20a, ASI^{+}20b, BT24, BTA24, BJZ^{+}23, CFZL22, CAT20a, CAT20b, CAT24, DYL^{+}23, FSMM24, GGR^{+}24, GWL^{+}23, GTAK25, GMLB22, GMLB23, HLP22, HP23, HKM^{+}23, KS25a, KSS20, LCL^{+}20, LRH^{+}21, LFW^{+}22, MBF25, MGZ^{+}20, NFVN^{+}22, OSPP22, SS22, SBB^{+}20, SLR25, SWK20, SKS25, TMN24, TWW21, WLL^{+}23, WLKT23, WWS^{+}25, WLQ^{+}25, WS24, WGS^{+}25b, YCLZ24, YXL^{+}22, ZGS24, ZZW^{+}24, ZHL^{+}20]. Behavioral [CSvdA^{+}23, LHAE22, AZES21, GLM^{+}24, GTGHAG22, GKC22, HW25, JNKZ25, KYKI22, MBF25, MWR20, OG23, RDHSB22, RDHS23, STSV25, TLS23]. Behaviors [QXS^{+}22, BMM22, FK23, GPC^{+}22b, KKM^{+}25, KTH20, MD22, OCB21, OCB23, SM21, VB20, YHZ23, ZLDX25]. Behaviour [vdKWH20, BMH21, Cha24a, DS21, HSSPK21, MBFIdD21, Nie24, OMO20a, SRN24, ST25b, TN21, ZSGB^{+}22]. behavioural [AES22]. behaviours [KKR^{+}24, LLL^{+}22d]. Behind [GVEV22, PB25, BLC^{+}24, MSP^{+}20, NLW24]. Beholder [LLLB22]. BeiDou [ZY21]. belief [AR21, LQHZ25, MG24b, SAJP22]. beliefs [FSMM24]. believing [ASF^{+}23, BPL^{+}20, GHZS23]. benchmark [AAAT22, JLLG24]. Benchmarking [CDMR25, WTL^{+}24]. BenchMFC [JLLG24]. benefits [vdSF23]. Benford [FCA25]. benign [YCCZ21]. Berkle [LT25]. BERT [GSG23b]. best [CCE23]. better [GMP24, JJS^{+}24, RCA24]. Between [MTN^{+}23, AD23, OSTO20, SBJ^{+}24, WMC20, YAC25]. Beware [TDM^{+}21]. Beyond [CCL^{+}22, VCP25, YAC25, BAM^{+}24, DBB^{+}24, NR22]. BFGS [ZQCX24]. BFS2Adv [HLC^{+}24]. BGP [MAP23]. BHMDC [TQJ^{+}23]. Bi [LJO^{+}20, FYDH25, GTAK25, JDBB22, YZL^{+}20]. bi-attribute [FYDH25]. Bi-directional [LJO^{+}20, GTAK25]. bi-level [JDBB22, YZL^{+}20]. bias [GAC20, LHH23, RCD25]. Bicycle [HMB^{+}21]. bid [AN20]. BiDa [RMS^{+}21]. bidding [LLJ21, SCE21]. Bidirectional [LWL^{+}23, AESF23, CXL^{+}24, CJS^{+}21, YKY24, GSG23b]. big [AD21, GYL^{+}20, KEB23, LZ20, NNA^{+}23, PRTV22, PMA^{+}23, RMS^{+}21, SGS^{+}22, UHK^{+}21, ZWZX25]. BiGRU [ZGM21]. BiLSTM [SCW^{+}22, WJP^{+}23]. BiLSTM-attentiono [WJP^{+}23]. bimodal [GKC22]. BinAIV [GSK23]. binaries [GSYM22, GSK23, YPT22]. Binary [HQL^{+}22, JFQ^{+}22, BPCZ23, BSG23, CAFMB24, DCH^{+}22, JWG^{+}24, LCTZ25, LCC^{+}21, MKL21, TC23, WJP^{+}23, WLL24, YLPZ21, ZLC^{+}23a, ZTY^{+}20b]. binary-level [DCH^{+}22]. binding [HZG19, HZG20]. BINO [BPCZ23]. BinVulDet [WJP^{+}23]. bio [AC25b]. bio-inspired [AC25b]. biomechanics [GSLS21]. Biometric [GL22, LTU^{+}21, BH25b, DNB^{+}20, GR24, GYL^{+}24, GKC22, KZFM24, MC20, RDQL25, SK20, Sar21, SSC23, SYD^{+}25, TTP20, YWK^{+}22, YWWH22, ZSZ^{+}25]. biometric-based [GKC22, Sar21]. biometrics [ASA23, Gal20b, GBG20, RDHSB22, RDHS23, SCH^{+}20]. biometrics-based [SCH^{+}20]. biosecurity [SAK^{+}24]. biostatistics [RT20]. bipartite [LXC^{+}20]. bird [AR21]. Birds [JB22]. BiRe [LJO^{+}20]. Bit [ZLC^{+}23a]. BiTCN [XLS24]. BiTCN-TAEfficientNet [XLS24]. Bitcoin [AABE20, MSMH21]. bjCnet [HHL24]. Black [HLC^{+}24, RZ24, WYW^{+}22, ZCD^{+}22, AKWR21, BM24, DP25, QY22, RZW^{+}20, WYJ^{+}22, XSC^{+}23, XT24, YS22, ZHZ^{+}22, ZLH^{+}25, Kac25]. Black-box [HLC^{+}24, RZ24, WYW^{+}22, AKWR21, BM24, QY22, RZW^{+}20, WYJ^{+}22, XSC^{+}23, XT24, YS22, ZHZ^{+}22, ZLH^{+}25]. blanks [PSP^{+}22]. BLE [GGC24]. blessing [DBB^{+}24]. Blindfold [GMY22]. Block [OJL24]. Blockchain [SPZ^{+}20, AHSZ21, AKT21, ASB^{+}21, DCB^{+}21, HZX^{+}20, KAK21, KK20, LYW^{+}21, LLJ21, MZA^{+}20, PHH^{+}20, SCE21, Sar21, SJX^{+}20, SHL^{+}20, UHK^{+}21, WSC^{+}20, YCMM20, YL20, YPDC20, ZZW^{+}21, ZWX^{+}20]. blockchain-based [AHSZ21, HZX^{+}20, KAK21, MZA^{+}20, Sar21]. blockchain-enabled [YPDC20, ZZW^{+}21]. blockchains [LLZY20]. blocking [MDR20]. blocks [MBH20, SYZ^{+}23]. bloom [SA20, SLL23, DHW24, PC22]. board [SCBP24, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o]. boardroom [GBS22]. BoAu [YLY^{+}23]. body [CWC^{+}25, HAKK21, UWL22]. Bon [SN24]. Bon-APT [SN24]. books [MFA20]. Boolean [ZYGH24]. boosted [KAU^{+}23, SPN^{+}25]. Boosting [HDZ^{+}22, MLC^{+}23, WZW^{+}24b, XY24, ZDM^{+}25, VT25]. boot [ZLY23]. BopSkyline [WZW^{+}24b]. border [GRdAS23, KDBB25, OFIdD^{+}20, PdARD24]. boring [RCD21]. bot [RT20, RRMSM^{+}20, DCK^{+}23]. both [CEN^{+}25, CWZ^{+}24b]. Botnet [KAO^{+}23, AJPK24, BS24b, CCL^{+}21, DCK^{+}23, GS24, MMIS22, NAK23, SS22, SDA^{+}24b, WXD23, WWZ24, XSK23, ZPP20, BTS23b]. Botnets [TLT22, KSSL22, MGZ^{+}25, PCK20, PMF^{+}20]. boundary [LHZZ24, MZWM23, YLY^{+}23]. Bounding [KFHM24]. bounds [CXLV20]. box [AKWR21, BM24, CLS22, HLC^{+}24, QY22, RZW^{+}20, RZ24, SZW^{+}25, WYW^{+}22, WYJ^{+}22, XSC^{+}23, XT24, YS22, ZHZ^{+}22, ZLH^{+}25, ZCD^{+}22]. BPNN [PD25a]. brain [YK23]. brain-based [YK23]. brainprint [ZZW^{+}25]. Brainwave [TSKG23]. Brainwave-based [TSKG23]. braking [LHL^{+}24b]. branch [TDYQ23]. brand [BA20]. breach [AABLS24, MGTK22, TGOF22]. Breached [GBA22]. breaches [LdGW^{+}22, Lee23, LM23, MBU24, Rou22, SS23a, Wes20]. Breaking [HZG19, HZG20, JZQ^{+}22, dCPR25, PKPC24]. breath [YK23]. bridging [ADO23, YAC25]. bring [PNK20]. Bringing [WLC^{+}22]. broad [LZG^{+}22]. broadcast [TH21]. broader [CZK24]. Browser [GZS^{+}22, SRY^{+}25, HG20, RCKB24, Zha24]. Browser-based [SRY^{+}25]. browsers [FFPC23]. browsing [FFPC22, LCL^{+}20]. BSVD [YLPZ21]. budget [LSTL24]. buffer [LZD^{+}20]. bug [ZWL^{+}25, ZFW^{+}25]. bugging [MBSE22]. bugs [LWC^{+}25]. build [PLW^{+}23]. Building [CC20, GCS22, LCZW20, RCD25, RCA20, VS24, ZY24]. buildings [KEB23]. built [CDGS23]. burglary [Hod21]. bus [AA22a, DMRV21, LSG^{+}24, ZLC^{+}23a, ZL24]. Business [AKA25, BvtH^{+}23, HDS21, PFR20, SDG20]. businesses [CCE23, CZK24, Kem23, TRH21]. Bypassing [CXLV20]. Byte [YWL^{+}20, SAR24, TQJ^{+}23]. Byte-level [YWL^{+}20]. bytes [DSN^{+}22]. Byzantine [FSCM23].

C [DYL^{+}23]. C-BEDIM [DYL^{+}23]. CAA [HWFW25]. CACAO [TFMP25]. Cache [YLDW23, GYZ^{+}23, GYW^{+}23, ZYH^{+}21]. caches [DSR23]. CAE [CVL22]. CAESAR8 [LHYW22]. calculate [EAN^{+}22]. calculation [CZM^{+}25]. calculus [HP23]. call [AZ20, AZES21, AV25, BBD23, CMMST22, CZLZ24, DS21, JWY^{+}25, RGP23, STM^{+}20, WBC^{+}23, WGS^{+}25b]. call-based [BBD23]. calls [DSN^{+}22, SAB^{+}25, SN24, SLBG21]. camera [CRS^{+}22, LWW^{+}21, WWH^{+}22, ZWW^{+}21]. camouflage [CLS22]. camouflaged [LYHC24, WZC^{+}24]. campaign [CPP24, VBM23]. campaigns [Kem23, RG25]. Can [CDG22, CFRB25, MD22, SPV20, AMD^{+}21, LM24, LYL^{+}23, LCBC23, LCL^{+}25, SL20, SCSO20, TLPY21, WZZ^{+}22, WS24, ZCPB23, AA22a, DBG24, LM24, NPL^{+}23, PJL23, SQL^{+}24a, SQL^{+}24b, TLKW20, ZLC^{+}23a, ZL24, dASJ^{+}23]. CAN-ADF [TLKW20]. can-train-and-test [LM24]. Cancelable [ZXS^{+}25, KLZ24, SLL23, YWK^{+}22]. Cancellable [LTU^{+}21, TTP20]. CANova [NPL^{+}23]. CANSat [DBG24]. CANSat-IDS [DBG24]. can’t [LCBC23]. Capabilities [BHL^{+}23, FAAF23, BCP22, HRFS24, IB20, NNA^{+}23, SVP21, SSHP21, SLBG21]. Capability [BVB^{+}23, vdKWH20, ADO23, SL20]. capable [HHL^{+}22]. capacitive [CHH25]. capacity [CFRB25]. CAPEC [RHH^{+}22, SHM22a]. capitulation [MD22]. CAPRA [TZZY25]. CapsNet [ÇÜD21]. capsule [AC25b, WZW24a, ZCW^{+}24]. capsules [ZCW^{+}24]. CAPTCHA [AA20, DA20, HCLRM20, KOO24, OZW^{+}21, PA20, ZTK^{+}23]. CAPTCHA-based [KOO24]. CAPTCHAs [SRFC24, XY24, ZXXG24]. capture [SCVB21, ZY24]. Capturing [IFHP22, WJT^{+}22]. cards [AMAA21, JZQ^{+}22]. career [GAW^{+}23]. careers [HHP^{+}24]. CarePlus [HLX^{+}24]. carrying [LL23]. carving [FXL^{+}20]. cascaded [AB23]. Cascading [CCE23]. Case [GYW^{+}23, KBD^{+}23, LKS^{+}23, NTH25, ADM23, AGM20, AMD^{+}21, BFC^{+}25, BEB25, BLR23, BKV^{+}23, BQA^{+}23, CRdMM^{+}25, FFPC22, GKB21, KJPP23, KJB^{+}25, LHL^{+}24b, Mit20, NNA^{+}23, NSLM25, NDDH^{+}21, RDM20, SBL20, WW25a, WLLC24, YK22b, LGJW21]. case-study [BFC^{+}25]. cases [Jan22]. cat [PRPS25]. Catch [BTAK21]. CatchFuzz [KL24]. Catching [XZLG25]. categorical [WZW^{+}25]. categories [COH21, LWL21]. Categorization [DKSS20, CNTBG21]. categorized [ZYL^{+}20]. causal [GYZ^{+}25a, MGZ^{+}25]. Causality [SRT25, XCY^{+}25]. Causality-preserving [SRT25]. cause [SCSO20]. cave [AGAS25]. caves [YPT22]. CAVP [JLB22]. CCBHash [OJL24]. CCNOT [XLD25]. CD [CHWH25]. CD-Net [CHWH25]. CDDA [CTC^{+}25]. CDDA-MD [CTC^{+}25]. CDNs [GMY22]. cell [ZXH^{+}25]. Cellular [GC22]. censorship [DC23]. Center [ÖEÖ22, FF23, PTD20]. centered [RGACPA23]. centers [MBR23]. centralization [LWZ^{+}25]. centralized [KZYZ22]. centre [ACRB23]. centred [VBS^{+}25]. centric [ADÖU^{+}20, BQA^{+}23, KA23, MKC^{+}21, WPAG24, YC25, Zha24]. CEP [LLLB22]. CEP-based [LLLB22]. cepstral [VAAG23]. Certificate [ZSL25, AHSZ21, KJR^{+}20]. Certificate-based [ZSL25]. Certificateless [FMU22, GPK21, HSQ24, WBZ^{+}25, HSQH23, ZY21]. certification [CDN21, JJS^{+}24, KR23, AMNR20]. certifications [CSRA22]. Certified [RCA24, JCS^{+}25]. certs [JJS^{+}24]. CETP [LHG^{+}24]. CFG [JPL20]. CFI [FDK^{+}22, JPL20]. CFOA [LWH^{+}24]. Chain [HHS22, BS20, CKL^{+}25, KKGK25, KEM25b, TDM^{+}21]. chains [Rep23, SSTRD22, ZHC^{+}25b]. Challenge [AMAA21, GVJ23]. Challenge-response [AMAA21, GVJ23]. Challenges [ATA^{+}25, BCS^{+}21, CSRA22, DCB^{+}21, GBS22, GMLB22, GMLB23, GM24, GCV22, MAKH21, PCK20, PDWF25, AWVG20, AKT21, AKA25, BC23a, BBA^{+}21, BRMH22, BCKN22, BBA20, BCdGG20, CDF^{+}20, CG24, CZZ^{+}23, DCS^{+}22, DVAAM23, DKV^{+}24, GSSP23, GG22, GMB22, HAKK21, KKAS21, KJPP23, LX25, MNMS24, MAL22, MAF^{+}23, OLH^{+}23, PNF24, SSRM25, SDA25, SEGD^{+}22, SPZ^{+}20, SCVB21, VRPTB^{+}23, WIOO24, WASA24, WTL^{+}24, YW21, ZWDE24, ZMF^{+}25, ZYL^{+}25, ZLZX24, uS25]. challenging [NTH25, SCF^{+}25]. Chameleon [CSKD22]. change [ATMN20, Cha24a, MBF25]. changed [KSVY22]. changes [WHPL21]. channel [BEFI25, BH25a, CLL^{+}21, CDN21, DGS^{+}23, DCSY25, Gal20b, GSLS21, GYZ^{+}23, GYW^{+}23, HHSL20, HWD^{+}25, HWFW25, HXZ^{+}23, HZL^{+}23, JZQ^{+}22, LLLZ21, LWSW24, LWSW25, LCTZ25, MT21c, OYZ^{+}20, SMvH^{+}21, TTP23, WSQ^{+}22, WZDZ25, XGX^{+}24, Yan20, YZLS25, ZLH^{+}25]. Channels [SOZ^{+}23a, MVH^{+}21, MLM^{+}23, vDK22]. character [ZYL^{+}22]. character-level [ZYL^{+}22]. characterising [KDE20]. characteristic [CWYW25, CJJ^{+}22]. characteristics [AK20, CYC^{+}24, KSMW23, LWP^{+}25, TKS21, UWL22]. Characterization [GMBN21, SOZ^{+}23a, YCLZ24]. Characterizing [CJC^{+}24, GPRGMF25, XWZ^{+}20]. charging [HZX^{+}25, KHG24, NTBH^{+}22]. Charting [TKKF24]. ChatGPT [OUR^{+}23, RCA24]. checking [ASAA21, CXLV20, CJC^{+}24, CDM^{+}20]. checklists [LSMK25]. chief [Ram25, SV25]. Chimp [DP25]. China [LLL^{+}22d, TLS24b]. Chinese [GHZS23, YAC25]. Chinese-NLP [GHZS23]. Chip [CWHH23, AL23]. Chip-Free [CWHH23]. choice [TYZ^{+}24]. choices [BTDH20, MDR20]. CHOP [CXLV20]. Chrome [FFPC22, PBdL25]. CIFA [LPW25]. CIMA [CCO^{+}20]. ciphers [MLM^{+}23]. circuit [XLD25]. CIS [AdSJN25]. CIST [HHS22]. cities [MZA^{+}20, MEH^{+}24]. citizens [RvdSM23, vdWLM24]. citizenship [DAA24]. city [BCKN22, RKH^{+}22]. civil [WBZ^{+}25, ZY21]. CKDAN [YMZ^{+}23]. CL [RLL^{+}23]. CL-GAN [RLL^{+}23]. Class [AJPK24, LBW^{+}25, GJB22, MC22, RRMSM^{+}20, TC23, WMK23, YL23, ZXW^{+}25]. class-conditional [WMK23]. classes [BPCZ23]. classical [GYZ^{+}23, LBD^{+}23]. Classification [AFK^{+}24, ACA^{+}23, LBW^{+}25, NLA^{+}23, YWX^{+}22, ADDH25, ADAT25, AGI^{+}25, BLC^{+}24, BEAE25, BAD^{+}24, BTS23b, BS20, CCH^{+}23, ÇÜD21, CRP22, CHWH25, CKV22, DÇÜD22, DGS^{+}23, DSD^{+}25, DA20, DBG24, DZZ^{+}21, DYW24, GAJ25, GCZ21, GY23, GMP20, GTP^{+}25, GSS^{+}20b, GJCJ20, GS24, HRFS24, HLP22, HWFW25, JLLG24, yKK22, KAO^{+}23, KL22, KN23, KL23, LLL^{+}22a, LLJL23, LXS^{+}24, LHL^{+}24a, LWSW24, LWSW25, LBD^{+}23, LHG^{+}24, MZW^{+}23, MLXC24, MK24, MAAA20, MRL22, MGG^{+}20, NKA^{+}25, NFVN^{+}22, NMCRB21, NPL^{+}23, OS22, OFH^{+}25, Onw20, PFG^{+}24, RFdS23, RHH^{+}22, RRMSM^{+}20, RKG^{+}25, RNSS23, SAR24, SAA20, SASW23, SBB^{+}20, TQJ^{+}23, TY22, TLKK20, VAW^{+}20, VMS20, VCP25, Vu22, WZCP20, WTW21, WGRZ23, WML^{+}23, WZW24a, WXZ^{+}25, WLLL24a, WBC^{+}23, WGS^{+}25b, XGS^{+}21, XKJG25, XLS24, YSM^{+}21, YWL^{+}20, ZM22, ZX23b, ZWX^{+}24, ZJJS^{+}22, ZSGB^{+}22, ZCW^{+}24]. classification-by-retrieval [ADDH25]. classifications [BH25b]. Classifier [SPG^{+}25, ACKS22, BS24a, CLX21, RDPR25]. classifiers [AKAFQ22, AKA^{+}23, CCL^{+}22, RDQL25, RG23, SHK^{+}21, ZTYA25]. Classifying [TLT22, SPN^{+}25]. clean [ZXL^{+}25]. clean-label [ZXL^{+}25]. cleaning [HYZ25]. cleansing [XMK21]. clear [MSH^{+}25]. clearer [SBV24]. CLICKA [BHWE22]. client [AN21, ACGP25, LYD20, LJO^{+}20, SM24, VWB^{+}25]. client-controlled [LYD20]. client-server [SM24]. client-side [LJO^{+}20]. clients [CKT^{+}25]. climate [GW24]. Clipping [WR24]. cloaking [SM21]. close [CRS^{+}22]. close-Up [CRS^{+}22]. closed [HCCS22, HWML25]. closed-loop [HCCS22]. Cloud [ETAB22, LSXJ22, PSN^{+}22, TJ20, YSZ24, AHKJJ21, ATMN20, ATA^{+}25, Arc20, AOA22, BC23a, BSLK^{+}25, CB20, CZZ^{+}21, CWL^{+}22, CCS^{+}24, CKL^{+}25, CWC^{+}25, CBFH20, DL21, DSC20, FGWJ20, FMAC21, GSSC24, GSS20a, GLL^{+}24, GTAK25, GPK21, GXL25, HSQ24, HKY^{+}25, IE22, KG20, KR21, LYD20, LX25, LZG^{+}22, LPX20, LLL23d, MGL^{+}20, MM23, NHT^{+}25, NM21, OO21, OZW^{+}21, PMA^{+}23, PHT^{+}25, RYS25, RPR^{+}21, RW24, RLW^{+}20, SU21, Sun21, STH^{+}22, SPSK24, TXH^{+}23, TWW21, TZZ^{+}23, TWQ^{+}25, TSCM21, TV20, WCW^{+}23, WZW^{+}24b, XHB^{+}21, XYG25, YLZ^{+}21, ZKL^{+}23, ZJY^{+}23, ZLZC22, ZSL25]. cloud-Aided [LZG^{+}22]. cloud-assisted [GXL25]. cloud-based [IE22, MGL^{+}20, NHT^{+}25, RYS25, YSZ24]. cloud-edge [CWC^{+}25, LLL23d, PHT^{+}25]. cloud-native [CKL^{+}25, ZKL^{+}23]. cloud/edge [STH^{+}22]. clouds [SBV24]. CLTD [HWML25]. CLTD-FC [HWML25]. CLTracer [ZYH^{+}22]. cluster [PTD20]. Clustered [dCFA^{+}23, MYH25]. Clustering [HNZ^{+}23, NHC25, ÁFM25, DMD^{+}20, FNZ^{+}23, HMS21, KE25, LSWR20, LFC^{+}25, LYSB23, LSTL24, LCCZ24, LCZ^{+}25, LWZC23, PLW25, SWL24, SS23b, TYAH^{+}24, WWH^{+}22, WZX23, XHTZ20]. clusters [PTD20]. clutter [UÇ22]. cluttering [UÇ22]. CNN [AGI^{+}25, BKSA25, KAU^{+}23, KZYZ22, LMD^{+}22, LWZC23, LBD^{+}23, LZHL20, NN24, NJ22, PGRS24, VAW^{+}20, XXZ^{+}21, XKY25, ZYL^{+}22, ZH24].

CNN-AutoMIC [AGI^{+}25]. CNN-Based [NN24, KZYZ22]. CNN-LSTM [BKSA25]. CNNs [HLYZ21]. Co [ZD21, JK23, SRN24, ESTA^{+}25, CL20]. Co-Forest [ZD21]. co-occurrence [SRN24]. Co-presence [CL20]. co-relation [JK23]. CO-STOP [ESTA^{+}25]. coaching [KJPP23]. Coarse [FXL^{+}20, SLCS24]. coarse-grained [SLCS24]. Coarse-to-fine [FXL^{+}20]. COBIT [MSL^{+}24]. Code [OJL24, SKL23, WZY^{+}23, ACD25, AÇ23, AGAS25, BCM^{+}25, ESM24, FWF^{+}23, FHSQ20, FHZ^{+}22, HQL^{+}22, ILVR^{+}24, JFQ^{+}22, JWG^{+}24, LS25, LCB^{+}22, LWSW24, LWSW25, LWD^{+}25, LCC^{+}21, MFA20, MKL21, MZLK25, NGM^{+}24, SCW^{+}22, SCL^{+}21, TYZ^{+}24, TTP23, WJP^{+}23, WZW24a, WXT^{+}24, YPT22, ZQMC20, ZTY^{+}20a]. code-books [MFA20]. code-reuse [TTP23]. code2vec [SLDL23]. codebases [MFD25]. coded [hYAS^{+}24]. codes [Özt24, Wan21, ZUKH23]. coding [HZM21]. coefficients [VAAG23]. coefficients-based [VAAG23]. coercion [MC20]. coevolution [JAAS^{+}23]. coexistence [OCB23]. COFELET [KMG21]. COFELET-based [KMG21]. cognition [XLL^{+}25]. Cognitive [SARG23, ATJ^{+}22, CFRB25, FKP^{+}24, LDS24, MSH^{+}25, TKS21, Yan20]. Cognitively [LZCS22, KRN24]. collaboration [LHW^{+}20]. Collaborative [LBA25, NGM^{+}24, NAK23, CWC^{+}25, GPD^{+}23, HYZ25, HS25b, IE22, JZY^{+}25, LCC^{+}22, MGL^{+}20, PHT^{+}25, SS24b, TSW^{+}24, WTHL25, YWZ20]. collected [SZCZ24]. Collecting [BHWE22, GGR^{+}24]. collection [LWO^{+}25, LZ20, SSZ^{+}24, WZW^{+}25, ZNF^{+}23]. Collective [JB22, HG21]. collector [CSZ^{+}25]. collusion [SYH^{+}22]. collusive [WFY^{+}20]. combat [MKC^{+}23]. Combating [LYHC24, WYL^{+}24]. combination [CWHH23, WLLC24, ZZL^{+}22]. combinations [ZYJ^{+}23]. combinatorial [GZS^{+}22, NK21]. combined [DCSY25, DBG24, KR23, MYH25, NZZ^{+}22, YGG24]. Combining [AGI^{+}25, GT24a, SASW23, SDA^{+}24a, ZYL^{+}22, ZCZW24]. comic [WJB23]. comic-based [WJB23]. command [MB20a, YFdCdC^{+}24]. Comments [Sar21]. commerce [MT21b, REHS24, TLS24a, TLS24b, ZHL^{+}20]. Commercial [YCWW23, KSC23, RDS20, Sma25, XCC^{+}25]. commercializing [MSL^{+}24]. commitment [MT21c]. committee [LLZY20]. committee-based [LLZY20]. commodification [WB22]. Common [KRR23, CDN21, MBH20, JCS^{+}25, KJB^{+}25]. Communication [BRV^{+}22, JP25, ARR^{+}25, ARS^{+}25, DCS^{+}22, DZ22, FGH^{+}21, HS25b, KN20, MAL22, MAR^{+}23, MF25, NSLM25, RCA23, RRS21, STS25, Wan21, WMGH^{+}23, WJB23, XCSZ21, YXL^{+}22, dASJ^{+}23]. communication-efficient [FGH^{+}21, WMGH^{+}23]. communications [ARS^{+}25, ASH^{+}24, PdARD24, RMCT23, SZC^{+}21, TGOF22]. communities [FKL23, IE22, SBY^{+}22]. community [HSK22, KDDM20a, KDDM20b, WHW^{+}25]. community-based [KDDM20a, KDDM20b]. Communizer [IE22]. companies [CCE23, GMSA24, MGTK22, MBKD21]. companion [BA20]. company [GMBV21]. comparable [IB20]. Comparative [Alm25, BVB^{+}23, RSLM25, WLLL24a, ZJW^{+}22, KS25a, LP24, OG23, SEM^{+}23, SVPM21, SGVA21, WFT22]. comparison [ACD^{+}24, BvO23, BAD^{+}24, BAC^{+}23, LBD^{+}23, MCMMMA23, ZCPB23]. competence [BBJ20]. competency [GMPA23]. competition [Arc20, HLX^{+}24]. competitions [LHGB21]. Competitive [MS23]. Compile [ZTY^{+}20a]. Compile-time [ZTY^{+}20a]. Compiler [CCO^{+}20, AHuHS24]. Compiler-Enforced [CCO^{+}20]. compilers [LXW^{+}25a]. compiling [ATGK22]. completion [JZ23]. Complex [GMP24, GTGHAG22, QZP23, WWL25b, ZMF^{+}23]. complexities [ZMF^{+}25]. Compliance [MARB^{+}23, PNK20, AD23, AS24, ABA23, BTS23a, BBZS24, CSB^{+}25, CXC22, GGK25, GRdAS23, HKM^{+}23, HG20, HS25b, KKP22, MCJPS22, MSL^{+}24, MD22, NM24, Nie24, OXNA25, PAM^{+}24, SGT24, TN21, XWY21]. compliance-eliciting [KKP22]. compliant [KTH20, MLV^{+}21]. comply [ADM23]. Components [CYL^{+}23, SCSO20]. Composable [BHO23]. composed [MAR22]. composition [ASF^{+}23, LYHC24]. Compound [OJL24]. comprehension [TYZ^{+}24, WPAG24]. Comprehensive [ACA^{+}23, AN20, GM24, LSMK25, MVH^{+}21, PNFBH23, AAAT22, AZM24, ARS^{+}25, BCH24, BCGC25a, BCGC25b, BTE23, BAJ25, CYL^{+}24, DPK20, DPSS23, ESTA^{+}25, FWF^{+}23, FZZ^{+}25, GSSP23, HSB^{+}22, HXX^{+}24, JJT^{+}24, KKAS21, KJ23, LKW^{+}20, MJ25, MAL22, MBP20, RSLM25, RKKK25, SSR^{+}24, SDA25, SAK^{+}24, SMvH^{+}21, TMJ^{+}23, TKS21, VF22, XZY^{+}24, XZZ^{+}24]. compressed [SASW23, XYG25, ZWW^{+}21]. compression [CWQZ25, JZY^{+}25, LRH^{+}21]. compressive [DABB20]. Compromise [OO21, AKA25, AASC25, BEB25, CHA^{+}24b]. Compromise-resilient [OO21]. compromised [BSLK^{+}25, MAF^{+}23]. computation [BK22, GY23, JMU^{+}25, LLX^{+}20]. computational [RBM^{+}25]. Computer [RZ20a, ZHJ^{+}23, BTAK21, LGXZ22]. computerized [KKP22, RKG20]. Computers [ABAJ25a, BCGC25a, CAT24, DJG^{+}25, FFLQ24, GMLB23, GZ20, GZG20, HuRMMW22, HZG20, KLSD24a, KL23, LCH^{+}24a, LWSW25, Ske24, GSE20]. Computing [KFHM24, ZHJ^{+}23, CSZQ23, DL21, DXZZ25, FGWJ20, FMAC21, GSSC24, GTAK25, GG21, HZZ^{+}24, IM23, KKRP21, KS25b, KR21, LMZZ22, Mit20, OO21, PMA^{+}23, PSN^{+}22, RKES23, RW24, SRCP22, SC21, SPSK24, TV20, WCW^{+}23, WLL^{+}23, WLKHD25, YWZ20, ZZW^{+}21, ZJY^{+}23, ZHJ^{+}21, ZLY23, WQL25]. Con [AKA^{+}23]. Con-Detect [AKA^{+}23]. concealed [LZL^{+}22]. Concealing [QLL^{+}25]. concealment [GWF^{+}24]. Concept [GMLB22, AJPK24, ADL25, BBBB^{+}20, CTC^{+}25, FK22, FK24, JLLG24, LaTL^{+}25, GMLB23]. Concepts [BGM^{+}23, SSRM25]. Conceptual [GGK25, Ade21, KG20, OANA25, VF22]. conceptualization [MBF25]. concern [COH21, SSL20]. concerns [GSLS21, KLC23, MB20b, QS22, TLS24a, TLS24b]. concise [BDM^{+}24]. concolic [BCD21]. concurrency [ZFW^{+}25]. Conditional [PM25, GSSC24, WMK23]. Conditions [IMN22, EFI23, QZP23, RG23]. conferencing [KLC23]. Confidence [JWY^{+}25, MCJPS22, QLZ^{+}25]. Confidence-based [JWY^{+}25]. confidentiality [MT21a]. Configuration [LCZY23, CJL^{+}25, GDG21, MR25, PCR22]. configurations [RCKB24]. Confine [RGP23]. confinement [PJL23]. conflict [SCSO20]. confusion [LHZZ24]. congested [KUA^{+}25]. CoNN [LCC^{+}22]. CoNN-IDS [LCC^{+}22]. connected [ARR^{+}25, AFAS25, AZM24, HSOL23, MSP^{+}20, PX21]. Connection [XZZ^{+}25, DG25]. Consciousness [HG21]. Consensus [AKH20, AIJ23, ASB^{+}21, MLM20, WSC^{+}20]. consensus-based [MLM20]. consent [KK20, WB22, ZZ25]. considered [HMB^{+}21]. considering [OSTO20]. Consistency [RKSV20, NCN24]. consistent [RKA25]. Consolidating [BKV^{+}23]. constellation [RTBK21]. constellation-based [RTBK21]. Constrained [SYG^{+}21, CSvdA^{+}23, QPL^{+}23, ZKS25]. constraint [AJ20, DMM^{+}23]. constraints [BBBB^{+}20, Het21, WTL^{+}21, XWW^{+}20, Yan23]. construal [FK23]. Constructing [JYWS25, LMW^{+}25, YLH^{+}24]. construction [AASWC25, HZH^{+}24, OANA25, XWW^{+}20, ZDM^{+}25, ZJH^{+}24]. constructs [KS25a]. contact [HP23]. contactless [AMAA21]. Container [SSK23a, GKB21, JYZ^{+}23, MWL23, RGP23, TCM^{+}25]. containerized [AV25, DXZZ25]. Containers [WCOZ23, HLZ^{+}25]. containers-hybrid-deployment [HLZ^{+}25]. containing [SYF^{+}21]. Content [Dub24, YMZ^{+}23, CLS22, Lee23, LMZZ22, LWH^{+}24, WCW^{+}23, Wes20]. content-based [WCW^{+}23]. Context [ABK^{+}20, CL20, SM23, TZZY25, WYL^{+}24, CVZHN25, HLZ^{+}21, JLB22, LWC^{+}25, RYS25, SZW^{+}25, WJT^{+}23, ZZCD20]. Context-Aware [TZZY25, ABK^{+}20, WYL^{+}24, JLB22, SZW^{+}25]. Context-based [CL20]. context-sensitive [LWC^{+}25]. contexts [LK20]. Contextual [CVL22, DYW24, AZ20, CZH^{+}25, FJR23, MGZ^{+}25]. Contextualising [PFR20]. contingent [HS25b]. continual [RLL^{+}23, XCWC25]. Continuous [HZYT23, LZHL20, SSS^{+}21, SCTK23, TSCM21, AGBT23, BH25b, FEA22, GAZ25, KG20, LHAE22, SMCP21, TLHY25, WZDZ25, YMZ^{+}23, ZGS24]. Continuously [WGS^{+}25a]. contour [CZX^{+}21]. contours [CZCX20]. contract [XX20]. contracts [DSC20, HZL^{+}21, PHH^{+}20, SPZ^{+}20]. Contrasting [JSYK25]. Contrastive [YWX^{+}22, DHL^{+}23, GLX^{+}23, HHL24, LHG^{+}24, MLXC24, WXT^{+}24, WWZ24]. Control [ATGK22, DAZ20, GCS22, IMN22, KDBS22, LSXJ22, LWS^{+}25, PLB^{+}22, QYJ22, SSRM25, SJH^{+}21, TXY^{+}23, US20, AHuHS24, AJ20, ABK^{+}20, AASC25, APPN23, BZE^{+}20, CPT^{+}22, CZZ^{+}21, CHJ22, DXZZ25, DLC24a, DGD23, DS20, FYL^{+}24, FMU22, FDK^{+}22, GRKS24, GR23, GWY^{+}25, GKC22, HSW25, HRU24, HKM22, HZG^{+}23, HWML25, ISR^{+}23, JYL25, JPL20, KBK24, KMH^{+}22, LZL^{+}22, LQHZ25, LZZ^{+}21, LZX^{+}23, LXY^{+}24, LCSD20, MT24, MGGR22, MSD^{+}23, MB20a, MKC^{+}21, MKC^{+}23, MGZ^{+}20, MSZ20, NJ22, OFIdD^{+}20, PKK^{+}21, QJDD24, RYS25, Rep25, SUS25, SU21, SZL^{+}24, SLCS24, TLS24a, TFJ23, TH24, VLV25, WX23, WGS^{+}25a, WP21, XZWD25, Yan23, YLW24a, YGW^{+}25, ZLDX25, ZWB^{+}25, ZLZC22, dPRU25]. control- [GR23]. Control-flow [SSRM25, JPL20]. controllable [CWC^{+}25]. controlled [BBMW21, DZSCR24, HZL^{+}23, LYD20, ZC24]. controlled-channel [HZL^{+}23]. controller [KJI^{+}24, KBAG23, LKL23, SCOE23, TLKW20, TD21, XWW^{+}20, ZXB^{+}24, HMBY23]. controller-based [TD21, ZXB^{+}24]. controls [AdSJN25, BBZS24, BPL^{+}20, CB22a, JSYK25, SSHP21, SSG25b, UYMM21]. convergence [DMPT25]. Conversational [GSG23a]. Convex [KLNW21, CXLV20]. ConvLstm [CLL24]. Convolution [GYYK22, CLL24, LWLL25, SS22, WML^{+}23, YWK^{+}22, YKY24, ZYL^{+}22]. convolution-based [YWK^{+}22]. Convolutional [HZYT23, MMdSdS23, AWN^{+}23, AC25b, AGI^{+}25, BK22, CXL^{+}24, CXC^{+}24, FEA22, GCZ21, GJCJ20, GHZ^{+}24, HFLZ23, HZG^{+}23, JQYL22, KZFM24, LWSW24, LWSW25, LFHH22, LZA^{+}23, MV25, NRS21, PYT20, QMW^{+}20a, RKRS24, SWYL20, Tek21, TY22, TP25, YMR^{+}22]. Cookie [BBD^{+}24]. cooperation [LLL23d, Sun21]. Cooperative [ZHL^{+}20, JAAS^{+}23, ZLC^{+}23b]. coordinated [ESTA^{+}25, WS22]. coordinating [OZW^{+}21]. coordination [IE22]. coot [DP25]. copy [ZSL25]. copyright [CLCZ23]. CORAL [TCM^{+}25]. CORAS [SEM^{+}23]. Corbat [DHL^{+}23]. Cordon [DXZZ25]. core [KS25a]. coreLang [KBE^{+}24]. Corona [HP23]. Corona-Warn-App [HP23]. corporate [KN20, ZWDE24]. correct [Nie24]. Correlated [LSLL24, LWO^{+}25]. correlation [CZM^{+}25, GLX^{+}23, KMH^{+}22, LTL^{+}22, MFA20, PJ23]. correlations [LCL^{+}20]. correntropy [HMK^{+}20]. Corrigendum [ATMN23, ABAJ25a, BCGC25a, FFLQ24, GMLB23, GZ20, GZG20, HuRMMW22, HZG20, KLSD24a, LCH^{+}24a, Ske24, ZHJ^{+}23]. corruption [MLZ^{+}23, NCW^{+}23]. Cortex [CSLK24]. Cortex-M [CSLK24]. cosine [NMLN23, Özt24]. Cost [SVA21, ZCJ^{+}20, AMA22, DGH21, GJB22, LQL25b, UYMM21, WLL^{+}23]. Cost-based [SVA21]. Cost-effective [ZCJ^{+}20]. cost-effectiveness [AMA22]. cost-sensitive [GJB22]. costs [BMH21, FKvdA^{+}24]. COTS [YCWW23, JZQ^{+}22]. counter [HLX^{+}24]. counterfeit [VS24]. Countering [MTD^{+}24]. countermeasure [CAT24, YCCZ21]. Countermeasures [SGSCM23, CAT20a, CAT20b, HZKN22, JZQ^{+}22, KFS^{+}22, KJ23, LLL^{+}20a, PCK20, ZYL^{+}25]. counters [dNPM^{+}21]. counting [QZS^{+}25]. Countries [ÖEÖ22, HNH^{+}22]. country [AVR20, CVZHN25, CC20, LK20]. country-level [AVR20, LK20]. covariance [LFW^{+}25]. cover [MGGR22, ZXB^{+}24]. coverage [KL24, TDM^{+}21, UYMM21, WPZW25]. coverage-based [WPZW25]. coverage-guided [KL24]. Covert [HHSL20, LCTZ25, SOZ^{+}23a, WFKH25, CLL^{+}21, CDN21, MAL22, MVH^{+}21, OYZ^{+}20, PCK20, XGX^{+}24]. COVID [HvCSJ23, KSVY22, LSN^{+}21, SAJP22, vdWLM24]. COVID-19 [HvCSJ23, KSVY22, LSN^{+}21, SAJP22, vdWLM24]. CP [SYH^{+}22]. CP-ABE [SYH^{+}22]. CPAGs [BHO23]. CPID [RSSD23]. CPMTD [HZX^{+}21]. CPPS [VLV25]. CPS [CPV23]. CPS-GUARD [CPV23]. CPU [GR23]. CRA [Bel25, GVJ23]. CRA-RPL [GVJ23]. CRACK [RCA20]. Cracks [KKGK25]. crafted [LZ25]. Crafting [BSO25]. CRAMMTS [TKKF24]. crayfish [GSSC24]. creating [FKL23, GZG19, GZG20, HK25]. credential [CKT^{+}25, MBU24]. credentials [Sar21]. credibility [MCC^{+}24]. Crime [CB22b, HKM22, LSN^{+}21]. criminal [HHP^{+}24]. crisis [RRB25]. criteria [CDN21, ZMZ^{+}25, JCS^{+}25, KJB^{+}25]. critic [CKS^{+}24, DG25, LHD^{+}23]. Critical [SSN22, YLSJ23, AWVG20, BEAE25, CZCX20, DVAAM23, MAR^{+}23, MISO^{+}23, NTH25, NMT25, RTBK21, RA25, SS24a, SKG21, TDM^{+}21, WLLC24, WLKHD25, YWPC22]. CRO [MH25]. CRO-PUF [MH25]. Cross [HXZ^{+}23, KFHM24, KMH^{+}22, LXZ^{+}23, SOZ^{+}23a, WFLX23, ZYH^{+}22, ASM20, CPR21, CRP22, DWX^{+}23, GRdAS23, GMLB22, GMLB23, JWG^{+}24, KDBB25, LTZ23, LLJ^{+}23, LXC^{+}20, PFG^{+}24, PdARD24, PJ23, TSW^{+}23, TDM^{+}21, WTC^{+}22, ZMLP23]. cross-architecture [DWX^{+}23, JWG^{+}24, LXC^{+}20]. cross-border [GRdAS23, KDBB25, PdARD24]. cross-device [GMLB22, GMLB23]. Cross-domain [KMH^{+}22, LXZ^{+}23, WTC^{+}22]. Cross-Layer [KFHM24]. Cross-Ledger [ZYH^{+}22]. cross-modal [TSW^{+}23]. cross-organizational [PJ23]. cross-platform [LLJ^{+}23]. cross-regional [ASM20]. Cross-Router [SOZ^{+}23a]. cross-sectorial [TDM^{+}21]. Cross-site [HXZ^{+}23, CPR21, PFG^{+}24]. crosstalk [HZX^{+}25]. crowdsensing [SSZ^{+}24, ZXZ^{+}20]. crowdsourcing [AIJ23, GLL^{+}24]. CRUSOE [HSŠ^{+}22]. crypto [TML^{+}20]. cryptocurrencies [QS22, WNI20]. cryptocurrency [hPRPJ23, XWZ^{+}20]. Cryptographic [BAAM23, BRM21, LsJC^{+}22]. cryptography [PMNL21]. cryptosystem [PMNL21]. cryptosystems [SSC23]. CSE [GJB22]. CSE-IDS [GJB22]. CSec [GSS20a]. CSFuzzer [SZW^{+}25]. CSI [BH25a, CLM^{+}25]. CSRF [ABC22]. CTGAN [LHL^{+}24a]. CTGAN-SVM [LHL^{+}24a]. CTI [SL20, SVP21]. CTI-SOC2M2 [SVP21]. CTIMD [CZLZ24]. Cue [YK24]. Cue-based [YK24]. cues [BHWE22, BTH22]. Cultivating [KMOC20]. culture [Als20, GMSA24, MVSSZ24, SA22, ST25b, TFP21, UNBF21, WMC20, ZWDE24, dVABH20]. culture-behaviour [ST25b]. culture-Perspectives [dVABH20]. cumulative [CZZ^{+}21]. curated [LM24]. curation [PMF^{+}20]. Current [GG22, MM25, UNBF21, VRPTB^{+}23, AMA21, BWB^{+}21, BOS^{+}21, PvSvdB24, RSLM25, SDA25]. curriculum [LFHY23]. curriculum-guided [LFHY23]. curse [DBB^{+}24]. curve [LQY^{+}20, STM^{+}20]. curves [OS22]. custom [WTL^{+}21]. CVD [WXT^{+}24]. CVE [SHM22a]. CVFP [MS23]. CWE [SHM22a]. Cyber [ATMA23, AÇ23, AA23, BHO23, CEN^{+}25, CLN^{+}25, CCE23, CCO^{+}20, CZLZ24, Da 22, Da 23, DCS^{+}22, EAN^{+}22, FWMJ^{+}23, GKB21, HMPS20, Hod21, HZX^{+}21, KMG21, KBD^{+}23, LSN^{+}21, LKS^{+}23, MTN^{+}23, NN24, NSLM25, PT25, PS21, PNFBH23, DAAM24, RCA20, SSN22, SSTRD22, TRM22, VBF21, YKG20, ZKS25, ZFS25, ACRB23, AKH24, AHFL25, AJ20, ARS^{+}25, AWB24, ASI^{+}20a, ASI^{+}20b, FAAF23, BKSA25, BFW21, BRWF23, BTE23, BFR23, BAM^{+}24, BPL^{+}20, CS23, CPT^{+}22, CYL^{+}24, CPV23, CC20, CRC^{+}23, CR25, CYL^{+}21, CCW^{+}24, CPS25, Coj25, CSM^{+}24, DMRV21, DGH21, DHL^{+}23, DSBL23, DHZ24, EÖ25, ESBJ20, EA20, GSSC24, GMRB23, GWL^{+}23, GMP24, GSG23b, HNH^{+}22, Het21, HG21, HZ20b, HSŠ^{+}22, IB20, JQYL22, JJT^{+}24, JYWS25, JLS22, JSYK25, KKR^{+}24, KGK23, KDBB25, KLSW24, KFB^{+}25, KN20, KSVY22, KEM25b, LSWR20, LFC^{+}25, LM23, LHC24]. cyber [LL25, LZZ^{+}21, MJZ^{+}25, MB20a, MBB^{+}23, MISO^{+}23, MBR23, NKA^{+}25, NR24, NJ22, NCW^{+}20, NTH25, ND23, OXNA25, PNF23, PNF24, PLN23, PL23, PF20, PBA24, PDWF25, PJ23, QJDD24, RSSD23, RKKK25, RGG^{+}24, RSÁGVV23, RK24, RRB25, SBJ^{+}24, SPN^{+}25, SAK^{+}24, SMM22, Ske22, Ske24, Sma25, SBV24, SHM22b, SD25, SKG21, SHH^{+}24, TRH21, TAES20, TA25, TWLK22, TAA^{+}25, TDM^{+}21, TLY^{+}24, UNBF21, UYMM21, VLV25, WLH^{+}24, WHLS24, XZWD25, YKN21, YK22a, YK22b, YHS24, YWL25, YC25, YGB^{+}25, ZT21, ZMLP23, ZYL^{+}20, ZLY^{+}21, ZSS21, bOJW20, dPRU25, YGB^{+}25]. cyber-attack [AHFL25]. cyber-attacks [FAAF23, LSN^{+}21, MBB^{+}23, NJ22, RSÁGVV23, SPN^{+}25]. cyber-clustering [LFC^{+}25]. cyber-crime [LSN^{+}21]. Cyber-enabled [Hod21]. cyber-industrial [TLY^{+}24]. cyber-insurance [LL25, Ske22, Ske24]. cyber-persona [RSSD23]. cyber-phishing [BFW21]. Cyber-Physical [CEN^{+}25, CLN^{+}25, CCO^{+}20, KBD^{+}23, PT25, ZKS25, BHO23, HZX^{+}21, AHFL25, AJ20, BTE23, CS23, CYL^{+}24, CPV23, HZ20b, KEM25b, LZZ^{+}21, RKKK25, SMM22, SBV24, SKG21, SHH^{+}24, TA25, TWLK22, VLV25, XZWD25, YHS24, bOJW20]. cyber-physical-social [CPT^{+}22]. cyber-relevant [PF20]. Cyber-resilience [SSN22, DSBL23, ESBJ20]. cyber-resilient [dPRU25]. Cyber-Secure [NN24]. cyber-security [EA20, KSVY22, TRH21, YKN21]. Cyber-threat [ATMA23, DAAM24, VBF21]. Cyberattack [ADAT25, BPW^{+}20, BCP22, EL22, LWS^{+}25, RKH^{+}22, VM21, ZHC^{+}25b]. cyberattacks [FKvdA^{+}24, KTGDE20, RA25]. cyberbullying [BKA20]. Cybercrime [CTV21, HF25, vdWLM24, AAFF24, Kem23]. cybercrimes [BvtH^{+}23, HKM22]. cyberdefense [ATJ^{+}22]. CyberEntRel [AKH24]. Cybersecurity [Arc20, FAAF23, BVB^{+}23, BZE^{+}20, CMLM25, CZK24, FKL23, HR24, KKJ^{+}21, LKS^{+}23, MBF25, NB25a, OJH25, RD23, SEGD^{+}22, SCVB21, AM25, AHC22, ASA^{+}22, AA24a, Alm25, Als20, APK24, AdSJN25, ARS^{+}25, BTA24, BKS22, BFC^{+}25, BCKN22, BEB25, BHW24, BBJ20, BKV^{+}23, BSO25, BSR^{+}25, CRdMM^{+}25, CG24, CSB^{+}25, CVZHN25, Cha24a, CBB25, CGCY21, CAT20a, CAT20b, CKG22, CAT24, CFRB25, CFYBF23, ESA^{+}23, FSMM24, FRF24, FWR21, FKP^{+}24, Fur21, GBS22, GMB^{+}25, GAW^{+}23, JLC25, JSYK25, JNKZ25, KT23, KMOC20, KLJE24, KRN24, KYKI22, KIMJ23, KE25, Kil25, KKM^{+}25, LHGB21, LK20, Lee23, Les21, LX25, MWM^{+}25, MMvS20, MLS^{+}23, MSL^{+}24, MAAG22, NCM24, NNA^{+}23, NMT25, ND23, OLH^{+}23, OUR^{+}23, OXNA25, PAA^{+}24, PvSvdB24, PvSvdB25, Ram25, RCD21, RCD23, RCD25, RvdSM23, RZ24, SGSCM23, SS23a]. cybersecurity [SLZP24, STSV25, SPSK24, ST25b, TKKF24, TAA^{+}25, TFMP25, VB24, VBS^{+}25, VC25, VF22, WNF20, WASA24, WLL24, WJC23, XLL^{+}25, XQS25, YQL24, YWPC22, YLSJ23, YGB^{+}25, ZM25, ZJH^{+}24, vSRW^{+}20, vdKSCY22]. Cybersecurity-focused [OJH25]. CyberShapley [MWM^{+}25]. CyberShip [SEM^{+}23]. cyberspace [SK24a, vdB24]. Cyberthreat [SL20]. Cyberthreat-Intelligence [SL20]. CyFa [RCD25, RCD25]. cynicism [OJH25]. Cyphal [HMBY23]. Czech [DPSN20].

D [RJ24, CCS^{+}24, MBB^{+}23]. D-Score [MBB^{+}23]. D1 [ZY21]. D1Sec [ZY21]. DAC [WR24]. DACH [BSFB20]. DaE2 [OA25]. damage [HMB^{+}21]. DAmpADF [DHW24]. danger [JIG22]. Dangerous [LYL^{+}23]. Danish [KA23]. DAO [GVJ23]. Dark [BBD^{+}24, CRC^{+}23]. DARK-KERNEL [CRC^{+}23]. darknet [AHS25, LLL^{+}22a, MBS20, MRL22, ZHH^{+}25, RNSS23]. DarknetSec [LLL^{+}22a]. DaST [YS22]. Data [BGM^{+}23, BP24, DAZ20, DJG^{+}25, GYG^{+}23, GBA22, KFZ^{+}20, KD21, LSXJ22, MZW^{+}25, ÖEÖ22, RDQL25, SAK^{+}24, TWW21, TZZ^{+}23, US20, WZX^{+}22, WASA24, XX20, XHB^{+}21, XCL^{+}24, ZLH^{+}25, AGW25, AD21, AC25a, AV25, ACC^{+}20, ADC^{+}20, AABLS24, BC23a, BRWF23, BFR23, BL25, BLDB24, CEO^{+}25, CEN^{+}25, CPRV21, CSEJGUA25, CZ24, CQL^{+}21, CZZ^{+}23, CKL^{+}25, CWC^{+}25, CZM^{+}25, COH21, CJA25, CMCVGH^{+}23, DJKA25, DSS^{+}25, DMRV21, DDCK22, DJG^{+}23, DS20, DZSCR24, DVJ^{+}20, DTT^{+}22, ELD24, EA20, FKvdA^{+}24, FL21, GYL^{+}20, GR23, GPAF25, GSY^{+}20, GRJV24, GRdAS23, GPK21, GMB22, GLTH21, GPD^{+}23, GSE20, HFLZ23, HZX^{+}21, HXZ^{+}21, HHL^{+}22, HYZ25, HZX^{+}20, HSQ24, JZY^{+}25, KEB23, KKY^{+}23, KMAHU24, KF25, KS23a, KEK^{+}22, LT25, LdGW^{+}22, LS25, LWS^{+}20, LL22b, LM23, LLL^{+}23b, LSG^{+}24, LWL^{+}25, zLXZrxZ25, LYD20, LX21, LdSP21, LHH^{+}20, LRH^{+}21, LHS21, LWL21, LRL^{+}22, LSLL24, LWP^{+}25, LZS^{+}20]. data [LWLT21, LWT24, LWO^{+}25, LZ20, MT24, MJ25, MZA^{+}20, MS23, MPL^{+}25, MGTK22, MLV^{+}21, MVSSZ24, MC22, MR23, MDB24, NNA^{+}23, NCW^{+}23, NM21, OPK20, PRTV22, PT25, PMA^{+}23, PdARD24, PSP^{+}22, PMF^{+}20, QJQ22, QWLL23, RYS25, RMS^{+}21, Rou22, SASW23, SBY^{+}22, SU21, SVA21, SDA25, SGS^{+}22, SYH^{+}22, SSZ^{+}24, SBL20, SSK23b, SPL25, TGOF22, TJ20, TW20, TTRY20, TWLK22, TWQ^{+}25, TLKK20, UHK^{+}21, Vu22, WZCP20, WLC^{+}20, WTL^{+}21, WWH^{+}22, WTX^{+}22, WZW^{+}25, WTL^{+}24, WFY^{+}20, WHLS24, XCC^{+}25, YZL^{+}20, YL23, YGW^{+}25, YZP^{+}25, hYAS^{+}24, YLY20, YS22, YSZ24, ZNF^{+}23, ZTR^{+}23, ZJY^{+}23, ZZP^{+}25, ZYL^{+}20, ZW24, ZYGH24, ZS24, ZLZC22, ZSL25, ZWZX25, ZSGB^{+}22, vdKWH20, vdSFF20a]. Data-driven [WZX^{+}22, KF25, LdSP21, PMF^{+}20]. Data-free [RDQL25, YS22]. data-level [ZSGB^{+}22]. Data-oriented [KFZ^{+}20, GR23, WLC^{+}20]. data-selling [LL22b]. database [LPL25]. Databases [SHM22a, Alq22, PBA24, hYAS^{+}24]. Datadriven [TWLK22]. Dataset [GMBN21, JANR23, AD21, AMC25, BSG23, CHH25, CEO^{+}25, CKK^{+}23, DLC^{+}24b, FMA22, HK25, JLLG24, KAM^{+}21, LM24, NCN24, RHH^{+}22, SLR25, SZCZ24, VC25, ZPP20]. Datasets [ACA^{+}23, DV24, GCV22, AMD^{+}24, BvO23, CPRV21, CRP25, DMB21, FBR25, GC25, JZL^{+}24, KDE20, MF25, RDHSB22, SHSK20, WFT22, XXZ^{+}21, YLL^{+}22]. datasource [LYS^{+}24]. DATE [RJ24]. dating [KMP^{+}20]. DawnGNN [FGY^{+}24]. day [BCO^{+}22, CDJD24, GAJ25, HMK^{+}20, MRSP24, MBSE22, VRS^{+}25, WHW^{+}24]. DBMS [LWC^{+}25]. DBN [CFZL22]. DBN-LSTM [CFZL22]. DBWE [DHL^{+}23]. DBWE-Corbat [DHL^{+}23]. DC [TCYL21]. DC-Adam [TCYL21]. DCRNN [PD25b]. DDOFM [WTL^{+}23]. DDoS [KHG24, WIOO24, ARR^{+}22, AHC22, AN23, ADÖU^{+}20, AGAP24, ACGP25, AOA22, CCL^{+}21, DG20, DS23, DCS24, HNNN^{+}24, HSH24, HWML25, JGS^{+}24, KKAS21, KJ23, KR21, LRH^{+}21, LHS21, LJ22, MPM23, MKZ^{+}23, NN24, NHC25, PM25, Rs25, RFdS23, SB23, SBV24, WLQ20, WW25b, ZLZ^{+}24, ZCJ^{+}20, dNPM^{+}21]. De-cluttering [UÇ22]. De-CyFa [RCD25]. De-Wipimization [OPK20]. Dealing [ZM25, vdB24]. Deanonymizing [AABE20]. decade [VWB^{+}25]. decades [EL22, ZT21]. Deceive [DTK^{+}21]. December [Ano20c, Ano21c, Ano22c, Ano23c, Ano24c, Ano25c]. Decentralized [FFB^{+}23, LSXJ22, SJX^{+}20, DLC24a, DZSCR24, HNZ^{+}23, MCS22]. Decentralizing [LKS^{+}23]. Deception [MCC^{+}24, ZFS25, BTH22, CPP24, FWMJ^{+}23, HuRMMW22, HMMW22, JJT^{+}24, KZN^{+}25, LL22b, PBA24, RGG^{+}24, VWB^{+}25, ZT21]. deception-based [HuRMMW22, HMMW22]. Decision [AQAK21, SUS25, SHM22b, ATMA23, APK24, DPK20, FRF24, GSS^{+}20b, HuRMMW22, HMMW22, HSŠ^{+}22, KXZ^{+}23, KLSW24, LHZZ24, LZDZ21, MBK25, PBG22, PL23, RMVG25, SP20, SPN^{+}25, SAM25, SWK20, TSW^{+}24, ZMF^{+}23, ZM25, ZSJ^{+}25, ZCJ^{+}20, vdKSCY22]. decision-makers [APK24, DPK20, PL23]. decision-making [ATMA23, FRF24, LZDZ21, MBK25, TSW^{+}24, ZSJ^{+}25]. Decisions [CB22b, BLC^{+}24, MCvO24, SSL20]. decoder [KEM^{+}25a, ZWF^{+}24]. decoder-based [KEM^{+}25a]. Decoding [LP24]. decompiled [WJP^{+}23]. decomposition [CCH^{+}23, HSK22, LGL21, SW22, XLZ24, SG21]. Decoy [ZFS25]. DecoyPot [SB25]. decrypt [KKB^{+}25]. Decryption [FACHC21, YP21]. deduplication [LYD20, LMZZ22, YSZ24]. Deep [AYS20, AXAL24, CRS^{+}22, CRP22, CJA25, JSK^{+}23, KLZ24, KC22, KJI25, QYJ22, RDM20, RMS^{+}22, SJMP23, YQL24, YW21, ZWZX25, AOM24, AAAT22, AKH24, AKWR21, AHC22, AN23, AÇ23, AR21, AB23, AKA^{+}23, ADAT25, AES22, AC25b, BAC^{+}23, BGA24, CPV23, CKS^{+}24, CLS22, CTH^{+}22, CWX^{+}24, CLL24, CQC25, DGS^{+}23, FD23, GIO24, GMP20, GFM^{+}22, GLB^{+}22, GKS24, GJB22, HSX^{+}24, HK25, HNNN^{+}24, HGSB24, IMMS21, ISD24, JK21a, JKR^{+}21, KM24a, KS20, KAU^{+}23, LLL^{+}22a, LLL^{+}22b, LLL^{+}22c, LLF^{+}23, LZS^{+}23, LZNZ24, LXC^{+}20, LWD^{+}25, LBD^{+}23, LDS24, LBL^{+}20, LZDZ21, MM21, MG24b, MK24, MSG22, MMC24, MLQ^{+}23, MZLK25, NHT^{+}25, NMLN23, NZZ^{+}22, OSPP22, PRTV22, PGRS24, PKK^{+}21, PYT20, PGC^{+}24, PLP^{+}21, PD25b, PAA^{+}24, QZT^{+}22, RKES23, RPR^{+}21, Rs25, RVS24, RKH^{+}22, SMCP21, SGJ^{+}25, SSG21, SYL^{+}20, TK25, TYAH^{+}24, TSAG24, TMJ^{+}23, TC23, VS25]. deep [WLHC21, WCW^{+}23, WT23, WXD23, WWW^{+}22, WHW^{+}24, XMK21, XSK23, XHW^{+}22, YHS24, YK23, YWL^{+}20, YHH^{+}24, ZXXG24, ZYL^{+}22, ZCS^{+}23, ZLH^{+}25, ZCZZ24, ZCZW24, DP25, PMA^{+}23]. Deep-discovery [SJMP23]. deep-ELM [Rs25]. deep-feature [TYAH^{+}24]. Deep-Learning [KJI25]. deep-scale [AC25b]. deepdetectnet [Sin25]. Deepdom [SWYL20]. deepfake [GT24b, KM24a]. DeepStream [HMS21]. default [RCKB24]. Defeating [JN23]. defect [HHL24, ZGM21]. defence [MKM^{+}25, NGC^{+}25]. defences [AWJB21]. Defend [LKLZ24, CWW^{+}23, MT21c]. Defending [LHZZ24, RCKB24, Wan21, XLG^{+}23, ZZS^{+}23, DG25, WFKH25, YJK^{+}24]. Defense [BPW^{+}20, KWF25, LLL^{+}23a, LZX^{+}23, LLLC22, SYA^{+}21, SUC23, AIBKJ24, AEA24, AOA22, CZ24, CZCX20, DHW24, FEA22, GVEV22, Het21, HG21, HZX^{+}21, HXZ^{+}21, HZ20b, ISR^{+}23, JYCL25, KKAS21, KKJ^{+}21, LHGB21, LKA^{+}21, LLJL23, LZDZ21, LZZX21, MPM23, MQH^{+}23, MCM^{+}24, MSL^{+}25, NOVJ23, PAGDG24, PX21, QAA23, QJDD24, QYYH24, RS23, RKW^{+}23, RGG^{+}24, RFdS23, SGSS22, TZZ^{+}21, TSW^{+}24, TV20, VRS^{+}25, YKN21, YK22b, YGC25, ZT21, ZMF^{+}23, ZXZJ25, ZCJ^{+}20, ZHC25a, ZCZZ24]. DefenseNet [CQC25]. Defenses [VBCM23, ARR^{+}25, LXY^{+}24, MB20a, QLL^{+}25, SM24, VPM23, XZY^{+}24]. defensive [APK24, IB20, SMM22]. Defined [FEA22, DJKA25, DS23, DGD23, GIO24, KKY^{+}23, KBAG23, KJI25, LKW^{+}20, LWZ22, LLL23d, MHSK21, MKH24, MTD^{+}24, SJMP23, SBV24, YPDC20]. Defining [dVABH20, GW24]. definition [Ade21, PS21]. definitions [GGK25]. degree [ZNF21]. delay [HWML25]. delayed [OCMF23]. delays [BG25]. delegated [JZL^{+}24]. delegation [ZS24]. Deletion [XCL^{+}24, AL23, TW20]. delivering [LMW^{+}25]. delivery [LMZZ22]. Delphi [CKG22]. DeMal [HSK22]. demands [OJH25]. DeMarking [YGC25]. Dempster [QMC^{+}22]. demystified [DMD^{+}20]. Demystifying [CPRV21, GLB^{+}22, MRSV21]. Deniability [RKW^{+}22]. Denial [HZ22, KHG24, AWJB21, AOA22, FEA22, GIO24, IHJZ21, LWZ22, MAAA20]. Denial-of-Service [HZ22, KHG24]. Denoising [AAB22, GLX^{+}23, ZZCD20]. Dense [YLDW23, GTAK25]. Density [LWX^{+}22, MSL^{+}25]. Density-aware [LWX^{+}22]. deobfuscation [ZLLY25]. departments [ZWDE24]. Dependable [YHW^{+}23]. dependence [WZY^{+}23]. dependencies [ARR^{+}22, MVVL^{+}24]. dependency [LWZ^{+}25, SDG20, WHW^{+}25, ZXW^{+}25]. dependent [SBY^{+}22]. Deployment [BRV^{+}22, HLZ^{+}25, PCR22, PBA24]. depth [CBFH20, MHLZ22, NTBH^{+}22, PB25, SGT24, YP21]. deresponsibilization [RvdSM23]. descent [LQHZ25]. description [BS24a, NCN24, SLBG21, WHE^{+}23]. description-based [BS24a]. descriptions [Lee23, OMO20a]. descriptive [CCBGC25, PF20]. descriptor [JK23]. Design [BDC^{+}23, GHOS22, HWML25, JYZ^{+}23, KMG21, KRN24, LTZ23, OBK^{+}23, SS23b, SRM22, TFMP25, WBNT20, XLD25, ZLY23, AOA22, CNTBG21, HZG^{+}23, IJR^{+}25, RKG20, TN21, WSW^{+}22, ZLCA21, ZCJ^{+}24]. Design-knowledge [SRM22]. designated [TXH^{+}23]. Designing [ATJ^{+}22, AK24b, KSMW23, LXLL25, AIA23, GMP24]. designs [JNKZ25]. Detect [ZY24, ADDH25, AA21a, CSvdA^{+}23, DLQ^{+}21, GIO24, JIG22, LLZJ22, MBFIdD21, PT25, STS25, TML^{+}20, ZCPB23, AKA^{+}23]. detectability [MBB^{+}23]. Detecting [AKA^{+}23, ADC^{+}20, BWB^{+}21, CQLW25, CJJ^{+}22, DC23, EMS23, FHSQ20, GZ24, HF25, JQYL22, KJI^{+}24, KT25, LWC^{+}25, MAF^{+}23, MGZ^{+}25, MKZ^{+}23, NPCM24, RFdS23, TLT22, WJP^{+}23, WZC^{+}24, WW25b, WPS20, XGL20, XGX^{+}24, YZL^{+}20, YFdCdC^{+}24, YLW^{+}24b, YLZ^{+}25, YXL^{+}22, ZZWF21, ZG24, AB23, AKCS20, AMM^{+}21, BCO^{+}22, CQC25, DLC^{+}24b, HMK^{+}20, JYS21, JPLT20, KKR^{+}24, KTGDE20, KS23b, KUA^{+}25, KR21, LZW^{+}24, LDS24, MBPCC21, MSP^{+}20, MDB24, NCN24, RT20, RLL^{+}23, SPN^{+}25, SG21, SRM22, TZZY25, VT25, WXD23, ZHH^{+}25, ZGG^{+}25, KHG24]. Detection [AESF23, AFAkJS25, AFK^{+}24, AAB22, AYL25, AA23, AK24a, BSLK^{+}25, BAB^{+}25, BAJ25, CMFUA^{+}20, CWHH23, CWW^{+}25, CYL^{+}23, DGD23, DBG24, GMBN21, GM24, KE25, KC22, KJJ^{+}24, LLL^{+}23a, LBA25, LCZW20, LCH^{+}24a, LPW25, LWT24, LWS^{+}25, MLZ^{+}23, MDB^{+}20, MAR^{+}23, NN24, NLA^{+}23, NRKK23, OPK20, PAGDG24, PC22, QYJ22, Rou22, SB23, SN24, SOZ^{+}23a, SRT25, TXY^{+}23, WDJ^{+}23, XLZ^{+}22, YWX^{+}22, YLDW23, ZJC^{+}25, ZHJ^{+}23, AJPK24, ANAAK25, AHZL25, ALAN25, ADDH25, AHC22, AA22a, AN23, ABAM22, ADJS21, AD21, AASWC25, ASSW23, AÇ23, AA22b, ATMN20, AR21, ACKS22, AKA25, AWN^{+}23, AAAK23, AYS20, AZ20, AZES21, AC25b, AMD^{+}24, AGAP24, ACGP25, AV25, AXAL24, ADL25, AOA22, AK24b, BS24a, BS24b, BKA20, BT24, BKSA25, BFW21, BEFI25, BAAM23, BFC^{+}25, BL25, dCBJSdS23, BD23, BLDB24, BBD23, BAD^{+}24, BBMW21]. detection [BTS23b, BTM21, BM24, BG25, BAC^{+}23, BA20, BTAK21, BGA24, BSG23, BTH22, CLX21, CXL^{+}24, CTC^{+}25, CLN^{+}25, CCH^{+}23, CPV23, CDJD24, CJD25, CLHT24, CRP22, CKS^{+}24, CLL^{+}21, CFZL22, CZE^{+}22, CZZ^{+}23, CZLZ24, CQC^{+}24, CWX^{+}24, CXC^{+}24, CLL24, CCW^{+}24, CKL^{+}25, CBPW25, CXL^{+}25, CZH^{+}25, CZM^{+}25, CWQZ25, CVL22, CJS^{+}21, CSKD22, CGH^{+}23, CDM^{+}20, CL20, CLMU25, CZPX22, CMCVGH^{+}23, DP25, DSN^{+}22, DAZ20, DMB21, DCK^{+}23, DH25, DVAAM23, DKV^{+}24, DBB^{+}24, DYL^{+}23, DCS24, DPSS23, ESM24, ESTA^{+}25, EA20, FWF^{+}23, FHZ^{+}22, FYL^{+}24, FKS25, FWKB24, FFLQ23, FFLQ24, FBR25, FZZ^{+}25, FGY^{+}24, FCA25, FK22, FK24, FSN22, FCSP21, FFB^{+}23, GAJ25, GLM^{+}24, GCZ21, GDM^{+}24, GYZ^{+}25a, GDG21, GAZ25, GTP^{+}25, GT24b, GL21, GSK23, GZH^{+}24, GHMG24, GMLB22, GMB22, GMLB23, GCN^{+}24, GKS24, GHZ^{+}24, GJB22, GS24, HYW^{+}21, HRFS24, HHE22]. detection [HHSL20, HFLZ23, HCL^{+}23, HMS21, HKA24, HWW^{+}25, HYW^{+}20, HCCS22, HNNN^{+}24, HGSB24, HXZ^{+}23, HLX^{+}24, HLZ^{+}21, HZG^{+}23, HXX^{+}24, IMMS21, JDBB20, JDBB22, JRT^{+}25, kJLcL24, JKR^{+}21, JWG^{+}24, JXYF25, JLQ^{+}20, JZY^{+}25, JAAS^{+}23, KM24a, KT23, KJCL20, KLC25, KM24b, KK25b, KS20, KLSW24, KAU^{+}23, KBMM23, KLP^{+}24, KFP^{+}24, KD21, KSSL22, KBAG23, KS23a, KZFM24, KN23, LTL^{+}22, LM24, LLL^{+}22b, LCC^{+}22, LL22a, LsJC^{+}22, LFC^{+}25, LKK25, LWS^{+}20, LLYL21, LCY^{+}21, LWX^{+}22, LMD^{+}22, LQH^{+}22a, LLL^{+}22c, LCZY23, LXZ^{+}23, LHD^{+}23, LSG^{+}24, LCZ^{+}24, LCH^{+}24b, LZH^{+}24, LTZ^{+}24, LaTL^{+}25, LQHZ25, LZY^{+}25a, LCW^{+}22, LWZC23, LYLT24, LLLB22, LYS^{+}24, LCTZ25, LXW^{+}25b, LSAH21, LWZ^{+}23, LLLZ20, LGJW21, LGH21, LRH^{+}21, LHS21, LWL21, LFHH22, LWZ22, LFHX23, LKLZ24, LPZF24, LQL^{+}25a, LWP^{+}25, LFW^{+}25, LLH^{+}25, LJ22, LZA^{+}23, LLPL25, LZD^{+}20, LWLT21, LLL23d, LWLL25, MRSP24, MLM20, MSCJ21, MCS22, MG24a, MT24, MVA25, MAR22, MFA20, MG24b, MR25, MK24]. detection [MMSJ25, MSG22, MMC24, MCM^{+}24, MRG21, MV25, MC20, MKC^{+}21, MFD25, MSZ20, MSL^{+}25, MC22, MF25, MDB24, MMdSdS23, MMIS22, MZLK25, MHLZ22, NKA^{+}25, NH21, NAK23, NK21, NJ22, NRS21, NB25b, NHT^{+}25, NPL^{+}23, NMLN23, NGX^{+}23, OSPP22, OCMF23, OMO20b, OFIdD^{+}20, OX22, Özt24, PRTV22, PFG^{+}24, PLW25, PMA^{+}23, PM25, PGRS24, PSY^{+}22, PCG23, PSP^{+}22, PD25a, PCR22, PYT20, PHT^{+}25, PGP^{+}24, PGC^{+}24, PLP^{+}21, PTD^{+}25, PTD20, PC24, PD25b, PRPS25, QYW^{+}25, QMW20b, QMC^{+}22, QSA24, QMZ^{+}25, RBM^{+}25, RSSD23, RDPR25, RKA25, Rs25, RVS24, RKH^{+}22, RS23, RKRS24, RG25, RMVG25, Rep23, RSW^{+}21, RRMSM^{+}20, RK24, DAAM24, RDS20, RJ24, dCFA^{+}23, SAA20, SUS25, SKD23, SY24, SRY^{+}25, SLR25, SAB^{+}25, SLC^{+}21, SZL^{+}24, SDCL25, STPB25, SK24a, SMC^{+}21, SCW^{+}22, SFX24, SKM21, SGJ^{+}25, SVPM21, ST24, Sin25, ST25a]. detection [SKKG25, SSG21, SQL^{+}24a, SQL^{+}24b, SDA^{+}24b, SHSK20, SAM25, STM^{+}20, SG21, SGA^{+}23, SPL25, SWYL20, SCL^{+}21, SOZ^{+}23b, SHH^{+}24, SCL^{+}24, SZWH25, SCC^{+}25, TK25, TSAG24, TNN^{+}22, TCY^{+}20, TQJ^{+}23, TSW^{+}23, TLKW20, TTA23, Tek21, TA25, TCYL21, TP25, TAF24, TSCM21, TC23, TLS23, TLY^{+}24, TD24, VDL24, VLV25, VC25, VS25, VPM23, VDN24, WIOO24, WFK25, WLQ20, WLHC21, WYW^{+}22, WFT22, WJT^{+}22, WMK23, WTL^{+}23, WT23, WLL^{+}23, WJT^{+}23, WYL^{+}24, WYP^{+}24, WXT^{+}24, WWWS24, WWL^{+}25a, WXJ^{+}25, WWS^{+}25, WLQ^{+}25, WQL25, WRG^{+}21, WAPA25, WS25a, WLZ^{+}23, WWLH23, WFLX23, WHW^{+}24, WWZ24, WHLS24, XMK21, XXZ^{+}21, XZZ^{+}24, XCWY25, XLY^{+}20, XLZ24, XLW24, XLL23, XFY21, XT24, XZZ^{+}25, XQS25, XKY25, YMR^{+}22, YLPZ21, YLS^{+}22, YLL^{+}22, YL23, YHW^{+}23, YHS24, YWZW24, YLH^{+}24, YWZ^{+}24, YQL24, YZLS25, YLY20, YKY24, YKSK25, YWW24, YHC25, YW21, YLY^{+}23, YHY23b, YHH^{+}24, YAC25, ZKS25, ZWF^{+}24]. detection [ZDH^{+}23, ZMF^{+}25, ZXL^{+}25, ZLF^{+}20, ZJW^{+}22, ZM22, ZLC^{+}23a, ZCLW23, ZX23a, ZHG^{+}24, ZSLY25, ZLLY25, ZWB^{+}25, ZZHY25, ZGM21, ZHJ^{+}21, ZFT24, Zha24, ZFW^{+}25, ZYL^{+}22, ZLX^{+}24, ZZW^{+}24, ZSWT24, ZSJ^{+}25, ZLH23, ZLZX24, ZD21, ZH24, ZC24, ZXW^{+}25, ZYGL25, ZCD^{+}22, ZCZW24, ZWZX25, ZDM^{+}24, ZCPB23, dSVST23, uS25, vGCHZ24, HMT^{+}25, NCM24, SPG^{+}25]. detector [AMM23, LFCD21, QZT^{+}22, WSQ^{+}22, XGS^{+}20, JYS21]. Detectors [KDBS22, HK25, KDTR25, MCMMMA23, QZP23, RDM20, YPT22, ZHZ^{+}22]. Determination [FXZ22, DVJ^{+}20, FKP^{+}24, FK23, GDK^{+}21, KMOC20]. Determining [ZZJC20, PFHB21]. Deterministic [FSCM23, DGD23]. deterrence [CXC22, KTH20, WLDW24]. detrended [NH21]. devaluation [TMN24]. develop [AMD^{+}21]. Developer [AC25a, LP24]. Developer-friendly [AC25a]. Developing [Als20, UNBF21, vdKSCY22, CVZHN25, CC20, HNH^{+}22, HG21, WZG24, ZLCA21]. Development [BBR20, BPL^{+}20, DUK25, GMSA24, KBE^{+}24, AKH20, BBJ20, CDMO24, HvCSJ23, KSS20, MVA25, MAAG22, MEH^{+}24, NJ22, NB25a, OANA25, PLS20, TCJS22]. developments [BCKN22]. deviated [LWX^{+}22]. deviation [DYL^{+}23]. Device [RRS21, SSS^{+}21, CZM^{+}25, DMRV21, DCH^{+}22, GMLB22, GMLB23, HSSPK21, HW25, KKR^{+}24, KEM^{+}25a, MB20b, NAK23, PNK20, hPRPJ23, SMCP21, SCBP24, SYZ^{+}23, SCSO20]. device-level [NAK23]. Device-to-Device [RRS21, SSS^{+}21]. Devices [BWZ^{+}25, GC22, MAKH21, ARB20, AFAS25, AYS20, BEB25, CSZQ23, CPV23, CSvdA^{+}23, CGS22, DKSS20, FDK^{+}22, GGC24, GG21, ISD24, JCJ^{+}21, JKS^{+}25, KKR^{+}24, KFZ^{+}20, KS25b, LBWV23, LZW^{+}24, LPX20, MAF^{+}23, MK24, MSP^{+}20, RZT^{+}23, RMMP23, RK24, SMvH^{+}21, UWL22, VBS^{+}25, VAAG23, WSQ^{+}22, YM22, YCWW23, ZCL^{+}25, NJS^{+}24]. DevSecOps [KG20]. DFAID [LWX^{+}22]. DFirmSan [YGK^{+}25]. DGA [AFK^{+}24, CQC25, JRT^{+}25, LCW^{+}22, TLT22, ZPP20]. DGA-based [ZPP20]. DGDP [HZGM25]. DGIDS [SQL^{+}24a]. DHCP [TD21]. DHCPguard [TD21]. diagnosis [LLX^{+}20]. diagnostic [HLL24, dNPM^{+}21]. diagram [MG24a]. Diagrams [MZW^{+}25]. DIDS [TAF24]. DIFCS [LSXJ22]. difference [JK23]. differences [BJZ^{+}23, CFRB25]. different [COH21, DGH21, HSQH23, RKKG25]. Differential [CBK^{+}20, DMB21, HZG^{+}23, LBL^{+}20, LHO^{+}25, NHC25, TSAG24, AMN22, ACC^{+}20, CZS^{+}21, FNZ^{+}23, GT24a, HZGM25, HLL^{+}25, HNZ^{+}23, JMU^{+}25, KLNW21, KEK^{+}21, LLL^{+}23b, LSTL24, LXW^{+}25a, LFL^{+}25, zLXZrxZ25, LZC24, LSLL24, LWO^{+}25, MH25, NZT^{+}25, QYYH24, QLZ^{+}25, RZ20b, SY24, SGS^{+}22, SSZ^{+}24, TZZ^{+}21, TJD23, WLLS22, WZW^{+}25, XHTZ20, XRW^{+}25, YJK^{+}24, YHY23b, ZMF^{+}23, ZZP^{+}25, ZLZC22, ZHC25a, ZTW25]. Differentially [HX21, ISM22, LGL21, YZP^{+}25, ZNF21, ZYPT22, FFB^{+}23, GSS^{+}20b, GPD^{+}23, SVA21, SZL^{+}22, ZNF^{+}23, vDK22]. differentiated [CHP^{+}24, SSG^{+}25a]. Digestive [LKA^{+}21]. DIGFuPAS [DTK^{+}21]. Digital [CCBGC25, FFPC22, FFPC23, UÇ22, AAJP25, AVR20, AA21c, EÖ25, FXL^{+}20, HLL^{+}21, KR23, KA23, KLC23, LX25, NMLN23, PCR20, Rep23, SSN22, SSK22, Sha21, SZC^{+}21, WASA24, YL20, PPGC24]. Digitalization [GHOS22]. DIHBA [MZWM23]. Dimensional [AUIH25, CB20, JN23, SLZ^{+}25, WZCP20, ZNF^{+}23]. dioid [TA25]. direct [ADM23]. directed [HZGM25, JWY^{+}25, LWZ^{+}24]. direction [LLX25]. directional [GTAK25, LWZ22, LJO^{+}20]. Directions [BRV^{+}22, MAKH21, ATA^{+}25, BBA^{+}21, BRMH22, DKV^{+}24, GCS22, KKAS21, LGXZ22, PDWF25, SWW^{+}25, ZWH^{+}25]. directive [CSB^{+}25]. Directives [CBS25, HG20]. Dis [BSO25]. Dis-empowerment [BSO25]. disabilities [FHW22]. Disarm [Dub24]. Disarming [FFLQ23, FFLQ24]. disclaimers [BBD^{+}24]. Disclose [PLB^{+}22]. disclosing [vdSF23]. Disclosure [GOW24, COH21, Esm20, JYCL25, MB20b, WS22]. disclosures [AM25]. discounting [STSV25]. discourses [AKH20]. discover [ZWL^{+}25]. Discovering [CPR21, CYL^{+}21, LTCR25, MBPCC21, XT24, GZ24]. Discovery [GOW24, AGBB22, BRMH22, CCH^{+}22, DLL24, HSK22, KEK^{+}22, LHC24, LZ20, SJMP23, XCY^{+}25, ZXZ^{+}20]. discrete [LZNZ24, QMW20b, XLZ24]. discriminability [SARG23]. discrimination [LdSP21]. Discriminator [HX21]. discussions [PLN23]. displaying [MCvO24]. disposition [MT21b]. disruption [DANS20]. dissection [ZSGB^{+}22]. Distance [FCA25, KFHM24, MHE24, AABLS24, LX21, LLX25, LFW^{+}25, OSTO20, SYD^{+}25, ZZHY25, SQL^{+}24b]. Distance-based [FCA25, MHE24, LX21]. distillation [ANAAK25, CCC^{+}24, LFHY23, DAAM24, WWWS24, ZZS^{+}23]. Distilling [JCHS24]. DISTR [MGZ^{+}25]. Distributed [DFJ^{+}23, KHG24, KJB^{+}25, OCJ20, XHTZ20, ANG^{+}25, ÁFM25, AOA22, CMG20, DLQ^{+}21, DLC24a, DBG24, FEA22, GIO24, GLZ^{+}22, HYW^{+}20, IHJZ21, JQYL22, KLNW21, LWL^{+}25, LXH^{+}25, LZX^{+}23, RDM20, SPN^{+}25, TAF24, VK22, Vu22, Yan20, YLY20]. Distributing [PDML25]. distribution [AGVA21, DTT^{+}22, DMM^{+}23, DDW^{+}23, RYS25, XGS^{+}21, ZNF21]. DivaCAN [KJI^{+}24]. dive [AOM24]. divergence [LRH^{+}21]. diverse [HN25, MRG21, MLM^{+}23, OA25, PC24]. diversification [CAMTB23, CAFMB24]. Diversity [CHP^{+}24, OSTO20, RGG^{+}24]. diversity-based [RGG^{+}24]. DIVRS [TZZ^{+}23]. dK [HZGM25]. dK-DGDP [HZGM25]. DKaaS [CRC^{+}23]. DL [AYS20]. DL-Droid [AYS20]. DMalNet [LCZ^{+}22]. DMS [CWW^{+}25]. DMSCTS [HLZ^{+}25]. DNA [NRKK23, PMNL21]. DNN [CZX^{+}21, ZSL24]. DNN-based [ZSL24]. DNS [CLL^{+}21, DHW24, LCZ^{+}24, LWZC23, NBBS22, NOVJ23, PCK20, TLS23, WSQ^{+}22, XGL20, ZYH^{+}21]. do [AR25, HKM^{+}23, MGGR22, MW20, SV25, XLL^{+}25, KFZ^{+}20]. DO-RA [KFZ^{+}20]. Docker [MWL23]. document [DNB^{+}20]. Documentation [FGY^{+}24]. documents [KLA^{+}22, MGG^{+}20]. DoD [NTH25]. DoH [MHC^{+}25]. Doing [VBS^{+}25]. Domain [CZPX22, WDJ^{+}23, AKCS20, Alq22, BLDB24, CQC25, CNK25, DJKA25, KLC25, KZN^{+}25, KBE^{+}24, KMH^{+}22, LMD^{+}22, LCZY23, LXZ^{+}23, LCW^{+}22, LCL24, PSY^{+}22, SRY^{+}25, Sin25, SWYL20, WTL^{+}23, WTC^{+}22, WWLH23, XGX^{+}24, ZHG^{+}24, ZDZD24, ZZHY25]. domain-adaptive [LMD^{+}22]. domain-invariant [WWLH23]. DomainDynamics [CNK25]. domains [HvCSJ23, LFL^{+}25]. DomEye [XGX^{+}24]. Dominant [WZX^{+}22, XSL^{+}24]. domination [Da 23]. Don’t [MGGR22, PNF24, ZZ25, PLW^{+}23]. DoS-resilient [JP25]. DouBiGRU [ZGM21]. DouBiGRU-A [ZGM21]. double [LTZ23, MCC^{+}24, SA20, ZGM21, ZCZZ24]. double-cross-based [LTZ23]. DP [YJK^{+}24]. DP-GSGLD [YJK^{+}24]. DPO [LHO^{+}25]. DPO-Face [LHO^{+}25]. draw [BWB^{+}21]. drift [AJPK24, ADL25, CTC^{+}25, FK22, FK24, GMLB22, GMB22, GMLB23, JLLG24, LWP^{+}25, LLZJ22]. Driven [AA24b, MGL^{+}20, OBK^{+}23, CWW^{+}25, GWF^{+}24, GCN^{+}24, GYZ^{+}23, HMV^{+}25, KLC25, KF25, KT25, LdSP21, LJL^{+}25, MKC^{+}23, PMF^{+}20, SVP21, SB25, TSAG24, TKKF24, TRM22, WLC^{+}20, WZX^{+}22, WWL25b, YZC25, YLW^{+}24b, ZWL^{+}25]. driver [IFHP22]. drivers [FAAF23, GBS22]. drives [BWB^{+}21]. Driving [Cha24a, CWYW25, HKY^{+}25, WLL^{+}23]. Droid [AYS20]. DroidReach [BCM^{+}25]. DroidRL [WLZ^{+}23]. drone [dPRU25]. drones [MMSJ25]. dropped [GVJ23]. Dropping [SGVA21]. DROPSYS [KJJ^{+}24]. DSA [Bel25]. DTN [ATMN23]. DTW [ACD^{+}24]. DTW-based [ACD^{+}24]. Dual [GTAK25, DPSS23, DZZ^{+}22, GLL^{+}24, HLL^{+}25, LYW^{+}21, RYS25, SCW^{+}22, SHH^{+}24, XLZ24, YMZ^{+}23, ZLY23]. dual-architecture [ZLY23]. dual-blockchain [LYW^{+}21]. dual-layer [DPSS23]. dual-stage [DZZ^{+}22]. dual-strategy [HLL^{+}25]. dueling [ZCZZ24]. dumps [MPRB21]. Duopoly [BRWF23]. Duplicitous [vdSFF20a]. duration [XLW24]. during [HvCSJ23, LSN^{+}21, NSLM25, vdWLM24]. duty [RSLM25, Yan23]. DVGAN [SHH^{+}24]. Dynamic [CRS^{+}22, HLZ^{+}25, LLL^{+}20a, LWS^{+}20, LCZ^{+}22, MZWM23, MKC^{+}21, OFIdD^{+}20, PHT^{+}25, QLZ^{+}25, SQL^{+}24a, WTL^{+}23, WR24, ZTYA25, ZLH23, AZ20, CB20, CR25, DS21, DHL^{+}23, GDG21, GMB^{+}25, GYZ^{+}25b, GPK21, GKS24, HN25, HZ20b, HZZ^{+}24, JYZ^{+}23, LQH^{+}22a, LLL^{+}22c, LWL^{+}25, LCX^{+}25, LX21, MYW24, MYS^{+}25, MHC^{+}25, NNA^{+}23, QZS^{+}25, WNI20, WLQ20, WYF^{+}24, XCC^{+}25, YWL25, YGK^{+}25, ZXB^{+}24, ZSLY25, ZYL^{+}25]. Dynamically [PRPS25]. DynamicFuzz [JWY^{+}25]. Dynamics [SHM22a, BF20, BHWE22, CSEJGUA25, FKP^{+}24, MBK25, SCTK23, SRM22, VS24, ZSL25]. DynMark [QZS^{+}25].

E-commerce [ZHL^{+}20, MT21b, REHS24, TLS24a, TLS24b]. E-DIDS [TAF24]. e-government [NDDH^{+}21]. e-learning [OXNA25]. E-mail [BA20]. e-skimmers [REHS24]. E-tailing [SST21]. e-Voting [KAK21]. E-WebGuard [ZYGL25]. ear [GAS^{+}23]. Early [LKK25, ST24, AES22, CDJD24, CJD25, KJCL20, KSSL22, MHC^{+}25, PCR22, PGC^{+}24, ST25a, TG25]. Early-stage [LKK25]. easily [ZYJ^{+}23]. EaSTFLy [DCSW20]. eavesdropping [LKL23, Wan21]. EBDM [TC23]. eBPF [YHC25]. ECC [AL23]. ECDIS [KUA^{+}25]. ECG [DABB20]. echoes [SAJP22]. eclipse [XGS^{+}20]. eclipsed [XGS^{+}20]. eco [ARS^{+}25]. eco-system [ARS^{+}25]. economic [FKvdA^{+}24]. Economics [AWB24]. economy [MBS20]. ecosystem [CMMST22, SSN22, SGT24]. ecosystems [BAM^{+}24, LWL^{+}25]. ED [LXY^{+}24]. EDAS [ÖEÖ22]. EDCO [WWH^{+}22]. Edge [DH25, GYYK22, VRPTB^{+}23, ZHJ^{+}23, AJPK24, AWN^{+}23, CSZQ23, CWC^{+}25, HN25, HSH24, KSSL22, KS25b, LMZZ22, LLL23d, PHT^{+}25, RKES23, SC21, STH^{+}22, WLL^{+}23, WXD23, ZZW^{+}21, ZHJ^{+}21, ZHC25a, KK25b, MPM23]. Edge-assistance [GYYK22]. Edge-featured [DH25]. Edge-HetIoT [MPM23]. Editorial [AMB22, CS24, bOJW20, RZ20a, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o]. Education [HMPS20, ASA^{+}22, AMA21, BTDH20, FKP^{+}24, GSG23a, NCW^{+}20, SEGD^{+}22]. EEG [BBA20]. EETC [MZW^{+}23]. effect [ADM23, DHZ24, Esm20, FVF21, GMRB23, HSSPK21, LLL^{+}23b, PvSvdB25, TLS24a, WNI20, WSW^{+}22, XWY21]. Effective [GMBN21, LYL^{+}23, SYA^{+}21, ZCZZ24, ZCZW24, ATJ^{+}22, BAD^{+}24, CAFMB24, CLCZ23, CKG22, DS23, FBR25, GL21, GMLB22, GMB22, GMLB23, HYW^{+}21, KN20, LLJL23, LFHX23, UYMM21, WLL^{+}23, WLY^{+}24, WZG24, YS22, ZKAK23, ZLLY25, ZYJ^{+}23, ZCJ^{+}20]. Effectiveness [FACHC21, AMA22, AHKJJ21, PL23, WS22, WHPL21, WCL25, ZMF^{+}25]. Effects [BT24, DANS20, MGZ^{+}20, PLB^{+}22, BTH22, COH21, HSL23, LHGB21, SAJP22, vdSF21]. efficacies [LCL^{+}25]. efficacy [KLC23]. efficiency [yKK22, KMK^{+}24, RVJ25, SCH^{+}20, WLLC24, YGG24, ZYPT22]. Efficient [AA22b, ARAG23, BK22, CWW^{+}25, CSLK24, CFT22, DSR23, DVJ^{+}20, DCSW20, JPL20, LXJ^{+}25, LZC24, MGG^{+}20, NN24, NCW^{+}23, NR22, PCR20, PHH^{+}20, QYJ22, SLCS24, WCW^{+}23, XCL^{+}24, AIBKJ24, ASSW23, ACD^{+}24, BLSS22, CTC^{+}25, CFZL22, CXL^{+}25, CSZ^{+}25, DDW^{+}23, FGWJ20, FGH^{+}21, GG21, HYZ25, JYZ^{+}23, kJLcL24, JXYF25, JE21, KBMM23, LXH^{+}25, LHH^{+}20, LGJW21, LZZX21, LFHH22, MYW24, MK24, MDKR25, OA25, RZW^{+}20, SU25, SLC^{+}21, SKS21, TD24, WZB^{+}23, WMGH^{+}23, XSC^{+}23, XCC^{+}25, YWZ20, YLZ^{+}21, ZKAK23, ZXZ^{+}20, ZM20, ZTD21, ZX23b, ZYGH24, ZZCD20, ZFWL22, ZHC^{+}25b]. efficiently [GLZ^{+}22, WZZ^{+}22]. EfficientNet [YMR^{+}22]. effort [PSY^{+}22]. Egida [PCR22]. EII [HLP22]. EII-MBS [HLP22]. elaboration [XW20]. electric [NTBH^{+}22, SSN22, KHG24]. electromagnetic [BEFI25, ISD24]. electronic [KAM^{+}21, LSL20, SHL^{+}20]. element [HG20, RTBK21]. elements [SARG23]. eliciting [KKP22, Olu22]. elimination [KAO^{+}23]. ELM [Rs25]. Elucidating [MSH^{+}25]. eluding [SGSS22]. email [AA21a, AKA25, BMH21, DPSS23, MSA24, SSR^{+}24, SAM25, VWB^{+}25]. emails [BTH22, GGR^{+}24, VWB^{+}25]. embedded [BEFI25, BQA^{+}23, CGS22, CDN21, HZZ20, HXZ^{+}23, LZW^{+}24, LHW^{+}20, NNA^{+}23, TTP23, YFdCdC^{+}24]. embedding [BBBB^{+}20, DHL^{+}23, GL21, GZH^{+}24, JB21, KS23b, LHAE22, LXZ^{+}23, LCCZ24, LXH^{+}23, LWL21, LGG25, MHC^{+}25, TLK^{+}24, ZSWT24].

Embeddings [BDA23]. embodiment [DAA24]. emergencies [CFT22]. emergency [LHL^{+}24b]. emotion [BQA^{+}23]. emotion-aware [BQA^{+}23]. Empirical [KLJE24, KAK21, CCH^{+}22, DCB^{+}21, MBPCC21, MDR20, PL23, RBHW25, RGC24, RZ24, SSHP21, TKS21, TFP21, XSL^{+}24, YPWS20, ZMF^{+}25, ZYH^{+}21, vdKWH20, AM25]. Empirically [GMRB23]. employed [PLP^{+}21, SGSS22]. Employee [RCD21, YHZ23, ADM23, ABA23, Als20, BMM22, DAA24, FSMM24, HSW25, HHZ21, KSS20, RCD23]. employee-supervisor [DAA24]. employees [CVZHN25, KGIS21, MVSSZ24, OJH25, SA22]. employing [KK25a]. Empowered [CEN^{+}25, KJI25, PPGC24, PC24, ZHL25]. Empowering [WAPA25, XCL^{+}24]. empowerment [BSO25, DPTAM25]. EMR [YSZ24]. EmuID [CJJ^{+}22]. emulated [LFW^{+}22]. emulation [CJJ^{+}22, HS25a]. EMV [AMAA21]. Enabled [AA23, CYL^{+}24, DTK^{+}21, Hod21, KM24b, MK24, NMLN23, PRTV22, RVS24, SKG21, TK25, VBS^{+}25, YPDC20, ZZW^{+}21, ZLLY25]. enablers [HNH^{+}22]. Enabling [LLJ21, NNA^{+}23]. Enclave [RRM20, GMS20]. Enclave-based [RRM20]. enclaves [BRM21]. encoded [JCHS24, STPB25]. Encoder [AAB22, GSG23b, KEM^{+}25a, CVL22, TP25, YHZ^{+}22, ZWF^{+}24, LCZW20]. encoder-decoder [ZWF^{+}24]. encoders [WYF^{+}24]. encoding [TLS23]. encompassing [OAA^{+}22]. Encrypted [DMD^{+}20, LBW^{+}25, PCK20, ÁFM25, CYC^{+}24, DVJ^{+}20, GJCJ20, KKB^{+}25, LXS^{+}24, LPL25, LBD^{+}23, LHG^{+}24, LXW^{+}25b, MZW^{+}23, MLXC24, NOVJ23, OFH^{+}25, SASW23, SYG^{+}21, WFT22, WT23, WML^{+}23, XKJG25, ZYGH24]. encryption [BAAM23, LL22b]. End [BDC^{+}23, CKL^{+}25, Ogb21, PvSvdB25, EZLC21, hPRPJ23, RDPR25, ZXXG24]. End-to-end [CKL^{+}25, RDPR25, ZXXG24]. End-User [Ogb21]. End-users [PvSvdB25]. Endorsement [JKSS22]. Endpoint [SS24b, CZZ^{+}23, OZW^{+}21]. enemy [GSG23a]. Energy [MS23, GMPA23, KMAHU24, LHH^{+}20, OCJ20, SCH^{+}20, YWZ20, SPG^{+}25]. Energy-based [SPG^{+}25]. enforce [CDG22, HCCS22]. enforceable [CCBGC25]. Enforced [CCO^{+}20]. Enforcement [CCBGC25, Yan23, LL23, RKSV20, TAF24]. Enforcing [BDLP24]. engagement [HSL23, NCPVRT23, SB25]. Engine [AYL25, BKS22]. engineered [SASW23]. engineering [Alq22, LCZ^{+}22, NSA^{+}20, VS25, VBM23, WW25a, XMK21]. engineers [HP20]. engines [LL21a, XWJ^{+}24]. English [DPSN20]. Enhance [HXZ^{+}24, ASA23, CLMU25, KF25, YHH^{+}24]. EnhanceCTI [CPS25]. Enhanced [BD23, CPS25, kJLcL24, JIG22, KJR^{+}20, SYD^{+}25, ZXH^{+}25, ZYGL25, AASWC25, BS24a, BEAE25, BBMW21, CZLZ24, FSN22, GDM^{+}24, GSK23, HLP22, KLC25, KZN^{+}25, KJI25, LHL^{+}24a, LZY^{+}25a, SAA20, SAB^{+}25, WLL24, WQL25, XKY25, ZMQ21, ZJC^{+}25]. enhancement [ACD^{+}24, HWW^{+}25, KLP^{+}24, SK20, TSW^{+}23, YKK24]. Enhancing [AGW25, ASSW23, AdSJN25, AV25, CCL^{+}24, CHP^{+}24, CWQZ25, DXZZ25, GFM^{+}22, HMBY23, HMV^{+}25, HWWB24, JYCL25, KGIS21, KZFM24, LT25, LYSB23, LZNZ24, LCL^{+}25, LXW^{+}25b, LFW^{+}22, MCM^{+}24, MYH25, Ogb21, OXNA25, RBM^{+}25, RGG^{+}24, SK24a, SSG21, SPSK24, TAA^{+}25, VDN24, WR24, WPAG24, ZXB^{+}24, ZZW^{+}25, ZFT24, AIBKJ24, HS25b, PD25b, RKA25, ZW24, ZLH23]. Enimanal [DWX^{+}23]. ening [MTN^{+}23]. enough [GCV22]. Ensemble [KL22, NCM24, TC23, ADC^{+}20, BBD23, CHP^{+}24, CSKD22, DÇÜD22, GJB22, HRFS24, KAU^{+}23, KJI^{+}24, KS23b, MMC24, MRL22, NGX^{+}23, OA25, OZW^{+}21, PBG22, PD25a, PD25b, RDPR25, VAW^{+}20, VDN24, XQS25, ZLL24, KL23]. ensemble-based [ADC^{+}20, HRFS24, KS23b]. EnsembleFool [PLW^{+}21]. ensembles [MWR20]. Ensuring [YLW24a, IJR^{+}25, SHL^{+}20]. enterprise [DYL^{+}23, HHT23, JN23, LHYW22, WS25a]. enterprises [AAFF24, HR24, KFB^{+}25, UÇ22]. entities [AKH24]. Entity [MS25, ESM24, JLC25, JMU^{+}25, WLH^{+}24, ZHC^{+}25b]. entrepreneurs [BvtH^{+}23]. entropy [LHS21, LJ22, PM25, YWWH22, ZJJS^{+}22]. entropy-based [LHS21, PM25]. entry [BKPZ21, LQH22b, WCH^{+}25]. envelope [WWH^{+}22, WYP^{+}24]. envelope-based [WYP^{+}24]. Environment [KM22, AD21, AOA22, BS24a, BSLK^{+}25, BBMW21, CPR21, CB20, CDM^{+}20, DPSN20, GTAK25, HNNN^{+}24, KKRP21, LYSB23, LGJW21, MK24, NHFB25, OXNA25, RW24, RLW^{+}20, RJ24, SZWH25, TP25, WW25a, YLW24a, ZZW^{+}24, ZCJ^{+}24]. environmental [DHZ24]. environments [AC25a, ATA^{+}25, ASAA21, BC23a, BCH24, BS25, BAC^{+}23, CFT22, DXZZ25, FMAC21, HN25, HMBY23, LLL^{+}25, LLLB22, MHC^{+}25, MRRL23, OLH^{+}23, PSCT21, QSA24, SU21, SAB^{+}25, SWW^{+}25, TWQ^{+}25, WZZW20, WZDZ25, WLM25, ZWZX25]. EPFL [WR24]. EPFL-DAC [WR24]. epidemiological [LWL^{+}23]. epistemic [Ade21]. EPPM [XSL^{+}24]. EPRN [PD25a]. EPRN-WPS [PD25a]. epsilon [LLL^{+}23b]. EQL [LFW^{+}25]. equilibrium [BS24b, WZW24a]. equipment [HLL24, KJB^{+}25]. equivalence [ATGK22, CCL^{+}24, DVJ^{+}20]. era [AA25, AA21c, MTN^{+}23]. Erasure [LS25]. ERM [KLNW21, LLL^{+}23b]. erosion [DZZ^{+}22]. ERP [ZZW^{+}25]. Erratum [KL23]. error [kJLcL24, LLZJ22, PCR22]. error-level [kJLcL24]. errors [CZZ^{+}21, ZGG^{+}25]. espionage [HW20]. EspyDroid [GAL^{+}20a]. essential [BDM^{+}24]. Establish [BHL^{+}23]. establishing [ATGK22]. establishment [SCH^{+}20]. estimate [FKvdA^{+}24]. estimation [BKS22, GZ18, GZ20, KLR^{+}25, LFL^{+}25, LX21, LS21, MSZ20]. estimators [DPSN20]. Ethereum [LXH^{+}23, XGS^{+}20, ZTJ^{+}21]. Ethereum-based [ZTJ^{+}21]. Ethical [FRF24, RCA24, MBF25]. ethics [FWR21]. Ethiopian [ADM23]. EthReview [ZTJ^{+}21]. EU [Alm25]. EU-based [Alm25]. Europe [BKV^{+}23, LK20]. European [Bel25, RRB25]. eustress [AS24]. EV [KHG24]. evade [HSL20, LL21a, YPT22]. EvadeDroid [BM24]. Evading [MBSE22, QLL^{+}25, RG23]. evaluate [AMA21]. Evaluating [AA21b, AHKJJ21, BFC^{+}25, BMV24, BAM^{+}24, HHT23, KIMJ23, MFI20, MSL^{+}24, PFHB21, PLE^{+}23, RDHSB22, SRFC24, CHK24, GMRB23, GYW^{+}23, ZYJ^{+}23]. Evaluation [BCGC25b, EÖ25, KMG21, KDBS22, ÖEÖ22, PLB^{+}22, WS25a, AHL^{+}21, BGL^{+}20, CB20, CDMR25, DCB^{+}21, GMPA23, GAC20, GT24b, JS25, KLJE24, LLJ21, LZ20, MDR20, MCMMMA23, MR23, NGJ^{+}22, NGC^{+}25, OG23, PB25, PL23, RCKB24, SLZ^{+}25, SWL^{+}25, TFMP25, VS22, WK23, WLLL24a, vdSFF20a, BCGC25a]. evaluations [BCdGG20, BG25, IB20, MMC24]. evasion [AGAS25, BM24, CAMTB23, GWF^{+}24, MWR20, QMW^{+}20a]. evasive [GPC^{+}22b, SGSS22, VCP25, ZYL^{+}25]. event [ADAT25, CSZ^{+}25, HHZ21, KDE20, yKK22, KT25, LXX^{+}25, MKC^{+}23, WWW^{+}22, XZLG25, ZHC^{+}25b]. event-driven [KT25, MKC^{+}23]. events [ALD^{+}21, BHW24]. Everything [LZL^{+}22]. Evidence [XWY21, LLH^{+}25, MSA24, MAAG22, QMC^{+}22, SEGD^{+}22, ST25b]. evidence-based [SEGD^{+}22]. evident [SSGM21]. evidential [KUA^{+}25]. evil [LJO^{+}20, LPS23]. EvilModel [WLC^{+}22]. evolution [CGBK25, HZG^{+}23, LPZF24, SY24, SRCP22]. evolution-based [HZG^{+}23]. evolutionary [MR25, SZL^{+}24, SYL^{+}20, TH24]. Evolving [JAAS^{+}23, LXS^{+}24, LQL25b, MV25, PMF^{+}20]. EVSE [Kil25]. examination [DAA24, DNB^{+}20, PF20, SSR^{+}24]. examine [HKM^{+}23]. Examining [JJS^{+}24, MT21b, NJS^{+}24, RA25, TLS24a, TN21, WMC20, MGTK22]. Example [LCH^{+}24a, CTH^{+}22, KKK23, LCH^{+}24b, LaTL^{+}25, ZHZ^{+}22]. examples [AIBKJ24, CWW^{+}23, CJC^{+}24, DTT^{+}22, HLYZ21, LCL24, LJJ22, LWH^{+}24, PLW^{+}21, QMW^{+}20a, WZZ^{+}22, WMK23, YCCZ21]. excel [BTS23a]. Exceptional [Jan22]. Exchange [Esm20, XWZ^{+}20]. excited [AC25b]. excluding [MBH20]. executable [MMdSdS23]. executables [SGA^{+}23]. execute [LCBC23]. executing [YK22a]. Execution [KM22, AC25a, BS25, BCD21, DFJ^{+}23, DS21, DCSY25, GLC^{+}25, HSL20, LWC^{+}25, LLP25, MRRL23, SBB^{+}20, TWQ^{+}25, VCP25, ZWL^{+}25, ZCJ^{+}24]. EXEmples [KDTR25]. exempt [NPCM24]. exempt-by-reputation [NPCM24]. exercise [YK22a]. exercises [BBJ20, GMP24, YKN21]. exert [HSL23]. exfiltrate [ST24]. exfiltration [CSEJGUA25, GSE20]. Exhaustive [TAF24]. Existing [MAF^{+}23]. Expanding [HRFS24, Ske22, Ske24]. expectancy [ADO23]. expectancy-capability [ADO23]. expectation [ADO23]. expectations [FWMJ^{+}23, PFHB21]. expected [JSYK25]. Experience [LZY^{+}25b, DPSN20, TLS24a]. Experiences [FWMJ^{+}23, KFB^{+}25]. experiment [BT24, MB20b, SPV20]. experimental [AA20, CDMR25, DBB^{+}24, EFI23, PB25, RKW^{+}23, TN21]. experimentation [ATJ^{+}22]. experiments [BB20]. Expert [ZZ20, FWMJ^{+}23, GCN^{+}24, GHZ^{+}24, LZS^{+}23, MBB^{+}23, PLN23]. expert-based [MBB^{+}23]. expertise [Da 23]. experts [BTS23a, TAA^{+}25]. explain [MD22]. Explainability [GLM^{+}24, KFP^{+}24, SN24]. Explainable [GWL^{+}23, GKS24, SAR24, WFK25, XKJG25, ANAAK25, AHZL25, AMM23, AKAFQ22, GYZ^{+}25a, KLC25, SGJ^{+}25, GT24b]. Explaining [BvtH^{+}23, MVSSZ24, VB20, VBM23, SW22]. Explanation [MWM^{+}25, SL20, ZSJ^{+}25]. explanations [MMdR24]. Explanatory [KKM^{+}25, Zen23]. explicit [WJT^{+}22, XLW24]. exploit [LL25]. Exploitability [WCP^{+}23]. Exploitable [WCP^{+}23, YLZ^{+}25]. exploitation [GDK^{+}21, Jan22, LPS23, MM21, PKK^{+}21, RVJ25]. Exploiting [DMRV21, HSL20, LLLZ21, LLZJ22]. exploits [XLG^{+}23]. Exploration [WCP^{+}23, LET^{+}25, RPR^{+}21]. exploratory [HNH^{+}22, MT21b, PF20]. Exploring [AS24, AHuHS24, ASM20, APK24, BCM^{+}25, BSFB20, CHH25, FKP^{+}24, HS25b, JR24, KDBB25, Kem23, KSMW23, LWH^{+}24, MSA24, NM24, Ogb21, OCB21, BSG23, NNA^{+}23, VM21, ZZ25]. Exponential [PMA^{+}23]. Exposure [SKS25]. Expressive [IMN22]. ExpSSOA [PMA^{+}23]. ExpSSOA-Deep [PMA^{+}23]. Exsense [GLTH21]. Extended [JK23, KKY^{+}23, BvtH^{+}23, GMH^{+}23, MZW^{+}23, NLW24, NTH25, QAA23]. Extending [AMD^{+}24, GYZ^{+}23, RTBK21, WNF20, GOW24]. extensible [Rep25]. extension [AlD25, PBdL25]. extensions [KSC23, SLCS24]. extensive [JR24]. extortion [MCC^{+}24]. extra [FK23, YHZ23]. extra-role [FK23, YHZ23]. Extract [GLTH21]. Extracting [MB20b, ZYL^{+}20]. extraction [AKH24, CZL^{+}25, CPS25, GLH^{+}23, HLZ^{+}21, JLS22, KS20, LQH^{+}22a, LHZZ24, LP24, NKA^{+}25, NHT^{+}25, PD25b, SLR25, SWL24, SZCZ24, WLL24, WLH^{+}24, YWZ^{+}24, ZWB^{+}25]. extractions [MS25]. extractor [SYD^{+}25]. extractors [AFLR25]. Extreme [MC22, AAAK23, JHD^{+}20, KR21, SY24, VT25]. extremely [ZHH^{+}25]. eye [BT24]. eye-tracking [BT24]. eyes [RCD21].

Fabric [JKSS22, LZS^{+}20]. Face [CBK^{+}20, CCH^{+}23, CRS^{+}22, DMM^{+}23, kJLcL24, JK23, RMS^{+}22, RXFZ21, STH^{+}22, TLK^{+}24, WWLH23, LHO^{+}25]. Facebook [vdSFF20b]. faceted [ESTA^{+}25]. facial [BC23b, BQA^{+}23, GSLS21, LHO^{+}25, OFIdD^{+}20, SAAW21, TZZ22]. FACILE [ZCW^{+}24]. facilitate [RMMP23]. Facilitating [PKPC24]. facility [KJR^{+}20]. Facing [FL21]. Factor [WZX^{+}22, GGR^{+}24, GLX24, KK25a, KMK^{+}24, SCCZ20, WZZW20, WLKHD25, YK24]. factorial [HSW25]. factorization [LSAH21]. Factors [CBB25, MARB^{+}23, AWVG20, CPP24, DPK20, FJR23, LK20, LBL^{+}20, NM24, NLW24, NCW^{+}20, NMT25, PKPC24, RGC24, RA25, TFP21, YWPC22, YLSJ23]. facts [BDM^{+}24]. fail [BTS23a]. fails [ADO23]. failures [AAJP25, WZZW20]. Fair [BAC^{+}23, DSC20, LLZY20, MCMMMA23, WNF20]. fairly [ZYJ^{+}23]. Fairness [TZZ22, GTL^{+}22, LZRR23]. fake [KFP^{+}24, NCPVRT23, RDS20]. fallback [AA21b]. Falling [BJZ^{+}23, BTH22]. False [BP24, GSY^{+}20, HFLZ23, ASSW23, ADC^{+}20, HZX^{+}21, HXZ^{+}21, KHG24, LX21, MT24, PT25, RG23, TWLK22, VK22, YZL^{+}20]. FAMCF [ZWX^{+}24]. Families [TLT22]. Family [JANR23, BLC^{+}24, CDM^{+}20, CKV22, HLP22, IMMS21, JLLG24, SAR24, WTW21, YGC^{+}23, ZTYA25, ZLZ^{+}24, ZWX^{+}24]. fans [GSE20]. Fansmitter [GSE20]. farming [BAM^{+}24]. Fast [CAFMB24, CQC^{+}24, KXZ^{+}23, YS22, ADJS21, CBPW25, LGH21, PLW25]. Faster [HYZ25, HLYZ21]. FastReach [QWLL23]. FastText [CQC25]. fatigue [RCD23, RCD25, WYL^{+}24]. Fault [YGG24, FSCM23, PJL23, RLW^{+}20, SMvH^{+}21]. Fault-tolerant [YGG24, RLW^{+}20]. FaultMorse [HZL^{+}23]. faults [GVEV22]. favorable [ALD^{+}21]. FC [HWML25, ZWZX25]. FC-Trans [ZWZX25]. FCEVAL [ZYJ^{+}23]. FCSCNN [KZYZ22]. FD [dASJ^{+}23]. FDN [VS25]. FDN-SA [VS25]. FDT [LCL24]. FE [YS22]. FE-DaST [YS22]. fear [AS24, AAFF24, JDB^{+}23, VM21]. feasibility [KLR^{+}25]. feasible [CZG^{+}22]. feather [JB22]. Feature [BS24b, FK22, GYZ^{+}25b, KZYZ22, LdSP21, LXW^{+}24, SCW^{+}22, TP25, TRM22, WT23, WLZ^{+}23, ALAN25, AA22a, AASWC25, AAAAS20, AMN22, ACKS22, ADAT25, AC25b, AK24b, CLX21, CWW^{+}23, CCH^{+}23, CCC^{+}24, CPS25, CWW^{+}25, CSKD22, CRP25, CKV22, DYW24, DC23, FYL^{+}24, FCA25, GYZ^{+}25a, GDG21, GL21, HYW^{+}21, HLQD23, HWW^{+}25, HYW^{+}20, HWFW25, HXZ^{+}23, HLZ^{+}21, JK23, KM24a, KS20, KLP^{+}24, KAO^{+}23, LCZW20, LWX^{+}22, LCZ^{+}22, LQH^{+}22a, LHZZ24, LGJW21, LFHX23, LWH^{+}24, MBGF20, NKA^{+}25, NK21, NHT^{+}25, OX22, PRTV22, PTD20, RVS24, SA20, STB25, SRY^{+}25, STB^{+}20, SK24a, Sin25, SS23b, SSK23b, SCTK23, TYAH^{+}24, TSW^{+}23, TTA23, TLS23, TD24, WLQ20, WJT^{+}22, WTL^{+}23, WLY^{+}24, WXJ^{+}25, XLWC24, XKJG25, ZZHY25, OX22]. feature-based [CCC^{+}24, XKJG25]. feature-deviated [LWX^{+}22]. Feature-driven [TRM22, CWW^{+}25]. feature-level [SCTK23]. feature-momentum [HLQD23]. featured [DH25, GMBN21]. Features [KC22, NLA^{+}23, XLY^{+}20, ABAM22, AGI^{+}25, BKA20, DMM^{+}23, GW24, JFQ^{+}22, LLL^{+}22b, LLL^{+}22c, LYSB23, LTZ^{+}24, LSAH21, LLPL25, MBPCC21, NJS^{+}24, PGRS24, RT20, SASW23, TSKG23, WYP^{+}24, WWL^{+}25a, WWS^{+}25, YXL^{+}22, ZLF^{+}20, ZY24, ZSLY25, ZZL^{+}22, ZJJS^{+}22, ZCZW24]. February [KLSD24a, Ano20o, Ano21p, Ano22p, Ano23p, Ano24p, Ano25p]. FECC [LWZC23]. Federated [ACGP25, BAB^{+}25, DBB^{+}24, DCS24, GYG^{+}23, HMT^{+}25, LLL^{+}23a, LBA25, MG24a, NHC25, WR24, XRW^{+}25, YLDW23, dSVST23, AB23, AKT21, ADL25, dCBJSdS23, CZ24, CLCZ23, CWZL23, CQC^{+}24, CDLC24, DCSW20, FGWJ20, FGH^{+}21, FFB^{+}23, GTL^{+}22, GB23, GXZ23, HXZ^{+}24, HYW^{+}20, HXX^{+}24, HYZ^{+}24, ISM22, KEB23, LKA^{+}21, LYSB23, LZRR23, LLL^{+}25, LZC24, LLLC22, LLL^{+}23c, MBU24, NB25b, QJQ22, RZY^{+}24, dCFA^{+}23, SZL^{+}24, SM24, SZWH25, TCYL21, TSW^{+}21, UHK^{+}21, WZX23, WMGH^{+}23, YHW^{+}23, YJK^{+}24, YZL^{+}24, ZTR^{+}23, ZHG^{+}24, ZZS^{+}23, uS25, LWYG24, MLZ^{+}23]. federations [MBFIdD21]. FedGNN [ZHG^{+}24]. FedIMP [LWYG24]. FedMSE [NB25b]. FedRight [CLCZ23]. FedSIGN [GXZ23]. feedback [FWMJ^{+}23, HWML25, SZW^{+}25, WLQ20]. FELACS [ACGP25]. FePN [CWW^{+}23]. FeSA [FK22]. FeSAD [FK24]. Few [KT23, LMD^{+}22, PFG^{+}24, ZLX^{+}24, ADDH25, BFR23, CKV22, LXW^{+}24, WTW21, YLS^{+}22, YKSK25, ZWX^{+}24, ZJJS^{+}22, ZLX^{+}24]. Few-shot [KT23, LMD^{+}22, PFG^{+}24, ZLX^{+}24, ADDH25, BFR23, CKV22, LXW^{+}24, WTW21, YLS^{+}22, ZWX^{+}24, ZJJS^{+}22]. Few-VulD [ZLX^{+}24]. fewer [ZCW^{+}24]. FGACFS [LCSD20]. FGDA [KXZ^{+}23]. FGDA-GS [KXZ^{+}23]. FGMC [HMK^{+}20]. FGMC-HADS [HMK^{+}20]. FH [FDK^{+}22]. FH-CFI [FDK^{+}22]. Ficks [MV25]. Field [Kha21, MAR^{+}23, SPV20, WLLC24]. Field-Sensitive [Kha21]. fields [YGW^{+}25]. fifteen [Fur22]. file [AGAS25, CZZ^{+}21, DMB21, LCSD20, OPK20, WWL25b]. fileless [LPZF24, ST25a]. files [AlD25, Dub24, KKB^{+}25, NRKK23]. filesystem [LsJC^{+}22]. Fill [PSP^{+}22]. filling [zLXZrxZ25, OS22]. Filter [PC22, MCM^{+}24, SA20, SLL23]. filtered [WTC^{+}22]. filtering [CPS25, GPD^{+}23, KBMM23, LHX23, RGP23, SU25, XYZ^{+}25]. filters [DHW24, QLZ^{+}25]. Financial [Coj25, AS24, GPC^{+}22a, LdGW^{+}22, VBF21, vdKWH20]. find [MW20, WRG^{+}21]. Finder [LTCR25]. Fine [FDK^{+}22, RGP23, TW20, WHW^{+}25, AK20, AK24b, BFR23, CZL^{+}25, FXL^{+}20, JPL20, LCZ^{+}25, LCSD20, SDCL25, SLCS24, XLH^{+}25, XLG^{+}23, YSM^{+}21, YZC25]. Fine-grained [FDK^{+}22, RGP23, TW20, WHW^{+}25, AK20, JPL20, LCZ^{+}25, LCSD20, SDCL25, SLCS24, XLH^{+}25, XLG^{+}23, YSM^{+}21, YZC25]. fine-tuned [AK24b, CZL^{+}25]. fine-tuning [BFR23]. FineFool [CZX^{+}21]. FineGCP [WHW^{+}25]. finger [ZSL24]. fingerprint [DSS^{+}25, HABW24, RDPR25, SK20, SLL23, TTP20, WML^{+}23, XHWL20, YWK^{+}22]. Fingerprinting [LZ25, CSvdA^{+}23, GZS^{+}22, HW25, HHL^{+}22, KDDM20a, KDDM20b, LXW^{+}24, RDQL25, SYZ^{+}23, VAAG23, WLY^{+}24, XZY^{+}24, YZC25, Zha24]. fingerprints [LCZ^{+}25]. Finnish [KIAV22]. FinTem [HABW24]. FiPiBox [DUK25]. Fire [RVS24]. firefighter [ACC^{+}20]. firefly [SK24a]. firefly-based [SK24a]. Firefox [FFPC22]. firewall [ANG^{+}25, CDG22, CDG^{+}23, DUK25]. firewalls [ANG^{+}25]. firms [AM25]. Firmware [CYL^{+}23, ASF^{+}23, BWB^{+}21, LXC^{+}20, QPL^{+}23, YGK^{+}25, YFdCdC^{+}24, YLZ^{+}25, ZZJC20]. First [GHOS22, HZZ^{+}24, SAAW21, TK20, VMS20]. first- [VMS20]. fish [XYH^{+}20]. fish-swarm [XYH^{+}20]. fit [KDE20, MGGR22, YHZ23, LXC^{+}20]. Fix [CZE^{+}22]. fixing [CDF^{+}20]. FKR [ZM20]. FL [BAB^{+}25]. FLAD [DCS24]. FLADEN [HMT^{+}25]. flag [SCVB21]. flame [SDA^{+}24a]. Flash [MDB^{+}20, ZXB^{+}24]. flatness [TLPY21]. Flexible [VCD22, CRdMM^{+}25, Rep25, TXH^{+}23]. FlipIt [TZZ^{+}21]. flipping [LBA25]. flood [dNPM^{+}21]. Flooding [BGM^{+}23, PJL23, JGS^{+}24, KT25, MAR^{+}23, NH21, WFY^{+}20]. Flow [LSXJ22, MZW^{+}25, QYJ22, SPG^{+}25, AHuHS24, AA22b, ATGK22, CMCVGH^{+}23, DLC24a, FLLY24, FDK^{+}22, GLX^{+}23, JPL20, JXYF25, KMOC20, NCW^{+}23, QGW^{+}25, RKSV20, RFdS23, SSRM25, SLCS24, SYC^{+}24, WCH^{+}25, YGC25, Zha24]. flow-based [AA22b]. flow-centric [Zha24]. Flower [MS23]. flows [LL23, QZS^{+}25, TTRY20, YXL^{+}22]. FlowTracker [GLX^{+}23]. fluctuation [NH21, XGX^{+}24]. flux [ADJS21]. fly [GWY^{+}25]. Focus [IE22, JJW^{+}23, MLS^{+}23]. focused [HKM22, OJH25, vdSFF20b]. Focusing [GBA22, yKK22]. Foe [Gal20b].

Fog [AA24b, DKSS20, WQL25, FMU22, GG21, KKRP21, QSA24, SRCP22, SBV24, YHY23a, ZJY^{+}23, AMNR20, JGS^{+}24]. fog-assisted [SBV24]. fog-based [FMU22, YHY23a]. Fog-Driven [AA24b]. fog-to-cloud [ZJY^{+}23]. Foggier [SBV24]. FogHA [GG21]. follow [SSL20]. following [SS23a]. FoMO [vdSF23]. FoNAC [AMNR20]. fool [CDLC24, KFP^{+}24, WZZ^{+}22, ZM22]. Footsteps [FMU22]. forecast [KEB23]. Forecasting [ACC^{+}20, GSG23b, YPWS20, CLL24]. Forensic [ACG20, APPN23, hPRPJ23, ALZ^{+}20, AA21c, AOAA20, DNB^{+}20, FFPC22, FFPC23, HSL20, KMP^{+}20, SZC^{+}21, TK20, ZLCA21]. Forensics [ACG20, CMFUA^{+}20, SGC23, AN21, AJHA20, BD23, BTAK21, CWZ^{+}24b, FXL^{+}20, KKHK24, NDPC21, PCR20, WS21, WTC^{+}22, XWW^{+}20]. Forest [WFK25, ZD21, ÇÜD21, GSS^{+}20b, LCZW20]. forgery [NMLN23, WWLH23]. Formal [AIJ23, AJHA20, HRU24, MAK21, MSMH21, TO22, CCL^{+}24, WHE^{+}23, ZCAP23]. Formalization [SA25, SPZ^{+}20]. Formally [CDGS23]. fortify [Sin25]. Fortifying [WZG24]. Forums [CHP21]. forward [ESA^{+}23, MAF^{+}23, SLCS24]. forwarding [CZSW24, LLL23d]. foster [CMMST22]. fostered [SL20]. Fostering [DAA24, KMAHU24]. four [GLX24, KKS^{+}22]. four-factor [GLX24]. Fourier [LHS21]. fractal [ZUKH23]. fractional [ZUKH23]. Framework [JKSS22, KJI25, SHM22a, AJPK24, AAT23, AHZL25, ADDH25, AA25, AHFL25, AA22a, AIA23, AB23, AdSJN25, BEFI25, BCGC25a, BCGC25b, BK22, BD23, BH25b, BBJ20, BSR^{+}25, CRdMM^{+}25, CJD25, CGBK24, CHZ^{+}23, CHK24, CPS25, CXL^{+}25, CRP25, CAT20a, CAT20b, CAT24, CCBGC25, Coj25, DHW24, DG25, DS22, ESTA^{+}25, EA20, FK24, FSN22, FWR21, GMP20, GZH^{+}24, GYZ^{+}23, GXZ23, GLH^{+}23, GWY^{+}25, GS24, HN25, HHL24, HuRMMW22, HMMW22, HD25, HCCS22, HW25, HGSB24, HZ20a, HLX^{+}24, IE22, JKR^{+}21, JMU^{+}25, KKR^{+}24, Kil25, KN20, KLP^{+}24, KTGDE20, KS23b, KEK^{+}22, LKW^{+}20, LLYL21, LQH^{+}22a, LLL^{+}22c, LZY^{+}25b, LX25, LDS24, LHG^{+}24, LXH^{+}25, LGJW21, LJL^{+}25, LLH^{+}25, MYW24, MJZ^{+}25, MYS^{+}25, MAR22, MLB^{+}25, MZA^{+}20, MSG22, MMC24, MBK25, MKH24, MSL^{+}25, MCMMMA23, MEH^{+}24, NR24, NCN24, NPL^{+}23, Nie24, NM21, OS22, OFH^{+}25, OANA25]. framework [PMA^{+}23, hPRPJ23, PYT20, PGZB23, PC24, PD25b, QMW20b, QJDD24, RKES23, RCA23, Rep25, RZT^{+}23, RGPAF20, RGACPA23, SS24a, SCE21, SLC^{+}21, Sin25, SJH^{+}21, SBV24, SST21, SG21, SKS21, SWL^{+}25, ST25b, TLKW20, TGOF22, TFP21, VF22, WZB^{+}23, WLL24, WYJ^{+}24, WXD23, WBN^{+}20, WTC^{+}22, WHW^{+}24, XZZ^{+}24, XLL23, XWW^{+}20, XSC^{+}23, YLS^{+}22, YLSJ23, YGB^{+}25, YHH^{+}24, ZKS25, ZLCA21, ZYH^{+}22, ZMLP23, ZHG^{+}24, ZLX^{+}24, ZWX^{+}24, ZXW^{+}25, uS25, vGCHZ24]. frameworks [Alm25, CKG22, ESBJ20, MSL^{+}24, Olu22, TDM^{+}21, VRPTB^{+}23]. Framing [QXS23]. Fraud [Onw20, AIJ23, BBZS24, CYL^{+}21, CQLW25, SST21, VK22, WSW^{+}22]. fraud-resilience [AIJ23]. Frauds [ZTJ^{+}21]. Fréchet [GT24a]. Free [CWHH23, LZHL20, MQH^{+}23, RDQL25, XLG^{+}23, YS22]. free-text [LZHL20]. frequencies [HSQH23]. Frequency [LCH^{+}24a, HCL^{+}23, LCH^{+}24b, LFL^{+}25, LCL24, LLX^{+}23, TLY^{+}24, VAAG23, ZSL24, ZZZ^{+}23]. frequency-injection [ZSL24]. frequent [RLW^{+}20, WTL^{+}21]. fresh [DSR23]. friendly [AC25a, ZWX^{+}20]. friends [NPCM24]. fronting [XGX^{+}24]. Frontline [AASC25]. FS [YLS^{+}22]. FS-IDS [YLS^{+}22]. FsQCA [OXNA25, CVZHN25]. FTE [OYZ^{+}20]. FTE-based [OYZ^{+}20]. FTOA [WCH^{+}25]. FTOA-RP [WCH^{+}25]. Full [ZKL^{+}23, LLZJ22]. full-model [LLZJ22]. Full-stack [ZKL^{+}23]. fully [ÁFM25, Vu22]. Function [GAJ25, GYG^{+}23, PC22, CKL^{+}25, DS21, DTK^{+}21, LsJC^{+}22, TP25, WBC^{+}23, XYZ^{+}25, KEM^{+}25a]. function-call [DS21, WBC^{+}23]. function-preserving [DTK^{+}21]. functional [LHGB21]. Functionality [RSEK20, YLH^{+}24]. functionality-API [YLH^{+}24]. Functionality-based [RSEK20]. functions [BSLK^{+}25, BPCZ23, ISD24, LCC^{+}21, MT21a, Rep25, YKG20, YFdCdC^{+}24, YLZ^{+}25]. further [LCL23]. fusing [TLS23, ZGZL23, ZHC^{+}25b]. fusion [ASA23, AZES21, CLL24, GLH^{+}23, HLL^{+}25, HXZ^{+}23, HQL^{+}22, JWG^{+}24, JJW^{+}23, KM24a, KLP^{+}24, LQH22b, LGG25, MJZ^{+}25, PRTV22, PLW^{+}21, SCW^{+}22, SCTK23, TSKG23, TSW^{+}23, TTA23, WGRZ23, WLY^{+}24, WGS^{+}25b, XLWC24, XLS24, ZXH^{+}25, ZGS24]. fusion-based [SCW^{+}22, ZXH^{+}25]. Future [BRV^{+}22, GM24, LYLT24, MAKH21, AKT21, ATA^{+}25, BBA^{+}21, BRMH22, BvtH^{+}23, BCKN22, DKV^{+}24, GCS22, HSB^{+}22, KKAS21, LGXZ22, MAL22, MLC^{+}23, PNF23, PDWF25, SWW^{+}25, UNBF21, VRPTB^{+}23]. fuzz [LLP25]. FuzzAGG [LJL^{+}25]. fuzzer [KL24, SZW^{+}25, ZYJ^{+}23, WCZ^{+}24]. fuzzers [ZWH^{+}25]. Fuzzing [BRMH22, BS25, DCH^{+}22, LXW^{+}25a, MMSJ25, QY22, XWJ^{+}24, BCD21, CLHT24, EFI23, GSYM22, GLC^{+}25, JWY^{+}25, KL24, LZW^{+}24, hLHhLfW21, LWZ^{+}24, LXJ^{+}25, LJL^{+}25, PLW^{+}23, PCR20, QPL^{+}23, WCZW23, WCZ^{+}24, WPZW25, WLLC24, WWL25b, ZMZ^{+}25, ZFW^{+}25, ZG24, ZGG^{+}25]. Fuzzing-based [QY22]. fuzzing-driven [LJL^{+}25]. FUZZOLIC [BCD21]. Fuzzy [DKSS20, HMK^{+}20, KK25a, MSH^{+}25, ÖEÖ22, TK25, VS25, AFLR25, BC23a, EÖ25, OJL24, RBM^{+}25, RMS^{+}22, SDA^{+}24a, SYD^{+}25, SYG^{+}21, AZO^{+}23]. Fuzzy-AHP [AZO^{+}23]. Fuzzy-rule [TK25].

G [BBBB^{+}20]. GA [SSK23b]. GACN [LWL21]. gadgets [LMW^{+}25]. GADM [RDS20]. GAIL [CHZ^{+}23]. GAIL-PT [CHZ^{+}23]. gait [ACLA22, ACD^{+}24, TYAH^{+}24]. gait-based [TYAH^{+}24]. Galaxy [LFCD21]. GAMBD [LGZD23]. Game [GSS20a, HMPS20, HHS22, LYD20, QXS23, RKW^{+}23, FXZ22, JK21b, LL22b, LL25, LDW25, LZDZ21, LZZ^{+}21, MTD^{+}24, QYYH24, RZ20b, SYL^{+}20, Sun21, TZZ^{+}21, TSW^{+}24, WP21, ZMF^{+}23, ZCJ^{+}20, ZHC25a]. game-based [ZHC25a]. Game-theoretic [RKW^{+}23, MTD^{+}24]. games [CCL^{+}24, CHJ22, HZ20b, NB25a, WRG^{+}21, YKN21]. gamified [PSCT21]. Gamifying [LET^{+}25]. Gaming [HF25]. GAN [LYLT24, AMD^{+}24, DMM^{+}23, DTK^{+}21, LQL25b, RLL^{+}23, TZZ22, YCCZ21, YKK24]. GAN-based [YCCZ21, YKK24, RLL^{+}23, TZZ22]. GAN-like [AMD^{+}24]. GANs [HX21, SZCZ24]. gap [ADO23, GB21, SWK20, YAC25]. GAPBAS [LSTL24]. gapped [GSE20]. gaps [BOS^{+}21, RSLM25, WIOO24]. gas [MAR^{+}23]. Gaslight [PCR20]. gate [CBPW25, PLW25, XLD25]. gate-level [CBPW25, PLW25]. gated [GTAK25, RKRS24, ZGS24]. Gateway [AGVA21, KK25c]. gateways [FKL23]. gathering [MCJPS22]. Gaussian [HMK^{+}20, ZCLW23]. gaze [CKK^{+}23]. GC [FNZ^{+}23]. GC-NLDP [FNZ^{+}23]. GCN [CXC^{+}24, PHT^{+}25, ZLDX25]. GCN-MHSA [CXC^{+}24]. GConv [AWN^{+}23]. GDPR [Bel25, GRdAS23, MLV^{+}21, MR23, TSW^{+}21]. GDPR-compliant [MLV^{+}21]. GDroid [GCZ21]. Gender [GAC20, FKP^{+}24, GAW^{+}23, vdSFF20b]. gender-focused [vdSFF20b]. GenDroid [XSC^{+}23]. general [HLX^{+}24]. generalisation [SHK^{+}21]. Generalizable [LHGB21]. generalization [GPAF25, NZT^{+}25]. generalized [HKA24]. Generalizing [dCBJSdS23, WWLH23]. Generate [LLX25, KHG24, PLW^{+}21, QPL^{+}23, TLPY21, WP21]. generated [AKCS20, PMNL21, YXL^{+}22]. Generating [LGZD23, MWR20, MMdR24, SLX^{+}23, SLDL23, HLYZ21, RLL^{+}23, SLR25, TO22]. Generation [PLW^{+}23, PBA24, SHM22a, CJL^{+}25, DSS^{+}25, DHL^{+}23, GWL^{+}23, HD25, ILVR^{+}24, KLSD24a, KLSD24b, KS23a, LL21a, LWL21, LJL^{+}25, LQL25b, MLB^{+}25, MLS^{+}23, PA20, QZT^{+}22, RKKK25, RCA20, SSK23b, XYH^{+}20, YZP^{+}25, ZTYA25, ZHZ^{+}22]. Generation-based [PLW^{+}23]. Generative [KC22, ZZZ^{+}23, AESF23, CHZ^{+}23, GSSC24, KS23a, LYLT24, QMZ^{+}23, QSA24, RKA25, SHH^{+}24, WMK23, WS21, ZFT24]. generator [CLMU25, DMM^{+}23, PB25, PLW^{+}23]. Generic [RCD23, DSN^{+}22, ZCS^{+}23]. genes [HQL^{+}22]. Genetic [AASWC25, LSTL24, AA22a, AEA24, DMM^{+}23, FYL^{+}24, HYW^{+}21, JAAS^{+}23, SB23]. genetic-based [AEA24]. GeniGraph [AEA24]. genuine [QLL^{+}25]. geolocation [MW20]. geospatial [AHL^{+}21]. German [CSM^{+}24, HP23]. gesture [CHH25]. gestures [SCTK23, WHPL21]. get [MGGR22, ZZ25, RCD21]. GGT [CWX^{+}24]. GHGDroid [SFX24]. GhostEncoder [WYF^{+}24]. GhostNet [AC25b]. gifts [TDM^{+}21]. GJALLARHORN [BEFI25]. Gleaned [RD23]. Global [KC22, SFX24, ABA23, ARS^{+}25, CWZL23, LHAE22, LJJ22, RYS25, ZY24]. global-context [RYS25]. Global/Local [KC22]. GNN [GZH^{+}24]. GNN-based [GZH^{+}24]. GNSS [WYP^{+}24, WBZ^{+}25]. goals [GZG21]. Golden [CWHH23]. GONE [ZCS^{+}23]. Good [SGC23, TRH21, VS20]. Google [PBdL25]. Gordon [Ske24, Ske22]. governance [AWVG20, Coj25, SSK22, HR24]. Governing [GBS22, SSK22]. government [NDDH^{+}21]. GP [LL22a, SSK23b]. GPMT [SLX^{+}23]. GPS [KUA^{+}25, NLA^{+}23]. GPT [MLS^{+}23]. GPT-4 [MLS^{+}23]. GPU [MMIS22, PKK^{+}21]. GPU-based [MMIS22]. GQM [PFR20]. GQM-based [PFR20]. Gradient [CZ24, CJC^{+}24, HDZ^{+}22, KXZ^{+}23, LQHZ25, SPN^{+}25, VT25, ZTR^{+}23]. Gradient-based [CZ24]. GRAIN [XCY^{+}25]. grained [AK20, FDK^{+}22, JPL20, LCZ^{+}25, LCSD20, RGP23, SDCL25, SLCS24, TW20, WHW^{+}25, XLH^{+}25, XLG^{+}23, YSM^{+}21, YZC25]. gram [BDA23, HCL^{+}23, SQL^{+}24b, TQJ^{+}23]. GramBeddings [BDA23]. granted [OMO20a, WGS^{+}25a]. granular [MJZ^{+}25]. granularity [LLD^{+}25]. Graph [CWX^{+}24, TLPY21, XCY^{+}25, ZJC^{+}25, SHM22a, AAAAS20, ASAA21, AWN^{+}23, CLN^{+}25, CXC^{+}24, DWX^{+}23, DH25, DS21, DMZ24, FHZ^{+}22, FGY^{+}24, FCSP21, FNZ^{+}23, GCZ21, GZ24, GHMG24, GFH^{+}22, GHZ^{+}24, HFLZ23, HZH^{+}24, JQYL22, JYWS25, JZY^{+}25, KK25b, KEK^{+}22, LCZ^{+}22, LXZ^{+}23, LPL25, LFHH22, LJL^{+}25, LGG25, MWM^{+}25, MV25, OFH^{+}25, PB25, PFG^{+}24, PYT20, PJ23, QYW^{+}25, RKRS24, RG25, SDCL25, SFX24, SQL^{+}24a, SD25, SWYL20, SYG^{+}21, SCL^{+}24, TCY^{+}20, WML^{+}23, WHW^{+}25, WW25b, WZY^{+}23, WWZ24, XLZ24, Zen23, ZNF21, ZHG^{+}24, ZDM^{+}25, ZWB^{+}25, ZJH^{+}24, ZSWT24, ZLZX24, ZSGB^{+}22, dPR23]. graph-based [ASAA21, JZY^{+}25, KEK^{+}22, SDCL25, SFX24, SQL^{+}24a, ZSGB^{+}22]. Graph-guided [CWX^{+}24]. graph-theoretic [TCY^{+}20]. GraphCodeBERT [WXT^{+}24]. GraphFVD [SDCL25]. Graphical [NGC^{+}25, TLPY21, XLL^{+}25]. graphs [AEA24, BHO23, CHA^{+}24b, FWF^{+}23, HZGM25, HNZ^{+}23, JWY^{+}25, KLSD24a, KLSD24b, LCY^{+}21, MS25, PB25, PL23, SLZP24, SGY23, TCM^{+}25, WWS^{+}25, WBC^{+}23, XLH^{+}25, ZX23a, ZLDX25]. GraphSAGE [XZZ^{+}25]. GraphXSS [LFHH22]. GRC [MLS^{+}23]. greedy [GSS^{+}20b]. grey [SZW^{+}25]. grey-box [SZW^{+}25]. greybox [JWY^{+}25, WPZW25]. Grid [BP24, ARR^{+}22, ADC^{+}20, KMAHU24, KTGDE20, LYW^{+}21, LS21, SMM22, YZL^{+}20]. grids [HFLZ23]. Ground [JANR23]. Group [GC22, CYL^{+}21, GMBV21, HHP^{+}24, WCH^{+}25, YM22, ZLDX25]. group’-based [WCH^{+}25, CYL^{+}21, HHP^{+}24]. groups [JB22, LK20, RKKG25]. GRU [LBD^{+}23, TXY^{+}23, WQL25]. GRU-Autoencoder [WQL25]. GRU-Based [TXY^{+}23]. GS [KXZ^{+}23]. GSEDroid [GZH^{+}24]. GSGLD [YJK^{+}24]. GSPN [LHL^{+}24b]. GTM [GSS20a]. GTM-CSec [GSS20a]. guaranteeing [XHTZ20]. guarantees [GXZ23]. guard [MAP23, CPV23, XLG^{+}23]. guards [TA25]. Guessing [WZX^{+}22, YHZ^{+}22]. guidance [LWH^{+}24, WLDW24]. guide [Zen23, YKK24]. Guided [WCP^{+}23, CWX^{+}24, DZZ^{+}21, KXZ^{+}23, KL24, LFHY23, ZDH^{+}23, ZZHY25, ZSJ^{+}25, ZGG^{+}25]. Guiding [CSB^{+}25]. guilt [NPCM24]. guilt-by-association [NPCM24].

Haar [ZNF^{+}23]. habit [WJC23]. Habits [GBA22]. Hacker [RCA24, HHP^{+}24]. HADS [HMK^{+}20]. HAE [XKY25]. HAE-HRL [XKY25]. HAformer [JWG^{+}24]. HAGDetector [LCW^{+}22]. HAN [YLPZ21]. HAN-BSVD [YLPZ21]. handle [GJB22]. handler [LZW^{+}24]. handler-aware [LZW^{+}24]. Handling [HY21, Het21, HSŠ^{+}22, LWP^{+}25, MISO^{+}23]. HANDOM [WDJ^{+}23]. handover [GG21, NR22, YM22, ZMQ21]. hands [BDC^{+}22, CKK^{+}23]. hands-on [BDC^{+}22, CKK^{+}23]. handwritten [PA20]. Happn [KMP^{+}20]. Happning [KMP^{+}20]. hard [HLC^{+}24, MTN^{+}23, QY22]. hard-label [QY22]. hard-to-attack [HLC^{+}24]. Hardening [AWJB21, BRM21, HZX^{+}21]. harder [WS25b]. Hardware [CWHH23, DL21, HHS22, SYC^{+}24, CBPW25, CJS^{+}21, FDK^{+}22, GLC^{+}25, HLX^{+}24, PLW25, PLE^{+}23, SCBP24, SMC^{+}21, dNPM^{+}21]. hardware-assisted [FDK^{+}22]. Hardware-based [DL21]. harm [SCSO20]. harmful [HMB^{+}21]. harms [NTH25]. Harnessing [MLS^{+}23, GLZ^{+}22, GCN^{+}24]. hash [OJL24]. hashed [MT21c]. hashing [KLZ24, SK20, YWWH22]. Hawk [RVS24]. haystack [WRG^{+}21]. hazard [LHL^{+}24b]. HC [CZH^{+}25]. HC-NIDS [CZH^{+}25]. HClave [ZCJ^{+}24]. HDFEF [LQH^{+}22a]. HDL [AD21]. HDP [ZYL^{+}22]. HDP-CNN [ZYL^{+}22]. head [CXL^{+}24, CRS^{+}22, CXC^{+}24, RKRS24, XXZ^{+}21]. header [MMdSdS23]. health [KAM^{+}21, LM23, OANA25, SHL^{+}20, ZCL^{+}25, Esm20]. health-data [LM23]. healthcare [BSO25, CSZQ23, CFYBF23, HRU24, Lee23, MVSSZ24, NKA^{+}25, SC21, SHM22b, WZG24, ZS24]. heap [YLZ^{+}21]. HEART [KUA^{+}25]. heavy [RSLM25]. heavy-duty [RSLM25]. Hello [BMAP24]. HEOD [NCM24]. HER [LZY^{+}25b]. HER-PT [LZY^{+}25b]. herd [XW20]. Hessian [ZQCX24]. Heterogeneous [LCW^{+}22, WDJ^{+}23, ZJC^{+}25, AJPK24, dCBJSdS23, BBD23, CGS22, LZCS22, LZC24, QYW^{+}25, RZ20b, dCFA^{+}23, SFX24, SWYL20, ZGZL23, ZLY^{+}21]. HetIoT [MPM23]. Heuristic [dSFG20, GYZ^{+}25b, vSRW^{+}20]. Heuristic-based [dSFG20]. heuristics [TG25]. heuristics-based [TG25]. hex [JWG^{+}24, TQJ^{+}23]. HGAN4VD [ZJC^{+}25]. hidden [AABE20, CKT^{+}25, JHD^{+}20, LLL^{+}20b, SZY^{+}22, YHC25, ZMF^{+}25]. HIDIM [ZXW^{+}25]. Hiding [vDK22, ARAG23]. HIDS [KS23b]. HIE [Esm20]. Hierarchical [LBW^{+}25, AAAK23, BLC^{+}24, CYC^{+}24, KLP^{+}24, LQH^{+}22a, LZS^{+}23, MMdR24, SWL24, TSW^{+}24, YLPZ21, YXL^{+}22, ZXW^{+}25, ZYPT22, ZCW^{+}24, XLY^{+}20]. High [LCH^{+}24a, LLP25, ARAG23, CWZ^{+}24a, CWW^{+}25, GSS^{+}20b, HHE22, LYHC24, LCH^{+}24b, LXW^{+}25a, LaTL^{+}25, MZWM23, WZCP20, XLD25, ZYPT22]. high-dimensional [WZCP20]. High-Frequency [LCH^{+}24a, LCH^{+}24b]. high-level [LaTL^{+}25]. high-order [LYHC24]. high-quality [LXW^{+}25a]. high-speed [CWW^{+}25]. high-stability [XLD25]. High-trigger [LLP25]. higher [GZG21]. Highly [FGWJ20, ADJS21, Yan20, YLZ^{+}25]. highly-random [Yan20]. Highway [ZYL^{+}22]. Hijacking [WFKH25, LLG21]. Hilbert [WYP^{+}24]. Hindsight [LZY^{+}25b]. Historical [CZH^{+}25]. history [GLC^{+}25]. history-based [GLC^{+}25]. HMM [LWS^{+}20, STM^{+}20]. Hoc [GRJV24, KJR^{+}20, MBP20, RKA25]. HOG [BA20]. Hole [Kac25, DP25]. HoleMal [CXL^{+}25]. Holistic [NCW^{+}20, AKA^{+}23]. home [AWJB21, BEB25, LLC^{+}23, MSP^{+}20, VBS^{+}25]. homeland [NTH25]. homes [AMVG24, HZKN22, Hod21, ZMLP23]. homogeneous [BBD23]. homogenized [PS21]. homomorphic [AHL^{+}21, ÁFM25, MMJG24]. Honest [YCMM20]. honey [MGGR22, XRW^{+}25]. honeynet [RZ20b, SYF^{+}21, ZZ20]. honeypot [GSS20a, JJT^{+}24, SB25]. honeypots [MGGR22]. honeytokens [PBA24]. honeyword [GZG21]. honeywords [TLPY21]. Hooktracer [CMFUA^{+}20]. hop [DH25]. hope [NLW24]. hope-extended [NLW24]. hopping [Yan20]. hospitality [XWY21]. host [BL25, CXL^{+}25, JN23, MWR20, SG21]. host-based [BL25, MWR20]. host-level [CXL^{+}25]. hot [RCD21]. hotspot [GGZ25]. household [XSL^{+}24]. households [HSOL23]. HP [CWZ^{+}24a]. HP-MIA [CWZ^{+}24a]. HRL [XKY25]. HScheduler [GLC^{+}25]. HSMs [GMS20]. HSTF [XLY^{+}20]. HSTF-Model [XLY^{+}20]. HSTW [FLLY24]. HT2ML [WZB^{+}23]. HTLC [MT21c]. HTTP [CWW^{+}25, CKK^{+}23, DVAAM23, LZ25, WAPA25, XLY^{+}20, YXL^{+}22]. HTTP-based [DVAAM23, XLY^{+}20, YXL^{+}22]. HTTP/2 [CKK^{+}23, WAPA25]. HTTP/3 [CKK^{+}23]. HTTPS [AIBKJ24, ASM20, Sch24]. hull [CXLV20]. human [ATJ^{+}22, BQA^{+}23, CAT20a, CAT20b, CAT24, CPP24, GGR^{+}24, GAS^{+}23, HG20, KUA^{+}25, NCW^{+}20, VS22, VBS^{+}25, WLKHD25, NCM24]. Human-assisted [NCM24]. human-centred [VBS^{+}25]. human-centric [BQA^{+}23]. humans [VB20]. hunter [MSH^{+}25, PGRS24, PdARD24, TG25]. Hunting [KJI25, ABAJ25a, ABAJ25b, JHD^{+}20]. Hybrid [GMBN21, JMU^{+}25, KM24b, NHT^{+}25, PGRS24, QMC^{+}22, SWL^{+}25, AD21, ATA^{+}25, BKSA25, BTS23b, BBJ20, CFZL22, CQC25, CCL^{+}21, EÖ25, FLLY24, GYZ^{+}25b, GFH^{+}22, HLZ^{+}25, HZ20a, HYR^{+}22, KM24a, LLL^{+}22b, LGL21, hLHhLfW21, LWZ^{+}24, LJ22, MRSP24, MKM^{+}25, MGL^{+}20, MDKR25, NPL^{+}23, NMLN23, OS22, PRTV22, PM25, PCG23, QJDD24, RKES23, RLW^{+}20, SDA^{+}24a, SU21, SSN23, SS25, SGA^{+}23, SPSK24, TK25, VT25, WZB^{+}23, XZWD25, XKY25, YGC^{+}23, ZFWL22, ZFG^{+}24, ZZ20, vGCHZ24]. Hybrid-featured [GMBN21]. hybrid-NIDS [SS25]. HYDRA [GMP20]. hygiene [NCW^{+}20, STSV25]. Hyper [JYWS25]. HyperGo [LWZ^{+}24]. hypergraph [JYWS25]. Hyperledger [JKSS22, LZS^{+}20]. hyperparameter [MRG21]. Hypervisor [ATMN20, CDN21, MYW24, ZCJ^{+}24]. Hypervisor-based [ATMN20]. HyVulDect [GFH^{+}22].

I-MAD [LFCD21]. Ibn [RNA22]. IBV [JPL20]. IBV-CFI [JPL20]. icon [WSW^{+}22]. ICS [AMC25, MGGR22, RMMP23, WLLC24]. ICS-LTU2022 [AMC25]. ICT [AAJP25, KBE^{+}24, MBKD21, RVJ25]. ID [DG25, LRL^{+}22, RTBK21, SST21, ZWB^{+}25]. ID-based [DG25, LRL^{+}22]. identifiable [WPS20]. Identification [BD23, BDA23, CHP21, AGW25, ALZ^{+}20, BMH20, CZM^{+}25, GGC24, GWL^{+}23, GTAK25, IMMS21, KZYZ22, LLL^{+}22a, LWW^{+}21, MHE24, MHC^{+}25, RSSD23, RGPAF20, STB^{+}20, SKS21, TYAH^{+}24, WHPL21, WWH^{+}22, ZCAP23, ZSZ^{+}25, ZWW^{+}21, ZXL^{+}24, ZL24, ZLL24]. identifier [ADÖU^{+}20, LCZ^{+}25]. identifier-based [ADÖU^{+}20]. identify [LX25]. Identifying [DPKHP22, LTL^{+}22, MR23, NMT25, STS25, TH24, AAAAS20]. identities [SJX^{+}20]. Identity [MAKH21, Ogb21, AOM24, ADO23, BC23b, BHWE22, CWYW25, He25, JN23, KA23, MBU24, MBFIdD21, NGJ^{+}22, OG23, OJH25, Sar21, YL20, ZXH^{+}25, ZZW^{+}25]. identity-authentication [He25]. IDN [WZC^{+}24]. IDN-based [WZC^{+}24]. IDoS [HZ22]. IDPFilter [LB25]. IDS [YKK24, AAAT22, AA22a, DBG24, DTK^{+}21, FFB^{+}23, GYZ^{+}25a, GSS20a, GJB22, HM23, JGS^{+}24, JSK^{+}23, LCC^{+}22, LCL^{+}25, SGJ^{+}25, SSK23b, SPSK24, YLS^{+}22, ZWF^{+}24, uS25]. IEA [CWW^{+}25]. IEA-DMS [CWW^{+}25]. IEC [DMPT25]. IEEE [ZM20]. if [ST24]. IFAttn [JFQ^{+}22]. IFIP [Pit25]. IFTTT [AA25]. IID [LYSB23, uS25]. IIOT [DZ22, AA23, CWC^{+}25, DZSCR24, HWD^{+}25]. ILL [HM23]. ILL-IDS [HM23]. Image [ETAB22, OS22, VAW^{+}20, XGS^{+}21, AA20, AGI^{+}25, CZX^{+}21, DZZ^{+}21, DYW24, GJCJ20, JKR^{+}21, JJW^{+}23, NMLN23, NDPC21, OSPP22, WCW^{+}23, WLLL24a, WS21, WTC^{+}22, ZZCD20, ZCW^{+}24]. Image-Based [VAW^{+}20, OS22, XGS^{+}21, AA20, DYW24, JKR^{+}21]. image-behavior-based [OSPP22]. images [AÇ23, GJCJ20, MWL23, TZZ22, YWL^{+}20, ZWW^{+}21, ZX23a]. imbalance [AJPK24, GJB22, ZXW^{+}25]. imbalanced [AD21, ÇÜD21, DÇÜD22, FBR25, LHG^{+}24, MC22, XXZ^{+}21]. IMCEC [VAW^{+}20]. IMD [LSG^{+}24]. imitation [CHZ^{+}23]. immature [TZZY25]. Immersive [VBCM23]. immune [HLL^{+}21, JIG22]. immune-based [HLL^{+}21]. Impact [DJG^{+}25, JYF^{+}25, LBA25, TKS21, VLV25, ARR^{+}22, ALAN25, ABAJ25a, ABAJ25b, BG25, DJG^{+}23, GTP^{+}25, GAC20, HKM^{+}23, HHZ21, JZ23, KLC23, MVVL^{+}24, MGTK22, Mit20, MBH20, MR23, RA25, SRN24, TLS24b, TJD23, vdWLM24, SSL20]. impacted [LBL^{+}20]. impacting [CBB25, MAAG22]. impeding [PKPC24]. Imperceptible [DV24, QWW^{+}22, XMK21]. imperfect [YCMM20]. impersonation [JGS^{+}24]. impersonation-aware [JGS^{+}24]. implement [CZZ^{+}21]. Implementation [OBK^{+}23, RMS^{+}22, BDC^{+}23, HWML25, IO24, JYZ^{+}23, MT21a, SSK22, TFMP25, WBNT20, YCWW23, ZLY23]. Implementations [BGM^{+}23, CCVMZ24, LZ25, ZGG^{+}25]. Implementing [SSG25b, WNF20, WTHL25]. Implications [GMB22, CZK24, GMLB22, GMLB23, MHE24, MAP23, TRH21]. Implicit [KK25b, CHK24, WJT^{+}22]. I}mportance [LWYG24]. I}mportance-based [LWYG24]. improper [ZWL^{+}25]. Improve [LZRR23, JJT^{+}24, MPRB21, MKL21, WLLC24, XQS25, YQL24]. Improved [GLX^{+}23, MMJG24, WZW24a, BH25a, HLYZ21, JHD^{+}20, KLW^{+}23, MRG21, MAAA20, SCE21, Sin25, SQL^{+}24b, YWZW24, ZCLW23, ZCZZ24]. improvement [ASA^{+}22, ZJY^{+}23]. Improvements [SLDL23]. Improving [BKA20, BS20, CHA^{+}24b, DTT^{+}22, EFI23, FBL20, HLQD23, HKA24, HWD^{+}25, HYR^{+}22, JK21b, LL22a, LFHY23, LCL24, MMdSdS23, NG23, ZFG^{+}24, BMM22, CYL^{+}24, GSSC24, hLHhLfW21, MBR23]. impulsive [QYYH24]. imputation [MPL^{+}25]. In-band [PHT^{+}25]. In-depth [NTBH^{+}22, SGT24, CBFH20, MHLZ22, PB25, YP21]. in-network [KKY^{+}23, LXH^{+}25]. in-process [JYCL25]. in-role [YHZ23]. in-task [TN21]. In-vehicle [CZM^{+}25, AA22a, KJI^{+}24, LWLL25, PJL23, ZLC^{+}23a, ZL24]. incentives [BRWF23]. incentivized [DG20]. incidence [WX23]. Incident [BHL^{+}23, Het21, SJH^{+}21, AMD^{+}21, AWB24, HSŠ^{+}22, KLC23, MISO^{+}23, NNA^{+}23, NSLM25, RVJ25, RRB25, SVP21, SCBM23, TKS21, XWW^{+}20, vdKSCY22, PAM^{+}24]. incidents [ESA^{+}23, GMRB23, KKGK25, KN20, NTH25, PNF23, PNF24]. Including [dPR23, FBL20]. incomplete [BGL^{+}20, LZZ^{+}21]. inconsistencies [ASM20, GGK25, LTCR25]. Inconsistent [SOZ^{+}23b, TLS24a]. incorporating [HWFW25]. Incorrect [Nie24, SOZ^{+}23b]. increased [MBGF20]. increasing [MYH25]. Incremental [LXS^{+}24, LBW^{+}25, PCG23, HM23, LLLZ21, PC24, TT20]. independent [ACD25, AGW25, JCJ^{+}21]. Index [ÖEÖ22, FYDH25, PC24]. Indicators [AA21c, AASC25, BT24, CHA^{+}24b]. indirect [ADM23]. individual [BJZ^{+}23, CFRB25, KSMW23, LZRR23, LZH^{+}24, MT21b, SCBP24, SPV20, TYAH^{+}24, WMC20]. individuals [UÇ22]. indomitable [CCL^{+}21]. indoor [AK20, FMA22, SC23]. induced [LZD^{+}20]. inducement [PGP^{+}24]. induction [MAAA20, VK22]. Industrial [CS23, DAZ20, HW20, KWF25, KDBS22, LQHZ25, SJH^{+}21, TXY^{+}23, AA21c, AASC25, APPN23, BZE^{+}20, FYL^{+}24, FFB^{+}23, GDK^{+}21, HLL24, HZG^{+}23, ISR^{+}23, JSK^{+}23, KBK24, KMH^{+}22, KF25, LZZ^{+}21, LXY^{+}24, LJL^{+}25, MT24, MGGR22, MSZ20, NJ22, QJDD24, SUS25, SZL^{+}24, SYF^{+}21, SGJ^{+}25, TLY^{+}24, TH24, YHY23a, YZLS25, YGW^{+}25, ZLDX25, ZWB^{+}25, AHZL25, BBR20, LWS^{+}25, Sin25, ZKS25]. industries [TJD23]. Industry [JTF^{+}24, CPS25, FKvdA^{+}24, MV25, XWY21, dVABH20, CBB25, GMB^{+}25, MNMS24]. industry-specific [CPS25]. inertial [GAZ25]. Inevitable [BHL^{+}23]. infected [YP21]. Inference [GYYK22, CWZ^{+}24a, HXZ^{+}24, LKA^{+}21, LKL23, LLL^{+}20b, LYHC24, LXH^{+}25, LQH22b, LXZ^{+}22, PBG22, YGW^{+}25]. inferring [PMF^{+}20]. influence [Als20, CMA23, HR24, KSMW23, TJD23, WS24]. Influences [MCvO24]. Influencing [BTS23a, TCJS22, NM24]. Informal [HSW25]. Information [AWVG20, ABA23, Esm20, HDS21, JRP24, KKP22, LSXJ22, MD22, MBKD21, Ogb21, PLB^{+}22, SV25, SS23a, SSRK20, WMC20, ADM23, AD23, AS24, AGM20, ALZ^{+}20, AZO^{+}23, ADÖU^{+}20, ALD^{+}21, AKH20, BMM22, BT21, BBZS24, BSFB20, BGL^{+}20, CCE23, CXC22, CZH^{+}25, COH21, CFT22, CFYBF23, DPKHP22, DPTAM25, DLC24a, DANS20, DPK20, DKV^{+}24, FVF21, FBL20, FJR23, FF20, FVV23, GLTH21, GW24, HSW25, HKM^{+}23, HQL^{+}22, HS25b, ILB20, JZ23, JYCL25, KSS20, KGIS21, KAFDW22, yKK22, KSC23, KJJ^{+}24, KRR23, KIAV22, LTCR25, LZZ^{+}21, LL23, LQL^{+}25a, LHYW22, LET^{+}25, MAR22, MCJPS22, MGG^{+}25, MVSSZ24, MCvO24, MGG^{+}20, NM24, Nie24, OCB23, OXNA25, PLS20, Ram25, RKG20, RKKG25, SP20, SSHP21, SSL20, SL20, Spr23, SYC^{+}24, SCC^{+}25, TLS24b, TKS21, TMN24, TFP21, TN21, VBM23, Wes20, WPS20, XGS^{+}21]. information [XW20, XWY21, YHZ23, YSZ24, ZWDE24, ZJL^{+}20, ZZW^{+}25, ZZZ^{+}23, ZCW^{+}24, dVABH20, vdSF21]. information-gathering [MCJPS22]. information-theoretic [MAR22]. Informational [HZ22]. informative [AM25, MWM^{+}25]. informed [SCBM23]. infrastructure [BEAE25, HLFR20, KHG24, MAR^{+}23, NTH25, NMT25, RTBK21, TDM^{+}21, TSCM21, WLKHD25]. Infrastructures [SSN22, KJHL22, LET^{+}25, MISO^{+}23, RA25]. inhibiting [ABAJ25a, ABAJ25b]. initialization [PTD20]. initiatives [SEGD^{+}22]. injected [HZX^{+}21]. Injection [BP24, SKL23, ADDH25, AGAS25, ADC^{+}20, CMCVGH^{+}23, FBR25, GSY^{+}20, HFLZ23, HXZ^{+}21, LX21, MT24, PT25, SMvH^{+}21, SGVA21, TWLK22, VWJ24, YZL^{+}20, YFdCdC^{+}24, ZSL24, ZDZD24]. inline [BPCZ23]. Inner [SCSO20]. Innovation [MSSO^{+}25, MBKD21, QLZ^{+}25]. Innovative [LKS^{+}23, CDG^{+}23, EÖ25, JYL25]. input [OMO20b]. inputs [AKA^{+}23, ZG24]. insecure [LZX^{+}23]. Insecurity [SKL23]. insertion [GFM^{+}22]. Insider [EYYZ20, RSSD23, ASSW23, DYW24, FZZ^{+}25, GAC20, GSY^{+}20, HSW25, LLH^{+}25, PKPC24, PF20, QYW^{+}25, SH25, VB20, WRG^{+}21, XZZ^{+}24, YW21, dPR23]. insiders [RW24, VM21]. insightful [TSW^{+}21]. Insights [ASF^{+}23, DPKHP22, HG20, NMT25, RD23, SRCP22, TLS24b, VBM23, OAA^{+}22, RSLM25, WJC23]. Inspect [LXC^{+}20]. Inspecting [YCWW23]. inspector [ZL24]. inspired [AAAK23, AC25b, JRT^{+}25, KBAG23, YJK^{+}24]. instance [WLL^{+}23, ZHL^{+}22]. instantiation [LWC^{+}25]. institutions [LdGW^{+}22]. instruction [HLP22]. instructions [GFM^{+}22]. instrumental [CXC22]. instrumentation [DCH^{+}22, GSYM22]. insufficient [LWL21]. insurability [CSM^{+}24]. insurance [AWB24, CSM^{+}24, HSW25, HNH^{+}22, JSYK25, LM23, LL25, MTN^{+}23, Ske22, Ske24, UYMM21]. insure [LL25]. insurer [BRWF23]. integral [VF22]. integral-comprehensive [VF22]. Integrated [CFYBF23, MSSO^{+}25, RMS^{+}21, CXC22, NHT^{+}25, SS24b, SGJ^{+}25, WLHC21, ZLCA21]. Integrating [AdSJN25, LHL^{+}24b, SSN22, SAB^{+}25, WJC23, XW20, MBU24, SSG25b, ZXH^{+}25, ZH24]. Integration [AKT21, UHK^{+}21, CBK22, FK23, GS24, JYL25, KMOC20, LHW^{+}20, RBM^{+}25, SYH^{+}22, ZLF^{+}20]. integrative [BHW24]. Integrity [LWL^{+}25, hYAS^{+}24, AAS24, DS21, FDK^{+}22, GG22, GPK21, JPL20, KM24a, KBC21, LT25, LRL^{+}22, LPX20, LZS^{+}20, MYW24, MT21a, NGM^{+}24, NCW^{+}23, SLCS24, TJ20, TWW21, TZZ^{+}23, XHB^{+}21]. Intel [KBC21, RRM20]. intellectual [AAS24]. Intelligence [APA^{+}24, SL20, WLC^{+}20, ATMA23, BFR23, BAM^{+}24, CTV21, CRC^{+}23, CZLZ24, CZL^{+}25, CPS25, GLH^{+}23, Het21, HZH^{+}24, ILVR^{+}24, JYWS25, JLS22, KMAHU24, KRR23, LHC24, MJZ^{+}25, MS25, PJ23, RBM^{+}25, SVP21, VRPTB^{+}23, WLH^{+}24, YWL25, ZGZL23, ZYL^{+}20, ZHC25a, ZSS21]. Intelligence-driven [WLC^{+}20, SVP21]. Intelligent [ALZ^{+}20, BFW21, BTS23b, ATMN23, BKSA25, CGS22, CHZ^{+}23, DG25, FWKB24, GS24, HFK^{+}24, HLL24, IFHP22, LYW^{+}21, LZS^{+}23, LZY^{+}25b, MLQ^{+}23, MF25, SYF^{+}21, VPM23, YLL23, ZHL25]. intensified [PRPS25]. intensity [RZY^{+}24]. Intent [OBK^{+}23, LZY^{+}25a, YLW^{+}24b]. Intent-Driven [OBK^{+}23, YLW^{+}24b]. intention [ADM23, Esm20, JZ23, LZL^{+}22, SWK20]. intention-behavior [SWK20]. intention-concealed [LZL^{+}22]. intentional [JZ23]. Intentions [HHZ21, BTS23a, VBM23]. inter [ARR^{+}25, RMCT23, SU25, MF25]. inter-app [RMCT23]. inter-arrival [SU25]. inter-UAVs [MF25]. inter-vehicular [ARR^{+}25]. interacting [BDLP24]. interaction [QMZ^{+}25]. Interactions [QXS23, KSMW23, VB20, WJT^{+}22]. Interactive [RXFZ21, KJPP23, LSMK25]. intercultural [IFHP22]. interdependent [AEA24, LB25]. interdisciplinary [GMP24]. interest [GGZ25, KT25, MLQ^{+}23, WFY^{+}20]. interference [Wan21]. intermittently [AFAS25]. international [ZWDE24]. Internet [AHZL25, BBR20, RNA22, Sin25, DC23, GLX24, STZ^{+}22, WLKT23, AKT21, ASG^{+}25, AZDF20, BTS23a, CRP22, CGBK24, CZH^{+}25, CGCY21, FGH^{+}21, GPRGMF25, GXL25, HLFR20, HXX^{+}24, HYR^{+}22, JSK^{+}23, JCJ^{+}21, KK25b, KEM^{+}25a, KFS^{+}22, KJI25, LXX^{+}25, LL21b, LPX20, MKM^{+}25, MM25, NAK23, NHC25, OANA25, OAA^{+}22, PMF^{+}20, PNFBH23, SASW23, SM23, SAB^{+}25, SvSZ21, SWK20, SWL^{+}25, SCC^{+}25, TFJ23, UHK^{+}21, YHY23a]. Internet-of-Things [JSK^{+}23, LPX20, PMF^{+}20]. interoperation [YLW24a]. Interpretability [DV24, KLW^{+}23]. Interpretable [HWFW25, LFCD21, LaTL^{+}25, TXY^{+}23, BLC^{+}24, IMMS21, JFQ^{+}22, CWW^{+}25]. interrogating [Ade21]. intersection [JZL^{+}24]. intersections [PJ23]. interval [FSN22, HHSL20, WWW^{+}22]. interventions [ACC^{+}20, MBF25]. Intra [AGAS25, MF25, ARR^{+}25, LWL21, NG23]. intra- [ARR^{+}25]. intra-kernel [NG23]. Intra-section [AGAS25]. intrinsic [LLL^{+}22c, OG23]. introspection [DSR23, TML^{+}20]. introspection-based [TML^{+}20]. intruders [SLR25]. Intrusion [AYL25, BTE23, CPV23, LLL^{+}23a, LBA25, LCC^{+}22, LCZW20, LQL^{+}25a, PRTV22, SCOE23, SVPM21, SPG^{+}25, SRT25, WLHC21, YWZW24, ANAAK25, ALAN25, AHC22, AA22a, ABAM22, AD21, ATMN20, AWN^{+}23, AAAK23, AC25b, AV25, AK24b, BKSA25, BFC^{+}25, BL25, dCBJSdS23, CPRV21, CKS^{+}24, CQC^{+}24, CZH^{+}25, DCK^{+}23, DH25, DBB^{+}24, FYL^{+}24, FWKB24, FFB^{+}23, GYZ^{+}25a, GC25, GYZ^{+}25b, GL21, GJB22, HYW^{+}21, HRFS24, HFK^{+}24, HHE22, HCL^{+}23, HKA24, HGSB24, HZG^{+}23, HXX^{+}24, JN23, JXYF25, JLQ^{+}20, KK25b, KS20, KDE20, KLSW24, KJI^{+}24, KBAG23, KS23a, LM24, LLL^{+}22b, LFC^{+}25, LLYL21, LWX^{+}22, LQH^{+}22a, LHD^{+}23, LQHZ25, LLLB22, LGH21, LFW^{+}25, LZA^{+}23, MAR22, MFA20, MR25, MV25, MSZ20, MC22, MF25, MDB24, NK21, NB25b, NPL^{+}23, OCMF23, PMA^{+}23, PGRS24, PCG23, PD25a, PTD20, PRPS25, QMC^{+}22, QSA24, RBM^{+}25, RKA25, RVS24, RKRS24, STB25, SUS25, SKD23]. intrusion [SY24, SLR25, SZL^{+}24, SMM22, Sin25, SSG21, SQL^{+}24a, SQL^{+}24b, SG21, SZWH25, TK25, TP25, TAF24, TC23, TD24, VPM23, VDN24, WFK25, WJT^{+}22, WJT^{+}23, WXJ^{+}25, XLL23, XZZ^{+}25, XKY25, YLS^{+}22, YLL^{+}22, YHW^{+}23, YZLS25, YHH^{+}24, ZLF^{+}20, ZJW^{+}22, ZCLW23, ZWB^{+}25, ZZHY25, ZFT24, ZLH23, ZLZX24, ZXW^{+}25, ZCD^{+}22, ZWZX25, ZDM^{+}24, ZCPB23, dSVST23, uS25, DBG24, PAGDG24]. intrusions [BAD^{+}24, GVEV22, LDS24, VT25]. intrusive [dNPM^{+}21]. invariant [AJPK24, WWLH23, ZC24]. invariants [SUS25]. invasive [BCP22]. Inventory [BRV^{+}22, GMSA24]. inverse [GBG20]. Inversion [JCHS24]. Inverted [GYG^{+}23]. Invertible [TTP20, HABW24]. Investigating [ALAN25, ASG^{+}25, AABLS24, CDF^{+}20, GPC^{+}22a, HSSPK21, KFB^{+}25, LBA25, LdGW^{+}22, OUR^{+}23, PTD^{+}25, OPK20]. Investigation [MF25, AA21c, AOAA20, BJZ^{+}23, BD23, CSM^{+}24, LCY^{+}21, LBL^{+}20, LWL^{+}23, hPRPJ23, RBHW25, SWS^{+}25, SZC^{+}21, WHW^{+}25, WLQ^{+}25]. investigative [AHS25]. investment [FAAF23, SSL20]. invisible [MZWM23]. Invoice [KLA^{+}22]. involved [SS23b]. IoHT [MAKH21]. IoMT [KK25a, ZZP^{+}25]. IoRT [NHC25]. iOS [KS25a, KMP^{+}20]. IoT [CGCY21, MM25, PMF^{+}20, SWK20, MB20b, AJPK24, AAAT22, AA25, AL23, ASF^{+}23, ARB20, ACA^{+}23, AB23, AA24b, ACGP25, AWJB21, AK24b, BWZ^{+}25, BS24b, BRV^{+}22, BAB^{+}25, BTS23b, BAC^{+}23, BLR23, CSZQ23, CCVMZ24, CPV23, CGS22, CGBK25, CWL^{+}22, CXL^{+}25, CYL^{+}23, CRP25, CSRA22, DKSS20, DUK25, DWX^{+}23, DS22, DCH^{+}22, ESTA^{+}25, FFB^{+}23, FDK^{+}22, GC22, HN25, HMT^{+}25, HSSPK21, HW25, ISD24, JGS^{+}24, JKS^{+}25, KKR^{+}24, KJCL20, KLC25, KAU^{+}23, KR23, KFZ^{+}20, KSSL22, KK25c, KS25b, KJ23, LMD^{+}22, LXLL25, LDS24, LLLB22, LLG20, MRSP24, MLM20, MNMS24, MAF^{+}23, MK24, MS23, MV25, MSP^{+}20, MBB^{+}23, MGL^{+}20, MGZ^{+}25, NB25b, PM25, PBG22, PC22, Rs25, RKH^{+}22, RGG^{+}24, RZT^{+}23, RK24, SDA^{+}24a, SS22, dCFA^{+}23, SYZ^{+}23, STB^{+}20, SBV24, SS25, SKG21, SGA^{+}23, SZWH25, SWW^{+}25, TK25, VAAG23, WSC^{+}20, WLM25, WBNT20, WXD23]. IoT [WBC^{+}23, XLZ24, YHW^{+}23, YWZ20, YLZ^{+}25, ZKAK23, ZTYA25, ZDM^{+}24]. IoT-based [RKH^{+}22, SWW^{+}25]. IoT-BOTNET [BTS23b]. IoT-DDoS [SBV24]. IoT-Fog [JGS^{+}24]. IoT-PRIDS [ZDM^{+}24]. IoT-related [MBB^{+}23]. IoT/Non [MK24]. IoT/Non-IoT [MK24]. IoTPredictor [KKR^{+}24]. IoTSecSim [CGBK24]. IP [GSSP23]. IPA [LL22a]. IPv6 [XZG^{+}25]. iQUIC [DG25]. IR [GYZ^{+}25a]. IR-IDS [GYZ^{+}25a]. iris [MDKR25, ZXS^{+}25, CG24, SA20]. IRL [LL21a]. IRL-based [LL21a]. IRP [DLQ^{+}21]. irreversible [GBG20]. ISA [SLCS24]. ISC [MVSSZ24]. ISO [CRdMM^{+}25, MSL^{+}24]. isolated [LCTZ25, ZCJ^{+}24]. isolating [JYS21]. isolation [CZG^{+}22, JYCL25]. ISP [GGK25]. issue [CS24, RZ20a, bOJW20]. Issues [CZZ^{+}23, MDB^{+}20, SSR^{+}24, WS22, AMC23, CZE^{+}22, CBFH20, LB25, SCTK23, SWW^{+}25, TMJ^{+}23, ZCL^{+}25]. IT-OT [ARR^{+}22]. IT/OT [LET^{+}25]. Italian [Sch24]. ITC [SCOE23]. items [GGK25]. itemset [RLW^{+}20]. itemsets [ARAG23].

jacking [NTBH^{+}22]. Jadeite [OSPP22]. Jamming [BCP22, CDGS23, LLJL23, Yan20]. jamming-resilient [Yan20]. January [Ano20p, Ano21q, Ano22q, Ano23q, Ano24q, Ano25q]. Java [OSPP22].

JavaScript [AR21, CWQZ25, FHSQ20, FHZ^{+}22, HLZ^{+}21, LXW^{+}25a, LFHX23, XWJ^{+}24]. JIT [LXW^{+}25a]. Job [OJH25, BKV^{+}23, RPR^{+}21]. Joint [AKH24, BT21, WLL24, WWW^{+}22, CLX21]. JointGAN [PSP^{+}22]. JOWMDroid [CLX21]. JSContana [HLZ^{+}21]. JStrong [FHZ^{+}22]. judiciary [AdSJN25]. Juliet [BSG23]. July [ABAJ25a, Ano20q, Ano21r, Ano22r, Ano23r, Ano24r, Ano25r]. Jump [XZZ^{+}25]. Jump-Knowledge [XZZ^{+}25]. June [LCH^{+}24a, Ano20r, Ano21s, Ano22s, Ano23s, Ano24s, Ano25s]. jure [AKH20]. justice [ADM23, AABLS24]. justification [JZ23, KJR^{+}20]. justify [MGTK22].

K-SAD [KE25]. Kaminsky [ZYH^{+}21]. Kaminsky-style [ZYH^{+}21]. Keep [LCB^{+}22, BTH22, LXS^{+}24]. Keeping [GMY22]. Kellect [CSZ^{+}25]. Kernel [CSZ^{+}25, YHC25, ZWF^{+}24, CZG^{+}22, DXZZ25, kJLcL24, LWSW24, LWSW25, LMW^{+}25, LWS^{+}25, MSCJ21, NG23, ZQMC20, ZWL^{+}25, CRC^{+}23]. Kernel-based [CSZ^{+}25]. Kernel-level [YHC25, DXZZ25]. Key [AA24b, BCKN22, HWW^{+}25, HLFR20, HLZ^{+}21, KKB^{+}25, LWO^{+}25, NCW^{+}23, NMT25, SCH^{+}20, SYZ^{+}23, TFP21]. key-blocks [SYZ^{+}23]. Keypad [LQH22b]. keys [GMY22]. Keystroke [ASA23, CMFUA^{+}20, BF20, BHWE22, LKL23, LZHL20, MBK25, SCTK23, YMZ^{+}23]. keyword [SYG^{+}21, XGS25, YLZ^{+}21]. kill [BS20]. kill-chain [BS20]. Kirkpatrick [KIMJ23]. Knapsack [ASAA21]. KNN [AGI^{+}25]. KNN-based [AGI^{+}25]. Know [GSG23a, LLC^{+}23, SV25, SJ21]. KnowCTI [WLH^{+}24]. Knowledge [KEK^{+}22, WLH^{+}24, ANAAK25, AJHA20, BWZ^{+}25, BTA24, BMM22, BOS^{+}21, CYL^{+}24, CCC^{+}24, DDCK22, GLZ^{+}22, HZH^{+}24, KLC25, LZS^{+}23, LPL25, LGG25, MS25, NR24, DAAM24, SLZP24, SVWJ25, SAM25, SLX^{+}23, SRM22, SCVB21, TFMP25, WWWS24, WLDW24, YL20, ZLDX25, ZJH^{+}24, XZZ^{+}25]. Knowledge-based [WLH^{+}24]. knowledge-driven [KLC25]. knows [LCL23]. Korea [PKPC24]. KPDFI [NCW^{+}23]. KPointer [LCB^{+}22]. KronoDroid [GMBN21]. KRTunnel [WSQ^{+}22]. KRYSTAL [KEK^{+}22]. Kubernetes [ZCAP23].

LaAeb [FZZ^{+}25]. LAAKA [AA24b]. Label [LBA25, BG25, HRFS24, QY22, RZW^{+}20, RHH^{+}22, YGC^{+}23, ZXL^{+}25]. Label-flipping [LBA25]. label-only [RZW^{+}20]. labeled [BLDB24, XLH^{+}25]. labeling [GTGHAG22, GCV22, OCMF23, Özt24, PSY^{+}22, TLPY21]. labelled [CEO^{+}25]. Labels [JANR23, YLY^{+}23]. lack [BBD^{+}24, BLDB24]. ladder [MSD^{+}23]. landscape [GSG23b]. Language [DGH21, ZDM^{+}25, AA21a, AKAFQ22, AKA^{+}23, BDM^{+}24, BSG23, CCW^{+}24, CZL^{+}25, DDW^{+}23, GAJ25, HZH^{+}24, KJHL22, KBE^{+}24, KLJE24, MB20a, MSL^{+}24, MISO^{+}23, NCN24, SB25, WHE^{+}23, WPAG24]. Large [CHP21, CYL^{+}23, CHH25, CCW^{+}24, CZL^{+}25, CBPW25, CPP24, DDW^{+}23, GMBV21, GJCJ20, HZH^{+}24, LWSW24, LWSW25, MSL^{+}24, NCN24, PLW25, dCFA^{+}23, SB25, SJ21, SCC^{+}25, WCL25, XGL20, YWW24, ZTR^{+}23, ZDM^{+}25]. Large-Scale [CHP21, CYL^{+}23, CPP24, GJCJ20, PLW25, SJ21, SCC^{+}25, WCL25, XGL20, YWW24]. LARM [ACCS25]. last [GVEV22]. late [ST24]. Latent [LK20, AFAkJS25, DC23, GB23, ZYH^{+}21, ZLH22]. Lateral [DYL^{+}23, SKKG25]. Latest [HSB^{+}22]. Lattice [SU21]. laundering [AKWR21, YLL23]. law [FCA25, MV25]. laws [Olu22]. Layer [KFHM24, BAB^{+}25, BBR20, DPSS23, DZ22, GS24, He25, HWW^{+}25, JHD^{+}20, LLL^{+}22d, RSÁGVV23, SB23, ZHL25, ZCS^{+}23]. layered [RMVG25]. layers [SLR25]. LBS [HLL^{+}25]. LCDA [SSS^{+}21]. LDCDroid [LWP^{+}25]. LDP [SYH^{+}22]. LDPSR [zLXZrxZ25]. leaders [DPKHP22]. leadership [Ram25]. leads [BPL^{+}20]. leak [XZLG25]. leakage [CZ24, HMB^{+}21, HZX^{+}25, LZH^{+}24, XT24, YJK^{+}24, vdKWH20]. LeakFocus [XZLG25]. Leaking [KSC23, CZE^{+}22]. leaks [CKT^{+}25, LLLZ21]. learn [AGI^{+}25, Gur25]. Learning [ACA^{+}23, AXAL24, BAB^{+}25, BPW^{+}20, DCS24, ESA^{+}23, GM24, GYG^{+}23, HMT^{+}25, JSK^{+}23, JTF^{+}24, KMG21, KJI25, LLL^{+}23a, LBA25, LBW^{+}25, LWP^{+}25, NLA^{+}23, NHC25, PNF23, RKW^{+}22, SUC23, TYAH^{+}24, WR24, WLLL24b, WWLH23, YLDW23, AOM24, AAAT22, AKH24, AHC22, AN23, AÇ23, AA22b, ATMN23, AB23, AA21a, AKT21, AKCS20, AA20, APK24, AAAK23, AYS20, ADAT25, AES22, AZDF20, AGAP24, ACGP25, AWJB21, AQAK21, AV25, ADC^{+}20, BKA20, BCO^{+}22, BFR23, BFC^{+}25, BAD^{+}25, dCBJSdS23, BBD23, BM24, BAC^{+}23, BTAK21, BGA24, BQA^{+}23, CRS^{+}22, CPR21, CDJD24, CLHT24, CRP22, CZ24, CKS^{+}24, CZS^{+}21, CLS22, CTH^{+}22, CLCZ23, CHZ^{+}23, CWZL23, CQC^{+}24, CDLC24, CQC25, CJA25, CNTBG21, DGS^{+}23, DA20, DBB^{+}24, DCSW20, DDW^{+}23, DHL^{+}23, DC23, ELD24, FGWJ20, FGH^{+}21, FK24, FD23, FSN22, FO25, FJR23, FFB^{+}23, GIO24, GPC^{+}22a]. learning [GLZ^{+}22, GMP20, GMPMS21, GFM^{+}22, GTP^{+}25, GTL^{+}22, GB23, GLX^{+}23, GKS24, GXZ23, GJB22, HHL24, HLP22, HRU24, HDS21, HXZ^{+}24, HYW^{+}20, HK25, HGSB24, HM23, HXX^{+}24, HYZ^{+}24, IMMS21, ISM22, ISD24, ISR^{+}23, JHD^{+}20, JK21a, JE21, KLC25, KMOC20, KS20, KEB23, KAU^{+}23, KJI^{+}24, KEM^{+}25a, KAO^{+}23, KSSL22, KS25b, KR21, LLL^{+}22a, LLL^{+}22b, LKA^{+}21, LLJL23, LCZ^{+}22, LLF^{+}23, LYSB23, LZS^{+}23, LZRR23, LHD^{+}23, LLJ^{+}23, LZNZ24, LWYG24, LXS^{+}24, LTZ^{+}24, LLL^{+}25, LXC^{+}20, LHZZ24, LWD^{+}25, LBD^{+}23, LDS24, LBL^{+}20, LXH^{+}23, LYS^{+}24, LZC24, LZG^{+}22, LLH^{+}25, LLLC22, LLL^{+}23c, LWS^{+}25, MLZ^{+}23, MLXC24, MG24a, MM21, MPM23, MK24, MSG22, MLQ^{+}23, MH25, MCMMMA23, MDB24, MMIS22, MZLK25, MGG^{+}20, MHLZ22, NB25b, NHT^{+}25, NHFB25, NMLN23, NZZ^{+}22, NGX^{+}23, NDPC21, OSPP22, OCMF23, OA25, OXNA25, PRTV22, PGRS24, PKK^{+}21, PNF24, PYT20, PGC^{+}24, PMF^{+}20, PC24, PD25b, PAA^{+}24, QJQ22]. learning [QZT^{+}22, QY22, RKES23, RPR^{+}21, RKH^{+}22, RZW^{+}20, RLL^{+}23, RZY^{+}24, RG25, RHH^{+}22, RMCT23, RK24, RNSS23, dCFA^{+}23, SMCP21, SY24, SZL^{+}24, SB23, SM24, SGJ^{+}25, SMM22, SARG23, Sin25, SHK^{+}21, SKKG25, SSG21, SYL^{+}20, SZWH25, SRM22, TK25, TSAG24, TSW^{+}24, TMJ^{+}23, TCYL21, TSW^{+}21, TD24, UHK^{+}21, VDL24, VDN24, WLHC21, WFT22, WJT^{+}22, WT23, WZX23, WMGH^{+}23, WYL^{+}24, WYF^{+}24, WXT^{+}24, WXD23, WXZ^{+}25, WTC^{+}22, WWW^{+}22, WLZ^{+}23, WHW^{+}24, WWZ24, XCY^{+}25, XSK23, XQS25, XCWC25, YLS^{+}22, YL23, YHW^{+}23, YHS24, YJK^{+}24, YZL^{+}24, YWW24, YWL^{+}20, YW21, YHH^{+}24, YLW^{+}24b, YPT22, ZKAK23, ZWW^{+}21, ZJW^{+}22, ZTR^{+}23, ZHG^{+}24, ZLY^{+}21, ZLC^{+}23b, ZLX^{+}24, ZHZ^{+}22, ZFWL22, ZCD^{+}22, ZJJS^{+}22, ZZS^{+}23, ZWZX25, ZCPB23, dSVST23, PAGDG24, WZB^{+}23, YQL24]. Learning-based [LLL^{+}23a, AZDF20, CLHT24, CJA25, DA20, FFB^{+}23, GTP^{+}25, GKS24, HHL24, HRU24, ISR^{+}23, KSSL22, LLJL23, MLXC24, RKH^{+}22, SGJ^{+}25, TSAG24, WJT^{+}22, YWW24, YHH^{+}24]. Ledger [ZYH^{+}22]. LEGO [ZFWL22]. length [HHSL20, HMB^{+}21]. lens [BSG23, CSEJGUA25, CKK^{+}23]. Lerning [YWX^{+}22]. less [PSY^{+}22]. Lessons [CC20]. Let [LXS^{+}24]. level [AVR20, BFR23, CZZ^{+}23, CBPW25, CXL^{+}25, CJS^{+}21, DXZZ25, DCH^{+}22, FK23, GMH^{+}23, JDBB22, kJLcL24, KJB^{+}25, LK20, LCY^{+}21, LaTL^{+}25, MCJPS22, NAK23, NM21, PLW25, SSHP21, SEGD^{+}22, SYH^{+}22, SCTK23, YZL^{+}20, YHC25, YWL^{+}20, ZYL^{+}22, ZSJ^{+}25, ZSGB^{+}22, ZZ20]. levels [SSK23a]. Levenshtein [SQL^{+}24b]. Lever [DANS20]. Leveraging [GR23, LT25, MFD25, SU25, TWQ^{+}25, VCP25, ZJC^{+}25, ZDM^{+}24, BHWE22, CJD25, OA25, SRY^{+}25]. Leviathan [Da 22]. liabilities [HZG19, HZG20]. library [JYF^{+}25, LCC^{+}21, YFdCdC^{+}24]. license [QMW^{+}20a]. licensing [HSW25]. LiDAR [HKY^{+}25]. LiDL [KJCL20]. life [KKGK25]. lifecycle [CNK25, LYS^{+}24, MJ25]. lifecycle-based [CNK25]. lifetime [HM23]. LIGAA [ZZZ^{+}23]. Light [LYL^{+}23, ZKS25, WPS20, LYL^{+}23]. Light-weight [ZKS25, WPS20]. LightGBM [JLQ^{+}20, LGH21]. lightning [vDK22]. Lightweight [GVJ23, GZ18, GZ20, PC22, SSS^{+}21, TLKK20, ZHJ^{+}23, AÇ23, AMM23, AWN^{+}23, ASB^{+}21, ASH^{+}24, AK24b, CZSW24, CXL^{+}25, GDM^{+}24, GZH^{+}24, LYW^{+}21, LZZX21, LWLL25, SS25, SYC^{+}24, WWWS24, WWS^{+}25, WXD23, XYZ^{+}25, YHZ^{+}22, YGK^{+}25, ZHH^{+}25, ZHJ^{+}21, AA24b, SP20]. Like [CCC^{+}24, AMD^{+}24]. likelihood [KLR^{+}25, XW20]. LIME [MMdR24]. limitation [LXLL25]. Limitations [MDB^{+}20, DVAAM23, GC25, MAL22]. limited [AMD^{+}24, BMH20, ZQCX24]. limits [AFAS25]. LINDDUN [RGPAF20, SVWJ25]. LINDDUN-based [RGPAF20]. line [GVEV22]. linear [ABAM22, AGI^{+}25, AK24b, TLK^{+}24, XHWL20, YWK^{+}22, ZUKH23]. lines [VB24]. link [DMZ24]. linkage [MHE24, WLLS22]. links [GZ24]. LinTracer [ZHC^{+}25b]. Linux [ACCS25, DCH^{+}22, FFPC22, GSK23, HMK^{+}20, LMW^{+}25, YGK^{+}25, YFdCdC^{+}24, YLZ^{+}25, ZWL^{+}25]. Linux-based [DCH^{+}22, YGK^{+}25, YFdCdC^{+}24, YLZ^{+}25]. LiSRA [SP20]. listing [SWK20]. Literature [JTF^{+}24, MVA25, MARB^{+}23, AA21b, ALD^{+}21, BB20, BOS^{+}21, CMLM25, CTV21, EL22, GG22, GGK25, HW20, IO24, KGK23, KBK24, KGIS21, MJ25, MMSJ25, NM24, NPF^{+}23, OK23, PNK20, PVFM^{+}21, SV25, VB24, YLL^{+}22]. little [SLX^{+}23]. live [KKHK24]. liveness [KZFM24, MC20]. lizard [BC23a, SDA^{+}24a]. LKRepNet [LWSW25, LWSW24]. LLM [HZH^{+}24, WTHL25, YFdCdC^{+}24]. LLM-based [YFdCdC^{+}24]. LLM-supported [WTHL25]. LLM-TIKG [HZH^{+}24]. LLMs [MFD25]. load [KEB23]. Local [AMN22, CCS^{+}24, DMZ24, KC22, ACC^{+}20, CJC^{+}24, CYHW25, DLQ^{+}21, FNZ^{+}23, GB23, HNZ^{+}23, LFL^{+}25, zLXZrxZ25, LWO^{+}25, LWS^{+}25, SVA21, SSZ^{+}24, XHTZ20, ZY24]. local-differential-privacy-based [ACC^{+}20]. Localization [KJCL20, TYZ^{+}24, YHY23b]. Locally [ZNF^{+}23, TLK^{+}24]. Location [NLA^{+}23, QIN25, AK20, AHL^{+}21, BT21, CMG20, FMA22, HYR^{+}22, KEK^{+}21, LQY^{+}20, LWL^{+}23, QWLL23, ZJL^{+}20, ZTW25]. location-Based [KEK^{+}21, AK20, LQY^{+}20]. locator [WZY^{+}23]. LocAuth [AK20]. Lock [MT21c]. locking [GSSP23]. Loeb [Ske24, Ske22]. log [BDM^{+}24, CEO^{+}25, CSZ^{+}25, FZZ^{+}25, LSWR20, PEGA23, SWS^{+}25, WHLS24, YKY24]. log-text [FZZ^{+}25]. LogDoS [ADÖU^{+}20]. Loggers [CMFUA^{+}20]. logging [ADÖU^{+}20, KLSW24, SSGM21]. logging-based [ADÖU^{+}20]. Logic [DKSS20, GSSP23, BC23a, KK25a, LXX^{+}25, MSD^{+}23, SDA^{+}24a, XWW^{+}20, ZGG^{+}25]. Logical [DAZ20, TCM^{+}25]. Login [ABC22, KKHK24, MCvO24, VRS^{+}25]. logins [GZ24, TG25]. logistic [LSLL24]. logo [BA20]. LogoSENSE [BA20]. LogPrécis [BDM^{+}24]. logs [ADAT25, FO25, IM23, RSW^{+}21, XGL20]. long [AOA22, GMB22, QSA24]. long-term [GMB22]. longest [HZL^{+}23]. longevity [RTBK21]. longformer [BS24a]. Longitudinal [MWL23, BMM22, GPC^{+}22b, HP23, ZZW^{+}25]. look [SAAW21, TK20]. loop [HCCS22, HWML25, LZD^{+}20]. LOPA [XHWL20]. Loss [GYG^{+}23, HX21, QLZ^{+}25, WTC^{+}22]. lossless [CSZ^{+}25]. Low [LRH^{+}21, LLX^{+}23, MZWM23, RG23, WLL^{+}23, WCZ^{+}24, WW25b, ZHH^{+}25, ZZZ^{+}23]. low-frequency [ZZZ^{+}23]. low-proportion [ZHH^{+}25]. Low-rate [LRH^{+}21, WW25b]. low-repetition [WCZ^{+}24]. LowPTor [ZHH^{+}25]. LP [ZQCX24]. LP-BFGS [ZQCX24]. LPSE [GZ20, GZ18]. LR [GB23, WW25b]. LR-BA [GB23]. LR-STGCN [WW25b]. LSTM [AOA22, BKSA25, CLL^{+}21, CFZL22, FBR25, HWFW25, LXY^{+}24, LPW25, XKY25]. LSTM-CNN-based [XKY25]. LSTM-ED [LXY^{+}24]. LSTMs [RSW^{+}21]. LTL [AHFL25]. LTL-based [AHFL25]. LTU2022 [AMC25]. LUT [CJS^{+}21]. LUT-level [CJS^{+}21].

m [OO21, CSLK24, DS20, JGS^{+}24, TD24]. M-MultiSVM [TD24]. M-PIVAD [DS20]. M-RL [JGS^{+}24]. m-times [OO21]. M2FD [ADL25]. M3D [BAB^{+}25]. M3D-FL [BAB^{+}25]. MA [MNMS24]. Machine [AOM24, ACA^{+}23, BPW^{+}20, DA20, ETAB22, GM24, JTF^{+}24, KSSL22, LLJL23, NLA^{+}23, PAGDG24, PAA^{+}24, RPR^{+}21, RKW^{+}22, RRS21, SUC23, WFT22, AA22b, ATMN23, AA21a, AKCS20, AA20, APK24, AAAK23, AZDF20, AWJB21, AV25, ADC^{+}20, ASH^{+}24, BKA20, BFC^{+}25, BAD^{+}25, BBD23, BM24, BGA24, BQA^{+}23, CLHT24, DSR23, ELD24, FK24, FO25, FJR23, FSCM23, GPC^{+}22a, GLZ^{+}22, GMPMS21, GTP^{+}25, HRU24, HDS21, ISR^{+}23, JHD^{+}20, JWG^{+}24, JE21, KAO^{+}23, KR21, LBD^{+}23, LYS^{+}24, MSCJ21, MH25, MCMMMA23, MDB24, MMIS22, MGG^{+}20, MHLZ22, NDPC21, OA25, PMNL21, QY22, RZW^{+}20, RHH^{+}22, RNSS23, SMCP21, SB23, SMM22, SHK^{+}21, TD24, UHK^{+}21, VDL24, WT23, WXZ^{+}25, YSM^{+}21, YPT22, ZKAK23, ZJW^{+}22, ZFWL22, ZCD^{+}22, WZB^{+}23]. machine-to-machine [ASH^{+}24]. machines [KBC21, WSW24]. MAD [LFCD21]. MADONNA [SRY^{+}25]. MaDroid} [DLC^{+}24b]. MAG [ISD24]. MAG-PUFs [ISD24]. MageCart [REHS24]. MAGIC [SRN24]. Mahalanobis [LX21]. mail [BA20]. major [AAJP25]. makers [APK24, DPK20, PL23]. Making [QJQ22, ATMA23, FRF24, LZDZ21, MBK25, SWK20, TSW^{+}24, ZSJ^{+}25]. Mal [LTZ^{+}24]. Maladaptive [BMH21, NLW24]. Malaysian [HNH^{+}22, NMT25]. Malbert [XFY21]. MalConv [LGZD23]. MaldomDetector [AKCS20]. malice [JAAS^{+}23]. Malicious [AAB22, BAB^{+}25, FHZ^{+}22, GTAK25, HLZ^{+}21, RJ24, SWYL20, WDJ^{+}23, XQS25, YLY^{+}23, AR21, ANG^{+}25, BDM^{+}24, CXL^{+}24, CTC^{+}25, CMMST22, CXC^{+}24, CWQZ25, CDM^{+}20, DZSL25, EMS23, FHSQ20, FO25, HYW^{+}20, JDBB20, KKR^{+}24, KLA^{+}22, LTL^{+}22, LWSW24, LZY^{+}25a, LWSW25, LXW^{+}25b, MW20, MBPCC21, MMdR24, MZLK25, NRKK23, OA25, PSY^{+}22, RW24, SM21, SRY^{+}25, STB^{+}20, SK24a, SCW^{+}22, SKM21, SGA^{+}23, SLX^{+}23, SCC^{+}25, TG25, WFT22, WTL^{+}23, WT23, WZW24a, XCWY25, XLW24, XT24, YXL^{+}22, ZUKH23, ZY24, ZHG^{+}24, ZLLY25, ZHL^{+}20, ZZCD20, ZCD^{+}22, MLZ^{+}23]. malicious-aware [DZSL25]. maliciousness [DLC^{+}24b]. maliciousness-aware [DLC^{+}24b]. MAlign [SAR24]. malleability [KAK21]. MalSPM [NFVN^{+}22]. Malware [GMBN21, GM24, HLP22, JANR23, KC22, LTZ^{+}24, MDB^{+}20, OJL24, PLP^{+}21, QYJ22, RSW^{+}21, RKG^{+}25, SRN24, VBCM23, WLC^{+}22, YWX^{+}22, AMM23, ASG^{+}25, AYS20, AZ20, AZES21, AES22, AGI^{+}25, AGAS25, ADL25, BS24a, BLC^{+}24, BLDB24, BBD23, BM24, BCS^{+}21, BG25, BTAK21, BGA24, CAMTB23, CLX21, ÇÜD21, CRP22, CZLZ24, CXL^{+}25, CKV22, CLMU25, DLQ^{+}21, DÇÜD22, DWX^{+}23, DGS^{+}23, DLC^{+}24b, FFLQ23, FFLQ24, FGY^{+}24, FCA25, FSN22, FCSP21, GLM^{+}24, GPC^{+}22b, GCZ21, GDM^{+}24, GTGHAG22, GDG21, GWF^{+}24, GMP20, GMPMS21, GFM^{+}22, GTP^{+}25, GZH^{+}24, GMLB22, GMB22, GMLB23, HLX^{+}24, HSK22, IMMS21, JYL25, JHD^{+}20, JDBB20, JDBB22, JPLT20, JKR^{+}21, JLLG24, KDDM20a, KDDM20b, KAU^{+}23, KBMM23, KZYZ22, KDTR25, LL21a, LFCD21, LCZ^{+}22, LLL^{+}22c, LGZD23, LHL^{+}24a, LBD^{+}23, LSAH21, LWZ^{+}23, LdSP21, LLLZ20, LFW^{+}22, LKLZ24, LQL25b, LWP^{+}25, MMC24, MRG21, MCMMMA23]. malware [MHLZ22, NBBS22, NPCM24, NFVN^{+}22, NHT^{+}25, NZZ^{+}22, OS22, OSPP22, OMO20b, OX22, PYT20, PBdL25, PTD^{+}25, QZT^{+}22, QMZ^{+}25, RBHW25, RS23, RG23, SAR24, SBB^{+}20, SGSS22, STPB25, SFX24, SN24, ST25a, SGVA21, STM^{+}20, SSK^{+}24, TSAG24, TQJ^{+}23, TTA23, TY22, VAW^{+}20, VMS20, VCP25, WTW21, WLZ^{+}23, WBC^{+}23, WGS^{+}25b, XZWD25, XGS^{+}21, XSL^{+}24, XFY21, XLS24, YMR^{+}22, YGC^{+}23, YLH^{+}24, YWZ^{+}24, YWL^{+}20, YPT22, ZDH^{+}23, ZXL^{+}25, ZTYA25, ZSLY25, ZYL^{+}25, ZHZ^{+}22, ZWX^{+}24, ZCW^{+}24, BCO^{+}22]. Malware-SMELL [BCO^{+}22]. Malware2ATT&CK [SSK^{+}24]. Man [VBCM23, BDC^{+}23]. man-at-the-End [BDC^{+}23]. Man-in-the-Room [VBCM23]. Management [BHL^{+}23, MAKH21, AOM24, AHSZ21, AMD^{+}21, AZO^{+}23, AGBB22, BDC^{+}23, BSFB20, CYL^{+}24, CCE23, CZZ^{+}23, CGCY21, DHZ24, HZ20a, HZ22, JS25, yKK22, LLD^{+}25, LKLZ24, MLM20, MMvS20, MKH24, MBKD21, NTBH^{+}22, NSLM25, OANA25, QLZ^{+}25, RRB25, SBJ^{+}24, Sar21, SS23a, SAK^{+}24, TKS21, TFMP25, TLHY25, VBF21, XLG^{+}23, YL20, ZSS21, dPR23, PAM^{+}24]. manager [MBU24, NLW24, SSL20]. managerial [GAC20]. managers [CKT^{+}25, vdKSCY22]. Managing [BHL^{+}23, LHYW22]. mandatory [BRWF23]. MANET [DP25, JIG22, RKRS24, SY24]. mangle [VB20]. manifold [AFAkJS25, BTAK21, LHS21, ZLH22]. manifolds [GT24a]. Manipulate [ZLC^{+}23b]. manipulation [CCH^{+}23, KD21]. manned [YGG24]. manned-unmanned [YGG24]. manoeuvers [SGSS22]. Manual [RDS20]. manufacturing [ASF^{+}23, AAS24, CYL^{+}24, KF25, KEM25b]. many [Spr23]. map [CYL^{+}24]. MapperDroid [SLBG21]. Mapping [BCKN22, CLX21, GHZ^{+}24, HY21, LHC24, LWSW24, LWSW25, NR24, SGSCM23, SSK^{+}24, TV20, YLH^{+}24]. maps [NRS21]. March [FFLQ24, Ano20s, Ano21t, Ano22t, Ano23t, Ano24t, Ano25t]. margin [HMTC22]. MARISMA [MSSO^{+}25, RMS^{+}21]. MARISMA-BiDa [RMS^{+}21]. maritime [KDBB25, MSSO^{+}25, TKKF24, TAA^{+}25, YGB^{+}25]. MARK [PSP^{+}22]. market [ALD^{+}21, CSM^{+}24, VS20, WNI20]. marketing [AMA21]. marketplace [DZSCR24]. Markov [HuRMMW22, HMMW22, LLL^{+}20b, PGC^{+}24, YWL^{+}20, ZCJ^{+}20]. MARTE [HZZ20]. MASA [SLC^{+}21]. masculinity [Da 23]. Mask [DZZ^{+}21, ZDH^{+}23]. Mask-guided [DZZ^{+}21, ZDH^{+}23]. masked [RG25, WJT^{+}23]. masking [ATJ^{+}22]. massive [RRS21, ZLZ^{+}24]. masterkeys [GL22]. matching [FO25, LXC^{+}20, MBH20, ZZCD20, ZZJC20]. mathematical [ZUKH23]. matrices [PGC^{+}24, PL23]. matrix [CZZ^{+}21, LSAH21, MMJG24, NGJ^{+}22, Onw20]. matters [SRT25]. mature [SVP21]. Maturity [BHL^{+}23, SSHP21, AA21c, CRdMM^{+}25, KLSW24, NSA^{+}20, YLSJ23]. Max [BP24]. Max-Min [BP24]. maximizing [YWWH22]. maximum [HMTC22, LLX25]. maximum-margin [HMTC22]. maxout [PMA^{+}23, RVS24]. May [Ano20t, Ano21u, Ano22u, Ano23u, Ano24u, Ano25u]. MBS [HLP22]. MCTVD [DGS^{+}23]. MD [CTC^{+}25]. MDADroid [YLH^{+}24]. MDD [ZHG^{+}24]. MDD-FedGNN [ZHG^{+}24]. MDLC [BS24a]. me [BMAP24, HSL23, LWZ^{+}25, MBSE22]. Mealy [PMNL21]. meaningful [HSL23]. meaningful-engagement [HSL23]. Means [XHTZ20, GT24a, LSTL24]. measure [RCD25]. Measurement [PNFBH23, DYL^{+}23, GGK25, HLZ^{+}25, HZZ^{+}24, LLL^{+}22d, LRH^{+}21, SOZ^{+}23b]. measurements [KBC21]. measures [BvtH^{+}23, SWW^{+}25, WW25a, WFY^{+}20]. Measuring [CJKR21, CPP24, LWZ^{+}25, SGT24, ACRB23, GMH^{+}23, RCD23, SBL20]. mechanism [ADÖU^{+}20, CXL^{+}24, CWHH23, CXC^{+}24, DZ22, He25, HXZ^{+}23, JLQ^{+}20, KLZ24, LTZ23, LFL^{+}25, zLXZrxZ25, LCL^{+}20, LZX^{+}23, LJ22, LJO^{+}20, LLL23d, RGG^{+}24, RSÁGVV23, SCW^{+}22, SMM22, SLL23, WLDW24, YC25, ZGM21]. Mechanisms [ACA^{+}23, JR24, MK24, STB^{+}20, ZSWT24]. media [AGM20, ALZ^{+}20, AZO^{+}23, KM24a, NCPVRT23, ZLZC22, vdSFF20a, vdSF23]. median [WTC^{+}22]. Mediating [vdSF21, ADO23, OJH25, SS23a, WLKT23, vdSF23]. mediation [SAM25]. mediator [XSL^{+}24]. Medical [DV24, GXL25, GJCJ20, GLX24, HZX^{+}20, LLX^{+}20, SBL20, RNA22]. medium [AAFF24, CZK24, HR24, KFB^{+}25]. medium-sized [KFB^{+}25]. meet [BMAP24]. Mel [VAAG23]. Mel-frequency [VAAG23]. Meltdown [CZG^{+}22]. Meltdown-type [CZG^{+}22]. MeMalDet [MMC24]. MEMBER [LLL^{+}22b]. membership [CWZ^{+}24a, FYDH25, HXZ^{+}24]. memorable [GZG19, GZG20]. Memorizing [YKSK25]. Memory [CMFUA^{+}20, CCO^{+}20, HK25, AMM23, AN21, AOA22, BTAK21, DS20, DS21, JXYF25, MMC24, MPRB21, PKK^{+}21, QSA24, VAAG23, WS24, XLG^{+}23, YGK^{+}25]. memory-based [AMM23]. mental [XLL^{+}25]. MeNU [YKSK25]. MER [ZLDX25]. MER-GCN [ZLDX25]. Merkle [HZG20, HZG19]. message [LCZ^{+}25, XW20, YHY23a, ZY21]. messages [WBZ^{+}25]. Meta [PvSvdB25, AD23, GYZ^{+}25b, KTH20, WLLL24b, WHE^{+}23, ZJJS^{+}22, ZCPB23]. Meta-analysis [PvSvdB25, AD23, KTH20]. meta-heuristic [GYZ^{+}25b]. meta-learning [WLLL24b, ZJJS^{+}22, ZCPB23]. Metadata [SJ21, BS20]. metal [JKS^{+}25]. Metamorphic [NFVN^{+}22, GMPMS21, LSAH21]. Metaverse [RMG25, QAA23, SSG^{+}25a, VBCM23]. meters [GZ18, GZ20]. Method [CWW^{+}25, RDM20, TRM22, WIOO24, ACRB23, AEA24, AZ20, ATGK22, CTC^{+}25, CDJD24, CLL^{+}21, CWHH23, CXC^{+}24, CWZ^{+}24b, CKG22, CMG20, DLQ^{+}21, DGS^{+}23, FLLY24, GDM^{+}24, GYZ^{+}25a, GYZ^{+}25b, GGZ25, HYW^{+}21, HHSL20, HLL^{+}21, HLL^{+}25, HZZ20, HLYZ21, HZG^{+}23, HXX^{+}24, JXYF25, JZY^{+}25, KS20, KFP^{+}24, LLL^{+}22a, Lee23, LL21a, LC21, LHX23, LHL^{+}24a, LCZ^{+}25, LQHZ25, hLHhLfW21, LZC24, LLLZ20, LRH^{+}21, LHS21, LCC^{+}21, LFHX23, LGG25, LJ22, MLXC24, MFA20, MBB^{+}23, MGL^{+}20, MLQ^{+}23, NK21, NJ22, PLW25, PLW^{+}21, PGP^{+}24, PGC^{+}24, QZT^{+}22, QIN25, QMZ^{+}25, RZ20b, SWL24, SGS^{+}22, SGA^{+}23, TQJ^{+}23, TSW^{+}24, TKKF24, WLQ20, WTX^{+}22, WCZW23, WTL^{+}23, WCZ^{+}24, WLZ^{+}24, WCH^{+}25, WWS^{+}25, WLQ^{+}25, WQL25, WWZ24, XLWC24, XFY21, XHWL20, YZL^{+}20, YWZW24, YLH^{+}24, YWZ^{+}24, YWL25, YGW^{+}25, YCLZ24, ZWF^{+}24, ZCLW23, ZMF^{+}23, ZXH^{+}25, ZHH^{+}25, ZSLY25, ZLLY25, ZWB^{+}25, ZZW^{+}24, ZZZ^{+}23]. method-based [CKG22]. Methodological [JS25, CS23]. methodologies [BAD^{+}24, MPL^{+}25]. Methodology [BVB^{+}23, CHP21, OBK^{+}23, SBJ^{+}24, CDMO24, CR25, CFYBF23, FFPC22, HD25, KMH^{+}22, LZZ^{+}21, PFR20, dNPM^{+}21]. Methods [ZLZX24, BBA20, BKPZ21, BAC^{+}23, CZ24, CKS^{+}24, CSKD22, CCH^{+}22, EFI23, GWF^{+}24, GMB^{+}25, KKS^{+}22, Les21, PvSvdB24, STB25, SLZP24, SHSK20, TZZ22, WLHC21, WYP^{+}24, WRG^{+}21, YLL^{+}22, YK23, ZWDE24, ZJW^{+}22, ZWZX25]. metric [AQAK21, LZNZ24]. metrics [AIA23, EFI23, FF23, GYW^{+}23, OCJ20, PFR20, SRT25]. MFMCNS [ACKS22]. MFT [JXYF25]. MFXSS [LFHX23]. MGA [AA22a]. MGA-IDS [AA22a]. MHSA [CXC^{+}24]. micro [CCE23, CJL^{+}25]. micro-segmentation [CJL^{+}25]. microarchitectural [CJJ^{+}22, LLP25]. microbiome [MHE24]. Microgrid [KBD^{+}23]. micropayments [KK20]. microscope [BDC^{+}22]. Microscopic [NMCRB21]. microservice [PVFM^{+}21]. microservice-based [PVFM^{+}21]. microservices [JQYL22, MM23]. Microsoft [Dub24, KLA^{+}22]. MIDAS [RMVG25]. middle [LLG20]. migration [GMS20]. MIL [DMRV21]. MIL-STD-1553 [DMRV21]. MILS [CDN21]. mimic [LZCS22]. mimic-based [LZCS22]. Min [BP24, YWWH22]. min-entropy [YWWH22]. Mind [PKK^{+}21]. miner [YCMM20]. minimal [KLW^{+}23, ZDH^{+}23]. Minimized [GSYM22, CZZ^{+}21]. Mining [WTL^{+}21, GFH^{+}22, GWY^{+}25, KLW^{+}23, NFVN^{+}22, RLW^{+}20, SWL24, SAK^{+}24, TT20, WTL^{+}23, WT23, YCMM20]. minority [MC22]. MinSIB [GSYM22]. minutes [BMV24]. MIPS [ZZJC20]. MIRDETECTOR [LZY^{+}25a]. Misinformation [SAJP22]. mislabeled [KBMM23]. Misleading [ZM22]. missing [TRH21, WJC23]. Mission [GM24]. MitID [KA23]. mitigate [HSH24, PT25, RJ21, REHS24]. Mitigating [CSEJGUA25, HS25b, LB25, WZX23, ZTJ^{+}21, HWML25, MBU24, RFdS23, VBS^{+}25, WW25b]. Mitigation [KK25c, NN24, NPF^{+}23, WCOZ23, WFY^{+}20, ACD25, AK24a, DS23, ESTA^{+}25, FVF21, GR23, HNNN^{+}24, IHJZ21, JJS^{+}24, LJ22, MLS^{+}23, MAR^{+}23, Rs25, RCD25, SSK23a, SBV24, SDG20, TD21]. mitigator [PJL23]. MiTM [LPS23, AAS24]. MITRE [PL23, WS25a, ZFS25]. mix [MLM20]. mixed [DMB21, ZWDE24]. mixed-methods [ZWDE24]. Mixing [BCD21]. mixture [HMK^{+}20, HHE22, ZCLW23]. mixture-based [HMK^{+}20]. ML [JRT^{+}25, LFC^{+}25, LCL^{+}25, LLLZ20, RS23]. ML-based [JRT^{+}25, LFC^{+}25, LCL^{+}25, LLLZ20, RS23]. MLP [WLQ20]. MLP-based [WLQ20]. MMF [XLWC24]. MMM [HHE22]. MMM-RF [HHE22]. Mobile [ADL25, AA22b, ADO23, CHWH25, CDM^{+}20, GRJV24, GG21, IMMS21, JK21b, KJB^{+}25, KJR^{+}20, KKS^{+}22, LT25, LMZZ22, LQH22b, LPX20, MT21a, MB20b, OO21, RKA25, RCA23, RDHSB22, RDHS23, RSEK20, SMvH^{+}21, SSZ^{+}24, WSQ^{+}22, WCW^{+}23, XGL20, XQS25, YWZ20, YCWW23, ZLCA21]. MobileNet [RDPR25]. mobility [JGS^{+}24]. modal [CHK24, RDHSB22, TSW^{+}23, AYL25]. modality [FKS25]. Modbus [RMMP23, SPL25, WCZ^{+}24]. Modbus/TCP [RMMP23]. mode [FFPC23, TH21]. Model [AAB22, BAB^{+}25, BHL^{+}23, CDM^{+}20, JCHS24, JNKZ25, KC22, LCH^{+}24a, LCH^{+}24b, LWT24, OCMF23, TAES20, WDJ^{+}23, ZLF^{+}20, AIBKJ24, AAT23, AGM20, AZO^{+}23, AHC22, AES22, AJHA20, AOAA20, AXAL24, BvtH^{+}23, BC23b, BTS23b, BLR23, BS20, CXL^{+}24, CB22a, CDMO24, ÇÜD21, CLL^{+}21, CLCZ23, CWZL23, CLL24, CZL^{+}25, CWYW25, DZSL25, DLC24a, DPK20, DS21, EÖ25, FBR25, FF20, GYZ^{+}25a, GHZS23, GSS20a, GTL^{+}22, HHE22, HP23, HDS21, HWW^{+}25, HLL24, HZH^{+}24, IMMS21, JSK^{+}23, JRT^{+}25, JLB22, KK25a, Kem23, KLSW24, yKK22, KBMM23, KMAHU24, KG20, KS23a, LLL^{+}22b, LL22a, LZS^{+}23, LLJ^{+}23, LXS^{+}24, LTZ^{+}24, LXX^{+}25, LDW25, LCW^{+}22, LWZC23, LHZZ24, LGJW21, LFHX23, LWP^{+}25, LZA^{+}23, LJO^{+}20, LLZJ22, LWLT21, LWLL25, MSCJ21, MCS22, MLZ^{+}23, MG24b, MGG^{+}25, MHSK21, MKH24, MDR20, MV23, MAAA20, MRL22, MGG^{+}20, NSA^{+}20, PM25, PLW^{+}21, QJQ22, QMZ^{+}23, Rs25, RLL^{+}23, SKD23]. model [SB25, SZY^{+}22, SSG^{+}25a, SYF^{+}21, Ske22, Ske24, SS23b, Sun21, SSK^{+}24, SZCZ24, SCL^{+}24, TK25, TYAH^{+}24, TZZ^{+}21, TY22, TP25, VT25, VB24, Vu22, WNF20, WLHC21, WYW^{+}22, WX23, WK23, XLY^{+}20, XW20, XWJ^{+}24, YKN21, YHZ^{+}22, YZW^{+}20, ZUKH23, ZWW^{+}21, ZLH22, ZCLW23, ZZWF21, ZCZW24, ZSS21, dSVST23, vdKWH20, LWYG24, KIMJ23, XLY^{+}20]. Model-agnostic [LCH^{+}24a, LCH^{+}24b]. Model-based [JNKZ25, TAES20, AAT23, BLR23, CDMO24]. model-driven [SB25]. modeled [AEA24]. Modeling [CKG22, KBD^{+}23, KG20, NLA^{+}23, QYYH24, SS22, TA25, WCOZ23, YK22a, ZYH^{+}21, ZGNZ21, tBLLV21, HZZ20, KJHL22, KBE^{+}24, KLJE24, KBK24, KKM^{+}25, OUR^{+}23, RCA23, SHSK20, VF22, WTW21, WLZ^{+}24, XZWD25]. Modelling [ARR^{+}22, ATMN23, LM23, MISO^{+}23, CGBK24, HuRMMW22, HMMW22, MSMH21, NGJ^{+}22, NGC^{+}25]. Models [IMN22, SUC23, WLC^{+}22, ATJ^{+}22, AÇ23, AV25, BLDB24, BDM^{+}24, CCW^{+}24, CBFH20, DÇÜD22, DDCK22, DDW^{+}23, FXZ22, FBL20, GAJ25, GPC^{+}22a, GT24b, HMK^{+}20, JJW^{+}23, JAAS^{+}23, KDBB25, MSL^{+}24, NCN24, PSN^{+}22, PBG22, QY22, RZW^{+}20, SCBP24, TC23, VK22, WZZ^{+}22, WPAG24, XLL^{+}25, XT24, YMZ^{+}23, ZM22, ZDM^{+}25]. Moderated [SAM25]. moderating [LHH23, TLS24a, VM21, XWY21]. moderation [SAJP22]. modern [AAT23, IJR^{+}25, KJHL22, MBSE22, MLV^{+}21]. modern-day [MBSE22]. modification [BWB^{+}21]. modified [HFLZ23, WS25b, XYH^{+}20]. modular [RZT^{+}23, TA25]. modularity [CBPW25]. modulation [AL23]. Module [HSK22, ZXB^{+}24]. modules [MPRB21]. Modulo [MVVL^{+}24]. moments [RKG^{+}25]. momentum [HLQD23, MLC^{+}23]. MONDEO [SDA^{+}24b]. MONDEO-Tactics5G [SDA^{+}24b]. money [YLL23]. Monitor [Kha21, ACCS25]. monitoring [CSZQ23, DABB20, LCL^{+}20, Rep23, TLHY25, WGS^{+}25a, ZKAK23, ZCL^{+}25]. monolithic [NG23]. months [BMV24]. Moore [LQY^{+}20]. mooring [LBWV23]. moral [HSW25]. Morpheus [MYW24]. morphing [kJLcL24]. morphological [Onw20]. MOSDroid [STPB25]. MoSFPAD [RDPR25]. MOSTO [RMMP23]. MOTIF [JANR23]. motion [ZGS24]. Motivation [GOW24, vdKWH20, AS24, ADO23, BvtH^{+}23, CXC22, KS25a, KIMJ23, KKM^{+}25, NLW24, OCB23, SA22, VM21]. Motivation-Behaviour [vdKWH20]. motivations [BvtH^{+}23, OG23]. motives [OG23]. movement [DYL^{+}23, SKKG25]. Moving [NNA^{+}23, TV20, HZX^{+}21, HXZ^{+}21, MKM^{+}25, RGG^{+}24, RFdS23, TZZ^{+}21, ZCJ^{+}20]. MPC [JYL25]. MPDroid [ZSLY25]. MPDS [SYH^{+}22]. MPDS-RCA [SYH^{+}22]. MPSAutodetect [AAB22]. MPTCP [KD21]. MQTT [CFT22, MVH^{+}21, STS25]. MR [AC25a]. MRC [TYZ^{+}24]. MRC-VulLoc [TYZ^{+}24]. MSNFuzz [ZMZ^{+}25]. MTA [WCZ^{+}24]. MTC [YM22]. MTD [AHKJJ21]. MTProto [MV23]. MTS [SHH^{+}24]. MTS-DVGAN [SHH^{+}24]. MUD [MLZ^{+}23]. MUD-PQFed [MLZ^{+}23]. Multi [AZES21, BFR23, BAB^{+}25, CB20, DG20, FKS25, GTGHAG22, KZN^{+}25, KKK23, LFL^{+}25, LLL^{+}22d, RKA25, RKRS24, RZY^{+}24, RMVG25, SLZ^{+}25, SYH^{+}22, WXJ^{+}25, WGS^{+}25b, XCWY25, ZMZ^{+}25, ZGS24, ATA^{+}25, ACKS22, BBR20, CXL^{+}24, CTV21, CVZHN25, CXC^{+}24, CHK24, CKL^{+}25, DJKA25, DH25, ESTA^{+}25, GY23, GPK21, GS24, HRFS24, HFLZ23, HWW^{+}25, JN23, JJW^{+}23, JZY^{+}25, JMU^{+}25, KLP^{+}24, KJB^{+}25, LLL^{+}22b, LLL^{+}20a, LLX^{+}20, LZW^{+}24, LLD^{+}25, LWC^{+}25, LYS^{+}24, LFHX23, LLH^{+}25, LJO^{+}20, MJZ^{+}25, MLQ^{+}23, MGZ^{+}25, MR23, OZW^{+}21, RDHSB22, RSÁGVV23, RHH^{+}22, RJ24, SMCP21, SLC^{+}21, Sin25, SCCZ20, SWS^{+}25, TDYQ23, TSW^{+}24, TYZ^{+}24, TTA23, TWQ^{+}25, TSCM21, TC23, WZZW20, WTW21, WLL^{+}23, WYL^{+}24, WYP^{+}24, WYJ^{+}24, WWL^{+}25a, WZDZ25, WS21, XXZ^{+}21, XCY^{+}25, XGS25, ZNF^{+}23, ZLC^{+}23b, ZLH23, ZCJ^{+}20, ZZWF21, ZSL25]. multi-agent [TSW^{+}24, ZLC^{+}23b]. multi-attack [RKRS24]. multi-attributed [SLC^{+}21]. multi-audio [TTA23]. multi-behavioral [GTGHAG22]. multi-channel [WZDZ25]. multi-choice [TYZ^{+}24]. multi-class [TC23]. multi-classifier [ACKS22]. multi-cloud [ATA^{+}25, TSCM21]. multi-copy [ZSL25]. Multi-criteria [ZMZ^{+}25]. multi-datasource [LYS^{+}24]. multi-device [SMCP21]. Multi-dimensional [CB20, SLZ^{+}25, JN23, ZNF^{+}23]. Multi-domain [KZN^{+}25, DJKA25, Sin25]. Multi-domains [LFL^{+}25]. multi-environment [RJ24]. multi-faceted [ESTA^{+}25]. multi-factor [SCCZ20, WZZW20]. multi-feature [ACKS22, LFHX23]. multi-features [WYP^{+}24, WWL^{+}25a]. multi-focus [JJW^{+}23]. multi-granular [MJZ^{+}25]. multi-granularity [LLD^{+}25]. multi-graph [HFLZ23]. multi-head [CXL^{+}24, CXC^{+}24, XXZ^{+}21]. multi-hop [DH25]. multi-instance [WLL^{+}23]. Multi-intensity [RZY^{+}24]. multi-interest [MLQ^{+}23]. multi-keyword [XGS25]. multi-label [HRFS24, RHH^{+}22]. Multi-labeling [GTGHAG22]. Multi-layer [BAB^{+}25, LLL^{+}22d, BBR20, GS24, HWW^{+}25, RSÁGVV23]. Multi-layered [RMVG25]. Multi-level [BFR23, SYH^{+}22, KJB^{+}25]. multi-modal [CHK24, RDHSB22]. Multi-modality [FKS25]. multi-model [LJO^{+}20]. Multi-motion [ZGS24]. multi-objective [ZCJ^{+}20]. multi-operation [WS21]. multi-parameter [JN23]. multi-party [GY23, JMU^{+}25, LLX^{+}20]. multi-path [LLL^{+}20a]. Multi-Perspective [AZES21, WGS^{+}25b]. multi-plan [LWC^{+}25]. Multi-probability [XCWY25]. multi-prototype [WTW21]. multi-replica [GPK21, TWQ^{+}25]. multi-scale [ZLH23]. multi-server [WZZW20]. multi-source [CKL^{+}25, SWS^{+}25]. multi-stage [CVZHN25, MGZ^{+}25, WYJ^{+}24, ZZWF21]. multi-staged [JN23]. multi-stakeholder [MR23]. multi-step [LLH^{+}25, WYL^{+}24, XCY^{+}25]. Multi-strategy [WXJ^{+}25]. Multi-targeted [KKK23]. multi-task [LLL^{+}22b]. multi-type [KLP^{+}24, LZW^{+}24]. multi-user [CHK24]. Multi-view [RKA25, JZY^{+}25, OZW^{+}21]. multi-vocal [CTV21]. Multiclass [VMS20, DÇÜD22, SPN^{+}25]. multicontroller [SCOE23]. Multidimensional [WZW^{+}25, HXZ^{+}21]. multidisciplinary [CGCY21]. multidomain [YLW24a]. multifeatured [DLC^{+}24b]. Multifractal [NH21]. multigranularity [HQL^{+}22]. multilevel [XLWC24]. multilingual [MFI20]. Multimedia [ZHJ^{+}23, ZHJ^{+}21]. Multimodal [LTU^{+}21, ZZW^{+}24, GAS^{+}23, GHZS23, GMP20, XLWC24, ZSLY25]. Multimodal-based [ZZW^{+}24]. multinomial [HHE22]. multiobjective [CPR21, TH24]. multiparty [BK22]. multipath [KD21]. multiple [DZSL25, KAM^{+}21, KS23b, LWS^{+}25, MLM20, MLM^{+}23, WTX^{+}22, ZZL^{+}22]. multiple-mix-attack [MLM20]. multiple-source [WTX^{+}22]. multiple-type [ZZL^{+}22]. multiplication [MMJG24]. MultiRHM [JN23]. multiscale [GRJV24, LFW^{+}25]. multisets [STPB25]. Multistage [SDA^{+}24b, LGJW21]. MultiSVM [TD24]. MultiTLS [MLM^{+}23]. Multiuser [DZ22]. Multivariate [CLL24, GHMG24, TXY^{+}23, ATMN20, CVL22, TLY^{+}24, XLZ24, YWW24]. Multivocal [IO24, PVFM^{+}21, BOS^{+}21]. municipalitie [Sch24]. Municipality2HTTPS [Sch24]. Mutate [CAFMB24]. Mutation [WCZ^{+}24, LXJ^{+}25, LQL25b, WCZW23, WLLC24]. mutation-based [LQL25b]. Mutual [KKRP21, AMAA21, BWZ^{+}25, OO21]. mutually [SBY^{+}22]. MVX [ACD25]. MVX-based [ACD25]. my [AR25, CDG22, RCD21, SGJ^{+}25]. Myanmar [CC20].

n [HCL^{+}23]. n-gram [HCL^{+}23]. NAISS [REHS24]. naïve [GL21, Vu22, ZX23b]. name [LCZY23, LCW^{+}22]. named [ESM24, JLC25, WFY^{+}20]. names [AKCS20, CQC25, CNK25, SOZ^{+}23b]. NAmpKeeper [DHW24]. Narrative [TRH21, Gal20b]. narratives [HSL23]. Nash [WZW24a]. NAT [MSP^{+}20]. National [SEGD^{+}22, ASA^{+}22, LHGB21, TDM^{+}21, WLKHD25]. nations [MAAG22]. native [BCM^{+}25, CKL^{+}25, ZKL^{+}23]. natural [AA21a, AKAFQ22, AKA^{+}23]. naturalistic [BJZ^{+}23]. Nature [JRT^{+}25, KBAG23, Ade21, AAAK23, VF22, vdWLM24]. Nature-inspired [JRT^{+}25, KBAG23, AAAK23]. Navigating [BCH24, NTH25, ZWDE24, HR24]. navigation [DCS^{+}22, WBZ^{+}25, ZY21]. NB [ZX23b]. NCMFuzzer [WLLC24]. NDIF [LXH^{+}25]. NDN [KT25, LPW25]. NE-GConv [AWN^{+}23]. nearest [SYG^{+}21]. Need [QXS^{+}22, APK24, BCdGG20, SV25, SJ21, XLL^{+}25]. need-to-know [SJ21]. needle [WRG^{+}21]. needs [MSH^{+}25, SL20, UNBF21]. Negated [IMN22]. Negative [IMN22, ASSW23, YFO^{+}21]. neighbor [XZG^{+}25]. Neo [MT21c]. Neo4j [VWJ24]. NERO [MRSP24]. Net [MV25, KM24b, KLZ24, CHWH25, NMLN23]. NetFlow [CRP25, CJA25]. Netherlands [vdWLM24]. netlists [CBPW25, PLW25]. NetObfu [LZZX21]. Network [BBMW21, BAC^{+}23, BDA23, DP25, EYYZ20, GPRGMF25, GC25, GC22, GYYK22, HCL^{+}23, HXZ^{+}21, KWF25, LZDZ21, LXZ^{+}22, LZA^{+}23, MAK21, PHT^{+}25, PLB^{+}22, SPG^{+}25, WLC^{+}22, WDJ^{+}23, XLZ^{+}22, YLDW23, ZSGB^{+}22, AKWR21, AD21, AR21, AMVG24, ACKS22, AWN^{+}23, AC25b, AGI^{+}25, AJHA20, AGBB22, AOA22, BCH24, BD23, BTM21, CXL^{+}24, CEN^{+}25, CLN^{+}25, CWW^{+}23, CPRV21, CFZL22, CWX^{+}24, CXC^{+}24, CYC^{+}24, CZH^{+}25, CSKD22, CMCVGH^{+}23, DPKHP22, DCK^{+}23, DH25, DSD^{+}25, DYL^{+}23, DHL^{+}23, DZZ^{+}22, DC23, FHZ^{+}22, FWKB24, FGY^{+}24, FLLY24, FEA22, FF20, GSSC24, GCZ21, GDM^{+}24, GYZ^{+}25a, GYL^{+}20, GYZ^{+}25b, GRJV24, GTAK25, GHMG24, GCV22, GJCJ20, GFH^{+}22, GJB22, HHE22, HFLZ23, HZGM25, HZM21, HKA24, He25, HNNN^{+}24, HGSB24, JPLT20, kJLcL24, JKR^{+}21, JCJ^{+}21, KLC25, KDDM20a, KDDM20b, KK25b, KJI^{+}24, KKY^{+}23, KJB^{+}25, KSSL22, KS23a, KJ23, KZFM24, LT25, LLL^{+}22b, LWX^{+}22, LHD^{+}23, LYHC24]. network [LWSW24, LWSW25, LYLT24, LLL^{+}22d, LXH^{+}23, LXH^{+}25, LXJ^{+}25, LGH21, LZZX21, LFHH22, LWZ22, LFW^{+}25, LHW^{+}20, LLL23d, MSCJ21, MZW^{+}23, MLXC24, MNMS24, MG24b, MLQ^{+}23, MHC^{+}25, MVH^{+}21, MC22, MYH25, MDB24, MKZ^{+}23, NK21, NB25b, OFH^{+}25, PB25, PMA^{+}23, PGRS24, PSP^{+}22, PD25a, PGP^{+}24, PLP^{+}21, PRPS25, QYW^{+}25, QMZ^{+}23, QGW^{+}25, QZS^{+}25, QMW20b, QIN25, QSA24, RJ21, RYS25, RKA25, RVS24, RKRS24, SS22, dCFA^{+}23, SLR25, SMM22, SSG21, SZW^{+}25, SS25, SDG20, SGA^{+}23, SK24b, SYG^{+}21, SHH^{+}24, TK25, TSW^{+}24, TLKW20, Tek21, TY22, TP25, TC23, TD24, UWL22, VPM23, WNF20, WJT^{+}22, WGRZ23, WJT^{+}23, WML^{+}23, WZC^{+}24, WWWS24, WW25b, WXJ^{+}25, WTL^{+}24, WHW^{+}24, XCY^{+}25, XGX^{+}24, XLL23, XKY25, YLPZ21, YCMM20, YLL^{+}22, YWZW24, YCLZ24, YKY24, YGC25, YLZ^{+}25, ZMZ^{+}25, ZLY^{+}20, ZJW^{+}22, ZY24, ZLLY25, ZZHY25, ZFT24, ZZL^{+}22, ZYL^{+}22, ZWH^{+}25, ZLH23, ZHL^{+}20, ZXW^{+}25, ZCD^{+}22]. network [ZJJS^{+}22, ZCZZ24, ZWZX25, ZCW^{+}24, dSVST23, vDK22, HMBY23, LXW^{+}24, MMdSdS23, SRY^{+}25]. Network-based [HXZ^{+}21, ACKS22, BD23, DC23, GJB22, dSVST23]. network-edge [KSSL22]. network-independent [JCJ^{+}21]. networked [BCH24]. networking [DS23, DGD23, MHSK21, MKH24, WFY^{+}20, YPDC20]. Networks [HMT^{+}25, KJCL20, MBK^{+}22, QYJ22, ZJC^{+}25, AKWR21, AA22a, ADÖU^{+}20, AESF23, AK20, AB23, AGVA21, AWJB21, AOAA20, AK24b, BAB^{+}25, BK22, dCBJSdS23, CWZL23, CWW^{+}25, CWC^{+}25, DJKA25, DUK25, DWX^{+}23, DKV^{+}24, DTK^{+}21, ESTA^{+}25, GIO24, GRJV24, GLB^{+}22, GHZ^{+}24, HAKK21, HSX^{+}24, HZG^{+}23, ISM22, JQYL22, JGS^{+}24, JR24, KZN^{+}25, KKY^{+}23, KAO^{+}23, KJR^{+}20, KF25, KBAG23, LTL^{+}22, LKW^{+}20, LKA^{+}21, LCC^{+}22, LZL^{+}22, LYLT24, LWLL25, MLM20, MT21a, MBP20, MRG21, MBGF20, MT21c, MTD^{+}24, NRS21, NMLN23, NR22, PM25, PBG22, PYT20, PHT^{+}25, PGZB23, QMW^{+}20a, RKA25, Rs25, RHDF21, RFdS23, RZT^{+}23, RJ24, dCFA^{+}23, SJMP23, SLC^{+}21, SGJ^{+}25, SDA^{+}24b, SD25, SWYL20, WTX^{+}22, WMK23, WCW^{+}23, WML^{+}23, WZW24a, WP21, WS21, XCSZ21, XMK21, XLZ24, XLW24, XHW^{+}22, YMR^{+}22, YM22, Yan20, YMZ^{+}23, ZUKH23, ZM20, ZMQ21, ZMF^{+}23, ZLY^{+}21, ZFT24, ZGS24, ZCS^{+}23]. networks [ZLZX24, ZLH^{+}25, ZDM^{+}24, dPR23, dASJ^{+}23, LWS^{+}25, MBGF20]. networks-based [YMR^{+}22]. networks-CAN [AA22a]. Neural [AKWR21, BDA23, GYYK22, MMdSdS23, QYJ22, SRY^{+}25, WLC^{+}22, AGI^{+}25, BK22, BD23, CWX^{+}24, CXC^{+}24, DWX^{+}23, DH25, FHZ^{+}22, FGY^{+}24, FEA22, GDM^{+}24, GLB^{+}22, GJCJ20, GFH^{+}22, HSX^{+}24, HNNN^{+}24, HZG^{+}23, ISM22, JPLT20, JKR^{+}21, KZFM24, LKA^{+}21, LCC^{+}22, LZNZ24, LXH^{+}25, MRG21, MBGF20, MT21b, MKZ^{+}23, NRS21, PLP^{+}21, PRPS25, QYW^{+}25, QMW^{+}20a, RYS25, SJMP23, SZL^{+}24, SK24b, Tek21, TY22, TC23, VS25, WGRZ23, WCW^{+}23, WZC^{+}24, WW25b, XMK21, XCY^{+}25, XWJ^{+}24, XHW^{+}22, YMR^{+}22, ZY24, ZLLY25, ZZL^{+}22, ZYL^{+}22, ZCS^{+}23, ZLZX24, ZLH^{+}25, ZYGL25, ZJJS^{+}22, MRSP24]. neural-stacked [VS25]. neuroimaging [MT21b]. Neuronal [BCP22]. Neurosymbolic [KLC25]. neutralization [SPV20, STSV25, ZZ25]. Never [BOS^{+}21]. news [KFP^{+}24, MGG^{+}20]. Next [HD25, RCA20, ZTYA25]. next-generation [ZTYA25]. NFV [LLYL21]. NIDS [CZH^{+}25, FWKB24, HKA24, SS25]. NIDS-Vis [HKA24]. NIOM [JRT^{+}25]. NIOM-DGA [JRT^{+}25]. NIS [CSB^{+}25]. NKAB [Kac25]. NLDP [FNZ^{+}23]. NLP [GHZS23, SZY^{+}22, SZC^{+}21]. NLP-based [SZC^{+}21]. No [TMN24]. Node [AMNR20, AWN^{+}23, HZZ^{+}24, XSK23, ZSGB^{+}22]. node.js [ZG24]. nodes [JIG22, LTL^{+}22, XCWY25]. noise [AMN22, DZZ^{+}21, GSE20, HX21, HHL^{+}22, HWWB24, KLW^{+}23, YLY^{+}23, ZCS^{+}23]. noisy [ZW24]. Non [RXFZ21, TTP20, AD23, ABAM22, AGI^{+}25, BTM21, CWL^{+}22, DGD23, DS20, DPSN20, FSMM24, FYDH25, GGK25, HABW24, HSL20, KLNW21, LWZ22, MD22, PLN23, Sar21, TAA^{+}25, WLLC24, ZUKH23, dNPM^{+}21, LYSB23, MBKD21, uS25]. non- [GGK25]. non-/compliance [GGK25]. non-compliance [AD23, MD22]. non-control [DS20]. non-convex [KLNW21]. non-critical [WLLC24]. non-deterministic [DGD23]. non-directional [LWZ22]. non-English [DPSN20]. non-expert [PLN23]. non-experts [TAA^{+}25]. non-ICT [MBKD21]. Non-IID [LYSB23, uS25]. Non-Interactive [RXFZ21]. non-intrusive [dNPM^{+}21]. Non-Invertible [TTP20, HABW24]. Non-IoT [MK24]. non-IT [FSMM24]. non-linear [ABAM22, AGI^{+}25, ZUKH23]. non-parametric [BTM21]. non-repudiation [CWL^{+}22]. non-sensitive [FYDH25]. non-transferable [Sar21]. non-uniform [HSL20]. noncompliance [Nie24]. nonmalicious [JZ23]. nonnegative [LSAH21]. nonparametric [MCS22]. nonproprietary [MB20a]. NOP [GFM^{+}22]. normality [YKSK25]. normalizing [ISM22]. normative [PF20]. NoSQL [VWJ24]. notice [CAT24, DJG^{+}25, LWSW25]. Novel [ADÖU^{+}20, FEA22, GVJ23, GYG^{+}23, GBA22, AIBKJ24, AA25, AZO^{+}23, AEA24, APA^{+}24, BTM21, BS20, CR25, CZCX20, CZX^{+}21, CXC^{+}24, CWZ^{+}24a, DNB^{+}20, EÖ25, GLL^{+}24, GYL^{+}24, GKC22, HHE22, IB20, JKR^{+}21, JXYF25, KLZ24, KE25, LLL^{+}22a, LKA^{+}21, LLYL21, LLL^{+}22c, LXX^{+}25, LX25, LHG^{+}24, LLLZ20, LWL21, LGG25, LPS23, MSCJ21, MCS22, MT24, MSG22, MV25, MSP^{+}20, NKA^{+}25, NK21, NGC^{+}25, NGX^{+}23, OSPP22, PMNL21, PD25b, QMZ^{+}23, RJ24, SPG^{+}25, SLL23, Tek21, TY22, WTW21, XFY21, XKY25, YCCZ21, YLH^{+}24, YWZ^{+}24, YC25, ZLY^{+}20, ZLZ^{+}24, ZW24, ZXW^{+}25, ALZ^{+}20]. November [Ano20u, Ano21v, Ano22v, Ano23v, Ano24v, Ano25v]. novice [RCKB24]. NPFTaint [YLZ^{+}25]. NS [YFO^{+}21]. NSAPs [ZLY^{+}20]. NTFS [OPK20]. NTLFlowLyzer [SLR25]. NTP [TH21]. Nudges [QXS^{+}22]. Nudging [GZGG20, PSCT21, QXS23]. number [ACC^{+}20]. numerical [ZUKH23]. numerically [HHL^{+}22]. nursing [MVSSZ24]. nWFH [JRP24].

O [KJB^{+}25]. O-RAN [KJB^{+}25]. O2A [LLF^{+}23]. O2O [RDS20]. OASoSIS [KAFDW22]. OAuth [ABC22]. obfuscate [GFM^{+}22]. Obfuscated [KC22, AESF23, AMM23]. obfuscated-malware [AMM23]. Obfuscation [STPB25, AHuHS24, BT21, CLM^{+}25, CMG20, LYHC24, LZZX21, LXZ^{+}22, LHO^{+}25, ZXB^{+}24, ZTY^{+}20b]. Obfuscation-resilient [STPB25]. Object [Kha21, MQH^{+}23, AFLR25, CZCX20, CZX^{+}21, QZP23, ZM22]. Object-free [MQH^{+}23]. Object-Oriented [Kha21]. objective [ZCJ^{+}20]. objectives [PFR20]. objects [EZLC21, LMW^{+}25]. Oblivious [ANG^{+}25, AC25a, RRM20]. Observational [CCL^{+}24]. observer [RCD25]. observers [YHY23b]. occurrence [FBL20, SRN24]. October [Ano20v, Ano21w, Ano22w, Ano23w, Ano24w, Ano25w]. odometers [QLZ^{+}25]. ODRE [CCBGC25]. Off [CKS^{+}24, YCWW23, SBJ^{+}24, SYL^{+}20]. Off-policy [CKS^{+}24]. Off-The-Shelf [YCWW23]. Offensive [MDK^{+}23, ILVR^{+}24]. Office [Dub24, KLA^{+}22]. officer [Ram25, SV25]. offloading [HSH24, YWZ20]. offs [GTL^{+}22, MR23]. offset [XHWL20]. oil [MAR^{+}23]. OLE [Dub24]. omnidirectional [AGBT23]. On-chip [AL23]. On-manifold [ZLH22, AFAkJS25]. on-the-fly [GWY^{+}25]. One [YL23, GR24, RRMSM^{+}20, ZLC^{+}23b]. One-class [YL23, RRMSM^{+}20]. one-time [GR24]. One4All [ZLC^{+}23b]. Online [HF25, KDBS22, QXS23, TCM^{+}25, AJPK24, ATMN20, AJHA20, AOAA20, BFW21, CMG20, GG22, JR24, JB22, OA25, PLN23, PSCT21, SCE21, SCCZ20, SZC^{+}21, SZL^{+}22, WSW^{+}22, WTX^{+}22, WLZ^{+}24, WXZ^{+}25, ZNL^{+}20]. Onlooker [FVF21]. only [LLC^{+}23, RZW^{+}20]. Ontological [CBS25]. ontology [CSB^{+}25, DS22, HMV^{+}25]. ontology-based [HMV^{+}25]. OOD [LHZZ24]. opcode [MMdR24, STPB25]. Open [CYL^{+}23, GM24, HK25, MDB^{+}20, MBK^{+}22, WWH^{+}22, AMC23, AGVA21, DVAAM23, GYW^{+}23, KMAHU24, KRR23, KG20, NGM^{+}24, NDDH^{+}21, RCD21, SSRM25, SZY^{+}22, SPG^{+}25, SPL25, TZZY25, TLHY25, ZYL^{+}25, dPRU25, CCBGC25].

Open-Set [WWH^{+}22]. Open-Source [CYL^{+}23, GYW^{+}23, KG20, NGM^{+}24, NDDH^{+}21, TZZY25]. OpenC2 [MB20a]. operating [AlD25, HZZ^{+}24, KS23b, SGVA21]. operation [WS21]. operational [AGBB22, CFYBF23]. Operationalizing [TFMP25]. Operations [SCBM23, ACRB23, BRM21, FF23, MBR23, PGP^{+}24, SVP21, YQL24]. Operations-informed [SCBM23]. opinion [DPKHP22]. opinions [FWMJ^{+}23]. OPMonitor [WGS^{+}25a]. opportunistic [LT25, RHDF21]. opportunities [CBFH20, DCB^{+}21, HAKK21, LX25, MSL^{+}24, SDA25, YW21]. Opportunity [KLR^{+}25, vdKWH20, HR24]. optical [LWZ22]. Optimal [AA22a, GDG21, TZZ^{+}21, AC25b, AK24a, CYHW25, GPAF25, JYL25, JSYK25, LQL25b, WX23, ZCZW24]. OptimalLink [Rs25]. Optimisation [UYMM21, BK22, RMVG25, TH24, ZMLP23]. optimised [GSSC24, JRT^{+}25]. optimising [MMvS20, MC20]. Optimism [FSMM24, LHH23]. Optimization [BP24, DP25, LS20, WSW24, YPT22, AHuHS24, AAAK23, CLX21, CXLV20, CSKD22, EÖ25, GSSC24, GYZ^{+}25a, GYZ^{+}25b, Kac25, KM24b, LFL^{+}25, LWH^{+}24, NK21, PRTV22, PMA^{+}23, PGRS24, PRPS25, SDA^{+}24a, WWH^{+}22, WXJ^{+}25, BS24b, RVS24]. Optimization-based [LS20, WSW24]. optimize [CDG^{+}23]. Optimized [CSKD22, JZL^{+}24, KR21, MT21a, RYS25, STB25, BC23a, BTS23b, CZZ^{+}21, CZS^{+}21, KM24a, MMJG24, MV25, PRPS25, SY24, SK24a, SK24b, TK25, SRY^{+}25]. optimizer [YJK^{+}24]. Optimizing [PD25a, SBB^{+}20, SASW23]. Optiwords [GZG19, GZG20]. oracle [HCLRM20, LXW^{+}25a]. oracle-based [HCLRM20]. Orchestration [CBK22, SGSS22, TLHY25]. order [LYHC24, VMS20]. ordering [CDG^{+}23]. organisation [RKKG25]. organisational [NR24, PNF24, dVABH20]. organisations [AMA21, KGIS21]. Organization [APA^{+}24, ZWX^{+}20]. organization-friendly [ZWX^{+}20]. Organization-Specific [APA^{+}24]. Organizational [SA22, AMA22, ADM23, AABLS24, DHZ24, HHT23, HG21, JRP24, LX25, PF20, PJ23, VM21]. Organizations [MDK^{+}23, AMD^{+}21, BBZS24, DPKHP22, DPTAM25, ESA^{+}23, HNH^{+}22, NMT25, PNK20, vdKWH20]. organizing [NRS21]. Oriented [Kha21, DS22, GR23, KFZ^{+}20, MVSSZ24, OMO20a, STZ^{+}22, WLC^{+}20, ZHL25]. OSINT [TRM22]. OSS [CDF^{+}20]. OSSIntegrity [NGM^{+}24]. OSTIS [APA^{+}24]. OT [ARR^{+}22, GMB^{+}25, LWL^{+}23, LET^{+}25]. OTA [IJR^{+}25]. other [WT23]. Otupy [Rep25]. outcomes [HSL23, OJH25]. Outlier [NCM24, CPV23]. outlier-aware [CPV23]. OUTliers [IM23]. outlook [ZT21]. outsourced [JZL^{+}24, hYAS^{+}24, ZX23b]. outsourcing [CQL^{+}21, QYYH24, RXFZ21, RLW^{+}20, TLK^{+}24, ZZCD20]. over-granted [WGS^{+}25a]. overcome [LXLL25, SPV20]. Overcoming [BLDB24, uS25]. overconfidence [FJR23]. overflow [LZD^{+}20, WCH^{+}25, ZZL^{+}22]. Overlapping [XLZ^{+}22]. overload [DPTAM25]. overloading [KTGDE20]. oversampling [LGH21]. overview [ISR^{+}23]. OWASP [WK23]. own [PNK20]. Owners [XCL^{+}24, vdWLM24]. ownership [HSQ24]. Oxley [Wes20].

P [YHZ23]. P-E [YHZ23]. P2P [BC23a, CWZL23, WWZ24, XSK23]. P4 [AFK^{+}24, ESTA^{+}25, XZG^{+}25, YPDC20]. P4-based [XZG^{+}25]. P4-powered [ESTA^{+}25]. P4-to-blockchain [YPDC20]. P4Drop [XYZ^{+}25]. P4NSA [XZG^{+}25]. P4Pi [DUK25]. PAC [WP21]. packet [CZSW24, CQLW25, FLLY24, QZS^{+}25, YPDC20, ZDM^{+}24]. packets [XYZ^{+}25]. packing [GTP^{+}25, VDL24]. page [BA20, CZG^{+}22, VRS^{+}25]. page-Table [CZG^{+}22]. Pages [BDA23, GZ20, GZG20]. pairwise [AQAK21]. Palm [LWZL25]. PalmNet [KLZ24]. palmprint [KLZ24]. pandemic [HvCSJ23, LSN^{+}21, vdWLM24]. panels [AWB24]. paperless [KDBB25]. papers [Pit25]. paradigms [CRS^{+}22]. paradox [ARB20, LLC^{+}23, MT21b, WP21]. Paradoxical [SSK22]. parallel [JLQ^{+}20, KLP^{+}24, PLW25, VDN24]. Parameter [LWYG24, JN23]. parameters [CLX21, CZLZ24, SVA21]. parametric [BTM21]. parser [YPDC20]. parsers [WWL25b]. parsimonious [SU25]. parsing [YLZ^{+}25]. Partial [YHY23a, STM^{+}20]. partially [LLL^{+}25]. participation [FKL23]. particle [CSKD22]. partitioning [DLL24, KDDM20a, KDDM20b, SCC^{+}25, WHW^{+}25]. party [GY23, JYF^{+}25, JMU^{+}25, LLX^{+}20, LB25]. pass [GRKS24]. passcode [DSS^{+}25]. Passive [ASI^{+}20a, ASI^{+}20b, KFHM24, MT24]. Passive- [ASI^{+}20a, ASI^{+}20b]. passive-active [MT24]. Passphrase [BF20, MFI20]. PASSVM [ADJS21]. Password [GBA22, WZX^{+}22, CKT^{+}25, DPSN20, Fur22, GZ18, GZG19, GZ20, GZG20, GZGG20, HMB^{+}21, LZCS22, LP24, NLW24, SW22, SPV20, WS24, XRW^{+}25]. password-strength [GZ18, GZ20]. passwordless [KA23, MBU24]. passwords [GZG19, WS25b, YHZ^{+}22]. PasswordTensor [SW22]. past [CCH^{+}22, VWB^{+}25]. Patch [DV24, MCM^{+}24, QWW^{+}22, TZZY25, ZXZJ25]. patches [FKS25, PDML25, SCL^{+}21, SKS25, WCL25]. PatchView [FKS25]. patents [CMA23]. Path [AA23, ADÖU^{+}20, AK24a, LLL^{+}20a, hLHhLfW21, MLB^{+}25, SLDL23, SYC^{+}24, WLQ^{+}25, XKJG25]. path-based [WLQ^{+}25]. paths [SKG21, XKJG25]. PathWatcher [WLQ^{+}25]. patient [Esm20, RNA22]. pattern [JK23, LKL23, NFVN^{+}22, RMS^{+}21, SLR25, SKM21, YK23]. pattern-based [SKM21]. patterns [BBD^{+}24, GHZ^{+}24, JDBB20, LP24, QMZ^{+}25, WTL^{+}21, HK25]. Paying [RSÁGVV23]. payload [HHSL20, LFHH22]. Payment [CB22b, AVR20, FXZ22, MSMH21, MT21c, vDK22]. payments [DSC20, vDK22]. PCA [BFC^{+}25, MMJG24, RXFZ21]. PCA-Based [RXFZ21, BFC^{+}25]. PCaaD [GDK^{+}21]. PDCleaner [JZY^{+}25]. PdGAT [ZWB^{+}25]. PdGAT-ID [ZWB^{+}25]. PDHF [ZCZW24]. PDSMV3 [PD25b]. PDSMV3-DCRNN [PD25b]. PE [AGAS25, FSN22, GWF^{+}24, LWZ^{+}23, NRKK23]. Peculiarities [CDF^{+}20].

Pedometer [DJG^{+}25, DJG^{+}23]. PEDR [LLZJ22]. peer [WMGH^{+}23]. peer-to-peer [WMGH^{+}23]. PeerG [WWZ24]. PeerRemove [XSK23]. Penetration [HSOL23, CHZ^{+}23, LSMK25, LZS^{+}23, LZY^{+}25b, SA25, WLZ^{+}24]. penetrations [WFKH25]. PenGym [NHFB25]. Penterep [LSMK25]. pentesting [NHFB25]. people [BTH22, LBWV23]. perceived [KMK^{+}24, vdSF23]. Perception [LBW^{+}25, KKS^{+}22, PL23, VBF21]. perceptions [APK24, LdGW^{+}22, NJS^{+}24, PF20, RCD21, RCD23, SAAW21, TLS24b]. perceptron [RSÁGVV23]. Performance [BAD^{+}24, MYS^{+}25, ACRB23, BAC^{+}23, CFRB25, DCB^{+}21, FF23, HLX^{+}24, ISR^{+}23, JJT^{+}24, LHGB21, MAP23, OCMF23, PTD^{+}25, RMS^{+}22, SK20, SCBP24, ZFT24, dNPM^{+}21]. Perhaps [RvdSM23]. periodic [ZWB^{+}25]. permission [HZ20a, LKK25, LL23, MCvO24, NCN24, OMO20a, WJB23, XLG^{+}23]. permission-based [HZ20a]. permission-carrying [LL23]. permission-related [MCvO24]. permissioned [ASB^{+}21, DCB^{+}21, SJX^{+}20]. permissionless [LLZY20]. permissions [MB20b, SLBG21, WGS^{+}25a]. perpetrator [XZLG25]. Persistence [VHMGRR22, MG24a, WPZW25]. Persistent [AASWC25, HZ20b, CZZ^{+}23, CZPX22, ISR^{+}23, QYYH24, RBHW25, WPZW25, XGL20, XLWC24, ZXH^{+}25, ZGNZ21]. person [MCM^{+}24]. persona [RSSD23]. Personal [KOO24, SGS^{+}22, CSEJGUA25, COH21, GRdAS23, MHE24, PdARD24, WP21]. personality [FF20, GZGG20]. Personalized [QXS^{+}22, CH25, GZGG20, LC21, LFL^{+}25, SSZ^{+}24, SZWH25, ZW24]. personally [WPS20]. Perspective [AZES21, LKS^{+}23, AS24, ASM20, ABA23, Als20, BDC^{+}22, BBZS24, CBB25, FKP^{+}24, HHZ21, HQL^{+}22, IFHP22, KKHK24, KLP^{+}24, LBWV23, Les21, MBKD21, NLW24, OUR^{+}23, RKKK25, RZ24, SSG25b, TSW^{+}21, VB20, WGS^{+}25b, YHZ23, ZWW^{+}21, ZLZ^{+}24, ZZ25]. Perspectives [GM24, PLN23, dVABH20, AAFF24, HSB^{+}22, HKY^{+}25, QS22, SSK23a]. persuasion [WSW^{+}22]. persuasiveness [XW20]. perturbation [CZX^{+}21, GPD^{+}23, ZDH^{+}23]. perturbations [LJJ22, LLX25, MBSE22, VS22, WSW24, WLLL24a, WLLL24b]. perturbed [AKA^{+}23]. phantom [HYR^{+}22]. Phase [DKSS20, LLL^{+}23a, LLZJ22]. Phenomenon [XLZ^{+}22]. phishers [LKK25]. PhishHunter [WZC^{+}24]. Phishing [BDA23, TJD23, XXZ^{+}21, dSFG20, AA21a, AXAL24, AMM^{+}21, BT24, BFW21, BMH21, BMV24, BJZ^{+}23, BAJ25, BA20, BTH22, CPP24, DPSS23, FF20, GMBV21, GGR^{+}24, HHT23, HvCSJ23, KM24b, LKK25, LHH23, LXH^{+}23, LGJW21, MG24b, MSA24, NPF^{+}23, PC24, PD25b, RGC24, SARG23, SAM25, TCY^{+}20, VRS^{+}25, VWB^{+}25, VS25, WZC^{+}24, WLDW24, YAC25, ZYL^{+}22, ZCZW24, vGCHZ24]. PhishingRTDS [AXAL24]. PhiUSIIL [PC24]. phone [ZXH^{+}25]. photoplethysmography [WZDZ25]. PHP [XLW24]. PHP-based [XLW24]. PHRiMA [HZ20a]. phylogenetic [CDM^{+}20]. Physical [AA23, CEN^{+}25, CLN^{+}25, CCO^{+}20, He25, KBD^{+}23, LYL^{+}23, LKS^{+}23, PT25, PC22, SMvH^{+}21, YGW^{+}25, ZKS25, AHFL25, AJ20, AMVG24, BHO23, BTE23, CS23, CPT^{+}22, CYL^{+}24, CPV23, DZ22, GRKS24, GKC22, HZX^{+}21, HZ20b, HLYZ21, ISD24, KGK23, Kil25, KEM25b, LZZ^{+}21, RKKK25, SAAW21, SMM22, SBV24, SKG21, SHH^{+}24, TAES20, TA25, TWLK22, VLV25, XZWD25, XHW^{+}22, YHS24, YC25, ZHL25, bOJW20, NJS^{+}24]. Physical-layer [He25]. Physical-world [LYL^{+}23]. physically [CCS^{+}24, KEM^{+}25a]. PIGNUS [JSK^{+}23]. pillars [Bel25, EÖ25]. PIN [BKPZ21]. PIN-entry [BKPZ21]. pipeline [KBMM23]. pipelines [BG25, PTD^{+}25]. PISA [WYJ^{+}22]. pitfalls [BCdGG20, BCS^{+}21]. PIVAD [DS20]. Pixel [WYJ^{+}22]. pixels [Gur25, ZQCX24]. PIXHELL [Gur25]. PKI4IoT [HLFR20]. PKIs [GMY22]. place [MTN^{+}23, WK20]. plan [LLL^{+}20b, LWC^{+}25]. plane [KKY^{+}23, WLC^{+}20]. planning [CMG20, MGZ^{+}20, WLZ^{+}24]. plant [RDM20, SRM22]. plants [SRM22]. plate [QMW^{+}20a]. platform [Arc20, LLJ^{+}23, QIN25, SZC^{+}21, ZKL^{+}23, ZYJ^{+}23, CBK22]. Platforms [BVB^{+}23, AZO^{+}23, DCB^{+}21, RDS20, ZSS21]. Plausible [RKW^{+}22]. play [BHW24]. playbooks [SCBM23, TFMP25]. PLC [MSD^{+}23, XWW^{+}20]. PLC-SEIFF [XWW^{+}20]. PLC-VBS [MSD^{+}23]. PLS [CVZHN25, OXNA25]. PLS-SEM [CVZHN25, OXNA25]. PM [XGS25]. point [CCS^{+}24, HKY^{+}25, LCB^{+}22, LLZJ22, MV25, TMN24]. pointers [LCB^{+}22]. Points [CYHW25, BDM^{+}24, GGZ25]. poison [MZWM23, ZLC^{+}23b]. Poisoning [GYG^{+}23, KDBS22, LLL^{+}23a, BAD^{+}25, HYZ^{+}24, MDB24, XHWL20, YHW^{+}23, ZYH^{+}21, LWYG24]. poker [RCD21]. polar [HZM21]. Policies [MARB^{+}23, ANG^{+}25, ABA23, BDLP24, CCBGC25, GZGG20, HN25, HCCS22, KSMW23, LLL^{+}22d, MLS^{+}23, MYMC20, Nie24, ND23, PNK20, RKSV20, RKG20, RKKG25, SST21, WGS^{+}25a, WP21, ZCAP23, ZZ25]. Policy [GBA22, ADM23, AD23, AS24, BBZS24, CPT^{+}22, CDG22, CKS^{+}24, CZZ^{+}21, CXC22, CJL^{+}25, Esm20, GZG19, GZG20, GWY^{+}25, HSW25, HKM^{+}23, JZ23, KKP22, LL23, MFI20, MD22, MHSK21, MKH24, MAAG22, NM24, PLS20, SGT24, SWL24, SPV20, TN21, WCH^{+}25, WPAG24, XWY21, vdSPR24]. Politeness [HG20]. political [DANS20]. Pollination [MS23]. Pollution [YLDW23, ZG24]. polymethodological [CYL^{+}21]. Pool [YCLZ24]. poor [BPL^{+}20]. population [KIAV22]. port [GKB21]. portable [MMdSdS23, Rep25]. portfolio [LL22b]. position [ACD25, SAB^{+}25, YLZ^{+}21]. position-heap-based [YLZ^{+}21]. position-independent [ACD25]. position-specific [SAB^{+}25]. positioning [SC23]. positive [ASSW23, RG23]. positives [VK22]. possibility [BKS22]. possible [MDR20]. post [MM21, WLZ^{+}24]. post-exploitation [MM21]. post-penetration [WLZ^{+}24]. potential [BFC^{+}25, CSRA22, NGJ^{+}22, ZMF^{+}23]. Power [IMN22, NTBH^{+}22, XRW^{+}25, AGVA21, DPKHP22, HZX^{+}21, HXZ^{+}21, KMAHU24, KTGDE20, LX21, MH25, SSN22, TWLK22, VLV25, ZLLY25]. Power-ASTNN [ZLLY25]. powered [ESTA^{+}25]. powerful [PCR20]. PowerShell [ZLLY25, AAB22]. PPG [LCP^{+}23]. PPIS [GJCJ20]. PPT [CWZL23]. PQFed [MLZ^{+}23]. Practical [ABC22, KDBS22, QXS^{+}22, SZL^{+}22, ZXL^{+}25, BM24, CQC^{+}24, JYL25, LLL^{+}23b, NOVJ23, SLX^{+}23]. Practically [WTHL25]. Practice [ATMN23, AMD^{+}21, ATMA23, Als20, CCE23, Da 23, DSBL23, FKL23, JJS^{+}24, SHK^{+}21, VB20, WJC23]. practices [BSFB20, DPTAM25, Fur22, LP24, MMvS20, MEH^{+}24, PNF24, UNBF21]. practitioner [SSG25b]. practitioners [SSHP21, SL20]. Pragmatism [GHOS22]. Pre [MPRB21, CJD25, DÇÜD22, LHG^{+}24, WYF^{+}24, XFY21, YMZ^{+}23, ZSLY25]. pre-attack [CJD25]. Pre-processing [MPRB21]. pre-trained [DÇÜD22, WYF^{+}24, YMZ^{+}23]. pre-training [LHG^{+}24, XFY21, ZSLY25]. Precaution [SSK23a]. precautions [GMRB23, LHH23]. Précis [BDM^{+}24]. Precise [GAL^{+}20a, LCX^{+}25, LFC^{+}25]. Precision [DLL24, LCX^{+}25, CWZ^{+}24a, He25, JPL20, ZYGL25]. Precursor [LZH^{+}24, OCB21]. predict [ASI^{+}20a, ASI^{+}20b, CFRB25, FJR23, RGC24]. Predicting [HHP^{+}24, KKR^{+}24, Wes20, ZLY^{+}21, ZZL^{+}22]. Prediction [EYYZ20, SD25, AHFL25, AZ20, AES22, CWZ^{+}24a, CJA25, DMZ24, HHL24, HDS21, KRR23, SK24b, TWW21, WYJ^{+}24, WBN^{+}20, WWW^{+}22, YZL^{+}24, ZLY^{+}20, ZTY^{+}20b, dSFG20]. predictions [CHP^{+}24, JCHS24, STB25]. predictive [FBL20, KKM^{+}25]. preference [ZLY^{+}21]. preferences [AR25, IFHP22]. prefix [ZW24]. preparedness [CZK24, LK20]. preprocessing [ZSGB^{+}22]. presence [CJJ^{+}22, CL20]. presentation [OFIdD^{+}20, RDPR25]. preservation [FGWJ20, LZC24, LHH^{+}20, MJ25, RKES23, TZZ22, TSW^{+}21, VS20, WTX^{+}22, ZNF21]. preserve [GYL^{+}20]. Preserving [CBK^{+}20, CCL^{+}21, ELD24, GYYK22, JKSS22, KEB23, LZG^{+}22, STH^{+}22, TLKK20, ZTR^{+}23, AN20, ÁFM25, ARAG23, CSZQ23, CWZL23, CLM^{+}25, DSS^{+}25, DDCK22, DLL24, DTK^{+}21, FGH^{+}21, GY23, GRKS24, GLL^{+}24, GJCJ20, GLX24, GXL25, HLL^{+}25, HZX^{+}20, HSQH23, HYZ^{+}24, ILB20, JPL20, JMU^{+}25, KAM^{+}21, LLX^{+}20, LYW^{+}21, LQY^{+}20, LRL^{+}22, LLL^{+}23c, MLZ^{+}23, MMJG24, MZA^{+}20, MZLK25, NZT^{+}25, PPGC24, PHH^{+}20, PJ23, QWLL23, RHDF21, RNA22, SMCP21, SBY^{+}22, SJX^{+}20, SYH^{+}22, SGY23, SS23b, SRT25, TLK^{+}24, Vu22, WZCP20, WK20, WZB^{+}23, WZW^{+}24b, XGS25, XYG25, YM22, YWZ20, YLZ^{+}21, YZL^{+}24, YFO^{+}21, ZSZ^{+}25, ZXZ^{+}20, ZNL^{+}20, ZX23b, ZLZC22, ZZCD20, ZFWL22, ZTW25, uS25]. pressure [BTH22, CAT20a, CAT20b, CAT24, JZ23, MKL21]. pressure-based [MKL21]. prevails [ADO23]. Prevalence [vdWLM24]. prevent [HKM22]. Preventing [TH21, JYS21]. Prevention [HKM22, ADÖU^{+}20, BAJ25, Kem23, LLLB22, LLG21, MG24b, RKA25, RW24, SST21, vdKWH20]. Preventive [LL22b, MBKD21]. prey [PGRS24]. prices [LLJ21]. pricing [LYW^{+}21, SGS^{+}22]. PRIDS [ZDM^{+}24]. principles [FRF24, KRN24, MBF25]. principlist [FWR21]. printer [NMCRB21]. printing [NMCRB21]. prior [LZS^{+}23, NTH25, SLX^{+}23]. prioritisation [TCJS22]. prioritization [CKS^{+}24, JLB22, MWM^{+}25, SSN23, WYL^{+}24]. Prioritizing [TTRY20]. priority [hLHhLfW21]. Privacify [WPAG24]. Privacy [ÁFM25, CSZQ23, CBK^{+}20, CLM^{+}25, DSS^{+}25, DJG^{+}23, DJG^{+}25, DLL24, FGH^{+}21, GRKS24, GT24a, GPAF25, GTL^{+}22, GJCJ20, GYYK22, HSQH23, JKSS22, KEB23, LLX^{+}20, LQY^{+}20, LLL^{+}22d, LZG^{+}22, MLZ^{+}23, MJ25, MZLK25, NZT^{+}25, NHC25, OSTO20, OANA25, PJ23, RGACPA23, SGY23, STH^{+}22, TLK^{+}24, TSW^{+}21, Vu22, WZCP20, WR24, XGS25, YWZ20, ZNL^{+}20, ZZCD20, vdSFF20b, vdSPR24, AIBKJ24, AR25, AGM20, AA25, AZO^{+}23, AVR20, ARB20, ATA^{+}25, AMN22, ASAA21, AN20, AZDF20, ACC^{+}20, ARAG23, BC23b, Bel25, BCKN22, BT21, CZS^{+}21, CWZL23, CYHW25, COH21, DDCK22, DZSCR24, ELD24, Esm20, FGWJ20, FMA22, FNZ^{+}23, GY23, GYL^{+}20, GLL^{+}24, GSLS21, GCN^{+}24, GGZ25, GXZ23, GYL^{+}24, GLX24, GXL25, HZGM25, HP23, HLL^{+}25, HSB^{+}22, HSSPK21, HNZ^{+}23, HHL^{+}22, HLL24, HZX^{+}20, HYZ^{+}24, HYR^{+}22, ILB20, IFHP22, JB22, JMU^{+}25, KLNW21, KAM^{+}21, KEK^{+}21, KSMW23, KLC23]. privacy [LYW^{+}21, LC21, LCA^{+}23, LLL^{+}23b, LLC^{+}23, LSTL24, LZH^{+}24, LFL^{+}25, zLXZrxZ25, LCL^{+}20, LBL^{+}20, LZC24, LHH^{+}20, LRL^{+}22, LWL^{+}23, LSLL24, LB25, LLL^{+}23c, LHO^{+}25, LWO^{+}25, MCS22, MMJG24, MVA25, MZA^{+}20, MHE24, MB20b, MVSSZ24, MT21b, Olu22, PEGA23, PPGC24, PHH^{+}20, PLN23, PFHB21, QLZ^{+}25, QIN25, QWLL23, RKES23, RCKB24, RMG25, RHDF21, RNA22, RGPAF20, RSEK20, SS24a, SMCP21, SBY^{+}22, SC23, SGT24, SDA25, SJX^{+}20, SSG^{+}25a, SGS^{+}22, SHL^{+}20, SRCP22, SVWJ25, SYH^{+}22, SSZ^{+}24, SS23b, SvSZ21, SYL^{+}20, SWK20, SWW^{+}25, TSAG24, TLS24a, TLS24b, TMJ^{+}23, TGOF22, TZZ22, TLKK20, VS20, WK20, WTX^{+}22, WLLS22, WZB^{+}23, WZW^{+}24b, WASA24, WZW^{+}25, WPAG24, WP21, XHTZ20, XT24, XYG25, YM22, YJK^{+}24, YLZ^{+}21, YZL^{+}24, YC25, YFO^{+}21, YHY23b, ZSZ^{+}25, ZJL^{+}20, ZXZ^{+}20, ZTR^{+}23, ZCL^{+}25, ZZP^{+}25, ZX23b, ZW24, ZLZC22, ZCS^{+}23, ZFWL22, ZZ25, ZTW25, bOJW20, uS25, vdSFF20a]. privacy-aware [DZSCR24, MCS22]. privacy-based [KEK^{+}21]. Privacy-centered [RGACPA23]. privacy-driven [GCN^{+}24]. Privacy-Preserving [GYYK22, JKSS22, KEB23, LZG^{+}22, STH^{+}22, ÁFM25, CSZQ23, CLM^{+}25, DLL24, FGH^{+}21, GRKS24, GJCJ20, HSQH23, LLX^{+}20, LQY^{+}20, MLZ^{+}23, MZLK25, NZT^{+}25, PJ23, SGY23, TLK^{+}24, Vu22, WZCP20, XGS25, YWZ20, ZZCD20, TLKK20, AN20, CWZL23, DDCK22, GY23, GLL^{+}24, GXL25, HZX^{+}20, HYZ^{+}24, ILB20, JMU^{+}25, LYW^{+}21, LRL^{+}22, LLL^{+}23c, MMJG24, MZA^{+}20, PPGC24, PHH^{+}20, QWLL23, RHDF21, RNA22, SMCP21, SBY^{+}22, SJX^{+}20, SYH^{+}22, WK20, WZB^{+}23, WZW^{+}24b, XYG25, YM22, YLZ^{+}21, YZL^{+}24, YFO^{+}21, ZSZ^{+}25, ZXZ^{+}20, ZX23b, ZLZC22, ZFWL22, uS25]. privacy-privilege [SSG^{+}25a]. private [BQA^{+}23, FFPC22, FFPC23, FFB^{+}23, GMY22, GSS^{+}20b, GPD^{+}23, HX21, HYZ25, ISM22, JZL^{+}24, KGIS21, KKB^{+}25, LGL21, LLL^{+}23b, PJ23, RPR^{+}21, SVA21, STZ^{+}22, SZL^{+}22, YZP^{+}25, YSZ24, YHY23b, ZNF21, ZNF^{+}23, ZYPT22, vDK22]. privilege [SSG^{+}25a]. PRIVIUM [SSG^{+}25a]. PrivRS [YZP^{+}25]. PrivySharing [MZA^{+}20]. Proactive [RGG^{+}24, WS25a, ZFS25, CR25, HZ20b, CBK22]. Probabilistic [HMMW22, FSCM23, KJHL22, LLL^{+}20b, LDW25, SY24, HuRMMW22]. Probability [LWZ^{+}24, KT25, XCWY25, ZSJ^{+}25]. Probability-based [LWZ^{+}24, KT25]. probing [LZ25, WCH^{+}25]. problem [JDBB22, LLC^{+}23, LWP^{+}25, TT20]. problems [LCA^{+}23]. proces [HuRMMW22]. Process [ETAB22, JKSS22, CDF^{+}20, HDS21, HMMW22, JYCL25, LHH23, MMvS20, MSH^{+}25, MEH^{+}24, PAM^{+}24, SDG20, SRM22, WTHL25, WWS^{+}25]. processes [KF25, KS23b, RGPAF20, SG21, ZCJ^{+}20]. processing [AA21a, CXL^{+}25, FF20, MPRB21, MLV^{+}21]. processor [GYW^{+}23]. processors [CCVMZ24, DCSY25, KSC23]. procmonML [MWR20]. procurement [LLJ21, NM21]. Product [ZTJ^{+}21, RVJ25]. products [JCS^{+}25]. profile [NR24, PC24]. profiles [BKV^{+}23]. Profiling [Sma25, GPAF25, RSSD23, SLR25, ZPP20]. profitability [MCC^{+}24]. program [CGH^{+}23, HSL20, WJP^{+}23, WZG24, XWJ^{+}24]. Programmable [AFK^{+}24, XYZ^{+}25, XWW^{+}20]. programme [WS22]. Programming [CZE^{+}22, AASWC25, AC25a, JAAS^{+}23, WLC^{+}20]. Programs [Kha21, AMA21, JWY^{+}25, LL23, VS24]. progress [Ano20p]. project [MVVL^{+}24, NDDH^{+}21, TCJS22]. Projects [GHOS22, CDF^{+}20]. promise [SEGD^{+}22]. prompt [XCWC25]. Proof [BWZ^{+}25, TXH^{+}23, WSC^{+}20, BBBB^{+}20, TO22, XCC^{+}25, YL20, YZW^{+}20]. proof-of-concept [BBBB^{+}20]. proof-of-X [YZW^{+}20]. propagation [ARR^{+}22, JYL25, UWL22, WX23, XZWD25, ZUKH23, ZGNZ21]. propensity [TLS24a]. properties [WHLS24]. Property [SDCL25, AAS24, NCW^{+}23]. proportion [ZHH^{+}25]. proposals [BSR^{+}25]. Proposed [AN21]. proposition [CRdMM^{+}25]. proprietary [SYZ^{+}23, YGW^{+}25]. prosocial [CXC22, PKPC24]. prospect [JSYK25]. prospects [KFS^{+}22, SSR^{+}24]. protect [AR25, AZDF20]. Protecting [BvtH^{+}23, BAD^{+}25, BHW24, WK20, ZW24, IE22, KBAG23, QMZ^{+}23, SU21, hYAS^{+}24, ZCS^{+}23]. Protection [Da 23, GOW24, OG23, AS24, ADO23, BDC^{+}23, BvtH^{+}23, BT21, CS23, CLCZ23, CH25, CSLK24, CBFH20, FMA22, GPAF25, GGZ25, GYL^{+}24, HABW24, HD25, HCLRM20, HHL^{+}22, HLL24, KS25a, KIMJ23, KLZ24, KKM^{+}25, LC21, LWZL25, LBL^{+}20, LLL^{+}22d, LWL^{+}23, LSLL24, MSSO^{+}25, MDKR25, MR23, NLW24, OCB23, QIN25, SA20, SA22, SYD^{+}25, SWW^{+}25, VCD22, VM21, XZG^{+}25, XX20, YSZ24, ZXH^{+}25]. protection-capable [HHL^{+}22]. protective [AABLS24, BvtH^{+}23, BMM22]. protects [BMV24]. Protocol [LCCZ24, OYZ^{+}20, ZY21, AIJ23, AMAA21, ASB^{+}21, CWZL23, DZSCR24, GLX24, KOO24, LQY^{+}20, LLZY20, LPS23, MS23, MSMH21, NR22, PJL23, PHH^{+}20, RNA22, SCH^{+}20, SYZ^{+}23, Sch24, SZW^{+}25, VRS^{+}25, WSC^{+}20, WCZW23, WLLC24, YM22, YGW^{+}25, ZMZ^{+}25, ZGG^{+}25, ZWH^{+}25]. Protocols [ARR^{+}25, BRV^{+}22, KK25c, ATMN23, AN20, AFLR25, BCH24, BLR23, CCL^{+}24, CDGS23, He25, LXX^{+}25, MM25, ZSZ^{+}25]. prototype [QZT^{+}22, WTW21, ZG24]. prove [HZG19, HZG20]. Provenance [HP20, RG25, DJKA25, JZY^{+}25, LCY^{+}21, WWS^{+}25, SRT25]. Provenance-based [RG25, SRT25]. ProvGOutLiner [WWS^{+}25]. provider [Lee23]. providing [KFB^{+}25, KJB^{+}25]. provision [UHK^{+}21]. proximity [AAAAS20]. Proxy [OYZ^{+}20, REHS24]. PrSLoc [YHY23b]. pruning [LXJ^{+}25, ZTR^{+}23]. pseudo [WJP^{+}23]. PSI [HYZ25]. PSO [RJ24, SK24a]. PSO-D-SEM [RJ24]. PSOGSARFC [BBMW21]. psychological [AABLS24, BKA20, DPTAM25]. psychology [WJC23]. psychometrics [GMSA24]. PT [CHZ^{+}23, LZY^{+}25b]. PTB [XHW^{+}22]. public [AOA22, CPRV21, CKK^{+}23, GPK21, HLFR20, HSQH23, ILB20, Kem23, KDE20, KGIS21, Kil25, RDHSB22, SSRK20, VS20, ZJY^{+}23]. public-good [VS20]. publicly [WCL25]. publishing [DDCK22, NZT^{+}25, WZCP20, WK20, ZNF21, ZW24]. PUF [DKSS20, KK25a, LL21b, MH25]. PUF-based [LL21b]. Pufferfish [LCL^{+}20]. PUFs [ISD24, PHH^{+}20]. pull [LBWV23]. Pulse [GAJ25]. pupil [CCC^{+}24]. puppet [LMW^{+}25]. purchase [HSSPK21]. purchasers [LLC^{+}23]. purification [CWW^{+}23]. purpose [KDE20]. push [LBWV23]. push-pull-mooring [LBWV23]. PVFL [YZL^{+}24]. PVIDM [ZZCD20]. pwnable [KJPP23, KJPP23]. Pwnable-Sherpa [KJPP23]. Pyramid [YCLZ24, WTL^{+}21, ZYL^{+}22]. Python [ESM24, MFD25].

Q [LZDZ21, ZCZZ24]. Q-network [LZDZ21, ZCZZ24]. QoS [SYL^{+}20]. QPUF [XLD25]. QR [AÇ23]. QRNG [XLD25]. QSEE [KM22]. quad [LGL21]. quad-tree [LGL21]. quadratic [DLL24]. Qualcomm [KM22]. Qualifications [Ram25]. Qualifying [SBL20]. quality [BRWF23, CHA^{+}24b, LLJ21, LXW^{+}25a, SLZP24, YWL25]. Quantifiable [IB20]. quantification [SRN24, YQL24]. Quantifying [HMB^{+}21, PBG22, MR23, XT24]. quantile [HKM^{+}23]. Quantitative [LZZ^{+}21, tBLLV21, AR25, Kil25, MDR20, WK23, ZYJ^{+}23]. Quantized [MLZ^{+}23]. quantum [BCH24, BCGC25a, BCGC25b, BFC^{+}25, Mit20]. quantum-safe [BCH24, BCGC25a, BCGC25b]. queries [PCK20, QWLL23, SYG^{+}21]. Query [DDW^{+}23, RZW^{+}20, AHL^{+}21, DFJ^{+}23, LQY^{+}20, PEGA23, SGY23, VWJ24, WZW^{+}24b, XSC^{+}23, XGS25, YLZ^{+}21, ZYGH24]. Query-efficient [DDW^{+}23, RZW^{+}20, XSC^{+}23]. querying [LWZ^{+}25]. quest [SVP21]. Questioning [WS25b, XLL^{+}25]. queuing [MYS^{+}25]. QUIC [DG25]. quo [BSFB20].

R [HLYZ21, XGS25]. R-CNNs [HLYZ21]. RA [KFZ^{+}20]. rabbits [GRJV24]. RADAMS [HZ22]. RADIANT [KWF25]. Radio [MBK^{+}22, Yan20]. RAF [MLB^{+}25]. RAF-AG [MLB^{+}25]. RAIDS [SKD23]. raising [SEGD^{+}22]. RAM [RRM20]. RAN [KJB^{+}25, TLHY25, CDJD24]. Random [ÇÜD21, WFK25, XYH^{+}20, HWWB24, LCZW20, SLL23, Yan20]. randomization [WLC^{+}20, JKS^{+}25]. randomized [KLZ24]. randomly [PMNL21]. Randpay [KK20]. Range [GPD^{+}23, ZYGH24, DHL^{+}23, GMP24, LLZJ22]. Range-based [GPD^{+}23]. Ranges [RCA20, YKG20, YK22a, YK22b]. ranking [MDKR25]. RanSMAP [HK25]. RansoGuard [CJD25]. RansomSpector [TML^{+}20]. Ransomware [ACCS25, BBA^{+}21, BGM^{+}23, CSEJGUA25, CB22b, FACHC21, MBS20, AM25, BAAM23, BvtH^{+}23, CSvdA^{+}23, CDJD24, CJD25, DMB21, FXZ22, FK22, FK24, GAJ25, GKS24, HK25, JAAS^{+}23, KKB^{+}25, LL22b, MKC^{+}21, MKC^{+}23, MLS^{+}23, MCC^{+}24, MMdSdS23, MTN^{+}23, ST24, TML^{+}20, YP21, ZJJS^{+}22, HK25]. Ransomware-as-a-Service [MBS20]. ransomworm [ACKS22]. rapid [ST25b]. rat [PGRS24]. rate [LRH^{+}21, MZWM23, WX23, WCZ^{+}24, WW25b, ZKS25]. rates [ASSW23, RBM^{+}25]. Rating [ZTJ^{+}21]. ratio [MZWM23]. Rationality [Het21]. raw [SAR24, SASW23]. raw-byte [SAR24]. RBD24 [CEO^{+}25]. RCA [CZCX20, SYH^{+}22]. RCA-SOC [CZCX20]. Rcryptect [LsJC^{+}22]. RCVaR [FKvdA^{+}24]. RDPCF [GPD^{+}23]. re [ZWW^{+}21]. re-compressed [ZWW^{+}21]. reachability [BCM^{+}25, JYF^{+}25, QWLL23, SGY23, XGS25]. REACT [HFK^{+}24]. reactions [ALD^{+}21]. Reactive [KWF25]. read [LCBC23]. readiness [AA21c, APPN23, JRP24, NMT25]. Reading [GHZS23, GGR^{+}24, TYZ^{+}24]. ready [SCF^{+}25]. Real [AYL25, ACLA22, JLQ^{+}20, LsJC^{+}22, MVA25, AYS20, AXAL24, BMAP24, BTM21, JZQ^{+}22, KKGK25, LGJW21, LZD^{+}20, MFA20, Mit20, SBV24, VDN24, WW25a, XHW^{+}22, YGC25, ZM22]. real-life [KKGK25]. Real-time [AYL25, JLQ^{+}20, LsJC^{+}22, MVA25, AXAL24, BTM21, MFA20, SBV24, VDN24, YGC25]. Real-world [ACLA22, JZQ^{+}22, LZD^{+}20, Mit20]. Realistic [DCK^{+}23, HS25a, NHFB25, VPM23, BG25, MCMMMA23, SB25]. realities [AAFF24]. Reality [VBCM23, MBR23, OLH^{+}23, QAA23, WRG^{+}21]. realizable [CCS^{+}24]. realization [SA25]. really [HKM^{+}23, KMP^{+}20]. realm [KGK23]. Reasoning [ZLDX25, KUA^{+}25, LGG25, MRSP24]. recipient [KK20]. reciprocity [BH25a]. ReckDroid [CQLW25]. RecMaL [YGC^{+}23]. Recognition [CBK^{+}20, ACLA22, ACD^{+}24, BC23b, BPCZ23, BA20, BQA^{+}23, CZX^{+}21, CWZ^{+}24b, CWYW25, CKV22, ESM24, JLC25, KXZ^{+}23, KKK23, KN23, LLL^{+}20b, QMW^{+}20a, RXFZ21, SAAW21, STH^{+}22, TLK^{+}24, WTW21, ZHL^{+}20]. Recognize [DJG^{+}25, DJG^{+}23]. recommendation [MGL^{+}20, MLQ^{+}23, RSEK20, SVA21, ZD21, ZH24]. recommendations [GC25, US20, WS22]. recommender [HSB^{+}22]. reconfigurable [CLN^{+}25, LZCS22, ZHL25]. reconnaissance [QJDD24]. Reconsidering [STSV25]. Reconstruct [DMM^{+}23]. Reconstruction [Dub24, BH25a, CLL24, WJT^{+}23, XCY^{+}25, XYG25, ZGZL23]. record [HSL20, SHL^{+}20]. record/replay [HSL20]. records [KAM^{+}21, MHE24]. recover [YCCZ21]. recovering [CWZ^{+}24b]. recovery [AAJP25, ABAJ25a, ABAJ25b, LZCS22, LLL23d]. Rectify [YGC^{+}23]. Recurrent [JPLT20, RK24, GTAK25, LZDZ21, PRPS25, RYS25, XLW24]. recurring [HZL^{+}23]. recursive [KAO^{+}23, LWZ^{+}25]. red [CQLW25, RCD21, WLZ^{+}24]. red-hot [RCD21]. Redefining [VM21]. Reducing [BBZS24, CB22b, VK22, WW25a, RVJ25]. reduction [RGP23, SMC^{+}21, ZLH^{+}25]. redundancy [AHKJJ21]. redundant [LXJ^{+}25, ZCAP23]. Reentrancy [GHZ^{+}24]. Refactoring [SKL23]. Reference [JANR23, ZZJC20]. references [JCS^{+}25]. referencing [AHL^{+}21]. refining [PBG22]. reflected [CPR21]. Reflection [LJO^{+}20, CJA25, GAL^{+}20a]. refocusing [CZCX20]. RegGuard [GR23]. regime [BRWF23]. region [ACC^{+}20, BSFB20]. regional [ASM20, LHGB21]. regions [GGZ25, LHO^{+}25]. registers [GR23]. Registration [GC22]. regression [HKM^{+}23, KDTR25, KIAV22, LSLL24, SRN24]. regularization [LMD^{+}22]. regulation [CXC22, Coj25]. regulations [Olu22]. regulatory [MSL^{+}24]. Reinforcement [ZHZ^{+}22, CPR21, CKS^{+}24, CZS^{+}21, CTH^{+}22, GFM^{+}22, KS25b, LLF^{+}23, LZS^{+}23, LHD^{+}23, MM21, NHFB25, RMCT23, RK24, SS24b, TSW^{+}24, WYL^{+}24, WLZ^{+}23, XCY^{+}25, XSK23, YHS24, ZLC^{+}23b]. reinterpreted [WBC^{+}23]. Related [CHP21, CMA23, DAA24, HHZ21, LdSP21, MBB^{+}23, MCvO24, Nie24]. relation [GY23, JK23, MS25, WLH^{+}24, YWWH22]. relation-based [YWWH22]. relational [LGG25, PBA24, WLL24]. relations [AKH24]. relationship [CZM^{+}25, DAA24, MYMC20, SWL24, WMC20]. relationships [WNI20, ZYH^{+}22]. relativity [GMB22]. release [zLXZrxZ25]. releases [ZWL^{+}25]. releasing [SKS25]. relevant [PF20]. reliability [AUIH25, JS25, KUA^{+}25, WQL25]. Reliable [KL24, LFHY23, WFK25, dSVST23]. Religions [RD23]. RELM [PD25a]. relocation [SLL23]. remain [Spr23]. remediation [WTHL25]. remember [WS25b]. Remote [AFLR25, GC22, KBC21, ND23, PGC^{+}24, KFS^{+}22, Rep25]. removal [XSK23]. remove [AMN22]. Removing [AKWR21]. rendezvous [Yan20]. renewable [OCJ20]. Repackage [MRSV21]. Repackaging [MRSV21]. reparameterisation [LWSW24, LWSW25]. repeat [LLC^{+}23]. repetition [WCZ^{+}24]. replace [LBWV23]. replacement [GGZ25, WCH^{+}25, XHB^{+}21]. replay [GR24, HSL20, LZY^{+}25b]. replica [GPK21, TWQ^{+}25]. replication [CPRV21, FSCM23]. Report [MLB^{+}25, GOW24]. reported [WCL25]. reports [FKvdA^{+}24, SOZ^{+}23b]. Representation [CBS25, WJT^{+}22, AV25, DYW24, DC23, FWF^{+}23, FHZ^{+}22, LCZ^{+}25, LZY^{+}25a, LXH^{+}23, MWM^{+}25, OFH^{+}25, RG25, WLL24, WWLH23, WWZ24, ZLH23]. Representations [GSG23b, GB23, XLL^{+}25, ZYL^{+}22, ZDM^{+}24]. repudiation [CWL^{+}22]. reputation [JS25, NPCM24, SSL20]. repute [WSC^{+}20]. request [OMO20b, WJB23]. requests [LKK25, LZ25, OMO20a]. require [KK20].

Requirements [HW25, MARB^{+}23, RKG20, AR25, FVV23, HMV^{+}25, KKP22, NSA^{+}20, Olu22]. requiring [BCP22]. rerouting [MAP23]. rerouting-resistant [MAP23]. Research [BRV^{+}22, SLZP24, ZHJ^{+}21, ATMA23, AA21b, BBA^{+}21, BCS^{+}21, BOS^{+}21, GB21, GGK25, GCS22, HAKK21, JDB^{+}23, KKAS21, KYKI22, KMAHU24, Lee23, MM25, PNF23, RMG25, RSLM25, SSK23a, SV25, WIOO24, WJC23, WZG24, ZJW^{+}22, vdSPR24, WLDW24, ZHJ^{+}23]. residual [CLL24, RYS25, WGS^{+}25a, XKY25]. residuals [LWLL25]. Resilience [CCL^{+}22, CCO^{+}20, LKS^{+}23, MH25, AAJP25, AIJ23, Coj25, DSBL23, DHZ24, ESBJ20, PBG22, SSN22, XCSZ21]. Resilient [BTM21, HZ22, HGSB24, JP25, MWR20, OO21, STPB25, WLC^{+}20, Yan20, dPRU25]. Resist [WZX^{+}22]. resistant [BKPZ21, MAP23]. resisting [LWZL25, SYH^{+}22]. ResNet [MZW^{+}23]. ResNeXt [CWHH23]. resolution [BCP22, JMU^{+}25, LCZY23, LCZ^{+}24, LWZ^{+}25]. Resource [ZZW^{+}21, AEA24, CSvdA^{+}23, HLX^{+}24, JSYK25, JE21, OCJ20, XX20, ZKS25, ZKAK23]. resource-constrained [CSvdA^{+}23, ZKS25]. resource-efficient [JE21]. responders [AASC25]. Response [BHL^{+}23, DANS20, GVJ23, SJH^{+}21, AMD^{+}21, AMAA21, AWB24, BTE23, BMH21, HFK^{+}24, KLC25, KLC23, MGTK22, NNA^{+}23, NTH25, RVJ25, Rep23, SVP21, SCBM23]. responses [FVF21, HSW25]. responsibilities [PF20, Ram25]. restoration [WTC^{+}22]. restriction [DZZ^{+}21]. restrictions [DFJ^{+}23]. restructuring [SDG20]. result [SGY23]. results [SEGD^{+}22, XZWD25]. resurrection [CCL^{+}21]. Rethinking [AASC25, HMTC22]. Retraction [DJG^{+}25, LWSW25]. retrievability [TXH^{+}23]. retrieval [ADDH25, WCW^{+}23]. Retrospect [ZT21]. reusable [SVWJ25]. reuse [ACD25, PLW^{+}23, TTP23]. reused [WS25b]. Reveal [QXS^{+}22]. Revealing [AAFF24, YAC25]. Reverse [XMK21, HP20, REHS24]. Reversing [GBG20]. Review [GM24, JTF^{+}24, KKJ^{+}21, Les21, MVA25, MARB^{+}23, SM23, YW21, ZTJ^{+}21, AAJP25, ATMA23, AWVG20, AA21b, AMC23, ALD^{+}21, AKA25, BKPZ21, BHW24, BB20, BOS^{+}21, CMLM25, CS23, CCVMZ24, CTV21, CLHT24, DVAAM23, EL22, ESBJ20, Gal20b, GB21, GG22, GGK25, GMB^{+}25, HSB^{+}22, HKM22, HW20, IHJZ21, IO24, JR24, KGK23, KBK24, KGIS21, LYLT24, MMSJ25, MSA24, MHLZ22, NM24, NPF^{+}23, NRS21, NB25a, OK23, OANA25, PNK20, PNF23, PVFM^{+}21, PAA^{+}24, PvSvdB24, RMG25, RSLM25, RKKK25, RDS20, SV25, SGSCM23, SDA25, SAK^{+}24, SM24, SL20, SHSK20, ST25b, TNN^{+}22, TRH21, TMJ^{+}23, VB24, YLL^{+}22, ZSZ^{+}25, ZG24, vdSPR24]. Reviewer [Ano21x, Ano22x]. Reviewing [WTL^{+}24, EMS23]. reviews [EMS23]. Revisited [AA23, PCR20]. Revisiting [JCS^{+}25]. revocable [DSS^{+}25]. revocation [AHSZ21, KJR^{+}20]. revolution [AA21c]. rewards [BMH21]. rewriting [MYMC20]. RF [HHE22, KFHM24, LBD^{+}23]. RFE [KAO^{+}23]. RFP [LMD^{+}22]. RFP-CNN [LMD^{+}22]. RGB [XLS24]. RHEMAT [GAS^{+}23]. rhetorical [JDB^{+}23]. Rhysida [KKB^{+}25]. richer [ZCW^{+}24]. Riemannian [LHS21]. right [JK23, LCB^{+}22]. right-angle [JK23]. Rights [CCBGC25, PF20]. RIME [WXJ^{+}25]. ring [TZZ^{+}23]. Ripple2Detect [LLH^{+}25]. RISC [BF23, DCSY25, GYW^{+}23]. RISC-V [BF23, DCSY25, GYW^{+}23]. Rise [VBCM23]. Risk [BSFB20, EYYZ20, EAN^{+}22, SP20, TCM^{+}25, tBLLV21, vSRW^{+}20, AM25, AZO^{+}23, AVR20, AJ20, ARS^{+}25, ASI^{+}20a, ASI^{+}20b, BT24, BCGC25a, BCGC25b, BDC^{+}23, BKS22, BAD^{+}25, CB22a, CEO^{+}25, CCE23, CR25, CNK25, CGCY21, DGH21, DHZ24, EÖ25, FSMM24, FBL20, GMB^{+}25, GKB21, HLL^{+}21, HZ20a, HSH24, JNKZ25, KAFDW22, KTH20, LdGW^{+}22, LCZY23, LET^{+}25, MWL23, MR23, NGJ^{+}22, NSLM25, RVJ25, RMS^{+}21, SEM^{+}23, SGSCM23, SS23a, SAK^{+}24, Sma25, SDG20, TZZY25, TAES20, TLS24a, TKKF24, VBF21, WW25a, WNF20, WLM25, YGB^{+}25, ZMLP23, dPR23, vdSFF20a, vdSFF20b]. risk-aware [AJ20]. risk-based [BAD^{+}25, CB22a, MWL23]. Riskio [HMPS20]. risks [BCH24, HZKN22, LHYW22, MSL^{+}24, MDR20, PGZB23, WLKHD25, XT24, vdSF23]. risky [MDR20]. RisQFLan [tBLLV21]. RL [DG25, JYL25, JGS^{+}24]. RNN [AK24a, CJD25, LZHL20, SLDL23]. RNN-based [CJD25]. RNNIDS [SSG21]. road [SYG^{+}21]. roadmap [JYL25]. robot [LJL^{+}25]. Robotic [NHC25, YZLS25]. robotics [RKW^{+}23]. robots [LZX^{+}23]. Robust [AES22, AGAP24, CHWH25, GAS^{+}23, GR24, LWW^{+}21, MRL22, QYJ22, SKD23, SVWJ25, SUC23, SZWH25, SCC^{+}25, WJT^{+}23, WZDZ25, XHW^{+}22, CWW^{+}23, DSS^{+}25, ESTA^{+}25, FLLY24, GRJV24, GW24, HGSB24, JYCL25, KAM^{+}21, KDDM20a, KDDM20b, KK25b, LLL^{+}25, LX21, LXW^{+}25b, MT21a, PAGDG24, QJQ22, QZS^{+}25, SS25, VC25, WMGH^{+}23, YK23, ZKAK23, ZXL^{+}24]. robustness [AMVG24, BAC^{+}23, CCL^{+}22, CHP^{+}24, GXZ23, HMBY23, HMTC22, HKA24, ISM22, KDTR25, LYSB23, LFHY23, LZNZ24, MAR22, MPL^{+}25, MCM^{+}24, MBGF20, NG23, PTD^{+}25, SCH^{+}20, YHH^{+}24]. rock [MTN^{+}23]. rogue [LLZJ22]. Role [ARS^{+}25, ASA^{+}22, ADO23, CCE23, CPRV21, CGH^{+}23, DPTAM25, FK23, HS25b, KLW^{+}23, LHGB21, LHH23, NNA^{+}23, OJH25, SS23a, TLS24a, TMN24, TN21, TT20, VM21, WLKT23, YZP^{+}25, YHZ23, vdSF23]. Roles [GHOS22, Ogb21, AABLS24, BMH21, ND23, XLL^{+}25]. ROMEO [BSG23]. Room [VBCM23]. root [LCZ^{+}24]. rootkit [YHC25]. ROP [KJJ^{+}24]. ROSTAM [MBU24]. rotation [CRS^{+}22]. route [CMG20]. Router [SOZ^{+}23a, QPL^{+}23]. routine [LdGW^{+}22]. routing [ATMN23, AK24a, DP25, DZSL25, HYR^{+}22, LT25, MS23, OCJ20, RJ21, XZLG25]. Royale [BS24b, GRJV24]. RP [WCH^{+}25]. RPFUZZ [LXJ^{+}25]. RPL [GVJ23, SVPM21]. RRIoT [RK24]. RSSI [GGC24]. RSSI-based [GGC24]. rule [CDG^{+}23, CLMU25, LQHZ25, MAAA20, Nie24, PKPC24, TK25, VK22]. rule-related [Nie24]. Run [MM23]. Run-time [MM23]. runtime [AHFL25, CXLV20, KFZ^{+}20, LKK25, ZCJ^{+}24]. Rust [BS25].

S [DYL^{+}23, GMSA24, RJ24]. S-BEDIM [DYL^{+}23]. S-DATE [RJ24]. SA [VS25]. SABDTM [HZZ^{+}24]. SAD [KE25]. SAE [LSG^{+}24]. safe [BCH24, BCGC25a, BCGC25b, uS25]. SAFE-IDS [uS25]. Safeguarding [ARR^{+}25, KM24a, SK24a, MYW24, MSSO^{+}25]. safeguards [PEGA23]. Safety [CCO^{+}20, DMPT25, ESA^{+}23, OANA25, OXNA25, PSCT21, SBJ^{+}24, SSG^{+}25a]. Safety-security [DMPT25]. sailfishos [TK20]. saliency [DTT^{+}22]. Salp [KE25]. sample [CWX^{+}24, LXY^{+}24, QZT^{+}22]. samples [CWZ^{+}24b, DTK^{+}21, GTGHAG22, HWWB24, SZY^{+}22]. sampling [SLL23, XCWY25, ZSJ^{+}25]. sampling-based [XCWY25]. samplings [BMH20]. sanctions [HKM^{+}23]. sanctity [ATGK22]. sand [PRPS25]. sandbox [SGSS22, VCP25]. sandboxes [LFW^{+}22]. sandboxing [NG23]. sanitizable [LRL^{+}22]. sanitizer [YGK^{+}25]. Sarbanes [Wes20]. SatCom [CG24]. satellite [ARS^{+}25]. satellites [DBG24]. satisfaction [KMK^{+}24]. Satisfiability [MVVL^{+}24]. saturated [WX23]. Saudi [ARB20]. SBA [WAPA25]. SBTDDL [MSG22]. SCA [SCF^{+}25]. SCADA [AMC23, SKM21, US20, WFKH25, dCPR25]. scalability [PB25]. Scalable [JP25, KDDM20a, KDDM20b, BHO23, LZS^{+}20, SWYL20]. Scale [CHP21, CJKR21, AC25b, BPL^{+}20, CBPW25, CYL^{+}23, CPP24, GSG23a, GJCJ20, HHT23, PLW25, dCFA^{+}23, SJ21, SCC^{+}25, WCL25, XGL20, YWW24, ZLH23]. scaling [UÇ22]. scams [BTH22, XWZ^{+}20]. SCAN [HNZ^{+}23]. scanner [BLSS22, ZLC^{+}23a]. scanners [AHS25]. scanning [MSD^{+}23]. scapegoat [HSX^{+}24, QLL^{+}25]. SCASS [dCPR25]. scenario [LLYL21, XCY^{+}25, ZGZL23]. Scenarios [YKG20, CWYW25, JRP24, SCF^{+}25, YKN21, YK22a]. scenes [PB25]. SCFI [SLCS24]. SCFL [WZX23]. Scheduling [ZWL^{+}25, GLC^{+}25, LLYL21, ZXB^{+}24]. Scheduling-driven [ZWL^{+}25]. Scheme [AA24b, KFHM24, LTU^{+}21, QXS^{+}22, XCL^{+}24, ZFS25, AL23, AMNR20, BA20, CB20, CZS^{+}21, CWL^{+}22, CWZ^{+}24a, CWC^{+}25, DCSY25, GR24, GLL^{+}24, GPK21, GYL^{+}24, GXL25, GKC22, HLZ^{+}25, HZX^{+}20, HZZ^{+}24, HYZ^{+}24, ILB20, KLZ24, KJR^{+}20, LSL20, LLX^{+}20, LYW^{+}21, LLD^{+}25, LWL^{+}25, LWZL25, LL21b, LZZX21, LRL^{+}22, LWZ22, LWL^{+}23, LPX20, LZS^{+}20, LHW^{+}20, NKA^{+}25, NZT^{+}25, OO21, PPGC24, QZS^{+}25, RJ21, RRS21, SBY^{+}22, SLL23, TLK^{+}24, TJ20, TW20, XZZ^{+}25, XCC^{+}25, YL20, YGG24, YLY20, ZLY^{+}20, ZXZ^{+}20, ZM20, ZJY^{+}23, ZTY^{+}20b, ZS24]. schemes [LMZZ22, SYD^{+}25, WZZW20, YZW^{+}20]. science [ESA^{+}23, ZLCA21]. SCIRAS [CSEJGUA25]. SCL [WXT^{+}24]. SCL-CVD [WXT^{+}24]. scoping [MSA24, OANA25, vdSPR24]. score [AGM20, KN23, MPRB21, HMV^{+}25, MBB^{+}23]. scores [TO22]. scoring [JAAS^{+}23, KRR23, SAB^{+}25, SSN23]. SCOUT [IM23]. SCP [HKM22]. scream [Gur25]. Script [AAB22]. scripting [CPR21, HXZ^{+}23, PFG^{+}24, WFLX23]. Scripts [ZLLY25]. scrutiny [Bel25]. SD [ESTA^{+}25]. SD-IoT [ESTA^{+}25]. SDN [CZSW24, DTK^{+}21, HNNN^{+}24, KKAS21, LJ22, MNMS24, MK24, NN24, RFdS23, SCOE23, TK25, TD21, WIOO24, WW25b, XCWY25]. SDN-enabled [DTK^{+}21, MK24]. SDN-IoT [MNMS24]. sealed [AN20, LLJ21]. sealed-bid [AN20]. SealFS [SSGM21]. seaports [GKB21]. search [BC23a, GJCJ20, LLLZ21, LLJ^{+}23, LPL25, MRG21, SDA^{+}24a, SZL^{+}24, TDYQ23, YWZW24, ZNL^{+}20]. searching [GJCJ20, hLHhLfW21, LCC^{+}21]. SEC [Pit25, LHL^{+}24b, SEM^{+}23, JJS^{+}24]. sec-certs [JJS^{+}24]. SecChecker [YCWW23]. SecKG2vec [LGG25]. second [VMS20]. second-order [VMS20]. SecOnto [CBS25]. SecRASP [HD25]. secret [Wan21, HMV^{+}25]. SECRET-SCORE [HMV^{+}25]. SecTEP [LLJ21]. section [AGAS25, XGS^{+}21]. sector [AS24, GMPA23, KMAHU24, Lee23, MBKD21, SST21, VBF21, WASA24, Coj25]. sectorial [TDM^{+}21]. Secure [AA24b, BWZ^{+}25, CDMO24, CWC^{+}25, DABB20, ETAB22, FYDH25, GJCJ20, JP25, KKRP21, KM22, LMZZ22, LHH^{+}20, LSXJ22, NN24, OBK^{+}23, RHDF21, XCL^{+}24, YWWH22, ZX23b, AL23, AA20, ARS^{+}25, AFLR25, AK24a, BK22, BF23, BRM21, BQA^{+}23, CZZ^{+}21, DP25, DCSW20, GY23, GRJV24, HABW24, HYZ25, HZX^{+}20, IJR^{+}25, JMU^{+}25, KLZ24, KR23, KKS^{+}22, LT25, LLX^{+}20, LLJ21, zLXZrxZ25, LL21b, LZS^{+}20, LLL^{+}23c, LZ20, MZA^{+}20, MDKR25, MLM^{+}23, NSA^{+}20, OCJ20, OLH^{+}23, PPGC24, RJ21, RKSV20, RPR^{+}21, RXFZ21, RLW^{+}20, RRS21, SBY^{+}22, SCE21, STZ^{+}22, TWW21, UHK^{+}21, VRPTB^{+}23, WCW^{+}23, YLW24a, YPDC20, ZTK^{+}23, ZJY^{+}23, ZXB^{+}24, ZLZC22, ZYGH24]. Secured [ASH^{+}24, LHW^{+}20]. Securely [GLZ^{+}22]. SecureQwen [MFD25]. SecureRC [GY23]. Securing [CG24, CGS22, KS25b, NHC25, SC21, ANG^{+}25, GVJ23]. Security [ACA^{+}23, ATA^{+}25, ABAJ25a, AGVA21, CJKR21, CBS25, CAT24, DJKA25, DJG^{+}25, FFLQ24, GCS22, GMLB23, GHOS22, GZ20, GZG20, HMPS20, HZG20, HZZ^{+}24, IM23, KMG21, Kha21, KLSD24a, KEM25b, KL23, LCA^{+}23, LCH^{+}24a, LWSW25, MBP20, MGL^{+}20, MVSSZ24, MBK^{+}22, MAP23, Ogb21, OMO20a, ÖEÖ22, PVFM^{+}21, PNFBH23, QXS^{+}22, QXS23, SK20, SV25, SM23, SP20, Sha21, Ske24, SvSZ21, SAM25, TMJ^{+}23, TD21, WZX^{+}22, WLM25, WMC20, WCOZ23, ZJY^{+}23, ZMF^{+}23, ZHJ^{+}23, dASJ^{+}23, tBLLV21, AMA22, ADM23, AD23, AS24, ACRB23, AL23, ATMA23, AEA24, AWVG20, ASM20, ALD^{+}21, AUIH25, AAS24, AMVG24, ABA23, AMA21, AKH20, AZM24, AFLR25, ASI^{+}20a, ASI^{+}20b, AASC25, BC23a, BCH24, BCGC25a, BCGC25b, BH25a, BvO23, Bel25, BF20, BBBB^{+}20, BLR23, BTDH20, BBZS24, BSFB20, BDLP24, BPL^{+}20, CCL^{+}24, CB22a, CEN^{+}25, CYL^{+}24, CCVMZ24]. security [CCE23, CBK22, CC20, CMA23, CKK^{+}23, CGBK24, CZE^{+}22, CXC22, CSZ^{+}25, CGH^{+}23, CSRA22, CBFH20, CDGS23, CDN21, Da 22, DMPT25, DPKHP22, DPTAM25, DCS^{+}22, DAA24, DXZZ25, DPK20, DZ22, DHZ24, EÖ25, EA20, FKS25, FVF21, FBL20, FF23, FJR23, FK23, FVV23, GSSC24, GMRB23, GSSP23, GB21, GG22, GMPA23, GSG23a, GAC20, GSS20a, GMSA24, GSG23b, GKB21, GZG21, GW24, HSW25, HAKK21, HN25, HMBY23, HDS21, HuRMMW22, HMMW22, HZM21, HWD^{+}25, HD25, HKM^{+}23, HCCS22, HSB^{+}22, HMV^{+}25, HSSPK21, HG20, HZZ20, HZX^{+}21, HHZ21, HZL^{+}21, HS25b, IO24, IJR^{+}25, JQYL22, JJS^{+}24, JRP24, JK21b, JR24, JZ23, JIG22, JB21, KK25a, KKR^{+}24, KZN^{+}25, KGK23, KKGK25, KS25a, KSS20, KKP22, KDBB25, KLSW24, KLZ24, KFB^{+}25, KGIS21, KAFDW22, yKK22, KKY^{+}23, KA23, KMAHU24, KN20, KJB^{+}25, KMK^{+}24, KF25, KG20]. security [KSVY22, KTH20, KZFM24, KIAV22, LSN^{+}21, LSWR20, LKW^{+}20, LS25, LHL^{+}24b, LLD^{+}25, LDW25, LDS24, LYLT24, LCL^{+}25, LLG20, LZZ^{+}21, LS21, LL23, LGG25, LHYW22, LET^{+}25, MGGR22, MNMS24, MV25, MZW^{+}25, MD22, MGG^{+}25, MHSK21, MKH24, MGZ^{+}20, MWL23, MM23, MBKD21, MM25, Mit20, MSMH21, MEH^{+}24, MISO^{+}23, MRRL23, MBR23, NM24, NJS^{+}24, NTBH^{+}22, NR24, NDDH^{+}21, NGC^{+}25, Nie24, NG23, NM21, OCB21, OG23, OCB23, OANA25, Olu22, OAA^{+}22, OA25, OXNA25, PLS20, PNK20, PSN^{+}22, PNF23, PNF24, PLN23, PD25a, PCR22, PX21, PFR20, PDML25, PC24, QMW20b, QS22, RVJ25, RMG25, RSLM25, RKA25, Ram25, RKKK25, RZ20a, RGG^{+}24, Rep25, RSEK20, RMMP23, RMCT23, RKG20, RKKG25, RRB25, SSK23a, SDA^{+}24a, SBJ^{+}24, SS24a, SCH^{+}20, SU21, SC23, SSK22, SVP21, SSHP21, SPN^{+}25, SSR^{+}24, SS23a, SSL20]. security [SSG^{+}25a, SMC^{+}21, SS24b, SHL^{+}20, SLZ^{+}25, SRFC24, SL20, SRCP22, SVPM21, SSG25b, SHM22b, SS25, SH25, SCTK23, SK24b, SOZ^{+}23b, SLL23, SWL^{+}25, SWW^{+}25, SSTRD22, SSRK20, TRH21, TAES20, TKS21, TGOF22, TMN24, TTRY20, TAF24, TFP21, TCJS22, TDM^{+}21, TN21, TLHY25, UNBF21, US20, UYMM21, VB20, VF22, WZZW20, WLKT23, WBNT20, WAPA25, WK23, Wes20, WCL25, WS24, WS25b, WWW^{+}22, WHLS24, XZG^{+}25, XYZ^{+}25, XW20, XWW^{+}20, XWY21, YKG20, YKN21, YK22a, YGG24, YPWS20, YSM^{+}21, YHZ23, YCWW23, ZKAK23, ZWDE24, ZLY^{+}20, ZXH^{+}25, ZZW^{+}25, ZHL25, ZCL^{+}25, ZGG^{+}25, ZY21, ZCJ^{+}24, ZCZZ24, bOJW20, dVABH20, vdSF21, CBK22, LDW25, dCPR25]. Security-aware [DJKA25, JB21]. security-by-design [IJR^{+}25]. security-compliant [KTH20]. Security-Driven [MGL^{+}20]. security-efficiency [YGG24]. Security-first [HZZ^{+}24]. Security-oriented [OMO20a]. security-related [DAA24, HHZ21]. security-risk [KTH20]. SEDML [GLZ^{+}22]. seed [GLC^{+}25, QPL^{+}23]. Seeing [ASF^{+}23]. Seeking [JDB^{+}23, VBM23]. segment [CG24]. Segmentation [MAK21, PA20, CJL^{+}25, MQH^{+}23, SN24, ZHL^{+}22]. Segmentation-validation [PA20]. Segmented [JKS^{+}25]. SEIFF [XWW^{+}20]. selected [Pit25]. selecting [dNPM^{+}21]. Selection [GBA22, ZFS25, ALAN25, AA22a, ABAM22, AMN22, ACGP25, AK24a, BS24b, CSKD22, FYL^{+}24, FCA25, FK22, GYZ^{+}25a, GYZ^{+}25b, HYW^{+}21, KLSW24, LCZW20, LLL^{+}20a, LLZY20, LP24, MAP23, NK21, PTD20, RVS24, STB25, STB^{+}20, SK24a, SS23b, SSK23b, TZZ^{+}21, TD24, UYMM21, WLQ20, WXJ^{+}25, WLZ^{+}23]. Selective [LLL23d, kJLcL24]. selectors [AK24b]. Self [GSSC24, MBGF20, AMD^{+}24, BSO25, CXL^{+}24, CXC22, CXC^{+}24, FWKB24, FKP^{+}24, FK23, GWL^{+}23, GRJV24, HXZ^{+}23, ISM22, IE22, KMOC20, KIAV22, LLL^{+}22a, LLX^{+}20, NGJ^{+}22, NRS21, PEGA23, RKRS24, RSÁGVV23, WJT^{+}23, WYF^{+}24, XXZ^{+}21, vdSF23, LCC^{+}21]. self-adjustment [PEGA23]. self-adversarial [GWL^{+}23]. self-assessed [KIAV22]. self-assessment [BSO25]. Self-attention [GSSC24, CXL^{+}24, CXC^{+}24, GRJV24, HXZ^{+}23, RKRS24, RSÁGVV23, XXZ^{+}21]. self-attentive [LLL^{+}22a]. self-determination [FKP^{+}24, FK23, KMOC20]. self-disclosing [vdSF23]. Self-feature-squeezing [MBGF20]. self-normalizing [ISM22]. self-organizing [NRS21]. self-protecting [IE22]. self-regulation [CXC22]. self-serviced [LLX^{+}20]. self-sovereign [NGJ^{+}22]. self-supervised [WJT^{+}23, WYF^{+}24]. self-supervision [AMD^{+}24]. self-sustaining [FWKB24]. selfish [JIG22, YCMM20]. SELinux [RKSV20]. selling [LL22b]. SEM [CVZHN25, OXNA25, RJ24]. Semantic [GSK23, JWG^{+}24, OSTO20, ADAT25, BT21, CPS25, CH25, FXL^{+}20, FHSQ20, FHZ^{+}22, GZH^{+}24, GFH^{+}22, HLP22, LLJ^{+}23, LPL25, LLH^{+}25, LGG25, MQH^{+}23, SAA20, SWS^{+}25, YGW^{+}25]. semantic-aware [CH25]. semantic-based [SAA20]. Semantic-enhanced [GSK23]. Semantics [ZTY^{+}20b, ZTW25, SCL^{+}24, YWZ^{+}24, ZHC^{+}25b]. Semantics-aware [ZTY^{+}20b, ZTW25]. Semi [AHS25, KFHM24, NB25b, Rs25, YKY24, ZD21, AOAA20, LHG^{+}24, NGX^{+}23, Vu22, YGB^{+}25]. semi-automated [AOAA20, YGB^{+}25]. semi-fully [Vu22]. Semi-passive [KFHM24]. Semi-supervised [AHS25, NB25b, Rs25, YKY24, ZD21, LHG^{+}24, NGX^{+}23]. Sender [ZL24]. SENSE [LLJ^{+}23]. sensemaking [DSBL23]. sensing [CLM^{+}25, DABB20, XYG25, YLY20]. Sensitive [Kha21, ARAG23, CJD25, FYDH25, GLTH21, GJB22, KAM^{+}21, LWC^{+}25, LHO^{+}25, OSTO20, OX22, WK20, ZMZ^{+}25, ZFW^{+}25]. sensitivity [TLS24b, WLDW24]. Sensor [HZYT23, WXZ^{+}25, GYL^{+}20, LTL^{+}22, LLD^{+}25, LQH22b, LQL^{+}25a, LLL23d, MSG22, MSL^{+}25, PRPS25, RJ21, SKS21, YKSK25, ZUKH23, ZXH^{+}25, ZGS24]. Sensor-based [HZYT23, MSG22]. SensorRE [HP20]. sensors [GAZ25]. sentiment [OUR^{+}23]. separation [Yan23]. September [Ano20w, Ano21y, Ano22y, Ano23x, Ano24x, Ano25x]. sequence [AZ20, AZES21, CWQZ25, DS21, FLLY24, HZL^{+}23, KD21, LLL^{+}22c, NZZ^{+}22, QMW20b, SAR24, STS25, WGS^{+}25b, XLS24, Yan20, YLW^{+}24b, ZLC^{+}23a, ZZWF21]. sequence-based [YLW^{+}24b]. sequence-timing [FLLY24]. sequence-to-sequence [ZZWF21]. sequences [CZLZ24, GZS^{+}22, MMdR24, STPB25, STM^{+}20]. Sequential [ZSJ^{+}25, HGSB24, NFVN^{+}22, WLLL24b, XZZ^{+}24]. Series [TXY^{+}23, ÁFM25, CLL24, CVL22, GPC^{+}22a, GHMG24, LHGB21, LLPL25, LWO^{+}25, TLY^{+}24, WTL^{+}21, WAPA25, XLZ24, YZL^{+}20, YHS24, YWW24, ZZL^{+}22]. Serious [HMPS20, HHS22, YKN21, LDW25, NB25a]. server [GLL^{+}24, LCZY23, LCZ^{+}24, MBU24, SM24, STH^{+}22, WZZW20]. serverless [BSLK^{+}25]. servers [dNPM^{+}21]. Service [HZ22, KHG24, LZW^{+}24, MBS20, AABE20, AWJB21, AABLS24, AOA22, CRC^{+}23, CKL^{+}25, DS22, FEA22, GIO24, IHJZ21, LWL^{+}25, LQY^{+}20, LXJ^{+}25, LWZ22, MAAA20, PSN^{+}22, Rep23, STZ^{+}22, Sun21, TLHY25, WZW^{+}24b]. service-based [PSN^{+}22]. service-oriented [STZ^{+}22]. serviced [LLX^{+}20]. services [ASG^{+}25, AV25, BF23, DSC20, KEK^{+}21, KKY^{+}23, MGL^{+}20, NCPVRT23, OO21, TLS24a, TLS24b, TD21, WQL25, ZNL^{+}20]. Session [CJKR21]. Set [WWH^{+}22, JZL^{+}24, SPG^{+}25, SCC^{+}25, XMK21]. SETA [HHZ21, WZG24]. sets [BL25]. Setting [Ogb21, ABA23, ZZCD20]. settings [AR25, LHAE22, MBGF20, MRL22]. SEV [CDMR25]. seven [MAAG22]. severity [RA25]. SFE [LWL21]. SFE-GACN [LWL21]. SG [WP21]. SG-PAC [WP21]. SGBA [HSX^{+}24]. SGX [CDMR25, GMS20, KBC21, RRM20]. SHA [GCN^{+}24]. SHA-256 [GCN^{+}24]. Shadow [DPTAM25, RZY^{+}24, CSLK24]. shadows [ASG^{+}25, XZZ^{+}24]. Shafer [QMC^{+}22]. Shall [MRSV21, SSL20]. SHAP [GS24]. shape [ZZCD20]. shaped [PTD20]. ShapeShifter [HLYZ21, WLC^{+}20]. shaping [FRF24]. shared [GPK21]. Sharing [LSXJ22, BRWF23, CZZ^{+}21, CWC^{+}25, CFT22, HZX^{+}20, ILB20, LYW^{+}21, LHH^{+}20, MZA^{+}20, SYH^{+}22, YSZ24, ZZP^{+}25, ZS24, ZLZC22]. sHDP [LWS^{+}20]. sHDP-HMM [LWS^{+}20]. Shelf [YCWW23]. shepherd [PMA^{+}23]. Sherpa [KJPP23]. SHFuzz [LZW^{+}24]. shield [XYH^{+}20]. short [AOA22, HLC^{+}24, QSA24]. short-term [AOA22, QSA24]. shortest [LLX25]. shortest-distance [LLX25]. shot [ADDH25, BCO^{+}22, BFR23, CDJD24, CKV22, KT23, LMD^{+}22, LXW^{+}24, PFG^{+}24, WTW21, WLY^{+}24, YLS^{+}22, ZLX^{+}24, ZWX^{+}24, ZJJS^{+}22]. Should [HKM^{+}23]. Shoulder [BB20, BKPZ21]. shoulder-surfing [BKPZ21]. SHRIMPS [CHK24]. shuffle [AHKJJ21]. Shuffled [PMA^{+}23, WZW^{+}25]. shuttles [BCKN22]. Siamese [KZYZ22, WZC^{+}24, ZJJS^{+}22]. side [BEFI25, DCSY25, Gal20b, GSLS21, GYZ^{+}23, GYW^{+}23, HWD^{+}25, JZQ^{+}22, LLLZ21, LJO^{+}20, SMvH^{+}21, TTP23, YZLS25, ZLH^{+}25]. side-channel [BEFI25, DCSY25, Gal20b, GSLS21, GYZ^{+}23, GYW^{+}23, HWD^{+}25, JZQ^{+}22, LLLZ21, SMvH^{+}21, TTP23, YZLS25, ZLH^{+}25]. SIDiLDNG [SQL^{+}24b]. siege [AMVG24]. SIEM [BS20, MLV^{+}21]. sight [GMY22]. Sigmoid [ZWF^{+}24]. sign [GXZ23, MBU24, MCvO24]. sign-based [GXZ23]. sign-on [MBU24, MCvO24]. Signal [ACD^{+}24, KD21, MSZ20, NPL^{+}23, UWL22, WSW^{+}22, WYP^{+}24]. signaling [CHJ22]. signals [WZDZ25]. signature [LRL^{+}22, SRN24, TZZ^{+}23, XKJG25, ZY21]. signs [HLYZ21, KXZ^{+}23]. silence [LKK25]. similarity [FCSP21, JFQ^{+}22, LWZL25, LLH^{+}25, MPRB21, OJL24, Özt24, PC24, SQL^{+}24b, SCL^{+}21, VRS^{+}25, WCZW23, YZP^{+}25, ZSWT24]. similarity-based [Özt24, SQL^{+}24b, VRS^{+}25]. simple [YHH^{+}24]. simplification [XLW24]. Simulation [ATMN23, BJZ^{+}23, CGBK24, CGBK25, KLR^{+}25, KJHL22]. simulation-based [KLR^{+}25]. simulations [EL22]. Sina [RNA22]. sine [NMLN23]. Single [SCBP24, MBU24, MCvO24]. Single-board [SCBP24]. singular [SG21]. sinks [TTRY20]. site [CPR21, HXZ^{+}23, PFG^{+}24, WFLX23]. sites [FF20, MW20]. Situation [Ogb21, AMD^{+}21, OK23]. situational [HSŠ^{+}22, MBR23, TAA^{+}25, HKM22]. sized [KFB^{+}25]. skeletal [KXZ^{+}23]. sketch [CZSW24]. sketches [AFLR25]. skies [SBV24]. skills [Alm25, Fur21, KIAV22, SCVB21]. skimmers [REHS24]. skipping [WYJ^{+}22]. skipping-based [WYJ^{+}22]. SKT [ZWF^{+}24]. SKT-IDS [ZWF^{+}24]. skyline [WZW^{+}24b]. slice [BBBB^{+}20, JB21, ZL24]. sliced [WZY^{+}23]. SlicedLocator [WZY^{+}23]. slicing [ZXS^{+}25]. SLIFER [PTD^{+}25]. Slovenian [VBM23]. Slow [CWW^{+}25, ZKS25]. slow-rate [ZKS25]. SLR [HW20]. small [AAFF24, CCE23, CZZ^{+}21, CZK24, HR24, KFB^{+}25, SZCZ24, TRH21]. small- [KFB^{+}25]. small-to-medium [CZK24]. Smart [AMVG24, BP24, ARR^{+}22, AR25, ASAA21, AWJB21, ADC^{+}20, BEB25, BAM^{+}24, CMLM25, CYL^{+}24, DSC20, GSY^{+}20, HZKN22, HFLZ23, HRU24, Hod21, HZL^{+}21, KTGDE20, LBWV23, LYW^{+}21, LLC^{+}23, LPL25, LS21, LQH22b, LHW^{+}20, MZA^{+}20, MK24, MEH^{+}24, PHH^{+}20, RKH^{+}22, SBY^{+}22, SCSO20, SPZ^{+}20, SC21, SMM22, XX20, XGS^{+}20, YZL^{+}20, ZMLP23, bOJW20]. smarter [YZC25]. Smartphone [GSLS21, HZYT23, AGW25, ACLA22, BTDH20, KS25a, LTZ23, MGZ^{+}20]. smartphone-based [ACLA22]. smartphones [ASA23, MSG22, ZGS24]. smartwatch [GAZ25, LCL23]. SMCD [MZLK25]. SME [vdWLM24]. smell [AES22, BCO^{+}22]. smooth [LJJ22]. smoothing [DMZ24, GLB^{+}22]. SMOTE [XQS25]. SMOTEBoost [LPW25]. SMS [AMD^{+}24, XQS25]. SMT [MVVL^{+}24]. SNDMI [PGP^{+}24]. Sniff [CDJD24]. Snooping [HZX^{+}25]. SOAR [BRO^{+}23, WW25a]. SOC [CZCX20, GVEV22]. SOC2M2 [SVP21]. Social [AGM20, PLB^{+}22, ALZ^{+}20, AZO^{+}23, AMA21, AJHA20, AOAA20, CPT^{+}22, CQL^{+}21, CCL^{+}21, DPKHP22, DS22, EMS23, FKP^{+}24, FF20, HZGM25, HNZ^{+}23, HYR^{+}22, JR24, JB22, LT25, LZL^{+}22, NCPVRT23, QIN25, TG25, VS25, VBM23, WW25a, WTX^{+}22, WP21, YWZW24, ZYL^{+}20, ZGNZ21, vdSFF20a, vdSF23, TG25]. Social-Hunter [TG25]. socio [HR24, MMvS20]. socio-technical [HR24, MMvS20]. Sociodemographic [LK20]. SOCs [OK23]. soft [CCVMZ24, LHD^{+}23, LLX25, SAJP22]. soft-processors [CCVMZ24]. softmax [HMTC22]. Software [FD23, HWW^{+}25, MARB^{+}23, TYZ^{+}24, ZGM21, AUIH25, Alq22, BDC^{+}23, CDMO24, DJKA25, DS23, DGD23, FEA22, FVV23, GIO24, GMSA24, HHL24, HP20, HZZ20, HQL^{+}22, IE22, KKP22, KKY^{+}23, KG20, KBAG23, KJI25, LKW^{+}20, LWZ22, LLL23d, MVA25, MVVL^{+}24, MHSK21, MKH24, MWL23, MEH^{+}24, MTD^{+}24, NDDH^{+}21, RKKG25, SJMP23, SSN23, SBV24, SK24b, SOZ^{+}23b, SKS25, TZZY25, TCJS22, VCD22, WGRZ23, WPZW25, XSL^{+}24, XCWC25, YLPZ21, YPDC20, ZLX^{+}24]. software-Defined [FEA22, DJKA25, DGD23, GIO24, KKY^{+}23, KBAG23, KJI25, LKW^{+}20, LWZ22, LLL23d, MKH24, MTD^{+}24, SJMP23, SBV24]. SOHO [QPL^{+}23]. SoK [DA20, MM23, RBHW25, VWB^{+}25, VPM23]. solicitation [XZG^{+}25]. solid [BWB^{+}21, GMSA24]. SOLID-S [GMSA24]. solution [DG20, KA23, MBU24, WZDZ25, ZMQ21]. solutions [AA25, AZDF20, CZE^{+}22, DL21, KKAS21, KLSW24, KEM25b, NBBS22, SSRM25]. solve [SCTK23]. solvers [DA20]. Solving [BBBB^{+}20]. sophisticated [SSK^{+}24]. sound [SGVA21]. Source [CYL^{+}23, HYR^{+}22, SKL23, ZWW^{+}21, CKL^{+}25, ESM24, GYW^{+}23, KMAHU24, KRR23, KG20, LWD^{+}25, LWW^{+}21, LXW^{+}25b, NGM^{+}24, NDDH^{+}21, NMCRB21, Özt24, SWS^{+}25, SPL25, TYZ^{+}24, TZZY25, WWH^{+}22, WTX^{+}22]. Sources [TRM22, BAM^{+}24, KRR23]. South [PKPC24]. sovereign [NGJ^{+}22]. space [AHuHS24, AFAkJS25, DKV^{+}24, LsJC^{+}22, LWL21, LKLZ24, OS22, SMC^{+}21, ZLH22]. space-filling [OS22]. spaces [CPT^{+}22]. Spain [GPRGMF25]. Spain-Based [GPRGMF25]. Spam [NRS21, AMD^{+}24, DPSS23, SAA20]. Spark [PRTV22]. SparkShield [VC25]. sparse [LLX^{+}23, MV25, MMdR24, TP25, ZCLW23]. SparSFA [WMGH^{+}23]. sparsification [LLL^{+}23c]. Spartan [MBGF20]. Spatial [WWWS24, GLL^{+}24, GHMG24, LQY^{+}20, WW25b, ZLF^{+}20, ZYGH24, YCLZ24]. Spatial-temporal [WWWS24, GHMG24, WW25b, ZLF^{+}20]. Spatio [CLN^{+}25, XLY^{+}20, LQL^{+}25a]. Spatio-Temporal [CLN^{+}25, XLY^{+}20, LQL^{+}25a]. spatiotemporal [NZT^{+}25, ZWB^{+}25]. SPDM [AFAS25]. Speaker [ZLL24, AR25]. Speaker-specific [ZLL24]. spear [BAJ25]. Special [CS24, bOJW20, RZ20a]. specialists [APK24]. specialization [LHGB21]. specialized [BFR23]. Specific [APA^{+}24, CPS25, CRP25, DS21, SAB^{+}25, UWL22, ZLL24]. Specification [CPT^{+}22, HMV^{+}25, LCCZ24]. spectral [LWLL25]. speculative [LLP25]. speech [GSLS21, KKK23, KN23]. speed [CWW^{+}25]. Spherical [ÖEÖ22]. Spider [AR21]. spills [ZXL^{+}24]. split [KEM^{+}25a]. split-learning [KEM^{+}25a]. Spoofing [NLA^{+}23, HKY^{+}25, JK23, KUA^{+}25, MTD^{+}24, SSR^{+}24, WYP^{+}24, XZG^{+}25, XYZ^{+}25]. sports [BHW24]. Spot [QMW^{+}20a, LYL^{+}23]. SPP [YCLZ24]. spread [SYF^{+}21]. spreading [Wan21]. SPrivAD [SBY^{+}22]. spurious [ALZ^{+}20]. spy [LZX^{+}23]. spy-robots [LZX^{+}23]. Spyware [PGP^{+}24, WWL^{+}25a]. SQL [CMCVGH^{+}23, FBR25]. squeeze [AC25b, KM24b, NMLN23]. squeezing [MBGF20]. SRC [STH^{+}22]. SRC-based [STH^{+}22]. SSAE [LJ22]. SSH [LL22a, ZZ20]. Stability [WX23, XLD25]. stabilized [PRPS25]. Stack [JKS^{+}25, CSLK24, LCB^{+}22, ZKL^{+}23]. Stacked [AAB22, dCBJSdS23, MMC24, TP25, VT25, VS25, ZCLW23]. stacked-unsupervised [dCBJSdS23]. Stacking [KS23b, MRL22, OZW^{+}21]. stage [CVZHN25, CDG^{+}23, DZZ^{+}22, FXL^{+}20, GAW^{+}23, KSS20, KBMM23, LKK25, LYS^{+}24, MGZ^{+}25, ST24, ST25a, SZCZ24, WYJ^{+}24, ZZWF21]. staged [JN23, MKC^{+}23]. stakeholder [MR23]. stakeholders [TKKF24]. standard [AKH20, BF23, WK23]. standardised [APK24]. standards [BCKN22]. State [PLS20, QS22, AMC23, BWB^{+}21, FSCM23, KDE20, LLF^{+}23, LDW25, LX21, LWZ^{+}23, LS21, PAA^{+}24, SMC^{+}21, SZW^{+}25, VRPTB^{+}23, ZMZ^{+}25, ZLY^{+}20, ZYH^{+}21]. state-based [LLF^{+}23]. State-of-the-art [QS22, AMC23, LWZ^{+}23, PAA^{+}24]. state-sensitive [ZMZ^{+}25]. statement [ZSJ^{+}25]. statement-level [ZSJ^{+}25]. statements [BDM^{+}24]. States [WCP^{+}23]. Static [DSN^{+}22, LZD^{+}20, QPL^{+}23, BLSS22, GMB^{+}25, GMPMS21, GTP^{+}25, GSYM22, LL23, MCMMMA23, NDDH^{+}21, OX22, SAR24, TDYQ23, WPS20, ZXL^{+}25, ZSLY25]. station [NTBH^{+}22]. Stationary [RDHS23]. Statistical [GBA22, HCLRM20, SMM22, ABAM22, ATMN20, BTM21, WHLS24, XZZ^{+}24]. statistics [zLXZrxZ25, VMS20, ZZJC20]. status [AHSZ21, BSFB20]. STD [DMRV21]. Stealing [RG23]. stealth [WFKH25]. Stealthy [LYL^{+}23, WYF^{+}24, ZC24, ADC^{+}20, APPN23, HSX^{+}24]. Step [GZS^{+}22, GKC22, LLH^{+}25, WYL^{+}24, XCY^{+}25]. Steps [DJG^{+}25, DJG^{+}23]. stewardship [OCB21, OCB23]. STFT [TLY^{+}24]. STFT-TCAN [TLY^{+}24]. STGCN [WW25b]. Still [EZLC21, CZG^{+}22]. STIX [CHA^{+}24b]. STL [AD21]. STL-HDL [AD21]. Stochastic [SPN^{+}25, LZDZ21, WP21]. Stock [ALD^{+}21, MGTK22]. Stop [MBSE22, HLYZ21, ESTA^{+}25]. Storage [ETAB22, HK25, SSGM21, GPK21, HSQ24, LXLL25, LPX20, RYS25, TXH^{+}23, TWW21, TZZ^{+}23, TWQ^{+}25, XHB^{+}21, XCC^{+}25, YSZ24, ZJY^{+}23, ZSL25]. Storage-based [SSGM21]. Stories [BLC^{+}24]. story [HSL23]. STPA [LHL^{+}24b, SEM^{+}23, YGB^{+}25]. STPA-Cyber [YGB^{+}25]. STPA-Sec [LHL^{+}24b, SEM^{+}23]. StratDef [RS23]. Strategic [AA24a, CHJ22, GHOS22, RS23, AZO^{+}23]. Strategies [BPW^{+}20, FKL23, LCL^{+}25, WCOZ23, AAJP25, ATJ^{+}22, ASA^{+}22, ATA^{+}25, GG22, HZ20b, ISR^{+}23, JYL25, NPF^{+}23, OCMF23, QAA23, RCD25, SSK23a, ZCZZ24]. strategy [BCO^{+}22, BCGC25a, BCGC25b, GWF^{+}24, GLC^{+}25, HZM21, HLL^{+}25, HZ22, LKA^{+}21, LLF^{+}23, LSTL24, PLW^{+}21, QLL^{+}25, SCC^{+}25, TZZ^{+}21, WSW^{+}22, WXJ^{+}25, XSK23, ZHC25a, dSFG20]. strategy-driven [GWF^{+}24]. stream [HMS21, LHS21]. streams [DSN^{+}22, LWX^{+}22, PEGA23]. strength [DPSN20, GZ18, GZ20, MFI20, SW22]. strengthened [GCN^{+}24]. Strengthening [AAJP25, VC25, ZHC25a, CZCX20, MNMS24]. stress [AD23, HS25b, TN21]. STRIDE [AAT23, SEM^{+}23]. strike [LCX^{+}25]. string [ZZJC20]. stripped [LCC^{+}21]. strong [FGWJ20, GZG19, GZG20]. stronger [MLM^{+}23]. Structural [LSAH21, JNKZ25, LTZ^{+}24, LGG25]. structure [CWHH23, LWSW24, LWSW25, MYH25, RHDF21, SCE21, SRT25]. structure-based [SCE21]. structures [HR24]. studies [GYW^{+}23, NTH25]. Study [HZM21, KBD^{+}23, LKS^{+}23, SYF^{+}21, ZYH^{+}21, AR25, AMD^{+}21, ARB20, AA20, Alq22, BCH24, BMM22, BFC^{+}25, BLR23, BKV^{+}23, BQA^{+}23, CG24, CPRV21, CZK24, CKG22, CCH^{+}22, DBB^{+}24, FFPC22, GPC^{+}22b, GB21, GKB21, HNH^{+}22, HRU24, HP23, KS25a, KYKI22, KJPP23, KKHK24, KJ23, LHL^{+}24b, LYD20, MW20, MBPCC21, Mit20, MT21b, NNA^{+}23, NSLM25, NDDH^{+}21, QYYH24, RDM20, RKW^{+}23, RGC24, RZ24, Sch24, SBL20, TKS21, TV20, TN21, VWJ24, WW25a, WFT22, XSL^{+}24, YK22b, ZUKH23, ZZW^{+}25, vdSF23]. Studying [MPL^{+}25, ESA^{+}23]. stung [MGGR22]. style [JCHS24, SAM25, ZYH^{+}21]. styles [RCD25, XLL^{+}25]. Sub [STM^{+}20]. Sub-curve [STM^{+}20]. subgraph [OX22]. subgraph-based [OX22]. Sublinear [LPL25]. subsequences [AGW25]. subset [AA22a]. subspace [GHZ^{+}24, LFW^{+}25, ZC24]. substitute [YS22, ZLH22]. substring [YLZ^{+}21]. substring-of-keyword [YLZ^{+}21]. subsystem [HLZ^{+}25]. success [AWVG20, MZWM23, YWPC22, YLSJ23, ZSS21]. SUETA [ZLL24]. suite [GYW^{+}23]. SuM [CSLK24]. summary [BDM^{+}24]. SuperLearner [AIBKJ24]. Supervised [RKW^{+}22, WXT^{+}24, AHS25, KT23, LHG^{+}24, MLXC24, NB25b, NGX^{+}23, Rs25, WJT^{+}23, WYF^{+}24, YKY24, ZD21, ZCPB23]. supervision [AMD^{+}24, XWY21, YLL23]. supervisor [DAA24]. Supervisory [US20]. Superword [GZG21]. supplier [LSLL24, RVJ25]. suppliers [TDM^{+}21]. Supply [HHS22, KKGK25, KEM25b, SSTRD22, TDM^{+}21]. support [HP20, HSŠ^{+}22, KKP22, KFB^{+}25, MSCJ21, MR25, MSH^{+}25, SP20, SHM22b, WSW24, ZM25, vdKSCY22, RDPR25]. supported [WTHL25]. supporting [CHP^{+}24, CFYBF23, HSQH23, HSQ24, XCC^{+}25, ZSL25, dNPM^{+}21]. suppression [SHK^{+}21]. Surf [HZX^{+}25]. Surf-Snooping [HZX^{+}25]. surface [AGBB22, RGP23, ZHL25]. surfaces [NGJ^{+}22]. surfing [BKPZ21, BB20]. surrounding [NTH25]. surveillance [DCS^{+}22, SJ21, vdSFF20a]. Survey [AOM24, ACA^{+}23, BRV^{+}22, BRMH22, BAAM23, GCS22, HZKN22, KEK^{+}21, KLSD24a, KLSD24b, KEM25b, LCP^{+}23, LWD^{+}25, PNFBH23, ZWH^{+}25, ASF^{+}23, AN20, AZDF20, AZM24, FAAF23, BTE23, BZE^{+}20, BBA20, BAJ25, BTDH20, CZE^{+}22, CCW^{+}24, CL20, DCB^{+}21, DKV^{+}24, EMS23, FSCM23, GWF^{+}24, GC25, GBG20, HSW25, HAKK21, HY21, IB20, JJT^{+}24, KKAS21, KFS^{+}22, LSWR20, LCY^{+}21, LCA^{+}23, LWZ^{+}23, LPZF24, LGXZ22, MJ25, MAL22, MBP20, MM25, MRRL23, NDPC21, OAA^{+}22, PSN^{+}22, PX21, PDWF25, QGW^{+}25, QS22, RDHS23, SS24a, SC23, SSC23, SHL^{+}20, SCCZ20, SWW^{+}25, TKKF24, TFJ23, TSW^{+}21, WLM25, YFO^{+}21, ZYL^{+}25, ZHL25, ZCL^{+}25, ZJH^{+}24, ZLZX24, dSFG20]. survey-driven [TKKF24]. Survivable [FMAC21]. Surviving [GVEV22]. Susceptibility [FF20, MVH^{+}21, RGC24, TJD23, VWB^{+}25]. sustainable [BSR^{+}25, EÖ25, MEH^{+}24]. sustaining [FWKB24]. SVD [WZX23]. SVDD [LWLT21, LWT24]. SVM [AK24b, GL21, GYL^{+}24, LHL^{+}24a, LJ22]. SVulDetector [ZSWT24]. Swarm [KE25, SY24, AR21, CSKD22, PGRS24, PRPS25, XYH^{+}20]. Swedish [VBF21]. SwiftIDS [JLQ^{+}20]. swipe [ASA23, CHH25]. Switches [AFK^{+}24, XYZ^{+}25]. switching [XCWY25]. SWMAT [VAAG23]. sybil [KJCL20, PC22, YHY23b]. symbolic [LXX^{+}25, MV23, SBB^{+}20, VCP25]. symmetric [SK20]. SYN [LJO^{+}20, NH21]. synchronisation [HZM21, He25]. synchronization [TH21]. SynDroid [LHL^{+}24a]. Syntax [RKG^{+}25, XLW24, XWJ^{+}24]. syntax-aware [XWJ^{+}24]. synthesis [BMAP24, YWPC22, ZQMC20]. synthesized [NRKK23]. Synthetic [KS23a, LGH21, SSK23b, WTL^{+}24, YZP^{+}25]. System [APA^{+}24, CZZ^{+}23, KBD^{+}23, LSWR20, OBK^{+}23, TXY^{+}23, ZTJ^{+}21, SHM22a, ANAAK25, ALAN25, AHSZ21, ABAM22, ADJS21, AD21, AlD25, AK20, AKCS20, ACKS22, ABAJ25a, ABAJ25b, AAAK23, ARS^{+}25, AV25, AXAL24, AASC25, AOA22, BKSA25, BBD23, CYL^{+}24, CDG22, CGS22, CQC^{+}24, CZH^{+}25, CBFH20, DP25, DH25, DNB^{+}20, DZ22, EAN^{+}22, FWKB24, FFB^{+}23, GY23, GYL^{+}20, GCN^{+}24, GZG21, GFH^{+}22, HFK^{+}24, HKA24, HG21, HGSB24, HZX^{+}21, HHZ21, HXZ^{+}21, HLX^{+}24, HZZ^{+}24, HWML25, IE22, JK21a, JIG22, JLQ^{+}20, JCJ^{+}21, JNKZ25, KS20, KDBB25, KAFDW22, KJPP23, KJJ^{+}24, KKHK24, KRR23, KBAG23, KS23b, KL23, LCC^{+}22, LS25, LFC^{+}25, LYW^{+}21, LLYL21, LCY^{+}21, LZCS22, LWYG24, LHL^{+}24b, LQHZ25, LDW25, LGH21, LZDZ21, LZG^{+}22, LQL^{+}25a, LCSD20, LHW^{+}20, MT24, MGGR22, MR25, MV25, MYMC20, MSZ20, OPK20, PAM^{+}24, PAGDG24, QMC^{+}22, QWLL23, RKKK25, RSEK20, RGP23, SEM^{+}23, SUS25, SCE21]. system [SKD23, SAB^{+}25, SSN23, SGJ^{+}25, SC21, SQL^{+}24a, SQL^{+}24b, SBL20, SG21, SPL25, TAF24, TFMP25, TD24, VLV25, VS24, WFK25, WJT^{+}22, WLL^{+}23, XHWL20, XKY25, YWK^{+}22, YZLS25, YC25, YHH^{+}24, ZLF^{+}20, ZZW^{+}21, ZTK^{+}23, ZLDX25, ZWX^{+}20, ZHC^{+}25b, ZLL24, ZZ20, DBG24, LCZW20]. System-level [CZZ^{+}23, LCY^{+}21]. Systematic [AWVG20, BvO23, JTF^{+}24, MVA25, MARB^{+}23, SM23, SH25, ACRB23, AA21b, ALD^{+}21, AKA25, BKPZ21, BB20, BSR^{+}25, CMLM25, CTV21, CLHT24, DA20, EL22, ESBJ20, GG22, GGK25, HY21, HKM22, HW20, KGK23, KBK24, KGIS21, MMSJ25, NM24, NPF^{+}23, NB25a, OK23, PNK20, PNF23, PvSvdB24, QAA23, QS22, RMG25, SGSCM23, TNN^{+}22, TV20, VB24, YLL^{+}22, YWPC22, ZHL25]. Systematical [GPC^{+}22b]. Systems [AA24b, AA23, CEN^{+}25, CLN^{+}25, CCO^{+}20, DAZ20, GCS22, KBD^{+}23, KFHM24, KDBS22, LTU^{+}21, LKS^{+}23, PT25, SJH^{+}21, dCPR25, AOM24, AHFL25, AVR20, AJ20, AEA24, AAS24, AGVA21, FAAF23, AK24b, APPN23, BEFI25, BHO23, BTE23, BFC^{+}25, BEAE25, BZE^{+}20, BDLP24, CS23, CRS^{+}22, CPV23, CKS^{+}24, CHK24, CKL^{+}25, CFYBF23, DCK^{+}23, DLC24a, DANS20, DBB^{+}24, ELD24, EMS23, FYL^{+}24, FFLQ23, FFLQ24, GLM^{+}24, GTP^{+}25, GSY^{+}20, GDK^{+}21, GJB22, HMK^{+}20, HYW^{+}21, HHE22, HRU24, HSB^{+}22, HZ20b, HZG^{+}23, ILB20, ISR^{+}23, IFHP22, JS25, JZY^{+}25, KBK24, KAFDW22, Kil25, KKK23, KMH^{+}22, KS23b, KL22, KN23, LLJ21, LQH^{+}22a, LXLL25, LLLB22, LZZ^{+}21, LXY^{+}24, LJL^{+}25, MT24, MYS^{+}25, MAR22, MFA20, MBU24, MSSO^{+}25, MKM^{+}25, MC20, MLV^{+}21, MF25, MDB24, NKA^{+}25, NTBH^{+}22, NJ22, OFIdD^{+}20, PCR22, PVFM^{+}21, QAA23, QMW^{+}20a, QJDD24]. systems [RBM^{+}25, RPR^{+}21, RKA25, RCA23, SZL^{+}24, SS24b, SHL^{+}20, SKM21, SVPM21, SBV24, SSG21, Spr23, SKG21, SGVA21, SJ21, SHH^{+}24, TAES20, TA25, TWLK22, TTP23, TLY^{+}24, TH24, US20, VK22, WFKH25, WSC^{+}20, WS25a, XZWD25, YHS24, YWW24, YP21, ZWB^{+}25, ZLZX24, bOJW20, PAGDG24, SRT25, ZKS25].

Table [CZG^{+}22, WCH^{+}25]. tabular [YL23]. tactical [KEK^{+}22]. tactics [CGBK25, LHC24, SDA^{+}24b]. Tactics5G [SDA^{+}24b]. TAEfficientNet [XLS24]. tag [JYZ^{+}23, XHB^{+}21]. tagging [WLL24]. tags [TJ20]. tailing [SST21]. tailored [KZFM24]. tailoring [RKKG25]. Taint [WCP^{+}23, JYZ^{+}23, TDYQ23, XLH^{+}25, YFdCdC^{+}24, ZTD21]. Taint-Guided [WCP^{+}23]. take [BvtH^{+}23]. Taking [BMV24, LHH23]. talk [BCdGG20]. Talking [DGH21]. Tamp [AKAFQ22]. Tamp-X [AKAFQ22]. tamper [SSGM21]. tamper-evident [SSGM21]. tampered [AKAFQ22]. tampering [CZM^{+}25]. tap [CZE^{+}22]. Target [WCZ^{+}24, HZX^{+}21, HXZ^{+}21, MKM^{+}25, RGG^{+}24, RFdS23, RKKG25, TZZ^{+}21, TV20, ZCJ^{+}20]. targeted [AM25, DSD^{+}25, HvCSJ23, KKK23, WLLL24b, JCHS24]. targeting [KL22, KL23]. Targets [AYL25]. task [ÇÜD21, HSH24, LLL^{+}22b, TN21, YWZ20, LXW^{+}24]. Task-adaptive [LXW^{+}24]. tasks [BCP22]. taught [SCVB21]. taxonomic [AAFF24, OLH^{+}23]. taxonomies [SGSCM23]. Taxonomy [ACA^{+}23, BBR20, DANS20, KKAS21, LGXZ22, VHMGRR22, ZYL^{+}25, ZWH^{+}25, DVAAM23, GB21, KEM25b, NTH25, Onw20, PS21, WLM25]. TCAN [TLY^{+}24]. TCN [SS22, LWT24, TLY^{+}24]. TCN-attention [TLY^{+}24]. TCN-Transformer-SVDD [LWT24]. TCP [KD21, RMMP23, WCZ^{+}24, XYZ^{+}25]. TDSF [HLL^{+}25]. TDX [CDMR25]. teacher [CCC^{+}24]. Team [LHGB21, SA22, WTHL25, WLZ^{+}24]. team-based [WTHL25]. teaming [YGG24]. teams [RVJ25]. Technical [FF23, BPL^{+}20, CPP24, GSG23b, HR24, KKGK25, MMvS20]. technique [ALZ^{+}20, CTC^{+}25, CZX^{+}21, CZL^{+}25, CBPW25, GIO24, GAS^{+}23, GVJ23, HABW24, HYR^{+}22, ILVR^{+}24, LLJL23, LYS^{+}24, LLLZ20, Özt24, PT25, ZKAK23]. Techniques [PNFBH23, VHMGRR22, ABAM22, ATMN23, AHKJJ21, AUIH25, BRMH22, BAM^{+}24, CZE^{+}22, CQC25, CL20, DVAAM23, DKV^{+}24, GSSP23, HKM22, HWWB24, JJT^{+}24, KK25a, KEK^{+}21, KL24, LHC24, LPZF24, MVA25, MJ25, MPM23, MAL22, MAF^{+}23, MW20, NDPC21, PX21, PDWF25, RBHW25, RHH^{+}22, SPV20, STSV25, SKKG25, SSK^{+}24, TJD23, VWB^{+}25, WIOO24, ZMF^{+}25, ZT21, ZWH^{+}25]. technological [AABLS24, CMA23]. Technologies [BRV^{+}22, SWW^{+}25, AABLS24, MCJPS22]. technology [AGBB22, BPL^{+}20, CDMR25, KK20, RDQL25, XZG^{+}25]. TEE [AC25a, CDMR25, WZB^{+}23]. TEE-MR [AC25a]. TEEMRDA [TWQ^{+}25]. TEEnder [GMS20]. Telegram [MV23]. telehealth [GSG23b]. Telemetry [PHT^{+}25]. Telescope [GPRGMF25]. Tell [HSL23]. template [BPCZ23, HABW24, HMV^{+}25, KLZ24, LWZL25, MDKR25, SA20, SK20, SYD^{+}25, TTP20, ZXS^{+}25]. template-driven [HMV^{+}25]. templates [GR24]. Temporal [CLN^{+}25, SKM21, CXL^{+}24, GHMG24, HFLZ23, HMS21, LWS^{+}20, LYHC24, LTZ^{+}24, LCL^{+}20, LQL^{+}25a, LZA^{+}23, MMC24, QYW^{+}25, SS22, SN24, WWWS24, WW25b, XLY^{+}20, YKY24, ZLF^{+}20, ZSGB^{+}22]. temporospatial [TZZ^{+}21]. ten [CCH^{+}22]. tendencies [ASI^{+}20a, ASI^{+}20b]. tender [LLJ21]. tensions [DSBL23, SSK22]. tensor [FWF^{+}23, SW22]. TensorFlow [FD23]. TensorFlow-based [FD23]. term [AOA22, GMB22, QSA24]. terminal [LPX20]. ternary [DCSW20, JK23]. terrain [NTH25]. terrorism [PS21]. terrorists [AAAAS20]. tertiary [KYKI22]. test [KJB^{+}25, LM24, LXW^{+}25a, WLLC24, vdKWH20]. testbed [WBNT20]. testbeds [KGK23, YKG20]. TestCloudIDS [VC25]. Testing [AFAS25, BRO^{+}23, XSL^{+}24, CDMO24, CHZ^{+}23, CWX^{+}24, HSOL23, LSMK25, LZS^{+}23, LZY^{+}25b, LLP25, MEH^{+}24, SA25, SK24b, WLZ^{+}24, WRG^{+}21]. tests [BTM21, FYDH25]. Text [ALZ^{+}20, FZZ^{+}25, JLS22, KL22, KL23, LZHL20, SDA25, ZXXG24]. texts [HLC^{+}24]. Textual [JLC25, SYA^{+}21, OMO20a]. texture [VMS20]. TFAN [LXW^{+}24]. tfidfvectorizer [SG21]. TGPrint [WML^{+}23]. theft [LZX^{+}23, ZXH^{+}25]. their [Ade21, IHJZ21, KEK^{+}21, KJ23, LBWV23, SGSCM23, Sar21, SSC23, TLT22]. them [BTAK21]. theoretic [GSS20a, MYS^{+}25, MAR22, MTD^{+}24, RKW^{+}23, TCY^{+}20, WLLS22, ZJL^{+}20]. Theoretical [CAT20a, CAT20b, CAT24, Coj25, LYD20, SL20]. Theories [MVVL^{+}24]. Theory [ATMN23, GOW24, BvtH^{+}23, DPKHP22, FKP^{+}24, FK23, HHZ21, JIG22, JDB^{+}23, JSYK25, KS25a, KSS20, KKM^{+}25, KTH20, LL22b, LL25, MD22, QMC^{+}22, XW20]. there [MM23, PNF24]. Things [CGCY21, NHC25, SWK20, GLX24, STZ^{+}22, AHZL25, AKT21, AZDF20, BBR20, CRP22, CGBK24, CZH^{+}25, FGH^{+}21, GXL25, HLFR20, HYR^{+}22, JSK^{+}23, JCJ^{+}21, KK25b, KEM^{+}25a, KFS^{+}22, KJI25, LXX^{+}25, LL21b, LPX20, MKM^{+}25, MM25, NAK23, OANA25, OAA^{+}22, PMF^{+}20, RNA22, SASW23, SM23, SAB^{+}25, Sin25, SvSZ21, SCC^{+}25, TFJ23, UHK^{+}21, YHY23a]. think [PNF24]. third [JYF^{+}25, LB25]. third-party [JYF^{+}25, LB25]. Thread [ZFW^{+}25]. Thread-sensitive [ZFW^{+}25]. Threat [ABAJ25a, ABAJ25b, APA^{+}24, CLN^{+}25, EYYZ20, HZH^{+}24, KBD^{+}23, KBK24, yKK22, KJI25, LCY^{+}21, MDK^{+}23, RCA23, TRM22, VHMGRR22, WCOZ23, AJPK24, AAT23, ATMA23, ASSW23, BFR23, BC23b, BAM^{+}24, CTV21, CRC^{+}23, CZZ^{+}23, CZLZ24, CCW^{+}24, CPS25, CZPX22, DYW24, FZZ^{+}25, FBL20, GWL^{+}23, GSY^{+}20, GLH^{+}23, HRU24, Het21, HS25a, JYWS25, JLS22, KBE^{+}24, KLJE24, KFP^{+}24, LHC24, LLH^{+}25, MJZ^{+}25, MSH^{+}25, MS25, NGC^{+}25, PDWF25, PJ23, QAA23, QYW^{+}25, QYYH24, RSSD23, RBHW25, RGPAF20, RK24, DAAM24, SU25, SGJ^{+}25, SVWJ25, TMN24, TCYL21, TSCM21, VBF21, VC25, VM21, WLKT23, WLH^{+}24, WRG^{+}21, WS25a, XLWC24, XZZ^{+}24, YWL25, YW21, ZGZL23, ZYL^{+}20, ZSS21, ZZ20, dPR23, vdKSCY22]. Threat-Hunting [KJI25]. threat-occurrence [FBL20]. threats [BMH21, CGBK25, GAC20, GSG23b, HSW25, HWD^{+}25, ISR^{+}23, JRP24, KDBB25, MSG22, MV25, MZW^{+}25, SH25, SK24b, XGL20, ZGNZ21, dPR23, AASWC25, HZ20b, KWF25]. Three [CCH^{+}23, ZT21, DGS^{+}23, VB24]. three-channel [DGS^{+}23]. Three-classification [CCH^{+}23]. threshold [KE25, ZS24]. threshold-based [KE25]. throughput [XGX^{+}24]. Thwarting [TTP23]. tier [UWL22]. tier-two [UWL22]. TIKG [HZH^{+}24]. Time [CAT20a, CAT20b, GMBN21, LWO^{+}25, TXY^{+}23, UÇ22, WAPA25, YHS24, AYL25, ÁFM25, AXAL24, BTM21, BTH22, CLL24, CVL22, FSN22, GPC^{+}22a, GR24, GHMG24, GMB22, HHSL20, HCL^{+}23, HZM21, He25, HSL20, HWML25, JLQ^{+}20, KMK^{+}24, LL22a, LsJC^{+}22, LLPL25, MVA25, MFA20, MM23, MT21c, NZZ^{+}22, Rou22, SU25, SBV24, TA25, TH21, TLY^{+}24, VDN24, WHPL21, WTL^{+}21, WWW^{+}22, XLZ24, YZL^{+}20, YWW24, YGC25, ZTY^{+}20a, ZZL^{+}22, CAT24]. time-aware [HCL^{+}23]. Time-based [GMBN21]. time-efficiency [KMK^{+}24]. time-frequency [TLY^{+}24]. time-interval-based [FSN22]. time-Lock [MT21c]. time-satisfaction [KMK^{+}24]. time-series [CVL22, LLPL25]. timed [TA25]. timeline [LSN^{+}21]. times [OO21]. TIMFuser [MJZ^{+}25]. TIMiner [ZYL^{+}20]. Timing [CDN21, QXS23, FLLY24, HHSL20, JYS21, LCTZ25, MYW24]. timing-based [MYW24]. TLS [GZS^{+}22, TO22]. TLS-Based [GZS^{+}22]. TOCTOU [BWZ^{+}25, KR23]. TOCTOU-secure [KR23]. today [VWB^{+}25]. together [LBD^{+}23]. Tokenizer [ZZHY25]. Tokenless [LTU^{+}21]. tolerant [FSCM23, RLW^{+}20, SCOE23, YGG24]. tomography [LXZ^{+}22]. tomography-based [LXZ^{+}22]. TOMSAC [SBJ^{+}24]. too [ARB20, ST24]. Tool [SCF^{+}25, ACG20, BKS22, JCJ^{+}21, KLJE24, MSD^{+}23, MHC^{+}25, SPL25, YKN21]. tool-based [JCJ^{+}21]. toolkit [RMMP23, ZFWL22]. Tools [FACHC21, BRO^{+}23, CBK22, CFYBF23, GMPMS21, KLC23, PCR20, RKG20, SCF^{+}25, YKG20]. toolset [HSŠ^{+}22]. Top [LLL^{+}23c, LLL^{+}22d, SS23a, SYG^{+}21, XGS25]. Top- [LLL^{+}23c, SYG^{+}21, XGS25]. Topic [CMA23, GWL^{+}23, OUR^{+}23, SH25, WZZ^{+}22]. topological [MDB24, ZLH23]. Topology [WLLS22, SHM22a, LYHC24, LZZX21, LXZ^{+}22, LZX^{+}23]. Topology-theoretic [WLLS22]. tor [GMH^{+}23, AESF23, AABE20, AN21, LXW^{+}24, MAP23, XZY^{+}24]. touch [AGBT23, SCTK23]. touch-based [AGBT23]. touchscreen [WHPL21]. tower [ZGS24]. Toy [SCF^{+}25]. TPL [ZMF^{+}25]. TPM [LHW^{+}20]. Traceability [SSTRD22, AHS25, HQL^{+}22]. traceable [SJX^{+}20, ZS24]. Traces [QYJ22, OPK20]. Tracing [PdARD24, ZYH^{+}22, HP23, QZS^{+}25]. Tracking [LXH^{+}23, ATMN20, BT24, MW20, MCJPS22, SYC^{+}24, ZHC^{+}25b]. trade [GTL^{+}22, KDBB25, MR23, SBJ^{+}24, SYL^{+}20]. trade-off [SBJ^{+}24, SYL^{+}20]. trade-offs [GTL^{+}22, MR23]. tradeoff [CYHW25]. Traffic [GPRGMF25, LBW^{+}25, ZXL^{+}24, ZHJ^{+}23, AHS25, AIBKJ24, AESF23, CXL^{+}24, CTC^{+}25, CPRV21, CXC^{+}24, CYC^{+}24, CHWH25, CXL^{+}25, DSD^{+}25, DBG24, DHL^{+}23, DMD^{+}20, GCV22, JQYL22, KSSL22, LLL^{+}22a, LLLZ21, LWX^{+}22, LXS^{+}24, LCCZ24, LBD^{+}23, LHG^{+}24, LXW^{+}25b, LWZ22, LLPL25, LWT24, MSCJ21, MZW^{+}23, MLXC24, MGZ^{+}25, MHC^{+}25, MYH25, MRL22, NOVJ23, NZZ^{+}22, OFH^{+}25, PGP^{+}24, PGC^{+}24, QZS^{+}25, QMZ^{+}25, RNSS23, RJ24, SASW23, SYZ^{+}23, SLR25, STB^{+}20, SLX^{+}23, WFT22, WT23, WML^{+}23, WWWS24, XKJG25, YCLZ24, YLY^{+}23, YXL^{+}22, ZHH^{+}25, ZHJ^{+}21, ZCD^{+}22, ZSGB^{+}22]. Traffics [XLY^{+}20]. train [LM24]. trained [DÇÜD22, PGRS24, WYF^{+}24, YMZ^{+}23]. Training [BVB^{+}23, BLDB24, KMG21, AMA21, CWZL23, CKG22, FVV23, GSG23a, HHT23, HSL23, KRN24, KIMJ23, LHGB21, LCC^{+}22, LZRR23, LZNZ24, LDW25, LHG^{+}24, MSA24, NHFB25, PvSvdB24, PvSvdB25, RKH^{+}22, RCD21, RCD23, SEGD^{+}22, SARG23, XMK21, XLL23, XFY21, YS22, ZSLY25]. Trajectory [HLL^{+}25, CH25, GGZ25, HHP^{+}24, LC21, zLXZrxZ25, NZT^{+}25, WK20, ZW24]. Trajectory-preserving [HLL^{+}25]. trans [MV25, ZWZX25]. trans-Net [MV25]. Transaction [LXH^{+}23, KAK21, WTL^{+}21]. transactions [AABE20, KK20, RPR^{+}21]. transductive [ZLY^{+}21]. transfer [AGAP24, HSQ24, KL22, KL23, ZWW^{+}21, ZLL24]. transferability [ALAN25, DTT^{+}22, HWWB24, LCL24, RZ24, XY24, ZFG^{+}24]. Transferable [DDW^{+}23, LJJ22, WZZ^{+}22, HLQD23, LWH^{+}24, Sar21, WLLL24b]. Transferring [CCC^{+}24]. transfers [GRdAS23, PdARD24]. transform [AESF23, FEA22, LHS21, ZNF^{+}23]. Transformation [GWF^{+}24, AFAkJS25, CBB25, LX25, LCL24, MLC^{+}23, MHSK21, MKH24, SA20, Sha21, TTRY20, WASA24, ZXZJ25, ZWF^{+}24]. transformed [HDZ^{+}22]. Transformer [ANAAK25, HZYT23, LWT24, WYJ^{+}24, ZXXG24, DÇÜD22, HCL^{+}23, JXYF25, MG24b, SCBP24, WCZW23, WAPA25, WWL25b, ZZHY25, ZGS24, GAJ25, LFCD21, WCZ^{+}24]. Transformer-based [ANAAK25, WYJ^{+}24, ZXXG24]. transformer-driven [WWL25b]. Transformers [GSG23b, GZ24, KK25b, DAAM24]. transient [DCSY25]. transition [BCGC25b, LDW25, BCGC25a]. transitions [HR24]. transmission [GRJV24, HKY^{+}25, WPS20]. transnational [CYL^{+}21]. Transparency [PLB^{+}22, BBD^{+}24, DPTAM25, RW24, SBL20, VRS^{+}25]. transparent [CDMR25, SGJ^{+}25]. transport [SLR25]. transportation [CVL22, IFHP22]. traps [HCLRM20]. treatment [RDM20, SGSCM23, SRM22]. tree [AQAK21, CCL^{+}22, LGL21, MWR20, NGJ^{+}22, PBG22, SUS25, SCE21, VK22, XLW24, ZW24, ZSWT24, ZYPT22, LT25]. tree-based [CCL^{+}22, VK22, ZSWT24]. trees [BGL^{+}20, KLSD24a, KLSD24b, LCTZ25, SPN^{+}25, RKG^{+}25]. Trends [CGBK25, CZZ^{+}23, GSSP23, GSG23b, HSB^{+}22, SDA25, ZLZX24]. tri [ZWW^{+}21, AYL25]. Tri-modal [AYL25]. tri-transfer [ZWW^{+}21]. triage [MWM^{+}25]. TriCh [LWSW25, LWSW24]. TriCh-LKRepNet [LWSW25, LWSW24]. trick [WJC23]. TRIDENT [AYL25]. trigger [LCX^{+}25, LLP25, ZDZD24, ZTYA25, CZE^{+}22]. Trigger-Action [CZE^{+}22]. trigger-based [ZTYA25]. triggering [WLDW24]. triggers [SZY^{+}22, WYF^{+}24]. trilateral [ZCJ^{+}20]. tripartite [Sun21]. triple [LWSW24, LWSW25, WLL24]. triple-channel [LWSW24, LWSW25]. triplet [SCL^{+}24, ZXXG24]. trojan [PGC^{+}24, CWHH23, CBPW25, LZ25, PLW25, WWL^{+}25a, XLY^{+}20]. TrojanProbe [LZ25]. Trojans [CJS^{+}21]. true [LZ20]. Trust [SSS^{+}21, SGJ^{+}25, AMA22, AZM24, BOS^{+}21, BPL^{+}20, CB20, DZSL25, FMAC21, IO24, JS25, JR24, LLD^{+}25, LZ20, MLM20, MYW24, MS23, OCJ20, PLE^{+}23, SS24b, Sun21, WFLX23, YLSJ23, ZZW^{+}21, vdSF23]. TrustBuilder [CWL^{+}22]. TrustCNAV [WBZ^{+}25]. Trusted [ETAB22, AC25a, BS25, DL21, HLZ^{+}25, HYW^{+}20, HZZ^{+}24, LLL^{+}25, MRRL23, TWQ^{+}25, ZLY23]. trustworthy [BQA^{+}23, JLLG24, NM21, YQL24]. trustzone [JK21b]. Truth [JANR23, DLL24, ZXZ^{+}20]. TS [LTZ^{+}24]. TS-Mal [LTZ^{+}24]. TSCHA [KSC23]. TTSAD [LWT24]. tuned [AK24b, CZL^{+}25]. tuning [BFR23, XCWC25]. tuning-based [XCWC25]. tunnel [LWZC23, LZ25, MHC^{+}25]. tunnelling [TLS23]. tunnels [MYH25]. Tunter [WCP^{+}23]. turbulence [DHZ24]. turn [GKC22]. turns [KFP^{+}24]. twin [LJO^{+}20, WSW24, LPS23, PPGC24]. twins [SSN22]. Twitter [AAAAS20, BKA20, PLN23, RRMSM^{+}20, SAJP22, SKS25]. Two [DKSS20, EL22, GZS^{+}22, KMK^{+}24, LLL^{+}23a, LYS^{+}24, CDG^{+}23, FXL^{+}20, HXZ^{+}23, JHD^{+}20, KK25a, KBMM23, LSL20, SZCZ24, UWL22, YK24, ZGS24]. two-channel [HXZ^{+}23]. Two-factor [KMK^{+}24, KK25a]. two-hidden-layer [JHD^{+}20]. Two-Phase [DKSS20, LLL^{+}23a]. Two-stage [LYS^{+}24, CDG^{+}23, FXL^{+}20, KBMM23, SZCZ24]. Two-Step [GZS^{+}22]. two-tower [ZGS24]. Type [RRS21, AABLS24, ÇÜD21, CZG^{+}22, HZL^{+}21, KLP^{+}24, LZW^{+}24, Özt24, ZZL^{+}22]. types [CYC^{+}24, ZLZ^{+}24]. typical [CWYW25]. typology [BTA24]. TZMon [JK21b].

UAF [XLG^{+}23]. UAF-GUARD [XLG^{+}23]. UASN [LZX^{+}23]. UAV [SGC23, YKK24, LLD^{+}25, NLA^{+}23, YKSK25]. UAVs [MF25, TAF24, WXZ^{+}25]. Ubiquitous [KFHM24, OO21]. UCFL [DKSS20]. UCRF [QPL^{+}23]. UDP [JGS^{+}24]. UK [FAAF23, CCE23, Coj25]. Ultra [YLDW23]. Ultra-Dense [YLDW23]. ultrasonic [ZXH^{+}25]. UMUDGA [ZPP20]. Unaddressed [SGC23]. unauthorized [MVSSZ24]. Unawareness [XT24]. uncertainties [HCLRM20]. uncertainty [Ade21, YQL24, ZM25, vdB24]. Unchanged [Fur22]. Unclonable [KEM^{+}25a, PC22, ISD24]. uncontrolled [WZDZ25]. uncover [FO25, PdARD24]. Uncovering [NZZ^{+}22]. under-constrained [QPL^{+}23]. Underground [CHP21]. Undermining [PKK^{+}21]. Understanding [AAFF24, FK23, KS25a, KLC23, LLL^{+}23b, LLG20, OK23, Ram25, RCD23, TLS24b, WZZW20, AKA25, AZ20, GZGG20]. underwater [LZX^{+}23]. undiscovered [Spr23]. unexplored [OAA^{+}22]. unfavorable [ALD^{+}21]. Unified [YZW^{+}20, IB20]. uniform [BL25, HSL20]. uninformed [ZZ25]. Union [Bel25, RRB25, XRW^{+}25]. unique [DSS^{+}25, WS25b]. Unit [KJB^{+}25]. Universal [AHL^{+}21, CDLC24, DMZ24, DSD^{+}25, LLX25, MCM^{+}24, VS22, WLLL24a, WLLL24b, YCCZ21]. Unknown [ZWF^{+}24, CKV22, JZQ^{+}22, LCCZ24, LWL21, TG25, WTW21, YXL^{+}22, ZCPB23]. Unleashing [BDM^{+}24, ILVR^{+}24]. unlock [LTZ23]. unmanned [HMBY23, YGG24, SGC23]. Unmasking [CKT^{+}25, OA25]. Unpacking [KLC23, SWK20]. Unravelling [VB24]. unreliable [JWY^{+}25]. unstructured [AGM20, GLTH21, JLS22, LHC24, LTCR25, OMO20b, WTX^{+}22]. Unsupervised [PSY^{+}22, PTD20, AHZL25, dCBJSdS23, KDDM20a, KDDM20b, LMD^{+}22, LLJ^{+}23, QMW20b, SKKG25, WJT^{+}23, ZCPB23]. unusable [RCD23]. Unveiling [CGBK25, CYC^{+}24, HWD^{+}25, LHH23, ST25a, TSAG24, XZZ^{+}24, XKJG25, AKH20, TG25]. update [IJR^{+}25, OCMF23]. updates [XCC^{+}25, dSVST23]. Updating [KDTR25, CHWH25]. URL [BDA23, PC24, PD25b, dSFG20]. URL-based [dSFG20]. URLs [OA25]. usability [MKL21, SCTK23]. Usable [BF20, KKS^{+}22, GWY^{+}25, JCJ^{+}21, LCL23, NJS^{+}24]. usage [DLC24a, DZSCR24, LKL23, LHH^{+}20, Sch24]. usage-controlled [DZSCR24]. USB [HZX^{+}25]. Use [YK22b, Alq22, AABLS24, BRM21, BRO^{+}23, HP23, JDBB20, JE21, KKK23, MCJPS22, NLW24, XLG^{+}23, vdSFF20b]. use-after-free [XLG^{+}23]. used [DVAAM23]. usefulness [MZW^{+}25]. User [DKSS20, GBA22, KA23, MLZ^{+}23, Ogb21, PPGC24, QXS^{+}22, AR25, AGW25, BKA20, BTS23a, BTDH20, CHH25, CG24, CHK24, GGR^{+}24, GPAF25, GZGG20, GPD^{+}23, GKC22, HP23, KSMW23, LsJC^{+}22, LHAE22, LZH^{+}24, LFW^{+}22, MBFIdD21, MKC^{+}21, MD22, MBK25, MB20b, NJS^{+}24, OMO20a, OMO20b, PFHB21, PAA^{+}24, RCKB24, SU21, SAAW21, SSG^{+}25a, SLZ^{+}25, SKS21, TWW21, UWL22, VWB^{+}25, WPAG24, YC25, KKS^{+}22]. User-centric [KA23, MKC^{+}21, WPAG24]. User-empowered [PPGC24]. user-granted [OMO20a]. user-space [LsJC^{+}22]. user-specific [UWL22]. Users [HZYT23, AABE20, EMS23, FHW22, KS25a, LCL23, PLN23, PSCT21, PvSvdB25, XSL^{+}24]. userspace [LCSD20]. USIM [JZQ^{+}22]. Using [AFK^{+}24, CMFUA^{+}20, DLQ^{+}21, FO25, FJR23, GJB22, HZYT23, KC22, MRG21, QYJ22, TTA23, WLLC24, WRG^{+}21, AGW25, AKH24, AL23, AN23, ABAM22, AÇ23, AA22b, ATMN23, AK20, AYS20, ADAT25, AK24a, ADC^{+}20, AXAL24, AK24b, BS24a, BS24b, BKA20, BvtH^{+}23, BLDB24, BTM21, CHH25, CEO^{+}25, CEN^{+}25, CLN^{+}25, CCH^{+}23, CPV23, CPR21, CVZHN25, CZSW24, CLL^{+}21, CZZ^{+}21, CFZL22, CZLZ24, CHA^{+}24b, CSKD22, CJA25, CKV22, DP25, DZSL25, DAZ20, DKSS20, DUK25, DWX^{+}23, DG25, DS21, DMM^{+}23, DSC20, DHL^{+}23, FMA22, FBR25, FGY^{+}24, FCA25, FK24, FEA22, FKvdA^{+}24, GY23, GZS^{+}22, GR24, GSG23b, GTAK25, GL21, GB23, GZH^{+}24, GMS20, GKS24, GJCJ20, HDS21, HuRMMW22, HMMW22, He25, HNNN^{+}24, HW25, HLZ^{+}21, HYR^{+}22, JQYL22, kJLcL24, JIG22, JSYK25, KAM^{+}21, KDDM20a, KDDM20b, KJI^{+}24, KEM^{+}25a, KJJ^{+}24, KKM^{+}25, KFP^{+}24, KG20, KSSL22, KS23b]. using [LL22a, LLL^{+}20b, LFCD21, LXZ^{+}23, LLC^{+}23, LHL^{+}24b, LTZ^{+}24, LGH21, LRH^{+}21, LPW25, LQL^{+}25a, LWS^{+}25, MMJG24, MYS^{+}25, MPM23, MG24b, MWM^{+}25, MSG22, MMC24, MKM^{+}25, MV25, MBSE22, MC20, MBK25, MB20b, MMdR24, MMdSdS23, MLM^{+}23, MGG^{+}20, MKZ^{+}23, MTD^{+}24, NFVN^{+}22, NRS21, NCN24, NRKK23, NZZ^{+}22, OSPP22, OMO20a, OMO20b, ÖEÖ22, Özt24, PRTV22, PMA^{+}23, PGRS24, PJL23, PHH^{+}20, PHT^{+}25, PSCT21, QSA24, RBM^{+}25, RSSD23, RRM20, RYS25, Rs25, RDHSB22, RHH^{+}22, RSW^{+}21, RMMP23, RTBK21, RNSS23, RJ24, SA20, SK20, SS22, SAR24, SRY^{+}25, STB^{+}20, SJMP23, SB23, STPB25, SN24, SW22, SQL^{+}24b, SZW^{+}25, SD25, SYD^{+}25, SWK20, SLDL23, SHH^{+}24, SCL^{+}24, TSKG23, TG25, Tek21, TWW21, TH24, TD24, VT25, VAW^{+}20, WW25a, WLQ20, Wan21, WMK23, WZB^{+}23, WYL^{+}24, WXZ^{+}25, WS25a, WPS20, WPAG24, WHW^{+}24, XGS^{+}21, XLZ24, XQS25, YQL24, YK23]. using [YHY23b, YLW^{+}24b, ZFS25, ZLC^{+}23a, ZXB^{+}24, ZL24, ZLDX25, ZFT24, ZGS24, ZW24, ZSWT24, ZCJ^{+}20, ZZWF21, ZCD^{+}22, ZJJS^{+}22, ZCZZ24, HKM^{+}23]. utilising [JAAS^{+}23]. utility [ARAG23, CHJ22, CYHW25, GSS^{+}20b, JSYK25, LLL^{+}23b, LFL^{+}25, ZW24]. utilized [SRFC24]. Utilizing [CBK^{+}20, MKL21, EÖ25, HZH^{+}24, LDW25, LDS24, NTH25, XKY25]. utterance [ZLL24].

V [BF23, DCSY25, GYW^{+}23]. v2 [LFW^{+}25]. vaccine [SAJP22]. VADP [ZZP^{+}25]. VAE [LWLT21]. VAE-SVDD [LWLT21]. VAEPass [YHZ^{+}22]. valid [TG25]. validating [MZW^{+}25]. Validation [MMvS20, KBE^{+}24, KMAHU24, PA20]. validity [JDB^{+}23]. Value [EAN^{+}22, KLR^{+}25, KKP22, LWO^{+}25, MGTK22, SG21, XLH^{+}25]. Value-at-Risk [EAN^{+}22]. value-based [KKP22]. valued [OCB23]. values [FRF24, OSTO20, YKSK25]. values-based [FRF24]. vampire [AK24a]. VANET [CZS^{+}21, CQC^{+}24]. VANETs [DZSL25, HM23]. variable [GGK25, LHX23]. variables [WHLS24]. variance [LFW^{+}25]. variance-covariance [LFW^{+}25]. variant [MZW^{+}23]. variants [LQL25b]. variational [GHMG24, SHH^{+}24, YHZ^{+}22]. vault [RMS^{+}22]. vaults [XRW^{+}25]. VBS [MSD^{+}23]. VDExplainer [ZSJ^{+}25]. VDoTR [FWF^{+}23]. VDSimilar [SCL^{+}21]. VDTriplet [SCL^{+}24]. Vector [RDPR25, LaTL^{+}25, MSCJ21, WGRZ23, WSW24]. vectors [SLDL23]. Vehicle [KHG24, SGC23, AA22a, ARR^{+}25, CWYW25, CZM^{+}25, HMBY23, KJHL22, KJI^{+}24, LWLL25, NTBH^{+}22, PJL23, ZLC^{+}23a, ZL24]. VehicleLang [KJHL22]. Vehicles [NGC^{+}25, AAT23, AZM24, HFK^{+}24, IJR^{+}25, KKJ^{+}21, LQL^{+}25a, MSL^{+}25, PX21, RSLM25, ZC24, HXX^{+}24, SWL^{+}25]. Vehicular [JP25, ARR^{+}25, MBP20, WQL25]. veiled} [ST25a]. vein [LWZL25, ZSL24]. vendor [ASF^{+}23, SKS25]. veracity [AlD25]. VeracOS [AlD25]. VERI [CYL^{+}23]. verifiability [SGY23]. Verifiable [XCL^{+}24, YZL^{+}24, DSC20, HYZ^{+}24, JZL^{+}24, STH^{+}22, ZZCD20]. verification [AHSZ21, AIJ23, AHFL25, BC23b, CCL^{+}22, CPT^{+}22, CRS^{+}22, CZSW24, CQL^{+}21, CWYW25, DS21, GKC22, HZZ20, LHAE22, LWL^{+}25, LPX20, MHSK21, MV23, NGM^{+}24, RTBK21, SMC^{+}21, TXH^{+}23, TJ20, TZZ^{+}23, TO22, WK23, XHB^{+}21, YHY23a, ZQMC20, ZSL24]. verified [WGS^{+}25a]. verifier [HZL^{+}21]. verify [BOS^{+}21]. Verifying [SLBG21, CDGS23]. Verse [MS23]. Version [WWL25b]. versus [AKH20, JSYK25]. VERTFuzz [WWL25b]. vertical [CZ24, CDLC24, GB23, ZHG^{+}24]. vessel [GCN^{+}24, NSLM25]. via [AFAkJS25, BEFI25, BAC^{+}23, BWB^{+}21, CJC^{+}24, CHP^{+}24, CCC^{+}24, CJL^{+}25, CLM^{+}25, CNTBG21, DZZ^{+}22, DC23, FCSP21, GFM^{+}22, GZ24, GSE20, HLP22, HLQD23, HZL^{+}23, ISD24, KZFM24, LKL23, LZRR23, LCH^{+}24a, LZNZ24, LCH^{+}24b, LBW^{+}25, LaTL^{+}25, LL25, LLL^{+}22d, LXW^{+}25b, LXJ^{+}25, LWL21, LMW^{+}25, MLM20, NPCM24, RG25, SYZ^{+}23, SS24b, VMS20, WJP^{+}23, WZC^{+}24, WXT^{+}24, WZDZ25, XXZ^{+}21, XLY^{+}20, XLG^{+}23, YGC^{+}23, YHS24, YZP^{+}25, YWWH22, YXL^{+}22, ZGZL23, ZNF^{+}23, ZZS^{+}23]. vibration [LCL23]. Victim [CB22b, QMZ^{+}25, Sma25]. Victim-Attacker [QMZ^{+}25]. victimization [Kem23]. Video [HF25, BMV24, DMD^{+}20, FXL^{+}20, SJ21, ZXL^{+}24]. videos [KM24a]. Vietnamese [NCN24]. view [CG24, CYL^{+}21, CXC22, GMH^{+}23, JZY^{+}25, OMO20a, OZW^{+}21, RKA25, SJ21]. views [YCMM20]. vigilance [ND23]. VINCENT [DAAM24]. violation [FVF21, JZ23]. violations [HSW25, NM24]. Virtual [DS20, ETAB22, KEM^{+}25a, LKL23, MBR23, VBCM23, DSR23, HLL^{+}21, HZZ^{+}24, KBC21, OLH^{+}23, RPR^{+}21, YSM^{+}21]. virtualization [CBFH20, PDML25, TP25, ZXB^{+}24, ZTY^{+}20a, ZZW^{+}24]. Virtually [OLH^{+}23]. VirtualPatch [PDML25]. virus [SGSS22]. viruses [SYF^{+}21]. Vis [HKA24]. vision [AESF23, BTAK21, LGXZ22, DAAM24]. visit [LZL^{+}22]. Visitor [ZZP^{+}25]. Visitor-attribute-based [ZZP^{+}25]. visits [WK20]. visual [BT24, BCP22, LLD^{+}25, VRS^{+}25]. visualisation [TAA^{+}25]. visualization [CKV22, DGS^{+}23, FFLQ23, FFLQ24, LLLZ20, LGXZ22, PLP^{+}21, SD25]. visualization-based [FFLQ23, FFLQ24]. Visually [QWW^{+}22]. vitality [CCL^{+}21]. VMIFresh [DSR23]. vocal [CTV21]. Voice [BMAP24, LKK25, LCA^{+}23]. VoIP [CMMST22]. Voltage [ZL24]. Volume [ABAJ25a, FFLQ24, GMLB23, GZ20, GZG20, HZG20, KLSD24a, LCH^{+}24a, Ske24]. volumetric [PBG22]. Voluntary [CXC22, OCB21]. vote [LSL20]. Voting [KAK21]. VPN [CKT^{+}25]. VPPFL [HYZ^{+}24]. VRKeyLogger [LKL23]. vs [DA20, MDR20]. Vul [LTCR25]. Vulcan [JLS22]. VulD [ZLX^{+}24]. VulGAI [ZX23a]. VulLoc [TYZ^{+}24]. Vulnerabilities [BRV^{+}22, EZLC21, HZKN22, SKL23, WCP^{+}23, AAT23, AMC23, AMC25, Alq22, BCO^{+}22, CPR21, CMA23, FD23, JYF^{+}25, LZW^{+}24, LXC^{+}20, RVJ25, RSLM25, SSN23, SPZ^{+}20, Spr23, SCL^{+}21, SKS25, TSAG24, TH24, YPWS20, YFdCdC^{+}24, YSM^{+}21, YLZ^{+}25, ZZL^{+}22]. Vulnerability [CLHT24, CYL^{+}23, FWF^{+}23, GOW24, KM22, MVVL^{+}24, MKM^{+}25, SCL^{+}21, SCL^{+}24, TSW^{+}23, US20, VS20, WS22, ZJC^{+}25, ZSWT24, ZG24, AIA23, BEFI25, BRMH22, BEAE25, BLR23, BLSS22, BSG23, CDF^{+}20, CGH^{+}23, CCH^{+}22, ESM24, GSK23, GFH^{+}22, GYW^{+}23, GHZ^{+}24, HDS21, HWW^{+}25, JJS^{+}24, JK21a, JYF^{+}25, JWG^{+}24, JLB22, KKHK24, KLP^{+}24, KRR23, KD21, LXZ^{+}23, LLJ^{+}23, LTCR25, LWD^{+}25, LFHX23, LLP25, MVA25, MSD^{+}23, MFD25, MGG^{+}25, NBBS22, Özt24, SDCL25, SMC^{+}21, SOZ^{+}23b, TYZ^{+}24, TZZY25, WGRZ23, WJP^{+}23, WXT^{+}24, WTHL25, WBN^{+}20, WZY^{+}23, XCWC25, YLPZ21, ZKL^{+}23, ZX23a, ZLX^{+}24, ZSJ^{+}25, vdSFF20b]. Vulnerable [PB25, MSP^{+}20]. VulTR [HWW^{+}25]. VWA [TLHY25]. VWA-6G [TLHY25]. VxWorks [CDN21].

WAF [JE21]. wake [LL21b]. wake-up [LL21b]. Walking [MSD^{+}23]. wall [CZG^{+}22]. wallet [hPRPJ23]. warfare [AKH20, CSM^{+}24]. Warn [HP23]. Warning [WLDW24, WSW^{+}22]. warnings [SAJP22]. Wasm [CAFMB24]. Wasm-Mutate [CAFMB24]. Wasmati [BLSS22]. Wasserstein [ZZHY25]. Waste [MRG21]. water [zLXZrxZ25, RDM20, SRM22]. water-filling [zLXZrxZ25]. watermarking [CDGS23, FLLY24, QMZ^{+}23, QGW^{+}25, QZS^{+}25, YGC25, ALZ^{+}20]. watermarks [AKWR21]. waters [TKKF24]. waterways [KUA^{+}25]. WaTrojan [ZDZD24]. Wave [BLR23]. Wavelet [BH25a, ZDZD24, FEA22, XLZ24]. Wavelet-based [BH25a]. way [ESA^{+}23, MAF^{+}23]. ways [GBS22]. Wazuh [WS25a]. WBAN [ABK^{+}20, SCH^{+}20]. WCGAN [SSK23b]. WCGAN-GP [SSK23b]. Wdt [HNZ^{+}23]. Wdt-SCAN [HNZ^{+}23]. weakly [KT23]. weakly-supervised [KT23]. Weaponizing [DANS20]. wearable [AGW25, ACD^{+}24, GYL^{+}20, OANA25, WZDZ25, ZCL^{+}25]. wearables [LQH22b]. web [AIBKJ24, BA20, CBK22, HD25, Lee23, LLLZ21, LCL^{+}20, LGJW21, MBU24, MW20, MCvO24, MGG^{+}20, PFHB21, RCKB24, RT20, SB25, Tek21, WK23, YSM^{+}21, ZYGL25, dNPM^{+}21, ASG^{+}25, BDA23, CJKR21, MCJPS22, RHH^{+}22]. web-based [Tek21, YSM^{+}21]. Web-tracking [MCJPS22]. WebAssembly [BLSS22, CAMTB23, CAFMB24]. WebGuard [ZYGL25]. webpage [YZC25]. webpages [TCY^{+}20]. webshell [HY21, XLW24]. website [Fur22, LGJW21, LXW^{+}24, MG24b, MCJPS22, SK24a, WLY^{+}24, XZY^{+}24, YAC25, ZYL^{+}22, CNTBG21]. websites [LLL^{+}22d, MBPCC21, SM21, Sch24, SRFC24, XXZ^{+}21, ZY24]. webview [EZLC21]. WebVR [LKL23]. WEDoHTool [MHC^{+}25]. weight [CLX21, WPS20, ZKS25]. weight-mapping [CLX21]. Weighted [KLNW21, PD25a, AK24a, WGRZ23, ZSWT24]. weighting [BEAE25, CLX21, KK25b]. we’re [PNF24]. west [YAC25]. Western [CZK24, YAC25]. WF [zLXZrxZ25]. WF-LDPSR [zLXZrxZ25]. WF3A [WLY^{+}24]. WGAN [LL22a, VDN24]. WGAN-GP [LL22a]. Where [BTS23a]. Which [RGC24, ZCPB23, KK20]. White [CLS22]. White-box [CLS22]. whitelist [AMM^{+}21]. whitelisting [ZLC^{+}23a]. Who [LWZ^{+}25, RCA24, LCL23]. Wickr [KKHK24]. wide [VLV25]. widespread [HMB^{+}21]. WiFi [CLM^{+}25]. wild [PMF^{+}20, RVJ25, SCCZ20]. will [LLC^{+}23]. Window [KLR^{+}25, AGW25]. windows [AGAS25, BS24a, CSZ^{+}25, FGY^{+}24, GPC^{+}22b, KKHK24, RSW^{+}21, KS23b, AZ20, CZPX22, KDTR25, LWZ^{+}23, MPRB21]. Wipimization [OPK20]. wiping [OPK20]. wireless [AK20, AMVG24, BH25a, CWC^{+}25, DZ22, GYL^{+}20, HAKK21, HZX^{+}25, KGK23, KS20, LTL^{+}22, LLL23d, PRPS25, RJ21, TC23, Wan21, ZUKH23]. wiretapping [MBSE22]. Withdrawal [CAT24]. within [AS24, CFT22, DMB21, GGK25, KAK21, MGGR22, MBFIdD21, MBS20, WWS^{+}25]. without [KKB^{+}25]. WOLVES [KLR^{+}25]. Word [MHC^{+}25, BEAE25, DHL^{+}23, KS23b, WGRZ23, ZYL^{+}22]. word-level [ZYL^{+}22]. work [JZ23]. workers [BTA24, ND23]. workflow [AN21]. workforce [Fur21, NR24]. workload [LLYL21]. World [RD23, ACLA22, JZQ^{+}22, LYL^{+}23, LZD^{+}20, Mit20, SAAW21, WW25a, XHW^{+}22, ZM22]. worm [WX23, ZGNZ21]. wormhole [AB23, KJCL20, MT21c]. worrying [TMN24]. Would [RvdSM23]. WPS [PD25a]. wrapper [KS20, MR25, STB^{+}20]. wrapper-based [STB^{+}20]. wristband [WZDZ25]. write [LMW^{+}25]. WSN [AK24a]. WSNs [CLM^{+}25, SPSK24].

X [AKAFQ22, WSC^{+}20, YZW^{+}20]. X-repute [WSC^{+}20]. x86 [GSK23]. XACML [MYMC20]. xAI [KFP^{+}24]. Xception [MMdSdS23]. XGBoost [SK24a]. XLMR4MD [NCN24]. XMal [AMM23]. XRan [GKS24]. XSS [CGS22, CTH^{+}22, LFHH22, LFHX23, WYW^{+}22].

YARA [CLMU25]. Years [GMBV21, CCH^{+}22, Fur22]. YOLACT [ZHL^{+}22]. YOLOv8 [GCN^{+}24]. YouTube [ZXL^{+}24].

Z [BLR23]. Z-Wave [BLR23]. Zero [BWZ^{+}25, GAJ25, YLSJ23, AMA22, AZM24, BCO^{+}22, BOS^{+}21, CDJD24, FMAC21, HMK^{+}20, IO24, SS24b, VRS^{+}25, WFLX23, WHW^{+}24, YL20, CDJD24, SSS^{+}21, MRSP24]. zero-day [BCO^{+}22, CDJD24, HMK^{+}20, VRS^{+}25, WHW^{+}24, MRSP24]. Zero-knowledge [BWZ^{+}25]. zero-knowledge-proof-based [YL20]. Zero-Ran [CDJD24]. zero-shot [BCO^{+}22, CDJD24]. zero-trust [AMA22, BOS^{+}21, IO24, SS24b]. ZipAST [CWQZ25]. ZKP [GYL^{+}24]. ZKSA [BWZ^{+}25]. Zone [ETAB22]. Zone-based [ETAB22]. zoom [WHPL21]. ZTA [SSS^{+}21]. ZTWeb [WFLX23].

References

Alqahtani:2020:IBC

[AA20] Fatmah H. Alqahtani and Fawaz A. Alsulaiman. Is imagebased CAPTCHA secure against attacks based on machine learning? An experimental study. Computers & Security, 88(??):Article 101635, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

              URL https://www.sciencedirect.com/science/article/
              pii/S0167404818312185.

Alhogail:2021:AML

[AA21a] Areej Alhogail and Afrah Alsabih. Applying machine learning and natural language processing to detect phishing email. Computers & Security, 110(??):Article 102414, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002388.

AlHusain:2021:EFA

[AA21b] Reem AlHusain and Ali Alkhalifah. Evaluating fallback authentication research: a systematic literature review. Computers & Security, 111(??):Article 102487, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003114.

Ariffin:2021:IMR

[AA21c] Khairul Akram Zainol Ariffin and Faris Hanif Ahmad. Indicators for maturity and readiness for digital forensic investigation in era of industrial revolution 4.0. Computers & Security, 105(??):Article 102237, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000614.

Aksu:2022:MIO

[AA22a] Dogukan Aksu and Muhammed Ali Aydin. MGA-IDS: Optimal feature subset selection for anomaly detection framework on in-vehicle networks-CAN bus based on genetic algorithm and intrusion detection approach. Computers & Security, 118(??):Article 102717, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001134.

Alani:2022:AEF

[AA22b] Mohammed M. Alani and Ali Ismail Awad. AdStop: Efficient flow-based mobile adware detection using machine learning. Computers & Security, 117(??):Article 102718, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001146.

Arat:2023:APD

[AA23] Ferhat Arat and Sedat Akleylek. Attack path detection for IIoT enabled cyber physical systems: Revisited. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000846.

AlDaajeh:2024:SC

[AA24a] Saleh AlDaajeh and Saed Alrabaee. Strategic cybersecurity. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001469.

Ali:2024:LLA

[AA24b] Hala Ali and Irfan Ahmed. LAAKA: Lightweight Anonymous Authentication and Key Agreement scheme for secure fog-driven IoT systems. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000713.

Ahmad:2025:NEA

[AA25] Wasim Ahmad and Aitizaz Ali. A new era of advanced privacy solutions with a novel IoT framework on IFTTT. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003645.

Aleroud:2020:GPF

[AAAAS20] Ahmed Aleroud, Nisreen Abu-Alsheeh, and Emad Al-Shawakfa. A graph proximity feature augmentation approach for identifying accounts of terrorists on Twitter. Computers & Security, 99(??):Article 102056, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303291.

Alzaqebah:2023:HID

[AAAK23] Abdullah Alzaqebah, Ibrahim Aljarah, and Omar Al-Kadi. A hierarchical intrusion detection system based on extreme learning machine and nature-inspired optimization. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003492.

Ahmad:2022:CDL

[AAAT22] Rasheed Ahmad, Izzat Alsmadi, Wasim Alhamdani, and Lo’ai Tawalbeh. A comprehensive deep learning benchmark for IoT IDS. Computers & Security, 114(??):Article 102588, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821004119.

Alahmadi:2022:MMP

[AAB22] Amal Alahmadi, Norah Alkhraan, and Wojdan BinSaeedan. MPSAutodetect: a malicious Powershell script detection model based on stacked denoising auto-encoder. Computers & Security, 116(??):Article 102658, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000578.

AlJawaheri:2020:DTH

[AABE20] Husam Al Jawaheri, Mashael Al Sabah, Yazan Boshmaf, and Aiman Erbad. Deanonymizing Tor hidden service users through Bitcoin transactions analysis. Computers & Security, 89(??):Article 101684, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818309908.

Ayaburi:2024:IUP

[AABLS24] Emmanuel W. Ayaburi, Francis Kofi Andoh-Baidoo, Jaeung Lee, and Mikko Siponen. Investigating the use of protective technologies after data breach: the roles of psychological distance, technological service type and organizational justice. Computers & Security, 143(??):??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002025.

Arroyabe:2024:RRC

[AAFF24] Marta F. Arroyabe, Carlos F. A. Arranz, Ignacio Fernandez De Arroyabe, and Juan Carlos Fernandez de Arroyabe. Revealing the realities of cybercrime in small and medium enterprises: Understanding fear and taxonomic perspectives. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001275.

Adel:2025:RMI

[AAJP25] Amr Adel, Noor H. S. Alani, Tony Jan, and Mukesh Prasad. A review of major ICT failures and recovery strategies: Strengthening digital resilience. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003670.

Alkofahi:2024:MAI

[AAS24] Hamza Alkofahi, Heba Alawneh, and Anthony Skjellum. MitM attacks on intellectual property and integrity of additive manufacturing systems: a security analysis. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001111.

Asiri:2025:FRR

[AASC25] Mohammed Asiri, Arjun Arunasalam, Neetesh Saxena, and Z. Berkay Celik. Frontline responders: Rethinking indicators of compromise for industrial control system security. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001105.

AlMamun:2025:GPE

[AASWC25] Abdullah Al Mamun, Harith Al-Sahaf, Ian Welch, and Seyit Camtepe. Genetic programming for enhanced detection of Advanced Persistent Threats through feature construction. Computers & Security, 149(??):??, February 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004905.

Abuabed:2023:STM

[AAT23] Zaina Abuabed, Ahmad Alsadeh, and Adel Taweel. STRIDE threat model-based framework for assessing the vulnerabilities of modern vehicles. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003012.

Alghamdi:2023:CFD

[AB23] Rubayyi Alghamdi and Martine Bellaiche. A cascaded federated deep learning based framework for detecting wormhole attacks in IoT networks. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004060.

Alraja:2023:ISP

[ABA23] Mansour Naser Alraja, Usman Javed Butt, and Maysam Abbod. Information security policies compliance in a global setting: an employee’s perspective. Computers & Security, 129 (??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001189.

Alsharabi:2025:CTH

[ABAJ25a] Naif Alsharabi, Akashdeep Bhardwaj, Abdulaziz Ayaba, and Amr Jadi. Corrigendum to “Threat hunting for adversary impact inhibiting system recovery” [Computers & Security, Volume 154, July 2025, 104464]. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002214. See [ABAJ25b].

Alsharabi:2025:THA

[ABAJ25b] Naif Alsharabi, Akashdeep Bhardwaj, Abdulaziz Ayaba, and Amr Jadi. Threat hunting for adversary impact inhibiting system recovery. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001531. See corrigendum

[ABAJ25a].

Al-Bakaa:2022:NID

[ABAM22] Aliaa Al-Bakaa and Bahaa Al-Musawi. A new intrusion detection system based on using non-linear statistical analysis and features selection techniques. Computers & Security, 122(??):Article 102906, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002991.

Arshad:2022:PAL

[ABC22] Elham Arshad, Michele Benolli, and Bruno Crispo. Practical attacks on login CSRF in OAuth. Computers & Security, 121(??):Article 102859, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200253X.

Arfaoui:2020:CAA

[ABK^{+}20] Amel Arfaoui, Omar Rafik Merad Boudia, Ali Kribeche, Sidi-Mohammed Senouci, and Mohamed Hamdi. Context-aware access control and anonymous authentication in WBAN. Computers & Security, 88(??):Article 101496, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818304802.

Alaca:2023:CAD

[AÇ23] Yusuf Alaca and Yüksel Çelik. Cyber attack detection with QR code images using lightweight deep learning models. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004576.

Alam:2025:TMD

[AC25a] A. K. M. Mubashwir Alam and Keke Chen. TEE-MR: Developer-friendly data oblivious programming for trusted execution environments. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004243.

Ammannamma:2025:BIO

[AC25b] Tamminina Ammannamma and A. S. N. Chakravarthy. A bioinspired optimal feature with convolutional GhostNet based squeeze excited deep-scale capsule network for intrusion detection. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005273.

Alex:2023:CSI

[ACA^{+}23] Christin Alex, Giselle Creado, Wesam Almobaideen, Orieb Abu Alghanam, and Maha Saadeh. A comprehensive survey for IoT security datasets taxonomy, classification and machine learning mechanisms. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001931.

Arcolezi:2020:FNF

[ACC^{+}20] Héber H. Arcolezi, Jean-François Couchot, Selene Cerna, Christophe Guyeux, Guillaume Royer, Béchara Al Bouna, and Xiaokui Xiao. Forecasting the number of firefighter interventions per region with local-differential-privacy-based data. Computers & Security, 96(??):Article 101888, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301619.

Anand:2025:LLA

[ACCS25] P. Mohan Anand, P. V. Sai Charan, Hrushikesh Chunduri, and Sandeep Kumar Shukla. LARM: Linux Anti Ransomware Monitor. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482500389X.

Avola:2024:SEE

[ACD^{+}24] Danilo Avola, Luigi Cinque, Maria De Marsico, Alessio Fagioli, Gian Luca Foresti, Maurizio Mancini, and Alessio Mecca. Signal enhancement and efficient DTW-based comparison for wearable gait recognition. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005527.

Abrath:2025:MBM

[ACD25] Bert Abrath, Bart Coppens, and Bjorn De Sutter. MVXbased mitigation of position-independent code reuse. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482500344X.

Anglano:2020:AFA

[ACG20] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. The Android Forensics Automator (AnForA): a tool for the automated forensic analysis of Android applications. Computers & Security, 88(??):Article 101650, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301944.

Anley:2025:FFL

[ACGP25] Mulualem Bitew Anley, Pasquale Coscia, Angelo Genovese, and Vincenzo Piuri. FELACS: Federated learning with adaptive client selection for IoT DDoS attack detection. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003311.

Almashhadani:2022:MMF

[ACKS22] Ahmad O. Almashhadani, Domhnall Carlin, Mustafa Kaiiali, and Sakir Sezer. MFMCNS: a multi-feature and multi-classifier network-based system for ransomworm detection. Computers & Security, 121(??):Article 102860, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002541.

Alobaidi:2022:RWS

[ACLA22] Hind Alobaidi, Nathan Clarke, Fudong Li, and Abdulrahman Alruban. Real-world smartphone-based gait recognition. Computers & Security, 113(??):Article 102557, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003813.

Agyepong:2023:SMM

[ACRB23] Enoch Agyepong, Yulia Cherdantseva, Philipp Reinecke, and Pete Burnap. A systematic method for measuring the performance of a cyber security operations centre analyst. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003510.

Al:2021:SHN

[AD21] Samed Al and Murat Dener. STL-HDL: a new hybrid network intrusion detection system for imbalanced dataset on big data environment. Computers & Security, 110(??):Article 102435, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002595.

Aggarwal:2023:ABS

[AD23] Akshay Aggarwal and Ram Kumar Dhurkari. Association between stress and information security policy non-compliance behavior: a meta-analysis. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003832.

Alzubi:2025:CEL

[ADAT25] Ahmad Alzu’bi, Omar Darwish, Amjad Albashayreh, and Yahya Tashtoush. Cyberattack event logs classification using deep learning with semantic feature analysis. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005285.

Ashrafuzzaman:2020:DSF

[ADC^{+}20] Mohammad Ashrafuzzaman, Saikat Das, Yacine Chakhchoukh, Sajjan Shiva, and Frederick T. Sheldon. Detecting stealthy false data injection attacks in the smart grid using ensemblebased machine learning. Computers & Security, 97(??): Article 101994, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302674.

Aharon:2025:CRF

[ADDH25] Udi Aharon, Ran Dubin, Amit Dvir, and Chen Hajaj. A classification-by-retrieval framework for few-shot anomaly detection to detect API injection. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005558.

Adesemowo:2021:TCD

[Ade21] A. Kayode Adesemowo. Towards a conceptual definition for IT assets through interrogating their nature and epistemic uncertainty. Computers & Security, 105(??):Article 102131, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304041.

Al-Duwairi:2021:PHA

[ADJS21] Basheer Al-Duwairi, Moath Jarrah, and Ahmed S. Shatnawi. PASSVM: a highly accurate fast flux detection system. Computers & Security, 110(??):Article 102431, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002558.

Augello:2025:MMM

[ADL25] Andrea Augello, Alessandra De Paola, and Giuseppe Lo Re. M2FD: Mobile malware federated detection under concept drift. Computers & Security, 152(??):??, May 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000501.

Aebissa:2023:DIE

[ADM23] Berhanu Aebissa, Gurpreet Dhillon, and Million Meshesha. The direct and indirect effect of organizational justice on employee intention to comply with information security policy: the case of Ethiopian banks. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300158X.

Alhelaly:2023:WEF

[ADO23] Yasser Alhelaly, Gurpreet Dhillon, and Tiago Oliveira. When expectation fails and motivation prevails: the mediating role of awareness in bridging the expectancy-capability gap in mobile identity protection. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003802.

Al-Duwairi:2020:LNL

[ADÖU^{+}20] Basheer Al-Duwairi, Öznur Özkasap, Ahmet Uysal, Ceren Kocaogullar, and Kaan Yildirim. LogDoS: a novel loggingbased DDoS prevention mechanism in path identifier-based information centric networks. Computers & Security, 99(??): Article 102071, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303448.

Alves:2025:ECJ

[AdSJN25] Renato Solimar Alves, Jady Pamella Barbacena da Silva, Luiz Antonio Ribeiro Junior, and Rafael Rabelo Nunes. Enhancing cybersecurity in the judiciary: Integrating additional controls into the CIS framework. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002731.

Al-Eiadeh:2024:GGB

[AEA24] Mohammad Ryiad Al-Eiadeh and Mustafa Abdallah. GeniGraph: a genetic-based novel security defense resource allocation method for interdependent systems modeled by attack graphs. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400230X.

Amer:2022:RDL

[AES22] Eslam Amer and Shaker El-Sappagh. Robust deep learning early alarm prediction model based on the behavioural smell for Android malware. Computers & Security, 116(??):Article 102670, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000694.

Al-Emari:2023:DOT

[AESF23] Salam Al-E’mari, Yousef Sanjalawe, and Salam Fraihat. Detection of obfuscated Tor traffic based on bidirectional generative adversarial networks and vision transform. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004224.

Al-Fawareh:2025:DMA

[AFAkJS25] Mohmmad Al-Fawa’reh, Jumana Abu-khalaf, Naeem Janjua, and Patryk Szewczyk. Detection of on-manifold adversarial attacks via latent space transformation. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825001208.

Alves:2025:TLS

[AFAS25] Renan C. A. Alves, Otávio F. Freitas, Bruno C. Albertini, and Marcos A. Simplicio. Testing the limits of SPDM: Authentication of intermittently connected devices. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004474.

AlSabeh:2024:DDC

[AFK^{+}24] Ali AlSabeh, Kurt Friday, Elie Kfoury, Jorge Crichigno, and Elias Bou-Harb. On DGA detection and classification using P4 programmable switches. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003122.

Arenas:2025:RSO

[AFLR25] Mónica P. Arenas, Georgios Fotiadis, Gabriele Lenzini, and Mohammadamin Rakeei. Remote secure object authentication: Secure sketches, fuzzy extractors, and security protocols. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400436X.

Alvarez:2025:PPD

[ÁFM25] Iván Abellán Álvarez, Joaquín Delgado Fernández, and Sergio Potenciano Menci. Privacy-preserving distributed clustering: a fully homomorphic encrypted approach for time series. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002688.

Anley:2024:RDA

[AGAP24] Mulualem Bitew Anley, Angelo Genovese, Davide Agostinello, and Vincenzo Piuri. Robust DDoS attack detection with adaptive transfer learning. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002670.

Aryal:2025:ISC

[AGAS25] Kshitiz Aryal, Maanak Gupta, Mahmoud Abdelsalam, and Moustafa Saleh. Intra-section code cave injection for adversarial evasion attacks on windows PE malware file. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003797.

Ashley:2022:AAS

[AGBB22] Travis Ashley, Sri Nikhil Gupta Gourisetti, Newton Brown, and Christopher Bonebrake. Aggregate attack surface management for network discovery of operational technology. Computers & Security, 123(??):Article 102939, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003315.

Aaby:2023:OAT

[AGBT23] Peter Aaby, Mario Valerio Giuffrida, William J. Buchanan, and Zhiyuan Tan. An omnidirectional approach to touch-based continuous authentication. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000561.

Andriani:2025:CAC

[AGI^{+}25] Simone Andriani, Stefano Galantucci, Andrea Iannacone, Antonio Maci, and Giuseppe Pirlo. CNN-AutoMIC: Combining convolutional neural network and autoencoder to learn non-linear features for KNN-based malware image classification. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001968.

Aghasian:2020:AMS

[AGM20] Erfan Aghasian, Saurabh Garg, and James Montgomery. An automated model to score the privacy of unstructured information — social media case. Computers & Security, 92(??):Article 101778, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300638.

Andrade:2021:SGP

[AGVA21] Eduardo Andrade, Jorge Granjal, João P. Vilela, and Carlos Arantes. A security gateway for power distribution systems in open networks. Computers & Security, 111(??):Article 102492, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003163.

Ahmadian:2025:EUI

[AGW25] Rouhollah Ahmadian, Mehdi Ghatee, and Johan Wahlström. Enhancing user identification through batch averaging of independent window subsequences using smartphone and wearable data. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005716.

Akgun:2022:NDA

[AHC22] Devrim Akgun, Selman Hizal, and Unal Cavusoglu. A new DDoS attacks intrusion detection model based on deep learning for cybersecurity. Computers & Security, 118(??):Article 102748, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001432.

Akande:2025:LBR

[AHFL25] Ayodeji James Akande, Zhe Hou, Ernest Foo, and Qinyi Li. LTL-based runtime verification framework for cyberattack anomaly prediction in cyber-physical systems. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001440.

Alavizadeh:2021:EES

[AHKJJ21] Hooman Alavizadeh, Jin B. Hong, Dong Seong Kim, and Julian Jang-Jaccard. Evaluating the effectiveness of shuffle and redundancy MTD techniques in the cloud. Computers & Security, 102(??):Article 102091, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303643.

Aloufi:2021:ULR

[AHL^{+}21] Asma Aloufi, Peizhao Hu, Hang Liu, Sherman S. M. Chow, and Kim-Kwang Raymond Choo. Universal location referencing and homomorphic evaluation of geospatial query. Computers & Security, 102(??):Article 102137, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304107.

Abduaziz:2025:SST

[AHS25] Kayumov Abduaziz, Chansu Han, and Ji Sun Shin. Semisupervised traceability analysis of investigative scanners of darknet traffic. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003700.

Adja:2021:BBC

[AHSZ21] Yves Christian Elloh Adja, Badis Hammi, Ahmed Serhrouchni, and Sherali Zeadally. A blockchain-based certificate revocation management and status verification system. Computers & Security, 104(??):Article 102209, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100033X.

Ahmed:2024:ECO

[AHuHS24] Hameeza Ahmed, Muhammad Faraz Hyder, Muhammad Fahim ul Haque, and Paulo Cesar Santos. Exploring compiler optimization space for control flow obfuscation. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000051.

Abudurexiti:2025:EUA

[AHZL25] Yilixiati Abudurexiti, Guangjie Han, Fan Zhang, and Li Liu. An explainable unsupervised anomaly detection framework for Industrial Internet of Things. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004358.

Albanese:2023:FDV

[AIA23] Massimiliano Albanese, Ibifubara Iganibo, and Olutola Adebiyi. A framework for designing vulnerability metrics. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002924.

Abolfathi:2024:TEW

[AIBKJ24] Masoumeh Abolfathi, Srivani Inturi, Farnoush Banaei-Kashani, and J. Haadi Jafarian. Toward enhancing web privacy on HTTPS traffic: a novel SuperLearner attack model and an efficient defense approach with adversarial examples. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005837.

Afzaal:2023:FVF

[AIJ23] Hamra Afzaal, Muhammad Imran, and Muhammad Umar Janjua. Formal verification of fraud-resilience in a crowdsourcing consensus protocol. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002006.

Akhuseyinoglu:2020:CRA

[AJ20] Nuray Baltaci Akhuseyinoglu and James Joshi. A constraint and risk-aware approach to attribute-based access control for cyber-physical systems. Computers & Security, 96(??):Article 101802, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820300869.

Arshad:2020:FKM

[AJHA20] Humaira Arshad, Aman Jantan, Gan Keng Hoon, and Isaac Oludare Abiodun. Formal knowledge model for online social network forensics. Computers & Security, 89(??): Article 101675, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302160.

A:2024:CIC

[AJPK24] Nitish A., Hanumanthappa J., S. P. Shiva Prakash, and Kirill Krinkin. Class imbalance and concept drift invariant online botnet threat detection framework for heterogeneous IoT edge. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001214.

Alawami:2020:LFG

[AK20] Mohsen A. Alawami and Hyoungshick Kim. LocAuth: a fine-grained indoor location-based authentication system using wireless networks characteristics. Computers & Security, 89(??):Article 101683, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302226.

Arunachalam:2024:DMV

[AK24a] Rajesh Arunachalam and Erode Dhanapal Ruby Kanmani. Detection and mitigation of vampire attacks with secure routing in WSN using weighted RNN and optimal path selection. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002967.

Azimjonov:2024:DAL

[AK24b] Jahongir Azimjonov and Taehong Kim. Designing accurate lightweight intrusion detection systems for IoT networks using fine-tuned linear SVM and feature selectors. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005084.

Ali:2023:DDA

[AKA^{+}23] Hassan Ali, Muhammad Suleman Khan, Amer AlGhadhban, Meshari Alazmi, Ahmed Alzamil, Khaled Al-utaibi, and Junaid Qadir. Con-Detect: Detecting adversarially perturbed natural language inputs to deep classifiers through holistic analysis. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002778.

Almutairi:2025:BEC

[AKA25] Amirah Almutairi, BooJoong Kang, and Nawfal Alhashimy. Business email compromise: a systematic review of understanding, detection, and challenges. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003190.

Ali:2022:TXA

[AKAFQ22] Hassan Ali, Muhammad Suleman Khan, Ala Al-Fuqaha, and Junaid Qadir. Tamp-X: Attacking explainable natural language classifiers through tampered activations. Computers & Security, 120(??):Article 102791, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001857.

Almashhadani:2020:MSD

[AKCS20] Ahmad O. Almashhadani, Mustafa Kaiiali, Domhnall Carlin, and Sakir Sezer. MaldomDetector: a system for detecting algorithmically generated domain names with machine learning. Computers & Security, 93(??):Article 101787, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300729.

Andersson:2020:CVW

[AKH20] Annika Andersson, Fredrik Karlsson, and Karin Hedström. Consensus versus warfare — unveiling discourses in de jure information security standard development. Computers & Security, 99(??):Article 102035, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303084.

Ahmed:2024:CJE

[AKH24] Kashan Ahmed, Syed Khaldoon Khurshid, and Sadaf Hina. CyberEntRel: Joint extraction of cyber entities and relations using deep learning. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004893.

Ali:2021:IBF

[AKT21] Mansoor Ali, Hadis Karimipour, and Muhammad Tariq. Integration of blockchain and federated learning for Internet of Things: Recent advances and future challenges. Computers & Security, 108(??):Article 102355, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001796.

Aiken:2021:NNL

[AKWR21] William Aiken, Hyoungshick Kim, Simon Woo, and Jungwoo Ryoo. Neural network laundering: Removing blackbox backdoor watermarks from deep neural networks. Computers & Security, 106(??):Article 102277, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001012.

Ahn:2023:ISC

[AL23] Na Young Ahn and Dong Hoon Lee. IoT security: On-chip secure deletion scheme using ECC modulation in IoT appliances. Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003322.

Adeke:2025:IIF

[ALAN25] James Msughter Adeke, Guangjie Liu, Lord Amoah, and Ogonna Joshua Nwali. Investigating the impact of feature selection on adversarial transferability in intrusion detection system. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000161.

Ali:2021:SMR

[ALD^{+}21] Syed Emad Azhar Ali, Fong-Woon Lai, P. D. D. Dominic, Nicholas James Brown, Paul Benjamin Benjamin Lowry, and Rao Faizan Ali. Stock market reactions to favorable and unfavorable information security events: a systematic literature review. Computers & Security, 110(??):Article 102451, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002753.

AlDuaij:2025:VOS

[AlD25] Naser AlDuaij. VeracOS: an operating system extension for the veracity of files. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002548.

Almeida:2025:CAE

[Alm25] Fernando Almeida. Comparative analysis of EU-based cybersecurity skills frameworks. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000185.

Alqahtani:2022:SUV

[Alq22] Sultan S. Alqahtani. A study on the use of vulnerabilities databases in software engineering domain. Computers & Security, 116(??):Article 102661, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000608.

Alshaikh:2020:DCC

[Als20] Moneer Alshaikh. Developing cybersecurity culture to influence employee behavior: a practice perspective. Computers & Security, 98(??):Article 102003, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302765.

Ahvanooey:2020:ANI

[ALZ^{+}20] Milad Taleby Ahvanooey, Qianmu Li, Xuefang Zhu, Mamoun Alazab, and Jing Zhang. ANiTW: a Novel Intelligent Text Watermarking technique for forensic identification of spurious information on social media. Computers & Security, 90(??):Article 101702, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302391.

Adams:2025:HIC

[AM25] Matthew Adams and Tyler Moore. How informative are cybersecurity risk disclosures? Empirical analysis of firms targeted by ransomware. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003153.

Alshaikh:2021:ASM

[AMA21] Moneer Alshaikh, Sean B. Maynard, and Atif Ahmad. Applying social marketing to evaluate current security education training and awareness programs in organisations. Computers & Security, 100(??):Article 102090, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303631.

Adahman:2022:AZT

[AMA22] Zillah Adahman, Asad Waqar Malik, and Zahid Anwar. An analysis of zero-trust architecture and its costeffectiveness for organizational security. Computers & Security, 122(??):Article 102911, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003042.

Al-Maliki:2021:CRM

[AMAA21] Ossama Al-Maliki and Hisham Al-Assam. Challenge-response mutual authentication protocol for EMV contactless cards. Computers & Security, 103(??):Article 102186, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000109.

Ahmad:2022:E

[AMB22] Atif Ahmad, Sean Maynard, and Richard Baskerville. Editorial. Computers & Security, 112(??):Article 102530, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003540.

Alanazi:2023:SVA

[AMC23] Manar Alanazi, Abdun Mahmood, and Mohammad Jabed Morshed Chowdhury. SCADA vulnerabilities and attacks: a review of the state-of-the-art and open issues. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004205.

Alanazi:2025:ILD

[AMC25] Manar Alanazi, Abdun Mahmood, and Mohammad Jabed Morshed Chowdhury. ICS-LTU2022: a dataset for ICS vulnerabilities. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004486.

Ahmad:2021:HCO

[AMD^{+}21] Atif Ahmad, Sean B. Maynard, Kevin C. Desouza, James Kotsias, Monica T. Whitty, and Richard L. Baskerville. How can organizations develop situation awareness for incident response: a case study of management practice. Computers & Security, 101(??):Article 102122, February 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303953.

Anidjar:2024:ELD

[AMD^{+}24] Or Haim Anidjar, Revital Marbel, Ran Dubin, Amit Dvir, and Chen Hajaj. Extending limited datasets with GAN-like self-supervision for SMS spam detection. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003031.

Azeez:2021:AAW

[AMM^{+}21] Nureni Ayofe Azeez, Sanjay Misra, Ihotu Agbo Margaret, Luis Fernandez-Sanz, and Shafi’i Muhammad Abdulhamid. Adopting automated whitelist approach for detecting phishing attacks. Computers & Security, 108(??):Article 102328, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001528.

Alani:2023:XLM

[AMM23] Mohammed M. Alani, Atefeh Mashatan, and Ali Miri. XMal: a lightweight memory-based explainable obfuscated-malware detector. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300319X.

Alishahi:2022:ANR

[AMN22] Mina Alishahi, Vahideh Moghtadaiee, and Hojjat Navidan. Add noise to remove noise: Local differential privacy for feature selection. Computers & Security, 123(??):Article 102934, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003261.

Aslam:2020:FAF

[AMNR20] Mudassar Aslam, Bushra Mohsin, Abdul Nasir, and Shahid Raza. FoNAC — an automated Fog Node Audit and Certification scheme. Computers & Security, 93(??):Article 101759, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300432.

Allen:2024:SHU

[AMVG24] Ashley Allen, Alexios Mylonas, Stilianos Vidalis, and Dimitris Gritzalis. Smart homes under siege: Assessing the robustness of physical security against wireless network attacks. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005977.

Alvarez:2020:CSP

[AN20] Ramiro Alvarez and Mehrdad Nojoumian. Comprehensive survey on privacy-preserving protocols for sealed-bid auctions. Computers & Security, 88(??):Article 101502, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818306631.

Alfosail:2021:TFP

[AN21] Malak Alfosail and Peter Norris. Tor forensics: Proposed workflow for client memory artefacts. Computers & Security, 106(??):Article 102311, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001358.

Aktar:2023:TDA

[AN23] Sharmin Aktar and Abdullah Yasin Nur. Towards DDoS attack detection using deep learning approach. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482300161X.

AL-Nomasy:2025:TBK

[ANAAK25] Nadiah AL-Nomasy, Abdulelah Alamri, Ahamed Aljuhani, and Prabhat Kumar. Transformer-based knowledge distillation for explainable intrusion detection system. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825001063.

Allami:2025:ODF

[ANG^{+}25] Ali Allami, Tyler Nicewarner, Ken Goss, Ashish Kundu, Wei Jiang, and Dan Lin. Oblivious and distributed firewall policies for securing firewalls from malicious attacks. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005066.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Computers & Security, 91(??):??, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Computers & Security, 95(??):??, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. Computers & Security, 99(??):??, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:EBb

[Ano20d] Anonymous. Editorial Board. Computers & Security, 89(??): Article 101695, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302329.

Anonymous:2020:EBc

[Ano20e] Anonymous. Editorial Board. Computers & Security, 90 (??):Article 101727, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820300146.

Anonymous:2020:EBd

[Ano20f] Anonymous. Editorial Board. Computers & Security, 91 (??):Article 101768, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820300535.

Anonymous:2020:EBe

[Ano20g] Anonymous. Editorial Board. Computers & Security, 92 (??):Article 101797, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820300821.

Anonymous:2020:EBf

[Ano20h] Anonymous. Editorial Board. Computers & Security, 93 (??):Article 101842, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301152.

Anonymous:2020:EBg

[Ano20i] Anonymous. Editorial Board. Computers & Security, 94 (??):Article 101903, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301796.

Anonymous:2020:EBh

[Ano20j] Anonymous. Editorial Board. Computers & Security, 95 (??):Article 101912, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301887.

Anonymous:2020:EBi

[Ano20k] Anonymous. Editorial Board. Computers & Security, 96(??): Article 101980, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302534.

Anonymous:2020:EBj

[Ano20l] Anonymous. Editorial Board. Computers & Security, 97 (??):Article 102015, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302881.

Anonymous:2020:EBk

[Ano20m] Anonymous. Editorial Board. Computers & Security, 98(??): Article 102052, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303254.

Anonymous:2020:EBl

[Ano20n] Anonymous. Editorial Board. Computers & Security, 99(??): Article 102102, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303758.

Anonymous:2020:F

[Ano20o] Anonymous. February 2020. Computers & Security, 89(??):??, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:PJ

[Ano20p] Anonymous. In progress (January 2020). Computers & Security, 88(??):??, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:Jb

[Ano20q] Anonymous. July 2020. Computers & Security, 94(??):??, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:Ja

[Ano20r] Anonymous. June 2020. Computers & Security, 93(??):??, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:Ma

[Ano20s] Anonymous. March 2020. Computers & Security, 90(??):??, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:Mb

[Ano20t] Anonymous. May 2020. Computers & Security, 92(??):??, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:N

[Ano20u] Anonymous. November 2020. Computers & Security, 98(??):??, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:O

[Ano20v] Anonymous. October 2020. Computers & Security, 97(??):??, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2020:S

[Ano20w] Anonymous. September 2020. Computers & Security, 96(??): ??, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Computers & Security, 103(??):??, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Computers & Security, 107(??): ??, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Computers & Security, 111 (??):??, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Computers & Security, 100 (??):Article 102128, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820304016.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Computers & Security, 101 (??):Article 102183, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000079.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Computers & Security, 102 (??):Article 102193, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000171.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Computers & Security, 103 (??):Article 102234, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000584.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Computers & Security, 104 (??):Article 102255, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000791.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Computers & Security, 105 (??):Article 102295, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482100119X.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Computers & Security, 106 (??):Article 102333, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821001577.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Computers & Security, 107 (??):Article 102363, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821001875.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Computers & Security, 108(??): Article 102409, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821002339.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Computers & Security, 109 (??):Article 102428, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821002522.

Anonymous:2021:EBk

[Ano21n] Anonymous. Editorial Board. Computers & Security, 110(??): Article 102477, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003011.

Anonymous:2021:EBl

[Ano21o] Anonymous. Editorial Board. Computers & Security, 111(??): Article 102520, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003448.

Anonymous:2021:F

[Ano21p] Anonymous. February 2021. Computers & Security, 101(??): ??, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. January 2021. Computers & Security, 100(??): ??, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:J

[Ano21r] Anonymous. July 2021. Computers & Security, 106(??):??, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Computers & Security, 105(??):??, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021. Computers & Security, 102(??):??, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021. Computers & Security, 104(??):??, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:N

[Ano21v] Anonymous. November 2021. Computers & Security, 110 (??):??, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021. Computers & Security, 109(??): ??, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2021:RA

[Ano21x] Anonymous. Reviewer acknowledgement 2020. Computers & Security, 100(??):Article 102139, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304120.

Anonymous:2021:S

[Ano21y] Anonymous. September 2021. Computers & Security, 108 (??):??, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Computers & Security, 115(??):??, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Computers & Security, 119(??): ??, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Computers & Security, 123 (??):??, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Computers & Security, 112 (??):Article 102565, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003898.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Computers & Security, 113 (??):Article 102592, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821004156.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Computers & Security, 114 (??):Article 102618, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000177.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Computers & Security, 115 (??):Article 102649, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000487.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Computers & Security, 116 (??):Article 102701, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000992.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Computers & Security, 117 (??):Article 102736, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001316.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Computers & Security, 118 (??):Article 102767, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001626.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Computers & Security, 119 (??):Article 102795, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001894.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Computers & Security, 120(??): Article 102852, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822002462.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Computers & Security, 121 (??):Article 102881, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822002759.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Computers & Security, 122(??): Article 102927, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822003194.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Computers & Security, 123(??): Article 102971, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822003637.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Computers & Security, 113(??): ??, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Computers & Security, 112(??): ??, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Computers & Security, 118(??):??, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Computers & Security, 117(??):??, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Computers & Security, 114(??):??, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Computers & Security, 116(??):??, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Computers & Security, 122 (??):??, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. Computers & Security, 121(??): ??, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2022:RA

[Ano22x] Anonymous. Reviewer acknowledgement 2021. Computers & Security, 113(??):Article 102593, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004168.

Anonymous:2022:S

[Ano22y] Anonymous. September 2022. Computers & Security, 120 (??):??, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Computers & Security, 131(??): ??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Computers & Security, 135 (??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004242.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Computers & Security, 125(??): ??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004692.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Computers & Security, 126(??): ??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000214.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Computers & Security, 127(??): ??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000615.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001426.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Computers & Security, 129(??): ??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001840.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300216X.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Computers & Security, 131(??): ??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002432.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Computers & Security, 132 (??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823003073.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Computers & Security, 133(??): ??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003528.

Anonymous:2023:EBk

[Ano23n] Anonymous. Editorial Board. Computers & Security, 134 (??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823004042.

Anonymous:2023:EBl

[Ano23o] Anonymous. Editorial Board. Computers & Security, 135 (??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823004613.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Computers & Security, 125(??): ??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Computers & Security, 134 (??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Computers & Security, 133(??): ??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Computers & Security, 132 (??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Computers & Security, 143(??): ??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. Computers & Security, 147 (??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:EBb

[Ano24d] Anonymous. Editorial Board. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005205.

Anonymous:2024:EBc

[Ano24e] Anonymous. Editorial Board. Computers & Security, 137(??): ??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005916.

Anonymous:2024:EBd

[Ano24f] Anonymous. Editorial Board. Computers & Security, 138(??): ??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000233.

Anonymous:2024:EBe

[Ano24g] Anonymous. Editorial Board. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000592.

Anonymous:2024:EBf

[Ano24h] Anonymous. Editorial Board. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001342.

Anonymous:2024:EBg

[Ano24i] Anonymous. Editorial Board. Computers & Security, 141(??): ??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400169X.

Anonymous:2024:EBh

[Ano24j] Anonymous. Editorial Board. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002189.

Anonymous:2024:EBi

[Ano24k] Anonymous. Editorial Board. Computers & Security, 143(??): ??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002578.

Anonymous:2024:EBj

[Ano24l] Anonymous. Editorial Board. Computers & Security, 144 (??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003171.

Anonymous:2024:EBk

[Ano24m] Anonymous. Editorial Board. Computers & Security, 145(??): ??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003511.

Anonymous:2024:EBl

[Ano24n] Anonymous. Editorial Board. Computers & Security, 146 (??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003924.

Anonymous:2024:EBm

[Ano24o] Anonymous. Editorial Board. Computers & Security, 147 (??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824004103.

Anonymous:2024:F

[Ano24p] Anonymous. February 2024. Computers & Security, 137(??): ??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Ja

[Ano24q] Anonymous. January 2024. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Jc

[Ano24r] Anonymous. July 2024. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Jb

[Ano24s] Anonymous. June 2024. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Ma

[Ano24t] Anonymous. March 2024. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:Mb

[Ano24u] Anonymous. May 2024. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:N

[Ano24v] Anonymous. November 2024. Computers & Security, 146 (??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:O

[Ano24w] Anonymous. October 2024. Computers & Security, 145(??): ??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2024:S

[Ano24x] Anonymous. September 2024. Computers & Security, 144 (??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Aa

[Ano25a] Anonymous. April 2025. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Ab

[Ano25b] Anonymous. August 2025. Computers & Security, 155(??): ??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:D

[Ano25c] Anonymous. December 2025. Computers & Security, 159 (??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:EBa

[Ano25d] Anonymous. Editorial Board. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005431.

Anonymous:2025:EBb

[Ano25e] Anonymous. Editorial Board. Computers & Security, 149(??): ??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005649.

Anonymous:2025:EBc

[Ano25f] Anonymous. Editorial Board. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000227.

Anonymous:2025:EBd

[Ano25g] Anonymous. Editorial Board. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000653.

Anonymous:2025:EBe

[Ano25h] Anonymous. Editorial Board. Computers & Security, 152(??): ??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000963.

Anonymous:2025:EBf

[Ano25i] Anonymous. Editorial Board. Computers & Security, 153(??): ??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001166.

Anonymous:2025:EBg

[Ano25j] Anonymous. Editorial Board. Computers & Security, 154(??): ??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001725.

Anonymous:2025:EBh

[Ano25k] Anonymous. Editorial Board. Computers & Security, 155(??): ??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002093.

Anonymous:2025:EBi

[Ano25l] Anonymous. Editorial Board. Computers & Security, 156 (??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825002421.

Anonymous:2025:EBj

[Ano25m] Anonymous. Editorial Board. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825003244.

Anonymous:2025:EBk

[Ano25n] Anonymous. Editorial Board. Computers & Security, 158 (??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003530.

Anonymous:2025:EBl

[Ano25o] Anonymous. Editorial Board. Computers & Security, 159 (??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825004079.

Anonymous:2025:F

[Ano25p] Anonymous. February 2025. Computers & Security, 149(??): ??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Ja

[Ano25q] Anonymous. January 2025. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Jc

[Ano25r] Anonymous. July 2025. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Jb

[Ano25s] Anonymous. June 2025. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Ma

[Ano25t] Anonymous. March 2025. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:Mb

[Ano25u] Anonymous. May 2025. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:N

[Ano25v] Anonymous. November 2025. Computers & Security, 158 (??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:O

[Ano25w] Anonymous. October 2025. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:2025:S

[Ano25x] Anonymous. September 2025. Computers & Security, 156 (??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Aydin:2022:LST

[AOA22] Hakan Aydin, Zeynep Orman, and Muhammed Ali Aydin. A long short-term memory (LSTM)-based distributed denial of service (DDoS) detection and defense system design in public cloud network environment. Computers & Security, 118(??):Article 102725, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001201.

Arshad:2020:SAF

[AOAA20] Humaira Arshad, Esther Omlara, Isaac Oludare Abiodun, and Abdulhai Aminu. A semi-automated forensic investigation model for online social networks. Computers & Security, 97(??):Article 101946, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302224.

Aboukadri:2024:MLI

[AOM24] Sara Aboukadri, Aafaf Ouaddah, and Abdellatif Mezrioui. Machine learning in identity and access management systems: Survey and deep dive. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000300.

Arikkat:2024:ONO

[APA^{+}24] Dincy R. Arikkat, Vinod P., Rafidha Rehiman K. A., Serena Nicolazzo, Antonino Nocera, Georgiana Timpau, and Mauro Conti. OSTIS: a novel organization-specific threat intelligence system. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002955.

Alshaikh:2024:EPD

[APK24] Omar Alshaikh, Simon Parkinson, and Saad Khan. Exploring perceptions of decision-makers and specialists in defensive machine learning cybersecurity applications: the need for a standardised approach. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823006041.

Azzam:2023:FRI

[APPN23] Mazen Azzam, Liliana Pasquale, Gregory Provan, and Bashar Nuseibeh. Forensic readiness of industrial control systems under stealthy attacks. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004023.

Appiah:2021:DTP

[AQAK21] Benjamin Appiah, Zhiguang Qin, Ayidzoe Mighty Abra, and Ansuura JohnBosco Aristotle Kanpogninge. Decision tree pairwise metric learning against adversarial attacks. Computers & Security, 106(??):Article 102268, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000924.

Alex:2021:SBS

[AR21] Scaria Alex and T. Dhiliphan Rajkumar. Spider bird swarm algorithm with deep belief network for malicious JavaScript detection. Computers & Security, 107(??):Article 102301, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001255.

Acosta:2025:AHD

[AR25] Luca Hernández Acosta and Delphine Reinhardt. “Alexa, how do you protect my privacy?” A quantitative study of user preferences and requirements about smart speaker privacy settings. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006084.

Ashraf:2023:EPP

[ARAG23] Mohamed Ashraf, Sherine Rady, Tamer Abdelkader, and Tarek F. Gharib. Efficient privacy preserving algorithms for hiding sensitive high utility itemsets. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002705.

Aleisa:2020:PPA

[ARB20] Noura Aleisa, Karen Renaud, and Ivano Bongiovanni. The privacy paradox applies to IoT devices too: a Saudi Arabian study. Computers & Security, 96(??):Article 101897, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301711.

Arce:2020:CPC

[Arc20] Daniel G. Arce. Cybersecurity and platform competition in the cloud. Computers & Security, 93(??):Article 101774, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300584.

Acarali:2022:MSG

[ARR^{+}22] Dilara Acarali, K. Rajesh Rao, Muttukrishnan Rajarajan, Doron Chema, and Mark Ginzburg. Modelling smart grid IT-OT dependencies for DDoS impact propagation. Computers & Security, 112(??):Article 102528, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003527.

Aledhari:2025:SCA

[ARR^{+}25] Mohammed Aledhari, Rehma Razzak, Mohamed Rahouti, Abbas Yazdinejad, Reza M. Parizi, Basheer Qolomany, Mohsen Guizani, Junaid Qadir, and Ala Al-Fuqaha. Safeguarding connected autonomous vehicle communication: Protocols, intra- and inter-vehicular attacks and defenses. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000410.

Ansong:2025:RCS

[ARS^{+}25] Samuel Ansong, Windhya Rankothge, Somayeh Sadeghi, Hesamodin Mohammadian, Farrukh Bin Rashid, and Ali Ghorbani. Role of cybersecurity for a secure global communication eco-system: a comprehensive cyber risk assessment for satellite communications. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004619.

Aggarwal:2024:EEF

[AS24] Akshay Aggarwal and Shashi Kant Srivastava. Exploring eustress and fear: a new perspective on protection motivation in information security policy compliance within the financial sector. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001585.

AlDaajeh:2022:RNC

[ASA^{+}22] Saleh AlDaajeh, Heba Saleous, Saed Alrabaee, Ezedin Barka, Frank Breitinger, and Kim-Kwang Raymond Choo. The role of national cybersecurity strategies on the improvement of cybersecurity education. Computers & Security, 119(??):Article 102754, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001493.

Al-Saraireh:2023:KSB

[ASA23] Jaafer Al-Saraireh and Mohammad Rasool AlJa’afreh. Keystroke and swipe biometrics fusion to enhance smartphones authentication. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200414X.

Alom:2021:KGB

[ASAA21] Zulfikar Alom, Bikash Chandra Singh, Zeyar Aung, and Mohammad Abdul Azim. Knapsack graph-based privacy checking for smart environments. Computers & Security, 105(??):Article 102240, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100064X.

Altarawneh:2021:AAP

[ASB^{+}21] Amani Altarawneh, Fei Sun, Richard R. Brooks, Owulakemi Hambolu, Lu Yu, and Anthony Skjellum. Availability analysis of a permissioned blockchain with a lightweight consensus protocol. Computers & Security, 102(??):Article 102098, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303710.

Akiyama:2023:SAB

[ASF^{+}23] Mitsuaki Akiyama, Shugo Shiraishi, Akifumi Fukumoto, Ryota Yoshimoto, Eitaro Shioji, and Toshihiro Yamauchi. Seeing is not always believing: Insights on IoT manufacturing from firmware composition analysis and vendor survey. Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002997.

Allegretta:2025:WSI

[ASG^{+}25] Mauro Allegretta, Giuseppe Siracusano, Roberto González, Marco Gramaglia, and Juan Caballero. Web of shadows: Investigating malware abuse of Internet services. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004875.

Ashrif:2024:SLA

[ASH^{+}24] Fatma Foad Ashrif, Elankovan A. Sundararajan, Mohammad Kamrul Hasan, Rami Ahmad, Salwani Abdullah, and Raniyah Wazirali. Secured lightweight authentication for 6LoWPANs in machine-to-machine communications. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003079.

Arend:2020:PARa

[ASI^{+}20a] Isabel Arend, Asaf Shabtai, Tali Idan, Ruty Keinan, and Yoella Bereby-Meyer. Passive- and not active-risk tendencies predict cyber security behavior. Computers & Security, 96(??):Article 101929, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302066.

Arend:2020:PARb

[ASI^{+}20b] Isabel Arend, Asaf Shabtai, Tali Idan, Ruty Keinan, and Yoella Bereby-Meyer. Passive- and not active-risk tendencies predict cyber security behavior. Computers & Security, 97(??):Article 101964, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301735.

Alashwali:2020:EHS

[ASM20] Eman Salem Alashwali, Pawel Szalachowski, and Andrew Martin. Exploring HTTPS security inconsistencies: a crossregional perspective. Computers & Security, 97(??):Article 101975, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302480.

AlSlaiman:2023:EFN

[ASSW23] Muhanned AlSlaiman, Mohammed I. Salman, Mariam M. Saleh, and Bin Wang. Enhancing false negative and positive rates for efficient insider threat detection. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004588.

Ali:2025:SPM

[ATA^{+}25] Sijjad Ali, Dhani Bux Talpur, Adeel Abro, Khulud Salem Alshudukhi, Ghadah Naif Alwakid, Mamoona Humayun, Farhan Bashir, Shuaib Ahmed Wadho, and Asadullah Shah. Security and privacy in multi-cloud and hybrid cloud environments: Challenges, strategies, and future directions. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002883.

Awadhutkar:2022:CFE

[ATGK22] Payas Awadhutkar, Ahmed Tamrawi, Ryan Goluch, and Suresh Kothari. Control flow equivalence method for establishing sanctity of compiling. Computers & Security, 115 (??):Article 102608, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000074.

Aggarwal:2022:DEM

[ATJ^{+}22] Palvi Aggarwal, Omkar Thakoor, Shahin Jabbari, Edward A. Cranford, Christian Lebiere, Milind Tambe, and Cleotilde Gonzalez. Designing effective masking strategies for cyberdefense through human experimentation and cognitive models. Computers & Security, 117(??):Article 102671, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000700.

Ainslie:2023:CTI

[ATMA23] Scott Ainslie, Dean Thompson, Sean Maynard, and Atif Ahmad. Cyber-threat intelligence for security decisionmaking: a review and research agenda for practice. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002626.

Aldribi:2020:HBC

[ATMN20] Abdulaziz Aldribi, Issa Traoré, Belaid Moa, and Onyekachi Nwamuo. Hypervisor-based cloud intrusion detection through online multivariate statistical change tracking. Computers & Security, 88(??):Article 101646, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301907.

Alaoui:2023:CST

[ATMN23] El Arbi Abdellaoui Alaoui, Stephane Cedric Koumetio Tekouabou, Yassine Maleh, and Anand Nayyar. Corrigendum to “Towards to intelligent routing for DTN protocols using machine learning techniques [Simulation Modelling Practice and Theory 117 (2022) 102475]. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004047.

Ali:2025:ASS

[AUIH25] Mohammad Ali, Ahsan Ullah, Md. Rashedul Islam, and Rifat Hossain. Assessing of software security reliability: Dimensional security assurance techniques. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005364.

Araujo:2025:EID

[AV25] Iury Araujo and Marco Vieira. Enhancing intrusion detection in containerized services: Assessing machine learning models and an advanced representation for system call data. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001270.

Akanfe:2020:ACL

[AVR20] Oluwafemi Akanfe, Rohit Valecha, and H. Raghav Rao. Assessing country-level privacy risk for digital payment systems. Computers & Security, 99(??):Article 102065, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303382.

Arce:2024:EIR

[AWB24] Daniel Arce, Daniel W. Woods, and Rainer Böhme. Economics of incident response panels in cyber insurance. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000439.

Anthi:2021:HML

[AWJB21] Eirini Anthi, Lowri Williams, Amir Javed, and Pete Burnap. Hardening machine learning denial of service (DoS) defences against adversarial attacks in IoT smart home networks. Computers & Security, 108(??):Article 102352, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001760.

Altaf:2023:GLN

[AWN^{+}23] Tanzeela Altaf, Xu Wang, Wei Ni, Ren Ping Liu, and Robin Braun. NE-GConv: a lightweight node edge graph convolutional network for intrusion detection. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001955.

AlGhamdi:2020:ISG

[AWVG20] Sultan AlGhamdi, Khin Than Win, and Elena Vlahu-Gjorgievska. Information security governance challenges and critical success factors: Systematic review. Computers & Security, 99(??):Article 102030, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303035.

Asiri:2024:PRT

[AXAL24] Sultan Asiri, Yang Xiao, Saleh Alzahrani, and Tieshan Li. PhishingRTDS: a real-time detection system for phishing attacks using a deep learning model. Computers & Security, 141 (??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001445.

Alla:2025:TTM

[AYL25] Ildi Alla, Selma Yahia, and Valeria Loscri. TRIDENT: Trimodal Real-time Intrusion Detection Engine for New Targets. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003657.

Alzaylaee:2020:DDD

[AYS20] Mohammed K. Alzaylaee, Suleiman Y. Yerima, and Sakir Sezer. DL-Droid: Deep learning based Android malware detection using real devices. Computers & Security, 89(??):Article 101663, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819300161.

Amer:2020:DWM

[AZ20] Eslam Amer and Ivan Zelinka. A dynamic Windows malware detection and prediction method based on contextual understanding of API call sequence. Computers & Security, 92(??):Article 101760, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300444.

Amiri-Zarandi:2020:SML

[AZDF20] Mohammad Amiri-Zarandi, Rozita A. Dara, and Evan Fraser. A survey of machine learning-based solutions to protect privacy in the Internet of Things. Computers & Security, 96(??):Article 101921, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301978.

Amer:2021:MPM

[AZES21] Eslam Amer, Ivan Zelinka, and Shaker El-Sappagh. A multi-perspective malware detection approach through behavioral fusion of API call sequence. Computers & Security, 110(??):Article 102449, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482100273X.

Annabi:2024:TZT

[AZM24] Malak Annabi, Abdelhafid Zeroual, and Nadhir Messai. Towards zero trust security in connected vehicles: a comprehensive survey. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003237.

Ahvanooey:2023:ANF

[AZO^{+}23] Milad Taleby Ahvanooey, Mark Xuefang Zhu, Shiyan Ou, Hassan Dana Mazraeh, Wojciech Mazurczyk, Kim-Kwang Raymond Choo, and Chuan Li. AFPr-AM: a novel Fuzzy-AHP based privacy risk assessment model for strategic information management of social media platforms. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001736.

Bozkir:2020:LCH

[BA20] Ahmet Selman Bozkir and Murat Aydos. LogoSENSE: a companion HOG based logo detection scheme for phishing web page and e-mail brand recognition. Computers & Security, 95(??):Article 101855, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301279.

Begovic:2023:CRE

[BAAM23] Kenan Begovic, Abdulaziz Al-Ali, and Qutaibah Malluhi. Cryptographic ransomware encryption detection: Survey. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002596.

BenAtia:2025:MFM

[BAB^{+}25] Okba Ben Atia, Mustafa Al Samara, Ismail Bennis, Abdelhafid Abouaissa, Jaafar Gaber, and Pascal Lorenz. M3D-FL: Multi-layer malicious model detection for federated learning in IoT networks. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001336.

Bovenzi:2023:NAD

[BAC^{+}23] Giampaolo Bovenzi, Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, Valerio Persico, and Antonio Pescapé. Network anomaly detection methods in IoT environments via deep learning: a fair comparison of performance and robustness. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000779.

Bingu:2024:PCA

[BAD^{+}24] Rajesh Bingu, Salina Adinarayana, Jagjit Singh Dhatterwal, Sadam Kavitha, Eswar Patnala, and Hrushikesava Raju Sangaraju. Performance comparison analysis of classification methodologies for effective detection of intrusions. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001950.

Bena:2025:PML

[BAD^{+}25] Nicola Bena, Marco Anisetti, Ernesto Damiani, Chan Yeob Yeun, and Claudio A. Ardagna. Protecting machine learning from poisoning attacks: a risk-based approach. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001579.

Birthriya:2025:DPS

[BAJ25] Santosh Kumar Birthriya, Priyanka Ahlawat, and Ankit Kumar Jain. Detection and prevention of spear phishing attacks: a comprehensive survey. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000069.

Bui:2024:ABE

[BAM^{+}24] Hang Thanh Bui, Hamed Aboutorab, Arash Mahboubi, Yansong Gao, Nazatul Haque Sultan, Aufeef Chauhan, Mohammad Zavid Parvez, Michael Bewong, Rafiqul Islam, Zahid Islam, Seyit A. Camtepe, Praveen Gauravaram, Dineshkumar Singh, M. Ali Babar, and Shihao Yan. Agriculture 4.0 and beyond: Evaluating cyber threat intelligence sources and techniques in smart farming ecosystems. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000555.

Bosnjak:2020:SSE

[BB20] Leon Bosnjak and Bostjan Brumen. Shoulder surfing experiments: a systematic literature review. Computers & Security, 99(??):Article 102023, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302960.

Bidgoly:2020:SMC

[BBA20] Amir Jalaly Bidgoly, Hamed Jalaly Bidgoly, and Zeynab Arezoumand. A survey on methods and challenges in EEG based authentication. Computers & Security, 93(??):Article 101788, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300730.

Beaman:2021:RRA

[BBA^{+}21] Craig Beaman, Ashley Barkworth, Toluwalope David Akande, Saqib Hakak, and Muhammad Khurram Khan. Ransomware: Recent advances, analysis, challenges and future research directions. Computers & Security, 111(??):Article 102490, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100314X.

Boutigny:2020:SSC

[BBBB^{+}20] François Boutigny, Stéphane Betgé-Brezetz, Gregory Blanc, Antoine Lavignotte, Hervé Debar, and Houda Jmila. Solving security constraints for 5G slice embedding: a proof-of-concept. Computers & Security, 89(??):Article 101662, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404819302044.

Bhat:2023:SCB

[BBD23] Parnika Bhat, Sunny Behal, and Kamlesh Dutta. A system call-based Android malware detection approach with homogeneous and heterogeneous ensemble machine learning. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001876.

Berens:2024:CDD

[BBD^{+}24] Benjamin Maximilian Berens, Mark Bohlender, Heike Dietmann, Chiara Krisam, Oksana Kulyk, and Melanie Volkamer. Cookie disclaimers: Dark patterns and lack of transparency. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004170.

Brilingaite:2020:FCD

[BBJ20] Agne Brilingaite, Linas Bukauskas, and Ausrius Juozapavicius. A framework for competence development and assessment in hybrid cybersecurity exercises. Computers & Security, 88(??):Article 101607, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404819301580.

Boahen:2021:NAD

[BBMW21] Edward Kwadwo Boahen, Brunel Elvire Bouya-Moko, and Changda Wang. Network anomaly detection in a controlled environment based on an enhanced PSOGSARFC. Computers & Security, 104(??):Article 102225, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000493.

Berger:2020:AII

[BBR20] Stephan Berger, Olga Bürger, and Maximilian Röglinger. Attacks on the Industrial Internet of Things — development of a multi-layer taxonomy. Computers & Security, 93(??):Article 101790, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300754.

Brown:2024:RFO

[BBZS24] Dennis Brown, Gunjan Batra, Humayun Zafar, and Khawaja Saeed. Reducing fraud in organizations through information security policy compliance: an information security controls perspective. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002633.

B:2023:ALS

[BC23a] Judy Flavia B. and Balika J. Chelliah. Artificial lizard search optimized fuzzy logic approach to addressing authentication and data security challenges in P2P cloud environments. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003851.

Beltran:2023:PTM

[BC23b] Marta Beltrán and Miguel Calvo. A privacy threat model for identity verification based on facial recognition. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002341.

Borzacchiello:2021:FMF

[BCD21] Luca Borzacchiello, Emilio Coppa, and Camil Demetrescu. FUZZOLIC: Mixing fuzzing and concolic execution. Computers & Security, 108(??):Article 102368, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001929.

Botacin:2020:WNT

[BCdGG20] Marcus Botacin, Fabricio Ceschin, Paulo de Geus, and André Grégio. We need to talk about antiviruses: challenges and pitfalls of AV evaluations. Computers & Security, 95(??):Article 101859, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301310.

Baseri:2025:CEF

[BCGC25a] Yaser Baseri, Vikas Chouhan, Ali Ghorbani, and Aaron Chow. Corrigendum to ”Evaluation framework for quantum security risk assessment: a comprehensive strategy for quantum-safe transition” [Computers & Security, 150, 104272]. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825003086.

Baseri:2025:EFQ

[BCGC25b] Yaser Baseri, Vikas Chouhan, Ali Ghorbani, and Aaron Chow. Evaluation framework for quantum security risk assessment: a comprehensive strategy for quantum-safe transition. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005789.

Baseri:2024:NQS

[BCH24] Yaser Baseri, Vikas Chouhan, and Abdelhakim Hafid. Navigating quantum security risks in networked environments: a comprehensive study of quantum-safe network protocols. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001846.

Benyahya:2022:ACS

[BCKN22] Meriem Benyahya, Anastasija Collen, Sotiria Kechagia, and Niels Alexander Nijdam. Automated city shuttles: Mapping the key challenges in cybersecurity, privacy and standards to future developments. Computers & Security, 122(??):Article 102904, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002978.

Borzacchiello:2025:DER

[BCM^{+}25] Luca Borzacchiello, Matteo Cornacchia, Davide Maiorca, Giorgio Giacinto, and Emilio Coppa. DroidReach++: Exploring the reachability of native code in Android applications.

Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003463.

Barros:2022:MSZ

[BCO^{+}22] Pedro H. Barros, Eduarda T. C. Chagas, Leonardo B. Oliveira, Fabiane Queiroz, and Heitor S. Ramos. Malware-SMELL: a zero-shot learning strategy for detecting zero-day vulnerabilities. Computers & Security, 120(??):Article 102785, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001808.

Bernal:2022:NJC

[BCP22] Sergio López Bernal, Alberto Huertas Celdrán, and Gregorio Martínez Pérez. Neuronal jamming cyberattack over invasive BCIs affecting the resolution of tasks requiring visual capabilities. Computers & Security, 112(??):Article 102534, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003588.

Botacin:2021:CPM

[BCS^{+}21] Marcus Botacin, Fabricio Ceschin, Ruimin Sun, Daniela Oliveira, and André Grégio. Challenges and pitfalls in malware research. Computers & Security, 106(??):Article 102287, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001115.

Bhardwaj:2023:ENN

[BD23] Sonam Bhardwaj and Mayank Dave. Enhanced neural network-based attack investigation framework for network forensics: Identification, detection, and analysis of the attack. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004315.

Bozkir:2023:GNN

[BDA23] Ahmet Selman Bozkir, Firat Coskun Dalgic, and Murat Aydos. GramBeddings: a new neural network for URL based identification of phishing Web pages through N-gram embeddings. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200356X.

Botacin:2022:AUM

[BDC^{+}22] Marcus Botacin, Felipe Duarte Domingues, Fabrício Ceschin, Raphael Machnicki, Marco Antonio Zanata Alves, Paulo Lício de Geus, and André Grégio. AntiViruses under the microscope: a hands-on perspective. Computers & Security, 112(??):Article 102500, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003242.

Basile:2023:DIA

[BDC^{+}23] Cataldo Basile, Bjorn De Sutter, Daniele Canavese, Leonardo Regano, and Bart Coppens. Design, implementation, and automation of a risk management approach for man-at-the-end software protection. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002316.

Buccafurri:2024:ESP

[BDLP24] Francesco Buccafurri, Vincenzo De Angelis, Sara Lazzaro, and Andrea Pugliese. Enforcing security policies on interacting authentication systems. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000725.

Boffa:2024:LUL

[BDM^{+}24] Matteo Boffa, Idilio Drago, Marco Mellia, Luca Vassio, Danilo Giordano, Rodolfo Valentim, and Zied Ben Houidi. LogPrécis: Unleashing language models for automated malicious log analysis: Précis: a concise summary of essential points, statements, or facts. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001068.

BenYahya:2025:BBW

[BEAE25] Aissa Ben Yahya, Hicham El Akhal, El Mehdi Ismaili Alaoui, and Abdelbaki El Belrhiti El Alaoui. Bayes-based word weighting for enhanced vulnerability classification in critical infrastructure systems. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001403.

Bonaventura:2025:CSD

[BEB25] Davide Bonaventura, Sergio Esposito, and Giampaolo Bella. A case of smart devices that compromise home cybersecurity. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005923.

Barredo:2025:GFV

[BEFI25] Jorge Barredo, Maialen Eceiza, Jose Luis Flores, and Mikel Iturbe. GJALLARHORN: a framework for vulnerability detection via electromagnetic side-channel analysis in embedded systems. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003815.

Beltran:2025:AAU

[Bel25] Marta Beltrán. AI algorithms under scrutiny: GDPR, DSA, AI act and CRA as pillars for algorithmic security and privacy in the European union. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003177.

Bhana:2020:PKD

[BF20] Bhaveer Bhana and Stephen Flowerday. Passphrase and keystroke dynamics authentication: Usable security. Computers & Security, 96(??):Article 101925, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302017.

Bove:2023:BSS

[BF23] Davide Bove and Julian Funk. Basic secure services for standard RISC-V architectures. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003255.

Bellante:2025:EPQ

[BFC^{+}25] Armando Bellante, Tommaso Fioravanti, Michele Carminati, Stefano Zanero, and Alessandro Luongo. Evaluating the potential of quantum machine learning in cybersecurity: a case-study on PCA-based intrusion detection systems. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000306.

Bayer:2023:MLF

[BFR23] Markus Bayer, Tobias Frey, and Christian Reuter. Multilevel fine-tuning, data augmentation, and few-shot learning for specialized cyber threat intelligence. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003401.

Barraclough:2021:ICP

[BFW21] P. A. Barraclough, G. Fehringer, and J. Woodward. Intelligent cyber-phishing detection for online. Computers & Security, 104(??):Article 102123, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303965.

Botacin:2025:TMR

[BG25] Marcus Botacin and Heitor Gomes. Towards more realistic evaluations: the impact of label delays in malware detection pipelines. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004279.

Brown:2024:AML

[BGA24] Austin Brown, Maanak Gupta, and Mahmoud Abdelsalam. Automated machine learning for deep learning based malware detection. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004923.

Buldas:2020:AEA

[BGL^{+}20] Ahto Buldas, Olga Gadyatskaya, Aleksandr Lenin, Sjouke Mauw, and Rolando Trujillo-Rasua. Attribute evaluation on attack trees with incomplete information. Computers & Security, 88(??):Article 101630, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301774.

Berardi:2023:DFA

[BGM^{+}23] Davide Berardi, Saverio Giallorenzo, Andrea Melis, Simone Melloni, Loris Onori, and Marco Prandini. Data flooding against ransomware: Concepts and implementations. Computers & Security, 131(??):??, August 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002055.

Basha:2025:WBC

[BH25a] Nora Basha and Bechir Hamdaoui. Wavelet-based CSI reconstruction for improved wireless security through channel reciprocity. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001129.

Boshoff:2025:CFC

[BH25b] Dutliff Boshoff and Gerhard P. Hancke. A classifications framework for continuous biometric authentication (2018– 2024). Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005911.

Bitzer:2023:MIM

[BHL^{+}23] Michael Bitzer, Björn Häckel, Daniel Leuthe, Joshua Ott, Bastian Stahl, and Jacqueline Strobel. Managing the inevitable — a maturity model to establish incident response management capabilities. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004424.

Barrere:2023:CPA

[BHO23] Martín Barrère, Chris Hankin, and Dean O’Reilly. Cyberphysical attack graphs (CPAGs): Composable and scalable attack graphs for cyber-physical systems. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002584.

Bongiovanni:2024:PPI

[BHW24] Ivano Bongiovanni, David M. Herold, and Simon J. Wilde. Protecting the play: an integrative review of cybersecurity in and for sports events. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003699.

Buckley:2022:CCL

[BHWE22] Oliver Buckley, Duncan Hodges, Jonathan Windle, and Sally Earl. CLICKA: Collecting and leveraging identity cues with keystroke dynamics. Computers & Security, 120(??):Article 102780, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001754.

Beu:2023:FPA

[BJZ^{+}23] N. Beu, A. Jayatilaka, M. Zahedi, A. Babar, L. Hartley, W. Lewinsmith, and I. Baetu. Falling for phishing attempts: an investigation of individual differences that are associated with behavior in a naturalistic phishing simulation. Computers & Security, 131(??):??, August 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002237.

Berry:2022:EOF

[BK22] Cate Berry and Nikos Komninos. Efficient optimisation framework for convolutional neural networks with secure multiparty computation. Computers & Security, 117(??):Article 102679, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000773.

Balakrishnan:2020:ICD

[BKA20] Vimala Balakrishnan, Shahzaib Khan, and Hamid R. Arabnia. Improving cyberbullying detection using Twitter users’ psychological features and machine learning. Computers & Security, 90(??):Article 101710, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302470.

Binbeshr:2021:SRP

[BKPZ21] Farid Binbeshr, M. L. Mat Kiah, Lip Yee Por, and A. A. Zaidan. A systematic review of PIN-entry methods resistant to shoulder-surfing attacks. Computers & Security, 101(??):Article 102116, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303898.

Baskerville:2022:CRE

[BKS22] Richard L. Baskerville, Jongwoo Kim, and Carl Stucke. The cybersecurity risk estimation engine: a tool for possibility based risk analysis. Computers & Security, 120(??):Article 102752, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200147X.

Bamber:2025:HCL

[BKSA25] Sukhvinder Singh Bamber, Aditya Vardhan Reddy Katkuri, Shubham Sharma, and Mohit Angurala. A hybrid CNN-LSTM approach for intelligent cyber intrusion detection system. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004516.

Budde:2023:CCE

[BKV^{+}23] Carlos E. Budde, Anni Karinsalo, Silvia Vidor, Jarno Salonen, and Fabio Massacci. Consolidating cybersecurity in Europe: a case study on job profiles assessment. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004746.

Bergner:2025:UAH

[BL25] Kevin Bergner and Dieter Landes. A uniform assessment of host-based intrusion detection data sets. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001919.

Bao:2024:SBD

[BLC^{+}24] Huaifeng Bao, Wenhao Li, Huashan Chen, Han Miao, Qiang Wang, Zixian Tang, Feng Liu, and Wen Wang. Stories behind decisions: Towards interpretable malware family classification with hierarchical attention. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002487.

Bhardwaj:2024:OLL

[BLDB24] Sonam Bhardwaj, Adrian Shuai Li, Mayank Dave, and Elisa Bertino. Overcoming the lack of labeled data: Training malware detection models using adversarial domain adaptation. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000701.

Braghin:2023:MBA

[BLR23] Chiara Braghin, Mario Lilli, and Elvinia Riccobene. A modelbased approach for vulnerability analysis of IoT security protocols: the Z-Wave case study. Computers & Security, 127(??): ??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004291.

Brito:2022:WES

[BLSS22] Tiago Brito, Pedro Lopes, Nuno Santos, and José Fragoso Santos. Wasmati: an efficient static vulnerability scanner for WebAssembly. Computers & Security, 118(??):Article 102745, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001407.

Bostani:2024:EPE

[BM24] Hamid Bostani and Veelasha Moonsamy. EvadeDroid: a practical evasion attack on machine learning for black-box Android malware detection. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005862.

Bilika:2024:HMM

[BMAP24] Domna Bilika, Nikoletta Michopoulou, Efthimios Alepis, and Constantinos Patsakis. Hello me, meet the real me: Voice synthesis attacks on voice assistants. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005278.

Boran:2020:AIL

[BMH20] Tudor Boran, Muhamet Martinaj, and Md Shafaeat Hossain. Authorship identification on limited samplings. Computers & Security, 97(??):Article 101943, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302194.

Bax:2021:MBR

[BMH21] Samantha Bax, Tanya McGill, and Val Hobbs. Maladaptive behaviour in response to email phishing threats: the roles of rewards and response costs. Computers & Security, 106(??):Article 102278, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001024.

Belanger:2022:LSI

[BMM22] France Bélanger, Jürgen Maier, and Michaela Maier. A longitudinal study on improving employee information protective knowledge and behaviors. Computers & Security, 116 (??):Article 102641, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000402.

Berens:2024:TMP

[BMV24] Benjamin M. Berens, Mattia Mossano, and Melanie Volkamer. Taking 5 minutes protects you for 5 months: Evaluating an anti-phishing awareness video. Computers & Security, 137(??): ??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005308.

benOthmane:2020:ESI

[bOJW20] Lotfi ben Othmane, Doug Jacobson, and Edgar Weippl. Editorial: Special issue on security and privacy in smart cyberphysical systems. Computers & Security, 88(??):Article 101612, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301610.

Buck:2021:NTA

[BOS^{+}21] Christoph Buck, Christian Olenberger, André Schweizer, Fabiane Völter, and Torsten Eymann. Never trust, always verify: a multivocal literature review on current knowledge and research gaps of zero-trust. Computers & Security, 110(??):Article 102436, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002601.

Bhattar:2024:FDI

[BP24] Poornachandratejasvi Laxman Bhattar and Naran M. Pindoriya. False data injection attack with max-min optimization in smart grid. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000622.

Binosi:2023:BAR

[BPCZ23] Lorenzo Binosi, Mario Polino, Michele Carminati, and Stefano Zanero. BINO: Automatic recognition of inline binary functions from template classes. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002225.

Butavicius:2020:WBT

[BPL^{+}20] Marcus Butavicius, Kathryn Parsons, Meredith Lillie, Agata McCormac, Malcolm Pattinson, and Dragana Calic. When believing in technology leads to poor cyber security: Development of a trust in technical controls scale. Computers & Security, 98(??):Article 102020, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302935.

Bland:2020:MLC

[BPW^{+}20] John A. Bland, Mikel D. Petty, Tymaine S. Whitaker, Katia P. Maxwell, and Walter Alan Cantrell. Machine learning cyberattack and defense strategies. Computers & Security, 92(??):Article 101738, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404818309799.

Butt:2023:TSP

[BQA^{+}23] Muhammad Atif Butt, Adnan Qayyum, Hassan Ali, Ala Al-Fuqaha, and Junaid Qadir. Towards secure private and trustworthy human-centric embedded machine learning: an emotion-aware facial recognition case study. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004503.

Brandao:2021:HCO

[BRM21] André Brandão, João S. Resende, and Rolando Martins. Hardening cryptographic operations through the use of secure enclaves. Computers & Security, 108(??):Article 102327, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001516.

Beaman:2022:FVD

[BRMH22] Craig Beaman, Michael Redbourne, J. Darren Mummery, and Saqib Hakak. Fuzzing vulnerability discovery techniques: Survey, challenges and future directions. Computers & Security, 120(??):Article 102813, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002073.

Bridges:2023:TST

[BRO^{+}23] Robert A. Bridges, Ashley E. Rice, Sean Oesch, Jeffrey A. Nichols, Cory Watson, Kevin Spakes, Savannah Norem, Mike Huettel, Brian Jewell, Brian Weber, Connor Gannon, Olivia Bizovi, Samuel C. Hollifield, and Samantha Erwin. Testing SOAR tools in use. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001116.

Bang:2022:IIB

[BRV^{+}22] Ankur O. Bang, Udai Pratap Rao, Andrea Visconti, Alessandro Brighente, and Mauro Conti. An IoT inventory before deployment: a survey on IoT protocols, communication technologies, vulnerabilities, attacks, and future research directions. Computers & Security, 123(??):Article 102914, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003078.

Barreto:2023:DII

[BRWF23] Carlos Barreto, Olof Reinert, Tobias Wiesinger, and Ulrik Franke. Duopoly insurers’ incentives for data quality under a mandatory cyber data sharing regime. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300202X.

Bryant:2020:ISA

[BS20] Blake D. Bryant and Hossein Saiedian. Improving SIEM alert metadata aggregation with a novel kill-chain based classification model. Computers & Security, 94(??):Article 101817, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030095X.

Babu:2024:BMB

[BS24a] Sarath Babu and Virendra Singh. BD-MDLC: Behavior description-based enhanced malware detection for windows environment using longformer classifier. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003365.

Baker:2024:FSI

[BS24b] Qanita Bani Baker and Alaa Samarneh. Feature selection for IoT botnet detection using equilibrium and Battle Royale Optimization. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003651.

Blinowski:2025:FTE

[BS25] Grzegorz Blinowski and Michał Szaknis. Fuzzing trusted execution environments with Rust. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005017.

Brunner:2020:RMP

[BSFB20] Michael Brunner, Clemens Sauerwein, Michael Felderer, and Ruth Breu. Risk management practices in information security: Exploring the status quo in the DACH region. Computers & Security, 92(??):Article 101776, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300614.

Brust:2023:RBV

[BSG23] Clemens-Alexander Brust, Tim Sonnekalb, and Bernd Gruner. ROMEO: a binary vulnerability detection dataset for exploring Juliet through the lens of assembly language. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000755.

Ben-Shimol:2025:DCF

[BSLK^{+}25] Lavi Ben-Shimol, Danielle Lavi, Eitan Klevansky, Oleg Brodt, Dudu Mimran, Yuval Elovici, and Asaf Shabtai. Detection of compromised functions in a serverless cloud environment. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005674.

Burke:2025:DEE

[BSO25] Wendy Burke, Andrew Stranieri, and Taiwo Oseni. From disempowerment to empowerment: Crafting a healthcare cybersecurity self-assessment. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482400453X.

BustamanteV:2025:TSC

[BSR^{+}25] DiegoF BustamanteV, Luis E. Sánchez, David G. Rosado, Antonio Santos-Olmo, and Eduardo Fernández-Medina. Towards a sustainable cybersecurity framework for Agriculture 4.0 based on a systematic analysis of proposals. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003396.

Bostanipour:2021:JOL

[BT21] Behnaz Bostanipour and George Theodorakopoulos. Joint obfuscation of location and its semantic information for privacy protection. Computers & Security, 107(??):Article 102310, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001346.

Baltuttis:2024:EVR

[BT24] Dennik Baltuttis and Timm Teubner. Effects of visual risk indicators on phishing detection behavior: an eye-tracking experiment. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002451.

Baltuttis:2024:TCB

[BTA24] Dennik Baltuttis, Timm Teubner, and Marc T. P. Adam. A typology of cybersecurity behavior among knowledge workers. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000427.

Bozkir:2021:CTA

[BTAK21] Ahmet Selman Bozkir, Ersan Tahillioglu, Murat Aydos, and Ilker Kara. Catch them alive: a malware detection approach through memory forensics, manifold learning and computer vision. Computers & Security, 103(??):Article 102166, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304399.

Breitinger:2020:SSU

[BTDH20] Frank Breitinger, Ryan Tully-Doyle, and Courtney Hassenfeldt. A survey on smartphone user’s security choices, awareness and education. Computers & Security, 88(??):Article 101647, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404819301919.

Bashendy:2023:IRS

[BTE23] May Bashendy, Ashraf Tantawy, and Abdelkarim Erradi. Intrusion response systems for cyber-physical systems: a comprehensive survey. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003765.

Butavicius:2022:WPK

[BTH22] Marcus Butavicius, Ronnie Taib, and Simon J. Han. Why people keep falling for phishing scams: the effects of time pressure and deception cues on the detection of phishing emails. Computers & Security, 123(??):Article 102937, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003297.

Bollmann:2021:RRT

[BTM21] Chad A. Bollmann, Murali Tummala, and John C. McEachen. Resilient real-time network anomaly detection using novel non-parametric statistical tests. Computers & Security, 102 (??):Article 102146, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820304193.

Bansal:2023:WAF

[BTS23a] Gaurav Bansal, Jason Thatcher, and Sebastian Walter Schuetz. Where authorities fail and experts excel: Influencing Internet users’ compliance intentions. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000743.

Bojarajulu:2023:IIB

[BTS23b] Balaganesh Bojarajulu, Sarvesh Tanwar, and Thipendra Pal Singh. Intelligent IoT-BOTNET attack detection model with optimized hybrid classification model. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822004564.

Beuran:2023:CAM

[BVB^{+}23] Razvan Beuran, Jan Vykopal, Daniela Belajová, Pavel Čeleda, Yasuo Tan, and Yoichi Shinoda. Capability assessment methodology and comparative analysis of cybersecurity training platforms. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000305.

Bellman:2023:SAC

[BvO23] Christopher Bellman and Paul C. van Oorschot. Systematic analysis and comparison of security advice as datasets. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003819.

Bekkers:2023:PYB

[BvtH^{+}23] Luuk Bekkers, Susanne van ’t Hoff-de Goede, Ellen Misana ter Huurne, Ynze van Houten, Remco Spithoven, and Eric Rutger Leukfeldt. Protecting your business against ransomware attacks? explaining the motivations of entrepreneurs to take future protective measures against cybercrimes using an extended protection motivation theory model. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000093.

Brown:2021:DFM

[BWB^{+}21] Dane Brown, T. Owens Walker, Justin A. Blanco, Robert W. Ives, Hau T. Ngo, James Shey, and Ryan Rakvic. Detecting firmware modification on solid state drives via current draw analysis. Computers & Security, 102(??):Article 102149, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304223.

Bai:2025:ZSM

[BWZ^{+}25] Fenhua Bai, Zikang Wang, Kai Zeng, Chi Zhang, Tao Shen, Xiaohui Zhang, and Bei Gong. ZKSA: Secure mutual attestation against TOCTOU zero-knowledge proof based for IoT devices. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004413.

Bhamare:2020:CIC

[BZE^{+}20] Deval Bhamare, Maede Zolanvari, Aiman Erbad, Raj Jain, Khaled Khan, and Nader Meskin. Cybersecurity for industrial control systems: a survey. Computers & Security, 89(??):Article 101677, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302172.

Cabrera-Arteaga:2024:WMF

[CAFMB24] Javier Cabrera-Arteaga, Nicholas Fitzgerald, Martin Monperrus, and Benoit Baudry. Wasm-Mutate: Fast and effective binary diversification for WebAssembly. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000324.

Cabrera-Arteaga:2023:WDM

[CAMTB23] Javier Cabrera-Arteaga, Martin Monperrus, Tim Toady, and Benoit Baudry. WebAssembly diversification for malware evasion. Computers & Security, 131(??):??, August 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002067.

Chowdhury:2020:TPHa

[CAT20a] Noman H. Chowdhury, Marc T. P. Adam, and Timm Teubner. Time pressure in human cybersecurity behavior: Theoretical framework and countermeasures. Computers & Security, 97(??):Article 101931, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302078.

Chowdhury:2020:TPHb

[CAT20b] Noman H. Chowdhury, Marc T. P. Adam, and Timm Teubner. Time pressure in human cybersecurity behavior: Theoretical framework and countermeasures. Computers & Security, 97(??):Article 101963, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301747. See withdrawal notice

[CAT24].

Chowdhury:2024:WNT

[CAT24] Noman H. Chowdhury, Marc T. P. Adam, and Timm Teubner. Withdrawal notice to ‘Time pressure in human cybersecurity behavior: Theoretical framework and countermeasures’ [Computers and Security, 97 (2020) 101931]. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003036. See [CAT20b].

Challagidad:2020:MDD

[CB20] Praveen S. Challagidad and Mahantesh N. Birje. Multidimensional dynamic trust evaluation scheme for cloud environment. Computers & Security, 91(??):Article 101722, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300092.

Calvo:2022:MRB

[CB22a] Miguel Calvo and Marta Beltrán. A model for risk-based adaptive security controls. Computers & Security, 115(??): Article 102612, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000116.

Connolly:2022:RRC

[CB22b] Alena Yuryna Connolly and Hervé Borrion. Reducing ransomware crime: Analysis of victims’ payment decisions. Computers & Security, 119(??):Article 102760, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001559.

Chaudhuri:2025:FIC

[CBB25] Abhik Chaudhuri, Rajat Kumar Behera, and Pradip Kumar Bala. Factors impacting cybersecurity transformation: an Industry 5.0 perspective. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482400573X.

Compastie:2020:VSI

[CBFH20] Maxime Compastié, Rémi Badonnel, Olivier Festor, and Ruan He. From virtualization security issues to cloud protection opportunities: an in-depth analysis of system virtualization models. Computers & Security, 97(??):Article 101905, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301814.

Chamikara:2020:PPF

[CBK^{+}20] M. A. P. Chamikara, P. Bertok, I. Khalil, D. Liu, and S. Camtepe. Privacy preserving face recognition utilizing differential privacy. Computers & Security, 97(??):Article 101951, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302273.

Chahal:2022:PAA

[CBK22] Navdeep S. Chahal, Preeti Bali, and Praveen Kumar Khosla. A Proactive Approach to assess web application security through the integration of security tools in a Security Orchestration Platform. Computers & Security, 122(??):Article 102886, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002802.

Chen:2025:FMH

[CBPW25] Wei Chen, Zhiyuan Bai, Gaoyuan Pan, and Jian Wang. A fast modularity hardware Trojan detection technique for large scale gate-level netlists. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004164.

Castiglione:2025:SOR

[CBS25] Gianpietro Castiglione, Giampaolo Bella, and Daniele Francesco Santamaria. SecOnto: Ontological representation of security directives. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004553.

Chang:2020:BCS

[CC20] Lennon Y. C. Chang and Nicholas Coppel. Building cyber security awareness in a developing country: Lessons from Myanmar. Computers & Security, 97(??):Article 101959, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302352.

Cimmino:2025:ODR

[CCBGC25] Andrea Cimmino, Juan Cano-Benito, and Raúl García-Castro. Open Digital Rights Enforcement framework (ODRE): From descriptive to enforceable policies. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005881.

Chen:2024:LTL

[CCC^{+}24] Jinyin Chen, Zhiqi Cao, Ruoxi Chen, Haibin Zheng, Xiao Li, Qi Xuan, and Xing Yang. Like teacher, like pupil: Transferring backdoors via feature-based knowledge distillation. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003468.

Cartwright:2023:CIB

[CCE23] Anna Cartwright, Edward Cartwright, and Esther Solomon Edun. Cascading information on best practice: Cyber security risk management in UK micro and small businesses and the role of IT companies. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001980.

Cui:2022:ESV

[CCH^{+}22] Lei Cui, Jiancong Cui, Zhiyu Hao, Lun Li, Zhenquan Ding, and Yongji Liu. An empirical study of vulnerability discovery methods over the past ten years. Computers & Security, 120(??):Article 102817, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002115.

Cao:2023:TCF

[CCH^{+}23] Yungui Cao, Jiazhen Chen, Liqing Huang, Tianqian Huang, and Feng Ye. Three-classification face manipulation detection using attention-based feature decomposition. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004163.

Chew:2021:PID

[CCL^{+}21] Chit-Jie Chew, Ying-Chin Chen, Jung-San Lee, Chih-Lung Chen, and Kuo-Yu Tsai. Preserving indomitable DDoS vitality through resurrection social hybrid botnet. Computers & Security, 106(??):Article 102284, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001085.

Calzavara:2022:BRR

[CCL^{+}22] Stefano Calzavara, Lorenzo Cazzaro, Claudio Lucchese, Federico Marcuzzi, and Salvatore Orlando. Beyond robustness: Resilience verification of tree-based classifiers. Computers & Security, 121(??):Article 102843, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002371.

Cai:2024:OES

[CCL^{+}24] Liujia Cai, Guangying Cai, Siqi Lu, Guangsong Li, and Yongjuan Wang. Observational equivalence and security games: Enhancing the formal analysis of security protocols. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000865.

Chekole:2020:CCE

[CCO^{+}20] Eyasu Getahun Chekole, Sudipta Chattopadhyay, Martín Ochoa, Huaqun Guo, and Unnikrishnan Cheramangalath. CIMA: Compiler-enforced resilience against memory safety attacks in cyber-physical systems. Computers & Security, 94 (??):Article 101832, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301061.

Chen:2024:LAP

[CCS^{+}24] Zhiyu Chen, Feng Chen, Yiming Sun, Mingjie Wang, Shangdong Liu, and Yimu Ji. Local aggressive and physically realizable adversarial attacks on 3D point cloud. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004492.

Caraveo-Cacep:2024:RSI

[CCVMZ24] Miguel Antonio Caraveo-Cacep, Rubén Vázquez-Medina, and Antonio Hernández Zavala. A review on security implementations in soft-processors for IoT applications. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005874.

Chen:2024:SLL

[CCW^{+}24] Yiren Chen, Mengjiao Cui, Ding Wang, Yiyang Cao, Peian Yang, Bo Jiang, Zhigang Lu, and Baoxu Liu. A survey of large language models for cyber threat detection. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003213.

Canfora:2020:IVF

[CDF^{+}20] Gerardo Canfora, Andrea Di Sorbo, Sara Forootani, Antonio Pirozzi, and Corrado Aaron Visaggio. Investigating the vulnerability fixing process in OSS projects: Peculiarities and challenges. Computers & Security, 99(??):Article 102067, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303400.

Ceragioli:2022:CMF

[CDG22] Lorenzo Ceragioli, Pierpaolo Degano, and Letterio Galletta. Can my firewall system enforce this policy? Computers & Security, 117(??):Article 102683, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000815.

Coscia:2023:ITS

[CDG^{+}23] Antonio Coscia, Vincenzo Dentamaro, Stefano Galantucci, Antonio Maci, and Giuseppe Pirlo. An innovative twostage algorithm to optimize firewall rule ordering. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003334.

Costa:2023:FVS

[CDGS23] Gabriele Costa, Pierpaolo Degano, Letterio Galletta, and Simone Soderi. Formally verifying security protocols built on watermarking and jamming. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000433.

Cen:2024:ZRS

[CDJD24] Mingcan Cen, Xizhen Deng, Frank Jiang, and Robin Doss. Zero-Ran Sniff: a zero-day ransomware early detection method based on zero-shot learning. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001500.

Chen:2024:UAB

[CDLC24] Peng Chen, Xin Du, Zhihui Lu, and Hongfeng Chai. Universal adversarial backdoor attacks to fool vertical federated learning. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005114.

Cimino:2020:MCM

[CDM^{+}20] Mario G. C. A. Cimino, Nicoletta De Francesco, Francesco Mercaldo, Antonella Santone, and Gigliola Vaglini. Model checking for malicious family detection and phylogenetic analysis in mobile environment. Computers & Security, 90(??): Article 101691, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302287.

Casola:2024:SSD

[CDMO24] Valentina Casola, Alessandra De Benedictis, Carlo Mazzocca, and Vittorio Orbinato. Secure software development and testing: a model-based methodology. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005497.

Coppolino:2025:EET

[CDMR25] Luigi Coppolino, Salvatore D’Antonio, Giovanni Mazzeo, and Luigi Romano. An experimental evaluation of TEE technology: Benchmarking transparent approaches based on SGX, SEV, and TDX. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001464.

Cotroneo:2021:TCC

[CDN21] Domenico Cotroneo, Luigi De Simone, and Roberto Natella. Timing covert channel analysis of the VxWorks MILS embedded hypervisor under the common criteria security certification. Computers & Security, 106(??):Article 102307, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001310.

Canonico:2025:ECP

[CEN^{+}25] Roberto Canonico, Giovanni Esposito, Annalisa Navarro, Simon Pietro Romano, Giancarlo Sperlì, and Andrea Vignali. Empowered cyber-physical systems security using both network and physical data. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000719.

Calvo:2025:RLD

[CEO^{+}25] Albert Calvo, Santiago Escuder, Nil Ortiz, Josep Escrig, and Maxime Compastié. RBD24: a labelled dataset with risk activities using log application data. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005960.

Cole:2025:CID

[CFRB25] Kelly A. Cole, Alexander L. Francis, Marcus K. Rogers, and Joe Balazs. Can individual differences in cognitive capacity predict cybersecurity performance? Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001853.

Colombo:2022:EAB

[CFT22] Pietro Colombo, Elena Ferrari, and Engin Deniz Tümer. Efficient ABAC based information sharing within MQTT environments under emergencies. Computers & Security, 120(??):Article 102829, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002231.

Coutinho:2023:ICM

[CFYBF23] Bruno Coutinho, João Ferreira, Iryna Yevseyeva, and Vitor Basto-Fernandes. Integrated cybersecurity methodology and supporting tools for healthcare operational information systems. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000998.

Chen:2022:ENB

[CFZL22] Aiguo Chen, Yang Fu, Xu Zheng, and Guoming Lu. An efficient network behavior anomaly detection using a hybrid DBN-LSTM network. Computers & Security, 114(??):Article 102600, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821004235.

Casaril:2024:SSU

[CG24] Francesco Casaril and Letterio Galletta. Securing SatCom user segment: a study on cybersecurity challenges in view of IRIS^{2}. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001007.

Chee:2024:IFM

[CGBK24] Kok Onn Chee, Mengmeng Ge, Guangdong Bai, and Dan Dongseong Kim. IoTSecSim: a framework for modelling and simulation of security in Internet of Things. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004443.

Chee:2025:UEI

[CGBK25] Kok Onn Chee, Mengmeng Ge, Guangdong Bai, and Dan Dongseong Kim. Unveiling the evolution of IoT threats: Trends, tactics, and simulation analysis. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002263.

Choo:2021:MAI

[CGCY21] Kim-Kwang Raymond Choo, Keke Gai, Luca Chiaraviglio, and Qing Yang. A multidisciplinary approach to Internet of Things (IoT) cybersecurity and risk management. Computers & Security, 102(??):Article 102136, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304090.

Cifuentes:2023:RPA

[CGH^{+}23] Cristina Cifuentes, François Gauthier, Behnaz Hassanshahi, Padmanabhan Krishnan, and Davin McCall. The role of program analysis in security vulnerability detection: Then and now. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003735.

Chaudhary:2022:SHE

[CGS22] Pooja Chaudhary, Brij B. Gupta, and A. K. Singh. Securing heterogeneous embedded devices against XSS attack in intelligent IoT system. Computers & Security, 118(??):Article 102710, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001080.

Chen:2025:PSA

[CH25] Yong-Yi Chen and Yu-Ling Hsueh. A personalized and semantic-aware approach for trajectory protection. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002974.

Chaudhary:2024:DBC

[Cha24a] Sunil Chaudhary. Driving behaviour change with cybersecurity awareness. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001597.

Chen:2024:IQI

[CHA^{+}24b] Sheng-Shan Chen, Ren-Hung Hwang, Asad Ali, Ying-Dar Lin, Yu-Chih Wei, and Tun-Wen Pai. Improving quality of indicators of compromise using STIX graphs. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002773.

C:2025:ECS

[CHH25] Kiran K. C., Md Shafaeat Hossain, and Carl Haberfeld. Exploring capacitive swipe gesture for user authentication using a new large dataset. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001610.

Chen:2022:SSU

[CHJ22] Jianan Chen, Qin Hu, and Honglu Jiang. Strategic signaling for utility control in audit games. Computers & Security, 118(??):Article 102721, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200116X.

Chen:2024:SFE

[CHK24] Jiayi Chen, Urs Hengartner, and Hassan Khan. SHRIMPS: a framework for evaluating multi-user, multi-modal implicit authentication systems. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005047.

Cabrero-Holgueras:2021:MLS

[CHP21] José Cabrero-Holgueras and Sergio Pastrana. A methodology for large-scale identification of related accounts in underground forums. Computers & Security, 111(??):Article 102489, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003138.

Chen:2024:DSR

[CHP^{+}24] Xi Chen, Wei Huang, Ziwen Peng, Wei Guo, and Fan Zhang. Diversity supporting robustness: Enhancing adversarial robustness via differentiated ensemble predictions. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001627.

Chen:2025:CNR

[CHWH25] Yanan Chen, Botao Hou, Bin Wu, and Hao Hu. CD-Net: Robust mobile traffic classification against apps updating. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005200.

Chen:2023:GPI

[CHZ^{+}23] Jinyin Chen, Shulong Hu, Haibin Zheng, Changyou Xing, and Guomin Zhang. GAIL-PT: an intelligent penetration testing framework with generative adversarial imitation learning. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004473.

Chuah:2025:DLB

[CJA25] Edward Chuah, Arshad Jhumka, and Aladdin Ayesh. Deep learning-based prediction of reflection attacks using NetFlow data. Computers & Security, 156(??):??, September 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002160.

Chen:2024:ACA

[CJC^{+}24] Ruoxi Chen, Haibo Jin, Jinyin Chen, Haibin Zheng, Shilian Zheng, Xiaoniu Yang, and Xing Yang. AdvCheck: Characterizing adversarial examples via local gradient checking. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004509.

Cen:2025:RRB

[CJD25] Mingcan Cen, Frank Jiang, and Robin Doss. RansoGuard: a RNN-based framework leveraging pre-attack sensitive APIs for early ransomware detection. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005996.

Choi:2022:EDP

[CJJ^{+}22] Yeseul Choi, Yunjong Jeong, Daehee Jang, Brent Byunghoon Kang, and Hojoon Lee. EmuID: Detecting presence of emulation through microarchitectural characteristic on ARM. Computers & Security, 113(??):Article 102569, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100393X.

Calzavara:2021:MWS

[CJKR21] Stefano Calzavara, Hugo Jonker, Benjamin Krumnow, and Alvise Rabitti. Measuring Web session security at scale. Computers & Security, 111(??):Article 102472, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002960.

Chen:2025:AAM

[CJL^{+}25] Andong Chen, Zhaoxuan Jin, Zhenyuan Li, Yan Chen, Yu Ning, and Ying Wang. AutoSeg: Automatic microsegmentation policy generation via configuration analysis.

Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002809.

Cho:2021:TBL

[CJS^{+}21] Mingi Cho, Jaedong Jang, Yezee Seo, Seyeon Jeong, Soochang Chung, and Taekyoung Kwon. Towards bidirectional LUT-level detection of hardware Trojans. Computers & Security, 104(??):Article 102223, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100047X.

Chowdhury:2022:MEC

[CKG22] Nabin Chowdhury, Sokratis Katsikas, and Vasileios Gkioulos. Modeling effective cybersecurity training frameworks: a Delphi method-based study. Computers & Security, 113(??): Article 102551, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003758.

Chatzoglou:2023:HGH

[CKK^{+}23] Efstratios Chatzoglou, Vasileios Kouliaridis, Georgios Kambourakis, Georgios Karopoulos, and Stefanos Gritzalis. A hands-on gaze on HTTP/3 security through the lens of HTTP/2 and a public dataset. Computers & Security, 125(??): ??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004436.

Chen:2025:EEA

[CKL^{+}25] Xuefei Chen, Jinfeng Kou, Haiqiang Li, Yuqi Zhang, Junchao Ma, Chen Li, and Bibo Tu. End-to-end anomaly detection of service function chain through multi-source data in cloud-native systems. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001506.

Chavali:2024:PAC

[CKS^{+}24] Lalitha Chavali, Abhinav Krishnan, Paresh Saxena, Barsha Mitra, and Aneesh Sreevallabh Chivukula. Off-policy actor-critic deep reinforcement learning methods for alert prioritization in intrusion detection systems. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400155X.

Chatzoglou:2025:UHC

[CKT^{+}25] Efstratios Chatzoglou, Vyron Kampourakis, Zisis Tsiatsikas, Georgios Karopoulos, and Georgios Kambourakis. Unmasking the hidden credential leaks in password managers and VPN clients. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006047.

Conti:2022:FSM

[CKV22] Mauro Conti, Shubham Khandhar, and P. Vinod. A fewshot malware classification approach for unknown family recognition using malware feature visualization. Computers & Security, 122(??):Article 102887, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002814.

Conti:2020:CBC

[CL20] Mauro Conti and Chhagan Lal. Context-based Co-presence detection techniques: a survey. Computers & Security, 88 (??):Article 101652, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819301968.

Chen:2023:FEM

[CLCZ23] Jinyin Chen, Mingjun Li, Yao Cheng, and Haibin Zheng. FedRight: an effective model copyright protection for federated learning. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004145.

Chafjiri:2024:VDT

[CLHT24] Sadegh Bamohabbat Chafjiri, Phil Legg, Jun Hong, and Michail-Antisthenis Tsompanas. Vulnerability detection through machine learning-based fuzzing: a systematic review. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002050.

Chen:2021:DCC

[CLL^{+}21] Shaojie Chen, Bo Lang, Hongyu Liu, Duokun Li, and Chuan Gao. DNS covert channel detection method using the LSTM model. Computers & Security, 104(??):Article 102095, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303680.

Chen:2024:MTS

[CLL24] Hongsong Chen, Xingyu Li, and Wenmao Liu. Multivariate time series anomaly detection by fusion of deep convolution residual autoencoding reconstruction model and ConvLstm forecasting model. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004911.

Chu:2025:PPW

[CLM^{+}25] Zhiming Chu, Guyue Li, Qingchun Meng, Haobo Li, and Yuwei Zeng. Privacy-preserving WiFi sensing in WSNs via CSI obfuscation. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002834.

Coscia:2025:AAB

[CLMU25] Antonio Coscia, Roberto Lorusso, Antonio Maci, and Giuseppe Urbano. APIARY: an API-based automatic rule generator for YARA to enhance malware detection. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000860.

Canonico:2025:TDR

[CLN^{+}25] Roberto Canonico, Francesco Lista, Annalisa Navarro, Giancarlo Sperlí, and Andrea Vignali. Threat detection in reconfigurable cyber-physical systems through spatio-temporal anomaly detection using graph attention network. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001981.

Chen:2022:WBC

[CLS22] Tianrong Chen, Jie Ling, and Yuping Sun. White-box content camouflage attacks against deep learning. Computers & Security, 117(??):Article 102676, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000748.

Cai:2021:JAM

[CLX21] Lingru Cai, Yao Li, and Zhi Xiong. JOWMDroid: Android malware detection based on feature weighting with joint optimization of weight-mapping and classifier parameters. Computers & Security, 100(??):Article 102086, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030359X.

Charmanas:2023:TIA

[CMA23] Konstantinos Charmanas, Nikolaos Mittas, and Lefteris Angelis. Topic and influence analysis on technological patents related to security vulnerabilities. Computers & Security, 128 (??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300038X.

Crespo-Martinez:2023:SIA

[CMCVGH^{+}23] Ignacio Samuel Crespo-Martínez, Adrián Campazas-Vega, Ángel Manuel Guerrero-Higueras, Virginia Riego-DelCastillo, Claudia Álvarez-Aparicio, and Camino Fernández-Llamas. SQL injection attack detection in network flow data. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0167404823000032.

Case:2020:HAD

[CMFUA^{+}20] Andrew Case, Ryan D. Maggio, Md Firoz-Ul-Amin, Mohammad M. Jalalzai, Aisha Ali-Gombe, Mingxuan Sun, and Golden G. Richard. Hooktracer: Automatic detection and analysis of keystroke loggers using memory forensics. Computers & Security, 96(??):Article 101872, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301450.

Corcoran:2020:DLO

[CMG20] Padraig Corcoran, Peter Mooney, and Andrei Gagarin. A distributed location obfuscation method for online route planning. Computers & Security, 95(??):Article 101850, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740481831071X.

Campoverde-Molina:2025:CSA

[CMLM25] Milton Campoverde-Molina and Sergio Luján-Mora. Cybersecurity in smart agriculture: a systematic literature review. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400590X.

Carrillo-Mondejar:2022:HVA

[CMMST22] J. Carrillo-Mondéjar, J. L. Martinez, and G. Suarez-Tangil. On how VoIP attacks foster the malicious call ecosystem. Computers & Security, 119(??):Article 102758, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001535.

Chiba:2025:DAL

[CNK25] Daiki Chiba, Hiroki Nakano, and Takashi Koide. Domain-Dynamics: Advancing lifecycle-based risk assessment of domain names. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000550.

Cohen:2021:WCD

[CNTBG21] Doron Cohen, Or Naim, Eran Toch, and Irad Ben-Gal. Website categorization via design attribute learning. Computers & Security, 107(??):Article 102312, August 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100136X.

Chua:2021:EDP

[COH21] Hui Na Chua, Jie Sheng Ooi, and Anthony Herbland. The effects of different personal data categories on information privacy concern and disclosure. Computers & Security, 110(??):Article 102453, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002777.

Cojocaru:2025:ARG

[Coj25] Andra Cojocaru. Aligning regulation and governance for cyber resilience: a theoretical framework for the UK Financial Sector. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825003165.

Chrysanthou:2024:ADM

[CPP24] Anargyros Chrysanthou, Yorgos Pantis, and Constantinos Patsakis. The anatomy of deception: Measuring technical and human factors of a large-scale phishing campaign. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000816.

Caturano:2021:DRC

[CPR21] Francesco Caturano, Gaetano Perrone, and Simon Pietro Romano. Discovering reflected cross-site scripting vulnerabilities using a multiobjective reinforcement learning environment. Computers & Security, 103(??):Article 102204, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000286.

Catillo:2021:DRP

[CPRV21] Marta Catillo, Antonio Pecchia, Massimiliano Rak, and Umberto Villano. Demystifying the role of public intrusion datasets: a replication study of DoS network traffic data. Computers & Security, 108(??):Article 102341, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001656.

Chen:2025:EES

[CPS25] Sheng-Shan Chen, Tun-Wen Pai, and Chin-Yu Sun. EnhanceCTI: Enhanced semantic filtering and feature extraction framework for industry-specific cyber threat intelligence. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003384.

Cao:2022:SAV

[CPT^{+}22] Yan Cao, Yuan Ping, Shaohua Tao, YongGang Chen, and YanXia Zhu. Specification and adaptive verification of access control policy for cyber-physical-social spaces. Computers & Security, 114(??):Article 102579, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100403X.

Catillo:2023:CGI

[CPV23] Marta Catillo, Antonio Pecchia, and Umberto Villano. CPS-GUARD: Intrusion detection for cyber-physical systems and IoT devices using outlier-aware deep autoencoders. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001207.

Chen:2024:FPI

[CQC^{+}24] Xiuzhen Chen, Weicheng Qiu, Lixing Chen, Yinghua Ma, and Jin Ma. Fast and practical intrusion detection system based on federated learning for VANET. Computers & Security, 142 (??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001822.

Chen:2025:HDD

[CQC25] Jiann-Liang Chen, Jian-Fu Qiu, and Yu-Hung Chen. A hybrid DGA DefenseNet for detecting DGA domain names based on FastText and deep learning techniques. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005388.

Chen:2021:AVS

[CQL^{+}21] Haowen Chen, Qiang Qu, Yexiong Lin, Xia Chen, and Keqin Li. Authenticity verification on social data outsourcing. Computers & Security, 100(??):Article 102077, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303503.

Cheng:2025:RDR

[CQLW25] Yu Cheng, Xiaofang Qi, Yanhui Li, and Yumeng Wang. ReckDroid: Detecting red packet fraud in Android apps. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400422X.

Cheimonidis:2025:NPD

[CR25] Pavlos Cheimonidis and Konstantinos Rantos. A novel proactive and dynamic cyber risk assessment methodology. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001282.

Charan:2023:DDK

[CRC^{+}23] P. V. Sai Charan, Goutham Ratnakaram, Hrushikesh Chunduri, P. Mohan Anand, and Sandeep Kumar Shukla. DKaaS: DARK-KERNEL as a service for active cyber threat intelligence. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002390.

Carneiro:2025:FIB

[CRdMM^{+}25] Fábio Dias Carneiro, Izabela Simon Rampasso, Sidney Luiz de Matos Mello, Tiago F. A. C. Sigahi, Hernán Lespay, and Rosley Anholon. A flexible ISO 27701-based framework for assessing cybersecurity maturity: a proposition and a case application. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003347.

Chaganti:2022:DLB

[CRP22] Rajasekhar Chaganti, Vinayakumar Ravi, and Tuan D. Pham. Deep learning based cross architecture Internet of Things malware detection and classification. Computers & Security, 120(??):Article 102779, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001742.

Choi:2025:ASF

[CRP25] Dayoung Choi, Joohong Rheey, and Hyunggon Park. Attackspecific feature analysis framework for NetFlow IoT datasets. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002251.

Castelblanco:2022:DFA

[CRS^{+}22] Alejandra Castelblanco, Esteban Rivera, Jesús Solano, Lizzy Tengana, Christian López, and Martín Ochoa. Dynamic face authentication systems: Deep learning verification for camera close-up and head rotation paradigms. Computers & Security, 115(??):Article 102629, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000281.

Canonico:2023:ICP

[CS23] Roberto Canonico and Giancarlo Sperlì. Industrial cyberphysical systems protection: a methodological review. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004418.

Coppa:2024:ESI

[CS24] Emilio Coppa and Sebastian Schrittwieser. Editorial: Special issue on ARES 2022. Computers & Security, 138(??): ??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005722.

Castiglione:2025:GCC

[CSB^{+}25] Gianpietro Castiglione, Daniele Francesco Santamaria, Giampaolo Bella, Laura Brisindi, and Gaetano Puccia. Guiding cybersecurity compliance: an ontology for the NIS 2 directive. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825003062.

Cevallos-Salas:2025:RDM

[CSEJGUA25] David Cevallos-Salas, José Estrada-Jiménez, Danny S. Guamán, and Luis Urquiza-Aguiar. Ransomware dynamics: Mitigating personal data exfiltration through the SCIRAS lens. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482500272X.

Chohra:2022:COF

[CSKD22] Aniss Chohra, Paria Shirani, ElMouatez Billah Karbab, and Mourad Debbabi. Chameleon: Optimized feature selection using particle swarm optimization and ensemble methods for network anomaly detection. Computers & Security, 117(??):Article 102684, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000827.

Choi:2024:SES

[CSLK24] Wonwoo Choi, Minjae Seo, Seongman Lee, and Brent Byunghoon Kang. SuM: Efficient shadow stack protection on ARM Cortex-M. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004789.

Cremer:2024:ICW

[CSM^{+}24] Frank Cremer, Barry Sheehan, Martin Mullins, Michael Fortmann, Barry J. Ryan, and Stefan Materne. On the insurability of cyber warfare: an investigation into the German cyber insurance market. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001871.

Cirne:2022:ISC

[CSRA22] André Cirne, Patrícia R. Sousa, João S. Resende, and Luís Antunes. IoT security certifications: Challenges and potential approaches. Computers & Security, 116(??):Article 102669, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000682.

Celdran:2023:BFD

[CSvdA^{+}23] Alberto Huertas Celdrán, Pedro Miguel Sánchez Sánchez, Jan von der Assen, Dennis Shushack, Ángel Luis Perales Gómez, Gérôme Bovet, Gregorio Martínez Pérez, and Burkhard Stiller. Behavioral fingerprinting to detect ransomware in resourceconstrained devices. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004200.

Chen:2025:KKB

[CSZ^{+}25] Tieming Chen, Qijie Song, Tiantian Zhu, Xuebo Qiu, Zhiling Zhu, and Mingqi Lv. Kellect: a kernel-based efficient and lossless event log collector for windows security. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482400508X.

Cao:2023:PPH

[CSZQ23] Wei Cao, Wenting Shen, Zhixiang Zhang, and Jing Qin. Privacy-preserving healthcare monitoring for IoT devices under edge computing. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003747.

Cai:2025:CME

[CTC^{+}25] Saihua Cai, Han Tang, Jinfu Chen, Yikai Hu, and Wuhao Guo. CDDA-MD: an efficient malicious traffic detection method based on concept drift detection and adaptation technique. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004267.

Chen:2022:XAE

[CTH^{+}22] Li Chen, Cong Tang, Junjiang He, Hui Zhao, Xiaolong Lan, and Tao Li. XSS adversarial example attacks based on deep reinforcement learning. Computers & Security, 120(??):Article 102831, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002255.

Cascavilla:2021:CTI

[CTV21] Giuseppe Cascavilla, Damian A. Tamburri, and Willem-Jan Van Den Heuvel. Cybercrime threat intelligence: a systematic multi-vocal literature review. Computers & Security, 105(??):Article 102258, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821000821.

Cayir:2021:RCF

[ÇÜD21] Aykut Çayir, Ugur Ünal, and Hasan Dag. Random CapsNet forest model for imbalanced malware type classification task. Computers & Security, 102(??):Article 102133, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304065.

Chevrot:2022:CCA

[CVL22] Antoine Chevrot, Alexandre Vernotte, and Bruno Legeard. CAE: Contextual auto-encoder for multivariate timeseries anomaly detection in air transportation. Computers & Security, 116(??):Article 102652, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000517.

Chanda:2025:ACA

[CVZHN25] Razib Chandra Chanda, Ali Vafaei-Zadeh, Haniruzila Hanifah, and Davoud Nikbin. Assessing cybersecurity awareness among bank employees: a multi-stage analytical approach using PLS-SEM, ANN, and fsQCA in a developing country context. Computers & Security, 149(??):??, February 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005145.

Chen:2025:SCC

[CWC^{+}25] Jiasheng Chen, Miao Wang, Zhenfu Cao, Xiaolei Dong, and Liwang Sun. Secure and controllable cloud-edge collaborative data sharing scheme for wireless body area networks in IIoT. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000781.

Chen:2023:DMG

[CWHH23] Shouhong Chen, Tao Wang, Zhentao Huang, and Xingna Hou. Detection method of Golden Chip-Free Hardware Trojan based on the combination of ResNeXt structure and attention mechanism. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003383.

Chen:2022:TNR

[CWL^{+}22] Fei Chen, Jiahao Wang, Jianqiang Li, Yang Xu, Cheng Zhang, and Tao Xiang. TrustBuilder: a non-repudiation scheme for IoT cloud applications. Computers & Security, 116(??):Article 102664, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000633.

Chen:2025:ZEM

[CWQZ25] Zixian Chen, Weiping Wang, Yan Qin, and Shigeng Zhang. ZipAST: Enhancing malicious JavaScript detection with sequence compression. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000793.

Cao:2023:FRF

[CWW^{+}23] Dongliang Cao, Kaimin Wei, Yongdong Wu, Jilian Zhang, Bingwen Feng, and Jinpeng Chen. FePN: a robust feature purification network to defend against adversarial examples. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003371.

Chen:2025:IDI

[CWW^{+}25] Jinfeng Chen, Hua Wu, Xiaohui Wang, Suyue Wang, Guang Cheng, and Xiaoyan Hu. IEA-DMS: an Interpretable feature-driven, Efficient and Accurate Detection Method for Slow HTTP DoS in high-speed networks. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005972.

Chen:2024:GGG

[CWX^{+}24] Zuohui Chen, Renxuan Wang, Jingyang Xiang, Yue Yu, Xin Xia, Shouling Ji, Qi Xuan, and Xiaoniu Yang. GGT: Graph-guided testing for adversarial sample detection of deep neural network. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000117.

Chen:2025:AIR

[CWYW25] Xing Chen, Jingsheng Wang, Song Yan, and Zuyin Wang. Anomalous identity recognition model based on vehicle driving characteristic verification in typical scenarios. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001646.

Chen:2024:HNM

[CWZ^{+}24a] Shi Chen, Wennan Wang, Yubin Zhong, Zuobin Ying, Weixuan Tang, and Zijie Pan. HP-MIA: a novel membership inference attack scheme for high membership prediction precision. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004819.

Chen:2024:MRA

[CWZ^{+}24b] Zigang Chen, Zhen Wang, Yuening Zhou, Fan Liu, Yuhong Liu, Tao Leng, and Haihua Zhu. A method for recovering adversarial samples with both adversarial attack forensics and recognition accuracy. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400292X.

Chen:2023:PPP

[CWZL23] Qian Chen, Zilong Wang, Wenjing Zhang, and Xiaodong Lin. PPT: a privacy-preserving global model training protocol for federated learning in P2P networks. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003583.

Chen:2022:VII

[CXC22] Yan Chen, Weidong Xia, and Karlene Cousins. Voluntary and instrumental information security policy compliance: an integrated view of prosocial motivation, self-regulation and deterrence. Computers & Security, 113(??):Article 102568, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003928.

Chen:2024:GMN

[CXC^{+}24] Jinfu Chen, Haodi Xie, Saihua Cai, Luo Song, Bo Geng, and Wuhao Guo. GCN-MHSA: a novel malicious traffic detection method based on graph convolutional neural network and multi-head self-attention mechanism. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003882.

Cai:2024:MNT

[CXL^{+}24] Saihua Cai, Han Xu, Mingjie Liu, Zhilin Chen, and Guofeng Zhang. A malicious network traffic detection model based on bidirectional temporal convolutional network with multi-head self-attention mechanism. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300490X.

Chen:2025:HLI

[CXL^{+}25] Ziqian Chen, Wei Xia, Zhen Li, Gang Xiong, Gaopeng Gou, Heng Zhang, Haikuo Li, and Junchao Xiao. HoleMal: a lightweight IoT malware detection framework based on efficient host-level traffic processing. Computers & Security, 152 (??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000495.

Chen:2020:CBR

[CXLV20] Yurong Chen, Hongfa Xue, Tian Lan, and Guru Venkataramani. CHOP: Bypassing runtime bounds checking through convex hull optimization. Computers & Security, 90(??): Article 101708, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302457.

Chen:2024:UET

[CYC^{+}24] Yufan Chen, Jiahai Yang, Susu Cui, Cong Dong, Bo Jiang, Yuling Liu, and Zhigang Lu. Unveiling encrypted traffic types through hierarchical network characteristics. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005540.

Chen:2025:PLO

[CYHW25] Zhenyu Chen, Lin Yao, Haibo Hu, and Guowei Wu. Points of the local optimal privacy utility tradeoff. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003116.

Chen:2021:DGB

[CYL^{+}21] Siqi Chen, Yufei Yuan, Xin (Robert) Luo, Jie Jian, and Yan Wang. Discovering group-based transnational cyber fraud actives: a polymethodological view. Computers & Security, 104(??):Article 102217, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000419.

Cheng:2023:VLS

[CYL^{+}23] Yiran Cheng, Shouguo Yang, Zhe Lang, Zhiqiang Shi, and Limin Sun. VERI: a large-scale open-source components vulnerability detection in IoT firmware. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004606.

Cao:2024:CKM

[CYL^{+}24] Yu Cao, Ang Yang, Hanning Li, Qingcheng Zeng, and Jing Gao. A comprehensive knowledge map for AI improving security management of cyber-physical system enabled smart manufacturing. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300559X.

Chang:2024:GBD

[CZ24] Wenhan Chang and Tianqing Zhu. Gradient-based defense methods for data leakage in vertical federated learning. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000452.

Chen:2020:RSN

[CZCX20] Jinyin Chen, Haibin Zheng, Ruoxi Chen, and Hui Xiong. RCA-SOC: a novel adversarial defense by refocusing on critical areas and strengthening object contours. Computers & Security, 96(??):Article 101916, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301929.

Chen:2022:FLT

[CZE^{+}22] Xuyang Chen, Xiaolu Zhang, Michael Elliot, Xiaoyin Wang, and Feng Wang. Fix the leaking tap: a survey of Trigger-Action Programming (TAP) security issues, detection techniques and solutions. Computers & Security, 120(??):Article 102812, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002061.

Cheng:2022:MTA

[CZG^{+}22] Yueqiang Cheng, Zhi Zhang, Yansong Gao, Zhaofeng Chen, Shengjian Guo, Qifei Zhang, Rui Mei, Surya Nepal, and Yang Xiang. Meltdown-type attacks are still feasible in the wall of kernel page-table isolation. Computers & Security, 113(??):Article 102556, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003801.

Chen:2025:HNH

[CZH^{+}25] Zijie Chen, Hailin Zou, Tao Hu, Xun Yuan, Xiaofen Fang, Yuanyuan Pan, and Jianqing Li. HC-NIDS: Historical contextual information based network intrusion detection system in Internet of Things. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000562.

Chidukwani:2024:CPS

[CZK24] Alladean Chidukwani, Sebastian Zander, and Polychronis Koutsakis. Cybersecurity preparedness of small-to-medium businesses: a Western Australia study with broader implications. Computers & Security, 145(??):??, October 2024.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003316.

Chen:2025:AAA

[CZL^{+}25] Minghao Chen, Kaijie Zhu, Bin Lu, Ding Li, Qingjun Yuan, and Yuefei Zhu. AECR: Automatic attack technique intelligence extraction based on fine-tuned large language model. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005194.

Chen:2024:CCT

[CZLZ24] Tieming Chen, Huan Zeng, Mingqi Lv, and Tiantian Zhu. CTIMD: Cyber threat intelligence enhanced malware detection using API call sequences with parameters. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004285.

Chen:2025:VDD

[CZM^{+}25] Zigang Chen, Hongwei Zhang, Qinyu Mu, Danlong Li, and Haihua Zhu. In-vehicle device data tampering detection: Accurate identification based on correlation calculation and data relationship. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003372.

Coulter:2022:DAW

[CZPX22] Rory Coulter, Jun Zhang, Lei Pan, and Yang Xiang. Domain adaptation for Windows advanced persistent threat detection. Computers & Security, 112(??):Article 102496, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003205.

Chen:2021:ODP

[CZS^{+}21] Xin Chen, Tao Zhang, Sheng Shen, Tianqing Zhu, and Ping Xiong. An optimized differential privacy scheme with reinforcement learning in VANET. Computers & Security, 110(??):Article 102446, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002704.

Chang:2024:LPF

[CZSW24] Heyu Chang, Xiaobing Zhang, Nianwen Si, and Ping Wu. A lightweight packet forwarding verification in SDN using sketch. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002086.

Chen:2021:FND

[CZX^{+}21] Jinyin Chen, Haibin Zheng, Hui Xiong, Ruoxi Chen, Tianyu Du, Zhen Hong, and Shouling Ji. FineFool: a novel DNN object contour attack on image recognition based on the attention perturbation adversarial technique. Computers & Security, 104(??):Article 102220, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000444.

Chen:2021:HIS

[CZZ^{+}21] E. Chen, Yan Zhu, Guizhen Zhu, Kaitai Liang, and Rongquan Feng. How to implement secure cloud file sharing using optimized attribute-based access control with small policy matrix and minimized cumulative errors. Computers & Security, 107(??):Article 102318, August 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001425.

Chen:2023:SLD

[CZZ^{+}23] Tieming Chen, Chenbin Zheng, Tiantian Zhu, Chunlin Xiong, Jie Ying, Qixuan Yuan, Wenrui Cheng, and Mingqi Lv. System-level data management for endpoint advanced persistent threat detection: Issues, challenges and trends. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003954.

DaSilva:2022:CSL

[Da 22] Joseph Da Silva. Cyber security and the Leviathan. Computers & Security, 116(??):Article 102674, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000724.

DaSilva:2023:PED

[Da 23] Joseph Da Silva. Protection, expertise and domination: Cyber masculinity in practice. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003188.

Dionysiou:2020:SMV

[DA20] Antreas Dionysiou and Elias Athanasopoulos. SoK: Machine vs. machine — a systematic classification of automated machine learning-based CAPTCHA solvers. Computers & Security, 97(??):Article 101947, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302236.

Davis:2024:FSR

[DAA24] Joshua M. Davis, Deepti Agrawal, and Rebekah Austin. Fostering security-related citizenship through the employeesupervisor relationship: an examination of supervisor security embodiment. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001986.

Rose:2024:VCT

[DAAM24] Luca De Rose, Giuseppina Andresini, Annalisa Appice, and Donato Malerba. VINCENT: Cyber-threat detection through vision transformers and knowledge distillation. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002293.

Djelouat:2020:SCS

[DABB20] Hamza Djelouat, Abbes Amira, Faycal Bensaali, and Issam Boukhennoufa. Secure compressive sensing for ECG monitoring. Computers & Security, 88(??):Article 101649, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301932.

Desouza:2020:WIS

[DANS20] Kevin C. Desouza, Atif Ahmad, Humza Naseer, and Munish Sharma. Weaponizing information systems for political disruption: the Actor, Lever, Effects, and Response Taxonomy (ALERT). Computers & Security, 88(??):Article 101606, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301579.

deAndrade:2023:SAA

[dASJ^{+}23] Ricardo de Andrade, Max Mauro Dias Santos, João Francisco Justo, Leopoldo Rideki Yoshioka, Hans-Joachim Hof, and João Henrique Kleinschmidt. Security architecture for automotive communication networks with CAN FD. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300113X.

Das:2020:ADI

[DAZ20] Tanmoy Kanti Das, Sridhar Adepu, and Jianying Zhou. Anomaly detection in industrial control systems using logical analysis of data. Computers & Security, 96(??):Article 101935, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302121.

Djaidja:2024:FLB

[DBB^{+}24] Taki Eddine Toufik Djaidja, Bouziane Brik, Abdelwahab Boualouache, Sidi Mohammed Senouci, and Yacine Ghamri-Doudane. Federated learning for 5G and beyond, a blessing and a curse — an experimental study on intrusion detection systems. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000087.

Driouch:2024:CIA

[DBG24] Otman Driouch, Slimane Bah, and Zouhair Guennoun. CANSat-IDS: an adaptive distributed Intrusion Detection System for satellites, based on combined classification of CAN traffic. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003389.

Duncan:2023:DNB

[DC23] Shawn P. Duncan and Hui Chen. Detecting network-based internet censorship via latent feature representation learning. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000482.

Dabbagh:2021:SEP

[DCB^{+}21] Mohammad Dabbagh, Kim-Kwang Raymond Choo, Amin Beheshti, Mohammad Tahir, and Nader Sohrabi Safa. A survey of empirical performance evaluation of permissioned blockchain platforms: Challenges and opportunities. Computers & Security, 100(??):Article 102078, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303515.

Bertoli:2023:GID

[dCBJSdS23] Gustavo de Carvalho Bertoli, Lourenço Alves Pereira Junior, Osamu Saotome, and Aldri Luiz dos Santos. Generalizing intrusion detection for heterogeneous networks: a stackedunsupervised federated learning approach. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000160.

Saez-de-Camara:2023:CFL

[dCFA^{+}23] Xabier Sáez de Cámara, Jose Luis Flores, Cristóbal Arellano, Aitor Urbieta, and Urko Zurutuza. Clustered federated learning architecture for network anomaly detection in large scale heterogeneous IoT networks. Computers & Security, 131(??): ??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002092.

Du:2022:FIF

[DCH^{+}22] Xuechao Du, Andong Chen, Boyuan He, Hao Chen, Fan Zhang, and Yan Chen. AflIot: Fuzzing on Linux-based IoT device with binary-level instrumentation. Computers & Security, 122(??):Article 102889, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002838.

Debicha:2023:ABR

[DCK^{+}23] Islam Debicha, Benjamin Cochez, Tayeb Kenaza, Thibault Debatty, Jean-Michel Dricot, and Wim Mees. Adv-Bot: Realistic adversarial botnet attacks against network intrusion detection systems. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300086X.

dAmbrosio:2025:SBS

[dCPR25] Nicola d’Ambrosio, Giulio Capodagli, Gaetano Perrone, and Simon Pietro Romano. SCASS: Breaking into SCADA Systems Security. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000045.

Dave:2022:CSC

[DCS^{+}22] Gaurav Dave, Gaurav Choudhary, Vikas Sihag, Ilsun You, and Kim-Kwang Raymond Choo. Cyber security challenges in aviation communication, navigation, and surveillance. Computers & Security, 112(??):Article 102516, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003400.

Doriguzzi-Corin:2024:FAF

[DCS24] Roberto Doriguzzi-Corin and Domenico Siracusa. FLAD: Adaptive federated learning for DDoS attack detection. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005072.

Dong:2020:EES

[DCSW20] Ye Dong, Xiaojun Chen, Liyan Shen, and Dakui Wang. EaST-FLy: Efficient and secure ternary federated learning. Computers & Security, 94(??):Article 101824, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300985.

Dong:2025:CSC

[DCSY25] Renhai Dong, Baojiang Cui, Yi Sun, and Jun Yang. A combined side-channel and transient execution attack scheme on RISC-V processors. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824006035.

Demirkiran:2022:EPT

[DÇÜD22] Ferhat Demirkiran, Aykut Çayir, Ugur Ünal, and Hasan Dag. An ensemble of pre-trained transformer models for imbalanced multiclass malware classification. Computers & Security, 121(??):Article 102846, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002401.

Desai:2022:BKA

[DDCK22] Nidhi Desai, Manik Lal Das, Payal Chaudhari, and Naveen Kumar. Background knowledge attacks in privacy-preserving data publishing models. Computers & Security, 122(??): Article 102874, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822002681.

Dong:2023:TAD

[DDW^{+}23] Huoyuan Dong, Jialiang Dong, Shaohua Wan, Shuai Yuan, and Zhitao Guan. Transferable adversarial distribution learning: Query-efficient adversarial attack against large language models. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003929.

DeCapitanidiVimercati:2023:DQE

[DFJ^{+}23] Sabrina De Capitani di Vimercati, Sara Foresti, Sushil Jajodia, Giovanni Livraga, Stefano Paraboschi, and Pierangela Samarati. Distributed query execution under access restrictions. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004485.

Dahiya:2020:MAA

[DG20] Amrita Dahiya and B. B. Gupta. Multi attribute auction based incentivized solution against DDoS attacks. Computers & Security, 92(??):Article 101763, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030047X.

Dey:2025:IIF

[DG25] Debasmita Dey and Nirnay Ghosh. iQUIC: an intelligent framework for defending QUIC connection ID-based DoS attack using advantage actor-critic RL. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740482500152X.

Desgeorges:2023:DAN

[DGD23] Loïc Desgeorges, Jean-Philippe Georges, and Thierry Divoux. Detection of anomalies of a non-deterministic software-defined networking control. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001384.

Derbyshire:2021:TDL

[DGH21] Richard Derbyshire, Benjamin Green, and David Hutchison. “Talking a different Language”: Anticipating adversary attack cost for cyber risk assessment. Computers & Security, 103(??):Article 102163, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304363.

Deng:2023:MMC

[DGS^{+}23] Huaxin Deng, Chun Guo, Guowei Shen, Yunhe Cui, and Yuan Ping. MCTVD: a malware classification method based on three-channel visualization and deep learning. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200476X.

Deng:2025:EFM

[DH25] Ping Deng and Yong Huang. Edge-featured multi-hop attention graph neural network for intrusion detection system. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004371.

Du:2023:DCB

[DHL^{+}23] Linfeng Du, Junjiang He, Tao Li, Yunpeng Wang, Xiaolong Lan, and Yunhua Huang. DBWE-Corbat: Background network traffic generation using dynamic word embedding and contrastive learning for cyber range. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001128.

Dai:2024:DFD

[DHW24] Yunwei Dai, Tao Huang, and Shuo Wang. DAmpADF: a framework for DNS amplification attack defense based on Bloom filters and NAmpKeeper. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000191.

Durst:2024:EET

[DHZ24] Susanne Durst, Christoph Hinteregger, and Malgorzata Zieba. The effect of environmental turbulence on cyber security risk management and organizational resilience. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005011.

Dietrich:2023:RYY

[DJG^{+}23] Aljoscha Dietrich, Kurunandan Jain, Georg Gutjahr, Bianca Steffes, and Christoph Sorge. I recognize you by your steps: Privacy impact of pedometer data. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822003868.

See retraction notice [DJG^{+}25].

Dietrich:2025:RNR

[DJG^{+}25] Aljoscha Dietrich, Kurunandan Jain, Georg Gutjahr, Bianca Steffes, and Christoph Sorge. Retraction notice to “I Recognize you by Your Steps: Privacy Impact of Pedometer Data” [Computers & Security 124 (2022) 102994]. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001944. See [DJG^{+}23].

Dam:2025:SAD

[DJKA25] Visal Dam, Fariha Tasmin Jaigirdar, Kallol Krishna Karmakar, and Adnan Anwar. Security-aware data provenance for multi-domain software-defined networks. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003669.

Das:2020:UUC

[DKSS20] Ayan Kumar Das, Sidra Kalam, Nausheen Sahar, and Ditipriya Sinha. UCFL: User categorization using fuzzy logic towards PUF based two-phase authentication of fog assisted IoT devices. Computers & Security, 97(??):Article 101938, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302145.

Diro:2024:ADS

[DKV^{+}24] Abebe Diro, Shahriar Kaisar, Athanasios V. Vasilakos, Adnan Anwar, Araz Nasirian, and Gaddisa Olani. Anomaly detection for space information networks: a survey of challenges, techniques, and future directions. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000063.

Demigha:2021:HBS

[DL21] Oualid Demigha and Ramzi Larguet. Hardware-based solutions for trusted cloud computing. Computers & Security, 103(??):Article 102117, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303904.

Denis:2024:DMU

[DLC24a] Nathanael Denis, Maryline Laurent, and Sophie Chabridon. A decentralized model for usage and information flow control in distributed systems. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002803.

Duan:2024:MMA

[DLC^{+}24b] Guoyun Duan, Haopeng Liu, Minjie Cai, Jianhua Sun, and Hao Chen. MaDroid: a maliciousness-aware multifeatured dataset for detecting Android malware. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002748.

Du:2024:PPQ

[DLL24] Ruizhong Du, Zhuang Liang, and Xiaoyan Liang. Privacypreserving quadratic truth discovery based on precision partitioning. Computers & Security, 146(??):??, November 2024.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003444.

Dai:2021:UIL

[DLQ^{+}21] Yusheng Dai, Hui Li, Yekui Qian, Yunling Guo, Ruipeng Yang, and Min Zheng. Using IRP and local alignment method to detect distributed malware. Computers & Security, 100(??):Article 102109, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303825.

Davies:2021:DAA

[DMB21] Simon R. Davies, Richard Macfarlane, and William J. Buchanan. Differential area analysis for ransomware attack detection within mixed file datasets. Computers & Security, 108(??):Article 102377, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002017.

Dvir:2020:EVT

[DMD^{+}20] Amit Dvir, Angelos K. Marnerides, Ran Dubin, Nehor Golan, and Chen Hajaj. Encrypted video traffic clustering demystified. Computers & Security, 96(??):Article 101917, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301930.

Dong:2023:RFF

[DMM^{+}23] Xingbo Dong, Zhihui Miao, Lan Ma, Jiajun Shen, Zhe Jin, Zhenhua Guo, and Andrew Beng Jin Teoh. Reconstruct face from features based on genetic algorithm using GAN generator as a distribution constraint. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004187.

DaSilva:2025:SSC

[DMPT25] Mike Da Silva, Stéphane Mocanu, Maxime Puys, and Pierre-Henri Thevenon. Safety-security convergence: Automation of IEC 62443-3-2. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001658.

DeSanto:2021:EMS

[DMRV21] D. De Santo, C. S. Malavenda, S. P. Romano, and C. Vecchio. Exploiting the MIL-STD-1553 avionic data bus with an active cyber device. Computers & Security, 100(??):Article 102097, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303709.

Ding:2024:LGS

[DMZ24] Hongli Ding, Zhao Ma, and Jing Zhu. Local graph smoothing for link prediction against universal attack. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002402.

Dhieb:2020:TNB

[DNB^{+}20] Thameur Dhieb, Sourour Njah, Houcine Boubaker, Wael Ouarda, Mounir Ben Ayed, and Adel M. Alimi. Towards a novel biometric system for forensic document examination. Computers & Security, 97(??):Article 101973, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302467.

doNascimento:2021:MSH

[dNPM^{+}21] Pablo Pessoa do Nascimento, Paulo Pereira, Jr., Marco Mialaret, Isac Ferreira, and Paulo Maciel. A methodology for selecting hardware performance counters for supporting nonintrusive diagnostic of flood DDoS attacks on web servers. Computers & Security, 110(??):Article 102434, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002583.

D:2025:SRB

[DP25] Sunitha D. and Latha PH. A secure routing and black hole attack detection system using coot Chimp Optimization Algorithm-based Deep Q network in MANET. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004711.

Diesch:2020:CMI

[DPK20] Rainer Diesch, Matthias Pfaff, and Helmut Krcmar. A comprehensive model of information security factors for decisionmakers. Computers & Security, 92(??):Article 101747, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300341.

Dang-Pham:2022:IIS

[DPKHP22] Duy Dang-Pham, Karlheinz Kautz, Ai-Phuong Hoang, and Siddhi Pittayachawan. Identifying information security opinion leaders in organizations: Insights from the theory of social power bases and social network analysis. Computers & Security, 112(??):Article 102505, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003291.

dAmbrosio:2023:IIT

[dPR23] Nicola d’Ambrosio, Gaetano Perrone, and Simon Pietro Romano. Including insider threats into risk management through Bayesian threat graph networks. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003206.

dAmbrosio:2025:CRO

[dPRU25] Nicola d’Ambrosio, Gaetano Perrone, Simon Pietro Romano, and Alberto Urraro. A cyber-resilient open architecture for drone control. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005108.

Doucek:2020:APS

[DPSN20] Petr Doucek, Lubos Pavlícek, Jirí Sedlácek, and Lea Nedomová. Adaptation of password strength estimators to a non-English environment — the Czech experience. Computers & Security, 95(??):Article 101757, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300419.

Doshi:2023:CDL

[DPSS23] Jay Doshi, Kunal Parmar, Raj Sanghavi, and Narendra Shekokar. A comprehensive dual-layer architecture for phishing and spam email detection. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002882.

Dang-Pham:2025:SIS

[DPTAM25] Duy Dang-Pham, Nik Thompson, Atif Ahmad, and Sean Maynard. Shadow information security practices in organizations: the role of information security transparency, overload, and psychological empowerment. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825002275.

Dileesh:2020:MPV

[DS20] E. D. Dileesh and A. P. Shanthi. M-PIVAD — virtual memory based approach against non-control data attacks. Computers & Security, 95(??):Article 101834, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818311374.

Dileesh:2021:ASD

[DS21] E. D. Dileesh and A. P. Shanthi. An application specific dynamic behaviour model using function-call sequence and memory access-graph for execution integrity verification. Computers & Security, 107(??):Article 102299, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001231.

Dhillon:2022:OOS

[DS22] Parveen Dhillon and Manpreet Singh. An ontology oriented service framework for social IoT. Computers & Security, 122(??):Article 102895, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002887.

Dayal:2023:AEM

[DS23] Neelam Dayal and Shashank Srivastava. Analyzing effective mitigation of DDoS attack with software defined networking. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001797.

Dupont:2023:TCR

[DSBL23] Benoît Dupont, Clifford Shearing, Marilyne Bernier, and Rutger Leukfeldt. The tensions of cyber-resilience: From sensemaking to practice. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002821.

Dorsala:2020:FPV

[DSC20] Mallikarjun Reddy Dorsala, V. N. Sastry, and Sudhakar Chapram. Fair payments for verifiable cloud services using smart contracts. Computers & Security, 90(??):Article 101712, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819302494.

Ding:2025:TTU

[DSD^{+}25] Ruiyang Ding, Lei Sun, Zhiyi Ding, Weifei Zang, and Leyu Dai. Towards targeted and universal adversarial attacks against network traffic classification. Computers & Security, 155(??): ??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001592.

daSilva:2020:HBS

[dSFG20] Carlo Marcelo Revoredo da Silva, Eduardo Luzeiro Feitosa, and Vinicius Cardoso Garcia. Heuristic-based strategy for phishing prediction: a survey of URL-based approach. Computers & Security, 88(??):Article 101613, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301622.

DOnghia:2022:ASD

[DSN^{+}22] Mario D’Onghia, Matteo Salvadore, Benedetto Maria Nespoli, Michele Carminati, Mario Polino, and Stefano Zanero. Apícula: Static detection of API calls in generic streams of bytes. Computers & Security, 119(??):Article 102775, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001705.

Dangl:2023:VEF

[DSR23] Thomas Dangl, Stewart Sentanoe, and Hans P. Reiser. VMIFresh: Efficient and fresh caches for virtual machine introspection. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004376.

Dash:2025:PPU

[DSS^{+}25] Priyabrata Dash, Debasis Samanta, Monalisa Sarma, Ashok Kumar Das, and Athanasios V. Vasilakos. Privacy preserving unique robust and revocable passcode generation from fingerprint data. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003876.

dosSantos:2023:FLR

[dSVST23] Roger R. dos Santos, Eduardo K. Viegas, Altair O. Santin, and Pietro Tedeschi. Federated learning for reliable model updates in network-based intrusion detection. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823003231.

Duy:2021:DDI

[DTK^{+}21] Phan The Duy, Le Khac Tien, Nghi Hoang Khoa, Do Thi Thu Hien, Anh Gia-Tuan Nguyen, and Van-Hau Pham. DIG-FuPAS: Deceive IDS with GAN and function-preserving on adversarial samples in SDN-enabled networks. Computers & Security, 109(??):Article 102367, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001917.

Dong:2022:ITA

[DTT^{+}22] Yansong Dong, Long Tang, Cong Tian, Bin Yu, and Zhenhua Duan. Improving transferability of adversarial examples by saliency distribution and data augmentation. Computers & Security, 120(??):Article 102811, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200205X.

Dubin:2024:CDR

[Dub24] Ran Dubin. Content disarm and reconstruction of Microsoft Office OLE files. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005564.

Datta:2025:FDF

[DUK25] Suvrima Datta, Venkanna U., and Aditya Kotha. FiPi-Box: Development of firewall for IoT networks using P4Pi. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002494.

Dunston:2024:AAI

[DV24] Snofy D. Dunston and Mary Anita Rajam V. AIPA: an adversarial imperceptible patch attack on medical datasets and its interpretability. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004844.

Diaz-Verdejo:2023:CRT

[DVAAM23] Jesús E. Díaz-Verdejo, Rafael Estepa Alonso, Antonio Estepa Alonso, and German Madinabeitia. A critical review of the techniques used for anomaly detection of HTTPbased attacks: taxonomy, limitations and open challenges. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003893.

daVeiga:2020:DOI

[dVABH20] Adéle da Veiga, Liudmila V. Astakhova, Adéle Botha, and Marlien Herselman. Defining organisational information security culture-perspectives from academia and industry. Computers & Security, 92(??):Article 101713, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300018.

Doctor:2020:EDE

[DVJ^{+}20] Jason N. Doctor, Jaideep Vaidya, Xiaoqian Jiang, Shuang Wang, Lisa M. Schilling, Toan Ong, Michael E. Matheny, Lucila Ohno-Machado, and Daniella Meeker. Efficient determination of equivalence for encrypted data. Computers & Security, 97(??):Article 101939, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302157.

Deng:2023:EAC

[DWX^{+}23] Liting Deng, Hui Wen, Mingfeng Xin, Hong Li, Zhiwen Pan, and Limin Sun. Enimanal: Augmented cross-architecture IoT malware analysis using graph neural networks. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300233X.

Deng:2025:CES

[DXZZ25] Qiqing Deng, Zhen Xu, Qihui Zhou, and Yan Zhang. Cordon: Enhancing security through kernel-level control in containerized computing environments. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003335.

Dong:2023:CBB

[DYL^{+}23] Cong Dong, Jiahai Yang, Song Liu, Zhi Wang, Yuling Liu, and Zhigang Lu. C-BEDIM and S-BEDIM: Lateral movement detection in enterprise network through behavior deviation measurement. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001773.

Duan:2024:CFR

[DYW24] Shu-Min Duan, Jian-Ting Yuan, and Bo Wang. Contextual feature representation for image-based insider threat classification. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000804.

Du:2022:MPL

[DZ22] Ruizhong Du and Lin Zhen. Multiuser physical layer security mechanism in the wireless communication system of the IIOT. Computers & Security, 113(??):Article 102559, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003837.

Dixit:2024:PAA

[DZSCR24] Akanksha Dixit, Bruno Bogaz Zarpelao, Max Smith-Creasey, and Muttukrishnan Rajarajan. A privacy-aware authentication and usage-controlled access protocol for IIoT decentralized data marketplace. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003559.

Dang:2025:TMV

[DZSL25] Xiaorui Dang, Guiqi Zhang, Ke Sun, and Yufeng Li. A trust model for VANETs using malicious-aware multiple routing. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004504.

Duan:2021:MGN

[DZZ^{+}21] Yexin Duan, Xingyu Zhou, Junhua Zou, Junyang Qiu, Jin Zhang, and Zhisong Pan. Mask-guided noise restriction adversarial attacks for image classification. Computers & Security, 100(??):Article 102111, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303849.

Duan:2022:AAD

[DZZ^{+}22] Yexin Duan, Junhua Zou, Xingyu Zhou, Wu Zhang, Zhengyun He, Dazhi Zhan, Jin Zhang, and Zhisong Pan. Adversarial attack via dual-stage network erosion. Computers & Security, 122(??):Article 102888, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002826.

Evangelou:2020:ADF

[EA20] Marina Evangelou and Niall M. Adams. An anomaly detection framework for cyber-security data. Computers & Security, 97(??):Article 101941, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302170.

Erola:2022:SCC

[EAN^{+}22] Arnau Erola, Ioannis Agrafiotis, Jason R. C. Nurse, Louise Axon, Michael Goldsmith, and Sadie Creese. A system to calculate cyber value-at-risk. Computers & Security, 113(??):Article 102545, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003692.

Eceiza:2023:IFA

[EFI23] Maialen Eceiza, Jose Luis Flores, and Mikel Iturbe. Improving fuzzing assessment methods through the analysis of metrics and experimental conditions. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003388.

Engstrom:2022:TDC

[EL22] Viktor Engström and Robert Lagerström. Two decades of cyberattack simulations: a systematic literature review. Computers & Security, 116(??):Article 102681, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000797.

ElMestari:2024:PDP

[ELD24] Soumia Zohra El Mestari, Gabriele Lenzini, and Huseyin Demirci. Preserving data privacy in machine learning systems. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005151.

Esposito:2023:DMR

[EMS23] Christian Esposito, Vincenzo Moscato, and Giancarlo Sperlì. Detecting malicious reviews and users affecting social reviewing systems: a survey. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003176.

Erdem:2025:ECS

[EÖ25] Mehmet Erdem and Akin Özdemir. Evaluation of cyber security risk pillars for a digital, innovative, and sustainable model utilizing a novel fuzzy hybrid optimization. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000835.

Ebert:2023:LSS

[ESA^{+}23] Nico Ebert, Thierry Schaltegger, Benjamin Ambuehl, Lorin Schöni, Verena Zimmermann, and Melanie Knieps. Learning from safety science: a way forward for studying cybersecurity incidents in organizations. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003450.

Estay:2020:SRC

[ESBJ20] Daniel A. Sepúlveda Estay, Rishikesh Sahay, Michael B. Barfod, and Christian D. Jensen. A systematic review of cyber-resilience assessment frameworks. Computers & Security, 97(??):Article 101996, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302698.

Esmaeilzadeh:2020:EPP

[Esm20] Pouyan Esmaeilzadeh. The effect of the privacy policy of Health Information Exchange (HIE) on patients’ information disclosure intention. Computers & Security, 95(??): Article 101819, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404818311763.

Ehrenberg:2024:PSC

[ESM24] Melanie Ehrenberg, Shahram Sarkani, and Thomas A. Mazzuchi. Python source code vulnerability detection with named entity recognition. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001032.

El-Sayed:2025:CSR

[ESTA^{+}25] Ameer El-Sayed, Ahmed A. Toony, Fayez Alqahtani, Yasser Alginahi, and Wael Said. CO-STOP: a robust P4-powered adaptive framework for comprehensive detection and mitigation of coordinated and multi-faceted attacks in SD-IoT networks. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000380.

ElBalmany:2022:SVM

[ETAB22] Chawki El Balmany, Zakariae Tbatou, Ahmed Asimi, and Mohamed Bamarouf. Secure virtual machine image storage process into a trusted zone-based cloud storage. Computers & Security, 120(??):Article 102815, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002097.

Elmrabit:2020:ITR

[EYYZ20] Nebrase Elmrabit, Shuang-Hua Yang, Lili Yang, and Huiyu Zhou. Insider threat risk prediction based on Bayesian network. Computers & Security, 96(??):Article 101908, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030184X.

El-Zawawy:2021:VAW

[EZLC21] Mohamed A. El-Zawawy, Eleonora Losiouk, and Mauro Conti. Vulnerabilities in Android webview objects: Still not the end! Computers & Security, 109(??):Article 102395, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002194.

Arroyabe:2023:CCC

[FAAF23] Ignacio Fernandez De Arroyabe, Carlos F. A. Arranz, Marta F. Arroyabe, and Juan Carlos Fernandez de Arroyabe. Cybersecurity capabilities and cyber-attacks as drivers of investment in cybersecurity systems: a UK survey for 2018 and 2019. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003467.

Filiz:2021:ERD

[FACHC21] Burak Filiz, Budi Arief, Orcun Cetin, and Julio Hernandez-Castro. On the effectiveness of ransomware decryption tools. Computers & Security, 111(??):Article 102469, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002935.

Figueira:2020:IIS

[FBL20] Pedro Tubío Figueira, Cristina López Bravo, and José Luis Rivas López. Improving information security risk analysis by including threat-occurrence predictive models. Computers & Security, 88(??):Article 101609, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301592.

Fathi:2025:ESI

[FBR25] Kholood Salah Fathi, Sherif Barakat, and Amira Rezk. An effective SQL injection detection model using LSTM for imbalanced datasets. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740482500080X.

Fernandes:2025:DBF

[FCA25] Pedro Fernandes, Séamus Ó Ciardhuáin, and Mário Antunes. Distance-based feature selection using Benford’s law for malware detection. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003141.

Frenklach:2021:AMD

[FCSP21] Tatiana Frenklach, Dvir Cohen, Asaf Shabtai, and Rami Puzis. Android malware detection via an app similarity graph. Computers & Security, 109(??):Article 102386, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002108.

Filus:2023:SVT

[FD23] Katarzyna Filus and Joanna Domańska. Software vulnerabilities in TensorFlow-based deep learning applications. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003406.

Fu:2022:FCF

[FDK^{+}22] Anmin Fu, Weijia Ding, Boyu Kuang, Qianmu Li, Willy Susilo, and Yuqing Zhang. FH-CFI: Fine-grained hardware-assisted control flow integrity for ARM-based IoT devices. Computers & Security, 116(??):Article 102666, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000657.

Fouladi:2022:NAD

[FEA22] Ramin Fadaei Fouladi, Orhan Ermis, and Emin Anarim. A novel approach for distributed denial of service defense using continuous wavelet transform and convolutional neural network for software-defined network. Computers & Security, 112(??):Article 102524, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003485.

Frauenstein:2020:SPS

[FF20] Edwin Donald Frauenstein and Stephen Flowerday. Susceptibility to phishing on social network sites: a personality information processing model. Computers & Security, 94(??):Article 101862, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301346.

Forsberg:2023:TPM

[FF23] Joonas Forsberg and Tapio Frantti. Technical performance metrics of a security operations center. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300439X.

Friha:2023:FID

[FFB^{+}23] Othmane Friha, Mohamed Amine Ferrag, Mohamed Benbouzid, Tarek Berghout, Burak Kantarci, and Kim-Kwang Raymond Choo. 2DF-IDS: Decentralized and differentially private federated learning-based intrusion detection system for industrial IoT. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300007X.

Fasci:2023:DVB

[FFLQ23] Lara Saidia Fascí, Marco Fisichella, Gianluca Lax, and Chenyi Qian. Disarming visualization-based approaches in malware detection systems. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0167404822004540. See corrigendum

[FFLQ24].

Fasci:2024:CSD

[FFLQ24] Lara Saidia Fascí, Marco Fisichella, Gianluca Lax, and Chenyi Qian. Corrigendum to “Disarming visualizationbased approaches in malware detection systems” [Computers & Security Volume 126, March 2023, 103062]. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002396. See [FFLQ23].

Fernandez-Fuentes:2022:DFA

[FFPC22] Xosé Fernández-Fuentes, Tomás F. Pena, and José C. Cabaleiro. Digital forensic analysis methodology for private browsing: Firefox and Chrome on Linux as a case study. Computers & Security, 115(??):Article 102626, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000256.

Fernandez-Fuentes:2023:DFA

[FFPC23] Xosé Fernández-Fuentes, Tomás F. Pena, and José C. Cabaleiro. Digital forensic analysis of the private mode of browsers on Android. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003358.

Fang:2021:PPC

[FGH^{+}21] Chen Fang, Yuanbo Guo, Yongjin Hu, Bowen Ma, Li Feng, and Anqi Yin. Privacy-preserving and communicationefficient federated learning in Internet of Things. Computers & Security, 103(??):Article 102199, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000237.

Fang:2020:HEF

[FGWJ20] Chen Fang, Yuanbo Guo, Na Wang, and Ankang Ju. Highly efficient federated learning with strong privacy preservation in cloud computing. Computers & Security, 96(??):Article 101889, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301620.

Feng:2024:DDA

[FGY^{+}24] Pengbin Feng, Le Gai, Li Yang, Qin Wang, Teng Li, Ning Xi, and Jianfeng Ma. DawnGNN: Documentation augmented windows malware detection using graph neural network. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000890.

Fang:2020:DMJ

[FHSQ20] Yong Fang, Cheng Huang, Yu Su, and Yaoyao Qiu. Detecting malicious JavaScript code based on semantic analysis. Computers & Security, 93(??):Article 101764, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300481.

Furnell:2022:AAA

[FHW22] Steven Furnell, Kirsi Helkala, and Naomi Woods. Accessible authentication: Assessing the applicability for users with disabilities. Computers & Security, 113(??):Article 102561, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003850.

Fang:2022:JMJ

[FHZ^{+}22] Yong Fang, Chaoyi Huang, Minchuan Zeng, Zhiying Zhao, and Cheng Huang. JStrong: Malicious JavaScript detection based on code semantic representation and graph neural network. Computers & Security, 118(??):Article 102715, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001110.

Frank:2023:UCF

[FJR23] Muriel Frank, Lennart Jaeger, and Lukas Manuel Ranft. Using contextual factors to predict information security overconfidence: a machine learning approach. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004382.

Fernando:2022:FFS

[FK22] Damien Warren Fernando and Nikos Komninos. FeSA: Feature selection architecture for ransomware detection under concept drift. Computers & Security, 116(??):Article 102659, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482200058X.

Frank:2023:UER

[FK23] Muriel Frank and Vanessa Kohn. Understanding extra-role security behaviors: an integration of self-determination theory and construal level theory. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002961.

Fernando:2024:FRD

[FK24] Damien Warren Fernando and Nikos Komninos. FeSAD ransomware detection framework with machine learning using adaption to concept drift. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005394.

Fisk:2023:CCP

[FKL23] Nathan Fisk, Nicholas M. Kelly, and Lori Liebrock. Cybersecurity communities of practice: Strategies for creating gateways to participation. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000986.

Francis:2024:EGD

[FKP^{+}24] Saneesh P. Francis, Vysakh Kani Kolil, Vipin Pavithran, Indrakshi Ray, and Krishnashree Achuthan. Exploring gender dynamics in cybersecurity education: a self-determination theory and social cognitive theory perspective. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002736.

Farhi:2025:PMM

[FKS25] Nitzan Farhi, Noam Koenigstein, and Yuval Shavitt. PatchView: Multi-modality detection of security patches. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825000458.

Franco:2024:REA

[FKvdA^{+}24] Muriel F. Franco, Fabian Künzler, Jan von der Assen, Chao Feng, and Burkhard Stiller. RCVaR: an economic approach to estimate cyberattacks costs using data from industry reports. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000385.

Fried:2021:FAA

[FL21]  Asaf Fried and Mark Last.  Facing airborne attacks

on ADS-B data with autoencoders. Computers & Security, 109(??):Article 102405, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002297.

Feng:2024:HRN

[FLLY24] Wangxin Feng, Xiangyang Luo, Tengyao Li, and Chunfang Yang. HSTW: a robust network flow watermarking method based on hybrid packet sequence-timing. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000026.

Fathalizadeh:2022:PPI

[FMA22] Amir Fathalizadeh, Vahideh Moghtadaiee, and Mina Alishahi. On the privacy protection of indoor location dataset using anonymization. Computers & Security, 117(??):Article 102665, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000645.

Ferretti:2021:SZT

[FMAC21] Luca Ferretti, Federico Magnanini, Mauro Andreolini, and Michele Colajanni. Survivable zero trust for cloud computing environments. Computers & Security, 110(??):Article 102419, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002431.

Frimpong:2022:FFC

[FMU22] Eugene Frimpong, Antonis Michalas, and Amjad Ullah. Footsteps in the fog: Certificateless fog-based access control. Computers & Security, 121(??):Article 102866, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002607.

Fu:2023:GNG

[FNZ^{+}23] Nan Fu, Weiwei Ni, Sen Zhang, Lihe Hou, and Dongyue Zhang. GC-NLDP: a graph clustering algorithm with local differential privacy. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003595.

Flynn:2025:UAM

[FO25] Rory Flynn and Oluwafemi Olukoya. Using approximate matching and machine learning to uncover malicious activity in logs. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482500001X.

Fenech:2024:EPS

[FRF24] Joseph Fenech, Deborah Richards, and Paul Formosa. Ethical principles shaping values-based cybersecurity decisionmaking. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000968.

Freitas:2023:DPS

[FSCM23] Tadeu Freitas, João Soares, Manuel E. Correia, and Rolando Martins. Deterministic or probabilistic? — A survey on Byzantine fault tolerant state machine replication. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823001104.

Fatoki:2024:OAR

[FSMM24] Jimoh G. Fatoki, Zixing Shen, and Carlo A. Mora-Monge. Optimism amid risk: How non-IT employees’ beliefs affect cybersecurity behavior. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001135.

Finder:2022:TIB

[FSN22] Ido Finder, Eitam Sheetrit, and Nir Nissim. A timeinterval-based active learning framework for enhanced PE malware acquisition and detection. Computers & Security, 121(??):Article 102838, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002322.

Furnell:2021:CWS

[Fur21] Steven Furnell. The cybersecurity workforce and skills. Computers & Security, 100(??):Article 102080, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303539.

Furnell:2022:AWP

[Fur22] Steven Furnell. Assessing website password practices — unchanged after fifteen years? Computers & Security, 120(??):Article 102790, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001845.

Farshadkhah:2021:OEA

[FVF21] Sahar Farshadkhah, Craig Van Slyke, and Bryan Fuller. Onlooker effect and affective responses in information security violation mitigation. Computers & Security, 100(??): Article 102082, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303552.

Fujs:2023:BST

[FVV23] Damjan Fujs, Simon Vrhovec, and Damjan Vavpotic. Balancing software and training requirements for information security. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003772.

Fan:2023:VVD

[FWF^{+}23] Yuanhai Fan, Chuanhao Wan, Cai Fu, Lansheng Han, and Hao Xu. VDoTR: Vulnerability detection based on tensor representation of comprehensive code graphs. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823001578.

Farrukh:2024:ANI

[FWKB24] Yasir Ali Farrukh, Syed Wali, Irfan Khan, and Nathaniel D. Bastian. AIS-NIDS: an intelligent and self-sustaining network intrusion detection system. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002876.

Ferguson-Walter:2023:CEF

[FWMJ^{+}23] Kimberly J. Ferguson-Walter, Maxine M. Major, Chelsea K. Johnson, Craig J. Johnson, Dakota D. Scott, Robert S. Gutzwiller, and Temmie Shade. Cyber expert feedback: Experiences, expectations, and opinions about cyber deception. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001785.

Formosa:2021:PFC

[FWR21] Paul Formosa, Michael Wilson, and Deborah Richards. A principlist framework for cybersecurity ethics. Computers & Security, 109(??):Article 102382, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002066.

Fang:2020:CFT

[FXL^{+}20] Junbin Fang, Guikai Xi, Rong Li, Qian Chen, Puxi Lin, Sijin Li, Zoe Lin Jiang, and Siu-Ming Yiu. Coarse-to-fine two-stage semantic video carving approach in digital forensics. Computers & Security, 97(??):Article 101942, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302182.

Fang:2022:DRP

[FXZ22] Rui Fang, Maochao Xu, and Peng Zhao. Determination of ransomware payment based on Bayesian game models. Computers & Security, 116(??):Article 102685, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000839.

Fu:2025:SBA

[FYDH25] Yue Fu, Qingqing Ye, Rong Du, and Haibo Hu. Secure biattribute index: Batch membership tests over the non-sensitive attribute. Computers & Security, 152(??):??, May 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000586.

Fang:2024:FSB

[FYL^{+}24] Yushan Fang, Yu Yao, Xiaoli Lin, Jiaxuan Wang, and Hao Zhai. A feature selection based on genetic algorithm for intrusion detection of industrial control systems. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005850.

Fei:2025:LCL

[FZZ^{+}25] Kexiong Fei, Jiang Zhou, Yucan Zhou, Xiaoyan Gu, Haihui Fan, Bo Li, Weiping Wang, and Yong Chen. LaAeb: a comprehensive log-text analysis based approach for insider threat detection. Computers & Security, 148(??):??, January 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004310.

Giddens:2020:GBI

[GAC20] Laurie Giddens, Laura C. Amo, and Dianna Cichocki. Gender bias and the impact on managerial evaluation of insider security threats. Computers & Security, 99(??):Article 102066, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303394.

Gaber:2025:ZDR

[GAJ25] Matthew Gaber, Mohiuddin Ahmed, and Helge Janicke. Zero day ransomware detection with pulse: Function classification with Transformer models and assembly language. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004723.

Gajrani:2020:EPR

[GAL^{+}20a] Jyoti Gajrani, Umang Agarwal, Vijay Laxmi, Bruhadeshwar Bezawada, Manoj Singh Gaur, Meenakshi Tripathi, and Akka Zemmari. EspyDroid+: Precise reflection analysis of Android apps. Computers & Security, 90(??):Article 101688, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819302251.

Galbally:2020:NFB

[Gal20b] Javier Galbally. A new foe in biometrics: a narrative review of side-channel attacks. Computers & Security, 96(??):Article 101902, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301784.

Ganapathi:2023:RRH

[GAS^{+}23] Iyyakutti Iyappan Ganapathi, Syed Sadaf Ali, Uttam Sharma, Pradeep Tomar, Muhammad Owais, and Naoufel Werghi. RHEMAT: Robust human ear based multimodal authentication technique. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002663.

Giboney:2023:BCC

[GAW^{+}23] Justin Scott Giboney, Bonnie Brinton Anderson, Geoffrey A. Wright, Shayna Oh, Quincy Taylor, Megan Warren, and Kylie Johnson. Barriers to a cybersecurity career: Analysis across career stage and gender. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002262.

Gholami:2025:ADB

[GAZ25] Arash Gholami, Furkan Alaca, and Mohammad Zulkernine. An anomaly detection based approach for continuous authentication with smartwatch inertial sensors. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003451.

Garg:2021:ASA

[GB21] Shivi Garg and Niyati Baliyan. Android security assessment: a review, taxonomy and research gap study. Computers & Security, 100(??):Article 102087, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303606.

Gu:2023:LBB

[GB23] Yuhao Gu and Yuebin Bai. LR-BA: Backdoor attack against vertical federated learning using local latent representations. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001037.

Guven:2022:NPP

[GBA22] Ebu Yusuf Güven, Ali Boyaci, and Muhammed Ali Aydin. A novel password policy focusing on altering user password selection habits: a statistical analysis on breached data. Computers & Security, 113(??):Article 102560, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003849.

Gomez-Barrero:2020:RIS

[GBG20] Marta Gomez-Barrero and Javier Galbally. Reversing the irreversible: a survey on inverse biometrics. Computers & Security, 90(??):Article 101700, March 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302378.

Gale:2022:GCB

[GBS22] Megan Gale, Ivano Bongiovanni, and Sergeja Slapnicar. Governing cybersecurity from the boardroom: Challenges, drivers, and ways ahead. Computers & Security, 121(??):Article 102840, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822002346.

Goswami:2022:RRG

[GC22] Hemangi Goswami and Hiten Choudhury. Remote registration and group authentication of IoT devices in 5G cellular network. Computers & Security, 120(??):Article 102806, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002000.

Goldschmidt:2025:NID

[GC25] Patrik Goldschmidt and Daniela Chudá. Network intrusion datasets: a survey, limitations, and recommendations. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001993.

Guggilam:2024:ESP

[GCN^{+}24] Naga Venkata Rishika Guggilam, Rupa Chiramdasu, Akhil Babu Nambur, Naveena Mikkineni, Yaodong Zhu, and Thippa Reddy Gadekallu. An expert system for privacy-driven vessel detection harnessing YOLOv8 and strengthened by SHA-256. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002049.

Graveto:2022:SBA

[GCS22] Vitor Graveto, Tiago Cruz, and Paulo Simöes. Security of building automation and control systems: Survey and future research directions. Computers & Security, 112(??):Article 102527, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003515.

Guerra:2022:DEC

[GCV22] Jorge Luis Guerra, Carlos Catania, and Eduardo Veas. Datasets are not enough: Challenges in labeling network traffic. Computers & Security, 120(??):Article 102810, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002048.

Gao:2021:GAM

[GCZ21] Han Gao, Shaoyin Cheng, and Weiming Zhang. GDroid: Android malware detection and classification with graph convolutional network. Computers & Security, 106(??):Article 102264, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000882.

Garcia:2021:OFC

[GDG21] David Escudero García and Noemí DeCastro-García. Optimal feature configuration for dynamic malware detection. Computers & Security, 105(??):Article 102250, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000742.

Green:2021:PTA

[GDK^{+}21] Benjamin Green, Richard Derbyshire, Marina Krotofil, William Knowles, Daniel Prince, and Neeraj Suri. PCaaD: Towards automated determination and exploitation of industrial systems. Computers & Security, 110(??):Article 102424, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002480.

Gao:2024:NAM

[GDM^{+}24] Caixia Gao, Yao Du, Fan Ma, Qiuyan Lan, Jianying Chen, and Jingjing Wu. A new adversarial malware detection method based on enhanced lightweight neural network. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003833.

Guo:2022:HHS

[GFH^{+}22] Wenbo Guo, Yong Fang, Cheng Huang, Haoran Ou, Chun Lin, and Yongyan Guo. HyVulDect: a hybrid semantic vulnerability mining system based on graph neural network. Computers & Security, 121(??):Article 102823, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002176.

Gibert:2022:EIN

[GFM^{+}22] Daniel Gibert, Matt Fredrikson, Carles Mateu, Jordi Planes, and Quan Le. Enhancing the insertion of NOP instructions to obfuscate malware via deep reinforcement learning. Computers & Security, 113(??):Article 102543, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003679.

Guo:2021:FEH

[GG21] Yimin Guo and Yajun Guo. FogHA: an efficient handover authentication for mobile devices in fog computing. Computers & Security, 108(??):Article 102358, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001826.

Garg:2022:SLR

[GG22] Manika Garg and Anita Goel. A systematic literature review on online assessment security: Current challenges and integrity strategies. Computers & Security, 113(??):Article 102544, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003680.

Gagnon:2024:RBA

[GGC24] Guillaume Gagnon, Sébastien Gambs, and Mathieu Cunche. RSSI-based attacks for identification of BLE devices. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003857.

Gerdin:2025:CIV

[GGK25] Marcus Gerdin, Åke Grönlund, and Ella Kolkowska. Conceptual inconsistencies in variable definitions and measurement items within ISP non-/compliance research: a systematic literature review. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000549.

Gallo:2024:HFP

[GGR^{+}24] Luigi Gallo, Danilo Gentile, Saverio Ruggiero, Alessio Botta, and Giorgio Ventre. The human factor in phishing: Collecting and analyzing user behavior when reading emails. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005813.

Gui:2025:TPP

[GGZ25] Ruowei Gui, Xiaolin Gui, and Xingjun Zhang. A trajectory privacy protection method based on the replacement of points of interest in hotspot regions. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005856.

Guan:2024:MTS

[GHMG24] Siwei Guan, Zhiwei He, Shenhui Ma, and Mingyu Gao. Multivariate time series anomaly detection with variational autoencoder and spatial-temporal graph network. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001780.

Guggenmos:2022:SFS

[GHOS22] Florian Guggenmos, Björn Häckel, Philipp Ollig, and Bastian Stahl. Security first, security by design, or security pragmatism — strategic roles of IT security in digitalization projects. Computers & Security, 118(??):Article 102747, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001420.

Guo:2024:RVD

[GHZ^{+}24] Longtao Guo, Huakun Huang, Lingjun Zhao, Peiliang Wang, Shan Jiang, and Chunhua Su. Reentrancy vulnerability detection based on graph convolutional networks and expert patterns under subspace mapping. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001962.

Ge:2023:RBM

[GHZS23] Zhaocheng Ge, Hanping Hu, Tengfei Zhao, and Dingmeng Shi. Reading is not believing: a multimodal adversarial attacker for Chinese-NLP model. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004448.

Gadallah:2024:DLT

[GIO24] Waheed G. Gadallah, Hosny M. Ibrahim, and Nagwa M. Omar. A deep learning technique to detect distributed denial of service attacks in software-defined networks. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004984.

Gupta:2022:CIU

[GJB22] Neha Gupta, Vinita Jindal, and Punam Bedi. CSE-IDS: Using cost-sensitive deep learning and ensemble algorithms to handle class imbalance in network-based intrusion detection systems. Computers & Security, 112(??):Article 102499, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003230.

Guo:2020:PPI

[GJCJ20] Cheng Guo, Jing Jia, Kim-Kwang Raymond Choo, and Yingmo Jie. Privacy-preserving image search (PPIS): Secure classification and searching using convolutional neural network over large-scale encrypted medical images. Computers & Security, 99(??):Article 102021, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302947.

Gunes:2021:CSR

[GKB21] Bunyamin Gunes, Gizem Kayisoglu, and Pelin Bolat. Cyber security risk assessment for seaports: a case study of a container port. Computers & Security, 103(??):Article 102196, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000201.

Gupta:2022:STN

[GKC22] Sandeep Gupta, Mouna Kacimi, and Bruno Crispo. Step and turn — a novel bimodal behavioral biometric-based user verification scheme for physical access control. Computers & Security, 118(??):Article 102722, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001171.

Gulmez:2024:XED

[GKS24] Sibel Gulmez, Arzu Gorgulu Kakisim, and Ibrahim Sogukpinar. XRan: Explainable deep learning-based ransomware detection using dynamic analysis. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400004X.

Gu:2021:EID

[GL21] Jie Gu and Shan Lu. An effective intrusion detection approach using SVM with naïve Bayes feature embedding. Computers & Security, 103(??):Article 102158, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304314.

Gernot:2022:BM

[GL22] Tanguy Gernot and Patrick Lacharme. Biometric masterkeys. Computers & Security, 116(??):Article 102642, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000414.

Grosse:2022:BSD

[GLB^{+}22] Kathrin Grosse, Taesung Lee, Battista Biggio, Youngja Park, Michael Backes, and Ian Molloy. Backdoor smoothing: Demystifying backdoor attacks on deep neural networks. Computers & Security, 120(??):Article 102814, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002085.

Guo:2025:HEH

[GLC^{+}25] Zihui Guo, Yin Lv, Ningning Cui, Liwei Chen, and Gang Shi. HScheduler: an execution history-based seed scheduling strategy for hardware fuzzing. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001671.

Guo:2023:FTI

[GLH^{+}23] Yongyan Guo, Zhengyu Liu, Cheng Huang, Nannan Wang, Hai Min, Wenbo Guo, and Jiayong Liu. A framework for threat intelligence extraction and fusion. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300281X.

Gong:2024:NDC

[GLL^{+}24] Zhimao Gong, Junyi Li, Yaping Lin, Lening Yuan, and Wen Gao. A novel dual cloud server privacy-preserving scheme in spatial crowdsourcing. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005680.

Galli:2024:EAB

[GLM^{+}24] Antonio Galli, Valerio La Gatta, Vincenzo Moscato, Marco Postiglione, and Giancarlo Sperlì. Explainability in AI-based behavioral malware detection systems. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001433.

Guo:2021:EES

[GLTH21] Yongyan Guo, Jiayong Liu, Wenwu Tang, and Cheng Huang. Exsense: Extract sensitive information from unstructured data. Computers & Security, 102(??):Article 102156, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304296.

Guan:2023:FIF

[GLX^{+}23] Zhong Guan, Chang Liu, Gang Xiong, Zhen Li, and Gaopeng Gou. FlowTracker: Improved flow correlation attacks with denoising and contrastive learning. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004102.

Guo:2024:PPF

[GLX24] Ping Guo, Wenfeng Liang, and Shuilong Xu. A privacy preserving four-factor authentication protocol for internet of medical things. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005424.

Gao:2022:SSE

[GLZ^{+}22] Yansong Gao, Qun Li, Yifeng Zheng, Guohong Wang, Jiannan Wei, and Mang Su. SEDML: Securely and efficiently harnessing distributed knowledge in machine learning. Computers & Security, 121(??):Article 102857, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002516.

Guerra-Manzanares:2024:MLA

[GM24] Alejandro Guerra-Manzanares. Machine learning for Android malware detection: Mission accomplished? A comprehensive review of open challenges and future perspectives. Computers & Security, 138(??):??, March 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005631.

Guerra-Manzanares:2022:RTI

[GMB22] Alejandro Guerra-Manzanares and Hayretdin Bahsi. On the relativity of time: Implications and challenges of data drift on long-term effective Android malware detection. Computers & Security, 122(??):Article 102835, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002292.

Ghoson:2025:RSD

[GMB^{+}25] Nourhan Halawi Ghoson, Vincent Meyrueis, Khaled Benfriha, Thomas Guiltat, and Stéphane Loubère. A review on the static and dynamic risk assessment methods for OT cybersecurity in Industry 4.0. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824006011.

Guerra-Manzanares:2021:KTB

[GMBN21] Alejandro Guerra-Manzanares, Hayretdin Bahsi, and Sven Nõmm. KronoDroid: Time-based hybrid-featured dataset for effective Android malware detection and characterization. Computers & Security, 110(??):Article 102399, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002236.

Gallo:2021:YAP

[GMBV21] Luigi Gallo, Alessandro Maiello, Alessio Botta, and Giorgio Ventre. 2 years in the anti-phishing group of a large company. Computers & Security, 105(??):Article 102259, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000833.

Gegenhuber:2023:EVM

[GMH^{+}23] Gabriel K. Gegenhuber, Markus Maier, Florian Holzbauer, Wilfried Mayer, Georg Merzdovnik, Edgar Weippl, and Johanna Ullrich. An extended view on measuring tor AS-level adversaries. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002122.

Guerra-Manzanares:2022:CDC

[GMLB22] Alejandro Guerra-Manzanares, Marcin Luckner, and Hayretdin Bahsi. Concept drift and cross-device behavior: Challenges and implications for effective Android malware detection. Computers & Security, 120(??):Article 102757, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001523. See corrigendum

[GMLB23].

Guerra-Manzanares:2023:CCD

[GMLB23] Alejandro Guerra-Manzanares, Marcin Luckner, and Hayretdin Bahsi. Corrigendum to Concept drift and cross-device behavior: Challenges and implications for effective Android malware detection, Computers and Security, Volume 120, 102757. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200390X. See [GMLB22].

Gibert:2020:HMD

[GMP20] Daniel Gibert, Carles Mateu, and Jordi Planes. HYDRA: a multimodal deep learning framework for malware classification. Computers & Security, 95(??):Article 101873, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301462.

Glas:2024:CYA

[GMP24] Magdalena Glas, Gerhard Messmann, and Günther Pernul. Complex yet attainable? An interdisciplinary approach to designing better cyber range exercises. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002700.

Georgiadou:2023:SAC

[GMPA23] Anna Georgiadou, Ariadni Michalitsi-Psarrou, and Dimitris Askounis. A security awareness and competency evaluation in the energy sector. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001098.

Gibert:2021:ASM

[GMPMS21] Daniel Gibert, Carles Mateu, Jordi Planes, and Joao Marques-Silva. Auditing static machine learning antimalware tools against metamorphic attacks. Computers & Security, 102(??):Article 102159, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304326.

Gandal:2023:EEE

[GMRB23] Neil Gandal, Tyler Moore, Michael Riordan, and Noa Barnir. Empirically evaluating the effect of security precautions on cyber incidents. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002900.

Guerreiro:2020:TSE

[GMS20] João Guerreiro, Rui Moura, and João Nuno Silva. TEEnder: SGX enclave migration using HSMs. Computers & Security, 96(??):Article 101874, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301474.

Glasauer:2024:DPS

[GMSA24] Christina Glasauer, Linda Maurer, Carina Spreitzer, and Rainer W. Alexandrowicz. Development and psychometrics of the SOLID-S — an inventory assessing software security culture in software development companies. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000543.

Galal:2022:BKP

[GMY22] Hisham Galal, Mohammad Mannan, and Amr Youssef. Blindfold: Keeping private keys in PKIs and CDNs out of sight. Computers & Security, 118(??):Article 102731, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001262.

Green:2024:RRE

[GOW24] Andrew William Green, DeJarvis Oliver, and Amy B. Woszczynski. To report or not to report? Extending protection motivation theory to vulnerability discovery and disclosure. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001810.

Gil:2025:PPA

[GPAF25] César Gil, Javier Parra-Arnau, and Jordi Forné. Privacy protection against user profiling through optimal data generalization. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004838.

Gallagher:2022:IML

[GPC^{+}22a] Michael Gallagher, Nikolaos Pitropakis, Christos Chrysoulas, Pavlos Papadopoulos, Alexios Mylonas, and Sokratis Katsikas. Investigating machine learning attacks on financial time series models. Computers & Security, 123(??):Article 102933, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200325X.

Galloro:2022:SLS

[GPC^{+}22b] Nicola Galloro, Mario Polino, Michele Carminati, Andrea Continella, and Stefano Zanero. A systematical and longitudinal study of evasive behaviors in windows malware. Computers & Security, 113(??):Article 102550, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003746.

Guo:2023:RRB

[GPD^{+}23] Taolin Guo, Shunshun Peng, Kai Dong, You Zhao, and Mingliang Zhou. RDPCF: Range-based differentially private user data perturbation for collaborative filtering. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003620.

Gudeme:2021:CMR

[GPK21] Jaya Rao Gudeme, Syam Kumar Pasupuleti, and Ramesh Kandukuri. Certificateless multi-replica public integrity auditing scheme for dynamic shared data in cloud storage. Computers & Security, 103(??):Article 102176, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304491.

Garcia-Penas:2025:CIB

[GPRGMF25] Rodolfo García-Peñas, Rafael A. Rodríguez-Gómez, and Gabriel Maciá-Fernández. Characterizing Internet background traffic from a Spain-based network telescope. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003827.

Geden:2023:RLC

[GR23] Munir Geden and Kasper Rasmussen. RegGuard: Leveraging CPU registers for mitigation of control- and dataoriented attacks. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001232.

Gernot:2024:RBS

[GR24] Tanguy Gernot and Christophe Rosenberger. Robust biometric scheme against replay attacks using one-time biometric templates. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004960.

Guaman:2023:AGC

[GRdAS23] Danny S. Guamán, David Rodriguez, Jose M. del Alamo, and Jose Such. Automated GDPR compliance assessment for cross-border personal data transfers in Android applications. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001724.

Gowtham:2024:ASA

[GRJV24] M. S. Gowtham, M. Ramkumar, S. Syed Jamaesha, and M. Vigenesh. Artificial self-attention rabbits battle royale multiscale network based robust and secure data transmission in mobile ad hoc networks. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001913.

Garcia-Rodriguez:2024:PPP

[GRKS24] Jesús García-Rodríguez, Stephan Krenn, and Daniel Slamanig. To pass or not to pass: Privacy-preserving physical access control. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004765.

Gupta:2024:IML

[GS24] Swechchha Gupta and Buddha Singh. An intelligent multilayer framework with SHAP integration for botnet detection and classification. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000841.

Guri:2020:FAD

[GSE20] Mordechai Guri, Yosef Solewicz, and Yuval Elovici. Fansmitter: Acoustic data exfiltration from air-gapped computers via fans noise. Computers & Security, 91(??):Article 101721, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300080.

Giboney:2023:KYE

[GSG23a] Justin Scott Giboney, Ryan M. Schuetzler, and G. Mark Grimes. Know your enemy: Conversational agents for security, education, training, and awareness at scale. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001177.

Govindarajan:2023:FCS

[GSG23b] Usharani Hareesh Govindarajan, Dhiraj Kumar Singh, and Hardik A. Gohel. Forecasting cyber security threats landscape and associated technical trends in telehealth using Bidirectional Encoder Representations from Transformers (BERT). Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003140.

Gu:2023:BSE

[GSK23] Yeming Gu, Hui Shu, and Fei Kang. BinAIV: Semanticenhanced vulnerability detection for Linux x86 binaries. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004182.

Griswold-Steiner:2021:SSP

[GSLS21] Isaac Griswold-Steiner, Zachary LeFevre, and Abdul Serwadda. Smartphone speech privacy concerns from sidechannel attacks on facial biomechanics. Computers & Security, 100(??):Article 102110, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303837.

Gill:2020:GCG

[GSS20a] Komal Singh Gill, Sharad Saxena, and Anju Sharma. GTM-CSec: Game theoretic model for cloud security based on IDS and honeypot. Computers & Security, 92(??):Article 101732, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300195.

Guan:2020:DPG

[GSS^{+}20b] Zhitao Guan, Xianwen Sun, Lingyun Shi, Longfei Wu, and Xiaojiang Du. A differentially private greedy decision forest classification algorithm with high utility. Computers & Security, 96(??):Article 101930, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302054.

G:2024:SAC

[GSSC24] Sahaya Stalin Jose. G., G. Sugitha, Ayshwarya Lakshmi. S., and Preethi Bangalore Chaluvaraj. Self-attention conditional generative adversarial network optimised with crayfish optimization algorithm for improving cyber security in cloud computing. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000749.

Gandhi:2023:LLI

[GSSP23] Jugal Gandhi, Diksha Shekhawat, M. Santosh, and Jai Gopal Pandey. Logic locking for IP security: a comprehensive analysis on challenges, techniques, and trends. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001062.

Gonen:2020:FDI

[GSY^{+}20] Serkan Gönen, H. Hüseyin Sayan, Ercan Nurcan Yilmaz, Furkan Üstünsoy, and Gökçe Karacayilmaz. False data injection attacks and the insider threat in smart systems. Computers & Security, 97(??):Article 101955, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302315.

Gu:2022:MMS

[GSYM22] Yeming Gu, Hui Shu, Pan Yang, and Rongkuan Ma. Min-SIB: Minimized static instrumentation for fuzzing binaries. Computers & Security, 122(??):Article 102894, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002875.

Garg:2024:PMC

[GT24a] Sonakshi Garg and Vicenç Torra. Privacy in manifolds: Combining k-anonymity with differential privacy on Fréchet means. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002888.

Gowrisankar:2024:AAA

[GT24b] Balachandar Gowrisankar and Vrizlynn L. L. Thing. An adversarial attack approach for eXplainable AI evaluation on deepfake detection models. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005941.

Goyal:2025:MBI

[GTAK25] Nandita Goyal, Kanika Taneja, Shivani Agarwal, and Harsh Khatter. Malicious behavior identification using dual attention based dense bi-directional gated recurrent network in the cloud computing environment. Computers & Security, 154(??): ??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001075.

Garcia-Teodoro:2022:MLC

[GTGHAG22] P. García-Teodoro, J. A. Gómez-Hernández, and A. Abellán-Galera. Multi-labeling of complex, multi-behavioral malware samples. Computers & Security, 121(??):Article 102845, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002395.

Gu:2022:PAM

[GTL^{+}22] Xiuting Gu, Zhu Tianqing, Jie Li, Tao Zhang, Wei Ren, and Kim-Kwang Raymond Choo. Privacy, accuracy, and model fairness trade-offs in federated learning. Computers & Security, 122(??):Article 102907, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003005.

Gibert:2025:AIP

[GTP^{+}25] Daniel Gibert, Nikolaos Totosis, Constantinos Patsakis, Quan Le, and Giulio Zizzo. Assessing the impact of packing on static machine learning-based malware detection and classification systems. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740482500183X.

Guri:2025:PWP

[Gur25] Mordechai Guri. PIXHELL: When pixels learn to scream. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002573.

Gouveia:2022:BLL

[GVEV22] Inês Pinto Gouveia, Marcus Völp, and Paulo Esteves-Verissimo. Behind the last line of defense: Surviving SoC faults and intrusions. Computers & Security, 123(??):Article 102920, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003121.

Goel:2023:CRN

[GVJ23] Shefali Goel, Abhishek Verma, and Vinod Kumar Jain. CRA-RPL: a novel lightweight challenge-response authenticationbased technique for securing RPL against dropped DAO attacks. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002560.

Gwebu:2024:DFR

[GW24] Kholekile L. Gwebu and Jing Wang. The defining features of a robust information security climate. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001937.

Geng:2024:SSD

[GWF^{+}24] Jiaxuan Geng, Junfeng Wang, Zhiyang Fang, Yingjie Zhou, Di Wu, and Wenhan Ge. A survey of strategy-driven evasion methods for PE malware: Transformation, concealment, and attack. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005059.

Ge:2023:ECT

[GWL^{+}23] Wenhan Ge, Junfeng Wang, Tongcan Lin, Binhui Tang, and Xiaohui Li. Explainable cyber threat behavior identification based on self-adversarial topic generation. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002791.

Guo:2025:FFU

[GWY^{+}25] Yunchuan Guo, Xiao Wang, Mingjie Yu, Fenghua Li, Zhen Pang, and Liang Fang. An on-the-fly framework for usable access control policy mining. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005170.

Guo:2025:CAA

[GXL25] Ping Guo, Shuilong Xu, and Wenfeng Liang. A cloud-assisted anonymous and privacy-preserving authentication scheme for Internet of Medical Things. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825003037.

Guo:2023:FSB

[GXZ23] Zhenyuan Guo, Lei Xu, and Liehuang Zhu. FedSIGN: a sign-based federated learning framework with privacy and robustness guarantees. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300384X.

Gao:2023:SSP

[GY23] Chen Gao and Jia Yu. SecureRC: a system for privacypreserving relation classification using secure multi-party computation. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000524.

Gupta:2023:NDP

[GYG^{+}23] Prajjwal Gupta, Krishna Yadav, Brij B. Gupta, Mamoun Alazab, and Thippa Reddy Gadekallu. A novel data poisoning attack in federated learning based on inverted loss function. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001803.

Ge:2020:PPB

[GYL^{+}20] Chunpeng Ge, Changchun Yin, Zhe Liu, Liming Fang, Juncen Zhu, and Huading Ling. A privacy preserve big data analysis system for wearable wireless sensor network. Computers & Security, 96(??):Article 101887, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301607.

Guo:2024:NBA

[GYL^{+}24] Chunjie Guo, Lin You, Xingyu Li, Gengran Hu, Shengguo Wang, and Chengtang Cao. A novel biometric authentication scheme with privacy protection based on SVM and ZKP. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003006.

Guo:2023:TMS

[GYW^{+}23] Pengfei Guo, Yingjian Yan, Junjie Wang, Jingxin Zhong, Yanjiang Liu, and Jinsong Xu. Towards a metrics suite for evaluating cache side-channel vulnerability: Case studies on an opensource RISC-V processor. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003905.

Guo:2022:PPC

[GYYK22] Li Guo, Jia Yu, Ming Yang, and Fanyu Kong. Privacypreserving convolution neural network inference with edgeassistance. Computers & Security, 123(??):Article 102910, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003030.

Guo:2023:ECS

[GYZ^{+}23] Pengfei Guo, Yingjian Yan, Fan Zhang, Chunsheng Zhu, Lichao Zhang, and Zibin Dai. Extending the classical side-channel analysis framework to access-driven cache attacks. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001657.

Gao:2025:IIN

[GYZ^{+}25a] Yazhuo Gao, Lin Yang, Ran Zhu, Yixuan Wu, Feng Yang, and Yining Cao. IR-IDS: a network intrusion detection method based on causal feature selection and explainable model optimization. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001841.

Gong:2025:FSM

[GYZ^{+}25b] XingYu Gong, Yi Yang, Yi Zhang, Na Li, Yu Guan, and RongKun Jiang. Feature selection method for network intrusion based on hybrid meta-heuristic dynamic optimization algorithm. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002019.

Guo:2018:LLP

[GZ18] Yimin Guo and Zhenfeng Zhang. LPSE: Lightweight passwordstrength estimation for password meters. Computers & Security, 73(??):507–518, March 2018. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404817301530.

See corrigendum [GZ20].

Guo:2020:CLL

[GZ20] Yimin Guo and Zhenfeng Zhang. Corrigendum to “LPSE: Lightweight password-strength estimation for password meters” [Computers and Security, Volume 73, 2018, Pages 507– 518]. Computers & Security, 94(??):Article 101879, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301528. See [GZ18].

Goncalves:2024:DAL

[GZ24] Luís Gonçalves and Cleber Zanchettin. Detecting abnormal logins by discovering anomalous links via graph transformers. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002499.

Guo:2019:ONP

[GZG19] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. Optiwords: a new password policy for creating memorable and strong passwords. Computers & Security, 85(??):423–435, August 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301105. See [GZG20].

Guo:2020:CON

[GZG20] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. Corrigendum to “Optiwords: a new password policy for creating memorable and strong password” [Computers and Security, Volume 85, 2019, Pages 423–435]. Computers & Security, 94(??):Article 101880, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030153X. See [GZG19].

Guo:2021:SHS

[GZG21] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. Superword: a honeyword system for achieving higher security goals. Computers & Security, 103(??):Article 101689, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302263.

Guo:2020:NPP

[GZGG20] Yimin Guo, Zhenfeng Zhang, Yajun Guo, and Xiaowei Guo. Nudging personalized password policies by understanding users’ personality. Computers & Security, 94(??):Article 101801, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300870.

Gu:2024:GGB

[GZH^{+}24] Jintao Gu, Hongliang Zhu, Zewei Han, Xiangyu Li, and Jianjin Zhao. GSEDroid: GNN-based Android malware detection framework using lightweight semantic embedding. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001081.

Garn:2022:TST

[GZS^{+}22] Bernhard Garn, Stefan Zauner, Dimitris E. Simos, Manuel Leithner, Richard Kuhn, and Raghu Kacker. A two-step TLS-Based browser fingerprinting approach using combinatorial sequences. Computers & Security, 114(??):Article 102575, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003990.

Hayat:2024:FSN

[HABW24] Amber Hayat, Syed Sadaf Ali, Ashok Kumar Bhateja, and Naoufel Werghi. FinTem: a secure and non-invertible technique for fingerprint template protection. Computers & Security, 142 (??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001779.

Hajar:2021:SWB

[HAKK21] Muhammad Shadi Hajar, M. Omar Al-Kadri, and Harsha Kumara Kalutarage. A survey on wireless body area networks: architecture, security challenges and research opportunities. Computers & Security, 104(??):Article 102211, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000353.

Henriques:2022:ACL

[HCCS22] João Henriques, Filipe Caldeira, Tiago Cruz, and Paulo Simões. An automated closed-loop framework to enforce security policies from anomaly detection. Computers & Security, 123(??):Article 102949, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003418.

Han:2023:NID

[HCL^{+}23] Xueying Han, Susu Cui, Song Liu, Chen Zhang, Bo Jiang, and Zhigang Lu. Network intrusion detection based on ngram frequency and time-aware transformer. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000810.

Hernandez-Castro:2020:AAU

[HCLRM20] Carlos Javier Hernández-Castro, Shujun Li, and María D. R-Moreno. All about uncertainties and traps: Statistical oraclebased attacks on a new CAPTCHA protection against oracle attacks. Computers & Security, 92(??):Article 101758, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300420.

He:2025:SNG

[HD25] Chenggang He and Chris H. Q. Ding. SecRASP: Next generation web application security protection methodology and framework. Computers & Security, 154(??):??, July 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001348.

Hariyanti:2021:ISV

[HDS21] Eva Hariyanti, Arif Djunaidy, and Daniel Siahaan. Information security vulnerability prediction based on business process model using machine learning approach. Computers & Security, 110(??):Article 102422, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002467.

He:2022:BAA

[HDZ^{+}22] Zhengyun He, Yexin Duan, Wu Zhang, Junhua Zou, Zhengfang He, Yunyun Wang, and Zhisong Pan. Boosting adversarial attacks with transformed gradient. Computers & Security, 118(??):Article 102720, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001158.

He:2025:PLI

[He25] Ting He. Physical-layer identity-authentication mechanism for network time synchronisation using network and precision time protocols. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002792.

Hettema:2021:RCC

[Het21] Hinne Hettema. Rationality constraints in cyber defense: Incident handling, attribution and cyber threat intelligence. Computers & Security, 109(??):Article 102396, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002200.

Higgs:2025:DCO

[HF25] James Higgs and Stephen Flowerday. Detecting cybercrime in online video gaming. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002172.

Hamad:2024:RAI

[HFK^{+}24] Mohammad Hamad, Andreas Finkenzeller, Michael Kühr, Andrew Roberts, Olaf Maennel, Vassilis Prevelakis, and Sebastian Steinhorst. REACT: Autonomous intrusion response system for intelligent vehicles. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003134.

Han:2023:FDI

[HFLZ23] Yinghua Han, Hantong Feng, Keke Li, and Qiang Zhao. False data injection attacks detection with modified temporal multi-graph convolutional network in smart grids. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004084.

House:2020:PSD

[HG20] Deanna House and Gabe Giordano. Politeness in security directives: Insights in browser compliance for the human element. Computers & Security, 99(??):Article 102007, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302807.

Ho:2021:CCD

[HG21] Shuyuan Mary Ho and Melissa Gross. Consciousness of cyber defense: a collective activity system for developing organizational cyber awareness. Computers & Security, 108(??):Article 102357, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001814.

Hore:2024:SDL

[HGSB24] Soumyadeep Hore, Jalal Ghadermazi, Ankit Shah, and Nathaniel D. Bastian. A sequential deep learning framework for a robust and resilient network intrusion detection system. Computers & Security, 144(??):??, September 2024.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002311.

Hammad:2022:MRN

[HHE22] Mohamed Hammad, Nabil Hewahi, and Wael Elmedany. MMM-RF: a novel high accuracy multinomial mixture model for network intrusion detection systems. Computers & Security, 120(??):Article 102777, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001729.

Hu:2022:TPP

[HHL^{+}22] Yun Hu, Aiqun Hu, Chunguo Li, Peng Li, and Chunyu Zhang. Towards a privacy protection-capable noise fingerprinting for numerically aggregated data. Computers & Security, 119(??):Article 102755, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200150X.

Han:2024:BCL

[HHL24] Jiaxuan Han, Cheng Huang, and Jiayong Liu. bjCnet: a contrastive learning-based framework for software defect prediction. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003298.

Hoffman:2024:PNH

[HHP^{+}24] Cameron John Hoffman, C. Jordan Howell, Robert C. Perkins, David Maimon, and Olena Antonaccio. Predicting new hackers’ criminal careers: a group-based trajectory approach. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005588.

Hart:2022:CSG

[HHS22] Stephen Hart, Basel Halak, and Vladimiro Sassone. CIST: a serious game for hardware supply chain. Computers & Security, 122(??):Article 102912, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003054.

Han:2020:CTC

[HHSL20] Jiaxuan Han, Cheng Huang, Fan Shi, and Jiayong Liu. Covert timing channel detection method based on time interval and payload length analysis. Computers & Security, 97(??):Article 101952, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302285.

Hillman:2023:EOP

[HHT23] Doron Hillman, Yaniv Harel, and Eran Toch. Evaluating organizational phishing awareness training on an enterprise scale. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002742.

Hu:2021:ISE

[HHZ21] Siqi Hu, Carol Hsu, and Zhongyun Zhou. The impact of SETA event attributes on employees’ security-related intentions: an event system theory perspective. Computers & Security, 109(??):Article 102404, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002285.

Hirano:2025:ROD

[HK25] Manabu Hirano and Ryotaro Kobayashi. RanSMAP: Open dataset of Ransomware Storage and Memory Access Patterns for creating deep learning based ransomware detectors. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005078.

He:2024:NVI

[HKA24] Ke He, Dan Dongseong Kim, and Muhammad Rizwan Asghar. NIDS-Vis: Improving the generalized adversarial robustness of network intrusion detection system. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482400333X.

Ho:2022:SCP

[HKM22] Heemeng Ho, Ryan Ko, and Lorraine Mazerolle. Situational Crime Prevention (SCP) techniques to prevent and control cybercrimes: a focused systematic review. Computers & Security, 115(??):Article 102611, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000104.

Hengstler:2023:SRD

[HKM^{+}23] Sebastian Hengstler, Stephan Kuehnel, Kristin Masuch, Ilja Nastjuk, and Simon Trang. Should i really do that? Using quantile regression to examine the impact of sanctions on information security policy compliance behavior. Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002808.

Hussain:2025:LPC

[HKY^{+}25] Tariq Hussain, Muhammad Nawaz Khan, Bailin Yang, Razaz Waheeb Attar, and Ahmed Alhomoud. LiDAR point cloud transmission: Adversarial perspectives of spoofing attacks in autonomous driving. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002330.

Han:2024:BBB

[HLC^{+}24] Xu Han, Qiang Li, Hongbo Cao, Lei Han, Bin Wang, Xuhua Bao, Yufei Han, and Wei Wang. BFS2Adv: Blackbox adversarial attack towards hard-to-attack short texts. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001184.

Hoglund:2020:PTP

[HLFR20] Joel Höglund, Samuel Lindemer, Martin Furuhed, and Shahid Raza. PKI4IoT: Towards public key infrastructure for the Internet of Things. Computers & Security, 89(??):Article 101658, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819302019.

Lin:2021:PBP

[hLHhLfW21] Pei hong Lin, Zheng Hong, Yi hao Li, and Li fa Wu. A priority based path searching method for improving hybrid fuzzing. Computers & Security, 105(??):Article 102242, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000663.

He:2021:IBR

[HLL^{+}21] Junjiang He, Tao Li, Beibei Li, Xiaolong Lan, Zhiyong Li, and Yunpeng Wang. An immune-based risk assessment method for digital virtual assets. Computers & Security, 102(??):Article 102134, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304077.

Hu:2024:IDM

[HLL24] Haibin Hu, Renxin Liu, and Rilong Liu. An intelligent diagnostic model for industrial equipment with privacy protection. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003419.

He:2025:TTP

[HLL^{+}25] Xianliang He, Junyi Li, Yaping Lin, Qiao Hu, and Xiehua Li. TDSF: Trajectory-preserving method of dual-strategy fusion with differential privacy in LBS. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003864.

Hao:2022:EMM

[HLP22] Jingwei Hao, Senlin Luo, and Limin Pan. EII-MBS: Malware family classification via enhanced adversarial instruction behavior semantic learning. Computers & Security, 122(??):Article 102905, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200298X.

He:2023:ITA

[HLQD23] Xianglong He, Yuezun Li, Haipeng Qu, and Junyu Dong. Improving transferable adversarial attack via feature-momentum. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000457.

Hu:2024:CGF

[HLX^{+}24] Yanfei Hu, Shuai Li, Wenchao Xue, Yanlong Zhao, and Yu Wen. CarePlus: a general framework for hardware performance counter based malware detection under system resource competition. Computers & Security, 143(??):??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001858.

Huang:2021:ISM

[HLYZ21] Shize Huang, Xiaowen Liu, Xiaolu Yang, and Zhaoxin Zhang. An improved ShapeShifter method of generating adversarial examples for physical attacks on stop signs against faster R-CNNs. Computers & Security, 104(??):Article 102120, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030393X.

Huang:2021:JMJ

[HLZ^{+}21] Yunhua Huang, Tao Li, Lijia Zhang, Beibei Li, and Xiaojie Liu. JSContana: Malicious JavaScript detection using adaptable context analysis and key feature extraction. Computers & Security, 104(??):Article 102218, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000420.

Han:2025:DDM

[HLZ^{+}25] Yufei Han, Chao Li, Jianbiao Zhang, Yifan Wang, Lehao Yu, Yihao Cao, Hong Shen, Weixing Hou, and Hailin Luo. DMSCTS: Dynamic measurement scheme for the containers-hybrid-deployment based on trusted subsystem. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004632.

Huang:2023:III

[HM23] Yunfan Huang and Maode Ma. ILL-IDS: an incremental lifetime learning IDS for VANETs. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822003844.

Harsha:2021:BAC

[HMB^{+}21] Benjamin Harsha, Robert Morton, Jeremiah Blocki, John Springer, and Melissa Dark. Bicycle attacks considered harmful: Quantifying the damage of widespread password length leakage. Computers & Security, 100(??):Article 102068, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303412.

Han:2023:ESR

[HMBY23] Kyusuk Han, Hyeran Mun, Malavika Balakrishnan, and Chan Yeob Yeun. Enhancing security and robustness of Cyphal on Controller Area Network in unmanned aerial vehicle environments. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003917.

Haider:2020:FHF

[HMK^{+}20] Waqas Haider, Nour Moustafa, Marwa Keshk, Amanda Fernandez, Kim-Kwang Raymond Choo, and Abdul Wahab. FGMC-HADS: Fuzzy Gaussian mixture-based correntropy models for detecting zero-day attacks from Linux systems. Computers & Security, 96(??):Article 101906, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301826.

Haseeb:2022:PMD

[HMMW22] Junaid Haseeb, Saif Ur Rehman Malik, Masood Mansoori, and Ian Welch. Probabilistic modelling of deception-based security framework using Markov decision process. Computers & Security, 115(??):Article 102599, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0167404821004223. See corrigendum [HuRMMW22].

Hart:2020:RSG

[HMPS20] Stephen Hart, Andrea Margheri, Federica Paci, and Vladimiro Sassone. Riskio: a serious game for cyber security awareness and education. Computers & Security, 95(??):Article 101827, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301012.

Harush:2021:DAB

[HMS21] Shimon Harush, Yair Meidan, and Asaf Shabtai. DeepStream: Autoencoder-based stream temporal clustering and anomaly detection. Computers & Security, 106(??):Article 102276, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001000.

Hendaoui:2025:FFL

[HMT^{+}25] Fatma Hendaoui, Rahma Meddeb, Lamia Trabelsi, Ahlem Ferchichi, and Rawia Ahmed. FLADEN: Federated learning for anomaly DEtection in IoT networks. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740482500135X.

Hassanin:2022:RMM

[HMTC22] Mohammed Hassanin, Nour Moustafa, Murat Tahtali, and Kim-Kwang Raymond Choo. Rethinking maximum-margin softmax for adversarial robustness. Computers & Security, 116(??):Article 102640, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000396.

Hnaini:2025:ESR

[HMV^{+}25] Hiba Hnaini, Raúl Mazo, Paola Vallejo, Andrés López, and Joël Champeau. Enhancing security requirements specification with SECRET-SCORE: a template-driven and ontologybased approach. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002949.

Halgamuge:2025:AES

[HN25] Malka N. Halgamuge and Dusit Niyato. Adaptive edge security framework for dynamic IoT security policies in diverse environments. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004334.

Hamid:2022:BEA

[HNH^{+}22] Nor Hasnul Azirah Abdul Hamid, Normalina Ibrahim @ Mat Nor, Fazlin Marini Hussain, Rajeswari Raju, Humza Naseer, and Atif Ahmad. Barriers and enablers to adoption of cyber insurance in developing countries: an exploratory study of Malaysian organizations. Computers & Security, 122(??):Article 102893, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002863.

Hnamte:2024:DAD

[HNNN^{+}24] Vanlalruata Hnamte, Ashfaq Ahmad Najar, Hong Nhung-Nguyen, Jamal Hussain, and Manohar Naik Sugali. DDoS attack detection and mitigation using deep neural network in SDN environment. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005710.

Hou:2023:WSC

[HNZ^{+}23] Lihe Hou, Weiwei Ni, Sen Zhang, Nan Fu, and Dongyue Zhang. Wdt-SCAN: Clustering decentralized social graphs with local differential privacy. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200428X.

Hodges:2021:CEB

[Hod21] Duncan Hodges. Cyber-enabled burglary of smart homes. Computers & Security, 110(??):Article 102418, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482100242X.

Henry:2020:SPS

[HP20] Wayne C. Henry and Gilbert L. Peterson. SensorRE: Provenance support for software reverse engineers. Computers & Security, 95(??):Article 101865, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301371.

Harborth:2023:PCM

[HP23] David Harborth and Sebastian Pape. A privacy calculus model for contact tracing apps: Analyzing the use behavior of the German Corona-Warn-App with a longitudinal user study. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002481.

Park:2023:FIF

[hPRPJ23] Ah hyun Park, Hyejin Ryu, Woobeen Park, and Doowon Jeong. Forensic investigation framework for cryptocurrency wallet in the end device. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003024.

Huang:2022:BCT

[HQL^{+}22] Yizhao Huang, Meng Qiao, Fudong Liu, Xingwei Li, Hairen Gui, and Chunyan Zhang. Binary code traceability of multigranularity information fusion from the perspective of software genes. Computers & Security, 114(??):Article 102607, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000062.

Hoong:2024:NCG

[HR24] Yang Hoong and Davar Rezania. Navigating Cybersecurity Governance: the influence of opportunity structures in socio-technical transitions for small and medium enterprises. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001536.

Hallaji:2024:EAC

[HRFS24] Ehsan Hallaji, Roozbeh Razavi-Far, and Mehrdad Saif. Expanding analytical capabilities in intrusion detection through ensemble-based multi-label classification. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000312.

Haque:2024:FTA

[HRU24] Nur Imtiazul Haque, Mohammad Ashiqur Rahman, and Selcuk Uluagac. Formal threat analysis of machine learningbased control systems: a study on smart healthcare systems. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000105.

Holm:2025:RBA

[HS25a] Hannes Holm and Teodor Sommestad. Realistic and balanced automated threat emulation. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000409.

Hwang:2025:MSS

[HS25b] Inho Hwang and Ribin Seo. Mitigating security stress: Exploring the contingent role of collaborative communication in enhancing information security compliance. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740482500015X.

Himeur:2022:LTS

[HSB^{+}22] Yassine Himeur, Shahab Saquib Sohail, Faycal Bensaali, Abbes Amira, and Mamoun Alazab. Latest trends of security and privacy in recommender systems: a comprehensive review and future perspectives. Computers & Security, 118(??):Article 102746, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001419.

Huang:2024:TOA

[HSH24] Haiou Huang, Bangyi Sun, and Liang Hu. A task offloading approach based on risk assessment to mitigate edge DDoS attacks. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000907.

Huang:2022:DMD

[HSK22] Yuyao Huang, Hui Shu, and Fei Kang. DeMal: Module decomposition of malware based on community discovery. Computers & Security, 117(??):Article 102680, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000785.

Hu:2020:ENU

[HSL20] Yang Hu, Mingshen Sun, and John C. S. Lui. Exploiting nonuniform program execution time to evade record/replay forensic analysis. Computers & Security, 88(??):Article 101516, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819300938.

Hull:2023:TMS

[HSL23] David Michael Hull, Sebastian Walter Schuetz, and Paul Benjamin Lowry. Tell me a story: the effects that narratives exert on meaningful-engagement outcomes in antiphishing training. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001621.

Heiding:2023:PTC

[HSOL23] Fredrik Heiding, Emre Süren, Johannes Olegård, and Robert Lagerström. Penetration testing of connected households. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200459X.

Huang:2024:CCS

[HSQ24] Yinghui Huang, Wenting Shen, and Jing Qin. Certificateless cloud storage auditing supporting data ownership transfer. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000397.

Huang:2023:PPC

[HSQH23] Yinghui Huang, Wenting Shen, Jing Qin, and Huiying Hou. Privacy-preserving certificateless public auditing supporting different auditing frequencies. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000913.

Husak:2022:CTC

[HSŠ^{+}22] Martin Husák, Lukáš Sadlek, Stanislav Špaček, Martin Laštovička, Michal Javorník, and Jana Komárková. CRU-SOE: a toolset for cyber situational awareness and decision support in incident handling. Computers & Security, 115 (??):Article 102609, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000086.

Ho-Sam-Sooi:2021:IES

[HSSPK21] Nick Ho-Sam-Sooi, Wolter Pieters, and Maarten Kroesen. Investigating the effect of security and privacy on IoT device purchase behaviour. Computers & Security, 102(??):Article 102132, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304053.

Haag:2025:ICR

[HSW25] Steffi Haag, Nils Siegfried, and Nane Winkler. Informal control responses to information security policy violations: a factorial survey on insurance employees’ moral licensing of insider threats. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002640.

He:2024:SSS

[HSX^{+}24] Ying He, Zhili Shen, Chang Xia, Jingyu Hua, Wei Tong, and Sheng Zhong. SGBA: a stealthy scapegoat backdoor attack against deep neural networks. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004339.

Haseeb:2022:CPM

[HuRMMW22] Junaid Haseeb, Saif ur Rehman Malik, Masood Mansoori, and Ian Welch. Corrigendum to ‘Probabilistic modelling of deception-based security framework using Markov decision process’ [Computers and Security 115 (2022)/102599]. Computers & Security, 117(??):Article 102689, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000876. See [HMMW22].

Hoheisel:2023:DPD

[HvCSJ23] Raphael Hoheisel, Guido van Capelleveen, Dipti K. Sarmah, and Marianne Junger. The development of phishing during the COVID-19 pandemic: an analysis of over 1100 targeted domains. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000688.

Hou:2020:IES

[HW20] Tie Hou and Victoria Wang. Industrial espionage — a systematic literature review (SLR). Computers & Security, 98(??):Article 102019, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302923.

Hofener:2025:RFI

[HW25] Ole Höfener and Qinghua Wang. Requirements framework for IoT device authentication using behavioral fingerprinting. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001488.

He:2025:IIS

[HWD^{+}25] Dalin He, Huanyu Wang, Tuo Deng, Jishi Liu, and Junnian Wang. Improving IIoT security: Unveiling threats through advanced side-channel analysis. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004401.

Hou:2025:ICC

[HWFW25] Yiting Hou, Xianglin Wei, Jianhua Fan, and Chao Wang. Interpretable CAA classification based on incorporating feature channel attention into LSTM. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005583.

Huang:2025:DIC

[HWML25] Kaijiao Huang, Lifei Wang, Faisal Mehmood, and Jianxun Liu. Design and implementation of a closed loop time delay feedback control (CLTD-FC) system for mitigating DDoS attacks. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000422.

He:2025:VSV

[HWW^{+}25] Haitao He, Sheng Wang, Yanmin Wang, Ke Liu, and Lu Yu. VulTR: Software vulnerability detection model based on multilayer key feature enhancement. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004449.

Huang:2024:ETA

[HWWB24] Jiahao Huang, Mi Wen, Minjie Wei, and Yanbing Bi. Enhancing the transferability of adversarial samples with random noise techniques. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004510.

Han:2021:DPG

[HX21] Chunling Han and Rui Xue. Differentially private GANs by adding noise to Discriminator’s loss. Computers & Security, 107(??):Article 102322, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001462.

Huang:2024:CID

[HXX^{+}24] Kun Huang, Rundong Xian, Ming Xian, Huimei Wang, and Lin Ni. A comprehensive intrusion detection method for the Internet of Vehicles based on federated learning architecture. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003729.

Hu:2021:NBM

[HXZ^{+}21] Yifan Hu, Peng Xun, Peidong Zhu, Yinqiao Xiong, Yufei Zhu, Weiheng Shi, and Chenxi Hu. Network-based multidimensional moving target defense against false data injection attack in power system. Computers & Security, 107(??):Article 102283, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001073.

Hu:2023:CSS

[HXZ^{+}23] Tianle Hu, Chonghai Xu, Shenwen Zhang, Shuangshuang Tao, and Luqun Li. Cross-site scripting detection with two-channel feature fusion embedded in self-attention mechanism. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003820.

He:2024:EMI

[HXZ^{+}24] Xinlong He, Yang Xu, Sicong Zhang, Weida Xu, and Jiale Yan. Enhance membership inference attacks in federated learning. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004455.

Hannousse:2021:HWA

[HY21] Abdelhakim Hannousse and Salima Yahiouche. Handling webshell attacks: a systematic mapping and survey. Computers & Security, 108(??):Article 102366, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001905.

Yeh:2024:ICD

[hYAS^{+}24] Jyh haw Yeh, Md Mashrur Arifin, Ning Shen, Ujwal Karki, Yi Xie, and Archana Nanjundarao. Integrity coded databases — protecting data integrity for outsourced databases. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004790.

Hussain:2022:ISL

[HYR^{+}22] Tariq Hussain, Bailin Yang, Haseeb Ur Rahman, Arshad Iqbal, Farman Ali, and Babar Shah. Improving source location privacy in social Internet of Things using a hybrid phantom routing technique. Computers & Security, 123(??):Article 102917, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003091.

Hei:2020:TFA

[HYW^{+}20] Xinhong Hei, Xinyue Yin, Yichuan Wang, Ju Ren, and Lei Zhu. A trusted feature aggregator federated learning for distributed malicious attack detection. Computers & Security, 99(??):Article 102033, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303060.

Halim:2021:EGA

[HYW^{+}21] Zahid Halim, Muhammad Nadeem Yousaf, Muhammad Waqas, Muhammad Sulaiman, Ghulam Abbas, Masroor Hussain, Iftekhar Ahmad, and Muhammad Hanif. An effective genetic algorithm-based feature selection method for intrusion detection systems. Computers & Security, 110(??):Article 102448, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002728.

Huang:2024:VVP

[HYZ^{+}24] Yuxian Huang, Geng Yang, Hao Zhou, Hua Dai, Dong Yuan, and Shui Yu. VPPFL: a verifiable privacy-preserving federated learning scheme against poisoning attacks. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004728.

Hu:2025:FSE

[HYZ25] Zhaowang Hu, Jun Ye, and Zhengqi Zhang. Faster secure and efficient collaborative private data cleaning based on PSI. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003906.

Hu:2020:PPB

[HZ20a] Xinwen Hu and Yi Zhuang. PHRiMA: a permission-based hybrid risk management framework for Android apps. Computers & Security, 94(??):Article 101791, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300766.

Huang:2020:DGA

[HZ20b] Linan Huang and Quanyan Zhu. A dynamic games approach to proactive defense strategies against Advanced Persistent Threats in cyber-physical systems. Computers & Security, 89(??):Article 101660, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302020.

Huang:2022:RRA

[HZ22] Linan Huang and Quanyan Zhu. RADAMS: Resilient and adaptive alert and attention management strategy against Informational Denial-of-Service (IDoS) attacks. Computers & Security, 121(??):Article 102844, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002383.

Hu:2019:BBA

[HZG19] Kexin Hu, Zhenfeng Zhang, and Kaiwen Guo. Breaking the binding: Attacks on the Merkle approach to prove liabilities and its applications. Computers & Security, 87(??):Article 101585, November 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/

pii/S0167404818314093. See corrigendum [HZG20].

Hu:2020:CBB

[HZG20] Kexin Hu, Zhenfeng Zhang, and Kaiwen Guo. Corrigendum to “Breaking the binding: Attacks on the Merkle approach to prove liabilities and its applications” [Computers and Security, Volume 87, 2019, 101585]. Computers & Security, 94(??):Article 101878, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301516. See [HZG19].

Huang:2023:DEB

[HZG^{+}23] Jia-Cheng Huang, Guo-Qiang Zeng, Guang-Gang Geng, Jian Weng, Kang-Di Lu, and Yu Zhang. Differential evolutionbased convolutional neural networks: an automatic architecture design method for intrusion detection in industrial control systems. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002201.

Hao:2025:DDD

[HZGM25] Fengrui Hao, Shiyi Zhao, Tianlong Gu, and Yan Ma. dK-DGDP: a differential privacy approach on directed social network graphs. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003360.

Hu:2024:LTT

[HZH^{+}24] Yuelin Hu, Futai Zou, Jiajia Han, Xin Sun, and Yilei Wang. LLM-TIKG: Threat intelligence knowledge graph construction utilizing large language model. Computers & Security, 145(??): ??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003043.

Hammi:2022:SSH

[HZKN22] Badis Hammi, Sherali Zeadally, Rida Khatoun, and Jamel Nebhen. Survey on smart homes: Vulnerabilities, risks, and countermeasures. Computers & Security, 117(??):Article 102677, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200075X.

Hu:2021:STV

[HZL^{+}21] Xinwen Hu, Yi Zhuang, Shang-Wei Lin, Fuyuan Zhang, Shuanglong Kan, and Zining Cao. A security type verifier for smart contracts. Computers & Security, 108(??):Article 102343, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100167X.

Hu:2023:FAC

[HZL^{+}23] Lifeng Hu, Fan Zhang, Ziyuan Liang, Ruyi Ding, Xingyu Cai, Zonghui Wang, and Wenguang Jin. FaultMorse: an automated controlled-channel attack via longest recurring sequence. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003959.

He:2021:SNT

[HZM21] Ting He, Yong Zheng, and Zherui Ma. Study of network time synchronisation security strategy based on polar coding. Computers & Security, 104(??):Article 102214, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000389.

Huang:2020:BBS

[HZX^{+}20] Haiping Huang, Peng Zhu, Fu Xiao, Xiang Sun, and Qinglong Huang. A blockchain-based scheme for privacypreserving and secure sharing of medical data. Computers & Security, 99(??):Article 102010, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302832.

Hu:2021:CCP

[HZX^{+}21] Yifan Hu, Peidong Zhu, Peng Xun, Bo Liu, Wenjie Kang, Yinqiao Xiong, and Weiheng Shi. CPMTD: Cyber-physical moving target defense for hardening the security of power system against false data injected attack. Computers & Security, 111(??):Article 102465, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002893.

Hou:2025:SSU

[HZX^{+}25] Yue Hou, Xinyan Zhou, Huakang Xia, Jian Wang, and Haiming Chen. Surf-Snooping: USB crosstalk leakage attacks on wireless charging. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001014.

Hu:2023:ASB

[HZYT23] Mingming Hu, Kun Zhang, Ruibang You, and Bibo Tu. AuthConFormer: Sensor-based continuous authentication of smartphone users using a convolutional transformer. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000329.

Hu:2020:SMV

[HZZ20] Xinwen Hu, Yi Zhuang, and Fuyuan Zhang. A security modeling and verification method of embedded software based on Z and MARTE. Computers & Security, 88(??):Article 101615, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404818304358.

Huang:2024:SSF

[HZZ^{+}24] Haoxiang Huang, Jianbiao Zhang, Lei Zhang, Jun Hu, and YiHao Cao. SABDTM: Security-first architecture-based dynamic trusted measurement scheme for operating system of the virtual computing node. Computers & Security, 137(??): ??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005576.

Iannacone:2020:QCE

[IB20] Michael D. Iannacone and Robert A. Bridges. Quantifiable and comparable evaluations of cyber defensive capabilities: a survey and novel, unified approach. Computers & Security, 96(??):Article 101907, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301838.

Iraqi:2022:CCC

[IE22] Omar Iraqi and Hanan El Bakkali. Communizer: a collaborative cloud-based self-protecting software communities framework — focus on the alert coordination system. Computers & Security, 117(??):Article 102692, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000906.

Islami:2022:CDP

[IFHP22] Lejla Islami, Simone Fischer-Hübner, and Panos Papadimitratos. Capturing drivers’ privacy preferences for intelligent transportation systems: an intercultural perspective. Computers & Security, 123(??):Article 102913, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003066.

Ismail:2021:RAB

[IHJZ21] Salih Ismail, Hani Ragab Hassen, Mike Just, and Hind Zantout. A review of amplification-based distributed denial of service attacks and their mitigation. Computers & Security, 109(??):Article 102380, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002042.

Iyieke:2025:ASD

[IJR^{+}25] Victormills Iyieke, Hesamaldin Jadidbonab, Abdur Rakib, Jeremy Bryans, Don Dhaliwal, and Odysseas Kosmas. An adaptable security-by-design approach for ensuring a secure Over the Air (OTA) update in modern vehicles. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005741.

Imine:2020:APP

[ILB20] Youcef Imine, Ahmed Lounis, and Abdelmadjid Bouabdallah. An accountable privacy-preserving scheme for public information sharing systems. Computers & Security, 93(??):Article 101786, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300717.

Iturbe:2024:UOA

[ILVR^{+}24] Eider Iturbe, Oscar Llorente-Vazquez, Angel Rego, Erkuden Rios, and Nerea Toledo. Unleashing offensive artificial intelligence: Automated attack technique code generation. Computers & Security, 147(??):??, December 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003821.

Ianni:2023:SSC

[IM23] Michele Ianni and Elio Masciari. SCOUT: Security by computing OUTliers on activity logs. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002651.

Iadarola:2021:TID

[IMMS21] Giacomo Iadarola, Fabio Martinelli, Francesco Mercaldo, and Antonella Santone. Towards an interpretable deep learning model for mobile malware detection and family identification. Computers & Security, 105(??):Article 102198, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000225.

Iyer:2022:EPN

[IMN22] Padmavathi Iyer, Amirreza Masoumzadeh, and Paliath Narendran. On the expressive power of negated conditions and negative authorizations in access control models. Computers & Security, 116(??):Article 102586, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004090.

Itodo:2024:MLR

[IO24] Cornelius Itodo and Murat Ozer. Multivocal literature review on zero-trust security implementation. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001287.

Ibrahim:2024:MPA

[ISD24] Omar Adel Ibrahim, Savio Sciancalepore, and Roberto Di Pietro. MAG-PUFs: Authenticating IoT devices via electromagnetic physical unclonable functions and deep learning. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002074.

Ibitoye:2022:DPS

[ISM22] Olakunle Ibitoye, M. Omair Shafiq, and Ashraf Matrawy. Differentially private self-normalizing neural networks for adversarial robustness in federated learning. Computers & Security, 116(??):Article 102631, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200030X.

Imran:2023:POM

[ISR^{+}23] Muhammad Imran, Hafeez Ur Rehman Siddiqui, Ali Raza, Muhammad Amjad Raza, Furqan Rustam, and Imran Ashraf.

A performance overview of machine learning-based defense strategies for advanced persistent threats in industrial control systems. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003553.

John:2023:EMS

[JAAS^{+}23] Taran Cyriac John, Muhammad Shabbir Abbasi, Harith Al-Sahaf, Ian Welch, and Julian Jang-Jaccard. Evolving malice scoring models for ransomware detection: an automated approach by utilising genetic programming and cooperative coevolution. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001256.

Jang:2022:PBE

[Jan22] Daehee Jang. Badaslr: Exceptional cases of ASLR aiding exploitation. Computers & Security, 112(??):Article 102510, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003345.

Joyce:2023:MMR

[JANR23] Robert J. Joyce, Dev Amlani, Charles Nicholas, and Edward Raff. MOTIF: a malware reference dataset with ground truth family labels. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003133.

Jmila:2021:TSA

[JB21] Houda Jmila and Gregory Blanc. Towards security-aware 5G slice embedding. Computers & Security, 100(??):Article 102075, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303485.

Jia:2022:BFC

[JB22] Haiyan Jia and Eric P. S. Baumer. Birds of a feather: Collective privacy of online social activist groups. Computers & Security, 115(??):Article 102614, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200013X.

Jeong:2024:TMI

[JCHS24] Hoyong Jeong, Kiwon Chung, Sung Ju Hwang, and Sooel Son. Targeted Model Inversion: Distilling style encoded in predictions. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002724.

Jung:2021:NIT

[JCJ^{+}21] Changhun Jung, Jinchun Choi, Rhongho Jang, David Mohaisen, and DaeHun Nyang. A network-independent toolbased usable authentication system for Internet of Things devices. Computers & Security, 108(??):Article 102338, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001620.

Janovsky:2025:RAR

[JCS^{+}25] Adam Janovsky, Lukasz Chmielewski, Petr Svenda, Jan Jancar, and Vashek Matyas. Revisiting the analysis of references among Common Criteria certified products. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000513.

Johnston:2023:SRV

[JDB^{+}23] Allen Johnston, Paul M. Di Gangi, France Bélanger, Robert E. Crossler, Mikko Siponen, Merrill Warkentin, and Tripti Singh. Seeking rhetorical validity in fear appeal research: an application of rhetorical theory. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004126.

Jerbi:2020:UAM

[JDBB20] Manel Jerbi, Zaineb Chelly Dagdia, Slim Bechikh, and Lamjed Ben Said. On the use of artificial malicious patterns for Android malware detection. Computers & Security, 92(??):Article 101743, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404818309994.

Jerbi:2022:AMD

[JDBB22] Manel Jerbi, Zaineb Chelly Dagdia, Slim Bechikh, and Lamjed Ben Said. Android malware detection as a bi-level problem. Computers & Security, 121(??):Article 102825, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482200219X.

Junior:2021:NWA

[JE21] Manoel Domingues Junior and Nelson F. F. Ebecken. A new WAF architecture with machine learning for resourceefficient use. Computers & Security, 106(??):Article 102290, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001140.

Jiang:2022:IBC

[JFQ^{+}22] Shuai Jiang, Cai Fu, Yekui Qian, Shuai He, Jianqiang Lv, and Lansheng Han. IFAttn: Binary code similarity analysis based on interpretable features with attention. Computers & Security, 120(??):Article 102804, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001985.

Javanmardi:2024:MRM

[JGS^{+}24] Saeed Javanmardi, Meysam Ghahramani, Mohammad Shojafar, Mamoun Alazab, and Antonio M. Caruso. M-RL: a mobility and impersonation-aware IDS for DDoS UDP flooding attacks in IoT-Fog networks. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000798.

Jahromi:2020:ITH

[JHD^{+}20] Amir Namavar Jahromi, Sattar Hashemi, Ali Dehghantanha, Kim-Kwang Raymond Choo, Hadis Karimipour, David Ellis Newton, and Reza M. Parizi. An improved two-hidden-layer extreme learning machine for malware hunting. Computers & Security, 89(??):Article 101655, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819301981.

Jim:2022:EMS

[JIG22] Lincy E. Jim, Nahina Islam, and Mark A. Gregory. Enhanced MANET security using artificial immune system based danger theory to detect selfish nodes. Computers & Security, 113(??):Article 102538, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100362X.

Janovsky:2024:SCE

[JJS^{+}24] Adam Janovsky, Jan Jancar, Petr Svenda, Łukasz Chmielewski, Jiri Michalik, and Vashek Matyas. sec-certs: Examining the security certification practice for better vulnerability mitigation. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001974.

Javadpour:2024:CSC

[JJT^{+}24] Amir Javadpour, Forough Ja’fari, Tarik Taleb, Mohammad Shojafar, and Chafika Benzaïd. A comprehensive survey on cyber deception techniques to improve honeypot performance. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000932.

Jin:2023:AAM

[JJW^{+}23] Xin Jin, Xin Jin, Ruxin Wang, Shin-Jye Lee, Qian Jiang, Shaowen Yao, and Wei Zhou. Adversarial attacks on multifocus image fusion models. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003656.

Jeon:2021:AAV

[JK21a] Sanghoon Jeon and Huy Kang Kim. AutoVAS: an automated vulnerability analysis system with a deep learning approach. Computers & Security, 106(??):Article 102308, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001322.

Jeon:2021:TIM

[JK21b] Sanghoon Jeon and Huy Kang Kim. TZMon: Improving mobile game security with ARM trustzone. Computers & Security, 109(??):Article 102391, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002157.

Jingade:2023:ERA

[JK23] Raghavendra Raghuram Jingade and Rajaram Sanjeev Kunte. Extended right-angle difference ternary co-relation pattern: a new feature descriptor for face anti-spoofing. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003310.

Jian:2021:NFI

[JKR^{+}21] Yifei Jian, Hongbo Kuang, Chenglong Ren, Zicheng Ma, and Haizhou Wang. A novel framework for image-based malware detection with a deep neural network. Computers & Security, 109(??):Article 102400, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002248.

Jung:2025:SSR

[JKS^{+}25] Junho Jung, BeomSeok Kim, Heeseung Son, Daehee Jang, Ben Lee, and Jinsung Cho. A Segmented Stack Randomization for bare-metal IoT devices. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000318.

J:2022:PPF

[JKSS22] Dharani J., Sundarakantham K., Kunwar Singh, and Mercy Shalinie S. A privacy-preserving framework for endorsement process in hyperledger fabric. Computers & Security, 116(??):Article 102637, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000360.

Jung:2022:CCA

[JLB22] Bill Jung, Yan Li, and Tamir Bechor. CAVP: a contextaware vulnerability prioritization model. Computers & Security, 116(??):Article 102639, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000384.

Jiang:2025:TAA

[JLC25] Tian Jiang, Yunqi Liu, and Xiaohui Cui. Textual adversarial attacks in cybersecurity named entity recognition. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005844.

Jiang:2024:BBD

[JLLG24] Yongkang Jiang, Gaolei Li, Shenghong Li, and Ying Guo. BenchMFC: a benchmark dataset for trustworthy malware family classification under concept drift. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000075.

Jin:2020:SRT

[JLQ^{+}20] Dongzi Jin, Yiqin Lu, Jiancheng Qin, Zhe Cheng, and Zhongshu Mao. SwiftIDS: Real-time intrusion detection system based on LightGBM and parallel intrusion detection mechanism. Computers & Security, 97(??):Article 101984, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302571.

Jo:2022:VAE

[JLS22] Hyeonseong Jo, Yongjae Lee, and Seungwon Shin. Vulcan: Automatic extraction and analysis of cyber threat intelligence from unstructured text. Computers & Security, 120(??):Article 102763, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001584.

Jnr:2025:HFD

[JMU^{+}25] Maxwell Dorgbefu Jnr, Yaw Marfo Missah, Najim Ussiph, Gaddafi Abdul-Salaam, Oliver Kornyo, and Joseph Mawulorm Mensah. Hybrid framework of differential privacy and secure multi-party computation for privacy-preserving entity resolution. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002925.

Jafarian:2023:MDM

[JN23] Jafar Haadi Jafarian and Amirreza Niakanlahiji. Multi-RHM: Defeating multi-staged enterprise intrusion attacks through multi-dimensional and multi-parameter host identity anonymization. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003509.

Jungebloud:2025:MBS

[JNKZ25] Tino Jungebloud, Nhung H. Nguyen, Dan Dongseong Kim, and Armin Zimmermann. Model-based structural and behavioral cybersecurity risk assessment in system designs. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002329.

Jin:2025:ASR

[JP25] Hongyu Jin and Panos Papadimitratos. Accountable, scalable and DoS-resilient secure vehicular communication. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001580.

Jang:2020:ICE

[JPL20] Hyerean Jang, Moon Chan Park, and Dong Hoon Lee. IBV-CFI: Efficient fine-grained control-flow integrity preserving CFG precision. Computers & Security, 94(??):Article 101828, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301024.

Jha:2020:RNN

[JPLT20] Sudan Jha, Deepak Prashar, Hoang Viet Long, and David Taniar. Recurrent neural network for detecting malware. Computers & Security, 99(??):Article 102037, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303102.

Jacob:2022:ADT

[JQYL22] Stephen Jacob, Yuansong Qiao, Yuhang Ye, and Brian Lee. Anomalous distributed traffic: Detecting cyber security attacks amongst microservices using graph convolutional networks. Computers & Security, 118(??):Article 102728, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001237.

Jethava:2024:EST

[JR24] Gordhan Jethava and Udai Pratap Rao. Exploring security and trust mechanisms in online social networks: an extensive review. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000919.

Jayarao:2024:IST

[JRP24] Guruprasad B. Jayarao, Sanjog Ray, and Prabin Kumar Panigrahi. Information security threats and organizational readiness in nWFH scenarios. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000464.

Jeremiah:2025:NDN

[JRT^{+}25] Daniel Jeremiah, Husnain Rafiq, Vinh Thong Ta, Muhammad Usman, Mohsin Raza, and Muhammad Awais. NIOM-DGA: Nature-inspired optimised ML-based model for DGA detection. Computers & Security, 157(??):??, October 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002500.

Janiszewski:2025:MRE

[JS25] Marek Janiszewski and Krzysztof Szczypiorski. Methodological reliability evaluation of trust and reputation management systems. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003098.

Jayalaxmi:2023:PDL

[JSK^{+}23] P. L. S. Jayalaxmi, Rahul Saha, Gulshan Kumar, Mamoun Alazab, Mauro Conti, and Xiaochun Cheng. PIGNUS: a deep learning model for IDS in industrial Internet-of-Things. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002250.

Joshi:2025:COR

[JSYK25] Chaitanya Joshi, Sergeja Slapnicar, Jinming Yang, and Ryan K. L. Ko. Contrasting the optimal resource allocation to cybersecurity controls and cyber insurance using prospect theory versus expected utility theory. Computers & Security, 154(??): ??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001397.

Jedrzejewski:2024:AML

[JTF^{+}24] Felix Viktor Jedrzejewski, Lukas Thode, Jannik Fischbach, Tony Gorschek, Daniel Mendez, and Niklas Lavesson. Adversarial machine learning in industry: a systematic literature review. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002931.

Jiang:2024:HSF

[JWG^{+}24] Xunzhi Jiang, Shen Wang, Yuxin Gong, Tingyue Yu, Li Liu, and Xiangzhan Yu. HAformer: Semantic fusion of hex machine code and assembly code for cross-architecture binary vulnerability detection. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003341.

Jiang:2025:DCB

[JWY^{+}25] Hao Jiang, Kang Wang, Yujie Yang, Shan Zhong, Shuai Zhang, Chengjie Liu, Xiarun Chen, and Weiping Wen. DynamicFuzz: Confidence-based directed greybox fuzzing for programs with unreliable call graphs. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003803.

Jiang:2025:MNM

[JXYF25] Xuefeng Jiang, Liuquan Xu, Li Yu, and Xianjin Fang. MFT: a novel memory flow transformer efficient intrusion detection method. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004796.

Jin:2025:EPI

[JYCL25] Hongjoo Jin, Sumin Yang, Haehyun Cho, and Dong Hoon Lee. Enhancing in-process isolation for robust defense against information disclosure attacks. Computers & Security, 152(??): ??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000598.

Jia:2025:IAT

[JYF^{+}25] Zhizhuang Jia, Chao Yang, Pengbin Feng, Xiaoyun Zhao, Xinghua Li, and Jianfeng Ma. Impact assessment of thirdparty library vulnerabilities through vulnerability reachability analysis. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002354.

Jafar:2025:IPR

[JYL25] Mousa Tayseer Jafar, Lu-Xing Yang, and Gang Li. An innovative practical roadmap for optimal control strategies in malware propagation through the integration of RL with MPC. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004917.

Javeed:2021:PDA

[JYS21] Arsalan Javeed, Cemal Yilmaz, and Erkay Savas. Detector +: an approach for detecting, isolating, and preventing timing attacks. Computers & Security, 110(??):Article 102454, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002789.

Jia:2025:HAG

[JYWS25] Junbo Jia, Li Yang, Yuchen Wang, and Anyuan Sang. Hyper attack graph: Constructing a hypergraph for cyber threat intelligence analysis. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004991.

Jia:2023:DIE

[JYZ^{+}23] Zhizhuang Jia, Chao Yang, Xiaoyun Zhao, Xinghua Li, and Jianfeng Ma. Design and implementation of an efficient container tag dynamic taint analysis. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004388.

Jiang:2023:IWP

[JZ23] Randi Jiang and Jianru Zhang. The impact of work pressure and work completion justification on intentional nonmalicious information security policy violation intention. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001633.

Jiang:2024:OVD

[JZL^{+}24] Guangshang Jiang, Hanlin Zhang, Jie Lin, Fanyu Kong, and Leyun Yu. Optimized verifiable delegated private set intersection on outsourced private datasets. Computers & Security, 141 (??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001238.

Jin:2022:BRW

[JZQ^{+}22] Chengbin Jin, Yongbin Zhou, Xinkuan Qiu, Qi Feng, and Qian Zhang. Breaking real-world COTS USIM cards with unknown side-channel countermeasures. Computers & Security, 113(??):Article 102531, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003552.

Jin:2025:PMV

[JZY^{+}25] Jiaobo Jin, Tiantian Zhu, Qixuan Yuan, Tieming Chen, Mingqi Lv, Chenbin Zheng, Jian-Ping Mei, and Xiang Pan. PDCleaner: a multi-view collaborative data compression method for provenance graph-based APT detection systems. Computers & Security, 152(??):??, May 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000483.

Kingo:2023:UCS

[KA23] Thomas Kingo and Diego F. Aranha. User-centric security analysis of MitID: the Danish passwordless digital identity solution. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002869.

Kacha:2025:NOA

[Kac25] Lynda Kacha. NKAB: an optimization approach for kanonymity based on Black Hole Algorithm. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825003013.

Ki-Aries:2022:ASS

[KAFDW22] Duncan Ki-Aries, Shamal Faily, Huseyin Dogan, and Christopher Williams. Assessing system of systems information security risk with OASoSIS. Computers & Security, 117(??):Article 102690, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000888.

Khan:2021:EAT

[KAK21] Kashif Mehboob Khan, Junaid Arshad, and Muhammad Mubashir Khan. Empirical analysis of transaction malleability within blockchain-based e-voting. Computers & Security, 100(??): Article 102081, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303540.

Kanwal:2021:RPP

[KAM^{+}21] Tehsin Kanwal, Adeel Anjum, Saif U. R. Malik, Haider Sajjad, Abid Khan, Umar Manzoor, and Alia Asheralieva. A robust privacy preserving approach for electronic health records using multiple dataset with multiple sensitive attributes. Computers & Security, 105(??):Article 102224, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000481.

Kornyo:2023:BAC

[KAO^{+}23] Oliver Kornyo, Michael Asante, Richard Opoku, Kwabena Owusu-Agyemang, Benjamin Tei Partey, Emmanuel Kwesi Baah, and Nkrumah Boadu. Botnet attacks classification in AMI networks with recursive feature elimination (RFE) and machine learning algorithms. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003668.

Khan:2023:NDB

[KAU^{+}23] Saddam Hussain Khan, Tahani Jaser Alahmadi, Wasi Ullah, Javed Iqbal, Azizur Rahim, Hend Khalid Alkahtani, Wajdi Alghamdi, and Alaa Omran Almagrabi. A new deep boosted CNN and ensemble learning based IoT malware detection. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300295X.

Kumar:2023:NII

[KBAG23] Chandan Kumar, Soham Biswas, Md. Sarfaraj Alam Ansari, and Mahesh Chandra Govil. Nature-inspired intrusion detection system for protecting software-defined networks controller. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003486.

Kucab:2021:RAI

[KBC21] Michał Kucab, Piotr Boryło, and Piotr Chołda. Remote attestation and integrity measurements with Intel SGX for virtual machines. Computers & Security, 106(??):Article 102300, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001243.

Khalil:2023:TMC

[KBD^{+}23] Shaymaa Mamdouh Khalil, Hayretdin Bahsi, Henry Ochieng’ Dola, Tarmo Korõtko, Kieran McLaughlin, and Vahur Kotkas. Threat modeling of cyber-physical systems — a case study of a microgrid system. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200342X.

Katsikeas:2024:DVC

[KBE^{+}24] Sotirios Katsikeas, Andrei Buhaiu, Mathias Ekstedt, Zeeshan Afzal, Simon Hacks, and Preetam Mukherjee. Development and validation of coreLang: a threat modeling language for the ICT domain. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003626.

Khalil:2024:TMI

[KBK24] Shaymaa Mamdouh Khalil, Hayretdin Bahsi, and Tarmo Korõtko. Threat modeling of industrial control systems: a systematic literature review. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004534.

Kishore:2023:ETS

[KBMM23] Pushkar Kishore, Swadhin Kumar Barisal, Durga Prasad Mohapatra, and Rajib Mall. An efficient two-stage pipeline model with filtering algorithm for mislabeled malware detection. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004091.

Kim:2022:OMD

[KC22] Jin-Young Kim and Sung-Bae Cho. Obfuscated malware detection using deep generative model based on global/local features. Computers & Security, 112(??):Article 102501, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003254.

Kumar:2021:DSS

[KD21] V. Anil Kumar and Debabrata Das. Data sequence signal manipulation in multipath TCP (MPTCP): the vulnerability, attack and its detection. Computers & Security, 103(??):Article 102180, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821000043.

Kayisoglu:2025:ECS

[KDBB25] Gizem Kayisoglu, Emre Duzenli, Pelin Bolat, and Aleksei Bondarenko. Exploring cyber security threats and security models in cross-border paperless maritime trade system. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002937.

Kravchik:2022:PEP

[KDBS22] Moshe Kravchik, Luca Demetrio, Battista Biggio, and Asaf Shabtai. Practical evaluation of poisoning attacks on online anomaly detectors in industrial control systems. Computers & Security, 122(??):Article 102901, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002942.

Karbab:2020:SRUa

[KDDM20a] ElMouatez Billah Karbab, Mourad Debbabi, Abdelouahid Derhab, and Djedjiga Mouheb. Scalable and robust unsupervised Android malware fingerprinting using communitybased network partitioning. Computers & Security, 96(??): Article 101932, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482030208X.

Karbab:2020:SRUb

[KDDM20b] ElMouatez Billah Karbab, Mourad Debbabi, Abdelouahid Derhab, and Djedjiga Mouheb. Scalable and robust unsupervised Android malware fingerprinting using communitybased network partitioning. Computers & Security, 97(??): Article 101965, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301437.

Kenyon:2020:PID

[KDE20] A. Kenyon, L. Deka, and D. Elizondo. Are public intrusion datasets fit for purpose characterising the state of the art in intrusion event datasets. Computers & Security, 99(??):Article 102022, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302959.

Kozak:2025:UWM

[KDTR25] Matous Kozak, Luca Demetrio, Dmitrijs Trizna, and Fabio Roli. Updating Windows malware detectors: Balancing robustness and regression against adversarial EXEmples. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001555.

Kilic:2025:KSS

[KE25] Vahide Nida Kiliç and Esra Saraç Essiz. K-Salp Swarm Anomaly Detection (K-SAD): a novel clustering and thresholdbased approach for cybersecurity applications. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825000148.

Khalil:2023:PPF

[KEB23] Maysaa Khalil, Moez Esseghir, and Leila Merghem Boulahia. Privacy-preserving federated learning: an application for big data load forecast in buildings. Computers & Security, 131(??): ??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001219.

Kim:2021:SDP

[KEK^{+}21] Jong Wook Kim, Kennedy Edemacu, Jong Seon Kim, Yon Dohn Chung, and Beakcheol Jang. A survey of differential privacy-based techniques and their applicability to locationbased services. Computers & Security, 111(??):Article 102464, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002881.

Kurniawan:2022:KKG

[KEK^{+}22] Kabul Kurniawan, Andreas Ekelhart, Elmar Kiesling, Gerald Quirchmayr, and A. Min Tjoa. KRYSTAL: Knowledge graph-based framework for tactical attack discovery in audit data. Computers & Security, 121(??):Article 102828, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200222X.

Kemp:2023:EPC

[Kem23] Steven Kemp. Exploring public cybercrime prevention campaigns and victimization of businesses: a Bayesian model averaging approach. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004813.

Khan:2025:EDB

[KEM^{+}25a] Raviha Khan, Hossien B. Eldeeb, Brahim Mefgouda, Omar Alhussein, Hani Saleh, and Sami Muhaidat. Encoder decoderbased virtual Physically Unclonable Function for Internet of Things device authentication using split-learning. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004693.

Kumar:2025:SCP

[KEM25b] Mahender Kumar, Gregory Epiphaniou, and Carsten Maple. Security of cyber-physical Additive Manufacturing supply chain: Survey, attack taxonomy and solutions. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002469.

Kuchar:2025:AAI

[KF25] Karel Kuchar and Radek Fujdiak. Analyzing anomalies in industrial networks: a data-driven approach to enhance security in manufacturing processes. Computers & Security, 153(??): ??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000847.

Khan:2025:IEP

[KFB^{+}25] Neeshe Khan, Steven Furnell, Maria Bada, Matthew Rand, and Jason R. C. Nurse. Investigating the experiences of providing cyber security support to small- and medium-sized enterprises. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001373.

Khalil:2024:CLR

[KFHM24] Fatima Mavra Khalil, Adnan Fazil, Muhammad Jawad Hussain, and Ammar Masood. Cross-layer RF distance bounding scheme for passive and semi-passive ubiquitous computing systems. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005436.

Kozik:2024:WET

[KFP^{+}24] Rafał Kozik, Massimo Ficco, Aleksandra Pawlicka, Marek Pawlicki, Francesco Palmieri, and Michał Choraś. When explainability turns into a threat — using xAI to fool a fake news detection method. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005096.

Kuang:2022:SRA

[KFS^{+}22] Boyu Kuang, Anmin Fu, Willy Susilo, Shui Yu, and Yansong Gao. A survey of remote attestation in Internet of Things: Attacks, countermeasures, and prospects. Computers & Security, 112(??):Article 102498, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003229.

Kuang:2020:DRD

[KFZ^{+}20] Boyu Kuang, Anmin Fu, Lu Zhou, Willy Susilo, and Yuqing Zhang. DO-RA: Data-oriented runtime attestation for IoT devices. Computers & Security, 97(??):Article 101945, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302212.

Kumar:2020:MCS

[KG20] Rakesh Kumar and Rinkaj Goyal. Modeling continuous security: a conceptual model for automated DevSecOps using open-source software over cloud (ADOC). Computers & Security, 97(??):Article 101967, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302406.

Khando:2021:EEI

[KGIS21] Khando Khando, Shang Gao, Sirajul M. Islam, and Ali Salman. Enhancing employees information security awareness in private and public organisations: a systematic literature review. Computers & Security, 106(??):Article 102267, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000912.

Kampourakis:2023:SLR

[KGK23] Vyron Kampourakis, Vasileios Gkioulos, and Sokratis Katsikas. A systematic literature review on wireless security testbeds in the cyber–physical realm. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002936.

Khakpour:2021:FSS

[Kha21] Narges Khakpour. A field-sensitive security monitor for object-oriented programs. Computers & Security, 108(??): Article 102349, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821001735.

Kim:2024:DDD

[KHG24] Yoonjib Kim, Saqib Hakak, and Ali Ghorbani. Detecting Distributed Denial-of-Service (DDoS) attacks that generate false authentications on Electric Vehicle (EV) charging infrastructure. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002943.

Kyytsonen:2022:SAI

[KIAV22] Maiju Kyytsönen, Jonna Ikonen, Anna-Mari Aalto, and Tuulikki Vehko. The self-assessed information security skills of the Finnish population: a regression analysis. Computers & Security, 118(??):Article 102732, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001274.

Kilic:2025:QFP

[Kil25] Ahmet Kilic. A quantitative framework for physical cybersecurity in public EVSE systems. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003748.

Khan:2023:EPM

[KIMJ23] Naurin Farooq Khan, Naveed Ikram, Hajra Murtaza, and Mehwish Javed. Evaluating protection motivation based cybersecurity awareness training on Kirkpatrick’s Model. Computers & Security, 125(??):??, February 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004412.

Kumari:2023:CSD

[KJ23] Pooja Kumari and Ankit Kumar Jain. A comprehensive study of DDoS attacks over IoT network and their countermeasures. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000068.

Krawiec:2025:PML

[KJB^{+}25] Piotr Krawiec, Robert Janowski, Jordi Mongay Batalla, Elżbieta Andrukiewicz, Waldemar Latoszek, and Constandinos X. Mavromoustakis. On providing multi-level security assurance based on Common Criteria for O-RAN mobile network equipment. A test case: O-RAN distributed unit. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005777.

Kaliyar:2020:LLE

[KJCL20] Pallavi Kaliyar, Wafa Ben Jaballah, Mauro Conti, and Chhagan Lal. LiDL: Localization with early detection of sybil and wormhole attacks in IoT networks. Computers & Security, 94(??):Article 101849, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030122X.

Katsikeas:2022:VPM

[KJHL22] Sotirios Katsikeas, Pontus Johnsson, Simon Hacks, and Robert Lagerström. VehicleLang: a probabilistic modeling and simulation language for modern vehicle IT infrastructures. Computers & Security, 117(??):Article 102705, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001031.

Khan:2024:DDV

[KJI^{+}24] Muneeb Hassan Khan, Abdul Rehman Javed, Zafar Iqbal, Muhammad Asim, and Ali Ismail Awad. DivaCAN: Detecting in-vehicle intrusion attacks on a controller area network using ensemble learning. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000130.

Kumar:2025:EDL

[KJI25] Prabhat Kumar, Alireza Jolfaei, and A. K. M. Najmul Islam. An enhanced deep-learning empowered threat-hunting framework for software-defined Internet of Things. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004140.

Kim:2024:DDR

[KJJ^{+}24] Seon Kwon Kim, Hongjoo Jin, Kyungho Joo, Jiwon Lee, and Dong Hoon Lee. DROPSYS: Detection of ROP attacks using system information. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001147.

Jia:2024:EFM

[kJLcL24] Cheng kun Jia, Min Long, and Yong chao Liu. Enhanced face morphing attack detection using error-level analysis and efficient selective kernel network. Computers & Security, 137(??): ??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005485.

Kim:2023:PSI

[KJPP23] Sung-Kyung Kim, Eun-Tae Jang, Hanjin Park, and Ki-Woong Park. Pwnable-Sherpa: an interactive coaching system with a case study of pwnable challenges. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004011.

Krishnan:2020:ECR

[KJR^{+}20] R. Santhana Krishnan, E. Golden Julie, Y. Harold Robinson, Raghvendra Kumar, Pham Huy Thong, and Le Hoang Son. Enhanced certificate revocation scheme with justification facility in mobile ad-hoc networks. Computers & Security, 97(??):Article 101962, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302388.

Konashevych:2020:RTB

[KK20] Oleksii Konashevych and Oleg Khovayko. Randpay: the technology for blockchain micropayments and transactions which require recipient’s consent. Computers & Security, 96(??):Article 101892, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301656.

Kalam:2025:AIS

[KK25a] Sidra Kalam and Ajit Kumar Keshri. Advancing IoMT security: a two-factor authentication model employing PUF and fuzzy logic techniques. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004437.

Karpagavalli:2025:EIW

[KK25b] C. Karpagavalli and M. Kaliappan. Edge Implicit Weighting with graph transformers for robust intrusion detection in Internet of Things network. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824006059.

Kumar:2025:AAM

[KK25c] K. Praveen Kumar and Dr. N. Suresh Kumar. Advanced attack mitigation in IoT gateway protocols. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002287.

Kaur:2021:CSD

[KKAS21] Sukhveer Kaur, Krishan Kumar, Naveen Aggarwal, and Gurdeep Singh. A comprehensive survey of DDoS defense solutions in SDN: Taxonomy, research challenges, and future directions. Computers & Security, 110(??):Article 102423, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002479.

Kim:2025:HDF

[KKB^{+}25] Giyoon Kim, Soojin Kang, Seungjun Baek, Kimoon Kim, and Jongsung Kim. How to decrypt files encrypted by Rhysida ransomware without the attacker’s private key. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482500029X.

Kampourakis:2025:CCT

[KKGK25] Vyron Kampourakis, Georgios Kavallieratos, Vasileios Gkioulos, and Sokratis Katsikas. Cracks in the chain: a technical analysis of real-life supply chain security incidents. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003621.

Kim:2024:SVW

[KKHK24] Giyoon Kim, Soojin Kang, Uk Hur, and Jongsung Kim. A study on vulnerability of the wickr login system in windows from a live forensics perspective. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823005825.

Kim:2021:CAV

[KKJ^{+}21] Kyounggon Kim, Jun Seok Kim, Seonghoon Jeong, Jo-Hee Park, and Huy Kang Kim. Cybersecurity for autonomous vehicles: Review of attacks and defense. Computers & Security, 103(??):Article 102150, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304235.

Ko:2023:MTA

[KKK23] Kyoungmin Ko, SungHwan Kim, and Hyun Kwon. Multitargeted audio adversarial example for use against speech recognition systems. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000780.

Kiran:2025:EPM

[KKM^{+}25] Uzma Kiran, Naurin Farooq Khan, Hajra Murtaza, Ali Farooq, and Henri Pirkkalainen. Explanatory and predictive modeling of cybersecurity behaviors using protection motivation theory. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824005091.

Karlsson:2022:ISP

[KKP22] Fredrik Karlsson, Ella Kolkowska, and Johan Petersson. Information security policy compliance-eliciting requirements for a computerized software to support value-based compliance analysis. Computers & Security, 114(??):Article 102578, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821004028.

Kalaria:2024:ISF

[KKR^{+}24] Rudri Kalaria, A. S. M. Kayes, Wenny Rahayu, Eric Pardede, and Ahmad Salehi S. IoTPredictor: a security framework for predicting IoT device behaviours and detecting malicious devices against cyber attacks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003420.

Kalaria:2021:SMA

[KKRP21] Rudri Kalaria, A. S. M. Kayes, Wenny Rahayu, and Eric Pardede. A secure mutual authentication approach to fog computing environment. Computers & Security, 111(??):Article 102483, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003072.

Kruzikova:2022:USU

[KKS^{+}22] Agata Kruzikova, Lenka Knapova, David Smahel, Lenka Dedkova, and Vashek Matyas. Usable and secure? User perception of four authentication methods for mobile banking. Computers & Security, 115(??):Article 102603, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000025.

Kim:2023:EDP

[KKY^{+}23] Jinwoo Kim, Yeonkeun Kim, Vinod Yegneswaran, Phillip Porras, Seungwon Shin, and Taejune Park. Extended data plane architecture for in-network security services in softwaredefined networks. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003686.

Kwon:2022:ETA

[KL22] Hyun Kwon and Sanghyun Lee. Ensemble transfer attack targeting text classification systems. Computers & Security, 117(??):Article 102695, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000931. See erratum [KL23].

Kwon:2023:EET

[KL23] Hyun Kwon and Sanghyun Lee. Erratum to ‘Ensemble transfer attack targeting text classification systems’ [Computers and Security 117 (2022) 1–8/102695]. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003364. See [KL22].

Kim:2024:CRA

[KL24] Hee Yeon Kim and Dong Hoon Lee. CatchFuzz: Reliable active anti-fuzzing techniques against coverage-guided fuzzer. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002062.

Koutsokostas:2022:IAA

[KLA^{+}22] Vasilios Koutsokostas, Nikolaos Lykousas, Theodoros Apostolopoulos, Gabriele Orazi, Amrita Ghosal, Fran Casino, Mauro Conti, and Constantinos Patsakis. Invoice #31415 attached: Automated analysis of malicious Microsoft Office documents. Computers & Security, 114(??):Article 102582, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821004053.

Kumar:2023:UAD

[KLC23] Kiran Kumar, Dapeng Liu, and Lemuria Carter. Understanding the adoption of digital conferencing tools: Unpacking the impact of privacy concerns and incident response efficacy. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002857.

Kalutharage:2025:NLD

[KLC25] Chathuranga Sampath Kalutharage, Xiaodong Liu, and Christos Chrysoulas. Neurosymbolic learning and domain knowledge-driven explainable AI for enhanced IoT network attack detection and response. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000070.

Katsikeas:2024:EET

[KLJE24] Sotirios Katsikeas, Engla Rencelj Ling, Pontus Johnsson, and Mathias Ekstedt. Empirical evaluation of a threat modeling language as a cybersecurity assessment tool. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000440.

Kang:2021:WDD

[KLNW21] Yilin Kang, Yong Liu, Ben Niu, and Weiping Wang. Weighted distributed differential privacy ERM: Convex and non-convex. Computers & Security, 106(??):Article 102275, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000997.

Kong:2024:MTV

[KLP^{+}24] Lingdi Kong, Senlin Luo, Limin Pan, Zhouting Wu, and Xinshuai Li. A multi-type vulnerability detection framework with parallel perspective fusion and hierarchical feature enhancement. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000889.

Kamtam:2025:WWO

[KLR^{+}25] Suraj Harsha Kamtam, Qian Lu, Abdur Rakib, Muhamad Azfar Ramli, Rakhi Manohar Mepparambath, Siraj Ahmed Shaikh, and Hoang Nga Nguyen. WOLVES: Window of opportunity attack feasibility likelihood value estimation through a simulation-based approach. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482500238X.

Konsta:2024:CSS

[KLSD24a] Alyzia-Maria Konsta, Alberto Lluch Lafuente, Beatrice Spiga, and Nicola Dragoni. Corrigendum to “Survey: Automatic generation of attack trees and attack graphs” [Computers and Security Volume 137, February 2024, 103602]. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005709. See [KLSD24b].

Konsta:2024:SAG

[KLSD24b] Alyzia-Maria Konsta, Alberto Lluch Lafuente, Beatrice Spiga, and Nicola Dragoni. Survey: Automatic generation of attack trees and attack graphs. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005126. See corri-

gendum [KLSD24a].

Kern:2024:LMD

[KLSW24] Manuel Kern, Max Landauer, Florian Skopik, and Edgar Weippl. A logging maturity and decision model for the selection of intrusion detection cyber security solutions. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001457.

Kang:2023:IMN

[KLW^{+}23] Hongzhaoning Kang, Gang Liu, Quan Wang, Qi Zhang, Jiamin Niu, and Nan Luo. An improved minimal noise role mining algorithm based on role interpretability. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482300010X.

Khan:2024:DSP

[KLZ24] Md Sabuj Khan, Hengjian Li, and Chuan Zhao. Deep secure PalmNet: a novel cancelable palmprint template protection scheme with deep attention net and randomized hashing security mechanism. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001640.

Khalid:2022:VAQ

[KM22] Fatima Khalid and Ammar Masood. Vulnerability analysis of Qualcomm Secure Execution Environment (QSEE). Computers & Security, 116(??):Article 102628, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200027X.

K:2024:SMI

[KM24a] Jayashre K. and Amsaprabhaa M. Safeguarding media integrity: a hybrid optimized deep feature fusion based deepfake detection in videos. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001615.

Kamble:2024:HOE

[KM24b] Naresh Kamble and Nilamadhab Mishra. Hybrid optimization enabled squeeze net for phishing attack detection. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002037.

Klauzer:2024:FSR

[KMAHU24] Anja Klauzer, Markus Maier, Lore Abart-Heriszt, and Johanna Ullrich. Fostering security research in the energy sector: a validation of open source intelligence for power grid model data. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400347X.

Katsantonis:2021:DEC

[KMG21] Menelaos N. Katsantonis, Ioannis Mavridis, and Dimitris Gritzalis. Design and evaluation of COFELET-based approaches for cyber security learning and training. Computers & Security, 105(??):Article 102263, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000870.

Koucham:2022:CDA

[KMH^{+}22] Oualid Koucham, Stéphane Mocanu, Guillaume Hiet, Jean-Marc Thiriet, and Frédéric Majorczyk. Cross-domain alert correlation methodology for industrial control systems. Computers & Security, 118(??):Article 102723, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001183.

Kruzikova:2024:TFA

[KMK^{+}24] Agata Kruzikova, Michal Muzik, Lenka Knapova, Lenka Dedkova, David Smahel, and Vashek Matyas. Two-factor authentication time: How time-efficiency and time-satisfaction are associated with perceived security and satisfaction. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005771.

Kam:2020:CCL

[KMOC20] Hwee-Joo Kam, Philip Menard, Dustin Ormond, and Robert E. Crossler. Cultivating cybersecurity learning: an integration of self-determination and flow. Computers & Security, 96(??):Article 101875, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301486.

Knox:2020:WRH

[KMP^{+}20] Shawn Knox, Steven Moghadam, Kenny Patrick, Anh Phan, and Kim-Kwang Raymond Choo. What’s really ‘ Happning’? A forensic analysis of Android and iOS Happn dating apps. Computers & Security, 94(??):Article 101833, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301073.

Knight:2020:FEC

[KN20] Richard Knight and Jason R. C. Nurse. A framework for effective corporate communication after cyber security incidents. Computers & Security, 99(??):Article 102036, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303096.

Kwon:2023:AAD

[KN23] Hyun Kwon and Seung-Hun Nam. Audio adversarial detection through classification score on speech recognition systems. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004539.

Krzyworzeka:2024:PCB

[KOO24] Natalia Krzyworzeka, Lidia Ogiela, and Marek R. Ogiela. Personal CAPTCHA-based authentication protocol. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005230.

Kushwah:2021:OEL

[KR21] Gopal Singh Kushwah and Virender Ranga. Optimized extreme learning machine for detecting DDoS attacks in cloud computing. Computers & Security, 105(??):Article 102260, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000845.

Khurshid:2023:AAT

[KR23] Anum Khurshid and Shahid Raza. AutoCert: Automated TOCTOU-secure digital certification for IoT with combined authentication and assurance. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003443.

Kavrestad:2024:DPC

[KRN24] Joakim Kävrestad, Jana Rambusch, and Marcus Nohlberg. Design principles for cognitively accessible cybersecurity training. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005400.

Kuhn:2023:CVS

[KRR23] Philipp Kühn, David N. Relke, and Christian Reuter. Common vulnerability scoring system prediction based on open source intelligence information sources. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001967.

Kasongo:2020:DLM

[KS20] Sydney Mambwe Kasongo and Yanxia Sun. A deep learning method with wrapper based feature extraction for wireless intrusion detection system. Computers & Security, 92(??):Article 101752, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300365.

Kumar:2023:SAD

[KS23a] Vikash Kumar and Ditipriya Sinha. Synthetic attack data generation model applying generative adversarial network for intrusion detection. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004461.

Kumar:2023:SEB

[KS23b] Yogendra Kumar and Basant Subba. Stacking ensemblebased HIDS framework for detecting anomalous system processes in Windows based operating systems using multiple word embedding. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003534.

Karayel:2025:USS

[KS25a] Tugçe Karayel and Metin Saygili. Understanding smartphone security behavior through the core constructs of protection motivation theory: a comparative study of iOS and Android users. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003414.

Kumar:2025:SID

[KS25b] Anit Kumar and Dhanpratap Singh. Securing IoT devices in edge computing through reinforcement learning. Computers & Security, 155(??):??, August 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001634.

Kim:2023:ATL

[KSC23] Suryeon Kim, Seungwon Shin, and Hyunwoo Choi. AVX-TSCHA: Leaking information through AVX extensions in commercial processors. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003474.

Kitkowska:2023:DPE

[KSMW23] Agnieszka Kitkowska, Yefim Shulman, Leonardo A. Martucci, and Erik Wästlund. Designing for privacy: Exploring the influence of affect and individual characteristics on users’ interactions with privacy policies. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003784.

Karjalainen:2020:TST

[KSS20] Mari Karjalainen, Mikko Siponen, and Suprateek Sarker. Toward a stage theory of the development of employees’ information security behavior. Computers & Security, 93(??):Article 101782, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300675.

Kumar:2022:MLB

[KSSL22] Ayush Kumar, Mrinalini Shridhar, Sahithya Swaminathan, and Teng Joon Lim. Machine learning-based early detection of IoT botnets using network-edge traffic. Computers & Security, 117(??):Article 102693, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000918.

Kumar:2022:WCC

[KSVY22] Rajesh Kumar, Siddharth Sharma, Chirag Vachhani, and Nitish Yadav. What changed in the cyber-security after COVID-19? Computers & Security, 120(??):Article 102821, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002152.

Kale:2023:FSW

[KT23] Rahul Kale and Vrizlynn L. L. Thing. Few-shot weaklysupervised cybersecurity anomaly detection. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001049.

Kumari:2025:DIF

[KT25] Matta Krishna Kumari and Nikhil Tripathi. Detecting interest flooding attacks in NDN: a probability-based event-driven approach. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004292.

Korba:2020:ABF

[KTGDE20] Abdelaziz Amara Korba, Nouredine Tamani, Yacine Ghamri-Doudane, and Nour El Islem Karabadji. Anomaly-based framework for detecting power overloading cyberattacks in smart grid AMI. Computers & Security, 96(??):Article 101896, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301693.

Kuo:2020:MAD

[KTH20] Kuang-Ming Kuo, Paul C. Talley, and Chi-Hsien Huang. A meta-analysis of the deterrence theory in security-compliant and security-risk behaviors. Computers & Security, 96(??): Article 101928, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302042.

Kurt:2025:AHR

[KUA^{+}25] Yasin Burak Kurt, Esma Uflaz, Emre Akyuz, Rafet Emek Kurt, and Osman Turan. Analysis of human reliability in detecting GPS spoofing on ECDIS in congested waterways under evidential reasoning and HEART approach. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000057.

Khan:2025:RRA

[KWF25] Irfan Khan, Syed Wali, and Yasir Ali Farrukh. RADIANT: Reactive Autoencoder Defense for Industrial Adversarial Network Threats. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000926.

Kang:2023:FGF

[KXZ^{+}23] Zi Kang, Hui Xia, Rui Zhang, Shuliang Jiang, Xiaolong Shi, and Zuming Zhang. FGDA-GS: Fast guided decision attack based on gradient signs for skeletal action recognition. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004327.

Khan:2022:CBR

[KYKI22] Naurin Farooq Khan, Amber Yaqoob, Muhammad Saud Khan, and Naveed Ikram. The cybersecurity behavioral research: a tertiary study. Computers & Security, 120(??):Article 102826, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002206.

Kuznetsov:2024:AEB

[KZFM24] Oleksandr Kuznetsov, Dmytro Zakharov, Emanuele Frontoni, and Andrea Maranesi. AttackNet: Enhancing biometric security via tailored convolutional neural network architectures for liveness detection. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001299.

Kamdem:2025:MDD

[KZN^{+}25] Priva Chassem Kamdem, Alain Zemkoho, Laurent Njilla, M. Nkenlifack, and Charles Kamhoua. Multi-domain deception for enhanced security in automotive networks. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002895.

Kong:2022:FFC

[KZYZ22] Ke Kong, Zhichao Zhang, Zi-Yuan Yang, and Zhaoxin Zhang. FCSCNN: Feature centralized Siamese CNN-based Android malware identification. Computers & Security, 112(??):Article 102514, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003382.

Li:2025:IAE

[LaTL^{+}25] Jiaxing Li, Yu an Tan, Xinyu Liu, Weizhi Meng, and Yuanzhang Li. Interpretable adversarial example detection via high-level concept activation vector. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824005248.

Liu:2025:IMI

[LB25] Shuaishuai Liu and Gergely Biczók. IDPFilter: Mitigating interdependent privacy issues in third-party apps. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000100.

Lavaur:2025:IIL

[LBA25] Léo Lavaur, Yann Busnel, and Fabien Autrel. Investigating the impact of label-flipping attacks against federated learning for collaborative intrusion detection. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001518.

Lichy:2023:WRB

[LBD^{+}23] Adi Lichy, Ofek Bader, Ran Dubin, Amit Dvir, and Chen Hajaj. When a RF beats a CNN and GRU, together — a comparison of deep learning and classical machine learning approaches for encrypted malware traffic classification. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003923.

Lin:2020:DPP

[LBL^{+}20] Ying Lin, Ling-Yan Bao, Ze-Minghui Li, Shu-Zheng Si, and Chao-Hsien Chu. Differential privacy protection over deep learning: an investigation of its impacted factors. Computers & Security, 99(??):Article 102061, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303345.

Li:2025:HPE

[LBW^{+}25] Zhiyuan Li, Lingbin Bu, Yifan Wang, Qiming Ma, Lin Tan, and Fanliang Bu. Hierarchical perception for encrypted traffic classification via class incremental learning. Computers & Security, 149(??):??, February 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005005.

Lenz:2023:WPR

[LBWV23] Julia Lenz, Zdravko Bozakov, Steffen Wendzel, and Simon Vrhovec. Why people replace their aging smart devices: a push-pull-mooring perspective. Computers & Security, 130 (??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001682.

Li:2021:PTP

[LC21] Jiachun Li and Guoqian Chen. A personalized trajectory privacy protection method. Computers & Security, 108(??):Article 102323, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001474.

Li:2023:SPP

[LCA^{+}23] Jingjin Li, Chao Chen, Mostafa Rahimi Azghadi, Hossein Ghodosi, Lei Pan, and Jun Zhang. Security and privacy problems in voice assistant applications: a survey. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003589.

Li:2022:KKC

[LCB^{+}22] YongGang Li, Yeh-Ching Chung, Yu Bao, Yi Lu, Shan-Qing Guo, and GuoYuan Lin. KPointer: Keep the code pointers on the stack point to the right code. Computers & Security, 120(??):Article 102781, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001766.

Li:2023:WYC

[LCBC23] YongGang Li, JiaZhen Cai, Yu Bao, and Yeh-Ching Chung. What you can read is what you can’t execute. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002870.

Liu:2021:SMS

[LCC^{+}21] Xueqian Liu, Shoufeng Cao, Zhenzhong Cao, Qu Gao, Lin Wan, and Fengyu Wang. SELF: a method of searching for library functions in stripped binary code. Computers & Security, 111(??):Article 102473, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002972.

Lee:2022:CII

[LCC^{+}22] Jung-San Lee, Ying-Chin Chen, Chit-Jie Chew, Chih-Lung Chen, Thu-Nguyet Huynh, and Chung-Wei Kuo. CoNN-IDS: Intrusion detection system based on collaborative neural networks and agile training. Computers & Security, 122(??):Article 102908, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003017.

Li:2024:PCU

[LCCZ24] Junchen Li, Guang Cheng, Zongyao Chen, and Peng Zhao. Protocol clustering of unknown traffic based on embedding of protocol specification. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004856.

Li:2024:CSM

[LCH^{+}24a] Qiao Li, Jing Chen, Kun He, Zijun Zhang, Ruiying Du, Jisi She, and Xinxin Wang. Corrigendum to “Model-agnostic Adversarial Example Detection via High-Frequency Amplification” [Computers & Security, Volume 141, June 2024, 103791]. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003110.

Li:2024:MAA

[LCH^{+}24b] Qiao Li, Jing Chen, Kun He, Zijun Zhang, Ruiying Du, Jisi She, and Xinxin Wang. Model-agnostic adversarial example detection via high-frequency amplification. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000920.

Liang:2020:PPM

[LCL^{+}20] Wenjuan Liang, Hong Chen, Ruixuan Liu, Yuncheng Wu, and Cuiping Li. A Pufferfish privacy mechanism for monitoring web browsing behavior under temporal correlations. Computers & Security, 92(??):Article 101754, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300389.

Lee:2023:VKW

[LCL23] Sunwoo Lee, Wonsuk Choi, and Dong Hoon Lee. The vibration knows who you are! A further analysis on usable authentication for smartwatch users. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004321.

Ling:2024:FIT

[LCL24] Jie Ling, Jinhui Chen, and Honglei Li. FDT: Improving the transferability of adversarial examples with frequency domain transformation. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002475.

Lin:2025:ECS

[LCL^{+}25] Ying-Dar Lin, Wei-Hsiang Chan, Yuan-Cheng Lai, Chia-Mu Yu, Yu-Sung Wu, and Wei-Bin Lee. Enhancing can security with ML-based IDS: Strategies and efficacies against adversarial attacks. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000112.

Li:2023:SPA

[LCP^{+}23] Lin Li, Chao Chen, Lei Pan, Leo Yu Zhang, Zhifeng Wang, Jun Zhang, and Yang Xiang. A survey of PPG’s application in authentication. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300398X.

Lovyagin:2020:FFG

[LCSD20] Nikita Yu. Lovyagin, George A. Chernishev, Kirill K. Smirnov, and Roman Yu. Dayneko. FGACFS: a fine-grained access control for *nix userspace file system. Computers & Security, 88(??):Article 101632, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819301798.

Lin:2025:CTC

[LCTZ25] Yuwei Lin, Yonghong Chen, Hui Tian, and Xiaolong Zhuang. Covert timing channel detection based on isolated binary trees. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005054.

Liang:2022:HHD

[LCW^{+}22] Jianbing Liang, Shuhui Chen, Ziling Wei, Shuang Zhao, and Wei Zhao. HAGDetector: Heterogeneous DGA domain name detection model. Computers & Security, 120(??): Article 102803, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001973.

Li:2025:PSP

[LCX^{+}25] Qingyun Li, Wei Chen, Xiaotang Xu, Yiting Zhang, and Lifa Wu. Precision strike: Precise backdoor attack with dynamic trigger. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004061.

Li:2021:TDI

[LCY^{+}21] Zhenyuan Li, Qi Alfred Chen, Runqing Yang, Yan Chen, and Wei Ruan. Threat detection and investigation with system-level provenance graphs: a survey. Computers & Security, 106(??):Article 102282, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001061.

Li:2022:DDM

[LCZ^{+}22] Ce Li, Zijun Cheng, He Zhu, Leiqi Wang, Qiujian Lv, Yan Wang, Ning Li, and Degang Sun. DMalNet: Dynamic malware analysis based on API feature engineering and graph learning. Computers & Security, 122(??):Article 102872, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002668.

Li:2024:DRS

[LCZ^{+}24] Chao Li, Jian Chen, Zhaoxin Zhang, Zhiping Li, Yanan Cheng, and Chendi Ma. DNS root server resolution anomaly detection. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002517.

Li:2025:FGM

[LCZ^{+}25] Degang Li, Xi Chen, Mingliang Zhu, Qingjun Yuan, and Chunxiang Gu. A fine-grained message clustering method based on message representation and identifier fingerprints. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003207.

Li:2020:BAE

[LCZW20] XuKui Li, Wei Chen, Qianru Zhang, and Lifa Wu. Building Auto-Encoder Intrusion Detection System based on random forest feature selection. Computers & Security, 95(??): Article 101851, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820301231.

Li:2023:CAD

[LCZY23] Chao Li, Yanan Cheng, Zhaoxin Zhang, and Ping Yu. Configuration anormaly detection and resolution risk assessment of authoritative domain name server. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300336X.

Lee:2022:IPA

[LdGW^{+}22] Jaeung Lee, Melchor C. de Guzman, Jingguo Wang, Manish Gupta, and H. Raghav Rao. Investigating perceptions about risk of data breaches in financial institutions: a routine activity-approach. Computers & Security, 121(??):Article 102832, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002267.

Lilhore:2024:CSF

[LDS24] Umesh Kumar Lilhore, Surjeet Dalal, and Sarita Simaiya. A cognitive security framework for detecting intrusions in IoT and 5G utilizing deep learning. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004704.

Liras:2021:FAD

[LdSP21] Luis Francisco Martín Liras, Adolfo Rodríguez de Soto, and Miguel A. Prada. Feature analysis for data-driven APT-related malware discrimination. Computers & Security, 104(??):Article 102202, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000262.

Li:2025:SAA

[LDW25] Tong Li, Fangqi Dong, and Chaoqun Wen. The Security Awareness Adventure: a serious game for security awareness training utilizing a state transition system and a probabilistic model. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001889.

Lee:2023:AWD

[Lee23] In Lee. Analyzing web descriptions of cybersecurity breaches in the healthcare provider sector: a content analytics research method. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000950.

Leszczyna:2021:RCA

[Les21]  Rafał Leszczyna.  Review of cybersecurity assessment

methods: Applicability perspective. Computers & Security, 108(??):Article 102376, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002005.

Luh:2025:GIS

[LET^{+}25] Robert Luh, Sebastian Eresheim, Paul Tavolato, Thomas Petelin, Simon Gmeiner, Andreas Holzinger, and Sebastian Schrittwieser. Gamifying information security: Adversarial risk exploration for IT/OT infrastructures. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824005935.

Lee:2025:MBI

[LFC^{+}25] Jung-San Lee, Yun-Yi Fan, Chia-Hao Cheng, Chit-Jie Chew, and Chung-Wei Kuo. ML-based intrusion detection system for precise APT cyber-clustering. Computers & Security, 149(??): ??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005157.

Li:2021:PMI

[LFCD21] Miles Q. Li, Benjamin C. M. Fung, Philippe Charland, and Steven H. H. Ding. I-MAD: Interpretable malware detector using Galaxy Transformer. Computers & Security, 108(??):Article 102371, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001954.

Liu:2022:PGE

[LFHH22] Zhonglin Liu, Yong Fang, Cheng Huang, and Jiaxuan Han. GraphXSS: an efficient XSS payload detection approach based on graph convolutional network. Computers & Security, 114(??):Article 102597, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100420X.

Liu:2023:MEX

[LFHX23] Zhonglin Liu, Yong Fang, Cheng Huang, and Yijia Xu. MFXSS: an effective XSS vulnerability detection method in JavaScript based on multi-feature model. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822004072.

Li:2023:IAR

[LFHY23] Jiawen Li, Kun Fang, Xiaolin Huang, and Jie Yang. Improving adversarial robustness through a curriculum-guided reliable distillation. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003218.

Li:2025:MDP

[LFL^{+}25] Yunfei Li, Xiaodong Fu, Li Liu, Jiaman Ding, Wei Peng, and Lianyin Jia. Multi-domains personalized local differential privacy frequency estimation mechanism for utility optimization. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005790.

Liu:2022:EMA

[LFW^{+}22] Songsong Liu, Pengbin Feng, Shu Wang, Kun Sun, and Jiahao Cao. Enhancing malware analysis sandboxes with emulated user behavior. Computers & Security, 115(??):Article 102613, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000128.

Liu:2025:MAN

[LFW^{+}25] Taotao Liu, Yu Fu, Kun Wang, Xueyuan Duan, and Qiuhan Wu. A multiscale approach for network intrusion detection based on variance-covariance subspace distance and EQL v2. Computers & Security, 148(??):??, January 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004784.

Liu:2025:SNS

[LGG25] Xiaojian Liu, Xinwei Guo, and Wen Gu. SecKG2vec: a novel security knowledge graph relational reasoning method based on semantic and structural fusion embedding. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004978.

Liu:2021:FNI

[LGH21] Jingmei Liu, Yuanbo Gao, and Fengjie Hu. A fast network intrusion detection system using adaptive synthetic oversampling and LightGBM. Computers & Security, 106(??):Article 102289, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001139.

Liu:2021:EMP

[LGJW21] Dong-Jie Liu, Guang-Gang Geng, Xiao-Bo Jin, and Wei Wang. An efficient multistage phishing website detection model based on the CASE feature framework: Aiming at the real web environment. Computers & Security, 110(??):Article 102421, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002455.

Li:2021:DPH

[LGL21] Shuyu Li, Yue Geng, and Yingle Li. A differentially private hybrid decomposition algorithm based on quad-tree. Computers & Security, 109(??):Article 102384, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100208X.

Long:2022:SAA

[LGXZ22] Teng Long, Qi Gao, Lili Xu, and Zhangbing Zhou. A survey on adversarial attacks in computer vision: Taxonomy, visualization and future directions. Computers & Security, 121(??):Article 102847, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002413.

Li:2023:GGA

[LGZD23] Kun Li, Wei Guo, Fan Zhang, and Jiayu Du. GAMBD: Generating adversarial malware against MalConv. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482300189X.

Levi:2022:BEC

[LHAE22] Matan Levi, Itay Hazan, Noga Agmon, and Sagi Eden. Behavioral embedding for continuous user verification in global settings. Computers & Security, 119(??):Article 102716, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001122.

Li:2024:ADM

[LHC24] Lingzi Li, Cheng Huang, and Junren Chen. Automated discovery and mapping ATT&CK tactics and techniques for unstructured cyber threat intelligence. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001160.

Li:2023:SAC

[LHD^{+}23] Zhengfa Li, Chuanhe Huang, Shuhua Deng, Wanyu Qiu, and Xieping Gao. A soft actor-critic reinforcement learning algorithm for network intrusion detection. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004121.

Lin:2024:CNS

[LHG^{+}24] Xinjie Lin, Longtao He, Gaopeng Gou, Jing Yu, Zhong Guan, Xiang Li, Juncheng Guo, and Gang Xiong. CETP: a novel semi-supervised framework based on contrastive pre-training for imbalanced encrypted traffic classification. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001949.

LaFleur:2021:TPS

[LHGB21] Claire La Fleur, Blaine Hoffman, C. Benjamin Gibson, and Norbou Buchler. Team performance in a series of regional and national US cybersecurity defense competitions: Generalizable effects of training and functional role specialization. Computers & Security, 104(??):Article 102229, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000535.

Liu:2020:SES

[LHH^{+}20] Jianghua Liu, Jingyu Hou, Xinyi Huang, Yang Xiang, and Tianqing Zhu. Secure and efficient sharing of authenticated energy usage data with privacy preservation. Computers & Security, 92(??):Article 101756, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300407.

Lei:2023:UPP

[LHH23] Wenjing Lei, Siqi Hu, and Carol Hsu. Unveiling the process of phishing precautions taking: the moderating role of optimism bias. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001591.

Li:2024:SAE

[LHL^{+}24a] Junhao Li, Junjiang He, Wenshan Li, Wenbo Fang, Geying Yang, and Tao Li. SynDroid: an adaptive enhanced Android malware classification method based on CTGAN-SVM. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300514X.

Li:2024:ISH

[LHL^{+}24b] Yufeng Li, Chengjian Huang, Qi Liu, Xiangyu Zheng, and Ke Sun. Integrating security in hazard analysis using STPA-Sec and GSPN: a case study of automatic emergency braking system. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001925.

Luo:2025:DFD

[LHO^{+}25] Yuling Luo, Tinghua Hu, Xue Ouyang, Junxiu Liu, Qiang Fu, Sheng Qin, Zhen Min, and Xiaoguang Lin. DPO-Face: Differential privacy obfuscation for facial sensitive regions. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001233.

Liu:2021:RMS

[LHS21] Zhen Liu, Changzhen Hu, and Chun Shan. Riemannian manifold on stream data: Fourier transform and entropybased DDoS attacks detection method. Computers & Security, 109(??):Article 102392, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002169.

Lu:2020:STI

[LHW^{+}20] Di Lu, Ruidong Han, Yue Wang, Yongzhi Wang, Xuewen Dong, Xindi Ma, Teng Li, and Jianfeng Ma. A secured TPM integration scheme towards smart embedded system based collaboration network. Computers & Security, 97(??):Article 101922, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301991.

Li:2023:VAA

[LHX23] Jiachun Li, Yuchao Hu, and Fei Xia. A variable adversarial attack method based on filtering. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003413.

Loft:2022:CAE

[LHYW22] Paul Loft, Ying He, Iryna Yevseyeva, and Isabel Wagner. CAESAR8: an agile enterprise architecture approach to managing information security risks. Computers & Security, 122(??):Article 102877, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002711.

Liang:2024:DAM

[LHZZ24] Chuang Liang, Jie Huang, Zeping Zhang, and Shuaishuai Zhang. Defending against model extraction attacks with OOD feature learning and decision boundary confusion. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300473X.

Long:2022:HME

[LJ22] Zhang Long and Wang Jinsong. A hybrid method of entropy and SSAE–SVM based DDoS detection and mitigation mechanism in SDN. Computers & Security, 115(??):Article 102604, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000037.

Liu:2022:TAE

[LJJ22] Yujia Liu, Ming Jiang, and Tingting Jiang. Transferable adversarial examples based on global smooth perturbations. Computers & Security, 121(??):Article 102816, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002103.

Liu:2025:FFD

[LJL^{+}25] Xiaosheng Liu, Wenqi Jiang, Zhongwei Li, Xianji Jin, Zihan Ma, and Qingyang Li. FuzzAGG: a fuzzing-driven attack graph generation framework for industrial robot systems. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005297.

Lu:2020:BCS

[LJO^{+}20] Qian Lu, Ruobing Jiang, Yuzhan Ouyang, Haipeng Qu, and Jiahui Zhang. BiRe: a client-side bi-directional SYN reflection mechanism against multi-model evil twin attacks. Computers & Security, 88(??):Article 101618, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301658.

Lee:2020:LGC

[LK20] Claire Seungeun Lee and Ji Hye Kim. Latent groups of cybersecurity preparedness in Europe: Sociodemographic factors and country-level contexts. Computers & Security, 97(??):Article 101995, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302686.

Lee:2021:DNN

[LKA^{+}21] Hongkyu Lee, Jeehyeong Kim, Seyoung Ahn, Rasheed Hussain, Sunghyun Cho, and Junggab Son. Digestive neural networks: a novel defense strategy against inference attacks in federated learning. Computers & Security, 109(??):Article 102378, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002029.

Lee:2025:SPE

[LKK25] Chanjong Lee, Bedeuro Kim, and Hyoungshick Kim. The silence of the phishers: Early-stage voice phishing detection with runtime permission requests. Computers & Security, 152(??): ??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000537.

Lee:2023:VVK

[LKL23] Jiyeon Lee, Hyosu Kim, and Kilho Lee. VRKeyLogger: Virtual keystroke inference attack via eavesdropping controller usage pattern in WebVR. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003711.

Liu:2024:DAA

[LKLZ24] Liang Liu, Xinyu Kuang, Lin Liu, and Lei Zhang. Defend against adversarial attacks in malware detection through attack space management. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001421.

Liang:2023:DCP

[LKS^{+}23] Xueping Liang, Charalambos Konstantinou, Sachin Shetty, Eranga Bandara, and Ruimin Sun. Decentralizing cyber physical systems for resilience: an innovative case study from a cybersecurity perspective. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003455.

Lee:2020:CSA

[LKW^{+}20] Seungsoo Lee, Jinwoo Kim, Seungwon Woo, Changhoon Yoon, Sandra Scott-Hayward, Vinod Yegneswaran, Phillip Porras, and Seungwon Shin. A comprehensive security assessment framework for software-defined networks. Computers & Security, 91(??):Article 101720, April 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300079.

Li:2021:IBM

[LL21a] Xintong Li and Qi Li. An IRL-based malware adversarial generation method to evade anti-malware engines. Computers & Security, 104(??):Article 102118, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303916.

Lin:2021:PBS

[LL21b] Hui-Tang Lin and Yuan-Yi Liang. A PUF-based secure wake-up scheme for Internet of Things. Computers & Security, 110(??):Article 102415, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100239X.

Lee:2022:ISD

[LL22a] Junwon Lee and Heejo Lee. Improving SSH detection model using IPA time and WGAN-GP. Computers & Security, 116(??):Article 102672, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000712.

Li:2022:PPA

[LL22b] Zhen Li and Qi Liao. Preventive portfolio against dataselling ransomware — a game theory of encryption and deception. Computers & Security, 116(??):Article 102644, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000438.

Liu:2023:PCS

[LL23] Xiaojian Liu and Kehong Liu. A permission-carrying security policy and static enforcement for information flows in Android programs. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004825.

Li:2025:IIH

[LL25] Zhen Li and Qi Liao. To insure or not to insure: How attackers exploit cyber-insurance via game theory. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002743.

Li:2023:WOK

[LLC^{+}23] Li Li, Tianfeng Li, Hua Cai, Jian Zhang, and Jianjun Wang. I will only know after using it: the repeat purchasers of smart home appliances and the privacy paradox problem. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000664.

Li:2025:AMG

[LLD^{+}25] Heqing Li, Xinde Li, Fir Dunkin, Zhentong Zhang, and Xiaoyan Lu. Adaptive multi-granularity trust management scheme for UAV visual sensor security under adversarial attacks. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004139.

Li:2023:AOS

[LLF^{+}23] Xiangjuan Li, Yang Li, Zhaowen Feng, Zhaoxuan Wang, and Quan Pan. ATS-O2A: a state-based adversarial attack strategy on deep reinforcement learning. Computers & Security, 129 (??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001694.

Liu:2020:USA

[LLG20]  Hui Liu, Juanru Li, and Dawu Gu.  Understanding

the security of app-in-the-middle IoT. Computers & Security, 97(??):Article 102000, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482030273X.

Lyvas:2021:AAH

[LLG21] Christos Lyvas, Costas Lambrinoudakis, and Dimitris Geneiatakis. On Android’s activity hijacking prevention. Computers & Security, 111(??):Article 102468, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002923.

Liu:2025:RSS

[LLH^{+}25] Hongle Liu, Ming Liu, Lansheng Han, Haili Sun, and Cai Fu. Ripple2Detect: a semantic similarity learning based framework for insider threat multi-step evidence detection. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000768.

Li:2021:SES

[LLJ21] Li Li, Jiayong Liu, and Peng Jia. SecTEP: Enabling secure tender evaluation with sealed prices and quality evaluation in procurement bidding systems over blockchain. Computers & Security, 103(??):Article 102188, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000122.

Li:2023:SUS

[LLJ^{+}23] Munan Li, Hongbo Liu, Xiangdong Jiang, Zheng Zhao, and Tianhao Zhang. SENSE: an unsupervised semantic learning model for cross-platform vulnerability search. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004108.

Lee:2023:MLB

[LLJL23] Sun-Jin Lee, Yu-Rim Lee, So-Eun Jeon, and Il-Gu Lee. Machine learning-based jamming attack classification and effective defense technique. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000792.

Li:2020:DCS

[LLL^{+}20a] Fenghua Li, Yongjun Li, Siyuan Leng, Yunchuan Guo, Kui Geng, Zhen Wang, and Liang Fang. Dynamic countermeasures selection for multi-path attacks. Computers & Security, 97(??):Article 101927, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302030.

Li:2020:APR

[LLL^{+}20b] Tun Li, Yutian Liu, Yanbing Liu, Yunpeng Xiao, and Nang An Nguyen. Attack plan recognition using hidden Markov and probabilistic inference. Computers & Security, 97(??):Article 101974, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302479.

Lan:2022:DNS

[LLL^{+}22a] Jinghong Lan, Xudong Liu, Bo Li, Yanan Li, and Tongtong Geng. DarknetSec: a novel self-attentive deep learning method for darknet traffic classification and application identification. Computers & Security, 116(??):Article 102663, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000621.

Lan:2022:MMT

[LLL^{+}22b] Jinghong Lan, Xudong Liu, Bo Li, Jie Sun, Beibei Li, and Jun Zhao. MEMBER: a multi-task learning model with hybrid deep features for network intrusion detection. Computers & Security, 123(??):Article 102919, December 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200311X.

Li:2022:NDF

[LLL^{+}22c] Ce Li, Qiujian Lv, Ning Li, Yan Wang, Degang Sun, and Yuanyuan Qiao. A novel deep framework for dynamic malware detection based on API sequence intrinsic features. Computers & Security, 116(??):Article 102686, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000840.

Lin:2022:PPC

[LLL^{+}22d] Xinjie Lin, Han Liu, Zhen Li, Gang Xiong, and Gaopeng Gou. Privacy protection of China’s top websites: a multilayer privacy measurement via network behaviours and privacy policies. Computers & Security, 114(??):Article 102606, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000050.

Lai:2023:TPD

[LLL^{+}23a] Yuan-Cheng Lai, Jheng-Yan Lin, Ying-Dar Lin, Ren-Hung Hwang, Po-Chin Lin, Hsiao-Kuang Wu, and Chung-Kuan Chen. Two-phase defense against poisoning attacks on federated learning-based intrusion detection. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001153.

Li:2023:TPD

[LLL^{+}23b] Yuzhe Li, Yong Liu, Bo Li, Weiping Wang, and Nan Liu. Towards practical differential privacy in data analysis: Understanding the effect of epsilon on utility in private ERM. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000573.

Lu:2023:TSS

[LLL^{+}23c] Shiwei Lu, Ruihu Li, Wenbin Liu, Chaofeng Guan, and Xiaopeng Yang. Top-k sparsification with secure aggregation for privacy-preserving federated learning. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003856.

Luo:2023:SFA

[LLL23d] Shiyao Luo, Yingxu Lai, and Jing Liu. Selective forwarding attack detection and network recovery mechanism based on cloud-edge cooperation in software-defined wireless sensor network. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004758.

Li:2025:RFL

[LLL^{+}25] Yong Li, TongTong Liu, HaiChao Ling, Wei Du, and XiangLin Ren. A robust federated learning algorithm for partially trusted environments. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004668.

Lima:2022:BCB

[LLLB22] Milton Lima, Ricardo Lima, Fernando Lins, and Michel Bonfim. Beholder — a CEP-based intrusion detection and prevention systems for IoT environments. Computers & Security, 120(??):Article 102824, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002188.

Lu:2022:DAB

[LLLC22] Shiwei Lu, Ruihu Li, Wenbin Liu, and Xuan Chen. Defense against backdoor attack in federated learning. Computers & Security, 121(??):Article 102819, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002139.

Liu:2020:NMM

[LLLZ20] Xinbo Liu, Yaping Lin, He Li, and Jiliang Zhang. A novel method for malware detection on ML-based visualization technique. Computers & Security, 89(??):Article 101682, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818314627.

Li:2021:ESC

[LLLZ21] Ding Li, Wei Lin, Bin Lu, and Yuefei Zhu. Exploiting sidechannel leaks in web traffic of incremental search. Computers & Security, 111(??):Article 102481, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003059.

Lu:2025:HTF

[LLP25] Chuan Lu, Senlin Luo, and Limin Pan. High-trigger fuzz testing for microarchitectural speculative execution vulnerability. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002561.

Lu:2025:AAB

[LLPL25] Hongyu Lu, Jiajia Liu, Jimin Peng, and Jiazhong Lu. Adversarial attacks based on time-series features for traffic detection. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004802.

Li:2020:PPS

[LLX^{+}20] Dong Li, Xiaofeng Liao, Tao Xiang, Jiahui Wu, and Junqing Le. Privacy-preserving self-serviced medical diagnosis scheme based on secure multi-party computation. Computers & Security, 90(??):Article 101701, March 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740481930238X.

Liu:2023:LFS

[LLX^{+}23] Jiyuan Liu, Bingyi Lu, Mingkang Xiong, Tao Zhang, and Huilin Xiong. Low frequency sparse adversarial attack. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002894.

Liu:2025:GUA

[LLX25] Dengbo Liu, Zhi Li, and Daoyun Xu. Generate universal adversarial perturbations by shortest-distance soft maximum direction attack. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004735.

Li:2021:NWS

[LLYL21] Chenxi Li, Jia Li, Jiahai Yang, and Jinlei Lin. A novel workload scheduling framework for intrusion detection system in NFV scenario. Computers & Security, 106(??):Article 102271, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100095X.

Lu:2022:PEP

[LLZJ22] Qian Lu, Shihao Li, Jiahui Zhang, and Ruobing Jiang. PEDR: Exploiting phase error drift range to detect fullmodel rogue access point attacks. Computers & Security, 114(??):Article 102581, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821004041.

Liu:2020:FSP

[LLZY20] Yizhong Liu, Jianwei Liu, Zongyang Zhang, and Hui Yu. A fair selection protocol for committee-based permissionless blockchains. Computers & Security, 91(??):Article 101718, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300055.

Li:2023:MHD

[LM23] Yuying Li and Rogemar Mamon. Modelling health-data breaches with application to cyber insurance. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822003558.

Lampe:2024:CTT

[LM24] Brooke Lampe and Weizhi Meng. can-train-and-test: a curated CAN dataset for automotive intrusion detection. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000786.

Li:2022:FSI

[LMD^{+}22] Kehong Li, Wengang Ma, Huawei Duan, Han Xie, and Juanxiu Zhu. Few-shot IoT attack detection based on RFP-CNN and adversarial unsupervised domain-adaptive regularization. Computers & Security, 121(??):Article 102856, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002504.

Liu:2025:CAW

[LMW^{+}25] Danjun Liu, Xuan Meng, Pengfei Wang, Xu Zhou, Wei Xie, and Baosheng Wang. Constructing arbitrary write via puppet objects and delivering gadgets in Linux kernel. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004942.

Lin:2022:SDS

[LMZZ22] Yu Lin, Yunlong Mao, Yuan Zhang, and Sheng Zhong. Secure deduplication schemes for content delivery in mobile edge computing. Computers & Security, 114(??):Article 102602, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000013.

Lykousas:2024:DDP

[LP24] Nikolaos Lykousas and Constantinos Patsakis. Decoding developer password patterns: a comparative analysis of password extraction and selection practices. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002797.

Li:2025:SSS

[LPL25] Jiaxin Li, Haipeng Peng, and Lixiang Li. Sublinear smart semantic search based on knowledge graph over encrypted database. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000082.

Louca:2023:NET

[LPS23] Constantinos Louca, Adamantini Peratikou, and Stavros Stavrou. A novel Evil Twin MiTM attack through 802.11v protocol exploitation. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001712.

Liu:2025:DCU

[LPW25] Liang Liu, Silin Peng, and Zhijun Wu. Detection of CIFA using SMOTEBoost and LSTM in NDN. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005571.

Lu:2020:IVS

[LPX20] Xiuqing Lu, Zhenkuan Pan, and Hequn Xian. An integrity verification scheme of cloud storage for Internet-of-Things mobile terminal devices. Computers & Security, 92(??):Article 101686, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740481930224X.

Liu:2024:SEF

[LPZF24] Side Liu, Guojun Peng, Haitao Zeng, and Jianming Fu. A survey on the evolution of fileless attacks and detection techniques. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300562X.

Li:2022:HHD

[LQH^{+}22a] Yanan Li, Tao Qin, Yongzhong Huang, Jinghong Lan, Zan-Hao Liang, and Tongtong Geng. HDFEF: a hierarchical and dynamic feature extraction framework for intrusion detection systems. Computers & Security, 121(??):Article 102842, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200236X.

Liu:2022:KEI

[LQH22b] Yuanzhen Liu, Umair Mujtaba Qureshi, and Gerhard Petrus Hancke. Keypad entry inference with sensor fusion from mobile and smart wearables. Computers & Security, 121(??): Article 102837, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822002310.

Li:2025:ICS

[LQHZ25] Jinyuan Li, Guangyu Qian, Wei He, and Wei Zhang. Industrial control system intrusion detection method based on belief rule base with gradient descent. Computers & Security, 155(??): ??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001762.

Liu:2025:IDS

[LQL^{+}25a] Qingxin Liu, Guihe Qin, Yanhua Liang, Jiaru Song, Wanning Liu, and Xue Zhou. Intrusion detection system for autonomous vehicles using sensor spatio-temporal information. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001907.

Liu:2025:GMB

[LQL25b] Xiaojian Liu, Zilin Qin, and Kehong Liu. μGAN: a mutation-based cost optimal adversarial malware generation approach against evolving Android malware variants. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003840.

Lian:2020:PPS

[LQY^{+}20] Huijuan Lian, Weidong Qiu, Di Yan, Jie Guo, Zhe Li, and Peng Tang. Privacy-preserving spatial query protocol based on the Moore curve for location-based service. Computers & Security, 96(??):Article 101845, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301188.

Liu:2021:LRD

[LRH^{+}21] Xinqian Liu, Jiadong Ren, Haitao He, Qian Wang, and Chen Song. Low-rate DDoS attacks detection method using data compression and behavior divergence measurement. Computers & Security, 100(??):Article 102107, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303801.

Liu:2022:IBS

[LRL^{+}22] Zhenpeng Liu, Lele Ren, Ruilin Li, Qiannan Liu, and Yonggang Zhao. ID-based sanitizable signature data integrity auditing scheme with privacy-preserving. Computers & Security, 121(??):Article 102858, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002528.

Liang:2020:OBA

[LS20] Yuting Liang and Reza Samavi. Optimization-based kanonymity algorithms. Computers & Security, 93(??):Article 101753, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300377.

Liu:2021:SAB

[LS21] Tian Liu and Tao Shu. On the security of ANN-based AC state estimation in smart grid. Computers & Security, 105 (??):Article 102265, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000894.

Lee:2025:ECB

[LS25] Rong-Teng Lee and Shan-Hsiang Shen. Erasure code backup system for data security. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000033.

Ling:2021:SFN

[LSAH21] Yeong Tyng Ling, Nor Fazlida Mohd Sani, Mohd. Taufik Abdullah, and Nor Asilah Wati Abdul Hamid. Structural features with nonnegative matrix factorization for metamorphic malware detection. Computers & Security, 104(??):Article 102216, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000407.

Li:2024:ADA

[LSG^{+}24] Huang Li, Yiqin Sang, Hongjuan Ge, Jie Yan, and Shijia Li. Anomaly detection of aviation data bus based on SAE and IMD. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005291.

Lee:2022:PRR

[LsJC^{+}22] Seungkwang Lee, Nam su Jho, Doyoung Chung, Yousung Kang, and Myungchul Kim. Rcryptect: Real-time detection of cryptographic function in the user-space filesystem. Computers & Security, 112(??):Article 102512, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003369.

Larriba:2020:TAE

[LSL20] Antonio M. Larriba, José M. Sempere, and Damián López. A two authorities electronic vote scheme. Computers & Security, 97(??):Article 101940, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302169.

Liu:2024:CDP

[LSLL24] Ming Liu, Xiao Song, Yong Li, and Wenxin Li. Correlated differential privacy based logistic regression for supplier data protection. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004522.

Lazarov:2025:PCP

[LSMK25] Willi Lazarov, Pavel Seda, Zdenek Martinasek, and Roman Kummel. Penterep: Comprehensive penetration testing with adaptable interactive checklists. Computers & Security, 154 (??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000884.

Lallie:2021:CSA

[LSN^{+}21] Harjinder Singh Lallie, Lynsay A. Shepherd, Jason R. C. Nurse, Arnau Erola, Gregory Epiphaniou, Carsten Maple, and Xavier Bellekens. Cyber security in the age of COVID-19: a timeline and analysis of cyber-crime and cyber-attacks during the pandemic. Computers & Security, 105(??):Article 102248, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000729.

Li:2024:GGA

[LSTL24] Yong Li, Xiao Song, Yuchun Tu, and Ming Liu. GAP-BAS: Genetic algorithm-based privacy budget allocation strategy in differential privacy K-means clustering algorithm. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823006077.

Landauer:2020:SLC

[LSWR20] Max Landauer, Florian Skopik, Markus Wurzenberger, and Andreas Rauber. System log clustering approaches for cyber security applications: a survey. Computers & Security, 92(??):Article 101739, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820300250.

Lu:2022:DSC

[LSXJ22] Jintian Lu, Jiakun Sun, Ruizhi Xiao, and Shuyuan Jin. DIFCS: a secure cloud data sharing approach based on decentralized information flow control. Computers & Security, 117(??):Article 102678, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000761.

Lakshmi:2025:EDI

[LT25] Vimitha R. Vidhya Lakshmi and Gireesh Kumar T. Enhancing data integrity in opportunistic mobile social network: Leveraging Berkle Tree and secure data routing against attacks. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004383.

Li:2025:VFD

[LTCR25] Qindong Li, Wenyi Tang, Xingshu Chen, and Hao Ren. Vul_{diff} Finder: Discovering inconsistencies in unstructured vulnerability information. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001361.

Lai:2022:IMN

[LTL^{+}22] Yingxu Lai, Liyao Tong, Jing Liu, Yipeng Wang, Tong Tang, Zijian Zhao, and Hua Qin. Identifying malicious nodes in wireless sensor networks based on correlation detection. Computers & Security, 113(??):Article 102540, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003643.

Lee:2021:TCS

[LTU^{+}21] Ming Jie Lee, Andrew Beng Jin Teoh, Andreas Uhl, Shiuan-Ni Liang, and Zhe Jin. A tokenless cancellable scheme for multimodal biometric systems. Computers & Security, 108(??):Article 102350, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001747.

Li:2023:DDC

[LTZ23] Wenjuan Li, Jiao Tan, and Nan Zhu. Design of doublecross-based smartphone unlock mechanism. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823001141.

Li:2024:TMM

[LTZ^{+}24] Wanyu Li, Hailiang Tang, Hailin Zhu, Wenxiao Zhang, and Chen Liu. TS-Mal: Malware detection model using temporal and structural features learning. Computers & Security, 140 (??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000531.

Li:2025:DDB

[LWC^{+}25] Jiaqi Li, Ke Wang, Yaoguang Chen, Yajin Zhou, Lei Wu, and Jiashui Wang. Detecting DBMS bugs with contextsensitive instantiation and multi-plan execution. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002536.

Liang:2025:SSC

[LWD^{+}25] Chen Liang, Qiang Wei, Jiang Du, Yisen Wang, and Zirui Jiang. Survey of source code vulnerability analysis based on deep learning. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004036.

Liu:2024:CET

[LWH^{+}24] Wanping Liu, Baojuan Wang, Dong Huang, Haolan Luo, and Ling Lu. CFOA: Exploring transferable adversarial examples by content feature optimization with attention guidance. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001834.

Liu:2021:SGN

[LWL21] Ao Liu, Yunpeng Wang, and Tao Li. SFE-GACN: a novel unknown attack detection under insufficient data via intra categories generation in embedding space. Computers & Security, 105(??):Article 102262, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000869.

Liu:2023:BLP

[LWL^{+}23] Xueyan Liu, Jing Wang, Qiong Liu, Xin Xiong, and Shufen Niu. Bidirectional location privacy protection scheme for epidemiological investigation based on OT. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003632.

Li:2025:IVS

[LWL^{+}25] Fang Li, Gang Wang, Guangjun Liu, Xiao Xue, and Deyu Zhou. Integrity verification scheme for distributed dynamic data in service ecosystems. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003608.

Luo:2025:ADM

[LWLL25] Feng Luo, Jiajia Wang, Zhihao Li, and Cheng Luo. An anomaly detection model for in-vehicle networks based on lightweight convolution with spectral residuals. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824006102.

Luo:2021:ABA

[LWLT21] Peng Luo, Buhong Wang, Tengyao Li, and Jiwei Tian. ADS-B anomaly data detection model based on VAE-SVDD. Computers & Security, 104(??):Article 102213, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000377.

Luo:2025:TSC

[LWO^{+}25] Yuling Luo, Yali Wan, Xue Ouyang, Junxiu Liu, Qiang Fu, Sheng Qin, Ziqi Yuan, and Tinghua Hu. Time series correlated key–value data collection with local differential privacy. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002998.

Liu:2025:LLD

[LWP^{+}25] Zhen Liu, Ruoyu Wang, Bitao Peng, Lingyu Qiu, Qingqing Gan, Changji Wang, and Wenbin Zhang. LDCDroid: Learning data drift characteristics for handling the model aging problem in Android malware detection. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482400600X.

Li:2020:DTA

[LWS^{+}20] Tengyao Li, Buhong Wang, Fute Shang, Jiwei Tian, and Kunrui Cao. Dynamic temporal ADS-B data attack detection based on sHDP-HMM. Computers & Security, 93(??):Article 101789, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300742.

Lv:2025:DCI

[LWS^{+}25] Fei Lv, Hangyu Wang, Rongkang Sun, Zhiwen Pan, Shuaizong Si, Meng Zhang, Weidong Zhang, Shichao Lv, and Limin Sun. Detection of cyberattack in Industrial Control Networks using multiple adaptive local kernel learning. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004577.

Li:2024:TLL

[LWSW24] Sicong Li, Jian Wang, Yafei Song, and Shuo Wang. TriCh-LKRepNet: a large kernel convolutional malicious code classification network for structure reparameterisation and triplechannel mapping. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002426. See retraction

notice [LWSW25].

Li:2025:RNT

[LWSW25] Sicong Li, Jian Wang, Yafei Song, and Shuo Wang. Retraction notice to “TriCh-LKRepNet: a large kernel convolutional malicious code classification network for structure reparameterisation and triple-channel mapping” [Computers & Security 144 (2024) 103937]. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824005133. See [LWSW24].

Luo:2024:TCR

[LWT24] Peng Luo, Buhong Wang, and Jiwei Tian. TTSAD: TCN-Transformer-SVDD model for anomaly detection in air traffic ADS-B data. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400141X.

Lin:2021:RSC

[LWW^{+}21] Hui Lin, Yan Wo, Yuanlu Wu, Ke Meng, and Guoqiang Han. Robust source camera identification against adversarial attacks. Computers & Security, 100(??):Article 102079, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303527.

Li:2022:DDA

[LWX^{+}22] Bin Li, Yijie Wang, Kele Xu, Li Cheng, and Zhiquan Qin. DFAID: Density-aware and feature-deviated active intrusion detection over network traffic streams. Computers & Security, 118(??):Article 102719, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001109.

Li:2024:FPM

[LWYG24] Xuan Li, Naiyu Wang, Shuai Yuan, and Zhitao Guan. FedIMP: Parameter I mportance-based M odel Poisoning attack against Federated learning system. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002414.

Liu:2022:TAD

[LWZ22] Tao Liu, He Wang, and Yuqing Zhang. A traffic anomaly detection scheme for non-directional denial of service attacks in software-defined optical network. Computers & Security, 112(??):Article 102467, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002911.

Ling:2023:AAA

[LWZ^{+}23] Xiang Ling, Lingfei Wu, Jiangyu Zhang, Zhenqing Qu, Wei Deng, Xiang Chen, Yaguan Qian, Chunming Wu, Shouling Ji, Tianyue Luo, Jingzheng Wu, and Yanjun Wu. Adversarial attacks against Windows PE malware detection: a survey of the state-of-the-art. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000445.

Lin:2024:HPB

[LWZ^{+}24] Peihong Lin, Pengfei Wang, Xu Zhou, Wei Xie, Kai Lu, and Gen Zhang. HyperGo: Probability-based directed hybrid fuzzing. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001524.

Luo:2025:WQM

[LWZ^{+}25] Meng Luo, Qiuyun Wang, Jianrong Zhang, Cheng Yu, Kai Zhou, Baojiang Cui, and Zhengwei Jiang. Who are querying for me? measuring the dependency and centralization in recursive resolution. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001890.

Liang:2023:FDT

[LWZC23] Jianbing Liang, Suxia Wang, Shuang Zhao, and Shuhui Chen. FECC: DNS tunnel detection model based on CNN and clustering. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000421.

Li:2025:PVT

[LWZL25] Yunpeng Li, Wei Wu, Yuan Zhang, and Chuanyang Li. Palm vein template protection scheme for resisting similarity attack. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005339.

Lin:2021:MDB

[LX21] Jing Lin and Kaiqi Xiong. Mahalanobis distance-based robust approaches against false data injection attacks on dynamic power state estimation. Computers & Security, 108(??):Article 102326, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001504.

Liang:2025:NFI

[LX25] Xueping Liang and Yilin Xu. A novel framework to identify cybersecurity challenges and opportunities for organizational digital transformation in the cloud. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825000288.

Liang:2020:FIV

[LXC^{+}20] Hongliang Liang, Zhuosi Xie, Yixiu Chen, Hua Ning, and Jianli Wang. FIT: Inspect vulnerabilities in cross-architecture firmware by deep learning and bipartite matching. Computers & Security, 99(??):Article 102032, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303059.

Lin:2023:TPE

[LXH^{+}23] Zhutian Lin, Xi Xiao, Guangwu Hu, Qing Li, Bin Zhang, and Xiapu Luo. Tracking phishing on Ethereum: Transaction network embedding approach for accounts representation learning. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003899.

Lin:2025:NDF

[LXH^{+}25] Shengrui Lin, Shaowei Xu, Binjie He, Hongyan Liu, Dezhang Kong, Xiang Chen, Dong Zhang, Chunming Wu, Ming Li, Xuan Liu, Yuqin Wu, and Muhammad Khurram Khan. NDIF: a distributed framework for efficient in-network neural network inference. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002822.

Lin:2025:REN

[LXJ^{+}25] Wenfeng Lin, Fangliang Xu, Zhiyuan Jiang, Gang Yang, Zhiwei Li, and Chaojing Tang. RPFUZZ: Efficient network service fuzzing via pruning redundant mutation. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003736.

Li:2025:DAI

[LXLL25] Jiacheng Li, Yang Xiao, Shuhui Li, and Tieshan Li. Designing accountable IoT systems to overcome IoT storage limitation. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004231.

Li:2024:LMK

[LXS^{+}24] Xiang Li, Jiang Xie, Qige Song, Yafei Sang, Yongzheng Zhang, Shuhao Li, and Tianning Zang. Let model keep evolving: Incremental learning for encrypted traffic classification. Computers & Security, 137(??):??, February 2024.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005345.

Lyu:2024:TTA

[LXW^{+}24] Qiuyun Lyu, Huihui Xie, Wei Wang, Yanyu Cheng, Yongqun Chen, and Zhen Wang. TFAN: a Task-adaptive Feature Alignment Network for few-shot website fingerprinting attacks on Tor. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002852.

Li:2025:FJJ

[LXW^{+}25a] Jizhe Li, Haoran Xu, Yongjun Wang, Zhiyuan Jiang, Huang Chun, Peidai Xie, Yongxin Chen, and Tian Xia. Fuzzing JavaScript JIT compilers with a high-quality differential test oracle. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003499.

Lin:2025:EET

[LXW^{+}25b] Wanshuang Lin, Chunhe Xia, Tianbo Wang, Mengyao Liu, and Yang Li. Enhancing encrypted traffic analysis via source APIs: a robust approach for malicious traffic detection. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002184.

Li:2025:NSM

[LXX^{+}25] Zehuan Li, Meihua Xiao, Yangping Xu, Fangping Chen, Huaibin Shao, and Sufen Yan. A novel symbolic model for analyzing Internet of Things protocols based on event logic. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001294.

Liu:2024:ASA

[LXY^{+}24] Yaru Liu, Lijuan Xu, Shumian Yang, Dawei Zhao, and Xin Li. Adversarial sample attacks and defenses based on LSTM-ED in industrial control systems. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000518.

Liu:2022:PAN

[LXZ^{+}22] Yaqun Liu, Changyou Xing, Guomin Zhang, Lihua Song, and Hongxiu Lin. AntiTomo: Network topology obfuscation against adversarial tomography-based topology inference. Computers & Security, 113(??):Article 102570, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003941.

Li:2023:CDV

[LXZ^{+}23] Xin Li, Yang Xin, Hongliang Zhu, Yixian Yang, and Yuling Chen. Cross-domain vulnerability detection using graph embedding and domain adaptation. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004096.

Liang:2020:GTS

[LYD20] Xueqin Liang, Zheng Yan, and Robert H. Deng. Game theoretical study on client-controlled cloud data deduplication. Computers & Security, 91(??):Article 101730, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300171.

Li:2024:CTC

[LYHC24] Xiaohui Li, Xiang Yang, Yizhao Huang, and Yue Chen. Combating temporal composition inference by high-order camouflaged network topology obfuscation. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824002864.

Li:2023:LCD

[LYL^{+}23] Yufeng Li, Fengyu Yang, Qi Liu, Jiangtao Li, and Chenhong Cao. Light can be dangerous: Stealthy and effective physical-world adversarial attack by Spot Light. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002559.

Lim:2024:FGA

[LYLT24] Willone Lim, Kelvin Sheng Chek Yong, Bee Theng Lau, and Colin Choon Lin Tan. Future of generative adversarial networks (GAN) for anomaly detection in network security: a review. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000348.

Lin:2024:TSM

[LYS^{+}24] Ying-Dar Lin, Shin-Yi Yang, Didik Sudyana, Fietyata Yudha, Yuan-Cheng Lai, and Ren-Hung Hwang. Two-stage multidatasource machine learning for attack technique and lifecycle detection. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001603.

Li:2023:EFL

[LYSB23] Yanli Li, Dong Yuan, Abubakar Sadiq Sani, and Wei Bao. Enhancing federated learning robustness in adversarial environment through clustering Non-IID features. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002298.

Li:2021:LPP

[LYW^{+}21] Kunchang Li, Yifan Yang, Shuhao Wang, Runhua Shi, and Jianbin Li. A lightweight privacy-preserving and sharing scheme with dual-blockchain for intelligent pricing system of smart grid. Computers & Security, 103(??):Article 102189, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000134.

Lv:2020:ASB

[LZ20] Denglong Lv and Shibing Zhu. Achieving secure big data collection based on trust evaluation and true data discovery. Computers & Security, 96(??):Article 101937, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302133.

Lv:2025:TFT

[LZ25] Liuying Lv and Peng Zhou. TrojanProbe: Fingerprinting Trojan tunnel implementations by actively probing crafted HTTP requests. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004528.

Lopes:2023:NID

[LZA^{+}23] Ivandro O. Lopes, Deqing Zou, Ihsan H. Abdulqadder, Saeed Akbar, Zhen Li, Francis Ruambo, and Wagner Pereira. Network intrusion detection based on the temporal convolutional model. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003759.

Ling:2024:EFL

[LZC24] Jie Ling, Junchang Zheng, and Jiahui Chen. Efficient federated learning privacy preservation method with heterogeneous differential privacy. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000166.

Li:2022:CRM

[LZCS22] Bin Li, Qinglei Zhou, Yan Cao, and Xueming Si. Cognitively reconfigurable mimic-based heterogeneous password recovery system. Computers & Security, 116(??):Article 102667, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000669.

Luo:2020:SDR

[LZD^{+}20] Peng Luo, Deqing Zou, Yajuan Du, Hai Jin, Changming Liu, and Jinan Shen. Static detection of real-world buffer overflow induced by loop. Computers & Security, 89(??):Article 101616, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819301634.

Liu:2021:NDD

[LZDZ21] Xiaohu Liu, Hengwei Zhang, Shuqin Dong, and Yuchen Zhang. Network defense decision-making based on a stochastic game system and a deep recurrent Q-network. Computers & Security, 111(??):Article 102480, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003047.

Liu:2022:PPC

[LZG^{+}22] Haiyang Liu, Hanlin Zhang, Li Guo, Jia Yu, and Jie Lin. Privacy-preserving cloud-aided broad learning system. Computers & Security, 112(??):Article 102503, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003278.

Li:2024:PPL

[LZH^{+}24] Xuefeng Li, Chensu Zhao, Yi Hu, Honglin Xie, Yuhang Wang, and Jingyang Zhao. Precursor of privacy leakage detection for individual user. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001809.

Lu:2020:CAF

[LZHL20] Xiaofeng Lu, Shengfei Zhang, Pan Hui, and Pietro Lio. Continuous authentication by free-text keystroke based on CNN and RNN. Computers & Security, 96(??):Article 101861, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301334.

Li:2022:EYC

[LZL^{+}22] Helin Li, Hui Zhu, Xiaodong Lin, Rongxing Lu, Zhipeng Yu, and Wei Lan. Everything you control is not everything: Achieving intention-concealed visit on social networks. Computers & Security, 119(??):Article 102778, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001730.

Li:2024:EAR

[LZNZ24] Chaofei Li, Ziyuan Zhu, Ruicheng Niu, and Yuting Zhao. Enhancing adversarial robustness for deep metric learning via neural discrete adversarial training. Computers & Security, 143(??):??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824002013.

Li:2023:IIF

[LZRR23] Jie Li, Tianqing Zhu, Wei Ren, and Kim-Kwang Raymond. Improve individual fairness in federated learning via adversarial training. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002468.

Lu:2020:SSD

[LZS^{+}20] Ning Lu, Yongxin Zhang, Wenbo Shi, Saru Kumari, and Kim-Kwang Raymond Choo. A secure and scalable data integrity auditing scheme based on hyperledger fabric. Computers & Security, 92(??):Article 101741, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300274.

Li:2023:HDR

[LZS^{+}23] Qianyu Li, Min Zhang, Yi Shen, Ruipeng Wang, Miao Hu, Yang Li, and Hao Hao. A hierarchical deep reinforcement learning model with expert prior knowledge for intelligent penetration testing. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002687.

Li:2024:SSH

[LZW^{+}24] Xixing Li, Lei Zhao, Qiang Wei, Zehui Wu, Weiming Shi, and Yunchao Wang. SHFuzz: Service handler-aware fuzzing for detecting multi-type vulnerabilities in embedded devices. Computers & Security, 138(??):??, March 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300528X.

Liu:2023:DAU

[LZX^{+}23] Linfeng Liu, Yaoze Zhou, Zhiyuan Xi, Jiagao Wu, and Jia Xu. Defense against underwater spy-robots: a distributed anti-theft topology control mechanism for insecure UASN. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001244.

Li:2025:MAM

[LZY^{+}25a] Hongmei Li, Tiantian Zhu, Jie Ying, Tieming Chen, Mingqi Lv, Jian-Ping Mei, Zhengqiu Weng, and Lili Shi. MIRDETEC-TOR: Applying malicious intent representation for enhanced APT anomaly detection. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002779.

Li:2025:HPI

[LZY^{+}25b] Mingda Li, Tiantian Zhu, Haoqi Yan, Tieming Chen, and Mingqi Lv. HER-PT: an intelligent penetration testing framework with Hindsight Experience Replay. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740482500046X.

Liu:2021:QCP

[LZZ^{+}21] Xiaoxue Liu, Jiexin Zhang, Peidong Zhu, Qingping Tan, and Wei Yin. Quantitative cyber-physical security analysis methodology for industrial control systems based on incomplete information Bayesian game. Computers & Security, 102(??):Article 102138, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304119.

Liu:2021:PNL

[LZZX21] Yaqun Liu, Jinlong Zhao, Guomin Zhang, and Changyou Xing. NetObfu: a lightweight and efficient network topology obfuscation defense scheme. Computers & Security, 110(??):Article 102447, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002716.

Mohammad:2020:IRI

[MAAA20] Rami Mustafa A. Mohammad, Mutasem K. Alsmadi, Ibrahim Almarashdeh, and Malek Alzaqebah. An improved rule induction based denial of service attacks classification model. Computers & Security, 99(??):Article 102008, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302819.

Mishra:2022:AIC

[MAAG22] Alok Mishra, Yehia Ibrahim Alzoubi, Memoona Javeria Anwar, and Asif Qumer Gill. Attributes impacting cybersecurity policy development: an evidence from seven nations. Computers & Security, 120(??):Article 102820, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002140.

Makhdoom:2023:DCI

[MAF^{+}23] Imran Makhdoom, Mehran Abolhasan, Daniel Franklin, Justin Lipman, Christian Zimmermann, Massimo Piccardi, and Negin Shariati. Detecting compromised IoT devices: Existing techniques, challenges, and a way forward. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002948.

Mhaskar:2021:FAN

[MAK21] Neerja Mhaskar, Mohammed Alabbad, and Ridha Khedri. A formal approach to network segmentation. Computers & Security, 103(??):Article 102162, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304351.

Mamdouh:2021:AIM

[MAKH21] Moustafa Mamdouh, Ali Ismail Awad, Ashraf A. M. Khalaf, and Hesham F. A. Hamed. Authentication and identity management of IoHT devices: Achievements, challenges, and future directions. Computers & Security, 111(??):Article 102491, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003151.

Makhdoom:2022:CSC

[MAL22] Imran Makhdoom, Mehran Abolhasan, and Justin Lipman. A comprehensive survey of covert communication techniques, limitations and future challenges. Computers & Security, 120(??):Article 102784, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001791.

Mitseva:2023:SPI

[MAP23] Asya Mitseva, Marharyta Aleksandrova, and Andriy Panchenko. Security and performance implications of BGP reroutingresistant guard selection algorithms for Tor. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002845.

Mages:2022:TIT

[MAR22] Tobias Mages, Magnus Almgren, and Christian Rohner. Towards an information-theoretic framework of intrusion detection for composed systems and robustness analyses. Computers & Security, 116(??):Article 102633, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000323.

Mohammed:2023:DMF

[MAR^{+}23] Abubakar Sadiq Mohammed, Eirini Anthi, Omer Rana, Neetesh Saxena, and Pete Burnap. Detection and mitigation of field flooding attacks on oil and gas critical infrastructure communication. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003996.

Mubarkoot:2023:SCR

[MARB^{+}23] Mohammed Mubarkoot, Jörn Altmann, Morteza Rasti-Barzoki, Bernhard Egger, and Hyejin Lee. Software compliance requirements, factors, and policies: a systematic literature review. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003777.

Mavroeidis:2020:NLC

[MB20a] Vasileios Mavroeidis and Joe Brule. A nonproprietary language for the command and control of cyber defenses — OpenC2. Computers & Security, 97(??):Article 101999, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302728.

Menard:2020:AIU

[MB20b] Philip Menard and Gregory J. Bott. Analyzing IOT users’ mobile device privacy concerns: Extracting privacy permissions using a disclosure experiment. Computers & Security, 95(??):Article 101856, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301280.

Meidan:2023:DSE

[MBB^{+}23] Yair Meidan, Daniel Benatar, Ron Bitton, Dan Avraham, and Asaf Shabtai. D-Score: an expert-based method for assessing the detectability of IoT-related cyber-attacks. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004655.

Mersinas:2025:CBC

[MBF25] Konstantinos Mersinas, Maria Bada, and Steven Furnell. Cybersecurity behavior change: a conceptualization of ethical principles for behavioral interventions. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824003304.

Martin:2021:ADU

[MBFIdD21] Alejandro G. Martín, Marta Beltrán, Alberto Fernández-Isabel, and Isaac Martín de Diego. An approach to detect user behaviour anomalies within identity federations. Computers & Security, 108(??):Article 102356, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001802.

Menet:2020:SNS

[MBGF20] François Menet, Paul Berthier, Michel Gagnon, and José M. Fernandez. Spartan Networks: Self-feature-squeezing neural networks for increased robustness in adversarial settings. Computers & Security, 88(??):Article 101537, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819300069.

Moia:2020:IEC

[MBH20] Vitor Hugo Galhardo Moia, Frank Breitinger, and Marco Aurélio Amaral Henriques. The impact of excluding common blocks for approximate matching. Computers & Security, 89(??):Article 101676, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819302159.

Mimran:2022:SOR

[MBK^{+}22] Dudu Mimran, Ron Bitton, Yehonatan Kfir, Eitan Klevansky, Oleg Brodt, Heiko Lehmann, Yuval Elovici, and Asaf Shabtai. Security of open radio access networks. Computers & Security, 122(??):Article 102890, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200284X.

Medvedev:2025:DMF

[MBK25] Viktor Medvedev, Arnoldas Budzys, and Olga Kurasova. A decision-making framework for user authentication using keystroke dynamics. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001828.

Mirtsch:2021:ISM

[MBKD21] Mona Mirtsch, Knut Blind, Claudia Koch, and Gabriele Dudek. Information security management in ICT and non-ICT sector companies: a preventive innovation perspective. Computers & Security, 109(??):Article 102383, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002078.

Malhi:2020:SVA

[MBP20] Avleen Kaur Malhi, Shalini Batra, and Husanbir Singh Pannu. Security of vehicular ad-hoc networks: a comprehensive survey. Computers & Security, 89(??):Article 101664, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818312872.

McGahagan:2021:DFD

[MBPCC21] John McGahagan, Darshan Bhansali, Ciro Pinto-Coelho, and Michel Cukier. Discovering features for detecting malicious websites: an empirical study. Computers & Security, 109(??):Article 102374, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482100198X.

Munsinger:2023:VRI

[MBR23] Brita Munsinger, Nicole Beebe, and Turquoise Richardson. Virtual reality for improving cyber situational awareness in security operations centers. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300278X.

Meland:2020:RSE

[MBS20] Per Håkon Meland, Yara Fareed Fahmy Bayoumy, and Guttorm Sindre. The ransomware-as-a-service economy within the darknet. Computers & Security, 92(??):Article 101762, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300468.

Mathov:2022:SBM

[MBSE22] Yael Mathov, Tal Ben Senior, Asaf Shabtai, and Yuval Elovici. Stop bugging me! evading modern-day wiretapping using adversarial perturbations. Computers & Security, 121(??):Article 102841, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002358.

Mahnamfar:2024:RPW

[MBU24] Amin Mahnamfar, Kemal Bicakci, and Yusuf Uzunay. ROS-TAM: a passwordless web single sign-on solution mitigating server breaches and integrating credential manager and federated identity systems. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000403.

Matthew:2020:AAO

[MC20] Peter Matthew and Susan Canning. An algorithmic approach for optimising biometric systems using liveness and coercion detection. Computers & Security, 94(??):Article 101831, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030105X.

Milosevic:2022:EMC

[MC22] Marija S. Milosevic and Vladimir M. Ciric. Extreme minority class detection in imbalanced data for network intrusion. Computers & Security, 123(??):Article 102940, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003327.

Meurs:2024:DDE

[MCC^{+}24] Tom Meurs, Edward Cartwright, Anna Cartwright, Marianne Junger, and Abhishta Abhishta. Deception in double extortion ransomware attacks: an analysis of profitability and credibility. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005801.

Martinez:2022:WTC

[MCJPS22] David Martínez, Eusebi Calle, Albert Jové, and Cristina Pérez-Solà. Web-tracking compliance: websites’ level of confidence in the use of information-gathering technologies. Computers & Security, 122(??):Article 102873, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200267X.

Mao:2024:ERP

[MCM^{+}24] Zimin Mao, Shuiyan Chen, Zhuang Miao, Heng Li, Beihao Xia, Junzhe Cai, Wei Yuan, and Xinge You. Enhancing robustness of person detection: a universal defense filter against adversarial patch attacks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003717.

Molina-Coronado:2023:TFC

[MCMMMA23] Borja Molina-Coronado, Usue Mori, Alexander Mendiburu, and Jose Miguel-Alonso. Towards a fair comparison and realistic evaluation framework of Android malware detectors based on static analysis and machine learning. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003881.

Ma:2022:NPA

[MCS22] Qian Ma, Baojiang Cui, and Cong Sun. A novel privacyaware model for nonparametric decentralized detection. Computers & Security, 117(??):Article 102688, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000864.

Morkonda:2024:IDP

[MCvO24] Srivathsan G. Morkonda, Sonia Chiasson, and Paul C. van Oorschot. Influences of displaying permission-related information on web single sign-on login decisions. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482300576X.

McLeod:2022:ISP

[MD22] Alexander McLeod and Diane Dolezel. Information security policy non-compliance: Can capitulation theory explain user behaviors? Computers & Security, 112(??):Article 102526, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003503.

Maiorca:2020:ADF

[MDB^{+}20] Davide Maiorca, Ambra Demontis, Battista Biggio, Fabio Roli, and Giorgio Giacinto. Adversarial detection of flash malware: Limitations and open issues. Computers & Security, 96(??):Article 101901, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301760.

Monkam:2024:TDA

[MDB24] Galamo F. Monkam, Michael J. De Lucia, and Nathaniel D. Bastian. A topological data analysis approach for detecting data poisoning attacks against machine learning based network intrusion detection systems. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002323.

Mirsky:2023:TOA

[MDK^{+}23] Yisroel Mirsky, Ambra Demontis, Jaidip Kotak, Ram Shankar, Deng Gelei, Liu Yang, Xiangyu Zhang, Maura Pintor, Wenke Lee, Yuval Elovici, and Battista Biggio. The threat of offensive AI to organizations. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003984.

Mohammed:2025:HRA

[MDKR25] Ali Hameed Yassir Mohammed, Rudzidatul Akmam Dziyauddin, Norshaliza Kamaruddin, and Fiza Abdul Rahim. A hybrid ranking algorithm for secure and efficient iris template protection. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005224.

Meyer:2020:AAP

[MDR20] Joachim Meyer, Omer Dembinsky, and Tal Raviv. Alerting about possible risks vs. blocking risky choices: a quantitative model and its empirical evaluation. Computers & Security, 97(??):Article 101944, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302200.

Mothanna:2024:ASP

[MEH^{+}24] Yusuf Mothanna, Wael ElMedany, Mustafa Hammad, Riadh Ksantini, and Mhd Saeed Sharif. Adopting security practices in software development process: Security testing framework for sustainable smart cities. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002906.

Mohammed:2025:IDT

[MF25] Ahmad Burhan Mohammed and Lamia Chaari Fourati. Investigation on datasets toward intelligent intrusion detection systems for intra and inter-UAVs communication systems. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005212.

Mahdavi:2020:RTA

[MFA20] Ehsan Mahdavi, Ali Fanian, and Fatima Amini. A real-time alert correlation method based on code-books for intrusion detection systems. Computers & Security, 89(??):Article 101661, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819302032.

Mechri:2025:SLL

[MFD25] Abdechakour Mechri, Mohamed Amine Ferrag, and Merouane Debbah. SecureQwen: Leveraging LLMs for vulnerability detection in Python codebases. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004565.

Maoneke:2020:ESM

[MFI20] Pardon Blessings Maoneke, Stephen Flowerday, and Naomi Isabirye. Evaluating the strength of a multilingual passphrase policy. Computers & Security, 92(??):Article 101746, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300304.

Ma:2024:FLB

[MG24a] Zihan Ma and Tianchong Gao. Federated learning backdoor attack detection with persistence diagram. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004674.

Majgave:2024:APW

[MG24b] Amol Babaso Majgave and Nitin L. Gavankar. Automatic phishing website detection and prevention model using transformer deep belief network. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003766.

Mulahuwaish:2020:ECM

[MGG^{+}20] Aos Mulahuwaish, Kevin Gyorick, Kayhan Zrar Ghafoor, Halgurd S. Maghdid, and Danda B. Rawat. Efficient classification model of web news documents using machine learning algorithms for accurate information. Computers & Security, 98(??):Article 102006, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302790.

Mejias:2025:MIS

[MGG^{+}25] Roberto J. Mejias, Joshua J. Greer, Gabrila C. Greer, Morgan M. Shepherd, and Raul Y. Reyes. A model for information security vulnerability awareness. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824006114.

Maesschalck:2022:DGS

[MGGR22] Sam Maesschalck, Vasileios Giotsas, Benjamin Green, and Nicholas Race. Don’t get stung, cover your ICS in honey: How do honeypots fit within industrial control system security. Computers & Security, 114(??):Article 102598, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821004211.

Meng:2020:SDH

[MGL^{+}20] Shunmei Meng, Zijian Gao, Qianmu Li, Hao Wang, Hong-Ning Dai, and Lianyong Qi. Security-driven hybrid collaborative recommendation method for cloud-based IoT services. Computers & Security, 97(??):Article 101950, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302261.

Masuch:2022:AJE

[MGTK22] Kristin Masuch, Maike Greve, Simon Trang, and Lutz M. Kolbe. Apologize or justify? Examining the impact of data breach response actions on stock value of affected companies? Computers & Security, 112(??):Article 102502, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003266.

Mi:2020:EPA

[MGZ^{+}20] Tianyue Mi, Mengke Gou, Guangyu Zhou, Yiqun Gan, and Ralf Schwarzer. Effects of planning and action control on smartphone security behavior. Computers & Security, 97(??):Article 101954, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302303.

Meng:2025:DDM

[MGZ^{+}25] Fanchao Meng, Jiaping Gui, Futai Zou, Yunbo Li, and Yue Wu. DISTR: Detecting multi-stage IoT botnets through contextual traffic and causal analytics. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002202.

Miah:2025:CPR

[MH25] Arafat Miah and Fakir Sharif Hossain. CRO-PUF: Resilience to machine learning and differential power attacks. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000021.

Miao:2025:WWE

[MHC^{+}25] Yang Miao, Xiaoyan Hu, Guang Cheng, Ruidong Li, Hua Wu, and Yang Meng. WEDoHTool: Word embedding based early identification of DoH tunnel tool traffic in dynamic network environments. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003694.

Mayer:2024:DBL

[MHE24] Rudolf Mayer, Markus Hittmeir, and Andreas Ekelhart. Distance-based linkage of personal microbiome records for identification and its privacy implications. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823004480.

Muzaffar:2022:DRM

[MHLZ22] Ali Muzaffar, Hani Ragab Hassen, Michael A. Lones, and Hind Zantout. An in-depth review of machine learning based Android malware detection. Computers & Security, 121(??):Article 102833, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002279.

Meng:2021:SPM

[MHSK21] Yunfei Meng, Zhiqiu Huang, Guohua Shen, and Changbo Ke. A security policy model transformation and verification approach for software defined networking. Computers & Security, 100(??):Article 102089, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030362X.

Mouratidis:2023:MLC

[MISO^{+}23] Haralambos Mouratidis, Shareeful Islam, Antonio Santos-Olmo, Luis E. Sanchez, and Umar Mukhtar Ismail. Modelling language for cyber security incident handling for critical infrastructures. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000494.

Mitchell:2020:IQC

[Mit20]  Chris J. Mitchell.  The impact of quantum comput-

ing on real-world security: a 5G case study. Computers & Security, 93(??):Article 101825, June 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300997.

Madhusudhanan:2025:PPT

[MJ25] Sheema Madhusudhanan and Arun Cyril Jose. Privacy preservation techniques through data lifecycle: a comprehensive literature survey. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001622.

Ma:2025:TMG

[MJZ^{+}25] Chunyan Ma, Zhengwei Jiang, Kai Zhang, Zhiting Ling, Jun Jiang, Yizhe You, Peian Yang, and Huamin Feng. TIM-Fuser: a multi-granular fusion framework for cyber threat intelligence. Computers & Security, 148(??):??, January 2025.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004462.

Malini:2024:EDL

[MK24] P. Malini and K. R. Kavitha. An efficient deep learning mechanisms for IoT/non-IoT devices classification and attack detection in SDN-enabled smart environment. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001196.

McIntosh:2021:DUC

[MKC^{+}21] Timothy McIntosh, A. S. M. Kayes, Yi-Ping Phoebe Chen, Alex Ng, and Paul Watters. Dynamic user-centric access control for detection of ransomware attacks. Computers & Security, 111(??):Article 102461, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002856.

McIntosh:2023:ASE

[MKC^{+}23] Timothy McIntosh, A. S. M. Kayes, Yi-Ping Phoebe Chen, Alex Ng, and Paul Watters. Applying staged event-driven access control to combat ransomware. Computers & Security, 128 (??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000706.

Meng:2024:MTB

[MKH24] Yunfei Meng, Changbo Ke, and Zhiqiu Huang. A model transformation based security policy automatic management framework for software-defined networking. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001512.

Meng:2021:UBC

[MKL21] Zhangyu Meng, Jun Kong, and Juan Li. Utilizing binary code to improve usability of pressure-based authentication. Computers & Security, 103(??):Article 102187, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000110.

Masud:2025:VDU

[MKM^{+}25] Mohammed Tanvir Masud, Marwa Keshk, Nour Moustafa, Benjamin Turnbull, and Willy Susilo. Vulnerability defence using hybrid moving target defence in Internet of Things systems. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000690.

Mustapha:2023:DDA

[MKZ^{+}23] Ali Mustapha, Rida Khatoun, Sherali Zeadally, Fadlallah Chbib, Ahmad Fadlallah, Walid Fahs, and Ali El Attar. Detecting DDoS attacks using adversarial neural network. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000275.

Mai:2025:RAR

[MLB^{+}25] Khang Mai, Jongmin Lee, Razvan Beuran, Ryosuke Hotchi, Sian En Ooi, Takayuki Kuroda, and Yasuo Tan. RAF-AG: Report analysis framework for attack path generation. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004309.

Mao:2023:BAA

[MLC^{+}23] Zhongshu Mao, Yiqin Lu, Zhe Cheng, Xiong Shen, Yang Zhang, and Jiancheng Qin. Boosting adversarial attacks with future momentum and future transformation. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000342.

Ma:2020:TMM

[MLM20] Zuchao Ma, Liang Liu, and Weizhi Meng. Towards multiplemix-attack detection via consensus-based trust management in IoT networks. Computers & Security, 96(??):Article 101898, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482030170X.

Moura:2023:MUM

[MLM^{+}23] Ricardo Moura, Ricardo Lopes, David R. Matos, Miguel L. Pardal, and Miguel Correia. MultiTLS: using multiple and diverse ciphers for stronger secure channels. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002523.

Meng:2023:IRM

[MLQ^{+}23] Shunmei Meng, Qianmu Li, Lianyong Qi, Xiaolong Xu, Rui Yuan, and Xuyun Zhang. An intelligent recommendation method based on multi-interest network and adversarial deep learning. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001748.

McIntosh:2023:HGG

[MLS^{+}23] Timothy McIntosh, Tong Liu, Teo Susnjak, Hooman Alavizadeh, Alex Ng, Raza Nowrozy, and Paul Watters. Harnessing GPT-4 for generation of cybersecurity GRC policies: a focus on ransomware attack mitigation. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003346.

Menges:2021:TGC

[MLV^{+}21] Florian Menges, Tobias Latzo, Manfred Vielberth, Sabine Sobola, Henrich C. Pöhls, Benjamin Taubmann, Johannes Köstler, Alexander Puchta, Felix Freiling, Hans P. Reiser, and Günther Pernul. Towards GDPR-compliant data processing in modern SIEM systems. Computers & Security, 103 (??):Article 102165, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820304387.

Ma:2024:BSC

[MLXC24] Yuxiang Ma, Zhaodi Li, Haoming Xue, and Jike Chang. A balanced supervised contrastive learning-based method for encrypted network traffic classification. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003286.

Ma:2023:MPT

[MLZ^{+}23] Hua Ma, Qun Li, Yifeng Zheng, Zhi Zhang, Xiaoning Liu, Yansong Gao, Said F. Al-Sarawi, and Derek Abbott. MUD-PQFed: Towards Malicious User Detection on model corruption in Privacy-preserving Quantized Federated learning. Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003164.

Maeda:2021:APE

[MM21] Ryusei Maeda and Mamoru Mimura. Automating postexploitation with deep reinforcement learning. Computers & Security, 100(??):Article 102108, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303813.

Minna:2023:SRT

[MM23] Francesco Minna and Fabio Massacci. SoK: Run-time security for cloud microservices. Are we there yet? Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000299.

Mishra:2025:CRI

[MM25] Raghavendra Mishra and Ankita Mishra. Current research on Internet of Things (IoT) security protocols: a survey. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006163.

Maniriho:2024:MMA

[MMC24] Pascal Maniriho, Abdun Naser Mahmood, and Mohammad Jabed Morshed Chowdhury. MeMalDet: a memory analysis-based malware detection framework using deep autoencoders and stacked ensemble under temporal evaluations. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001652.

Mitchell:2024:GSE

[MMdR24] Jeff Mitchell, Niall McLaughlin, and Jesus Martinez del Rincon. Generating sparse explanations for malicious Android opcode sequences using hierarchical LIME. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005473.

Moreira:2023:IRD

[MMdSdS23] Caio C. Moreira, Davi C. Moreira, and Claudomiro de S. de Sales, Jr. Improving ransomware detection based on portable executable header using Xception Convolutional Neural Network. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300175X.

Motylinski:2022:GBM

[MMIS22] Michal Motylinski, Áine MacDermott, Farkhund Iqbal, and Babar Shah. A GPU-based machine learning approach for detection of botnet attacks. Computers & Security, 123(??):Article 102918, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003108.

Ma:2024:IPP

[MMJG24] Xirong Ma, Chuan Ma, Yali Jiang, and Chunpeng Ge. Improved privacy-preserving PCA using optimized homomorphic matrix multiplication. Computers & Security, 138(??): ??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005679.

Malviya:2025:FDA

[MMSJ25] Vikas K. Malviya, Wei Minn, Lwin Khin Shar, and Lingxiao Jiang. Fuzzing drones for anomaly detection: a systematic literature review. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004620.

Malatji:2020:VST

[MMvS20] Masike Malatji, Annlizé Marnewick, and Suné von Solms. Validation of a socio-technical management process for optimising cybersecurity practices. Computers & Security, 95 (??):Article 101846, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482030119X.

Mageswari:2024:ASC

[MNMS24] R. Uma Mageswari, Zafar Ali Khan N., Gowthul Alam M. M., and Jerald Nirmal Kumar S. Addressing security challenges in Industry 4.0: AVA-MA approach for strengthening SDN-IoT network security. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002098.

Mangussi:2025:SRD

[MPL^{+}25] Arthur Dantas Mangussi, Ricardo Cardoso Pereira, Ana Carolina Lorena, Miriam Seoane Santos, and Pedro Henriques Abreu. Studying the robustness of data imputation methodologies against adversarial attacks. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002639.

Mahadik:2023:EHD

[MPM23] Shalaka S. Mahadik, Pranav M. Pawar, and Raja Muthalagu. Edge-HetIoT defense against DDoS attack using learning techniques. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002572.

Martin-Perez:2021:PPM

[MPRB21] Miguel Martín-Pérez, Ricardo J. Rodríguez, and Davide Balzarotti. Pre-processing memory dumps to improve similarity score of Windows modules. Computers & Security, 101(??):Article 102119, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303928.

Mao:2023:OFB

[MQH^{+}23] Jiaoze Mao, Yaguan Qian, Jianchang Huang, Zejie Lian, Renhui Tao, Bin Wang, Wei Wang, and Tengteng Yao. Object-free backdoor attack and defense on semantic segmentation. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002754.

Mollaeefar:2023:IQT

[MR23] Majid Mollaeefar and Silvio Ranise. Identifying and quantifying trade-offs in multi-stakeholder risk evaluation with applications to the data protection impact assessment of the GDPR. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001165.

Maldonado:2025:EWS

[MR25] Javier Maldonado and María Cristina Riff. An evolutionary wrapper to support intrusion detection system configuration. Computers & Security, 155(??):??, August 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482500166X.

Marques:2021:WUD

[MRG21] Pedro Marques, Matilda Rhode, and Ilir Gashi. Waste not: Using diverse neural networks from hyperparameter search for improved malware detection. Computers & Security, 108(??):Article 102339, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001632.

Mohanty:2022:RSE

[MRL22] Hardhik Mohanty, Arousha Haghighian Roudsari, and Arash Habibi Lashkari. Robust stacking ensemble model for darknet traffic classification under adversarial settings. Computers & Security, 120(??):Article 102830, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002243.

Munoz:2023:SST

[MRRL23] Antonio Muñoz, Ruben Ríos, Rodrigo Román, and Javier López. A survey on the (in)security of trusted execution environments. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000901.

M:2024:NNA

[MRSP24] Jesús F. Cevallos M., Alessandra Rizzardi, Sabrina Sicari, and Alberto Coen Porisini. NERO: NEural algorithmic reasoning for zeRO-day attack detection in the IoT: a hybrid approach. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002001.

Merlo:2021:YSR

[MRSV21] Alessio Merlo, Antonio Ruggia, Luigi Sciolla, and Luca Verderame. You shall not repackage! demystifying anti-repackaging on Android. Computers & Security, 103(??):Article 102181, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000055.

Manda:2023:CET

[MS23] Sridhar Manda and Charanjeet Singh. CVFP: Energy and trust aware data routing protocol based on Competitive Verse Flower Pollination algorithm in IoT. Computers & Security, 127(??): ??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004278.

Mouiche:2025:ERE

[MS25] Inoussa Mouiche and Sherif Saad. Entity and relation extractions for threat intelligence knowledge graphs. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004255.

Marshall:2024:EEE

[MSA24] Nina Marshall, Daniel Sturman, and Jaime C. Auton. Exploring the evidence for email phishing training: a scoping review. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823006053.

Ma:2021:NMA

[MSCJ21] Qian Ma, Cong Sun, Baojiang Cui, and Xiaohui Jin. A novel model for anomaly detection in network traffic based on kernel support vector machine. Computers & Security, 104(??):Article 102215, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000390.

Maesschalck:2023:WUL

[MSD^{+}23] Sam Maesschalck, Alexander Staves, Richard Derbyshire, Benjamin Green, and David Hutchison. Walking under the ladder logic: PLC-VBS: a PLC control logic vulnerability scanning tool. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000263.

Manimaran:2022:SNF

[MSG22] S. Manimaran, V. N. Sastry, and N. P. Gopalan. SBT-DDL: a novel framework for sensor-based threats detection on Android smartphones using deep learning. Computers & Security, 118(??):Article 102729, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001249.

Milani:2025:FCE

[MSH^{+}25] Alessandra Maciel Paz Milani, Arty Starr, Samantha Hill, Callum Curtis, Norman Anderson, David Moreno-Lumbreras, and Margaret-Anne Storey. Fuzzy to clear: Elucidating the threat hunter cognitive process and cognitive support needs. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003402.

McIntosh:2024:CIE

[MSL^{+}24] Timothy R. McIntosh, Teo Susnjak, Tong Liu, Paul Watters, Dan Xu, Dongwei Liu, Raza Nowrozy, and Malka N. Halgamuge. From COBIT to ISO 42001: Evaluating cybersecurity frameworks for opportunities, risks, and regulatory compliance in commercializing large language models. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002694.

Miao:2025:ASA

[MSL^{+}25] Zujia Miao, Cuiping Shao, Huiyun Li, Yunduan Cui, and Zhimin Tang. Adaptive sensor attack detection and defense framework for autonomous vehicles based on density. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004541.

Modesti:2021:FMS

[MSMH21] Paolo Modesti, Siamak F. Shahandashti, Patrick McCorry, and Feng Hao. Formal modelling and security analysis of Bitcoin’s payment protocol. Computers & Security, 107(??):Article 102279, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001036.

Meidan:2020:NAD

[MSP^{+}20] Yair Meidan, Vinay Sachidananda, Hongyi Peng, Racheli Sagron, Yuval Elovici, and Asaf Shabtai. A novel approach for detecting vulnerable IoT devices connected behind a home NAT. Computers & Security, 97(??):Article 101968, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302418.

Martinez:2025:IMP

[MSSO^{+}25] Ferney Martínez, Luis E. Sánchez, Antonio Santos-Olmo, David G. Rosado, and Eduardo Fernández-Medina. Integrated maritime protection: Innovation for the safeguarding of maritime systems based on MARISMA. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003888.

Miao:2020:ASE

[MSZ20] Kelei Miao, Xiufang Shi, and Wen-An Zhang. Attack signal estimation for intrusion detection in industrial control system. Computers & Security, 96(??):Article 101926, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302029.

Madani:2021:ORI

[MT21a] Mahdi Madani and Camel Tanougast. Optimized and robust implementation of mobile networks confidentiality and integrity functions. Computers & Security, 100(??):Article 102093, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303667.

Mohammed:2021:EPP

[MT21b] Zareef A. Mohammed and Gurvirender P. Tejay. Examining the privacy paradox through individuals’ neural disposition in e-commerce: an exploratory neuroimaging study. Computers & Security, 104(??):Article 102201, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000250.

Mohanty:2021:HNH

[MT21c] Susil Kumar Mohanty and Somanath Tripathy. n-HTLC: Neo hashed time-lock commitment to defend against wormhole attack in payment channel networks. Computers & Security, 106(??):Article 102291, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001152.

Ma:2024:NPA

[MT24] Yi-Wei Ma and Chia-Wei Tsou. A novel passive-active detection system for false data injection attacks in industrial control systems. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003018.

Mvah:2024:CAS

[MTD^{+}24] Fabrice Mvah, Vianney Kengne Tchendji, Clémentin Tayou Djamegni, Ahmed H. Anwar, Deepak K. Tosh, and Charles Kamhoua. Countering ARP spoofing attacks in softwaredefined networks using a game-theoretic approach. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823006065.

Mott:2023:BRH

[MTN^{+}23] Gareth Mott, Sarah Turner, Jason R. C. Nurse, Jamie MacColl, James Sullivan, Anna Cartwright, and Edward Cartwright. Between a rock and a hard(ening) place: Cyber insurance in the ransomware era. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300072X.

Miculan:2023:AVT

[MV23] Marino Miculan and Nicola Vitacolonna. Automated verification of Telegram’s MTProto 2.0 in the symbolic model. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004643.

Mathina:2025:AIS

[MV25] P. A. Mathina and K. Valarmathi. Advancing IoT security: a novel intrusion detection system for evolving threats in industry 4.0 using optimized convolutional sparse ficks law graph point trans-net. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824004747.

Madhushanie:2025:RTP

[MVA25] Nadisha Madhushanie, Sugandima Vidanagamachchi, and Nalin Arachchilage. Real-time privacy vulnerability detection techniques in software development: a systematic literature review. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003487.

Mileva:2021:CAM

[MVH^{+}21] Aleksandra Mileva, Aleksandar Velinov, Laura Hartmann, Steffen Wendzel, and Wojciech Mazurczyk. Comprehensive analysis of MQTT 5.0 susceptibility to network covert channels. Computers & Security, 104(??):Article 102207, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000316.

Mikuletic:2024:SPO

[MVSSZ24] Samanta Mikuletic, Simon Vrhovec, Brigita Skela-Savic, and Bostjan Zvanut. Security and privacy oriented information security culture (ISC): Explaining unauthorized access to healthcare data by nursing employees. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003991.

Marquez:2024:VIA

[MVVL^{+}24] A. Germán Márquez, Ángel Jesús Varela-Vaca, María Teresa Gómez López, José A. Galindo, and David Benavides. Vulnerability impact analysis in software project dependencies based on Satisfiability Modulo Theories (SMT).

Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005795.

Mansoori:2020:HDT

[MW20] Masood Mansoori and Ian Welch. How do they find us? A study of geolocation tracking techniques of malicious web sites. Computers & Security, 97(??):Article 101948, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302248.

Mills:2023:LRB

[MWL23] Alan Mills, Jonathan White, and Phil Legg. Longitudinal risk-based security assessment of Docker software container images. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003887.

Malach:2025:CEP

[MWM^{+}25] Alon Malach, Prasanna N. Wudali, Satoru Momiyama, Jun Furukawa, Toshinori Araki, Yuval Elovici, and Asaf Shabtai. CyberShapley: Explanation, prioritization, and triage of cybersecurity alerts using informative graph representation. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005765.

Mikhail:2020:PGE

[MWR20] Joseph W. Mikhail, Jamie C. Williams, and George R. Roelke. procmonML: Generating evasion resilient hostbased behavioral analytics from tree ensembles. Computers & Security, 98(??):Article 102002, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302753.

Mirzaei:2025:ENT

[MYH25] Reza Mirzaei, Nasser Yazdani, and Mohammad Sayad Haghighi. Enhancing network tunnels anonymity through increasing combined traffic in a clustered structure. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005303.

Mejri:2020:RSA

[MYMC20] Mohamed Mejri, Hamdi Yahyaoui, Azzam Mourad, and Mohamad Chehab. A rewriting system for the assessment of XACML policies relationship. Computers & Security, 97(??):Article 101957, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302339.

Madkaikar:2025:PAD

[MYS^{+}25] Gaurav Madkaikar, Karthikeya S. M. Yelisetty, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Performance analysis of dynamic ABAC systems using a queuing theoretic framework. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482500121X.

Ma:2024:MET

[MYW24] Yukun Ma, Chunlin Yu, and Chuliang Weng. Morpheus: an efficient timing-based attestation framework for safeguarding hypervisor integrity with dynamic trust. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002712.

Makhdoom:2020:PBB

[MZA^{+}20] Imran Makhdoom, Ian Zhou, Mehran Abolhasan, Justin Lipman, and Wei Ni. PrivySharing: a blockchain-based framework for privacy-preserving and secure data sharing in smart cities. Computers & Security, 88(??):Article 101653, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740481930197X.

Mu:2025:SPP

[MZLK25] Gaoli Mu, Hanlin Zhang, Jie Lin, and Fanyu Kong. SMCD: Privacy-preserving deep learning based malicious code detection. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005327.

Ma:2023:EEE

[MZW^{+}23] Xiuli Ma, Wenbin Zhu, Jieling Wei, Yanliang Jin, Dongsheng Gu, and Rui Wang. EETC: an extended encrypted traffic classification algorithm based on variant ResNet network. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000858.

Mbaka:2025:AUD

[MZW^{+}25] Winnie Bahati Mbaka, Xinran Zhang, Yunduo Wang, Tong Li, Fabio Massacci, and Katja Tuma. Assessing the usefulness of data flow diagrams for validating security threats. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001865.

Ma:2023:DDI

[MZWM23] Binhao Ma, Can Zhao, Dejun Wang, and Bo Meng. DIHBA: Dynamic, invisible and high attack success rate boundary backdoor attack with low poison ratio. Computers & Security, 129 (??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001220.

Nasir:2023:CDL

[NAK23] Muhammad Hassan Nasir, Junaid Arshad, and Muhammad Mubashir Khan. Collaborative device-level botnet detection for Internet of Things. Computers & Security, 129(??): ??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000822.

Ng:2025:CSG

[NB25a] Chiu Yeong Ng and Mohammad Khatim Bin Hasan. Cybersecurity serious games development: a systematic review. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006138.

Nguyen:2025:FSS

[NB25b] Van Tuan Nguyen and Razvan Beuran. FedMSE: Semisupervised federated learning approach for IoT network intrusion detection. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000264.

Nadler:2022:VAM

[NBBS22] Asaf Nadler, Ron Bitton, Oleg Brodt, and Asaf Shabtai. On the vulnerability of anti-malware solutions to DNS attacks. Computers & Security, 116(??):Article 102687, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000852.

Najafi:2024:HHA

[NCM24] Pejman Najafi, Feng Cheng, and Christoph Meinel. HEOD: Human-assisted Ensemble Outlier Detection for cybersecurity. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003456.

Nguyen:2024:XNV

[NCN24] Qui Ngoc Nguyen, Nguyen Tan Cam, and Kiet Van Nguyen. XLMR4MD: New Vietnamese dataset and framework for detecting the consistency of description and permission in Android applications using large language models. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001159.

Nevado-Catalan:2023:AFS

[NCPVRT23] David Nevado-Catalán, Sergio Pastrana, Narseo Vallina-Rodriguez, and Juan Tapiador. An analysis of fake social media engagement services. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004059.

Neigel:2020:HCH

[NCW^{+}20] Alexis R. Neigel, Victoria L. Claypoole, Grace E. Waldfogle, Subrata Acharya, and Gabriella M. Hancock. Holistic cyber hygiene education: Accounting for the human factors. Computers & Security, 92(??):Article 101731, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300183.

Nie:2023:KED

[NCW^{+}23] Xiaofan Nie, Liwei Chen, Haolai Wei, Yuantong Zhang, Ningning Cui, and Gang Shi. KPDFI: Efficient data flow integrity based on key property against data corruption attack. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000937.

Nwankpa:2023:RVR

[ND23] Joseph K. Nwankpa and Pratim Milton Datta. Remote vigilance: the roles of cyber awareness and cybersecurity policies among remote workers. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001761.

Nguyen-Duc:2021:ASA

[NDDH^{+}21] Anh Nguyen-Duc, Manh Viet Do, Quan Luong Hong, Kiem Nguyen Khac, and Anh Nguyen Quang. On the adoption of static analysis for software security assessment — a case study of an open-source e-government project. Computers & Security, 111(??):Article 102470, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002947.

Nowroozi:2021:SML

[NDPC21] Ehsan Nowroozi, Ali Dehghantanha, Reza M. Parizi, and Kim-Kwang Raymond Choo. A survey of machine learning techniques in adversarial image forensics. Computers & Security, 100(??):Article 102092, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303655.

Nawaz:2022:MMM

[NFVN^{+}22] M. Saqib Nawaz, Philippe Fournier-Viger, M. Zohaib Nawaz, Guoting Chen, and Youxi Wu. MalSPM: Metamorphic malware behavior analysis and classification using sequential pattern mining. Computers & Security, 118(??):Article 102741, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001365.

Novkovic:2023:IMK

[NG23] Bojan Novković and Marin Golub. Improving monolithic kernel security and robustness through intra-kernel sandboxing. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000147.

Nguyen:2025:GSM

[NGC^{+}25] Nhung H. Nguyen, Mengmeng Ge, Jin-Hee Cho, Terrence J. Moore, Seunghyun Yoon, Hyuk Lim, Frederica Nelson, Guangdong Bai, and Dan Dongseong Kim. Graphical security modelling for autonomous vehicles: a novel approach to threat analysis and defence evaluation. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005352.

Naik:2022:EPA

[NGJ^{+}22] Nitin Naik, Paul Grace, Paul Jenkins, Kshirasagar Naik, and Jingping Song. An evaluation of potential attack surfaces based on attack tree modelling and risk matrix applied to self-sovereign identity. Computers & Security, 120(??):Article 102808, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002024.

Nahum:2024:OCO

[NGM^{+}24] Mor Nahum, Edita Grolman, Inbar Maimon, Dudu Mimran, Oleg Brodt, Aviad Elyashar, Yuval Elovici, and Asaf Shabtai. OSSIntegrity: Collaborative open-source code integrity verification. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002827.

Niu:2023:NAD

[NGX^{+}23] Zequn Niu, Wenjie Guo, Jingfeng Xue, Yong Wang, Zixiao Kong, and Lu Huang. A novel anomaly detection approach based on ensemble semi-supervised active learning (ADESSA). Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001001.

Nashat:2021:MDF

[NH21] Dalia Nashat and Fatma A. Hussain. Multifractal detrended fluctuation analysis based detection for SYN flooding attack. Computers & Security, 107(??):Article 102315, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001395.

Nkoom:2025:SIR

[NHC25] Matilda Nkoom, Sena G. Hounsinou, and Garth V. Crosby. Securing the Internet of robotic things (IoRT) against DDoS attacks: a federated learning with differential privacy clustering approach. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001816.

Nguyen:2025:PRT

[NHFB25] Huynh Phuong Thanh Nguyen, Kento Hasegawa, Kazuhide Fukushima, and Razvan Beuran. PenGym: Realistic training environment for reinforcement learning pentesting agents.

Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004450.

Nguyen:2025:HFE

[NHT^{+}25] Pham Sy Nguyen, Tran Nhat Huy, Tong Anh Tuan, Pham Duy Trung, and Hoang Viet Long. Hybrid feature extraction and integrated deep learning for cloud-based malware detection. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400539X.

Niemimaa:2024:ICC

[Nie24] Marko Niemimaa. Incorrect compliance and correct noncompliance with information security policies: a framework of rule-related information security behaviour. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002918.

Nedeljkovic:2022:CBM

[NJ22] Dusan Nedeljkovic and Zivana Jakovljevic. CNN based method for the development of cyber-attacks detection algorithms in industrial control systems. Computers & Security, 114(??):Article 102585, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821004089.

Nanda:2024:EUS

[NJS^{+}24] Ashish Nanda, Jongkil Jay Jeong, Syed Wajid Ali Shah, Mohammad Nosouhi, and Robin Doss. Examining usable security features and user perceptions of Physical Authentication Devices. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005746.

Nazir:2021:NCO

[NK21] Anjum Nazir and Rizwan Ahmed Khan. A novel combinatorial optimization based feature selection method for network intrusion detection. Computers & Security, 102(??):Article 102164, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304375.

Nasayreh:2025:ADC

[NKA^{+}25] Ahmad Nasayreh, Haris M. Khalid, Hamza K. Alkhateeb, Jalal Al-Manaseer, Abdulla Ismail, and Hasan Gharaibeh. Automated detection of cyber attacks in healthcare systems: a novel scheme with advanced feature extraction and classification. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005947.

Nayfeh:2023:MLM

[NLA^{+}23] Mohammad Nayfeh, Yuchen Li, Khair Al Shamaileh, Vijay Devabhaktuni, and Naima Kaabouch. Machine learning modeling of GPS features with applications to UAV location spoofing detection and classification. Computers & Security, 126(??): ??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004771.

Nehme:2024:AMF

[NLW24] Alaa Nehme, Meng (Leah) Li, and Merrill Warkentin. Adaptive and maladaptive factors behind password manager use: a hope-extended protection motivation perspective. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002463.

Nugraha:2021:TFT

[NM21] Yudhistira Nugraha and Andrew Martin. Towards a framework for trustworthy data security level agreement in cloud procurement. Computers & Security, 106(??):Article 102266, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000900.

N:2024:EFI

[NM24] Balagopal N. and Saji K. Mathew. Exploring the factors influencing information security policy compliance and violations: a systematic literature review. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003675.

Nguyen:2021:MPA

[NMCRB21] Quoc-Thông Nguyen, An Mai, Lionel Chagas, and Nadège Reverdy-Bruas. Microscopic printing analysis and application for classification of source printer. Computers & Security, 108(??):Article 102320, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001449.

Nirmalapriya:2023:ASN

[NMLN23] G. Nirmalapriya, Balajee Maram, Ramanathan Lakshmanan, and M. Navaneethakrishnan. ASCA-squeeze net: Aquila sine cosine algorithm enabled hybrid deep learning networks for digital image forgery detection. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000652.

Noor:2025:IKF

[NMT25] Ahmad Fairuz Mohamed Noor, Sedigheh Moghavvemi, and Farzana Parveen Tajudeen. Identifying key factors of cybersecurity readiness in organizations: Insights from Malaysian critical infrastructure. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003633.

Najar:2024:CSS

[NN24] Ashfaq Ahmad Najar and S. Manohar Naik. Cyber-secure SDN: a CNN-based approach for efficient detection and mitigation of DDoS attacks. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000178.

Naseer:2023:MTA

[NNA^{+}23] Ayesha Naseer, Humza Naseer, Atif Ahmad, Sean B. Maynard, and Adil Masood Siddiqui. Moving towards agile cybersecurity incident response: a case study exploring the enabling role of big data analytics-embedded dynamic capabilities. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004352.

Niakanlahiji:2023:TPD

[NOVJ23] Amirreza Niakanlahiji, Soeren Orlowski, Alireza Vahid, and J. Haadi Jafarian. Toward practical defense against traffic analysis attacks on encrypted DNS traffic. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003935.

Najafi:2024:YYF

[NPCM24] Pejman Najafi, Wenzel Puenter, Feng Cheng, and Christoph Meinel. You are your friends: Detecting malware via guiltby-association and exempt-by-reputation. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823004297.

Naqvi:2023:MSA

[NPF^{+}23] Bilal Naqvi, Kseniia Perova, Ali Farooq, Imran Makhdoom, Shola Oyedeji, and Jari Porras. Mitigation strategies against the phishing attacks: a systematic literature review. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002973.

Nichelini:2023:CHI

[NPL^{+}23] Alessandro Nichelini, Carlo Alberto Pozzoli, Stefano Longari, Michele Carminati, and Stefano Zanero. CANova: a hybrid intrusion detection framework based on automatic signal classification for CAN. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000767.

Nyangaresi:2022:EHP

[NR22] Vincent Omollo Nyangaresi and Anthony Joachim Rodrigues. Efficient handover protocol for 5G and beyond networks. Computers & Security, 113(??):Article 102546, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003709.

Nautiyal:2024:FMO

[NR24] Lata Nautiyal and Awais Rashid. A framework for mapping organisational workforce knowledge profile in cyber security. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002281.

Ngwobia:2023:DMP

[NRKK23] Sunday Cosmos Ngwobia, Anca Ralescu, David Kapp, and Temesgen Kebede. Detection of malicious PE files using synthesized DNA artifacts. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300367X.

Neisari:2021:SRD

[NRS21] Ashraf Neisari, Luis Rueda, and Sherif Saad. Spam review detection using self-organizing maps and convolutional neural networks. Computers & Security, 106(??):Article 102274, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000985.

Niazi:2020:MMS

[NSA^{+}20] Mahmood Niazi, Ashraf Mohammed Saeed, Mohammad Alshayeb, Sajjad Mahmood, and Saad Zafar. A maturity model for secure requirements engineering. Computers & Security, 95(??):Article 101852, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301243.

Nganga:2025:CRC

[NSLM25] Allan Nganga, Joel Scanlan, Margareta Lützhöft, and Steven Mallam. Cyber risk communication during vessel incident management: a case study. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002962.

Nasr:2022:PJY

[NTBH^{+}22] Tony Nasr, Sadegh Torabi, Elias Bou-Harb, Claude Fachkha, and Chadi Assi. Power jacking your station: In-depth security analysis of electric vehicle charging station management systems. Computers & Security, 112(??):Article 102511, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003357.

Nolan:2025:NCT

[NTH25] Louis Nolan, Denise L. Tennant, and Deanna House. Navigating challenging terrain surrounding DoD response to homeland attacks on critical infrastructure: Case studies of prior incidents utilizing an extended taxonomy of cyber harms. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005030.

Niu:2025:PPS

[NZT^{+}25] Yutong Niu, Jian Zhang, Zhangguo Tang, Hao Yan, Min Zhu, and Huanzhou Li. Privacy-preserving spatiotemporal trajectory generalization publishing scheme with differential privacy. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002032.

Niu:2022:UAM

[NZZ^{+}22] Weina Niu, Jie Zhou, Yibin Zhao, Xiaosong Zhang, Yujie Peng, and Cheng Huang. Uncovering APT malware traffic using deep learning combined with time sequence and association analysis. Computers & Security, 120(??):Article 102809, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002036.

Omolara:2025:DUM

[OA25] Abiodun Esther Omolara and Moatsum Alawida. DaE2: Unmasking malicious URLs by leveraging diverse and efficient ensemble machine learning for online security. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004759.

Omolara:2022:ITS

[OAA^{+}22] Abiodun Esther Omolara, Abdullah Alabdulatif, Oludare Isaac Abiodun, Moatsum Alawida, Abdulatif Alabdulatif, Wafa’ Hamdan Alshoura, and Humaira Arshad. The Internet of Things security: a survey encompassing unexplored areas and new insights. Computers & Security, 112(??):Article 102494, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003187.

Okonkwo:2025:PSW

[OANA25] Chinedu Okonkwo, Ibukun Awolusi, Chukwuma Nnaji, and Oluwafemi Akanfe. Privacy and security of wearable Internet of Things: a scoping review and conceptual framework development for safety and health management in construction. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005819.

Ooi:2023:IDS

[OBK^{+}23] Sian En Ooi, Razvan Beuran, Takayuki Kuroda, Takuya Kuwahara, Ryosuke Hotchi, Norihito Fujita, and Yasuo Tan. Intent-driven secure system design: Methodology and implementation. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003479.

Ogbanufe:2021:ESP

[OCB21] Obi Ogbanufe, Robert E. Crossler, and David Biros. Exploring stewardship: a precursor to voluntary security behaviors. Computers & Security, 109(??):Article 102397, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002212.

Ogbanufe:2023:VCP

[OCB23] Obi Ogbanufe, Robert E. Crossler, and David Biros. The valued coexistence of protection motivation and stewardship in information security behaviors. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003522.

Obert:2020:DRE

[OCJ20] James Obert, Adrian Chavez, and Jay Johnson. Distributed renewable energy resource trust metrics and secure routing. Computers & Security, 88(??):Article 101620, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404819301671.

Olimpio:2023:MUI

[OCMF23] Gilberto Olímpio, Lasaro Camargos, Rodrigo Sanches Miani, and Elaine Ribeiro Faria. Model update for intrusion detection: Analyzing the performance of delayed labeling and active learning strategies. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003619.

Ozkan:2022:EAC

[ÖEÖ22] Baris Özkan, Mehmet Erdem, and Eren Özceylan. Evaluation of Asian countries using data center security index: a spherical fuzzy AHP-based EDAS approach. Computers & Security, 122(??):Article 102900, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002930.

Okonkwo:2025:GRF

[OFH^{+}25] Zulu Okonkwo, Ernest Foo, Zhe Hou, Qinyi Li, and Zahra Jadidi. A graph representation framework for encrypted network traffic classification. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004395.

Ortega:2020:DFP

[OFIdD^{+}20] David Ortega, Alberto Fernández-Isabel, Isaac Martín de Diego, Cristina Conde, and Enrique Cabello. Dynamic facial presentation attack detection for automated border control systems. Computers & Security, 92(??):Article 101744, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300298.

Ogbanufe:2023:CEB

[OG23] Obi Ogbanufe and Ling Ge. A comparative evaluation of behavioral security motives: Protection, intrinsic, and identity motivations. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000469.

Ogbanufe:2021:EEU

[Ogb21] Obi Ogbanufe. Enhancing end-user roles in information security: Exploring the setting, situation, and identity. Computers & Security, 108(??):Article 102340, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001644.

Ogbanufe:2025:JDI

[OJH25] Obi Ogbanufe, Mary C. Jones, and Julie I. Hancock. Job demands, identity, and outcomes: the mediating role of cynicism among cybersecurity-focused employees. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005832.

Onieva:2024:MSN

[OJL24] Jose A. Onieva, Pablo Pérez Jiménez, and Javier López. Malware similarity and a new fuzzy hash: Compound Code Block Hash (CCBHash). Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001573.

Ofte:2023:USA

[OK23] Håvard Jakobsen Ofte and Sokratis Katsikas. Understanding situation awareness in SOCs, a systematic literature review. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004618.

Odeleye:2023:VST

[OLH^{+}23] Blessing Odeleye, George Loukas, Ryan Heartfield, Georgia Sakellari, Emmanouil Panaousis, and Fotios Spyridonis. Virtually secure: a taxonomic assessment of cybersecurity challenges in virtual reality environments. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003431.

Olukoya:2022:AFE

[Olu22] Oluwafemi Olukoya. Assessing frameworks for eliciting privacy and security requirements from laws and regulations. Computers & Security, 117(??):Article 102697, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000955.

Olukoya:2020:SOV

[OMO20a] Oluwafemi Olukoya, Lewis Mackenzie, and Inah Omoronyia. Security-oriented view of app behaviour using textual descriptions and user-granted permission requests. Computers & Security, 89(??):Article 101685, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302238.

Olukoya:2020:TUU

[OMO20b] Oluwafemi Olukoya, Lewis Mackenzie, and Inah Omoronyia. Towards using unstructured user input request for malware detection. Computers & Security, 93(??):Article 101783, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300687.

Onwubiko:2020:FMM

[Onw20] Cyril Onwubiko. Fraud matrix: a morphological and analysisbased classification and taxonomy of fraud. Computers & Security, 96(??):Article 101900, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301759.

Olakanmi:2021:CRA

[OO21] Oladayo Olufemi Olakanmi and Kehinde Odeyemi. Compromiseresilient anonymous mutual authentication scheme for n by m-times ubiquitous mobile cloud computing services. Computers & Security, 108(??):Article 102369, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001930.

Oh:2020:WDD

[OPK20] Dong Bin Oh, Kyung Ho Park, and Huy Kang Kim. Dewipimization: Detection of data wiping traces for investigating NTFS file system. Computers & Security, 99(??): Article 102034, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820303072.

OShaughnessy:2022:IBM

[OS22] Stephen O’Shaughnessy and Stephen Sheridan. Image-based malware classification hybrid framework based on space-filling curves. Computers & Security, 116(??):Article 102660, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000591.

Obaidat:2022:PJN

[OSPP22] Islam Obaidat, Meera Sridhar, Khue M. Pham, and Phu H. Phung. Jadeite: a novel image-behavior-based approach for Java malware detection using deep learning. Computers & Security, 113(??):Article 102547, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003710.

Oishi:2020:SDP

[OSTO20] Keiichiro Oishi, Yuichi Sei, Yasuyuki Tahara, and Akihiko Ohsuga. Semantic diversity: Privacy considering distance between values of sensitive attribute. Computers & Security, 94(??):Article 101823, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300973.

Okey:2023:ICC

[OUR^{+}23] Ogobuchi Daniel Okey, Ekikere Umoren Udo, Renata Lopes Rosa, Demostenes Zegarra Rodríguez, and João Henrique Kleinschmidt. Investigating ChatGPT and cybersecurity: a perspective on topic modeling and sentiment analysis. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003863.

Ou:2022:PSS

[OX22] Fan Ou and Jian Xu. S^{3}Feature: a static sensitive subgraphbased feature for Android malware detection. Computers & Security, 112(??):Article 102513, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003370.

Oroni:2025:ECS

[OXNA25] Chrispus Zacharia Oroni, Fu Xianping, Daniela Daniel Ndunguru, and Arsenyan Ani. Enhancing cyber safety in elearning environment through cybersecurity awareness and information security compliance: PLS-SEM and FsQCA analysis. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005820.

Oakley:2020:PPF

[OYZ^{+}20] Jonathan Oakley, Lu Yu, Xingsi Zhong, Ganesh Kumar Venayagamoorthy, and Richard Brooks. Protocol proxy: an FTEbased covert channel. Computers & Security, 92(??):Article 101777, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300626.

Ozturk:2024:CSB

[Özt24] M. Maruf Öztürk. A cosine similarity-based labeling technique for vulnerability type detection using source codes. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400364X.

Ouyang:2021:CEC

[OZW^{+}21] Zhiyou Ouyang, Xu Zhai, Jinran Wu, Jian Yang, Dong Yue, Chunxia Dou, and Tengfei Zhang. A cloud endpoint coordinating CAPTCHA based on multi-view stacking ensemble. Computers & Security, 103(??):Article 102178, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482100002X.

Parvez:2020:SVB

[PA20] Mohammad Tanvir Parvez and Suliman A. Alsuhibany. Segmentation-validation based handwritten Arabic CAPTCHA generation. Computers & Security, 95(??):Article 101829, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301048.

Pritee:2024:MLD

[PAA^{+}24] Zinniya Taffannum Pritee, Mehedi Hasan Anik, Saida Binta Alam, Jamin Rahman Jim, Md Mohsin Kabir, and M. F. Mridha. Machine learning and deep learning for user authentication and authorization in cybersecurity: a stateof-the-art review. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000488.

Paya:2024:ARD

[PAGDG24] Antonio Paya, Sergio Arroni, Vicente García-Díaz, and Alberto Gómez. Apollon: a robust defense system against Adversarial Machine Learning attacks in Intrusion Detection Systems. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300456X.

Palma:2024:CAS

[PAM^{+}24] Alessandro Palma, Giacomo Acitelli, Andrea Marrella, Silvia Bonomi, and Marco Angelini. A compliance assessment system for Incident Management process. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003754.

Palma:2025:BSA

[PB25] Alessandro Palma and Silvia Bonomi. Behind the scenes of attack graphs: Vulnerable network generator for indepth experimental evaluation of attack graph scalability. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002652.

Prabhaker:2024:GDH

[PBA24] Nilin Prabhaker, Ghanshyam S. Bopche, and Michael Arock. Generation and deployment of honeytokens in relational databases for cyber deception. Computers & Security, 146 (??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003377.

Pereira:2025:AAG

[PBdL25] Gabriela Leite Pereira, Leonardo Silvino Brito, and Sidney Marlon Lopes de Lima. Antivirus applied to Google Chrome’s extension malware. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001543.

Pashamokhtari:2022:AQR

[PBG22] Arman Pashamokhtari, Gustavo Batista, and Hassan Habibi Gharakheili. AdIoTack: Quantifying and refining resilience of decision tree ensemble inference models against adversarial volumetric attacks on IoT networks. Computers & Security, 120(??):Article 102801, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200195X.

Pu:2022:LSA

[PC22] Cong Pu and Kim-Kwang Raymond Choo. Lightweight Sybil attack detection in IoT based on Bloom filter and physical unclonable function. Computers & Security, 113(??): Article 102541, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003655.

Prasad:2024:PDS

[PC24] Arvind Prasad and Shalini Chandra. PhiUSIIL: a diverse security profile empowered phishing URL detection framework based on similarity index and incremental learning. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004558.

Pasikhan:2023:IHI

[PCG23] Aryan Mohammadi Pasikhan, John A. Clark, and Prosanta Gope. Incremental hybrid intrusion detection for 6LoW-PAN. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003577.

Patsakis:2020:ECD

[PCK20] Constantinos Patsakis, Fran Casino, and Vasilios Katos. Encrypted and covert DNS queries for botnets: Challenges and countermeasures. Computers & Security, 88(??):Article 101614, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740481831321X.

Paruchuri:2020:GRE

[PCR20] Shravya Paruchuri, Andrew Case, and Golden G. Richard. Gaslight revisited: Efficient and powerful fuzzing of digital forensics tools. Computers & Security, 97(??):Article 101986, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302583.

Paya:2022:EAS

[PCR22] Antonio Paya, Alba Cotarelo, and Jose Manuel Redondo. Egida: Automated security configuration deployment systems with early error detection. Computers & Security, 116 (??):Article 102638, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000372.

Pavithra:2025:ONS

[PD25a] P. S. Pavithra and P. Durgadevi. Optimizing network security: Weighted average ensemble of BPNN and RELM in EPRN-WPS intrusion detection. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005959.

Prasad:2025:PDN

[PD25b] Y. Bhanu Prasad and Venkatesulu Dondeti. PDSMV3-DCRNN: a novel ensemble deep learning framework for enhancing phishing detection and URL extraction. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004280.

Pascual:2024:HTA

[PdARD24] Hugo Pascual, Jose M. del Alamo, David Rodriguez, and Juan C. Dueñas. Hunter: Tracing anycast communications to uncover cross-border personal data transfers. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482400124X.

Pizzi:2025:VDA

[PDML25] Simeone Pizzi, Samuele Doria, Nicholas Miazzo, and Eleonora Losiouk. VirtualPatch: Distributing Android security patches through Android virtualization. Computers & Security, 157 (??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825003049.

Prasad:2025:SCT

[PDWF25] Nilantha Prasad, Abebe Diro, Matthew Warren, and Mahesh Fernando. A survey of cyber threat attribution: Challenges, techniques, and future directions. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825002950.

Pamies-Estrems:2023:SAP

[PEGA23] David Pàmies-Estrems and Joaquin Garcia-Alfaro. On the self-adjustment of privacy safeguards for query log streams. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003607.

Posey:2020:EEO

[PF20] Clay Posey and Robert Folger. An exploratory examination of organizational insiders’ descriptive and normative perceptions of cyber-relevant rights and responsibilities. Computers & Security, 99(??):Article 102038, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303114.

Pan:2024:FSG

[PFG^{+}24] Hongyu Pan, Yong Fang, Wenbo Guo, Yijia Xu, and Changhui Wang. Few-shot graph classification on cross-site scripting attacks detection. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000506.

Pilton:2021:EPD

[PFHB21] Callum Pilton, Shamal Faily, and Jane Henriksen-Bulmer. Evaluating privacy — determining user privacy expectations on the web. Computers & Security, 105(??):Article 102241, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000651.

Philippou:2020:CAS

[PFR20] Eleni Philippou, Sylvain Frey, and Awais Rashid. Contextualising and aligning security metrics and business objectives: a GQM-based methodology. Computers & Security, 88(??):Article 101634, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404818305170.

Pi:2024:RAT

[PGC^{+}24] Ben Pi, Chun Guo, Yunhe Cui, Guowei Shen, Jialong Yang, and Yuan Ping. Remote access trojan traffic early detection method based on Markov matrices and deep learning. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005382.

Peng:2024:SSN

[PGP^{+}24] Jiaxin Peng, Chun Guo, Yuan Ping, Yunhe Cui, Yi Chen, and Guowei Shen. SNDMI: Spyware network traffic detection method based on inducement operations. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482400107X.

Parameswari:2024:HRS

[PGRS24] A. Parameswari, R. Ganeshan, V. Ragavi, and M. Shereesha. Hybrid rat swarm hunter prey optimization trained deep learning for network intrusion detection using CNN features. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005655.

Pohn:2023:FAA

[PGZB23] Daniela Pöhn, Nils Gruschka, Leonhard Ziegler, and Andre Büttner. A framework for analyzing authentication risks in account networks. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300425X.

Patil:2020:EPP

[PHH^{+}20] Akash Suresh Patil, Rafik Hamza, Alzubair Hassan, Nan Jiang, Hongyang Yan, and Jin Li. Efficient privacy-preserving authentication protocol using PUFs with blockchain smart contracts. Computers & Security, 97(??):Article 101958, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302340.

Pei:2025:DAD

[PHT^{+}25] Jinchuan Pei, Yuxiang Hu, Le Tian, Xinglong Pei, and Zihao Wang. Dynamic anomaly detection using in-band network telemetry and GCN for cloud-edge collaborative networks. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001117.

Pitropakis:2025:ISS

[Pit25] Nikolaos Pitropakis. IFIP SEC 2023 and 2024 selected papers. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002858.

Preuveneers:2023:PPC

[PJ23] Davy Preuveneers and Wouter Joosen. Privacy-preserving correlation of cross-organizational cyber threat intelligence with private graph intersections. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004157.

Park:2023:FAM

[PJL23] Sung Bum Park, Hyo Jin Jo, and Dong Hoon Lee. Flooding attack mitigator for in-vehicle CAN using fault confinement in CAN protocol. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000019.

Park:2021:MCA

[PKK^{+}21] Sang-Ok Park, Ohmin Kwon, Yonggon Kim, Sang Kil Cha, and Hyunsoo Yoon. Mind control attack: Undermining deep learning with GPU memory exploitation. Computers & Security, 102(??):Article 102115, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303886.

Park:2024:FIF

[PKPC24] Eun Hee Park, Jongwoo Kim, Young Soon Park, and Kyung Hee Chun. Facilitating and impeding factors to insiders’ prosocial rule breaking in South Korea. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004820.

Pirca:2023:EEE

[PL23] Ana Maria Pirca and Harjinder Singh Lallie. An empirical evaluation of the effectiveness of attack graphs and MITRE ATT&CK matrices in aiding cyber attack perception amongst decision-makers. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001645.

Pu:2022:DDE

[PLB^{+}22] Wenxi Pu, Siyuan Li, Gregory J. Bott, Marie Esposito, and Jason Bennett Thatcher. To disclose or not to disclose: an evaluation of the effects of information control and social network transparency. Computers & Security, 112(??):Article 102509, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003333.

Plappert:2023:EAH

[PLE^{+}23] Christian Plappert, Dominik Lorych, Michael Eckel, Lukas Jäger, Andreas Fuchs, and Ronald Heddergott. Evaluating the applicability of hardware trust anchors for automotive applications. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004248.

Pattnaik:2023:PNE

[PLN23] Nandita Pattnaik, Shujun Li, and Jason R. C. Nurse. Perspectives of non-expert users on cyber security and privacy: an analysis of online discussions on Twitter. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200400X.

Pinhero:2021:MDE

[PLP^{+}21] Anson Pinhero, Anupama M. L., Vinod P., C. A. Visaggio, Aneesh N., Abhijith S., and AnanthaKrishnan S. Malware detection employed by visualization and deep neural network. Computers & Security, 105(??):Article 102247, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000717.

Paananen:2020:SAI

[PLS20] Hanna Paananen, Michael Lapke, and Mikko Siponen. State of the art in information security policy development. Computers & Security, 88(??):Article 101608, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818313002.

Peng:2021:EMG

[PLW^{+}21] Wenyu Peng, Renyang Liu, Ruxin Wang, Taining Cheng, Zifeng Wu, Li Cai, and Wei Zhou. EnsembleFool: a method to generate adversarial examples based on model fusion strategy. Computers & Security, 107(??):Article 102317, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001413.

Pang:2023:GBF

[PLW^{+}23] Chengbin Pang, Hongbin Liu, Yifan Wang, Neil Zhenqiang Gong, Bing Mao, and Jun Xu. Generation-based fuzzing? Don’t build a new generator, reuse! Computers & Security, 129 (??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000883.

Pan:2025:FHT

[PLW25] Gaoyuan Pan, Huan Li, and Jian Wang. A fast hardware Trojan detection method with parallel clustering for large-scale gate-level netlists. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002597.

Pandey:2025:CEB

[PM25] Nimisha Pandey and Pramod Kumar Mishra. Conditional entropy-based hybrid DDoS detection model for IoT networks. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005042.

Pandey:2023:EDM

[PMA^{+}23] Bishwajeet Kumar Pandey, Veeramanickam M. R. M., Shabeer Ahmad, Ciro Rodriguez, and Doris Esenarro. ExpSSOA-Deep maxout: Exponential Shuffled shepherd optimization based Deep maxout network for intrusion detection using big data in cloud computing framework. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003674.

Pour:2020:DDC

[PMF^{+}20] Morteza Safaei Pour, Antonio Mangino, Kurt Friday, Matthias Rathbun, Elias Bou-Harb, Farkhund Iqbal, Sagar Samtani, Jorge Crichigno, and Nasir Ghani. On data-driven curation, learning, and analysis for inferring evolving Internet-of-Things (IoT) botnets in the wild. Computers & Security, 91 (??):Article 101707, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302445.

Pavithran:2021:NCB

[PMNL21] Pramod Pavithran, Sheena Mathew, Suyel Namasudra, and Pascal Lorenz. A novel cryptosystem based on DNA cryptography and randomly generated Mealy machine. Computers & Security, 104(??):Article 102160, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304338.

Patterson:2023:LCS

[PNF23] Clare M. Patterson, Jason R. C. Nurse, and Virginia N. L. Franqueira. Learning from cyber security incidents: a systematic review and future research agenda. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002195.

Patterson:2024:SDT

[PNF24] Clare M. Patterson, Jason R. C. Nurse, and Virginia N. L. Franqueira. “I don’t think we’re there yet”: the practices and challenges of organisational learning from cyber security incidents. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823006090.

Pour:2023:CSR

[PNFBH23] Morteza Safaei Pour, Christelle Nader, Kurt Friday, and Elias Bou-Harb. A comprehensive survey of recent Internet measurement techniques for cyber security. Computers & Security, 128 (??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000330.

Palanisamy:2020:CBY

[PNK20] Rathika Palanisamy, Azah Anir Norman, and Miss Laiha Mat Kiah. Compliance with bring your own device security policies in organizations: a systematic literature review. Computers & Security, 98(??):Article 101998, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302716.

Patel:2024:UES

[PPGC24] Chintan Patel, Aryan Pasikhani, Prosanta Gope, and John Clark. User-empowered secure privacy-preserving authentication scheme for Digital Twin. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000944.

Punitha:2025:DSR

[PRPS25] A. Punitha, P. Ramani, Ezhilarasi P., and Sridhar S. Dynamically stabilized recurrent neural network optimized with intensified sand cat swarm optimization for intrusion detection in wireless sensor network. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824003997.

P:2022:IDB

[PRTV22] Ramkumar M. P., P. V. Bhaskar Reddy, J. T. Thirukrishna, and Ch. Vidyadhari. Intrusion detection in big data using hybrid feature fusion and optimization enabled deep learning based on Spark architecture. Computers & Security, 116(??):Article 102668, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000670.

Plotnek:2021:CTH

[PS21] Jordan J. Plotnek and Jill Slay. Cyber terrorism: a homogenized taxonomy and definition. Computers & Security, 102(??):Article 102145, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304181.

Petrykina:2021:NUT

[PSCT21] Yelena Petrykina, Hadas Schwartz-Chassidim, and Eran Toch. Nudging users towards online safety using gamified environments. Computers & Security, 108(??):Article 102270, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000948.

Parast:2022:CCS

[PSN^{+}22] Fatemeh Khoda Parast, Chandni Sindhav, Seema Nikam, Hadiseh Izadi Yekta, Kenneth B. Kent, and Saqib Hakak. Cloud computing security: a survey of service-based models. Computers & Security, 114(??):Article 102580, March 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003977.

Patil:2022:MFB

[PSP^{+}22] Rajendra Patil, Vinay Sachidananda, Hongyi Peng, Akshay Sachdeva, and Mohan Gurusamy. MARK: Fill in the blanks through a JointGAN based data augmentation for network anomaly detection. Computers & Security, 119(??):Article 102759, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001547.

Park:2022:UMD

[PSY^{+}22] Kyung Ho Park, Hyun Min Song, Jeong Do Yoo, Su-Youn Hong, Byoungmo Cho, Kwangsoo Kim, and Huy Kang Kim. Unsupervised malicious domain detection with less labeling effort. Computers & Security, 116(??):Article 102662, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482200061X.

Padhan:2025:TDM

[PT25] Sushree Padhan and Ashok Kumar Turuk. A technique to detect and mitigate false data injection attacks in cyber-physical systems. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005595.

Prasad:2020:UFS

[PTD20] Mahendra Prasad, Sachin Tripathi, and Keshav Dahal. Unsupervised feature selection and cluster center initialization based arbitrary shaped clusters for intrusion detection. Computers & Security, 99(??):Article 102062, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303357.

Ponte:2025:SIP

[PTD^{+}25] Andrea Ponte, Dmitrijs Trizna, Luca Demetrio, Battista Biggio, Ivan Tesfai Ogbu, and Fabio Roli. SLIFER: Investigating performance and robustness of malware detection pipelines. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005704.

Pereira-Vale:2021:SMB

[PVFM^{+}21] Anelis Pereira-Vale, Eduardo B. Fernandez, Raúl Monge, Hernán Astudillo, and Gastón Márquez. Security in microservice-based systems: a multivocal literature review. Computers & Security, 103(??):Article 102200, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000249.

Prummer:2024:SRC

[PvSvdB24] Julia Prümmer, Tommy van Steen, and Bibi van den Berg. A systematic review of current cybersecurity training methods. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004959.

Prummer:2025:AEC

[PvSvdB25] Julia Prümmer, Tommy van Steen, and Bibi van den Berg. Assessing the effect of cybersecurity training on end-users: a meta-analysis. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740482400511X.

Pham:2021:SSA

[PX21] Minh Pham and Kaiqi Xiong. A survey on security attacks and defense techniques for connected and autonomous vehicles. Computers & Security, 109(??):Article 102269, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000936.

Pei:2020:ADL

[PYT20] Xinjun Pei, Long Yu, and Shengwei Tian. AMalNet: a deep learning framework based on graph convolutional networks for malware detection. Computers & Security, 93(??):Article 101792, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300778.

Qamar:2023:STA

[QAA23] Sara Qamar, Zahid Anwar, and Mehreen Afzal. A systematic threat analysis and defense strategies for the metaverse and extended reality systems. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000378.

Qiao:2025:ANF

[QGW^{+}25] Sibo Qiao, Qiang Guo, Min Wang, Haohao Zhu, Joel J. P. C. Rodrigues, and Zhihan Lyu. Advances in network flow watermarking: a survey. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003426.

Qing:2025:LPP

[QIN25] Haohua Qing, Roliana Ibrahim, and Hui Wen Nies. Location privacy protection method based on social network platform. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825003001.

Qin:2024:HCD

[QJDD24] Xingsheng Qin, Frank Jiang, Chengzu Dong, and Robin Doss. A hybrid cyber defense framework for reconnaissance attack in industrial control systems. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004169.

Qayyum:2022:MFL

[QJQ22] Adnan Qayyum, Muhammad Umar Janjua, and Junaid Qadir. Making federated learning robust to adversarial attacks by learning data and model association. Computers & Security, 121(??):Article 102827, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002218.

Qian:2025:EBD

[QLL^{+}25] Yaguan Qian, Zejie Lian, Yiming Li, Wei Wang, Zhaoquan Gu, Bin Wang, and Yanchun Zhang. Evading backdoor defenses: Concealing genuine backdoors through scapegoat strategy. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005315.

Qin:2025:DPL

[QLZ^{+}25] Zhongyuan Qin, Kefei Lu, Qunfang Zhang, Dinglian Wang, and Liquan Chen. Dynamic privacy loss management: the innovation of α-confidence differential privacy filters and odometers. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000276.

Qiu:2022:HID

[QMC^{+}22] Weicheng Qiu, Yinghua Ma, Xiuzhen Chen, Haiyang Yu, and Lixing Chen. Hybrid intrusion detection system based on Dempster–Shafer evidence theory. Computers & Security, 117(??):Article 102709, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001079.

Qian:2020:SEA

[QMW^{+}20a] Yaguan Qian, Danfeng Ma, Bin Wang, Jun Pan, Jiamin Wang, Zhaoquan Gu, Jianhai Chen, Wujie Zhou, and Jingsheng Lei. Spot evasion attacks: Adversarial examples for license plate recognition systems with convolutional neural networks. Computers & Security, 95(??):Article 101826, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301000.

Qin:2020:AUA

[QMW20b] Zhi-Quan Qin, Xing-Kong Ma, and Yong-Jun Wang. ADSAD: an unsupervised attention-based discrete sequence anomaly detection framework for network security analysis. Computers & Security, 99(??):Article 102070, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303436.

Qiao:2023:NMW

[QMZ^{+}23] Tong Qiao, Yuyan Ma, Ning Zheng, Hanzhou Wu, Yanli Chen, Ming Xu, and Xiangyang Luo. A novel model watermarking for protecting generative adversarial network. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000123.

Qu:2025:MTD

[QMZ^{+}25] Yanze Qu, Hailong Ma, Chaofan Zheng, Yiming Jiang, and Wenbo Wang. A malware traffic detection method based on victim-attacker interaction patterns. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825001750.

Qin:2023:USA

[QPL^{+}23] Chuan Qin, Jiaqian Peng, Puzhuo Liu, Yaowen Zheng, Kai Cheng, Weidong Zhang, and Limin Sun. UCRF: Static analyzing firmware to generate under-constrained seed for fuzzing SOHO router. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000676.

Quamara:2022:SSS

[QS22] Sidharth Quamara and Awadhesh Kumar Singh. A systematic survey on security concerns in cryptocurrencies:

State-of-the-art and perspectives. Computers & Security, 113(??):Article 102548, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003722.

Qu:2024:TID

[QSA24] Aiyan Qu, Qiuhui Shen, and Gholamreza Ahmadi. Towards intrusion detection in fog environments using generative adversarial network and long short-term memory network. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003092.

Quan:2023:FSP

[QWLL23] Hanyu Quan, Boyang Wang, Ming Li, and Iraklis Leontiadis. FastReach: a system for privacy-preserving reachability queries over location data. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004236.

Qian:2022:VIA

[QWW^{+}22] Yaguan Qian, Jiamin Wang, Haijiang Wang, Zhaoquan Gu, Bin Wang, Shaoning Zeng, and Wassim Swaileh. Visually imperceptible adversarial patch attacks. Computers & Security, 123(??):Article 102943, December 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003352.

Qu:2022:YBR

[QXS^{+}22] Leilei Qu, Ruojin Xiao, Wenchang Shi, Keman Huang, Bo Qin, and Bin Liang. Your behaviors reveal what you need: a practical scheme based on user behaviors for personalized security nudges. Computers & Security, 122(??):Article 102891, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002851.

Qu:2023:IFT

[QXS23] Leilei Qu, Ruojin Xiao, and Wenchang Shi. Interactions of framing and timing in nudging online game security. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003546.

Qin:2022:FBH

[QY22] Yi Qin and Chuan Yue. Fuzzing-based hard-label blackbox attacks against machine learning models. Computers & Security, 117(??):Article 102694, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200092X.

Qiang:2022:ERM

[QYJ22] Weizhong Qiang, Lin Yang, and Hai Jin. Efficient and robust malware detection based on control flow traces using deep neural networks. Computers & Security, 122(??):Article 102871, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002656.

Qi:2025:AAB

[QYW^{+}25] Yinhao Qi, Chuyi Yan, Zehui Wang, Chen Zhang, Song Liu, Zhigang Lu, and Bo Jiang. ATHITD: Attention-based temporal heterogeneous graph neural network for insider threat detection. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002767.

Qin:2024:MSD

[QYYH24] Yang Qin, Xiaofan Yang, Lu-Xing Yang, and Kaifan Huang. Modeling and study of defense outsourcing against advanced persistent threat through impulsive differential game approach. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003080.

Qin:2023:AAO

[QZP23] Yingxin Qin, Kejia Zhang, and Haiwei Pan. Adversarial attack for object detectors under complex conditions. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300370X.

Qiao:2025:DDP

[QZS^{+}25] Sibo Qiao, Haohao Zhu, Lin Sha, Min Wang, and Qiang Guo. DynMark: a dynamic packet counting watermarking scheme for robust traffic tracing in network flows. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002603.

Qiao:2022:AMS

[QZT^{+}22] Yanchen Qiao, Weizhe Zhang, Zhicheng Tian, Laurence T. Yang, Yang Liu, and Mamoun Alazab. Adversarial malware sample generation method based on the prototype of deep learning detector. Computers & Security, 119(??): Article 102762, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001572.

Roumani:2025:EFI

[RA25] Yaman Roumani and Mais Alraee. Examining the factors that impact the severity of cyberattacks on critical infrastructures. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824003791.

Ramezan:2025:UCI

[Ram25] Christopher A. Ramezan. Understanding the chief information security officer: Qualifications and responsibilities for cybersecurity leadership. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000525.

Rahman:2025:SEI

[RBHW25] Md Rayhanur Rahman, Setu Kumar Basak, Rezvan Mahdavi Hezaveh, and Laurie Williams. SoK: an empirical investigation of malware techniques in advanced persistent threat attacks. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825003074.

R:2025:EDR

[RBM^{+}25] Hannah Jessie Rani R, Amit Barve, Ashwini Malviya, Vivek Ranjan, Rubal Jeet, and Nilesh Bhosle. Enhancing detection rates in intrusion detection systems using fuzzy integration and computational intelligence. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002664.

Russo:2020:BNG

[RCA20] Enrico Russo, Gabriele Costa, and Alessandro Armando. Building next generation cyber ranges with CRACK. Computers & Security, 95(??):Article 101837, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301103.

Rao:2023:TMF

[RCA23] Siddharth Prakash Rao, Hsin-Yi Chen, and Tuomas Aura. Threat modeling framework for mobile communication systems. Computers & Security, 125(??):??, February 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004394.

Raman:2024:CBW

[RCA24] Raghu Raman, Prasad Calyam, and Krishnashree Achuthan. ChatGPT or Bard: Who is a better certified ethical hacker? Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001056.

Reeves:2021:GRH

[RCD21] A. Reeves, D. Calic, and P. Delfabbro. “Get a red-hot poker and open up my eyes, it’s so boring”^{1}: Employee perceptions of cybersecurity training. Computers & Security, 106(??):Article 102281, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100105X.

Reeves:2023:SGU

[RCD23] Andrew Reeves, Dragana Calic, and Paul Delfabbro. “Generic and unusable”: Understanding employee perceptions of cybersecurity training and measuring advice fatigue. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000470.

Reeves:2025:HCA

[RCD25] Andrew Reeves, Dragana Calic, and Paul Delfabbro. How to De-CyFa the actor-observer bias in cybersecurity fatigue: Building the CyFa measure of attribution styles and mitigation strategies. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740482400484X.

Radivojevic:2024:DNU

[RCKB24] Kristina Radivojevic, Nicholas Clark, Anna Klempay, and Paul Brenner. Defending novice user privacy: an evaluation of default web browser configurations. Computers & Security, 140 (??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000853.

Renaud:2023:CIG

[RD23] Karen Renaud and Marc Dupuis. Cybersecurity insights gleaned from world religions. Computers & Security, 132 (??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002365.

Ray-Dowling:2023:SMB

[RDHS23] Aratrika Ray-Dowling, Daqing Hou, and Stephanie Schuckers. Stationary mobile behavioral biometrics: a survey.

Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000949.

Ray-Dowling:2022:EMM

[RDHSB22] Aratrika Ray-Dowling, Daqing Hou, Stephanie Schuckers, and Abbie Barbir. Evaluating multi-modal mobile behavioral biometrics using public datasets. Computers & Security, 121(??):Article 102868, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002620.

Raman:2020:DAA

[RDM20] M. R. Gauthama Raman, Wenjie Dong, and Aditya Mathur. Deep autoencoders as anomaly detectors: Method and case study in a distributed water treatment plant. Computers & Security, 99(??):Article 102055, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030328X.

Rai:2025:MEE

[RDPR25] Anuj Rai, Somnath Dey, Pradeep Patidar, and Prakhar Rai. MoSFPAD: an end-to-end ensemble of MobileNet and Support Vector Classifier for fingerprint presentation attack detection. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824003742.

Ren:2025:DFF

[RDQL25] Ziting Ren, Yucong Duan, Qi Qi, and Lanhua Luo. Data-free fingerprinting technology for biometric classifiers. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000756.

Ruan:2020:GMF

[RDS20] Na Ruan, Ruoyu Deng, and Chunhua Su. GADM: Manual fake review detection for O2O commercial platforms. Computers & Security, 88(??):Article 101657, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302007.

Rus:2024:NRP

[REHS24] Adrian-Catalin Rus, Mohammed El-Hajj, and Dipti Kapoor Sarmah. NAISS: a reverse proxy approach to mitigate Mage-Cart’s e-skimmers in e-commerce. Computers & Security, 140 (??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000981.

Repetto:2023:AMD

[Rep23] Matteo Repetto. Adaptive monitoring, detection, and response for agile digital service chains. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002535.

Repetto:2025:OFP

[Rep25] Matteo Repetto. Otupy: a flexible, portable, and extensible framework for remote control of security functions. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482500286X.

Ribeiro:2023:DMD

[RFdS23] Marcos Aurélio Ribeiro, Mauro Sergio Pereira Fonseca, and Juliana de Santi. Detecting and mitigating DDoS attacks with moving target defense approach based on automated flow classification in SDN networks. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003723.

Rigaki:2023:SEM

[RG23] M. Rigaki and S. Garcia. Stealing and evading malware classifiers and antivirus at low false positive conditions. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001025.

Ren:2025:PBA

[RG25] Jiafeng Ren and Rong Geng. Provenance-based APT campaigns detection via masked graph representation learning. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004644.

Robles-Gonzalez:2023:PCA

[RGACPA23] Antonio Robles-González, Patricia Arias-Cabarcos, and Javier Parra-Arnau. Privacy-centered authentication: a new framework and analysis. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002638.

Ribeiro:2024:WFP

[RGC24] Liliana Ribeiro, Inês Sousa Guedes, and Carla Sofia Cardoso. Which factors predict susceptibility to phishing? an empirical study. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004686.

Rehman:2024:PDM

[RGG^{+}24] Zubaida Rehman, Iqbal Gondal, Mengmeng Ge, Hai Dong, Mark Gregory, and Zahir Tari. Proactive defense mechanism: Enhancing IoT security through diversity-based moving target defense and cyber deception. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005953.

Rostamipoor:2023:CFG

[RGP23] Maryam Rostamipoor, Seyedhamed Ghavamnia, and Michalis Polychronakis. Confine: Fine-grained system call filtering for container attack surface reduction. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002353.

Robles-Gonzalez:2020:LBF

[RGPAF20] Antonio Robles-González, Javier Parra-Arnau, and Jordi Forné. A LINDDUN-based framework for privacy threat analysis on identification and authentication processes. Computers & Security, 94(??):Article 101755, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300390.

Rashidibajgan:2021:SPP

[RHDF21] Samaneh Rashidibajgan, Thomas Hupperich, Robin Doss, and Anna Förster. Secure and privacy-preserving structure in opportunistic networks. Computers & Security, 104(??):Article 102208, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000328.

Riera:2022:NML

[RHH^{+}22] Tomás Sureda Riera, Juan-Ramón Bermejo Higuera, Javier Bermejo Higuera, José-Javier Martínez Herraiz, and Juan-Antonio Sicilia Montalvo. A new multi-label dataset for Web attacks CAPEC classification using machine learning techniques. Computers & Security, 120(??):Article 102788, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001833.

R:2021:SRS

[RJ21] Isaac Sajan R. and Jasper J. A secure routing scheme to mitigate attack in wireless adhoc sensor network. Computers & Security, 103(??):Article 102197, April 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000213.

Rustam:2024:MTD

[RJ24] Furqan Rustam and Anca Delia Jurcut. Malicious traffic detection in multi-environment networks using novel S-DATE and PSO-D-SEM approaches. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004741.

Rookard:2024:RRR

[RK24] Curtis Rookard and Anahita Khojandi. RRIoT: Recurrent reinforcement learning for cyber threat detection on IoT devices. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000877.

Rajkumar:2025:MVC

[RKA25] M. Rajkumar, J. Karthika, and S. S. Abinayaa. Multi-view consistent generative adversarial network for enhancing intrusion detection with prevention systems in mobile ad hoc networks against security attacks. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005480.

R:2023:HDL

[RKES23] Harine Rajashree R., Sundarakantham K., Sivasankar E., and Mercy Shalinie S. A hybrid deep learning framework for privacy preservation in edge computing. Computers & Security, 129 (??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001190.

Rostami:2020:RCT

[RKG20] Elham Rostami, Fredrik Karlsson, and Shang Gao. Requirements for computerized tools to design information security policies. Computers & Security, 99(??):Article 102063, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303369.

Rose:2025:MCT

[RKG^{+}25] Anthony J. Rose, Christine M. Schubert Kabban, Scott R. Graham, Wayne C. Henry, and Christopher M. Rondeau. Malware classification through Abstract Syntax Trees and Lmoments. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824003870.

Rashid:2022:ATD

[RKH^{+}22] Md. Mamunur Rashid, Joarder Kamruzzaman, Mohammad Mehedi Hassan, Tasadduq Imam, Santoso Wibowo, Steven Gordon, and Giancarlo Fortino. Adversarial training for deep learning-based cyberattack detection in IoTbased smart city applications. Computers & Security, 120(??): Article 102783, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482200178X.

Rostami:2025:TST

[RKKG25] Elham Rostami, Fredrik Karlsson, Ella Kolkowska, and Shang Gao. Towards software for tailoring information security policies to organisations’ different target groups. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003761.

Rani:2025:NGC

[RKKK25] Sita Rani, Aman Kataria, Sachin Kumar, and Vinod Karar. A new generation cyber-physical system: a comprehensive review from security perspective. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004000.

Reka:2024:MHS

[RKRS24] R. Reka, R. Karthick, R. Saravana Ram, and Gurkirpal Singh. Multi head self-attention gated graph convolutional network based multi-attack intrusion detection in MANET. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004364.

Radhika:2020:CAF

[RKSV20] B. S. Radhika, N. V. Narendra Kumar, R. K. Shyamasundar, and Parjanya Vyas. Consistency analysis and flow secure enforcement of SELinux policies. Computers & Security, 94(??):Article 101816, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300948.

Rass:2022:SML

[RKW^{+}22] Stefan Rass, Sandra König, Jasmin Wachter, Manuel Egger, and Manuel Hobisch. Supervised machine learning with plausible deniability. Computers & Security, 112(??):Article 102506, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003308.

Rass:2023:GTA

[RKW^{+}23] Stefan Rass, Sandra König, Jasmin Wachter, Víctor Mayoral-Vilches, and Emmanouil Panaousis. Game-theoretic APT defense: an experimental study on robotics. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002389.

Ren:2023:CGG

[RLL^{+}23] Yimo Ren, Hong Li, Peipei Liu, Jie Liu, Hongsong Zhu, and Limin Sun. CL-GAN: a GAN-based continual learning model for generating and detecting AGDs. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002274.

Rong:2020:TFT

[RLW^{+}20] Hong Rong, Jian Liu, Wei Wu, Jialu Hao, Huimei Wang, and Ming Xian. Toward fault-tolerant and secure frequent itemset mining outsourcing in hybrid cloud environment. Computers & Security, 98(??):Article 101969, November 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030242X.

Romdhana:2023:ASI

[RMCT23] Andrea Romdhana, Alessio Merlo, Mariano Ceccato, and Paolo Tonella. Assessing the security of inter-app communications in Android through reinforcement learning. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002213.

Rahartomo:2025:MSP

[RMG25] Argianto Rahartomo, Leonel Merino, and Mohammad Ghafari. Metaverse security and privacy research: a systematic review. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002913.

Rodriguez:2023:MTF

[RMMP23] Ricardo J. Rodríguez, Stefano Marrone, Ibai Marcos, and Giuseppe Porzio. MOSTO: a toolkit to facilitate security auditing of ICS devices using Modbus/TCP. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002833.

Rosado:2021:MBP

[RMS^{+}21] David G. Rosado, Julio Moreno, Luis E. Sánchez, Antonio Santos-Olmo, Manuel A. Serrano, and Eduardo Fernández-Medina. MARISMA-BiDa pattern: Integrated risk analysis for big data. Computers & Security, 102(??):Article 102155, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304284.

Rathgeb:2022:DFF

[RMS^{+}22] Christian Rathgeb, Johannes Merkle, Johanna Scholz, Benjamin Tams, and Vanessa Nesterowicz. Deep face fuzzy vault: Implementation and performance. Computers & Security, 113(??):Article 102539, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821003631.

Rendall:2025:MML

[RMVG25] Kieran Rendall, Alexios Mylonas, Stilianos Vidalis, and Dimitris Gritzalis. MIDAS: Multi-layered attack detection architecture with decision optimisation. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824004590.

Rasslan:2022:ISP

[RNA22] Mohamed Rasslan, Mahmoud M. Nasreldin, and Heba K. Aslan. Ibn Sina: a patient privacy-preserving authentication protocol in Medical Internet of Things. Computers & Security, 119(??):Article 102753, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001481.

Rust-Nguyen:2023:DTC

[RNSS23] Nhien Rust-Nguyen, Shruti Sharma, and Mark Stamp. Darknet traffic classification and adversarial attacks using machine learning. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000081.

Roumani:2022:DTD

[Rou22] Yaman Roumani. Detection time of data breaches. Computers & Security, 112(??):Article 102508, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003321.

Rajasoundaran:2021:MLB

[RPR^{+}21] S. Rajasoundaran, A. V. Prabu, Sidheswar Routray, S. V. N. Santhosh Kumar, Prince Priya Malla, Suman Maloji, Amrit Mukherjee, and Uttam Ghosh. Machine learning based deep job exploration and secure transactions in virtual private cloud systems. Computers & Security, 109(??):Article 102379, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002030.

Ruohonen:2025:CSI

[RRB25] Jukka Ruohonen, Kalle Rindell, and Simone Busetti. From cyber security incident management to cyber security crisis management in the European union. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404825003785.

Rachid:2020:EBO

[RRM20] Maan Haj Rachid, Ryan Riley, and Qutaibah Malluhi. Enclave-based oblivious RAM using Intel’s SGX. Computers & Security, 91(??):Article 101711, April 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302482.

Rodriguez-Ruiz:2020:OCC

[RRMSM^{+}20] Jorge Rodríguez-Ruiz, Javier Israel Mata-Sánchez, Raúl Monroy, Octavio Loyola-González, and Armando López-Cuevas. A one-class classification approach for bot detection on Twitter. Computers & Security, 91(??):Article 101715, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300031.

Roychoudhury:2021:SDD

[RRS21] Probidita Roychoudhury, Basav Roychoudhury, and Dilip K. Saikia. A secure device-to-device communication scheme for massive machine type communication. Computers & Security, 108(??):Article 102370, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001942.

Rashid:2023:SSD

[RS23] Aqib Rashid and Jose Such. StratDef: Strategic defense against adversarial attacks in ML-based malware detection. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003693.

Rajkumar:2025:SSD

[Rs25] K. Rajkumar and S. Mercy shalinie. Semi-supervised deep-ELM for DDoS attack detection and mitigation using the OptimalLink model in IoT networks. Computers & Security, 152 (??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000124.

Rendon-Segador:2023:PAC

[RSÁGVV23] Fernando J. Rendón-Segador, Juan A. Álvarez-García, and Angel Jesús Varela-Vaca. Paying attention to cyber-attacks: a multi-layer perceptron with self-attention mechanism. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002286.

Rocha:2020:FBM

[RSEK20] Thiago Rocha, Eduardo Souto, and Khalil El-Khatib. Functionality-based mobile application recommendation system with security and privacy awareness. Computers & Security, 97(??):Article 101972, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302455.

Rahimi:2025:CRS

[RSLM25] Narges Rahimi, Beth-Anne Schuelke-Leech, and Mitra Mirhassani. A comprehensive review of security vulnerabilities in heavy-duty vehicles: Comparative insights and current research gaps. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001415.

Racherache:2023:CIT

[RSSD23] Badis Racherache, Paria Shirani, Andrei Soeanu, and Mourad Debbabi. CPID: Insider threat detection using profiling and cyber-persona identification. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002602.

Ring:2021:MDW

[RSW^{+}21] Markus Ring, Daniel Schlör, Sarah Wunderlich, Dieter Landes, and Andreas Hotho. Malware detection on windows audit logs using LSTMs. Computers & Security, 109(??):Article 102389, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002133.

Rahman:2020:NBF

[RT20] Rizwan Ur Rahman and Deepak Singh Tomar. New biostatistics features for detecting web bot activity on web applications. Computers & Security, 97(??):Article 102001, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302741.

Rondeau:2021:ECI

[RTBK21] Christopher M. Rondeau, Michael A. Temple, J. Addison Betances, and Christine M. Schubert Kabban. Extending critical infrastructure element longevity using constellation-based ID verification. Computers & Security, 100(??):Article 102073, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303461.

Renaud:2023:WUC

[RvdSM23] Karen Renaud, Karl van der Schyff, and Stuart MacDonald. Would US citizens accept cybersecurity deresponsibilization? Perhaps not. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002110.

Radunovic:2025:EIS

[RVJ25] Vladimir Radunović, Mladen Veinović, and Aleksandar Jevremović. The efficiency of ICT suppliers’ product security incident response teams in reducing the risk of exploitation of vulnerabilities in the wild. Computers & Security, 152(??): ??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482500077X.

Rani:2024:AFH

[RVS24] B. Selva Rani, S. Vairamuthu, and Suresh Subramanian. Archimedes Fire Hawk Optimization enabled feature selection with deep maxout for network intrusion detection. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482400052X.

Rizvi:2024:ATM

[RW24] Syed Rizvi and Iyonna Williams. Analyzing transparency and malicious insiders prevention for cloud computing environment. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005321.

Ren:2021:NIS

[RXFZ21] Yanli Ren, Xiao Xu, Guorui Feng, and Xinpeng Zhang. Noninteractive and secure outsourcing of PCA-Based face recognition. Computers & Security, 110(??):Article 102416, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002406.

Rafi:2025:ODA

[RYS25] Shaik Mohammad Rafi, R. Yogesh, and M. Sriram. Optimized dual access control for cloud-based data storage and distribution using global-context residual recurrent neural network. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004887.

Rass:2020:CSS

[RZ20a] Stefan Rass and Quanyan Zhu. Computer and security special issue editorial. Computers & Security, 89(??):Article 101678, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819302184.

Ren:2020:DGM

[RZ20b] Jianguo Ren and Chunming Zhang. A differential game method against attacks in heterogeneous honeynet. Computers & Security, 97(??):Article 101870, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301425.

Roshan:2024:BBA

[RZ24] Khushnaseeb Roshan and Aasim Zafar. Black-box adversarial transferability: an empirical study in cybersecurity perspective. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001548.

Rizvi:2023:MFA

[RZT^{+}23] Syed Rizvi, Tatiana Zwerling, Benjamin Thompson, Shawn Faiola, Shakir Campbell, Stephen Fisanick, and Codi Hutnick. A modular framework for auditing IoT devices and networks. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002377.

Ren:2020:QEL

[RZW^{+}20] Yizhi Ren, Qi Zhou, Zhen Wang, Ting Wu, Guohua Wu, and Kim-Kwang Raymond Choo. Query-efficient label-only attacks against black-box machine learning models. Computers & Security, 90(??):Article 101698, March 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302354.

Ren:2024:SBA

[RZY^{+}24] Qixian Ren, Yu Zheng, Chao Yang, Yue Li, and Jianfeng Ma. Shadow backdoor attack: Multi-intensity backdoor attack against federated learning. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000415.

S:2020:ITP

[SA20] Ajish S. and K. S. AnilKumar. Iris template protection using double bloom filter based feature transformation. Computers & Security, 97(??):Article 101985, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302595.

Sharma:2022:OTC

[SA22] Shwadhin Sharma and Eduardo Aparicio. Organizational and team culture as antecedents of protection motivation among IT employees. Computers & Security, 120(??):Article 102774, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001699.

Skandylas:2025:APT

[SA25] Charilaos Skandylas and Mikael Asplund. Automated penetration testing: Formalization and realization. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001439.

Saidani:2020:SBC

[SAA20] Nadjate Saidani, Kamel Adi, and Mohand Saïd Allili. A semantic-based classification approach for an enhanced spam detection. Computers & Security, 94(??):Article 101716, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300043.

Seng:2021:FLU

[SAAW21] Sovantharith Seng, Mahdi Nasrullah Al-Ameen, and Matthew Wright. A first look into users’ perceptions of facial recognition in the physical world. Computers & Security, 105(??):Article 102227, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000511.

Shamim:2025:ISC

[SAB^{+}25] Nouman Shamim, Muhammad Asim, Thar Baker, Zeeshan Pervez, Ali Ismail Awad, and Albert Y. Zomaya. Integrating system calls and position-specific scoring for enhanced anomaly detection in Internet of Things environments. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003025.

Sharevski:2022:MWT

[SAJP22] Filipo Sharevski, Raniem Alsaadi, Peter Jachim, and Emma Pieroni. Misinformation warnings: Twitter’s soft moderation effects on COVID-19 vaccine belief echoes. Computers & Security, 114(??):Article 102577, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004016.

Shankar:2024:DMC

[SAK^{+}24] Deepa D. Shankar, Adresya Suresh Azhakath, Nesma Khalil, Sajeev J., Mahalakshmi T., and Sheeba K. Data mining for cyber biosecurity risk management — a comprehensive review. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005370.

Sturman:2025:SAD

[SAM25] Daniel Sturman, Jaime C. Auton, and Ben W. Morrison. Security awareness, decision style, knowledge, and phishing email detection: Moderated mediation analyses. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004346.

Sarier:2021:CBB

[Sar21] Neyire Deniz Sarier. Comments on biometric-based nontransferable credentials and their application in blockchainbased identity management. Computers & Security, 105 (??):Article 102243, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821000675.

Saha:2024:MES

[SAR24] Shoumik Saha, Sadia Afroz, and Atif Hasan Rahman. MAlign: Explainable static raw-byte based malware family classification using sequence alignment. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000154.

Singh:2023:CEL

[SARG23] Kuldeep Singh, Palvi Aggarwal, Prashanth Rajivan, and Cleotilde Gonzalez. Cognitive elements of learning and discriminability in anti-phishing training. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000159.

Saleh:2023:CRD

[SASW23] Mariam M. Saleh, Muhanned AlSlaiman, Mohammed I. Salman, and Bin Wang. Combining raw data and engineered features for optimizing encrypted and compressed Internet of Things traffic classification. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001979.

Sharif:2023:DAL

[SB23] Dyari Mohammed Sharif and Hakem Beitollahi. Detection of application-layer DDoS attacks using machine learning and genetic algorithms. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004212.

Sezgin:2025:DLL

[SB25] Anil Sezgin and Aytug Boyaci. DecoyPot: a large language model-driven web API honeypot for realistic attacker engagement. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001476.

Sebastio:2020:OSE

[SBB^{+}20] Stefano Sebastio, Eduard Baranov, Fabrizio Biondi, Olivier Decourbe, Thomas Given-Wilson, Axel Legay, Cassius Puodzius, and Jean Quilbeuf. Optimizing symbolic execution for malware behavior classification. Computers & Security, 93(??):Article 101775, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300602.

Sabaliauskaite:2024:TMT

[SBJ^{+}24] Giedre Sabaliauskaite, Jeremy Bryans, Hesamaldin Jadidbonab, Farhan Ahmad, Siraj Shaikh, and Paul Wooderson. TOMSAC — methodology for trade-off management between automotive safety and cyber security. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000993.

Spagnuelo:2020:QMT

[SBL20] Dayana Spagnuelo, Cesare Bartolini, and Gabriele Lenzini. Qualifying and measuring transparency: a medical data system case study. Computers & Security, 91(??):Article 101717, April 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740481830823X.

Snehi:2024:FSC

[SBV24] Manish Snehi, Abhinav Bhandari, and Jyoti Verma. Foggier skies, clearer clouds: a real-time IoT-DDoS attack mitigation framework in fog-assisted software-defined cyber-physical systems. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000038.

Sani:2022:SSP

[SBY^{+}22] Abubakar Sadiq Sani, Elisa Bertino, Dong Yuan, Ke Meng, and Zhao Yang Dong. SPrivAD: a secure and privacypreserving mutually dependent authentication and data access scheme for smart communities. Computers & Security, 115(??):Article 102610, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000098.

Singh:2021:SSH

[SC21] Ashish Singh and Kakali Chatterjee. Securing smart healthcare system with edge computing. Computers & Security, 108(??):Article 102353, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001772.

Sartayeva:2023:SIP

[SC23] Yerkezhan Sartayeva and Henry C. B. Chan. A survey on indoor positioning security and privacy. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002031.

Shaked:2023:OII

[SCBM23] Avi Shaked, Yulia Cherdantseva, Pete Burnap, and Peter Maynard. Operations-informed incident response playbooks. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003644.

Sanchez:2024:SBD

[SCBP24] Pedro Miguel Sánchez Sánchez, Alberto Huertas Celdrán, Gérôme Bovet, and Gregorio Martínez Pérez. Singleboard device individual authentication based on hardware performance and autoencoder transformer models. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005060.

Suo:2025:RSP

[SCC^{+}25] Yuhan Suo, Runqi Chai, Kaiyuan Chen, Senchun Chai, Wannian Liang, and Yuanqing Xia. Robust set partitioning strategy for malicious information detection in large-scale Internet of Things. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003773.

Sinigaglia:2020:SMF

[SCCZ20] Federico Sinigaglia, Roberto Carbone, Gabriele Costa, and Nicola Zannone. A survey on multi-factor authentication for online banking in the wild. Computers & Security, 95(??):Article 101745, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820300316.

Sarfaraz:2021:TSB

[SCE21] Aaliya Sarfaraz, Ripon K. Chakrabortty, and Daryl L. Essam. A tree structure-based improved blockchain framework for a secure online bidding system. Computers & Security, 102(??):Article 102147, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482030420X.

Shu:2025:TTS

[SCF^{+}25] Congyan Shu, Wentao Chen, Guisheng Fan, Huiqun Yu, Zijie Huang, and Yuguo Liang. Tool or toy: Are SCA tools ready for challenging scenarios? Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482500313X.

Sammoud:2020:NBB

[SCH^{+}20] Amal Sammoud, Mohamed Aymen Chalouf, Omessaad Hamdi, Nicolas Montavont, and Ammar Bouallegue. A new biometricsbased key establishment protocol in WBAN: energy efficiency and security robustness analysis. Computers & Security, 96(??):Article 101838, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301115.

Schiavone:2024:MSH

[Sch24] Antonio Giovanni Schiavone. Municipality2HTTPS: a study on HTTPS protocol’s usage in Italian municipalities’ websites. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005023.

Sun:2021:VVD

[SCL^{+}21] Hao Sun, Lei Cui, Lun Li, Zhenquan Ding, Zhiyu Hao, Jiancong Cui, and Peng Liu. VDSimilar: Vulnerability detection based on code similarity of vulnerabilities and patches. Computers & Security, 110(??):Article 102417, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002418.

Sun:2024:VVD

[SCL^{+}24] Hao Sun, Lei Cui, Lun Li, Zhenquan Ding, Siyuan Li, Zhiyu Hao, and Hongsong Zhu. VDTriplet: Vulnerability detection with graph semantics using triplet model. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000336.

Sanoussi:2023:IIT

[SCOE23] Nouhad Sanoussi, Kaouthar Chetioui, Ghizlane Orhanou, and Said El Hajji. ITC: Intrusion tolerant controller for multicontroller SDN architecture. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002614.

Shwartz:2020:ICH

[SCSO20] Omer Shwartz, Amir Cohen, Asaf Shabtai, and Yossi Oren. Inner conflict: How smart device components can cause harm. Computers & Security, 89(??):Article 101665, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404819302056.

Stylios:2023:CAF

[SCTK23] Ioannis Stylios, Sotirios Chatzis, Olga Thanou, and Spyros Kokolakis. Continuous authentication with feature-level fusion of touch gestures and keystroke dynamics to solve security and usability issues. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002730.

Svabensky:2021:CKS

[SCVB21] Valdemar Svábenský, Pavel Celeda, Jan Vykopal, and Silvia Brisáková. Cybersecurity knowledge and skills taught in capture the flag challenges. Computers & Security, 102 (??):Article 102154, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820304272.

Shen:2022:FFB

[SCW^{+}22] Gaoning Shen, Zhixiang Chen, Hui Wang, Heng Chen, and Shuqi Wang. Feature fusion-based malicious code detection with dual attention mechanism and BiLSTM. Computers & Security, 119(??):Article 102761, August 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001560.

Soylu:2025:PGV

[SD25] Mucahit Soylu and Resul Das. Prediction and graph visualization of cyber attacks using graph attention networks. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002238.

S:2024:AIS

[SDA^{+}24a] Senthil Pandi S, Roja Ramani D, Senthilselvi A, Dhanasekaran S, Kalpana B, and Alangudi Balaji N. Advancing IoT security with flame: a hybrid approach combining fuzzy logic and artificial lizard search optimization. Computers & Security, 145(??): ??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400289X.

Sousa:2024:MTM

[SDA^{+}24b] Bruno Sousa, Duarte Dias, Nuno Antunes, Javier Cámara, Ryan Wagner, Bradley Schmerl, David Garlan, and Pedro Fidalgo. MONDEO-Tactics5G: Multistage botnet detection and tactics for 5G/6G networks. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000695.

Shahriar:2025:CRC

[SDA25] Sakib Shahriar, Rozita Dara, and Rajen Akalu. A comprehensive review of current trends, challenges, and opportunities in text data privacy. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000471.

Shao:2025:GPG

[SDCL25] Miaomiao Shao, Yuxin Ding, Jing Cao, and Yilin Li. Graph-FVD: Property graph-based fine-grained vulnerability detection. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000392.

Stergiopoulos:2020:ANR

[SDG20] George Stergiopoulos, Panagiotis Dedousis, and Dimitris Gritzalis. Automatic network restructuring and risk mitigation through business process asset dependency analysis. Computers & Security, 96(??):Article 101869, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301413.

Shillair:2022:CEA

[SEGD^{+}22] Ruth Shillair, Patricia Esteve-González, William H. Dutton, Sadie Creese, Eva Nagyfejeo, and Basie von Solms. Cybersecurity education, awareness raising, and training initiatives: National level evidence-based results, challenges, and promise. Computers & Security, 119(??):Article 102756, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001511.

Sahay:2023:CRA

[SEM^{+}23] Rishikesh Sahay, D. A. Sepulveda Estay, Weizhi Meng, Christian D. Jensen, and Michael Bruhn Barfod. A comparative risk analysis on CyberShip system with STPA-Sec, STRIDE and CORAS. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000895.

Shen:2024:GGH

[SFX24] Lina Shen, Mengqi Fang, and Jian Xu. GHGDroid: Global heterogeneous graph-based Android malware detection. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001470.

Subba:2021:TSV

[SG21] Basant Subba and Prakriti Gupta. A tfidfvectorizer and singular value decomposition based host intrusion detection system framework for detecting anomalous system processes. Computers & Security, 100(??):Article 102084, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303576.

Subrahmanyam:2023:HMA

[SGA^{+}23] Singam Sai Bala Subrahmanyam, P. Goutham, Vasanth Kumar Reddy Ambati, C. V. Bijitha, and Hiran V. Nath. A hybrid method for analysis and detection of malicious executables in IoT network. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002493.

Studiawan:2023:UAV

[SGC23] Hudan Studiawan, George Grispos, and Kim-Kwang Raymond Choo. Unmanned Aerial Vehicle (UAV) forensics: the good, the bad, and the unaddressed. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300250X.

Shoukat:2025:TMI

[SGJ^{+}25] Shifa Shoukat, Tianhan Gao, Danish Javeed, Muhammad Shahid Saeed, and Muhammad Adil. Trust my IDS: an explainable AI integrated deep learning-based transparent threat detection system for industrial networks. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004966.

Shen:2022:PBD

[SGS^{+}22] Yuncheng Shen, Bing Guo, Yan Shen, Xuliang Duan, Xiangqian Dong, Hong Zhang, Chuanwu Zhang, and Yuming Jiang. Personal big data pricing method based on differential privacy. Computers & Security, 113(??):Article 102529, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003539.

Sanchez-Garcia:2023:CTT

[SGSCM23] Isaac D. Sánchez-García, Tomás San Feliu Gilabert, and Jose A. Calvo-Manzano. Countermeasures and their taxonomies for risk treatment in cybersecurity: a systematic mapping review. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000809.

Sharma:2022:OAM

[SGSS22] Amit Sharma, Brij B. Gupta, Awadhesh Kumar Singh, and V. K. Saraswat. Orchestration of APT malware evasive manoeuvers employed for eluding anti-virus and sandbox defense. Computers & Security, 115(??):Article 102627, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000268.

Shafei:2024:MPP

[SGT24] Hassan A. Shafei, Hongchang Gao, and Chiu C. Tan. Measuring privacy policy compliance in the Alexa ecosystem: Indepth analysis. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002682.

Stergiopoulos:2021:DMT

[SGVA21] George Stergiopoulos, Dimitris Gritzalis, Efstratios Vasilellis, and Argiro Anagnostopoulou. Dropping malware through sound injection: a comparative analysis on Android operating systems. Computers & Security, 105(??):Article 102228, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000523.

Song:2023:PPR

[SGY23] Yunjiao Song, Xinrui Ge, and Jia Yu. Privacy-preserving reachability query over graphs with result verifiability. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000020.

Stewart:2025:SAS

[SH25] Andrew Stewart and Christopher Hobbs. Systematic analysis of security advice on the topic of insider threats. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001002.

Shahim:2021:SDT

[Sha21] Abbas Shahim. Security of the digital transformation. Computers & Security, 108(??):Article 102345, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001693.

Sun:2024:MDA

[SHH^{+}24] Haili Sun, Yan Huang, Lansheng Han, Cai Fu, Hongle Liu, and Xiang Long. MTS-DVGAN: Anomaly detection in cyberphysical systems using a dual variational generative adversarial network. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004807.

Slijepcevic:2021:APH

[SHK^{+}21] Djordje Slijepčević, Maximilian Henzl, Lukas Daniel Klausner, Tobias Dam, Peter Kieseberg, and Matthias Zeppelzauer. k-anonymity in practice: How generalisation and suppression affect machine learning classifiers. Computers & Security, 111(??):Article 102488, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003126.

Shi:2020:ABE

[SHL^{+}20] Shuyun Shi, Debiao He, Li Li, Neeraj Kumar, Muhammad Khurram Khan, and Kim-Kwang Raymond Choo. Applications of blockchain in ensuring the security and privacy of electronic health record systems: a survey. Computers & Security, 97(??):Article 101966, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030239X.

nmez:2022:ADA

[SHM22a] Ferda Özdemir Sönmez, Chris Hankin, and Pasquale Malacaria. Attack dynamics: an automatic attack graph generation framework based on system topology, CAPEC, CWE, and CVE databases. Computers & Security, 123(??):Article 102938, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003303.

Sonmez:2022:DSH

[SHM22b] Ferda Özdemir Sönmez, Chris Hankin, and Pasquale Malacaria. Decision support for healthcare cyber security. Computers & Security, 122(??):Article 102865, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002590.

Stojanovic:2020:ADA

[SHSK20] Branka Stojanović, Katharina Hofer-Schmitz, and Ulrike Kleb. APT datasets and attack modeling for automated detection methods: a review. Computers & Security, 92(??):Article 101734, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300213.

Singh:2025:IIT

[Sin25] Kuldeep Singh. Industrial Internet of Things fortify: multidomain feature learning framework with deepdetectnet++ for improved intrusion detection. Computers & Security, 156 (??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825001956.

Sultan:2021:MBN

[SJ21] Shizra Sultan and Christian D. Jensen. Metadata based needto-know view in large-scale video surveillance systems. Computers & Security, 111(??):Article 102452, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002765.

Smith:2021:AIR

[SJH^{+}21] Richard Smith, Helge Janicke, Ying He, Fenia Ferra, and Adham Albakri. The Agile Incident Response for Industrial Control Systems (AIR4ICS) framework. Computers & Security, 109(??):Article 102398, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002224.

Shaji:2023:DDA

[SJMP23] Neena Susan Shaji, Tanushree Jain, Raja Muthalagu, and Pranav Mothabhau Pawar. Deep-discovery: Anomaly discovery in software-defined networks using artificial neural networks. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002304.

Shao:2020:ADT

[SJX^{+}20] Wei Shao, Chunfu Jia, Yunkai Xu, Kefan Qiu, Yan Gao, and Yituo He. AttriChain: Decentralized traceable anonymous identities in privacy-preserving permissioned blockchain. Computers & Security, 99(??):Article 102069, December 2020.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303424.

S:2020:SPE

[SK20] Ajish S. and K. S. Anil Kumar. Security and performance enhancement of fingerprint biometric template using symmetric hashing. Computers & Security, 90(??):Article 101714, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482030002X.

Sheikhi:2024:SCE

[SK24a] Saeid Sheikhi and Panos Kostakos. Safeguarding cyberspace: Enhancing malicious website detection with PSO optimized XGBoost and firefly-based feature selection. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400186X.

Suman:2024:ONN

[SK24b] Suman and Raees Ahmad Khan. An optimized neural network for prediction of security threats on software testing. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005369.

Sarikaya:2023:RRA

[SKD23] Alper Sarikaya, Banu Günel Kiliç, and Mehmet Demirci. RAIDS: Robust autoencoder-based intrusion detection system model against adversarial attacks. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823003930.

Skeoch:2022:EGL

[Ske22] Henry R. K. Skeoch. Expanding the Gordon–Loeb model to cyber-insurance. Computers & Security, 112(??):Article 102533, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821003576.

See corrigendum [Ske24].

Skeoch:2024:CSE

[Ske24] Henry R. K. Skeoch. Corrigendum to “Expanding the Gordon– Loeb model to cyber-insurance” [Computers and Security Volume 112 (2022) 102533]. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005667. See [Ske22].

Stellios:2021:AIE

[SKG21] Ioannis Stellios, Panayiotis Kotzanikolaou, and Christos Grigoriadis. Assessing IoT enabled cyber-physical attack paths against critical systems. Computers & Security, 107(??):Article 102316, August 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001401.

Smiliotopoulos:2025:ADL

[SKKG25] Christos Smiliotopoulos, Georgios Kambourakis, Constantinos Kolias, and Stefanos Gritzalis. Assessing the detection of lateral movement through unsupervised learning techniques. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004954.

Schuckert:2023:IRA

[SKL23] Felix Schuckert, Basel Katt, and Hanno Langweg. Insecurity refactoring: Automated injection of vulnerabilities in source code. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000317.

Shlomo:2021:TPB

[SKM21] Amit Shlomo, Meir Kalech, and Robert Moskovitch. Temporal pattern-based malicious activity detection in SCADA systems. Computers & Security, 102(??):Article 102153, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304260.

Sudhakar:2021:AEF

[SKS21] Sai Ram Vallam Sudhakar, Namrata Kayastha, and Kewei Sha. ActID: an efficient framework for activity sensor based user identification. Computers & Security, 108(??):Article 102319, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001437.

Syed:2025:ESV

[SKS25] Romilla Syed, Joti Kaur, and Leiser Silva. Exposure of software vulnerabilities on Twitter: Analyzing vendors’ behavior of releasing software patches. Computers & Security, 151(??): ??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000434.

Shin:2020:RTE

[SL20] Bongsik Shin and Paul Benjamin Lowry. A review and theoretical explanation of the ‘cyberthreat-intelligence (CTI) capability’ that needs to be fostered in information security practitioners and how this can be accomplished. Computers & Security, 92(??):Article 101761, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300456.

Solanki:2021:MVA

[SLBG21] Rajendra Kumar Solanki, Vijay Laxmi, Bruhadeshwar Bezawada, and Manoj Singh Gaur. MapperDroid: Verifying app capabilities from description to permissions and API calls. Computers & Security, 111(??):Article 102493, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003175.

Shao:2021:MEF

[SLC^{+}21] Minglai Shao, Jianxin Li, Yue Chang, Jun Zhao, and Xunxun Chen. MASA: an efficient framework for anomaly detection in multi-attributed networks. Computers & Security, 102(??):Article 102085, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303588.

She:2024:SEF

[SLCS24] Cairui She, Jinfeng Li, Liwei Chen, and Gang Shi. SCFI: Efficient forward fine-grained control flow integrity based on coarse-grained ISA extensions. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001019.

Sun:2023:ICG

[SLDL23] Xuekai Sun, Chunling Liu, Weiyu Dong, and Tieming Liu. Improvements to code2vec: Generating path vectors using RNN. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002328.

Sun:2023:NCF

[SLL23] Yanan Sun, Hengjian Li, and Nianqiang Li. A novel cancelable fingerprint scheme based on random security sampling mechanism and relocation bloom filter. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004138.

Shafi:2025:NTG

[SLR25] MohammadMoein Shafi, Arash Habibi Lashkari, and Arousha Haghighian Roudsari. NTLFlowLyzer: Towards generating an intrusion detection dataset and intruders behavior profiling through network and transport layers traffic analysis and pattern extraction. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004656.

Sun:2023:GGP

[SLX^{+}23] Peishuai Sun, Shuhao Li, Jiang Xie, Hongbo Xu, Zhenyu Cheng, and Rong Yang. GPMT: Generating practical malicious traffic based on adversarial attacks with little prior knowledge. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001670.

Shi:2025:MDA

[SLZ^{+}25] Guang Shi, Tao Li, Qinghe Zhou, Menglin Liu, and Yuan Feng. Multi-dimensional assessment for Android application security based on users’ evaluation. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001877.

Shi:2024:RQA

[SLZP24] Ze Shi, Hongyi Li, Di Zhao, and Chengwei Pan. Research on quality assessment methods for cybersecurity knowledge graphs. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001494.

Samarasinghe:2021:CBM

[SM21] Nayanamana Samarasinghe and Mohammad Mannan. On cloaking behaviors of malicious websites. Computers & Security, 101(??):Article 102114, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303874.

Saqib:2023:SSA

[SM23] Manasha Saqib and Ayaz Hassan Moon. A systematic security assessment and review of Internet of Things in the context of authentication. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200445X.

Sharma:2024:RCS

[SM24] Anee Sharma and Ningrinla Marchang. A review on client-server attacks and defenses in federated learning. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001020.

Smaga:2025:PVC

[Sma25] Paweł Smaga. Profiling the victim — cyber risk in commercial banks. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005807.

Shen:2021:AHS

[SMC^{+}21] Lixiang Shen, Dejun Mu, Guo Cao, Maoyuan Qin, Jiacheng Zhu, and Wei Hu. Accelerating hardware security verification and vulnerability detection through state space reduction. Computers & Security, 103(??):Article 102167, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304405.

Sanchez:2021:APP

[SMCP21] Pedro Miguel Sánchez Sánchez, Lorenzo Fernández Maimó, Alberto Huertas Celdrán, and Gregorio Martínez Pérez. Auth-CODE: a privacy-preserving and multi-device continuous authentication architecture based on machine and deep learning. Computers & Security, 103(??):Article 102168, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820304417.

Singh:2022:SML

[SMM22] Neeraj Kumar Singh, Mahshooq Abdul Majeed, and Vasundhara Mahajan. Statistical machine learning defensive mechanism against cyber intrusion in smart grid cyber-physical network. Computers & Security, 123(??):Article 102941, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003339.

Shepherd:2021:PFI

[SMvH^{+}21] Carlton Shepherd, Konstantinos Markantonakis, Nico van Heijningen, Driss Aboulkassimi, Clément Gaine, Thibaut Heckmann, and David Naccache. Physical fault injection and sidechannel attacks on mobile devices: a comprehensive analysis. Computers & Security, 111(??):Article 102471, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002959.

Shenderovitz:2024:BAD

[SN24] Gil Shenderovitz and Nir Nissim. Bon-APT: Detection, attribution, and explainability of APT malware using temporal segmentation of API calls. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001639.

Shvartzman:2023:CDC

[SOZ^{+}23a] Oren Shvartzman, Adar Ovadya, Kfir Zvi, Omer Shwartz, Rom Ogen, Yakov Mallah, Niv Gilboa, and Yossi Oren. Characterization and detection of cross-router covert channels. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000354.

Sun:2023:IMI

[SOZ^{+}23b] Hongyu Sun, Guoliang Ou, Ziqiu Zheng, Lei Liao, He Wang, and Yuqing Zhang. Inconsistent measurement and incorrect detection of software names in security vulnerability reports. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003875.

Schmitz:2020:LLS

[SP20] Christopher Schmitz and Sebastian Pape. LiSRA: Lightweight Security Risk Assessment for decision support in information security. Computers & Security, 90(??):Article 101656, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819301993.

Souza:2025:NOS

[SPG^{+}25] Manuela M. C. Souza, Camila T. Pontes, João J. C. Gondim, Luís P. F. Garcia, Luiz DaSilva, Eduardo F. M. Cavalcante, and Marcelo A. Marotta. A novel open set Energybased Flow Classifier for Network Intrusion Detection. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002585.

Suchorab:2025:ADS

[SPL25] Jakub Suchorab, Sebastian Plamowski, and Maciej Lawryńczuk. Anomaly detection system for Modbus data based on an open source tool. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002615.

Sekhar:2025:SGB

[SPN^{+}25] J C Sekhar, R Priyanka, Ashok Kumar Nanda, P Joel Josephson, M J D Ebinezer, and T Kalavathi Devi. Stochastic gradient boosted distributed decision trees security approach for detecting cyber anomalies and classifying multiclass cyberattacks. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000094.

Spring:2023:AHM

[Spr23] Jonathan M. Spring. An analysis of how many undiscovered vulnerabilities remain in information systems. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823001013.

Sundaramoorthy:2024:ECC

[SPSK24] K. Sundaramoorthy, K. E. Purushothaman, J. Jeba Sonia, and N. Kanthimathi. Enhancing cybersecurity in cloud computing and WSNs: a hybrid IDS approach. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003869.

Siponen:2020:CIN

[SPV20] Mikko Siponen, Petri Puhakainen, and Anthony Vance. Can individuals’ neutralization techniques be overcome? A field experiment on password policy. Computers & Security, 88(??):Article 101617, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404819301646.

Singh:2020:BSC

[SPZ^{+}20] Amritraj Singh, Reza M. Parizi, Qi Zhang, Kim-Kwang Raymond Choo, and Ali Dehghantanha. Blockchain smart contracts formalization: Approaches and challenges to address vulnerabilities. Computers & Security, 88(??):Article 101654, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404818310927.

Song:2024:DDG

[SQL^{+}24a] Jiaru Song, Guihe Qin, Yanhua Liang, Jie Yan, and Minghui Sun. DGIDS: Dynamic graph-based intrusion detection system for CAN. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400381X.

Song:2024:SSB

[SQL^{+}24b] Jiaru Song, Guihe Qin, Yanhua Liang, Jie Yan, and Minghui Sun. SIDiLDNG: a similarity-based intrusion detection system using improved Levenshtein Distance and N-gram for CAN. Computers & Security, 142(??):??, July 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001482.

Sicari:2022:ISP

[SRCP22] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. Insights into security and privacy towards fog computing evolution. Computers & Security, 120(??):Article 102822, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002164.

Shibbir:2024:ESC

[SRFC24] Md. Neyamul Islam Shibbir, Hasibur Rahman, Md Sadek Ferdous, and Farida Chowdhury. Evaluating the security of CAPTCHAs utilized on Bangladeshi websites. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824000750.

Sung:2022:DKL

[SRM22] Dillon Cheong Lien Sung, Gauthama Raman M. R., and Aditya P. Mathur. Design-knowledge in learning plant dynamics for detecting process anomalies in water treatment plants. Computers & Security, 113(??):Article 102532, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003564.

S:2024:MMB

[SRN24] Rohini S., Gowtham Ramesh, and Anand R. Nair. MAGIC: Malware behaviour analysis and impact quantification through signature co-occurrence and regression. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000361.

Suarez-Roman:2025:ASM

[SRT25] Manuel Suarez-Roman and Juan Tapiador. Attack structure matters: Causality-preserving metrics for Provenance-based Intrusion Detection Systems. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002676.

Senanayake:2025:MBB

[SRY^{+}25] Janaka Senanayake, Sampath Rajapaksha, Naoto Yanai, Harsha Kalutarage, and Chika Komiya. MADONNA: Browser-based malicious domain detection using Optimized Neural Network by leveraging AI and feature analysis. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000604.

Sadique:2022:MAA

[SS22] Farhan Sadique and Shamik Sengupta. Modeling and analyzing attacker behavior in IoT botnet using temporal convolution network (TCN). Computers & Security, 117(??):Article 102714, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001092.

Shaikh:2023:ISR

[SS23a] Faheem Ahmed Shaikh and Mikko Siponen. Information security risk assessments following cybersecurity breaches: the mediating role of top management attention to cybersecurity. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003662.

Srijayanthi:2023:DPP

[SS23b] S. Srijayanthi and T. Sethukarasi. Design of privacy preserving model based on clustering involved anonymization along with feature selection. Computers & Security, 126(??):??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004199.

Sadhya:2024:CSS

[SS24a] Debanjan Sadhya and Tanya Sahu. A critical survey of the security and privacy aspects of the Aadhaar framework. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400083X.

Shen:2024:ESR

[SS24b] Quan Shen and Yanming Shen. Endpoint security reinforcement via integrated zero-trust systems: a collaborative approach. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004479.

Srivastava:2025:AIA

[SS25] Arpita Srivastava and Ditipriya Sinha. ARLHNIDS-IoT: an accurate and robust lightweight hybrid-NIDS for IoT network security. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002044.

Sharma:2023:SBC

[SSC23] Shreyansh Sharma, Anil Saini, and Santanu Chaudhury. A survey on biometric cryptosystems and their applications. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003681.

Sohi:2021:REN

[SSG21] Soroush M. Sohi, Jean-Pierre Seifert, and Fatemeh Ganji. RN-NIDS: Enhancing network intrusion detection systems through deep learning. Computers & Security, 102(??):Article 102151, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304247.

Sharma:2025:PDP

[SSG^{+}25a] Saurabh Sharma, Jaiteg Singh, Ankur Gupta, Farman Ali, and Sukhjit Singh Sehra. PRIVIUM: a differentiated privacyprivilege model for user security and safety in the metaverse. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003475.

Sinan:2025:IIS

[SSG25b] Maysa Sinan, Mojtaba Shahin, and Iqbal Gondal. Implementing and integrating security controls: a practitioners’ perspective. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002056.

Soriano-Salvador:2021:SSB

[SSGM21] Enrique Soriano-Salvador and Gorka Guardiola-Múzquiz. SealFS: Storage-based tamper-evident logging. Computers & Security, 108(??):Article 102325, September 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001498.

Schmitz:2021:MLA

[SSHP21] Christopher Schmitz, Michael Schmid, David Harborth, and Sebastian Pape. Maturity level assessments of information security controls: an empirical analysis of practitioners assessment capabilities. Computers & Security, 108(??): Article 102306, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821001309.

Schinagl:2022:PTI

[SSK22] Stef Schinagl, Abbas Shahim, and Svetlana Khapova. Paradoxical tensions in the implementation of digital security governance: Toward an ambidextrous approach to governing digital security. Computers & Security, 122(??):Article 102903, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002966.

S:2023:CSP

[SSK23a] Devi Priya V. S., Sibi Chakkaravarthy Sethuraman, and Muhammad Khurram Khan. Container security: Precaution levels, mitigation strategies, and research perspectives. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004005.

Srivastava:2023:WGB

[SSK23b] Arpita Srivastava, Ditipriya Sinha, and Vikash Kumar. WCGAN-GP based synthetic attack data generation with GA based feature selection for IDS. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823003425.

Sun:2024:MSM

[SSK^{+}24] Huaqi Sun, Hui Shu, Fei Kang, Yuntian Zhao, and Yuyao Huang. Malware2ATT&CK: a sophisticated model for mapping malware to ATT&CK techniques. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000737.

Shao:2020:SWF

[SSL20] Xiuyan Shao, Mikko Siponen, and Fufan Liu. Shall we follow? Impact of reputation concern on information security managers’ investment decisions. Computers & Security, 97(??):Article 101961, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302376.

Salvi:2022:CRC

[SSN22] Andrea Salvi, Paolo Spagnoletti, and Nadia Saad Noori. Cyber-resilience of critical cyber infrastructures: Integrating digital twins in the electric power ecosystem. Computers & Security, 112(??):Article 102507, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100331X.

Sharma:2023:HSS

[SSN23] Abhishek Sharma, Sangeeta Sabharwal, and Sushama Nagpal. A hybrid scoring system for prioritization of software vulnerabilities. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001669.

Sethuraman:2024:CEE

[SSR^{+}24] Sibi Chakkaravarthy Sethuraman, Devi Priya V. S., Tarun Reddi, Mulka Sai Tharun Reddy, and Muhammad Khurram Khan. A comprehensive examination of email spoofing: Issues and prospects for email security. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005102.

Szczepaniuk:2020:ISA

[SSRK20] Edyta Karolina Szczepaniuk, Hubert Szczepaniuk, Tomasz Rokicki, and Bogdan Klepacki. Information security assessment in public administration. Computers & Security, 90 (??):Article 101709, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302469.

Sha:2025:CFA

[SSRM25] Zhanyu Sha, Carlton Shepherd, Amir Rafi, and Konstantinos Markantonakis. Control-flow attestation: Concepts, solutions, and open challenges. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005601.

Shah:2021:LLC

[SSS^{+}21] Syed W. Shah, Naeem F. Syed, Arash Shaghaghi, Adnan Anwar, Zubair Baig, and Robin Doss. LCDA: Lightweight continuous device-to-device authentication for a Zero Trust Architecture (ZTA). Computers & Security, 108(??):Article 102351, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001759.

Soomro:2021:FIF

[SST21] Zahoor Ahmed Soomro, Mahmood Hussain Shah, and Jason Thatcher. A framework for ID fraud prevention policies in etailing sector. Computers & Security, 109(??):Article 102403, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002273.

Syed:2022:TSC

[SSTRD22] Naeem Firdous Syed, Syed W. Shah, Rolando Trujillo-Rasua, and Robin Doss. Traceability in supply chains: a cyber security analysis. Computers & Security, 112(??):Article 102536, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003606.

Song:2024:AAP

[SSZ^{+}24] Haina Song, Hua Shen, Nan Zhao, Zhangqing He, Minghu Wu, Wei Xiong, and Mingwu Zhang. APLDP: Adaptive personalized local differential privacy data collection in mobile crowdsensing. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004273.

Singh:2024:TLI

[ST24] Narendra Singh and Somanath Tripathy. It’s too late if exfiltrate: Early stage Android ransomware detection. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001202.

Singh:2025:UVE

[ST25a] Narendra Singh and Somanath Tripathy. Unveiling the veiled: an early stage detection of fileless malware. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005376.

Sutton:2025:TCC

[ST25b] Anna Sutton and Lisa Tompson. Towards a cybersecurity culture-behaviour framework: a rapid evidence review. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004152.

Shafiq:2020:IMT

[STB^{+}20] Muhammad Shafiq, Zhihong Tian, Ali Kashif Bashir, Xiaojiang Du, and Mohsen Guizani. IoT malicious traffic identification using wrapper-based feature selection mechanisms. Computers & Security, 94(??):Article 101863, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301358.

S:2025:OIP

[STB25] Anagha A. S., Ciza Thomas, and N. Balakrishnan. Optimized intrusion predictions through feature selection methods. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002305.

Sun:2022:PPV

[STH^{+}22] Xin Sun, Chengliang Tian, Changhui Hu, Weizhong Tian, Hanlin Zhang, and Jia Yu. Privacy-preserving and verifiable SRC-based face recognition with cloud/edge server assistance. Computers & Security, 118(??):Article 102740, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001353.

Suaboot:2020:SCH

[STM^{+}20] Jakapan Suaboot, Zahir Tari, Abdun Mahmood, Albert Y. Zomaya, and Wei Li. Sub-curve HMM: a malware detection approach based on partial analysis of API call sequences. Computers & Security, 92(??):Article 101773, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300596.

Sharma:2025:MOR

[STPB25] Yogesh Kumar Sharma, Deepak Singh Tomar, R. K. Pateriya, and Shweta Bhandari. MOSDroid: Obfuscation-resilient Android malware detection using multisets of encoded opcode sequences. Computers & Security, 152(??):??, May 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000689.

Swain:2025:ICS

[STS25] Munmun Swain, Nikhil Tripathi, and Kamalakanta Sethi. Identifying communication sequence anomalies to detect DoS attacks against MQTT. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002159.

Siponen:2025:RNT

[STSV25] Mikko Siponen, Volkan Topalli, Wael Soliman, and Tiina Vestman. Reconsidering neutralization techniques in behavioral cybersecurity as cybersecurity hygiene discounting. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006126.

Showail:2022:ISP

[STZ^{+}22] Ahmad Showail, Rashid Tahir, Muhammad Fareed Zaffar, Muhammad Haris Noor, and Mohammed Al-Khatib. An internet of secure and private things: a service-oriented architecture. Computers & Security, 120(??):Article 102776, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001717.

Saravanan:2021:LBA

[SU21] N. Saravanan and A. Umamakeswari. Lattice based access control for protecting user data in cloud environments with hybrid security. Computers & Security, 100(??):Article 102074, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303473.

Sahin:2025:LIA

[SU25] Onur Sahin and Suleyman Uludag. Leveraging inter-arrival time for efficient threat filtering: a parsimonious approach. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001609.

Soremekun:2023:TBA

[SUC23] Ezekiel Soremekun, Sakshi Udeshi, and Sudipta Chattopadhyay. Towards backdoor attacks and defense in robust machine learning models. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000111.

Sun:2021:TGM

[Sun21] Pan Jun Sun. A tripartite game model of trust cooperation in cloud service. Computers & Security, 106(??):Article 102272, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000961.

Samiah:2025:DTB

[SUS25] Abdul Samiah, Muhammad Azmi Umer, and Shama Siddiqui. Decision tree based invariants for intrusion detection in industrial control system. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002007.

Sahin:2025:WDW

[SV25] Zeynep Sahin and Anthony Vance. What do we need to know about the Chief Information Security Officer? A literature review and research agenda. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824003687.

Shahani:2021:CBR

[SVA21] Snehkumar Shahani, R. Venkateswaran, and Jibi Abraham. Cost-based recommendation of parameters for local differentially private data aggregation. Computers & Security, 102(??):Article 102144, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482030417X.

Schlette:2021:PCS

[SVP21] Daniel Schlette, Manfred Vielberth, and Günther Pernul. CTI-SOC2M2 — the quest for mature, intelligence-driven security operations and incident response capabilities. Computers & Security, 111(??):Article 102482, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003060.

Simoglou:2021:IDS

[SVPM21] George Simoglou, George Violettas, Sophia Petridou, and Lefteris Mamatas. Intrusion detection systems for RPL security: a comparative analysis. Computers & Security, 104(??):Article 102219, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000432.

Strous:2021:SPI

[SvSZ21] Leon Strous, Suné von Solms, and André Zúquete. Security and privacy of the Internet of Things. Computers & Security, 102(??):Article 102148, March 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820304211.

Sion:2025:RRL

[SVWJ25] Laurens Sion, Dimitri Van Landuyt, Kim Wuyts, and Wouter Joosen. Robust and reusable LINDDUN privacy threat knowledge. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001087.

Shin:2022:PAE

[SW22] Youjin Shin and Simon S. Woo. PasswordTensor: Analyzing and explaining password strength using tensor decomposition. Computers & Security, 116(??):Article 102634, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000335.

Sun:2020:UIB

[SWK20] Qizhang Sun, Martijn C. Willemsen, and Bart P. Knijnenburg. Unpacking the intention-behavior gap in privacy decision making for the Internet of Things (IoT) using aspect listing. Computers & Security, 97(??):Article 101924, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302005.

Shang:2024:APM

[SWL24] Siyuan Shang, Xiaohan Wang, and Aodi Liu. ABAC policy mining method based on hierarchical clustering and relationship extraction. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400018X.

Sun:2025:HFS

[SWL^{+}25] Nan Sun, Wei Wang, Kexin Liu, Donghong Li, and Jinhu Lü. Hybrid framework for security evaluation in Internet of Vehicles. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000872.

Song:2025:MSL

[SWS^{+}25] Yubo Song, Kanghui Wang, Xin Sun, Zhongyuan Qin, Hua Dai, Weiwei Chen, Bang Lv, and Jiaqi Chen. A multisource log semantic analysis-based attack investigation approach. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006096.

Sun:2025:SPS

[SWW^{+}25] Panjun Sun, Yi Wan, Zongda Wu, Zhaoxi Fang, and Qi Li. A survey on privacy and security issues in IoT-based environments: Technologies, protection measures and future directions. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004024.

Sun:2020:DMD

[SWYL20] Xiaoqing Sun, Zhiliang Wang, Jiahai Yang, and Xinran Liu. Deepdom: Malicious domain detection with scalable and heterogeneous graph convolutional networks. Computers & Security, 99(??):Article 102057, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303308.

Sathiya:2024:SOD

[SY24] R. Sathiya and Dr. N. Yuvaraj. Swarm optimized differential evolution and probabilistic extreme learning based intrusion detection in MANET. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400275X.

Shao:2021:BED

[SYA^{+}21] Kun Shao, Junan Yang, Yang Ai, Hui Liu, and Yu Zhang. BDDR: an effective defense against textual backdoor attacks. Computers & Security, 110(??):Article 102433, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821002571.

Sun:2024:HIF

[SYC^{+}24] Haodong Sun, Zhi Yang, Xingyuan Chen, Hang Xu, and Zhanhui Yuan. Hardware information flow tracking based on lightweight path awareness. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003778.

Srinivas:2025:EBT

[SYD^{+}25] P. V. V. S. Srinivas, Nikhil Yadavalli, Venkata Durga, Karthik Kumar, and Prateesh Raju. Enhanced biometric template protection schemes using distance based fuzzy extractor. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002627.

Sheng:2021:SIH

[SYF^{+}21] Chuan Sheng, Yu Yao, Qiang Fu, Wei Yang, and Ying Liu. Study on the intelligent honeynet model for containing the spread of industrial viruses. Computers & Security, 111(??):Article 102460, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002844.

Sun:2021:CTN

[SYG^{+}21] Fangyuan Sun, Jia Yu, Xinrui Ge, Ming Yang, and Fanyu Kong. Constrained top-k nearest fuzzy keyword queries on encrypted graph in road network. Computers & Security, 111(??):Article 102456, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002807.

Song:2022:MRM

[SYH^{+}22] Haina Song, Fangfang Yin, Xinyu Han, Tao Luo, and Jianfeng Li. MPDS-RCA: Multi-level privacy-preserving data sharing for resisting collusion attacks based on an integration of CP-ABE and LDP. Computers & Security, 112(??):Article 102523, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003473.

Sun:2020:QPT

[SYL^{+}20] Zhe Sun, Lihua Yin, Chao Li, Weizhe Zhang, Ang Li, and Zhihong Tian. The QoS and privacy trade-off of adversarial deep learning: an evolutionary game approach. Computers & Security, 96(??):Article 101876, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301498.

Sang:2023:TID

[SYZ^{+}23] Yafei Sang, Jisong Yang, Yongzheng Zhang, Shuhao Li, and Peng Chang. Toward IoT device fingerprinting from proprietary protocol traffic via key-blocks aware approach. Computers & Security, 131(??):??, August 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300055X.

Sun:2021:NBD

[SZC^{+}21] Dongming Sun, Xiaolu Zhang, Kim-Kwang Raymond Choo, Liang Hu, and Feng Wang. NLP-based digital forensic investigation platform for online communications. Computers & Security, 104(??):Article 102210, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000341.

Sun:2024:TSM

[SZCZ24] Hui Sun, Tianqing Zhu, Wenhan Chang, and Wanlei Zhou. A two-stage model extraction attack on GANs with a small collected dataset. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005448.

Sun:2022:PDP

[SZL^{+}22] Jie Sun, Lingchen Zhao, Zhuotao Liu, Qi Li, Xinhao Deng, Qian Wang, and Yong Jiang. Practical differentially private online advertising. Computers & Security, 112(??):Article 102504, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100328X.

Shao:2024:AFL

[SZL^{+}24] Jun-Min Shao, Guo-Qiang Zeng, Kang-Di Lu, Guang-Gang Geng, and Jian Weng. Automated federated learning for intrusion detection of industrial control systems based on evolutionary neural architecture search. Computers & Security, 143(??): ??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002128.

Song:2025:CGB

[SZW^{+}25] Xiangpu Song, Yingpei Zeng, Jianliang Wu, Hao Li, Chaoshun Zuo, Qingchuan Zhao, and Shanqing Guo. CSFuzzer: a greybox fuzzer for network protocol using context-aware state feedback. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002706.

Sun:2025:RID

[SZWH25] Shimin Sun, Le Zhou, Ze Wang, and Li Han. Robust intrusion detection based on personalized federated learning for IoT environment. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001312.

Shao:2022:TON

[SZY^{+}22] Kun Shao, Yu Zhang, Junan Yang, Xiaoshuai Li, and Hui Liu. The triggers that open the NLP model backdoors are hidden in the adversarial samples. Computers & Security, 118(??):Article 102730, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001250.

Tiali:2025:MMD

[TA25] Bouthayna El Bouzaidi Tiali and Said Amari. Modeling and modular detection of time attacks in cyber-physical systems based on timed automata with guards and dioid algebra. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482500224X.

Too:2025:EMC

[TAA^{+}25] Dominic Too, Louise Axon, Ioannis Agrafiotis, Michael Goldsmith, and Sadie Creese. Enhancing maritime cyber situational awareness: a cybersecurity visualisation for nonexperts. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001221.

Tantawy:2020:MBR

[TAES20] Ashraf Tantawy, Sherif Abdelwahed, Abdelkarim Erradi, and Khaled Shaban. Model-based risk assessment for cyber physical systems security. Computers & Security, 96(??):Article 101864, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482030136X.

Tlili:2024:EDI

[TAF24] Fadhila Tlili, Samiha Ayed, and Lamia Chaari Fourati. Exhaustive distributed intrusion detection system for UAVs attacks detection and security enforcement (E-DIDS). Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001792.

terBeek:2021:QSR

[tBLLV21] Maurice H. ter Beek, Axel Legay, Alberto Lluch Lafuente, and Andrea Vandin. Quantitative security risk modeling and analysis with RisQFLan. Computers & Security, 109(??):Article 102381, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002054.

Tseng:2023:EEB

[TC23] Chinyang Henry Tseng and Ya-Ting Chang. EBDM: Ensemble binary detection models for multi-class wireless intrusion detection based on deep neural network. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003292.

Tondel:2022:ISP

[TCJS22] Inger Anne Tøndel, Daniela Soares Cruzes, Martin Gilje Jaatun, and Guttorm Sindre. Influencing the security prioritisation of an agile software development project. Computers & Security, 118(??):Article 102744, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001390.

Tayouri:2025:CCO

[TCM^{+}25] David Tayouri, Omri Sgan Cohen, Inbar Maimon, Dudu Mimran, Yuval Elovici, and Asaf Shabtai. CORAL: Container Online Risk Assessment with Logical attack graphs. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824006023.

Tan:2020:GTA

[TCY^{+}20] Choon Lin Tan, Kang Leng Chiew, Kelvin S. C. Yong, San Nah Sze, Johari Abdullah, and Yakub Sebastian. A graph-theoretic approach for the detection of phishing webpages. Computers & Security, 95(??):Article 101793, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030078X.

Tian:2021:TAF

[TCYL21] Pu Tian, Zheyi Chen, Wei Yu, and Weixian Liao. Towards asynchronous federated learning based threat detection: a DC-Adam approach. Computers & Security, 108(??):Article 102344, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821001681.

Tok:2021:SAS

[TD21] Mevlut Serkan Tok and Mehmet Demirci. Security analysis of SDN controller-based DHCP services and attack mitigation with DHCPguard. Computers & Security, 109(??): Article 102394, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821002182.

Turukmane:2024:MME

[TD24] Anil V. Turukmane and Ramkumar Devendiran. M-MultiSVM: an efficient feature selection assisted network intrusion detection system using machine learning. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004972.

Topping:2021:BSB

[TDM^{+}21] Colin Topping, Andrew Dwyer, Ola Michalec, Barnaby Craggs, and Awais Rashid. Beware suppliers bearing gifts!: Analysing coverage of supply chain cyber security in critical national infrastructure sectorial and cross-sectorial frameworks. Computers & Security, 108(??):Article 102324, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001486.

Tang:2023:AST

[TDYQ23] Chenghua Tang, Zheng Du, Mengmeng Yang, and Baohua Qiang. Android static taint analysis based on multi branch search association. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000962.

Tekerek:2021:NAW

[Tek21] Adem Tekerek. A novel architecture for web-based attack detection using convolutional neural network. Computers & Security, 100(??):Article 102096, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303692.

Trabelsi:2023:ACI

[TFJ23] Rahma Trabelsi, Ghofrane Fersi, and Mohamed Jmaiel. Access control in Internet of Things: a survey. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003826.

Tsirakis:2025:OCK

[TFMP25] Orestis Tsirakis, Konstantinos Fysarakis, Vasileios Mavroeidis, and Ioannis Papaefstathiou. Operationalizing cybersecurity knowledge: Design, implementation and evaluation of a knowledge management system for CACAO playbooks. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003852.

Tolah:2021:EAI

[TFP21] Alaa Tolah, Steven M. Furnell, and Maria Papadaki. An empirical analysis of the information security culture key factors framework. Computers & Security, 108(??):Article 102354, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001784.

Tang:2025:SHS

[TG25] Mingsheng Tang and Binbin Ge. Social-Hunter: a social heuristics-based approach to early unveiling unknown malicious logins using valid accounts. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824005753.

Thomas:2022:FDP

[TGOF22] Louise Thomas, Iqbal Gondal, Taiwo Oseni, and Selena (Sally) Firmin. A framework for data privacy and security accountability in data breach communications. Computers & Security, 116(??):Article 102657, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000566.

Tripathi:2021:PTS

[TH21] Nikhil Tripathi and Neminath Hubballi. Preventing time synchronization in NTP broadcast mode. Computers & Security, 102(??):Article 102135, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304089.

Tuptuk:2024:IVI

[TH24] Nilufer Tuptuk and Stephen Hailes. Identifying vulnerabilities of industrial control systems using evolutionary multiobjective optimisation. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005035.

Tian:2020:CDI

[TJ20] Junfeng Tian and Xuan Jing. Cloud data integrity verification scheme for associated tags. Computers & Security, 95(??):Article 101847, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301206.

Tian:2023:PSA

[TJD23] Chuan (Annie) Tian, Matthew L. Jensen, and Alexandra Durcikova. Phishing susceptibility across industries: the differential impact of influence techniques. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003978.

Tzvetanov:2020:FLF

[TK20] Krassimir Tzvetanov and Umit Karabiyik. A first look at forensic analysis of sailfishos. Computers & Security, 99(??):Article 102054, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303278.

T:2025:FRB

[TK25]  Johnpeter T. and Sakthisudhan Karuppanan.  Fuzzy-

rule based optimized hybrid deep learning model for network intrusion detection in SDN enabled IoT network. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000616.

Tatar:2024:CNW

[TKKF24] Unal Tatar, Bilge Karabacak, Omer F. Keskin, and Dominick P. Foti. Charting new waters with CRAMMTS: a survey-driven cybersecurity risk analysis method for maritime stakeholders. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003201.

Thangavelu:2021:ICI

[TKS21] Manisekaran Thangavelu, Venkataraghavan Krishnaswamy, and Mayank Sharma. Impact of comprehensive information security awareness and cognitive characteristics on security incident management — an empirical study. Computers & Security, 109(??):Article 102401, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100225X.

Tung:2025:VAA

[TLHY25] Yi-Chih Tung, En-Cheng Liou, Pen-Chih Hu, and Cheng-Han Yu. VWA-6G AI assisted continuous security monitoring over open RAN service management orchestration. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482500255X.

Tao:2024:PPO

[TLK^{+}24] Yunting Tao, Yuqun Li, Fanyu Kong, Yuliang Shi, Ming Yang, Jia Yu, and Hanlin Zhang. Privacy-preserving outsourcing scheme of face recognition based on locally linear embedding. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002347.

Tran:2020:LPP

[TLKK20] Ngoc Hong Tran, Nhien-An Le-Khac, and M-Tahar Kechadi. Lightweight privacy-preserving data classification. Computers & Security, 97(??):Article 101835, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301085.

Tariq:2020:CAC

[TLKW20] Shahroz Tariq, Sangyup Lee, Huy Kang Kim, and Simon S. Woo. CAN-ADF: the controller area network attack detection framework. Computers & Security, 94(??):Article 101857, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820301292.

Tian:2021:AFG

[TLPY21] Yanzhao Tian, Lixiang Li, Haipeng Peng, and Yixian Yang. Achieving flatness: Graph labeling can generate graphical honeywords. Computers & Security, 104(??):Article 102212, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000365.

Tu:2023:DTD

[TLS23] Yu Tu, Shuang Liu, and Qian Sun. DNS tunnelling detection by fusing encoding feature and behavioral feature. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002675.

Tao:2024:EIE

[TLS24a] Shouzheng Tao, Yezheng Liu, and Chunhua Sun. Examining the inconsistent effect of privacy control on privacy concerns in e-commerce services: the moderating role of privacy experience and risk propensity. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000956.

Tao:2024:UIS

[TLS24b] Shouzheng Tao, Yezheng Liu, and Chunhua Sun. Understanding information sensitivity perceptions and its impact on information privacy concerns in e-commerce services: Insights from China. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005552.

Tuan:2022:DCD

[TLT22] Tong Anh Tuan, Hoang Viet Long, and David Taniar. On detecting and classifying DGA botnets and their families. Computers & Security, 113(??):Article 102549, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003734.

Tu:2024:STT

[TLY^{+}24] Fei-Fan Tu, Dong-Jie Liu, Zhi-Wei Yan, Xiao-Bo Jin, and Guang-Gang Geng. STFT-TCAN: a TCN-attention based multivariate time series anomaly detection architecture with time-frequency analysis for cyber-industrial systems. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002669.

Tayyab:2023:CRD

[TMJ^{+}23] Muhammad Tayyab, Mohsen Marjani, N. Z. Jhanjhi, Ibrahim Abaker Targio Hashem, Raja Sher Afgun Usmani, and Faizan Qamar. A comprehensive review on deep learning algorithms: Security and privacy issues. Computers & Security, 131(??): ??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002079.

Tang:2020:RIB

[TML^{+}20] Fei Tang, Boyang Ma, Jinku Li, Fengwei Zhang, Jipeng Su, and Jianfeng Ma. RansomSpector: an introspectionbased approach to detect crypto ransomware. Computers & Security, 97(??):Article 101997, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302704.

Thompson:2024:SNP

[TMN24] Nik Thompson, Tanya McGill, and Nidhi Narula. “No point worrying” — the role of threat devaluation in information security behavior. Computers & Security, 143(??):??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001998.

Trang:2021:ERS

[TN21] Simon Trang and Ilja Nastjuk. Examining the role of stress and information security policy design in information security compliance behaviour: an experimental study of in-task behaviour. Computers & Security, 104(??):Article 102222, May 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000468.

Talib:2022:ABD

[TNN^{+}22] Manar Abu Talib, Qassim Nasir, Ali Bou Nassif, Takua Mokhamed, Nafisa Ahmed, and Bayan Mahfood. APT beaconing detection: a systematic review. Computers & Security, 122(??):Article 102875, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002693.

Tran:2022:FVT

[TO22] Duong Dinh Tran and Kazuhiro Ogata. Formal verification of TLS 1.2 by automatically generating proof scores. Computers & Security, 123(??):Article 102909, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003029.

Tikhe:2025:FAA

[TP25] Gajanan Nanaji Tikhe and Pushpinder Singh Patheja. Feature attention assisted convolutional stacked sparse auto-encoder model for intrusion detection in network function virtualization environment. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002846.

Tang:2023:BBH

[TQJ^{+}23] Yonghe Tang, Xuyan Qi, Jing Jing, Chunling Liu, and Weiyu Dong. BHMDC: a byte and hex n-gram based malware detection and classification method. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000287.

Tam:2021:GBM

[TRH21] Tracy Tam, Asha Rao, and Joanne Hall. The good, the bad and the missing: a narrative review of cyber-security implications for australian3 small businesses. Computers & Security, 109(??):Article 102385, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002091.

Tundis:2022:FDM

[TRM22] Andrea Tundis, Samuel Ruppert, and Max Mühlhäuser. A feature-driven method for automating the assessment of OSINT cyber threat sources. Computers & Security, 113(??):Article 102576, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821004004.

Taheri:2024:UVD

[TSAG24] Rahim Taheri, Mohammad Shojafar, Farzad Arabikhan, and Alexander Gegov. Unveiling vulnerabilities in deep learningbased malware detection: Differential privacy driven adversarial attacks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003407.

Torkura:2021:CAT

[TSCM21] K. A. Torkura, Muhammad I. H. Sukmana, Feng Cheng, and Christoph Meinel. Continuous auditing and threat detection in multi-cloud infrastructure. Computers & Security, 102(??):Article 102124, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303977.

TajDini:2023:BBA

[TSKG23] Mahyar TajDini, Volodymyr Sokolov, Ievgeniia Kuzminykh, and Bogdan Ghita. Brainwave-based authentication using features fusion. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001086.

Truong:2021:PPF

[TSW^{+}21] Nguyen Truong, Kai Sun, Siyao Wang, Florian Guitton, and YiKe Guo. Privacy preservation in federated learning: an insightful survey from the GDPR perspective. Computers & Security, 110(??):Article 102402, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002261.

Tao:2023:VDT

[TSW^{+}23] Wenxin Tao, Xiaohong Su, Jiayuan Wan, Hongwei Wei, and Weining Zheng. Vulnerability detection through crossmodal feature enhancement and fusion. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002511.

Tang:2024:MNA

[TSW^{+}24] Yunlong Tang, Jing Sun, Huan Wang, Junyi Deng, Liang Tong, and Wenhong Xu. A method of network attackdefense game and collaborative defense decision-making based on hierarchical multi-agent reinforcement learning. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482400172X.

Trnecka:2020:IAR

[TT20] Martin Trnecka and Marketa Trneckova. An incremental algorithm for the role mining problem. Computers & Security, 94(??):Article 101830, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301036.

Tarwireyi:2023:UMA

[TTA23] Paul Tarwireyi, Alfredo Terzoli, and Matthew O. Adigun. Using multi-audio feature fusion for Android malware detection. Computers & Security, 131(??):??, August 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300192X.

Trivedi:2020:NIC

[TTP20] Amit Kumar Trivedi, Dalton Meitei Thounaojam, and Shyamosree Pal. Non-invertible cancellable fingerprint template for fingerprint biometric. Computers & Security, 90 (??):Article 101690, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819302275.

Tsoupidi:2023:TCR

[TTP23] Rodothea Myrsini Tsoupidi, Elena Troubitsyna, and Panagiotis Papadimitratos. Thwarting code-reuse and side-channel attacks in embedded systems. Computers & Security, 133(??): ??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003152.

Tian:2020:PDF

[TTRY20] Ke Tian, Gang Tan, Barbara G. Ryder, and Danfeng (Daphne) Yao. Prioritizing data flows and sinks for app security transformation. Computers & Security, 92(??):Article 101750, May 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818306382.

Torquato:2020:MTD

[TV20] Matheus Torquato and Marco Vieira. Moving target defense in cloud computing: a systematic mapping study. Computers & Security, 92(??):Article 101742, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300286.

Tian:2020:FGA

[TW20] Junfeng Tian and Zhidan Wang. Fine-grained assured data deletion scheme based on attribute association. Computers & Security, 96(??):Article 101936, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030211X.

Tian:2022:DFD

[TWLK22] Jiwei Tian, Buhong Wang, Jing Li, and Charalambos Konstantinou. Datadriven false data injection attacks against cyber-physical power systems. Computers & Security, 121(??):Article 102836, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002309.

Tian:2025:TLT

[TWQ^{+}25] Hui Tian, Mengcheng Wang, Hanyu Quan, Chin-Chen Chang, and Athanasios V. Vasilakos. TEEMRDA: Leveraging trusted execution environments for multi-replica data auditing in cloud storage. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400556X.

Tian:2021:DIA

[TWW21] Junfeng Tian, Haoning Wang, and Meng Wang. Data integrity auditing for secure cloud storage using user behavior prediction. Computers & Security, 105(??):Article 102245, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000699.

Tan:2023:PRF

[TXH^{+}23] Xiao Tan, Qi Xie, Lidong Han, Shengbao Wang, and Wenhao Liu. Proof of retrievability with flexible designated verification for cloud storage. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003966.

Tang:2023:GBI

[TXY^{+}23] Chaofan Tang, Lijuan Xu, Bo Yang, Yongwei Tang, and Dawei Zhao. GRU-based interpretable multivariate time series anomaly detection in industrial control system. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000044.

Tekerek:2022:NMC

[TY22] Adem Tekerek and Muhammed Mutlu Yapici. A novel malware classification and augmentation model based on convolutional neural network. Computers & Security, 112(??):Article 102515, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003394.

Taha:2024:LDF

[TYAH^{+}24] Kamal Taha, Paul D. Yoo, Yousof Al-Hammadi, Sami Muhaidat, and Chan Yeob Yeun. Learning a deep-feature clustering model for gait-based individual identification. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004698.

Tang:2024:MVS

[TYZ^{+}24] Gaigai Tang, Lin Yang, Long Zhang, Hongyu Kuang, and Huiqiang Wang. MRC-VulLoc: Software source code vulnerability localization based on multi-choice reading comprehension. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001172.

Tan:2021:OTS

[TZZ^{+}21] Jinglei Tan, Hengwei Zhang, Hongqi Zhang, Hao Hu, Cheng Lei, and Zhenxiang Qin. Optimal temporospatial strategy selection approach to moving target defense: a FlipIt differential game model. Computers & Security, 108(??):Article 102342, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001668.

Tian:2022:FPP

[TZZ22] Huan Tian, Tianqing Zhu, and Wanlei Zhou. Fairness and privacy preservation for facial images: GAN-based methods. Computers & Security, 122(??):Article 102902, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002954.

Tian:2023:DDI

[TZZ^{+}23] Miao Tian, Yushu Zhang, Youwen Zhu, Liangmin Wang, and Yong Xiang. DIVRS: Data integrity verification based on ring signature in cloud storage. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003947.

Tang:2025:CCA

[TZZY25] Benxiao Tang, Shilin Zhang, Fei Zhu, and Aoshuang Ye. CAPRA: Context-aware patch risk assessment for detecting immature vulnerability in open-source software. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002299.

Ugur:2022:TCD

[UÇ22] Naciye Güliz Ugur and Kübra Çaliskan. Time for decluttering: Digital clutter scaling for individuals and enterprises. Computers & Security, 119(??):Article 102751, August 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001468.

Unal:2021:IFM

[UHK^{+}21] Devrim Unal, Mohammad Hammoudeh, Muhammad Asif Khan, Abdelrahman Abuarqoub, Gregory Epiphaniou, and Ridha Hamila. Integration of federated machine learning and blockchain for the provision of secure big data analytics for Internet of Things. Computers & Security, 109(??): Article 102393, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821002170.

Uchendu:2021:DCS

[UNBF21] Betsy Uchendu, Jason R. C. Nurse, Maria Bada, and Steven Furnell. Developing a cyber security culture: Current practices and future needs. Computers & Security, 109(??):Article 102387, October 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482100211X.

Upadhyay:2020:SSC

[US20] Darshana Upadhyay and Srinivas Sampalli. SCADA (supervisory control and data acquisition) systems: Vulnerability assessment and security recommendations. Computers & Security, 89(??):Article 101666, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302068.

ugli:2025:SIP

[uS25] Alimov Abdulboriy Abdulkhay ugli and Ji Sun Shin. SAFE-IDS: a privacy-preserving framework for overcoming non-IID challenges in federated intrusion detection. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825001804.

Umar:2022:ATT

[UWL22] Mubarak Umar, Zhenqiang Wu, and Xuening Liao. Authenticating tier-two body area network devices through user-specific signal propagation characteristics. Computers & Security, 120(??):Article 102800, September 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001948.

Uuganbayar:2021:OCI

[UYMM21] Ganbayar Uuganbayar, Artsiom Yautsiukhin, Fabio Martinelli, and Fabio Massacci. Optimisation of cyber insurance coverage with selection of cost effective security controls. Computers & Security, 101(??):Article 102121, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303941.

Vijayakanthan:2023:SMF

[VAAG23] Ramyapandian Vijayakanthan, Irfan Ahmed, and Aisha Ali-Gombe. SWMAT: Mel-frequency cepstral coefficientsbased memory fingerprinting for IoT devices. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002080.

Vasan:2020:IBM

[VAW^{+}20] Danish Vasan, Mamoun Alazab, Sobia Wassan, Babak Safaei, and Qin Zheng. Image-based malware classification using ensemble of CNN architectures (IMCEC). Computers & Security, 92(??):Article 101748, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030033X.

VanSlyke:2020:EIH

[VB20] Craig Van Slyke and France Belanger. Explaining the interactions of humans and artifacts in insider security behaviors: the mangle of practice perspective. Computers & Security, 99(??):Article 102064, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303370.

Valkenburg:2024:UTL

[VB24] Bert Valkenburg and Ivano Bongiovanni. Unravelling the three lines model in cybersecurity: a systematic literature review. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000099.

Vondracek:2023:RMI

[VBCM23] Martin Vondrácek, Ibrahim Baggili, Peter Casey, and Mehdi Mekni. Rise of the Metaverse’s immersive virtual reality malware and the man-in-the-room attack and defenses. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003157.

Varga:2021:CTP

[VBF21] Stefan Varga, Joel Brynielsson, and Ulrik Franke. Cyberthreat perception and risk management in the Swedish financial sector. Computers & Security, 105(??):Article 102239, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000638.

Vrhovec:2023:EIS

[VBM23] Simon Vrhovec, Igor Bernik, and Blaz Markelj. Explaining information seeking intentions: Insights from a Slovenian social engineering awareness campaign. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822004308.

Vasalou:2025:DCH

[VBS^{+}25] Asimina Vasalou, Laura Benton, Ana Serta, Andrea Gauthier, Ceylan Besevli, Sarah Turner, Rea Gill, Rachael Payler, Etienne Roesch, Kevin McAreavey, Kim Bauters, Weiru Liu, Hsueh-Ju Chen, Dennis Ivory, Manos Panaousis, and Georgios Loukas. Doing cybersecurity at home: a humancentred approach for mitigating attacks in AI-enabled home devices. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004176.

Vashishtha:2025:SCT

[VC25] Lalit Kumar Vashishtha and Kakali Chatterjee. Strengthening cybersecurity: TestCloudIDS dataset and SparkShield algorithm for robust threat detection. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S016740482400614X.

VandenBroeck:2022:FSP

[VCD22] Jens Van den Broeck, Bart Coppens, and Bjorn De Sutter. Flexible software protection. Computers & Security, 116(??):Article 102636, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822000359.

Vouvoutsis:2025:BSL

[VCP25] Vasilis Vouvoutsis, Fran Casino, and Constantinos Patsakis. Beyond the sandbox: Leveraging symbolic execution for evasive malware classification. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482400498X.

vandenBerg:2024:DUC

[vdB24] Bibi van den Berg. Dealing with uncertainty in cyberspace. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400244X.

vanDam:2022:HPL

[vDK22] Gijs van Dam and Rabiah Abdul Kadir. Hiding payments in lightning network with approximate differentially private payment channels. Computers & Security, 115(??):Article 102623, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000220.

vanderKleij:2022:DDS

[vdKSCY22] Rick van der Kleij, Jan Maarten Schraagen, Beatrice Cadet, and Heather Young. Developing decision support for cybersecurity threat and incident managers. Computers & Security, 113(??):Article 102535, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482100359X.

vanderKleij:2020:AET

[vdKWH20] Rick van der Kleij, Remco Wijn, and Tineke Hof. An application and empirical test of the capability opportunity motivation-behaviour model to data leakage prevention in financial organizations. Computers & Security, 97(??): Article 101970, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404820302431.

VanOuytsel:2024:AML

[VDL24] Charles-Henry Bertrand Van Ouytsel, Khanh Huu The Dam, and Axel Legay. Analysis of machine learning approaches to packing detection. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004467.

Vo:2024:AER

[VDN24] Hoang V. Vo, Hanh P. Du, and Hoa N. Nguyen. APELID: Enhancing real-time intrusion detection with augmented WGAN and parallel ensemble learning. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004777.

vanderSchyff:2021:MEI

[vdSF21] Karl van der Schyff and Stephen Flowerday. Mediating effects of information security awareness. Computers & Security, 106(??):Article 102313, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001371.

vanderSchyff:2023:MRP

[vdSF23] Karl van der Schyff and Stephen Flowerday. The mediating role of perceived risks and benefits when self-disclosing: a study of social media trust and FoMO. Computers & Security, 126(??): ??, March 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004631.

vanderSchyff:2020:DSM

[vdSFF20a] Karl van der Schyff, Stephen Flowerday, and Steven Furnell. Duplicitous social media and data surveillance: an evaluation of privacy risk. Computers & Security, 94(??):Article 101822, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820300961.

vanderSchyff:2020:PRU

[vdSFF20b] Karl van der Schyff, Stephen Flowerday, and Steven Furnell. Privacy risk and the use of Facebook apps: a gender-focused vulnerability assessment. Computers & Security, 96(??):Article 101866, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820301383.

vanderSchyff:2024:PPA

[vdSPR24] Karl van der Schyff, Suzanne Prior, and Karen Renaud. Privacy policy analysis: a scoping review and research agenda. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003705.

vandeWeijer:2024:CDC

[vdWLM24] Steve van de Weijer, Rutger Leukfeldt, and Asier Moneva. Cybercrime during the COVID-19 pandemic: Prevalence, nature and impact of cybercrime for citizens and SME owners in the Netherlands. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300603X.

Villalon-Fonseca:2022:NSC

[VF22] Ricardo Villalón-Fonseca. The nature of security: a conceptual framework for integral-comprehensive modeling of IT security and cybersecurity. Computers & Security, 120(??):Article 102805, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822001997.

vanGeest:2024:AHF

[vGCHZ24] R. J. van Geest, G. Cascavilla, J. Hulstijn, and N. Zannone. The applicability of a hybrid framework for automated phishing detection. Computers & Security, 139(??): ??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000373.

Villalon-Huerta:2022:TTA

[VHMGRR22] Antonio Villalón-Huerta, Hector Marco-Gisbert, and Ismael Ripoll-Ripoll. A taxonomy for threat actors’ persistence techniques. Computers & Security, 121(??):Article 102855, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002498.

Vorobyev:2022:RFP

[VK22] Ivan Vorobyev and Anna Krivitskaya. Reducing false positives in bank anti-fraud systems based on rule induction in distributed tree-based models. Computers & Security, 120(??):Article 102786, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200181X.

Varshini:2025:IDC

[VLV25] G. Y. Sree Varshini, S. Latha, and G. Y. Rajaa Vikhram. Impact and detection of cyber attacks in wide area control application of cyber-physical power system (CPPS). Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002366.

Vrhovec:2021:RTA

[VM21] Simon Vrhovec and Anze Mihelic. Redefining threat appraisals of organizational insiders and exploring the moderating role of fear in cyberattack protection motivation. Computers & Security, 106(??):Article 102309, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001334.

Verma:2020:MMC

[VMS20] Vinita Verma, Sunil K. Muttoo, and V. B. Singh. Multiclass malware classification via first- and second-order texture statistics. Computers & Security, 97(??):Article 101895, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301681.

Vitorino:2023:SRA

[VPM23] João Vitorino, Isabel Praça, and Eva Maia. SoK: Realistic adversarial attacks and defenses for intelligent network intrusion detection. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003437.

Villar-Rodriguez:2023:EIS

[VRPTB^{+}23] Esther Villar-Rodriguez, María Arostegi Pérez, Ana I. Torre-Bastida, Cristina Regueiro Senderos, and Juan López de Armentia. Edge intelligence secure frameworks: Current state and future challenges. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001888.

Varshney:2025:LPT

[VRS^{+}25] Gaurav Varshney, Akanksha Raj, Divya Sangwan, Sharif Abuadbba, Rina Mishra, and Yansong Gao. A login page transparency and visual similarity-based zero-day phishing defense protocol. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825002871.

Vakilinia:2020:VMP

[VS20] Iman Vakilinia and Shamik Sengupta. Vulnerability market as a public-good auction with privacy preservation. Computers & Security, 93(??):Article 101807, June 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300924.

Vadillo:2022:HEU

[VS22] Jon Vadillo and Roberto Santana. On the human evaluation of universal audio adversarial perturbations. Computers & Security, 112(??):Article 102495, January 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003199.

Vepraskas:2024:SDA

[VS24] Matthew J. Vepraskas and Benjamin W. Stirgwolt. A system dynamics approach to building counterfeit avoidance programs. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000129.

Vidyasri:2025:FSF

[VS25] P. Vidyasri and S. Suresh. FDN-SA: Fuzzy deep neural-stacked autoencoder-based phishing attack detection in social engineering. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004930.

vanSchaik:2020:RAA

[vSRW^{+}20] Paul van Schaik, Karen Renaud, Christopher Wilson, Jurjen Jansen, and Joseph Onibokun. Risk as affect: the affect heuristic in cybersecurity. Computers & Security, 90 (??):Article 101651, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404819301956.

V:2025:HMD

[VT25] Hari Vinayak M. V. and Jarin T. A hybrid model for detecting intrusions using stacked autoencoders and extreme gradient boosting. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005182.

Vu:2022:PPN

[Vu22] Duy-Hien Vu. Privacy-preserving Naive Bayes classification in semi-fully distributed data model. Computers & Security, 115(??):Article 102630, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822000293.

Veit:2025:SPD

[VWB^{+}25] Maxime Fabian Veit, Oliver Wiese, Fabian Lucas Ballreich, Melanie Volkamer, Douglas Engels, and Peter Mayer. SoK: the past decade of user deception in emails and today’s email clients’ susceptibility to phishing techniques. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005029.

VanLanduyt:2024:SNQ

[VWJ24] Dimitri Van Landuyt, Vincent Wijshoff, and Wouter Joosen. A study of NoSQL query injection in Neo4j. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300500X.

Wang:2021:DWC

[Wan21] Qinghua Wang. Defending wireless communication against eavesdropping attacks using secret spreading codes and artificial interference. Computers & Security, 103(??):Article 102175, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482030448X.

Wehbe:2025:ESS

[WAPA25] Nathalie Wehbe, Hyame Assem Alameddine, Makan Pourzandi, and Chadi Assi. Empowering 5G SBA security: Time series transformer for HTTP/2 anomaly detection. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482400419X.

Wang:2024:DPC

[WASA24] Shuang Wang, Muhammad Asif, Muhammad Farrukh Shahzad, and Muhammad Ashfaq. Data privacy and cybersecurity challenges in the digital transformation of the banking sector. Computers & Security, 147(??):??, December 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003560.

Woods:2022:CC

[WB22] Daniel W. Woods and Rainer Böhme. The commodification of consent. Computers & Security, 115(??):Article 102605, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000049.

Wu:2023:IMC

[WBC^{+}23] Chia-Yi Wu, Tao Ban, Shin-Ming Cheng, Takeshi Takahashi, and Daisuke Inoue. IoT malware classification based on reinterpreted function-call graphs. Computers & Security, 125(??): ??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004527.

Williams:2020:VAP

[WBN^{+}20] Mark A. Williams, Roberto Camacho Barranco, Sheikh Motahar Naim, Sumi Dey, M. Shahriar Hossain, and Monika Akbar. A vulnerability analysis and prediction framework. Computers & Security, 92(??):Article 101751, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300353.

Waraga:2020:DIA

[WBNT20] Omnia Abu Waraga, Meriem Bettayeb, Qassim Nasir, and Manar Abu Talib. Design and implementation of automated IoT security testbed. Computers & Security, 88(??):Article 101648, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404819301920.

Wu:2025:TCA

[WBZ^{+}25] Zhijun Wu, Yun Bai, Yuan Zhang, Liang Liu, and Meng Yue. TrustCNAV: Certificateless aggregate authentication of civil navigation messages in GNSS. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004772.

Wang:2025:FRG

[WCH^{+}25] Jinlong Wang, Yunhe Cui, Rongfei He, Yi Chen, Chun Guo, and Guowei Shen. FTOA-RP: a ‘group’-based flow entry replacement policy probing and flow table overflow attack method. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003189.

Woo:2025:LSA

[WCL25] Seunghoon Woo, Eunjin Choi, and Heejo Lee. A largescale analysis of the effectiveness of publicly reported security patches. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004863.

Wong:2023:SCT

[WCOZ23] Ann Yi Wong, Eyasu Getahun Chekole, Martín Ochoa, and Jianying Zhou. On the security of containers: Threat modeling, attack analysis, and mitigation strategies. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823000500.

Wang:2023:TAE

[WCP^{+}23] Ruipeng Wang, Kaixiang Chen, Zulie Pan, Yuwei Li, Qianyu Li, Yang Li, Min Zhang, and Chao Zhang. Tunter: Assessing exploitability of vulnerabilities with taint-guided exploitable states exploration. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200387X.

Wang:2023:ESC

[WCW^{+}23] Yu Wang, Liquan Chen, Ge Wu, Kunliang Yu, and Tianyu Lu. Efficient and secure content-based image retrieval with deep neural networks in the mobile cloud computing. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000731.

Wang:2024:MFL

[WCZ^{+}24] Wenpeng Wang, Zhixiang Chen, Ziyang Zheng, Hui Wang, and Junxing Luo. MTA Fuzzer: a low-repetition rate Modbus TCP fuzzing method based on Transformer and Mutation Target Adaptation. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002785.

Wang:2023:AFM

[WCZW23] Wenpeng Wang, Zhixiang Chen, Ziyang Zheng, and Hui Wang. An adaptive fuzzing method based on transformer and protocol similarity mutation. Computers & Security, 129(??): ??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001074.

Wang:2023:HHA

[WDJ^{+}23] Qing Wang, Cong Dong, Shijie Jian, Dan Du, Zhigang Lu, Yinhao Qi, Dongxu Han, Xiaobo Ma, Fei Wang, and Yuling Liu. HANDOM: Heterogeneous attention network model for malicious domain detection. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004515.

Westland:2020:ICS

[Wes20] J. Christopher Westland. The information content of Sarbanes–Oxley in predicting security breaches. Computers & Security, 90(??):Article 101687, March 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818302128.

Wali:2025:EAR

[WFK25] Syed Wali, Yasir Ali Farrukh, and Irfan Khan. Explainable AI and random forest based reliable intrusion detection system. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002317.

Wali:2025:CPA

[WFKH25] Syed Wali, Yasir Ali Farrukh, Irfan Khan, and John A. Hamilton. Covert penetrations: Analyzing and defending SCADA systems from stealth and hijacking attacks. Computers & Security, 156(??):??, September 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001385.

Wu:2023:ZCS

[WFLX23] Anbin Wu, Zhiyong Feng, Xiaohong Li, and Jianmao Xiao. ZTWeb: Cross site scripting detection based on zero trust. Computers & Security, 134(??):??, November 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003449.

Wang:2022:MLE

[WFT22] Zihao Wang, Kar Wai Fok, and Vrizlynn L. L. Thing. Machine learning for encrypted malicious traffic detection: Approaches, datasets and comparative study. Computers & Security, 113(??):Article 102542, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003667.

Wu:2020:MMC

[WFY^{+}20] Zhijun Wu, Wenzhi Feng, Meng Yue, Xinran Xu, and Liang Liu. Mitigation measures of collusive interest flooding attacks in named data networking. Computers & Security, 97(??):Article 101971, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302443.

Wang:2023:ACA

[WGRZ23] Qian Wang, Yuying Gao, Jiadong Ren, and Bing Zhang. An automatic classification algorithm for software vulnerability based on weighted word vector and fusion neural network. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200462X.

Wang:2025:OCM

[WGS^{+}25a] Xiao Wang, Yunchuan Guo, Zhe Sun, Mingjie Yu, Fenghua Li, and Liang Fang. OPMonitor: Continuously monitoring residual over-granted permissions in verified access control policies. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003128.

Wu:2025:MPA

[WGS^{+}25b] Peng Wu, Mohan Gao, Fuhui Sun, Xiaoyan Wang, and Li Pan. Multi-perspective API call sequence behavior analysis and fusion for malware classification. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004826.

Widel:2023:MAL

[WHE^{+}23] Wojciech Wideł, Simon Hacks, Mathias Ekstedt, Pontus Johnson, and Robert Lagerström. The meta attack language — a formal description. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823001943.

Wurzenberger:2024:ASP

[WHLS24] Markus Wurzenberger, Georg Höld, Max Landauer, and Florian Skopik. Analysis of statistical properties of variables in log data for advanced anomaly detection in cyber security. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005412.

Wang:2021:EZT

[WHPL21] Leran Wang, Md Shafaeat Hossain, Joshua Pulfrey, and Lisa Lancor. The effectiveness of zoom touchscreen gestures for authentication and identification and its changes over time. Computers & Security, 111(??):Article 102462, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821002868.

Wu:2024:ALF

[WHW^{+}24] Yali Wu, Yanghu Hu, Junhu Wang, Mengqi Feng, Ang Dong, and Yanxi Yang. An active learning framework using deep Q-network for zero-day attack detection. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824000142.

Wang:2025:FFG

[WHW^{+}25] Zhaoyang Wang, Yanfei Hu, Yu Wen, Boyang Zhang, Shuailou Li, Wenbo Wang, Zheng Liu, and Dan Meng. FineGCP: Finegrained dependency graph community partitioning for attack investigation. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006175.

Wabi:2024:DAD

[WIOO24] Abdullahi Aishatu Wabi, Ismaila Idris, Olayemi Mikail Olaniyi, and Joseph A. Ojeniyi. DDOS attack detection in SDN: Method of attacks, detection techniques, challenges and research gaps. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005618.

Watson:2023:CBA

[WJB23] Katie Watson, Mike Just, and Tessa Berg. A comicbased approach to permission request communication. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003340.

Weickert:2023:CRM

[WJC23] Tobias D. Weickert, Adam Joinson, and Barnaby Craggs. Is cybersecurity research missing a trick? integrating insights from the psychology of habit into research and practice. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000408.

Wang:2023:BDV

[WJP^{+}23] Yan Wang, Peng Jia, Xi Peng, Cheng Huang, and Jiayong Liu. BinVulDet: Detecting vulnerability in binary program via decompiled pseudo code and BiLSTM-attentiono. Computers & Security, 125(??):??, February 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004151.

Wang:2022:RLB

[WJT^{+}22] Wei Wang, Songlei Jian, Yusong Tan, Qingbo Wu, and Chenlin Huang. Representation learning-based network intrusion detection system by capturing explicit and implicit feature interactions. Computers & Security, 112(??):Article 102537, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821003618.

Wang:2023:RUN

[WJT^{+}23] Wei Wang, Songlei Jian, Yusong Tan, Qingbo Wu, and Chenlin Huang. Robust unsupervised network intrusion detection with self-supervised masked context reconstruction. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300041X.

Wang:2020:PSP

[WK20] Nana Wang and Mohan S. Kankanhalli. Protecting sensitive place visits in privacy-preserving trajectory publishing. Computers & Security, 97(??):Article 101949, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030225X.

Wen:2023:QSE

[WK23] Shao-Fang Wen and Basel Katt. A quantitative security evaluation and analysis model for web applications based on OWASP application security verification standard. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300442X.

Wang:2020:SID

[WLC^{+}20] Ye Wang, Qingbao Li, Zhifeng Chen, Ping Zhang, and Guimin Zhang. Shapeshifter: Intelligence-driven data plane randomization resilient to data-oriented programming attacks. Computers & Security, 89(??):Article 101679, February 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819302196.

Wang:2022:EBN

[WLC^{+}22] Zhi Wang, Chaoge Liu, Xiang Cui, Jie Yin, and Xutong Wang. EvilModel 2.0: Bringing neural network models into malware attacks. Computers & Security, 120(??):Article 102807, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002012.

Wang:2024:WDK

[WLDW24] Zhiying Wang, Guiyun Li, Yajing Diao, and Nianxin Wang. Warning deterrence or knowledge guidance? Research on triggering mechanism of phishing sensitivity. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001767.

Wang:2024:KKB

[WLH^{+}24] Gaosheng Wang, Peipei Liu, Jintao Huang, Haoyu Bin, Xi Wang, and Hongsong Zhu. KnowCTI: Knowledgebased cyber threat intelligence entity and relation extraction. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001251.

Wang:2021:IDM

[WLHC21] Zhendong Wang, Yaodi Liu, Daojing He, and Sammy Chan. Intrusion detection methods based on integrated deep learning model. Computers & Security, 103(??):Article 102177, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821000018.

Weir:2025:HFA

[WLKHD25] Charles Weir, Cecilia Loureiro-Koechlin, Lucy Hunt, and Louise Dennis. The human factor: Addressing computing risks for critical national infrastructure towards 2040. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002135.

Wang:2023:MRS

[WLKT23] Xuan Wang, Yaojie Li, Hanieh Javadi Khasraghi, and Cherie Trumbach. The mediating role of security anxiety in internet threat avoidance behavior. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003395.

Wang:2023:LCE

[WLL^{+}23] Zhihua Wang, Zhenyu Li, Zeminghui Li, Yingheng Xu, Fazhi Qi, and Jinsheng Kong. A low cost and effective multiinstance abnormal driving behavior detection system under edge computing. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002729.

Wang:2024:JRT

[WLL24] Xiaodi Wang, Zhonglin Liu, and Jiayong Liu. Joint relational triple extraction with enhanced representation and binary tagging framework in cybersecurity. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003067.

Wanyan:2024:NUN

[WLLC24] Hanxiao Wanyan, Yingxu Lai, Jing Liu, and Hao Chen. NCMFuzzer: Using non-critical field mutation and test case combination to improve the efficiency of ICS protocol fuzzing. Computers & Security, 141(??):??, June 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824001123.

Weng:2024:CER

[WLLL24a] Juanjuan Weng, Zhiming Luo, Dazhen Lin, and Shaozi Li. Comparative evaluation of recent universal adversarial perturbations in image classification. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004868.

Weng:2024:LTT

[WLLL24b] Juanjuan Weng, Zhiming Luo, Dazhen Lin, and Shaozi Li. Learning transferable targeted universal adversarial perturbations by sequential meta-learning. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004947.

Wang:2022:TTA

[WLLS22] Jincheng Wang, Zhuohua Li, John C. S. Lui, and Mingshen Sun. Topology-theoretic approach to address attribute linkage attacks in differential privacy. Computers & Security, 113(??):Article 102552, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100376X.

Waqdan:2025:SRA

[WLM25] Mofareh Waqdan, Habib Louafi, and Malek Mouhoub. Security risk assessment in IoT environments: a taxonomy and survey. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001452.

Wang:2020:DMB

[WLQ20] Meng Wang, Yiqin Lu, and Jiancheng Qin. A dynamic MLPbased DDoS attack detection method using feature selection and feedback. Computers & Security, 88(??):Article 101645, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404819301890.

Wang:2025:PPB

[WLQ^{+}25] Zehui Wang, Hao Li, Yinhao Qi, Wei Qiao, Song Liu, Chen Zhang, Bo Jiang, and Zhigang Lu. PathWatcher: a pathbased behavior detection method for attack detection and investigation. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002524.

Wang:2024:WSW

[WLY^{+}24] Zhiquan Wang, Tengyao Li, Meijuan Yin, Xiaoyun Yuan, Xiangyang Luo, and Lingling Li. WF3A: a N-shot website fingerprinting with effective fusion feature attention. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482400097X.

Wu:2023:DFS

[WLZ^{+}23] Yinwei Wu, Meijin Li, Qi Zeng, Tao Yang, Junfeng Wang, Zhiyang Fang, and Luyu Cheng. DroidRL: Feature selection for Android malware detection with reinforcement learning. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000366.

Wang:2024:RTA

[WLZ^{+}24] Zhenduo Wang, Saifei Li, Lijie Zhang, Chunduo Hu, and Lianshan Yan. A Red Team automated testing modeling and online planning method for post-penetration. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002505.

Wiley:2020:MTI

[WMC20] Ashleigh Wiley, Agata McCormac, and Dragana Calic. More than the individual: Examining the relationship between culture and information security awareness. Computers & Security, 88(??):Article 101640, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819301841.

Wang:2023:STR

[WMGH^{+}23] Han Wang, Luis Muñoz-González, Muhammad Zaid Hameed, David Eklund, and Shahid Raza. SparSFA: Towards robust and communication-efficient peer-to-peer federated learning. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000925.

Wang:2023:ADA

[WMK23] Hang Wang, David J. Miller, and George Kesidis. Anomaly detection of adversarial examples using class-conditional generative adversarial networks. Computers & Security, 124(??): ??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003480.

Wang:2023:TAF

[WML^{+}23] Leiqi Wang, Xiu Ma, Ning Li, Qiujian Lv, Yan Wang, Weiqing Huang, and Haiyan Chen. TGPrint: Attack fingerprint classification on encrypted network traffic based graph convolution attention networks. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003760.

Wang:2020:BNA

[WNF20] Jiali Wang, Martin Neil, and Norman Fenton. A Bayesian network approach for cybersecurity risk assessment implementing and extending the FAIR model. Computers & Security, 89(??):Article 101659, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404819300604.

Wajdi:2020:AED

[WNI20] Moussa Wajdi, Basty Nadia, and Ghazouani Ines. Asymmetric effect and dynamic relationships over the cryptocurrencies market. Computers & Security, 96(??):Article 101860, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301322.

Wu:2021:SPS

[WP21] Yu Wu and Li Pan. SG-PAC: a stochastic game approach to generate personal privacy paradox access-control policies in social networks. Computers & Security, 102(??):Article 102157, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820304302.

Woodring:2024:EPP

[WPAG24] Justin Woodring, Katherine Perez, and Aisha Ali-Gombe. Enhancing privacy policy comprehension through Privacify: a user-centric approach using advanced language models. Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400302X.

Wongwiwatchai:2020:DPI

[WPS20] Nattanon Wongwiwatchai, Phannawhat Pongkham, and Kunwadee Sripanidkulchai. Detecting personally identifiable information transmission in Android applications using light-weight static analysis. Computers & Security, 99(??):Article 102011, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302844.

Wang:2025:AAA

[WPZW25] Chiheng Wang, Jianshan Peng, Junhu Zhu, and Qingxian Wang. AugPersist: Automatically augmenting the persistence of coverage-based greybox fuzzing for persistent software. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004048.

Wang:2025:RAD

[WQL25] Yingqing Wang, Guihe Qin, and Yanhua Liang. A reliability anomaly detection method based on enhanced GRU-Autoencoder for Vehicular Fog Computing services. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005236.

Wei:2024:EDE

[WR24] Qianjin Wei and Gang Rao. EPFL-DAC: Enhancing privacy in federated learning with dynamic aggregation and clipping. Computers & Security, 143(??):??, August 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400213X.

Wasko:2021:UAR

[WRG^{+}21] Shannon Wasko, Rebecca E. Rhodes, Megan Goforth, Nathan Bos, Hannah P. Cowley, Gerald Matthews, Alice Leung, Satish Iyengar, and Jonathon Kopecky. Using alternate reality games to find a needle in a haystack: an approach for testing insider threat detection methods. Computers & Security, 107(??):Article 102314, August 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001383.

Wu:2021:TMO

[WS21] Jianyuan Wu and Wei Sun. Towards multi-operation image anti-forensics with generative adversarial networks. Computers & Security, 100(??):Article 102083, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303564.

Walshe:2022:CVD

[WS22] T. Walshe and A. C. Simpson. Coordinated Vulnerability Disclosure programme effectiveness: Issues and recommendations. Computers & Security, 123(??):Article 102936, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822003285.

Woods:2024:HMA

[WS24] Naomi Woods and Mikko Siponen. How memory anxiety can influence password security behavior. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004996.

Winkler:2025:PTD

[WS25a] Aidan M. Winkler and Prinkle Sharma. Proactive threat detection in enterprise systems using Wazuh: a MITRE ATT&CK evaluation. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003918.

Woods:2025:QSA

[WS25b] Naomi Woods and Mikko Siponen. Questioning a security assumption: Are unique passwords harder to remember than reused or modified passwords? Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002342.

Wang:2020:PXR

[WSC^{+}20] Eric Ke Wang, RuiPei Sun, Chien-Ming Chen, Zuodong Liang, Saru Kumari, and Muhammad Khurram Khan. Proof of Xrepute blockchain consensus protocol for IoT systems. Computers & Security, 95(??):Article 101871, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301449.

Wang:2022:KDC

[WSQ^{+}22] Senmiao Wang, Luli Sun, Sujuan Qin, WenMin Li, and Wentao Liu. KRTunnel: DNS channel detector for mobile devices. Computers & Security, 120(??):Article 102818, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002127.

Wang:2022:ESI

[WSW^{+}22] Jie Wang, Jiaming Shi, Xin Wen, Liang Xu, Ke Zhao, Fuyang Tao, Wenbiao Zhao, and Xiuying Qian. The effect of signal icon and persuasion strategy on warning design in online fraud. Computers & Security, 121(??):Article 102839, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002334.

Wang:2024:OBA

[WSW24] Wenjie Wang, Yuanhai Shao, and Yiju Wang. Optimizationbased adversarial perturbations against twin support vector machines. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004832.

Wang:2023:FME

[WT23] Zihao Wang and Vrizlynn L. L. Thing. Feature mining for encrypted malicious traffic detection with deep learning and other machine learning algorithms. Computers & Security, 128(??): ??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000536.

Wu:2022:ALF

[WTC^{+}22] Jianyuan Wu, Tianyao Tong, Yifang Chen, Xiangui Kang, and Wei Sun. An adversarial learning framework with crossdomain loss for median filtered image restoration and antiforensics. Computers & Security, 112(??):Article 102497, January 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003217.

Wang:2025:PIL

[WTHL25] Xiaoqing Wang, Yuanjing Tian, Keman Huang, and Bin Liang. Practically implementing an LLM-supported collaborative vulnerability remediation process: a team-based approach. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004188.

Wang:2021:MFP

[WTL^{+}21] Wei Wang, Jing Tian, Fang Lv, Guodong Xin, Yingfan Ma, and Bailing Wang. Mining frequent pyramid patterns from time series transaction data with custom constraints. Computers & Security, 100(??):Article 102088, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303618.

Wang:2023:DDM

[WTL^{+}23] Han Wang, Zhangguo Tang, Huanzhou Li, Jian Zhang, and Cheng Cai. DDOFM: Dynamic malicious domain detection method based on feature mining. Computers & Security, 130 (??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001700.

Wolf:2024:BSN

[WTL^{+}24] Maximilian Wolf, Julian Tritscher, Dieter Landes, Andreas Hotho, and Daniel Schlör. Benchmarking of synthetic network data: Reviewing challenges and approaches. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002980.

Wang:2021:NFS

[WTW21] Peng Wang, Zhijie Tang, and Junfeng Wang. A novel few-shot malware classification approach for unknown family recognition with multi-prototype modeling. Computers & Security, 106(??):Article 102273, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000973.

Wang:2022:PPM

[WTX^{+}22] Chenguang Wang, Zhu Tianqing, Ping Xiong, Wei Ren, and Kim-Kwang Raymond Choo. A privacy preservation method for multiple-source unstructured data in online social networks. Computers & Security, 113(??):Article 102574, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003989.

Waelchli:2025:RRS

[WW25a] Sandro Waelchli and Yoshija Walter. Reducing the risk of social engineering attacks using SOAR measures in a real world environment: a case study. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004425.

Wang:2025:LSD

[WW25b] Jin Wang and Liping Wang. LR-STGCN: Detecting and mitigating low-rate DDoS attacks in SDN based on spatialtemporal graph neural network. Computers & Security, 154 (??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482500149X.

Wang:2022:OSS

[WWH^{+}22] Bo Wang, Yue Wang, Jiayao Hou, Yi Li, and Yanqing Guo. Open-set source camera identification based on envelope of data clustering optimization (EDCO). Computers & Security, 113(??):Article 102571, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003953.

Wang:2025:AAT

[WWL^{+}25a] Siyu Wang, Haiyong Wu, Ning Lu, Wenbo Shi, and Zhiquan Liu. ATSDetector: an Android Trojan spyware detection approach with multi-features. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740482400525X.

Wen:2025:VVT

[WWL25b] Zhaoyu Wen, Zhiqiang Wang, and Biao Liu. VERTFuzz: Version transformer-driven fuzzing for complex file parsers. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482500330X.

Wu:2023:LDI

[WWLH23] Yuanlu Wu, Yan Wo, Caiyu Li, and Guoqiang Han. Learning domain-invariant representation for generalizing face forgery detection. Computers & Security, 130(??):??, July 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001906.

Wang:2025:PLA

[WWS^{+}25] Weiping Wang, Chenyu Wang, Hong Song, Kai Chen, and Shigeng Zhang. ProvGOutLiner: a lightweight anomaly detection method based on process behavior features within provenance graphs. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002780.

Wu:2022:JPS

[WWW^{+}22] Songyun Wu, Bo Wang, Zhiliang Wang, Shuhan Fan, Jiahai Yang, and Jia Li. Joint prediction on security event and time interval through deep learning. Computers & Security, 117(??):Article 102696, June 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000943.

Wang:2024:STK

[WWWS24] Xintong Wang, Zixuan Wang, Enliang Wang, and Zhixin Sun. Spatial-temporal knowledge distillation for lightweight network traffic anomaly detection. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005461.

Wu:2024:PPB

[WWZ24] Guangli Wu, Xingyue Wang, and Jing Zhang. PeerG: a P2P botnet detection method based on representation learning and graph contrastive learning. Computers & Security, 140(??): ??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824000762.

Wang:2023:SAO

[WX23] Ruiling Wang and Yakui Xue. Stability analysis and optimal control of worm propagation model with saturated incidence rate. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822004552.

Wei:2023:LDL

[WXD23] Chongbo Wei, Gaogang Xie, and Zulong Diao. A lightweight deep learning framework for botnet detecting at the IoT edge. Computers & Security, 129(??):??, June 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823001050.

Wang:2025:MSR

[WXJ^{+}25] Lan Wang, Jialing Xu, Liyun Jia, Tao Wang, Yujie Xu, and Xingchen Liu. Multi-strategy RIME optimization algorithm for feature selection of network intrusion detection. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000823.

Wang:2024:SCS

[WXT^{+}24] Rongcun Wang, Senlei Xu, Yuan Tian, Xingyu Ji, Xiaobing Sun, and Shujuang Jiang. SCL-CVD: Supervised contrastive learning for code vulnerability detection via Graph-CodeBERT. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002992.

Wei:2025:SAO

[WXZ^{+}25] Xiaomin Wei, Yizhen Xu, Haibin Zhang, Cong Sun, Xinghua Li, Fenghua Huang, and Jianfeng Ma. Sensor attack online classification for UAVs using machine learning. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005340.

Wang:2024:GSB

[WYF^{+}24] Qiannan Wang, Changchun Yin, Liming Fang, Zhe Liu, Run Wang, and Chenhao Lin. GhostEncoder: Stealthy backdoor attacks with dynamic triggers to pre-trained encoders in self-supervised learning. Computers & Security, 142(??): ??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001561.

Wang:2022:PPS

[WYJ^{+}22] Jie Wang, Zhaoxia Yin, Jing Jiang, Jin Tang, and Bin Luo. PISA: Pixel skipping-based attentional black-box adversarial attack. Computers & Security, 123(??):Article 102947, December 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200339X.

Wang:2024:TBF

[WYJ^{+}24] Wenbo Wang, Peng Yi, Junfang Jiang, Peng Zhang, and Xiang Chen. Transformer-based framework for alert aggregation and attack prediction in a multi-stage attack. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004431.

Wang:2024:CAF

[WYL^{+}24] Xiaoyu Wang, Xiaobo Yang, Xueping Liang, Xiu Zhang, Wei Zhang, and Xiaorui Gong. Combating alert fatigue with AlertPro: Context-aware alert prioritization using reinforcement learning for multi-step attack detection. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004935.

Wang:2024:HSE

[WYP^{+}24] Xiaoyan Wang, Jingjing Yang, Zixiao Peng, Shunfang Wang, and Ming Huang. Hilbert signal envelope-based multifeatures methods for GNSS spoofing detection. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002645.

Wang:2022:BBA

[WYW^{+}22] Qiuhua Wang, Hui Yang, Guohua Wu, Kim-Kwang Raymond Choo, Zheng Zhang, Gongxun Miao, and Yizhi Ren. Black-box adversarial attacks on XSS attack detection model. Computers & Security, 113(??):Article 102554, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003783.

Wang:2023:HEH

[WZB^{+}23] Qifan Wang, Lei Zhou, Jianli Bai, Yun Sing Koh, Shujie Cui, and Giovanni Russello. HT2ML: an efficient hybrid framework for privacy-preserving Machine Learning using HE and TEE. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004194.

Wang:2024:PDC

[WZC^{+}24] Maoli Wang, Xiaodong Zang, Jianbo Cao, Bowen Zhang, and Shengbao Li. PhishHunter: Detecting camouflaged IDN-based phishing attacks via Siamese neural network. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823005783.

Wang:2020:PPH

[WZCP20] Rong Wang, Yan Zhu, Chin-Chen Chang, and Qiang Peng. Privacy-preserving high-dimensional data publishing for classification. Computers & Security, 93(??):Article 101785, June 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820300705.

Wang:2025:RCA

[WZDZ25] Weiguang Wang, Xiao Zhang, Jinlian Du, and Wenbing Zhao. Robust continuous authentication via multi-channel photoplethysmography signals: a wearable wristband solution for uncontrolled environments. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482500375X.

Williams:2024:FHA

[WZG24] Jason A. Williams, Humayun Zafar, and Saurabh Gupta. Fortifying healthcare: an action research approach to developing an effective SETA program. Computers & Security, 138(??): ??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005643.

Wang:2024:ICN

[WZW24a] Meng Wang, Yahao Zhang, and WeiPing Wen. Improved capsule networks based on Nash equilibrium for malicious code classification. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823004133.

Wang:2024:BBP

[WZW^{+}24b] Weibo Wang, Yifeng Zheng, Songlei Wang, Zhongyun Hua, Lei Xu, and Yansong Gao. BopSkyline: Boosting privacy-preserving skyline query service in the cloud. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824001044.

Wang:2025:MCD

[WZW^{+}25] Ning Wang, Jian Zhuang, Zhigang Wang, Zhiqiang Wei, Yu Gu, Peng Tang, and Ge Yu. Multidimensional categorical data collection under shuffled differential privacy. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824006072.

Wang:2022:SDF

[WZX^{+}22] Chuanwang Wang, Junjie Zhang, Ming Xu, Haodong Zhang, and Weili Han. #Segments: a dominant factor of password security to resist against data-driven guessing. Computers & Security, 121(??):Article 102848, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002425.

Wang:2023:SMB

[WZX23] Yongkang Wang, Di-Hua Zhai, and Yuanqing Xia. SCFL: Mitigating backdoor attacks in federated learning based on SVD and clustering. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003243.

Wu:2023:SCV

[WZY^{+}23] Bolun Wu, Futai Zou, Ping Yi, Yue Wu, and Liang Zhang. SlicedLocator: Code vulnerability locator based on sliced dependence graph. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823003796.

Wang:2022:TAE

[WZZ^{+}22] Zhen Wang, Yitao Zheng, Hai Zhu, Chang Yang, and Tianyi Chen. Transferable adversarial examples can efficiently fool topic models. Computers & Security, 118(??):Article 102749, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001444.

Wang:2020:USF

[WZZW20] Ding Wang, Xizhe Zhang, Zijian Zhang, and Ping Wang. Understanding security failures of multi-factor authentication schemes for multi-server environments. Computers & Security, 88(??):Article 101619, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740481930166X.

Xu:2025:ECP

[XCC^{+}25] Zhenwu Xu, Xingshu Chen, Liangguo Chen, Xiao Lan, Hao Ren, and Changxiang Shen. An efficient and commercial proof of storage scheme supporting dynamic data updates. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002986.

Xu:2024:EDO

[XCL^{+}24] Zhenwu Xu, Xingshu Chen, Xiao Lan, Rui Tang, Shuyu Jiang, and Changxiang Shen. Empowering data owners: an efficient and verifiable scheme for secure data deletion. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002839.

Xia:2021:BAR

[XCSZ21] Yusheng Xia, Rongmao Chen, Jinshu Su, and Hongcheng Zou. Balancing anonymity and resilience in anonymous communication networks. Computers & Security, 101(??):Article 102106, February 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820303795.

Xue:2025:TPT

[XCWC25] Jiacheng Xue, Xiang Chen, Jiyu Wang, and Zhanqi Cui. Towards prompt tuning-based software vulnerability assessment with continual learning. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004899.

Xiao:2025:MPS

[XCWY25] Jingxu Xiao, Chaowen Chang, Ping Wu, and Lu Yuan. Multiprobability sampling-based detection of malicious switching nodes in SDN. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000136.

Xiao:2025:GGN

[XCY^{+}25] Fengrui Xiao, Shuangwu Chen, Jian Yang, Huasen He, Xiaofeng Jiang, Xiaobin Tan, and Dong Jin. GRAIN: Graph neural network and reinforcement learning aided causality discovery for multi-step attack scenario reconstruction. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004851.

Xu:2021:PMN

[XFY21] Zhifeng Xu, Xianjin Fang, and Gaoming Yang. Malbert: a novel pre-training method for malware detection. Computers & Security, 111(??):Article 102458, December 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002820.

Xiang:2020:DMA

[XGL20] Zongyuan Xiang, Dong Guo, and Qiang Li. Detecting mobile advanced persistent threats based on large-scale DNS logs. Computers & Security, 96(??):Article 101933, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820302091.

Xu:2020:ESD

[XGS^{+}20] Guangquan Xu, Bingjiang Guo, Chunhua Su, Xi Zheng, Kaitai Liang, Duncan S. Wong, and Hao Wang. Am I eclipsed? A smart detector of eclipse attacks for Ethereum. Computers & Security, 88(??):Article 101604, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404818313798.

Xiao:2021:IBM

[XGS^{+}21] Mao Xiao, Chun Guo, Guowei Shen, Yunhe Cui, and Chaohui Jiang. Image-based malware classification using section distribution information. Computers & Security, 110(??):Article 102420, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821002443.

Xu:2025:PPP

[XGS25] Ting Xu, Xinrui Ge, and Changheng Shao. PMkR: Privacypreserving multi-keyword top-k reachability query. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002147.

Xie:2024:DDN

[XGX^{+}24] Yibo Xie, Gaopeng Gou, Gang Xiong, Zhen Li, and Wei Xia. DomEye: Detecting network covert channel of domain fronting with throughput fluctuation. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002815.

Xu:2021:DTR

[XHB^{+}21] Guangwei Xu, Songhua Han, Yanke Bai, Xiangyang Feng, and Yanglan Gan. Data tag replacement algorithm for data integrity verification in cloud storage. Computers & Security, 103(??):Article 102205, April 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821000298.

Xia:2020:DMC

[XHTZ20] Chang Xia, Jingyu Hua, Wei Tong, and Sheng Zhong. Distributed K-means clustering guaranteeing local differential privacy. Computers & Security, 90(??):Article 101699, March 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404819302366.

Xue:2022:PRP

[XHW^{+}22] Mingfu Xue, Can He, Yinghao Wu, Shichang Sun, Yushu Zhang, Jian Wang, and Weiqiang Liu. PTB: Robust physical backdoor attacks against deep neural networks in real world. Computers & Security, 118(??):Article 102726, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001213.

Xue:2020:LLO

[XHWL20] Mingfu Xue, Can He, Jian Wang, and Weiqiang Liu. LOPA: a linear offset based poisoning attack method against adaptive fingerprint authentication system. Computers & Security, 99(??):Article 102046, December 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820303199.

Xu:2025:UTP

[XKJG25] Shi-Jie Xu, Kai-Chuan Kong, Xiao-Bo Jin, and Guang-Gang Geng. Unveiling traffic paths: Explainable path signature feature-based encrypted traffic classification. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005893.

Xue:2025:HHN

[XKY25] Yankun Xue, Chunying Kang, and Hongchen Yu. HAE-HRL: a network intrusion detection system utilizing a novel autoencoder and a hybrid enhanced LSTM-CNN-based residual network. Computers & Security, 151(??):??, April 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000173.

Xie:2025:DHS

[XLD25] Yuanfeng Xie, Hanqing Luo, and Aoxue Ding. Design of a high-stability QPUF and QRNG circuit based on CC-NOT gate. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003839.

Xu:2023:UGD

[XLG^{+}23] Guangquan Xu, Wenqing Lei, Lixiao Gong, Jian Liu, Hongpeng Bai, Kai Chen, Ran Wang, Wei Wang, Kaitai Liang, Weizhe Wang, Weizhi Meng, and Shaoying Liu. UAF-GUARD: Defending the use-after-free exploits via fine-grained memory permission management. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004400.

Xiang:2025:FGA

[XLH^{+}25] Dongming Xiang, Shuai Lin, Ke Huang, Zuohua Ding, Guanjun Liu, and Xiaofeng Li. A fine-grained approach for Android taint analysis based on labeled taint value graphs. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400467X.

Xiong:2023:AAT

[XLL23] Wen Ding Xiong, Kai Lun Luo, and Rui Li. AIDTF: Adversarial training framework for network intrusion detection. Computers & Security, 128(??):??, May 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000512.

Xie:2025:HDM

[XLL^{+}25] Yuntian Xie, Ting Lei, Zimo Li, Yujing Yang, Chunyin Chen, and Yuanyuan Long. How do mental models affect cybersecurity awareness? The roles of questioning styles, need for cognition, and graphical representations. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005984.

Xuan:2024:BTM

[XLS24] Bona Xuan, Jin Li, and Yafei Song. BiTCN-TAEfficientNet malware classification approach based on sequence and RGB fusion. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482400035X.

Xie:2024:PBM

[XLW24] Bailin Xie, Qi Li, and Yu Wang. PHP-based malicious webshell detection based on abstract syntax tree simplification and explicit duration recurrent networks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003547.

Xiao:2024:AMA

[XLWC24] Nan Xiao, Bo Lang, Ting Wang, and Yikai Chen. APT-MMF: an advanced persistent threat actor attribution method based on multimodal and multilevel feature fusion. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002657.

Xie:2020:HMH

[XLY^{+}20] Jiang Xie, Shuhao Li, Xiaochun Yun, Yongzheng Zhang, and Peng Chang. HSTF-Model: an HTTP-based Trojan detection model via the Hierarchical Spatio-temporal Features of Traffics. Computers & Security, 96(??):Article 101923, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482030198X.

Xie:2022:ADA

[XLZ^{+}22] Jiang Xie, Shuhao Li, Yongzheng Zhang, Peishuai Sun, and Hongbo Xu. Analysis and detection against network attacks in the overlapping phenomenon of behavior attribute. Computers & Security, 121(??):Article 102867, October 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822002619.

Xie:2024:ADM

[XLZ24] Shujiang Xie, Lian Li, and Yian Zhu. Anomaly detection for multivariate time series in IoT using discrete wavelet decomposition and dual graph attention networks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824003808.

Xiang:2021:REI

[XMK21] Zhen Xiang, David J. Miller, and George Kesidis. Reverse engineering imperceptible backdoor attacks on deep neural networks for detection and training set cleansing. Computers & Security, 106(??):Article 102280, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001048.

Xu:2025:MSD

[XQS25] Hongsheng Xu, Akeel Qadir, and Saima Sadiq. Malicious SMS detection using ensemble learning and SMOTE to improve mobile cybersecurity. Computers & Security, 154(??): ??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001324.

Xu:2025:PUF

[XRW^{+}25] Peng Xu, Tingting Rao, Wei Wang, Zhaojun Lu, and Kaitai Liang. Power of union: Federated honey password vaults against differential attack. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002810.

Xu:2023:GQE

[XSC^{+}23] Guangquan Xu, Hongfei Shao, Jingyi Cui, Hongpeng Bai, Jiliang Li, Guangdong Bai, Shaoying Liu, Weizhi Meng, and Xi Zheng. GenDroid: a query-efficient black-box Android adversarial attack framework. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002699.

Xing:2023:PAN

[XSK23] Ying Xing, Hui Shu, and Fei Kang. PeerRemove: an adaptive node removal strategy for P2P botnet based on deep reinforcement learning. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000391.

Xie:2024:TDM

[XSL^{+}24] Yitian Xie, Mikko Siponen, Gabriella Laatikainen, Gregory D. Moody, and Xiaosong Zheng. Testing the dominant mediator in EPPM: an empirical study on household anti-malware software users. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000774.

Xu:2024:UDD

[XT24] Jiacheng Xu and Chengxiang Tan. Unawareness detection: Discovering black-box malicious models and quantifying privacy leakage risks. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004753.

Xu:2020:IEL

[XW20] Feng Xu and Merrill Warkentin. Integrating elaboration likelihood model and herd theory in information security message persuasiveness. Computers & Security, 98(??):Article 102009, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302820.

Xu:2024:FJE

[XWJ^{+}24] Haoran Xu, Yongjun Wang, Zhiyuan Jiang, Shuhui Fan, Shaojing Fu, and Peidai Xie. Fuzzing JavaScript engines with a syntax-aware neural program model. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002529.

Xu:2020:PSP

[XWW^{+}20] Lijuan Xu, Bailing Wang, Lianhai Wang, Dawei Zhao, Xiaohui Han, and Shumian Yang. PLC-SEIFF: a programmable logic controller security incident forensics framework based on automatic construction of security constraints. Computers & Security, 92(??):Article 101749, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300328.

Xu:2021:MEA

[XWY21] Jian Xu, Xuequn Wang, and Lvxin Yan. The moderating effect of abusive supervision on information security policy compliance: Evidence from the hospitality industry. Computers & Security, 111(??):Article 102455, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002790.

Xia:2020:CCE

[XWZ^{+}20] Pengcheng Xia, Haoyu Wang, Bowen Zhang, Ru Ji, Bingyu Gao, Lei Wu, Xiapu Luo, and Guoai Xu. Characterizing cryptocurrency exchange scams. Computers & Security, 98(??):Article 101993, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302662.

Xiong:2020:DRP

[XX20] Wei Xiong and Li Xiong. Data resource protection based on smart contract. Computers & Security, 98(??):Article 102004, November 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302777.

Xiao:2021:PWD

[XXZ^{+}21] Xi Xiao, Wentao Xiao, Dianyan Zhang, Bin Zhang, Guangwu Hu, Qing Li, and Shutao Xia. Phishing websites detection via CNN and multi-head self-attention on imbalanced datasets. Computers & Security, 108(??):Article 102372, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001966.

Xu:2024:BTA

[XY24] Zisheng Xu and Qiao Yan. Boosting the transferability of adversarial CAPTCHAs. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003055.

Xu:2025:TPP

[XYG25] Kaidi Xu, Jia Yu, and Wenjing Gao. Towards privacypreserving compressed sensing reconstruction in cloud. Computers & Security, 151(??):??, April 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://

www.sciencedirect.com/science/article/pii/S0167404825000379.

Xin:2020:RAS

[XYH^{+}20] Ruishan Xin, Yidong Yuan, Jiaji He, Shuai Zhen, and Yiqiang Zhao. Random active shield generation based on modified artificial fish-swarm algorithm. Computers & Security, 88(??):Article 101552, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818312926.

Xiao:2025:PLS

[XYZ^{+}25] Junbi Xiao, Zhaoyu Yin, Yuhao Zhou, Kai Liu, Jian Wang, and Peiying Zhang. P4Drop: a lightweight security function for filtering TCP spoofing packets on programmable switches. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002901.

Xia:2025:PPB

[XZG^{+}25] Wenhao Xia, Liancheng Zhang, Yi Guo, Hongtao Zhang, and Lanxin Cheng. P4NSA: P4-based security protection technology for IPv6 neighbor solicitation and advertisement spoofing. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000896.

Xie:2025:LCP

[XZLG25] Yuancheng Xie, Zhaoxin Zhang, Ning Li, and Haoyang Gao. LeakFocus: Catching the perpetrator in routing leak event. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824006060.

Xiang:2025:NRM

[XZWD25] Huifang Xiang, Ruimei Zhang, Ziling Wang, and Di Dong. New results on modeling and hybrid control for malware propagation in cyber-physical systems. Computers & Security, 157(??): ??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002226.

Xiao:2024:CAW

[XZY^{+}24] Xi Xiao, Xiang Zhou, Zhenyu Yang, Le Yu, Bin Zhang, Qixu Liu, and Xiapu Luo. A comprehensive analysis of website fingerprinting defenses on Tor. Computers & Security, 136(??): ??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300487X.

Xiao:2024:USC

[XZZ^{+}24] Haitao Xiao, Yan Zhu, Bin Zhang, Zhigang Lu, Dan Du, and Yuling Liu. Unveiling shadows: a comprehensive framework for insider threat detection based on statistical and sequential analysis. Computers & Security, 138(??):??, March 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005758.

Xu:2025:AID

[XZZ^{+}25] Lijuan Xu, ZiCheng Zhao, Dawei Zhao, Xin Li, XiYu Lu, and DingYu Yan. AJSAGE: a intrusion detection scheme based on Jump-Knowledge Connection To Graph-SAGE. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824005698.

Yuan:2025:BWR

[YAC25] Ying Yuan, Giovanni Apruzzese, and Mauro Conti. Beyond the west: Revealing and bridging the gap between Western and Chinese phishing website detection. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004206.

Yang:2020:HRH

[Yan20] Bo Yang. A highly-random hopping sequence for jammingresilient channel rendezvous in distributed cognitive radio networks. Computers & Security, 96(??):Article 101809, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818309416.

Yang:2023:ESD

[Yan23] Benyuan Yang. Enforcement of separation of duty constraints in attribute-based access control. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404823002043.

Yogi:2025:NUC

[YC25] Manas Kumar Yogi and A. S. N. Chakravarthy. A novel user centric privacy mechanism in cyber physical system. Computers & Security, 149(??):??, February 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004681.

Yang:2021:NUG

[YCCZ21] Rui Yang, Tian-Jie Cao, Xiu-Qing Chen, and Feng-Rong Zhang. A novel and universal GAN-based countermeasure to recover adversarial examples to benign examples. Computers & Security, 111(??):Article 102457, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002819.

Yi:2024:ABC

[YCLZ24] Tao Yi, Xingshu Chen, Qindong Li, and Yi Zhu. An anomaly behavior characterization method of network traffic based on Spatial Pyramid Pool (SPP). Computers & Security, 141(??): ??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400110X.

Yang:2020:ABS

[YCMM20] Runkai Yang, Xiaolin Chang, Jelena Misić, and Vojislav B. Misić. Assessing blockchain selfish mining in an imperfect network: Honest and selfish miner views. Computers & Security, 97(??):Article 101956, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820302327.

Yu:2023:SIS

[YCWW23] Chuan Yu, Shuhui Chen, Ziling Wei, and Fei Wang. Sec-Checker: Inspecting the security implementation of 5G Commercial Off-The-Shelf (COTS) mobile devices. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002717.

Ye:2024:DCI

[YFdCdC^{+}24] Junjian Ye, Xincheng Fei, Xavier de Carné de Carnavalet, Lianying Zhao, Lifa Wu, and Mengyuan Zhang. Detecting command injection vulnerabilities in Linux-based embedded firmware with LLM-based taint analysis of library functions. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002761.

Yu:2021:AAP

[YFO^{+}21] Lu Yu, Yu Fu, Jonathan Oakley, Oluwakemi Hambolu, and Richard Brooks. On accuracy and anonymity of privacypreserving negative survey (NS) algorithms. Computers & Security, 105(??):Article 102206, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404821000304.

Yousaf:2025:SCS

[YGB^{+}25] Awais Yousaf, Sean Gunawan, Sunil Basnet, Victor Bolbot, Jianying Zhou, and Osiris A. Valdez Banda. STPA-Cyber: a semi-automated cyber risk assessment framework for maritime cybersecurity. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002482.

Yang:2023:RRM

[YGC^{+}23] Wang Yang, Mingzhe Gao, Ligeng Chen, Zhengxuan Liu, and Lingyun Ying. RecMaL: Rectify the malware family label via hybrid analysis. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000871.

Yuan:2025:DDN

[YGC25] Yali Yuan, Jian Ge, and Guang Cheng. DeMarking: a defense for network flow watermarking in real-time. Computers & Security, 152(??):??, May 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000446.

Yang:2024:FTS

[YGG24] Hanlin Yang, Yajun Guo, and Yimin Guo. Fault-tolerant security-efficiency combined authentication scheme for mannedunmanned teaming. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003572.

Yang:2025:DLD

[YGK^{+}25] Shanquan Yang, Yansong Gao, Boyu Kuang, Yixuan Yang, and Anmin Fu. DFirmSan: a lightweight dynamic memory sanitizer for Linux-based firmware. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001567.

Yang:2025:PSI

[YGW^{+}25] Yahui Yang, Yangyang Geng, Qiang Wei, Man Zhou, and Xin Che. Physical semantic inference method for industrial control proprietary protocol data fields. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404825003438.

Yu:2025:KLH

[YHC25] Yun-Che Yu, Ci-Yi Hung, and Li-Der Chou. Kernellevel hidden rootkit detection based on eBPF. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002718.

Yuan:2024:SFE

[YHH^{+}24] Xinwei Yuan, Shu Han, Wei Huang, Hongliang Ye, Xianglong Kong, and Fan Zhang. A simple framework to enhance the adversarial robustness of deep learning-based intrusion detection system. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005539.

Yang:2024:ATS

[YHS24] Xue Yang, Enda Howley, and Michael Schukat. ADT: Time series anomaly detection for cyber-physical systems via deep reinforcement learning. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824001263.

Yang:2023:DFL

[YHW^{+}23] Run Yang, Hui He, Yulong Wang, Yue Qu, and Weizhe Zhang. Dependable federated learning for IoT intrusion detection against poisoning attacks. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002912.

Yan:2023:PMV

[YHY23a] Haotian Yan, Haibo Hu, and Qingqing Ye. Partial message verification in fog-based industrial Internet of things. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004406.

Yuan:2023:PSA

[YHY23b] Yachao Yuan, Yu Huang, and Yali Yuan. PrSLoc: Sybil attack detection for localization with private observers using differential privacy. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001992.

Yang:2022:PVL

[YHZ^{+}22] Kunyu Yang, Xuexian Hu, Qihui Zhang, Jianghong Wei, and Wenfen Liu. VAEPass: a lightweight passwords guessing model based on variational auto-encoder. Computers & Security, 114(??):Article 102587, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004107.

Yin:2023:ERE

[YHZ23] Ya Yin, Carol Hsu, and Zhongyun Zhou. Employees’ inrole and extra-role information security behaviors from the P-E fit perspective. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003000.

Yang:2024:DGB

[YJK^{+}24] Chengyi Yang, Kun Jia, Deli Kong, Jiayin Qi, and Aimin Zhou. DP-GSGLD: a Bayesian optimizer inspired by differential privacy defending against privacy leakage in federated learning. Computers & Security, 142(??):??, July 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001408.

Yamin:2022:MEC

[YK22a] Muhammad Mudassar Yamin and Basel Katt. Modeling and executing cyber security exercise scenarios in cyber ranges. Computers & Security, 116(??):Article 102635, May 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000347.

Yamin:2022:UCA

[YK22b] Muhammad Mudassar Yamin and Basel Katt. Use of cyber attack and defense agents in cyber ranges: a case study. Computers & Security, 122(??):Article 102892, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002644.

Yousefi:2023:RBP

[YK23] Fares Yousefi and Hoshang Kolivand. A robust brain pattern for brain-based authentication methods using deep breath. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004303.

Yang:2024:CBT

[YK24] Zhenhua Yang and Jun Kong. Cue-based two factor authentication. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003730.

Yamin:2020:CRS

[YKG20] Muhammad Mudassar Yamin, Basel Katt, and Vasileios Gkioulos. Cyber ranges and security testbeds: Scenarios, functions, tools and architecture. Computers & Security, 88(??):Article 101636, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404819301804.

Kim:2022:TCM

[yKK22] Jae yeol Kim and Hyuk-Yoon Kwon. Threat classification model for security information event management focusing on model efficiency. Computers & Security, 120(??): Article 102789, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404822001821.

Yoo:2024:GGB

[YKK24] Jeong Do Yoo, Haerin Kim, and Huy Kang Kim. GUIDE: GAN-based UAV IDS Enhancement. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482400378X.

Yamin:2021:SGT

[YKN21] Muhammad Mudassar Yamin, Basel Katt, and Mariusz Nowostawski. Serious games as a tool to model attack and defense scenarios for cyber-security exercises. Computers & Security, 110(??):Article 102450, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002741.

Yoo:2025:MMN

[YKSK25] Jeong Do Yoo, Gang Min Kim, Min Geun Song, and Huy Kang Kim. MeNU: Memorizing normality for UAV anomaly detection with a few sensor values. Computers & Security, 150(??): ??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005546.

Yin:2024:SSL

[YKY24] Zhichao Yin, Xian Kong, and Chunyong Yin. Semi-supervised log anomaly detection based on bidirectional temporal convolution network. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824001093.

Yang:2020:ZKP

[YL20] Xiaohui Yang and Wenjie Li. A zero-knowledge-proof-based digital identity management scheme in blockchain. Computers & Security, 99(??):Article 102050, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303230.

Yang:2023:AOC

[YL23] Xiaohui Yang and Xiang Li. ATDAD: One-class adversarial learning for tabular data anomaly detection. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003590.

Yao:2023:DCP

[YLDW23] Lin Yao, Jia Li, Jing Deng, and Guowei Wu. Detection of cache pollution attack based on federated learning in ultradense network. Computers & Security, 124(??):??, January 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822003571.

Yang:2024:MNM

[YLH^{+}24] Jiyun Yang, Hanwei Li, Lijun He, Tao Xiang, and Yujie Jin. MDADroid: a novel malware detection method by constructing functionality-API mapping. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003663.

Yang:2022:SLR

[YLL^{+}22] Zhen Yang, Xiaodong Liu, Tong Li, Di Wu, Jinjiang Wang, Yunwei Zhao, and Han Han. A systematic literature review of methods and datasets for anomaly-based network intrusion detection. Computers & Security, 116(??):Article 102675, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000736.

Yang:2023:AML

[YLL23] Guangyi Yang, Xiaoxing Liu, and Beixin Li. Anti-money laundering supervision by intelligent algorithm. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002547.

Yan:2021:HBH

[YLPZ21] Han Yan, Senlin Luo, Limin Pan, and Yifei Zhang. HAN-BSVD: a hierarchical attention network for binary software vulnerability detection. Computers & Security, 108(??):Article 102286, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821001103.

Yang:2022:FIF

[YLS^{+}22] Jingcheng Yang, Hongwei Li, Shuo Shao, Futai Zou, and Yue Wu. FS-IDS: a framework for intrusion detection based on fewshot learning. Computers & Security, 122(??):Article 102899, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822002929.

Yeoh:2023:ZTC

[YLSJ23] William Yeoh, Marina Liu, Malcolm Shore, and Frank Jiang. Zero trust cybersecurity: Critical success factors and a maturity assessment framework. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482300322X.

Yang:2024:ESI

[YLW24a] Benyuan Yang, Lili Luo, and Zhimeng Wang. Ensuring secure interoperation of access control in a multidomain environment. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482300531X.

Yue:2024:DAA

[YLW^{+}24b] Hao Yue, Tong Li, Di Wu, Runzi Zhang, and Zhen Yang. Detecting APT attacks using an attack intent-driven and sequence-based learning approach. Computers & Security, 140 (??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482400049X.

Yin:2020:DSD

[YLY20] Chunyong Yin, Bo Li, and Zhichao Yin. A distributed sensing data anomaly detection scheme. Computers & Security, 97(??):Article 101960, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302364.

Yuan:2023:BMT

[YLY^{+}23] Qingjun Yuan, Chang Liu, Wentao Yu, Yuefei Zhu, Gang Xiong, Yongjuan Wang, and Gaopeng Gou. BoAu: Malicious traffic detection with noise labels based on boundary augmentation. Computers & Security, 131(??):??, August 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823002109.

Yin:2021:AEP

[YLZ^{+}21] Fan Yin, Rongxing Lu, Yandong Zheng, Jun Shao, Xue Yang, and Xiaohu Tang. Achieve efficient position-heap-based privacy-preserving substring-of-keyword query over cloud. Computers & Security, 110(??):Article 102432, November 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100256X.

Yue:2025:NDH

[YLZ^{+}25] Shudan Yue, Qingbao Li, Guimin Zhang, Xiaonan Li, Bocheng Xu, and Song Tian. NPFTaint: Detecting highly exploitable vulnerabilities in Linux-based IoT firmware with network parsing functions. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003682.

Yan:2022:PPH

[YM22] Xiaobei Yan and Maode Ma. A privacy-preserving handover authentication protocol for a group of MTC devices in 5G networks. Computers & Security, 116(??):Article 102601, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821004247.

Yadav:2022:ECN

[YMR^{+}22] Pooja Yadav, Neeraj Menon, Vinayakumar Ravi, Sowmya Vishvanathan, and Tuan D. Pham. EfficientNet convolutional neural networks-based Android malware detection. Computers & Security, 115(??):Article 102622, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000219.

Yang:2023:CCK

[YMZ^{+}23] Haitian Yang, Xiang Meng, Xuan Zhao, Yan Wang, Yuejun Liu, Xiaoyu Kang, Jiahui Shen, and Weiqing Huang. CKDAN: Content and keystroke dual attention networks with pre-trained models for continuous authentication. Computers & Security, 128(??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016740482300069X.

Yuste:2021:ARD

[YP21] Javier Yuste and Sergio Pastrana. Avaddon ransomware: an in-depth analysis and decryption of infected systems. Computers & Security, 109(??):Article 102388, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002121.

Yazdinejad:2020:PBS

[YPDC20] Abbas Yazdinejad, Reza M. Parizi, Ali Dehghantanha, and Kim-Kwang Raymond Choo. P4-to-blockchain: a secure blockchain-enabled packet parser for software defined networking. Computers & Security, 88(??):Article 101629, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404819301762.

Yuste:2022:OCC

[YPT22] Javier Yuste, Eduardo G. Pardo, and Juan Tapiador. Optimization of code caves in malware binaries to evade machine learning detectors. Computers & Security, 116(??):Article 102643, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822000426.

Yasasin:2020:FIS

[YPWS20] Emrah Yasasin, Julian Prester, Gerit Wagner, and Guido Schryen. Forecasting IT security vulnerabilities — an empirical analysis. Computers & Security, 88(??):Article 101610, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S016740481830854X.

Yang:2024:TTC

[YQL24] Tengfei Yang, Yuansong Qiao, and Brian Lee. Towards trustworthy cybersecurity operations using Bayesian Deep Learning to improve uncertainty quantification of anomaly detection. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002116.

Yu:2022:FDF

[YS22] Mengran Yu and Shiliang Sun. FE-DaST: Fast and effective data-free substitute training for black-box adversarial attacks. Computers & Security, 113(??):Article 102555, February 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821003795.

Yilmaz:2021:FGC

[YSM^{+}21] Fadi Yilmaz, Meera Sridhar, Abhinav Mohanty, Vasant Tendulkar, and Kevin W. Hamlen. A fine-grained classification and security analysis of web-based virtual machine vulnerabilities. Computers & Security, 105(??):Article 102246, June 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821000705.

Yu:2024:CSA

[YSZ24] Jingze Yu, Wenting Shen, and Xi Zhang. Cloud storage auditing and data sharing with data deduplication and private information protection for cloud-based EMR. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002359.

Yuan:2021:DLI

[YW21] Shuhan Yuan and Xintao Wu. Deep learning for insider threat detection: Review, challenges and opportunities. Computers & Security, 104(??):Article 102221, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000456.

Yang:2022:LCB

[YWK^{+}22] Wencheng Yang, Song Wang, James Jin Kang, Michael N. Johnstone, and Aseel Bedari. A linear convolution-based cancelable fingerprint biometric authentication system. Computers & Security, 114(??):Article 102583, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004065.

Yuan:2020:BLM

[YWL^{+}20] Baoguo Yuan, Junfeng Wang, Dong Liu, Wen Guo, Peng Wu, and Xuhua Bao. Byte-level malware classification based on Markov images and deep learning. Computers & Security, 92(??):Article 101740, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300262.

Yang:2025:ADQ

[YWL25] Libin Yang, Menghan Wang, and Wei Lou. An automated dynamic quality assessment method for cyber threat intelligence. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824003845.

Yeoh:2022:SSC

[YWPC22] William Yeoh, Shan Wang, Ales Popovic, and Noman H. Chowdhury. A systematic synthesis of critical success factors for cybersecurity. Computers & Security, 118(??):Article 102724, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001195.

Yu:2024:AAL

[YWW24] Rongwei Yu, Yong Wang, and Wang Wang. AMAD: Active learning-based multivariate time series anomaly detection for large-scale IT systems. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005138.

Yu:2022:SBH

[YWWH22] Lianyi Yu, Qiangjiang Wang, Yan Wo, and Guoqiang Han. Secure biometric hashing against relation-based attacks via maximizing min-entropy. Computers & Security, 118(??):Article 102750, July 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001456.

Yang:2022:AMD

[YWX^{+}22] Shaojie Yang, Yongjun Wang, Haoran Xu, Fangliang Xu, and Mantun Chen. An Android malware detection and classification approach based on contrastive lerning. Computers & Security, 123(??):Article 102915, December 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482200308X.

Yao:2020:PPE

[YWZ20] Yuanfan Yao, Ziyu Wang, and Pan Zhou. Privacy-preserving and energy efficient task offloading for collaborative mobile computing in IoT: an ADMM approach. Computers & Security, 96(??):Article 101886, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301590.

Yang:2024:NAM

[YWZ^{+}24] Hongyu Yang, Youwei Wang, Liang Zhang, Xiang Cheng, and Ze Hu. A novel Android malware detection method with API semantics extraction. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005606.

Yang:2024:IDM

[YWZW24] Zhongjun Yang, Qi Wang, Xuejun Zong, and Guogang Wang. Intrusion detection method based on improved social network search algorithm. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000828.

Yun:2022:DUH

[YXL^{+}22] Xiaochun Yun, Jiang Xie, Shuhao Li, Yongzheng Zhang, and Peishuai Sun. Detecting unknown HTTP-based malicious communication behavior via generated adversarial flows and hierarchical traffic features. Computers & Security, 121(??):Article 102834, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822002280.

Yuan:2025:ASA

[YZC25] Yali Yuan, Weiyi Zou, and Guang Cheng. Attack smarter: Attention-driven fine-grained webpage fingerprinting attacks. Computers & Security, 158(??):??, November 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003323.

Yang:2020:DBL

[YZL^{+}20] Liqun Yang, Xiaoming Zhang, Zhi Li, Zhoujun Li, and Yueying He. Detecting bi-level false data injection attack based on time series analysis method in smart grid. Computers & Security, 96(??):Article 101899, September 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301723.

Yin:2024:PVF

[YZL^{+}24] Benxin Yin, Hanlin Zhang, Jie Lin, Fanyu Kong, and Leyun Yu. PVFL: Verifiable federated learning and prediction with privacy-preserving. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000014.

Yang:2025:AAS

[YZLS25] Kai Yang, Yingjun Zhang, Ting Li, and Limin Sun. ASIDS: Acoustic side-channel based intrusion detection system for industrial robotic arms. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825002755.

Ye:2025:PDP

[YZP^{+}25] Xinxin Ye, Youwen Zhu, Jie Pan, Miao Zhang, and Hai Deng. PrivRS: Differentially private synthetic data generation via role similarity. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825003050.

Yu:2020:UAM

[YZW^{+}20] Guangsheng Yu, Xuan Zha, Xu Wang, Wei Ni, Kan Yu, J. Andrew Zhang, and Ren Ping Liu. A unified analytical model for proof-of-X schemes. Computers & Security, 96(??):Article 101934, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302108.

Zhou:2024:SAD

[ZC24] Man Zhou and Xin Che. Stealthy attack detection based on controlled invariant subspace for autonomous vehicles. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300545X.

Zahoor:2023:FAI

[ZCAP23] Ehtesham Zahoor, Maryam Chaudhary, Sabina Akhtar, and Olivier Perrin. A formal approach for the identification of redundant authorization policies in Kubernetes. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003838.

Zhu:2022:BBA

[ZCD^{+}22] Yiran Zhu, Lei Cui, Zhenquan Ding, Lun Li, Yongji Liu, and Zhiyu Hao. Black box attack and network intrusion detection using machine learning for malicious traffic. Computers & Security, 123(??):Article 102922, December 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003145.

Zhou:2020:CEM

[ZCJ^{+}20] Yuyang Zhou, Guang Cheng, Shanqing Jiang, Yuyu Zhao, and Zihan Chen. Cost-effective moving target defense against DDoS attacks using trilateral game and multi-objective Markov decision processes. Computers & Security, 97(??):Article 101976, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820302492.

Zhou:2024:HIE

[ZCJ^{+}24] Qihang Zhou, Wenzhuo Cao, Xiaoqi Jia, Shengzhi Zhang, Jiayun Chen, Nan Jiang, Weijuan Zhang, Haichao Du, Zhenyu Song, and Qingjia Huang. HClave: an isolated execution environment design for hypervisor runtime security. Computers & Security, 144(??):??, September 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824002256.

Zhang:2025:SSPb

[ZCL^{+}25] Bonan Zhang, Chao Chen, Ickjai Lee, Kyungmi Lee, and Kok-Leong Ong. A survey on security and privacy issues in wearable health monitoring devices. Computers & Security, 155(??):??, August 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825001427.

Zhang:2023:IDM

[ZCLW23] Tianyue Zhang, Wei Chen, Yuxiao Liu, and Lifa Wu. An intrusion detection method based on stacked sparse autoencoder and improved Gaussian mixture model. Computers & Security, 128 (??):??, May 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000548.

Zoppi:2023:WAC

[ZCPB23] Tommaso Zoppi, Andrea Ceccarelli, Tommaso Puccetti, and Andrea Bondavalli. Which algorithm can detect unknown attacks? Comparison of supervised, unsupervised and meta-learning algorithms for intrusion detection. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823000172.

Zheng:2023:GGN

[ZCS^{+}23] Haibin Zheng, Jinyin Chen, Wenchang Shangguan, Zhaoyan Ming, Xing Yang, and Zhijun Yang. GONE: a generic O(1) NoisE layer for protecting privacy of deep neural networks. Computers & Security, 135(??):??, December 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823003814.

Zou:2024:FCN

[ZCW^{+}24] Binghui Zou, Chunjie Cao, Longjuan Wang, Sizheng Fu, Tonghua Qiao, and Jingzhang Sun. FACILE: a capsule network with fewer capsules and richer hierarchical information for malware image classification. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823005163.

Zhu:2024:PEP

[ZCZW24] Erzhou Zhu, Kang Cheng, Zhizheng Zhang, and Huabin Wang. PDHF: Effective phishing detection model combining optimal artificial and automatic deep features. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004716.

Zhu:2024:EDS

[ZCZZ24] Zhengwei Zhu, Miaojie Chen, Chenyang Zhu, and Yanping Zhu. Effective defense strategies in network security using improved double dueling deep Q-network. Computers & Security, 136(??):??, January 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823004881.

Zhou:2021:SSR

[ZD21] Quanqiang Zhou and Liangliang Duan. Semi-supervised recommendation attack detection based on co-forest. Computers & Security, 109(??):Article 102390, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002145.

Zhan:2023:AAM

[ZDH^{+}23] Dazhi Zhan, Yexin Duan, Yue Hu, Lujia Yin, Zhisong Pan, and Shize Guo. AMGmal: Adaptive mask-guided adversarial attack against malware detection with minimal perturbation. Computers & Security, 127(??):??, April 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823000135.

Zohourian:2024:IPL

[ZDM^{+}24] Alireza Zohourian, Sajjad Dadkhah, Heather Molyneaux, Euclides Carlos Pinto Neto, and Ali A. Ghorbani. IoT-PRIDS: Leveraging packet representations for intrusion detection in IoT networks. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003390.

Zhang:2025:ABA

[ZDM^{+}25] Yongheng Zhang, Tingwen Du, Yunshan Ma, Xiang Wang, Yi Xie, Guozheng Yang, Yuliang Lu, and Ee-Chien Chang.

AttacKG+: Boosting attack graph construction with Large Language Models. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824005261.

Zhang:2024:WWD

[ZDZD24] Zhenghao Zhang, Jianwei Ding, Qi Zhang, and Qiyao Deng. WaTrojan: Wavelet domain trigger injection for backdoor attacks. Computers & Security, 140(??):??, May 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824000683.

Zenitani:2023:AGA

[Zen23] Kengo Zenitani. Attack graph analysis: an explanatory guide. Computers & Security, 126(??):??, March 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004734.

Zhu:2024:IAT

[ZFG^{+}24] Peican Zhu, Zepeng Fan, Sensen Guo, Keke Tang, and Xingyu Li. Improving adversarial transferability through hybrid augmentation. Computers & Security, 139(??):??, April 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005849.

Zambianco:2025:PDS

[ZFS25] Marco Zambianco, Claudio Facchinetti, and Domenico Siracusa. A proactive decoy selection scheme for cyber deception using MITRE ATT&CK. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004498.

Zhao:2024:ENI

[ZFT24] Xinxing Zhao, Kar Wai Fok, and Vrizlynn L. L. Thing. Enhancing network intrusion detection performance using generative adversarial networks. Computers & Security, 145(??):??, October 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003109.

Zhao:2025:TSF

[ZFW^{+}25] Jingwen Zhao, Yan Fu, Yanxia Wu, Jibin Dong, and Ruize Hong. Thread-sensitive fuzzing for concurrency bug detection. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004760.

Zhou:2022:LHT

[ZFWL22] Zhipeng Zhou, Qiang Fu, Qianjun Wei, and Qiang Li. LEGO: a hybrid toolkit for efficient 2PC-based privacy-preserving machine learning. Computers & Security, 120(??):Article 102782, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822001778.

Zhou:2024:DPP

[ZG24] Peng Zhou and Yuhan Gao. Detecting prototype pollution for node.js: Vulnerability review and new fuzzing inputs. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005357.

Zhao:2025:AAG

[ZGG^{+}25] Dongliang Zhao, Jiaxing Guo, Chunxiang Gu, Yonghui Zheng, and Xieli Zhang. AGLFuzz: Automata-Guided Fuzzing for detecting logic errors in security protocol implementations. Computers & Security, 149(??):??, February 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824002840.

Zhao:2021:PDS

[ZGM21] Jinxiong Zhao, Sensen Guo, and Dejun Mu. DouBiGRU-A: Software defect detection algorithm based on attention mechanism and double BiGRU. Computers & Security, 111(??): Article 102459, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404821002832.

Zhou:2021:MSW

[ZGNZ21] Peng Zhou, Xiaojing Gu, Surya Nepal, and Jianying Zhou. Modeling social worm propagation for advanced persistent threats. Computers & Security, 108(??):Article 102321, September 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404821001450.

Zhao:2024:MMS

[ZGS24] Chengmei Zhao, Feng Gao, and Zhihao Shen. Multimotion sensor behavior based continuous authentication on smartphones using gated two-tower transformer fusion networks. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823006089.

Zang:2023:ASR

[ZGZL23] Xiaodong Zang, Jian Gong, Xinchang Zhang, and Guiqing Li. Attack scenario reconstruction via fusing heterogeneous threat intelligence. Computers & Security, 133(??):??, October 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003309.

Zhou:2024:RAD

[ZH24] Quanqiang Zhou and Cheng Huang. A recommendation attack detection approach integrating CNN with bagging. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003353.

Zhao:2024:TFC

[Zha24] Rui Zhao. Toward the flow-centric detection of browser fingerprinting. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005515.

Zhou:2025:SED

[ZHC25a] Man Zhou, Lansheng Han, and Xin Che. Strengthening edge defense: a differential game-based edge intelligence strategy against APT attacks. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S016740482500269X.

Zhu:2025:LET

[ZHC^{+}25b] Tiantian Zhu, Wenya He, Tieming Chen, Jiabo Zhang, Mingqi Lv, Hongmei Li, Aohan Zheng, Jie Zheng, Mingjun Ma, Xiangyang Zheng, Zhengqiu Weng, and Shuying Wu. LinTracer: an efficient tracking system for cyberattack chains fusing entity and event semantics. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825001026.

Zhang:2024:MFV

[ZHG^{+}24] Sanfeng Zhang, Qingyu Hao, Zijian Gong, Fengzhou Zhu, Yan Wang, and Wang Yang. MDD-FedGNN: a vertical federated graph learning framework for malicious domain detection. Computers & Security, 147(??):??, December 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003985.

Zhang:2025:LLM

[ZHH^{+}25] Qiang Zhang, Cheng Huang, Jiaxuan Han, Shuyi Jiang, and Jiayong Liu. LowPTor: a lightweight method for detecting extremely low-proportion darknet traffic. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825001099.

Zhao:2021:RLA

[ZHJ^{+}21] Xu Zhao, Guangqiu Huang, Jin Jiang, Ling Gao, and Maozhen Li. Research on lightweight anomaly detection of multimedia traffic in edge computing. Computers & Security, 111(??):Article 102463, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S016740482100287X. See corrigendum [ZHJ^{+}23].

Zhao:2023:CRL

[ZHJ^{+}23] Xu Zhao, Guangqiu Huang, Jin Jiang, Ling Gao, and Maozhen Li. Corrigendum to ‘Research on Lightweight Anomaly Detection of Multimedia Traffic in Edge Computing’ Computer and Security, 111(2021) 102463. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404822003960. See [ZHJ^{+}21].

Zhou:2020:CMN

[ZHL^{+}20] Man Zhou, Lansheng Han, Hongwei Lu, Cai Fu, and Dezhi An. Cooperative malicious network behavior recognition algorithm in e-commerce. Computers & Security, 95(??):Article 101868, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301401.

Zhang:2022:AAY

[ZHL^{+}22] Zhaoxin Zhang, Shize Huang, Xiaowen Liu, Bingjie Zhang, and Decun Dong. Adversarial attacks on YOLACT instance segmentation. Computers & Security, 116(??):Article 102682, May 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822000803.

Zhang:2025:SSPa

[ZHL25] Shunliang Zhang, Weiqing Huang, and Yinlong Liu. A systematic survey on physical layer security oriented to reconfigurable intelligent surface empowered 6G. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482400405X.

Zhong:2022:RLB

[ZHZ^{+}22] Fangtian Zhong, Pengfei Hu, Guoming Zhang, Hong Li, and Xiuzhen Cheng. Reinforcement learning based adversarial malware example generation against black-box detectors. Computers & Security, 121(??):Article 102869, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002632.

Zhang:2025:HLH

[ZJC^{+}25] Yucheng Zhang, Xiaolin Ju, Xiang Chen, Misbahul Amin, and Zilong Ren. HGAN4VD: Leveraging heterogeneous graph attention networks for enhanced vulnerability detection. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002378.

Zhao:2024:SCK

[ZJH^{+}24] Xiaojuan Zhao, Rong Jiang, Yue Han, Aiping Li, and Zhichao Peng. A survey on cybersecurity knowledge graph construction. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004340.

Zhu:2022:FSM

[ZJJS^{+}22] Jinting Zhu, Julian Jang-Jaccard, Amardeep Singh, Ian Welch, Harith Al-Sahaf, and Seyit Camtepe. A few-shot meta-learning based Siamese neural network using entropy features for ransomware classification. Computers & Security, 117(??):Article 102691, June 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200089X.

Zhang:2020:ABL

[ZJL^{+}20] Wenjing Zhang, Bo Jiang, Ming Li, Ravi Tandon, Qiao Liu, and Hui Li. Aggregation-based location privacy: an information theoretic approach. Computers & Security, 97(??):Article 101953, October 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820302297.

Zhang:2022:CRN

[ZJW^{+}22] Chunying Zhang, Donghao Jia, Liya Wang, Wenjie Wang, Fengchun Liu, and Aimin Yang. Comparative research on network intrusion detection methods based on machine learning. Computers & Security, 121(??):Article 102861, October 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822002553.

Zhang:2023:SAI

[ZJY^{+}23] Wenfang Zhang, Heng Jiao, Zhuoqun Yan, Xiaomin Wang, and Muhammad Khurram Khan. Security analysis and improvement of a public auditing scheme for secure data storage in fog-to-cloud computing. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004114.

Zakariyya:2023:TRE

[ZKAK23] Idris Zakariyya, Harsha Kalutarage, and M. Omar Al-Kadri. Towards a robust, effective and resource efficient machine learning technique for IoT security monitoring. Computers & Security, 133(??):??, October 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823002985.

Zeng:2023:FSV

[ZKL^{+}23] Qingyang Zeng, Mohammad Kavousi, Yinhong Luo, Ling Jin, and Yan Chen. Full-stack vulnerability analysis of the cloudnative platform. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000834.

Zahid:2025:LWS

[ZKS25] Farzana Zahid, Matthew M. Y. Kuo, and Roopak Sinha. Light-weight slow-rate attack detection framework for resourceconstrained Industrial Cyber-Physical Systems. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S016740482500197X.

Zhang:2024:VIS

[ZL24] Guiqi Zhang and Yufeng Li. Voltage inspector: Sender identification for in-vehicle CAN bus using voltage slice.

Computers & Security, 145(??):??, October 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003225.

Zhang:2023:BSA

[ZLC^{+}23a] Guiqi Zhang, Qi Liu, Chenhong Cao, Jiangtao Li, and Yufeng Li. Bit scanner: Anomaly detection for in-vehicle CAN bus using binary sequence whitelisting. Computers & Security, 134(??):??, November 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823003462.

Zheng:2023:OMO

[ZLC^{+}23b] Haibin Zheng, Xiaohao Li, Jinyin Chen, Jianfeng Dong, Yan Zhang, and Changting Lin. One4All: Manipulate one agent to poison the cooperative multi-agent reinforcement learning. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822003972.

Zhang:2021:DSA

[ZLCA21] Xiaolu Zhang, Charles Zhechao Liu, Kim-Kwang Raymond Choo, and Jesus A. Alvarado. A design science approach to developing an integrated mobile app forensic framework. Computers & Security, 105(??):Article 102226, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482100050X.

Zhang:2025:MGR

[ZLDX25] Xiao Zhang, Yingxu Lai, Xinrui Dong, and Xinyu Xu. MER-GCN: Reasoning about attacking group behaviors using industrial control system attack knowledge graphs. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825002470.

Zhang:2020:MID

[ZLF^{+}20] Jianwu Zhang, Yu Ling, Xingbing Fu, Xiongkun Yang, Gang Xiong, and Rui Zhang. Model of the intrusion detection system based on the integration of spatial-temporal features. Computers & Security, 89(??):Article 101681, February 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404819302214.

Zhang:2022:MAA

[ZLH22] Chunkai Zhang, Xiaofeng Luo, and Peiyi Han. On-manifold adversarial attack based on latent space substitute model. Computers & Security, 120(??):Article 102770, September 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404822001651.

Zhong:2023:DMS

[ZLH23] Meihui Zhong, Mingwei Lin, and Zhu He. Dynamic multiscale topological representation for enhancing network intrusion detection. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823004261.

Zhou:2025:DRB

[ZLH^{+}25] Hanxun Zhou, Zhihui Liu, Yufeng Hu, Shuo Zhang, Longyu Kang, Yong Feng, Yan Wang, Wei Guo, and Cliff C. Zou. Data reduction for black-box adversarial attacks against deep neural networks based on side-channel attacks. Computers & Security, 153(??):??, June 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000902.

Zuo:2024:SSS

[ZLL24] Chu-Xiao Zuo, Jia-Yi Leng, and Wu-Jun Li. SUETA: Speakerspecific utterance ensemble based transfer attack on speaker identification system. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002530.

Zhang:2025:PAD

[ZLLY25] Sanfeng Zhang, Shangze Li, Juncheng Lu, and Wang Yang. Power-ASTNN: a deobfuscation and AST neural network enabled effective detection method for malicious PowerShell scripts. Computers & Security, 154(??):??, July 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825001300.

Zheng:2024:FVF

[ZLX^{+}24] Tianming Zheng, Haojun Liu, Hang Xu, Xiang Chen, Ping Yi, and Yue Wu. Few-VulD: a few-shot learning framework for software vulnerability detection. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002979.

Li:2025:WLL

[zLXZrxZ25] Yan zi Li, Li Xu, Jing Zhang, and Liao ru-xing Zhang. WF-LDPSR: a local differential privacy mechanism based on water-filling for secure release of trajectory statistics data. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400470X.

Zhan:2020:NNS

[ZLY^{+}20] Mengqi Zhan, Yang Li, Xinghua Yang, Wenjing Cui, and Yulin Fan. NSAPs: a novel scheme for network security state assessment and attack prediction. Computers & Security, 99(??):Article 102031, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303047.

Zhao:2021:APC

[ZLY^{+}21] Jun Zhao, Xudong Liu, Qiben Yan, Bo Li, Minglai Shao, Hao Peng, and Lichao Sun. Automatically predicting cyber attack preference with attributed heterogeneous attention networks and transductive learning. Computers & Security, 102(??):Article 102152, March 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404820304259.

Zheng:2023:DIT

[ZLY23] Chenming Zheng, Jun Li, and Xuanxia Yao. Design and implementation of trusted boot based on a new trusted computing dual-architecture. Computers & Security, 127(??):??, April 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823000056.

Zhao:2024:DFN

[ZLZ^{+}24] Ziming Zhao, Zhaoxuan Li, Zhihao Zhou, Jiongchi Yu, Zhuoxue Song, Xiaofei Xie, Fan Zhang, and Rui Zhang. DDoS family: a novel perspective for massive types of DDoS attacks. Computers & Security, 138(??):??, March 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005734.

Zheng:2022:TDA

[ZLZC22] Tengfei Zheng, Yuchuan Luo, Tongqing Zhou, and Zhiping Cai. Towards differential access control and privacypreserving for secure media data sharing in the cloud. Computers & Security, 113(??):Article 102553, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003771.

Zhong:2024:SGN

[ZLZX24] Meihui Zhong, Mingwei Lin, Chao Zhang, and Zeshui Xu. A survey on graph neural networks for intrusion detection systems: Methods, trends and challenges. Computers & Security, 141(??):??, June 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824001226.

Zhang:2020:FEA

[ZM20] Lyuye Zhang and Maode Ma. FKR: an efficient authentication scheme for IEEE 802.11ah networks. Computers & Security, 88(??):Article 101633, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404818313373.

Zhang:2022:MAC

[ZM22] Haotian Zhang and Xu Ma. Misleading attention and classification: an adversarial attack to fool object detection models in the real world. Computers & Security, 122(??):Article 102876, November 2022. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482200270X.

Zhang:2025:DUC

[ZM25] Yunxiao Zhang and Pasquale Malacaria. Dealing with uncertainty in cybersecurity decision support. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404824004589.

Zhang:2023:SDD

[ZMF^{+}23] Hengwei Zhang, Yan Mi, Yumeng Fu, Xiaohu Liu, Yuchen Zhang, Jindong Wang, and Jinglei Tan. Security defense decision method based on potential differential game for complex networks. Computers & Security, 129(??):??, June 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823000974.

Zhan:2025:HCA

[ZMF^{+}25] Lige Zhan, Jiang Ming, Jianming Fu, Guojun Peng, Letian Sha, and Lili Lan. The hidden complexities of Android TPL detection: an empirical analysis of techniques, challenges, and effectiveness. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482500361X.

Zhang:2023:CFC

[ZMLP23] Yunxiao Zhang, Pasquale Malacaria, George Loukas, and Emmanouil Panaousis. CROSS: a framework for cyber risk optimisation in smart homes. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823001608.

Zhang:2021:EHA

[ZMQ21] Lyuye Zhang, Maode Ma, and Yue Qiu. An enhanced handover authentication solution for 6LoWPAN networks. Computers & Security, 109(??):Article 102373, October 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001978.

Zhai:2025:MMC

[ZMZ^{+}25] Yuqi Zhai, Rui Ma, Zheng Zhang, Siqi Zhao, and Yuche Yang. MSNFuzz: Multi-criteria state-sensitive network protocol fuzzing. Computers & Security, 158(??):??, November 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003104.

Zhang:2021:DPG

[ZNF21] Sen Zhang, Weiwei Ni, and Nan Fu. Differentially private graph publishing with degree distribution preservation. Computers & Security, 106(??):Article 102285, July 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821001097.

Zhang:2023:LDP

[ZNF^{+}23] Dongyue Zhang, Weiwei Ni, Nan Fu, Lihe Hou, and Ruyu Zhang. Locally differentially private multi-dimensional data collection via Haar transform. Computers & Security, 130(??): ??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823002018.

Zhao:2020:PPS

[ZNL^{+}20] Yi Zhao, Jianting Ning, Kaitai Liang, Yanqi Zhao, Liqun Chen, and Bo Yang. Privacy preserving search services against online attack. Computers & Security, 95(??):Article 101836, August 2020. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404820301097.

Zago:2020:UDP

[ZPP20] Mattia Zago, Manuel Gil Pérez, and Gregorio Martínez Pérez. UMUDGA: a dataset for profiling DGA-based botnet. Computers & Security, 92(??):Article 101719, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300067.

Zhang:2024:LBA

[ZQCX24] Jiebao Zhang, Wenhua Qian, Jinde Cao, and Dan Xu. LP-BFGS attack: an adversarial attack based on the Hessian with limited pixels. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824000476.

Zhang:2020:AKC

[ZQMC20] Qiang Zhang, Jianzhong Qiao, Qingyang Meng, and Yu Chen. Automatic kernel code synthesis and verification. Computers & Security, 91(??):Article 101733, April 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300201.

Zhao:2024:TTD

[ZS24] Jing Zhao and Qianqian Su. A threshold traceable delegation authorization scheme for data sharing in healthcare. Computers & Security, 139(??):??, April 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005965.

Zola:2022:NTA

[ZSGB^{+}22] F. Zola, L. Segurola-Gil, J. L. Bruse, M. Galar, and R. Orduna-Urrutia. Network traffic analysis through node behaviour classification: a graph-based approach with temporal dissection and data-level preprocessing. Computers & Security, 115(??):Article 102632, April 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822000311.

Zheng:2025:VSD

[ZSJ^{+}25] Weining Zheng, Xiaohong Su, Yuan Jiang, Hongwei Wei, and Wenxin Tao. VDExplainer: Sequential decision-making and probability sampling guided statement-level explanation for vulnerability detection. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404825003591.

Zhang:2024:FIB

[ZSL24] Huijie Zhang, Weizhen Sun, and Ling Lv. A frequencyinjection backdoor attack against DNN-based finger vein verification. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016740482400261X.

Zhou:2025:CBM

[ZSL25] Hao Zhou, Wenting Shen, and Jinlu Liu. Certificate-based multi-copy cloud storage auditing supporting data dynamics. Computers & Security, 148(??):??, January 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404824004012.

Zhang:2025:MMP

[ZSLY25] Sanfeng Zhang, Heng Su, Hongxian Liu, and Wang Yang. MPDroid: a multimodal pre-training Android malware detection method with static and dynamic features. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824005686.

Zibak:2021:SMC

[ZSS21] Adam Zibak, Clemens Sauerwein, and Andrew Simpson. A success model for cyber threat intelligence management platforms. Computers & Security, 111(??):Article 102466, December 2021. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016740482100290X.

Zheng:2024:SVD

[ZSWT24] Weining Zheng, Xiaohong Su, Hongwei Wei, and Wenxin Tao. SVulDetector: Vulnerability detection based on similarity using tree-based attention and weighted graph embedding mechanisms. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824002335.

Zeng:2025:RPP

[ZSZ^{+}25] Li Zeng, Peisong Shen, Xiaojie Zhu, Xue Tian, and Chi Chen. A review of privacy-preserving biometric identification and authentication protocols. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824006151.

Zhang:2021:TDD

[ZT21] Li Zhang and Vrizlynn. L. L. Thing. Three decades of deception techniques in active cyber defense — retrospect and outlook. Computers & Security, 106(??):Article 102288, July 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404821001127.

Zhang:2021:EAT

[ZTD21] Jie Zhang, Cong Tian, and Zhenhua Duan. An efficient approach for taint analysis of Android applications. Computers & Security, 104(??):Article 102161, May 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030434X.

Zulfiqar:2021:EEB

[ZTJ^{+}21] Maryam Zulfiqar, Filza Tariq, Muhammad Umar Janjua, Adnan Noor Mian, Adnan Qayyum, Junaid Qadir, Falak Sher, and Muhammad Hassan. EthReview: an Ethereum-based product review system for mitigating rating frauds. Computers & Security, 100(??):Article 102094, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303679.

Zhang:2023:SAC

[ZTK^{+}23] Jie Zhang, Min-Yen Tsai, Kotcharat Kitchat, Min-Te Sun, Kazuya Sakai, Wei-Shinn Ku, Thattapon Surasak, and Tipajin Thaipisutikul. A secure annuli CAPTCHA system. Computers & Security, 125(??):??, February 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004175.

Zhang:2023:PDP

[ZTR^{+}23] Zhiqiu Zhang, Zhu Tianqing, Wei Ren, Ping Xiong, and Kim-Kwang Raymond Choo. Preserving data privacy in federated learning through large gradient pruning. Computers & Security, 125(??):??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482200431X.

Zou:2025:SAL

[ZTW25] Dikai Zou, Jun Tao, and Zuyan Wang. Semantics-aware location privacy preserving: a differential privacy approach. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404825000914.

Zhao:2020:CTC

[ZTY^{+}20a] Yujie Zhao, Zhanyong Tang, Guixin Ye, Dongxu Peng, Dingyi Fang, Xiaojiang Chen, and Zheng Wang. Compiletime code virtualization for Android applications. Computers & Security, 94(??):Article 101821, July 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300936.

Zhao:2020:SAO

[ZTY^{+}20b] Yujie Zhao, Zhanyong Tang, Guixin Ye, Dongxu Peng, Dingyi Fang, Xiaojiang Chen, and Zheng Wang. Semanticsaware obfuscation scheme prediction for binary. Computers & Security, 99(??):Article 102072, December 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482030345X.

Zhang:2025:DTB

[ZTYA25] Yefei Zhang, Sadegh Torabi, Jun Yan, and Chadi Assi. Dynamic trigger-based attacks against next-generation IoT malware family classifiers. Computers & Security, 149(??):??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004929.

Zarin:2023:NSN

[ZUKH23] Rahat Zarin, Niamat Ullah, Amir Khan, and Usa Wannasingha Humphries. A numerical study of a new non-linear fractal fractional mathematical model of malicious codes propagation in wireless sensor networks. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404823003942.

Zhao:2024:PPE

[ZW24] Yongxin Zhao and Chundong Wang. Protecting privacy and enhancing utility: a novel approach for personalized trajectory data publishing using noisy prefix tree. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404824002244.

Zhang:2025:PII

[ZWB^{+}25] Dongping Zhang, Mengting Wang, Yuzhen Bu, Jiabin Yu, and Li Yang. PdGAT-ID: an intrusion detection method for industrial control systems based on periodic extraction and spatiotemporal graph attention. Computers & Security, 149(??): ??, February 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824005169.

Zanke:2024:AIS

[ZWDE24] Anna Zanke, Thorsten Weber, Peter Dornheim, and Mathias Engel. Assessing information security culture: a mixedmethods approach to navigating challenges in international corporate IT departments. Computers & Security, 144(??):??, September 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002438.

Zha:2024:SIU

[ZWF^{+}24] Chao Zha, Zhiyu Wang, Yifei Fan, Xingming Zhang, Bing Bai, Yinjie Zhang, Sainan Shi, and Ruyun Zhang. SKT-IDS: Unknown attack detection method based on Sigmoid Kernel Transformation and encoder-decoder architecture. Computers & Security, 146(??):??, November 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404824003614.

Zheng:2025:SNP

[ZWH^{+}25] Chaoyang Zheng, Yunchao Wang, Huihui Huang, Yunfeng Wang, Haowen Chen, and Qiang Wei. Survey of network protocol fuzzers: Taxonomy, techniques, and directions. Computers & Security, 159(??):??, December 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404825003724.

Zhao:2025:ASD

[ZWL^{+}25] Jianzhou Zhao, Qiang Wei, Xingwei Li, Yunchao Wang, and Xixing Li. ASIRDetector: Scheduling-driven, asynchronous execution to discover asynchronous improper releases bug in Linux kernel. Computers & Security, 157(??):??, October 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002196.

Zhang:2021:SCI

[ZWW^{+}21] Guowen Zhang, Bo Wang, Fei Wei, Kaize Shi, Yue Wang, Xue Sui, and Meineng Zhu. Source camera identification for recompressed images: a model perspective based on tri-transfer learning. Computers & Security, 100(??):Article 102076, January 2021. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404820303497.

Zheng:2020:OFB

[ZWX^{+}20] Haibin Zheng, Qianhong Wu, Jan Xie, Zhenyu Guan, Bo Qin, and Zhiqiang Gu. An organization-friendly blockchain system. Computers & Security, 88(??):Article 101598, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404818314111.

Zhou:2024:FFS

[ZWX^{+}24] Fan Zhou, Dongxia Wang, Yanhai Xiong, Kun Sun, and Wenhai Wang. FAMCF: a few-shot Android malware family classification framework. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404824003328.

Zhu:2025:FTD

[ZWZX25] Yuedi Zhu, Yong Wang, Lin Zhou, and Yuan Xia. FC-Trans: Deep learning methods for network intrusion detection in big data environments. Computers & Security, 154(??):??, July 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404825000811.

Zhang:2023:VVD

[ZX23a] Chunyong Zhang and Yang Xin. VulGAI: vulnerability detection based on graphs and images. Computers & Security, 135(??):??, December 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016740482300411X.

Zhao:2023:SON

[ZX23b] Xueli Zhao and Zhihua Xia. Secure outsourced NB: Accurate and efficient privacy-preserving Naive Bayes classification. Computers & Security, 124(??):??, January 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822004035.

Zhang:2024:CEV

[ZXB^{+}24] Zheng Zhang, Jingfeng Xue, Thar Baker, Tian Chen, Yu an Tan, and Yuanzhang Li. COVER: Enhancing virtualization obfuscation through dynamic scheduling using flash controller-based secure module. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167404824003432.

Zhang:2025:ECP

[ZXH^{+}25] Jie Zhang, Lei Xie, Lang He, Zhongmin Wang, and Jing Chen. Enhanced cell phone security: an ultrasonic and sensor fusion-based persistent cell phone protection method integrating anti-theft & identity authentication. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404824004814.

Zhang:2024:TSB

[ZXL^{+}24] Xiyuan Zhang, Gang Xiong, Zhen Li, Chen Yang, Xinjie Lin, Gaopeng Gou, and Binxing Fang. Traffic spills the beans: a robust video identification attack against YouTube. Computers & Security, 137(??):??, February 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823005333.

Zhan:2025:PCL

[ZXL^{+}25] Dazhi Zhan, Kun Xu, Xin Liu, Tong Han, Zhisong Pan, and Shize Guo. Practical clean-label backdoor attack against static malware detection. Computers & Security, 150(??):??, March 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0167404824005868.

Zheng:2025:CIT

[ZXS^{+}25] Qianrong Zheng, Jianwen Xiang, Changtian Song, Rivalino Matias, Rui Hao, Songsong Liao, Xuemin Zhang, Meng Zhao, and Dongdong Zhao. Cancelable iris template based on slicing. Computers & Security, 152(??):??, May 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825000707.

Zhou:2025:HNF

[ZXW^{+}25] Weidong Zhou, Chunhe Xia, Tianbo Wang, Xiaopeng Liang, Wanshuang Lin, Xiaojian Li, and Song Zhang. HIDIM: a novel framework of network intrusion detection for hierarchical dependency and class imbalance. Computers & Security, 148(??): ??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004607.

Zhang:2024:TBE

[ZXXG24] Bo Zhang, Yu-Jie Xiong, Chunming Xia, and Yongbin Gao. Transformer-based end-to-end attack on text CAPTCHAs with triplet deep attention. Computers & Security, 146(??):??, November 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824003638.

Zhang:2020:EPP

[ZXZ^{+}20] Chuan Zhang, Chang Xu, Liehuang Zhu, Yanwei Li, Can Zhang, and Huishu Wu. An efficient and privacy-preserving truth discovery scheme in crowdsensing applications. Computers & Security, 97(??):Article 101848, October 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301218.

Zhang:2025:APT

[ZXZJ25] Xin Zhang, Shijie Xiao, Han Zhang, and Lixia Ji. Adaptive patch transformation for adversarial defense. Computers & Security, 153(??):??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404825000574.

Zhijun:2021:BDP

[ZY21]  Wu Zhijun and Yang Yiming.  BD-D1Sec: Protocol

of security authentication for BeiDou D1 civil navigation message based on certificateless signature. Computers & Security, 105(??):Article 102251, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000754.

Zhang:2024:DMW

[ZY24] Longwen Zhang and Qiao Yan. Detect malicious websites by building a neural network to capture global and local features of websites. Computers & Security, 137(??):??, February 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404823005503.

Zhao:2024:TES

[ZYGH24] Xuan Zhao, Jia Yu, Xinrui Ge, and Rong Hao. Towards efficient Secure Boolean Range Query over encrypted spatial data. Computers & Security, 136(??):??, January 2024. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404823004546.

Zhou:2025:WEN

[ZYGL25] Luchen Zhou, Wei-Chuen Yau, Y. S. Gan, and Sze-Teng Liong. E-WebGuard: Enhanced neural architectures for precision web attack detection. Computers & Security, 148(??):??, January 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404824004322.

Zhang:2021:SLS

[ZYH^{+}21] Haikuo Zhang, Jueyu Ye, Weihong Hu, Qian Wang, Xiali Yan, Qiaoli Yue, Wanbo Lv, Ming He, and Jue Wang. Study on the latent state of Kaminsky-style DNS cache poisoning: Modeling and empirical analysis. Computers & Security, 110(??):Article 102445, November 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821002698.

Zhang:2022:CCL

[ZYH^{+}22] Zongyang Zhang, Jiayuan Yin, Bin Hu, Ting Gao, Weihan Li, Qianhong Wu, and Jianwei Liu. CLTracer: a cross-ledger tracing framework based on address relationships. Computers & Security, 113(??):Article 102558, February 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003825.

Zhao:2023:FEQ

[ZYJ^{+}23] Xiaoyun Zhao, Chao Yang, Zhizhuang Jia, Yue Wang, and Jianfeng Ma. FCEVAL: an effective and quantitative platform for evaluating fuzzer combinations fairly and easily. Computers & Security, 132(??):??, September 2023. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016740482300264X.

Zhao:2020:TAE

[ZYL^{+}20] Jun Zhao, Qiben Yan, Jianxin Li, Minglai Shao, Zuti He, and Bo Li. TIMiner: Automatically extracting and analyzing categorized cyber threat intelligence from social data. Computers & Security, 95(??):Article 101867, August 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820301395.

Zheng:2022:PHC

[ZYL^{+}22] Faan Zheng, Qiao Yan, Victor C. M. Leung, F. Richard Yu, and Zhong Ming. HDP-CNN: Highway deep pyramid convolution neural network combining word-level and characterlevel representations for phishing website detection. Computers & Security, 114(??):Article 102584, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821004077.

Zhang:2025:SAD

[ZYL^{+}25] Dongni Zhang, Xiuzhang Yang, Side Liu, Yilin Zhou, Jianming Fu, and Guojun Peng. A survey on Android dynamic evasive malware: Taxonomy, countermeasures and open challenges. Computers & Security, 159(??):??, December 2025. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-

6208 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167404825003359.

Zhu:2022:DPH

[ZYPT22] Hui Zhu, Fan Yin, Shuangrong Peng, and Xiaohu Tang. Differentially private hierarchical tree with high efficiency. Computers & Security, 118(??):Article 102727, July 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404822001225.

Zuzcak:2020:ESA

[ZZ20] Matej Zuzcák and Milan Zenka. Expert system assessing threat level of attacks on a hybrid SSH honeynet. Computers & Security, 92(??):Article 101784, May 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820300699.

Zhu:2025:DGI

[ZZ25] Hou Zhu and Mingxin Zhang. “I don’t get it, but I accept it”: exploring uninformed consent to privacy policies: a neutralization perspective. Computers & Security, 153(??): ??, June 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825000859.

Zhou:2020:PPP

[ZZCD20] Jun Zhou, Meng Zheng, Zhenfu Cao, and Xiaolei Dong. PVIDM: Privacy-preserving verifiable shape context based image denoising and matching with efficient outsourcing in the malicious setting. Computers & Security, 88(??):Article 101631, January 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0167404819301786.

Zhang:2025:WDG

[ZZHY25] Hongpo Zhang, Zhaozhe Zhang, Haizhaoyang Huang, and Hehe Yang. Wasserstein distance guided feature Tokenizer transformer domain adaptation for network intrusion detection. Computers & Security, 157(??):??, October 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0167404825002512.

Zhu:2020:DBA

[ZZJC20] Xiaodong Zhu, Yi Zhang, Liehui Jiang, and Rui Chang. Determining the base address of MIPS firmware based on absolute address statistics and string reference matching. Computers & Security, 88(??):Article 101504, January 2020. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404819300860.

Zheng:2022:APM

[ZZL^{+}22] Zhangqi Zheng, Bing Zhang, Yongshan Liu, Jiadong Ren, Xuyang Zhao, and Qian Wang. An approach for predicting multiple-type overflow vulnerabilities based on combination features and a time series neural network algorithm. Computers & Security, 114(??):Article 102572, March 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821003965.

Zhang:2025:VVA

[ZZP^{+}25] Shaobo Zhang, Lujie Zhang, Tao Peng, Qin Liu, and Xiong Li. VADP: Visitor-attribute-based adaptive differential privacy for IoMT data sharing. Computers & Security, 156(??):??, September 2025. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404825002020.

Zhu:2023:ADB

[ZZS^{+}23] Chengcheng Zhu, Jiale Zhang, Xiaobing Sun, Bing Chen, and Weizhi Meng. ADFL: Defending backdoor attacks in federated learning via adversarial distillation. Computers & Security, 132(??):??, September 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167404823002766.

Zhang:2021:RAT

[ZZW^{+}21] Lejun Zhang, Yanfei Zou, Weizheng Wang, Zilong Jin, Yansen Su, and Huiling Chen. Resource allocation and trust computing for blockchain-enabled edge computing system. Computers & Security, 105(??):Article 102249, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000730.

Zheng:2024:MBA

[ZZW^{+}24] Luxin Zheng, Jian Zhang, Xiangyi Wang, Faxin Lin, and Zheng Meng. Multimodal-based abnormal behavior detection method in virtualization environment. Computers & Security, 143(??): ??, August 2024. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404824002104.

Zhang:2025:EIS

[ZZW^{+}25] Yufeng Zhang, Hongxin Zhang, Yijun Wang, Xiaorong Gao, and Chen Yang. Enhancing information security through brainprint: a longitudinal study on ERP identity authentication. Computers & Security, 150(??):??, March 2025. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S016740482400587X.

Zhou:2021:DMS

[ZZWF21] Peng Zhou, Gongyan Zhou, Dakui Wu, and Minrui Fei. Detecting multi-stage attacks using sequence-to-sequence model. Computers & Security, 105(??):Article 102203, June 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404821000274.

Zhu:2023:LGA

[ZZZ^{+}23] Hegui Zhu, Ying Zhu, Haoran Zheng, Yuchen Ren, and Wuming Jiang. LIGAA: Generative adversarial attack method based on low-frequency information. Computers & Security, 125(??): ??, February 2023. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167404822004497.